A Complete Bibliography of Publications in Conservation Biology: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

23 April 2025 Version 1.06

Title word cross-reference

$ [Bra15, Fie15]. $104.50 [Fie15]. $29.95 [Ano16x]. $55.00 [Ano16x]. ^{st} [RV15].

1 [MJG15].

2 [Ano16x]. 2.5-million-year [GBB^{+}15]. 20-year [DPGY16]. 2000 [BBM18, KHD^{+}15, MAM^{+}15]. 2010 [Ano11b, JCA^{+}11, ZBM^{+}10]. 2011 [Ano11a]. 2015 [Ano16a]. 2016 [Ano17a]. 2019 [Ano19b]. 21st [For11, LSK^{+}15, SPL^{+}11, TBM^{+}16, Car14]. 21st-century [TBM^{+}16].

3 [Bra15].

4 [Ano19w, Fie15]. 4th [Ano19w].

978 [Ano16x, Ano19w, Bra15, Fie15]. 978-0-19-964617-3 [Bra15]. 978-0-19-967958-4 [Fie15]. 978-0-226-11547-4 [Ano16x].

978-0-226-38621-8 [Ano19w]. 978-0-691-16137-2 [Ano16x]. 978-0-691-16500-4 [Ano19w]. 978-1-118-89527-6 [Ano19w]. 978-1-61091-796-4 [Ano19w].

Abalone [CCM12]. abandoned [WHG^{+}15]. abandonment [KK18]. abatable [KJWP19]. Abejas [JDA^{+}10]. Abejorros [MBGS11]. Abierto [CB10]. Abiertos [EMD^{+}10]. Ability [CEM^{+}13, LA13]. abiotic [AB15, BSH^{+}15]. aboveground [CML^{+}19]. Absence [SMH^{+}14]. abstract [AGS19]. abstracts [Ano18x]. Abundance [BBP12, CBM^{+}11, CLRV11, CPC^{+}10, CMNM10, EPD12, IUM10, KO12, KTN12, MRBL^{+}12, PPD^{+}12, RSPD13, TG13, WSHB11, WHB^{+}11, WKP11, BGO^{+}19, FHBK18, MPS^{+}16, RNE^{+}19, SA15, WLCW16]. Abundances [NCC13]. Abundancia [BBP12, CBM^{+}11, CLRV11, CPC^{+}10, CMNM10, EPD12, IUM10, KTN12, PPD^{+}12, WSHB11, WHB^{+}11, WKP11]. academic [Bur19, LET19]. Accelerating [CMG^{+}13]. Acceso [CB10]. Access [CB10, FLW14a, FLW14b, Löv16, Bur19, CBF^{+}19, RTGR17]. Acción [LSDK11]. Acciones [BGAS10, HM12, KCDC10]. Account [MB10, SGA^{+}19, TBF^{+}17]. accountability [ABMG19]. accounted [MCO^{+}17]. Accounting [CPC^{+}10, DKG^{+}15, MGMA^{+}14, TCP13, WHK^{+}17, FMV^{+}19]. Accumulation [GLNJ10]. Accuracy [CCBP14, JPT13]. accurate [EG17]. aceclofenac [GTC^{+}16]. Achieve [SBSF14, CGK^{+}19, CS17b]. Achieving [BGL^{+}14, CMS^{+}13, FLW14a, GHB^{+}13, Ric11, SKC^{+}12, MBV^{+}16]. Acid [MPH^{+}11]. Acidification [GFGR12, CE15]. Ácidos [MPH^{+}11]. Acknowledging [HAB^{+}11]. acknowledgment [RLL15]. Acknowledgments [Ano10g, Ano11i, Ano13e]. Acoustic [HCP^{+}12, PVP^{+}13, WRTB10, BMC^{+}18, JLCHNG^{+}17]. Acoustically [MRT^{+}12]. acquire [Arm18]. Across [CMB12, PSP11, ACS14, ALL^{+}12, BCA^{+}18, BS12, DSE15, GS16, GLNJ10, HH19b, JDA^{+}10, LWZL16, PCF^{+}19, Pek13, RCSK15, RK17]. Act [Fra11, CVN^{+}10, ENW^{+}18, Les15, LRMWG16, MAB^{+}13, RTT^{+}13, SRL^{+}13, TG15, Wil17, ZCCN13]. Acta [CVN^{+}10]. Action [BGAS10, CMS^{+}13, HDHL10, Joh10b, KCDC10, LW14, LSDK11, DSV17, GMG^{+}19, GTVC16, HPD16, LBB^{+}16, SSC^{+}19]. Action-Selection [BGAS10]. Actions [DFJ^{+}12, HODB10, HM12, NMMP14, BPH^{+}19, CM17]. active [SFKC16]. Actividad [DDPC12, HSAB10]. Actividades [CW10, LW11]. Activities [CW10, LW11, CPD^{+}15, HNO^{+}18]. Activity [BMA^{+}14, DDPC12, HSAB10, RMKM14]. Actores [BB10a, GZD11]. Actors [BB10a]. Actual [ZCCN13]. Actuales [Col10, MVB^{+}10, SKC^{+}12]. Acuática [MARV^{+}11]. Acuáticas [BFHP10, DC11]. Acuáticos [HMD^{+}10]. Acuerdos [HRC10]. Acumulación [GLNJ10]. Acústicas [WRTB10]. Ad [Ano12a, MO11]. Adams [SLH10]. Adapt [HH14, CBM16]. Adaptabilidad [SS12]. Adaptación [Han10, SM11]. Adaptation [CMG^{+}13, Han10, SM11, KP19, KLL^{+}16]. Adaptativo [CCBF12, HMM10, MB10, MFN^{+}11, MM10]. Adapting [Joh10a, PPM^{+}18]. Adaptive [CCBF12, HM16, HMM10, KHB18, MB10, MFN^{+}11, MM10, NS12, SAGSG^{+}14, Sto11, WJ13, AGFR18, FR16, JGS^{+}16, NBR^{+}15, RHWS18, WHK^{+}17]. Adaptive-Management [MFN^{+}11, MM10]. adaptively [HWB^{+}16]. Address [CMG^{+}13, DBR13, DSW16, NLM^{+}13, MBFB18, RKW19, WBB^{+}18, WHE^{+}17]. Addressing [GMV^{+}18, PS10a]. Adecuación [WB10]. Adecuadamente [HMD^{+}10]. Adecuados [PCTW10]. Adequately [HMD^{+}10]. adherence [PVO^{+}16]. Adición [HSAB10]. Adjustive [MB10]. ADN [HJSF^{+}12, KGS^{+}11]. Adopción [FKM^{+}11]. Adoption [FKM^{+}11]. advance [ABMG19, CAB^{+}18]. Advancing [SH11b]. Adventure [Nov16]. Advocacy [SR11, Wil12a, Wil12b, HPBP16, WGLL17]. Aerotransportado [SAC12]. Affect [Rud11a]. Affected [WSEM13]. Affecting [SBB^{+}14, ZAB16]. Afraid [Mih19]. Africa [BVAC17, CBM^{+}18, DWM^{+}19, ERC^{+}10, FWM^{+}12, HRC10, LW10, PCTW10, PRE^{+}13, VKKA13, PCTW10, FWM^{+}12, GSW10, LW10]. African [Ben15a, DHF16, EBL^{+}14, GT10, GSW10, KO12, MKR^{+}15, NBM^{+}12, VTT^{+}17, Wit11]. Africana [GT10]. Africanos [Wit11]. Afrotropical [PRM^{+}18]. After [Sch14b, BIP^{+}15, BJMT14, EDB12, HLC13, PGC^{+}13b, SS10a, VLGRS14]. against [BP13, GGM14, Par10, VMS^{+}12, YNR19]. Age [Gon13, McP11, GLdPL18, SRG^{+}15]. agencies [FR16, RKC^{+}19]. agency [LRMWG16]. Agenda [Fis12, OSF10, MAB^{+}15]. Agendas [SPL^{+}11]. agent [TT16]. Agents [SJP^{+}13]. aggregation [ACW18, BTTP17, WD17]. agri [BDKS15]. agri-environment [BDKS15]. Agrícola [MPB10]. Agrícolas [RWB11, GBG^{+}10, GZR12]. Agricultural [CKR^{+}13, EPB^{+}14, EBRS14, GBG^{+}10, GZR12, MG12, MPB10, Pek13, Sto12, Bra17, CPR^{+}16, DMF^{+}18, EKRW17, ERW18, HHB15, RK17]. Agriculture [ER14, KH16, MDJ14, ZFW^{+}11, DPS^{+}16, GUHE15, MAB^{+}15, SSW19]. Agriculture-Developed [ER14]. Agrobosques [GBG^{+}10]. Agroforests [GBG^{+}10, TQD^{+}18]. Aguas [VLRB12]. Aichi [MJG15, TUHK^{+}17]. Airborne [SAC12]. Aisladas [ZZW^{+}10]. Aislamiento [SPPN^{+}12]. Ajustativo [MB10]. Akcakaya [San19]. Aketajawe [SHA^{+}10a]. Aketajawe-Lolobata [SHA^{+}10a]. al [ABD^{+}17, Bai16, BE10, BBM15, CND10, Cat11, Cla15, Col10, CPC^{+}10, DW19, Dys15, ERBM17, FDBA17, FHK^{+}15, GLH12, Han10, HODB10, Har15, HRSK12, HFK^{+}10, IF17, KF18a, KNK13, LS19, LDC^{+}15, LET19, LDB^{+}10, MB10, Mat15, MTM12, MBW^{+}19, MH16, NJK^{+}19, NSBJ17b, OSF11, Oso15, PBH^{+}10b, PGLB^{+}12, Pet17, PI11, PHJ^{+}17, RM19, San19, Sch10, SSK^{+}11, SMB^{+}10, SHC^{+}11, TK15, TT16, WMS^{+}12, Wil12b, WRG^{+}12, vWWCF15]. Álamo [EDB12]. Alaska [AS13, BPA^{+}13, Lor13, PMM^{+}15, WHE^{+}17]. Alaskan [MDSF19]. albatrosses [WFC^{+}18]. Alcance [MC11a]. Alcázar [PS16]. Aldea [CAB^{+}10]. Alert [Str13]. Aleutian [ETB10]. Aleutiano [ETB10]. algal [BBH17]. Algas [ETB10, HRKMG12]. Algunos [LH10].

Alien [NKB10, SS10a]. Aligning [CS17b]. Alignment [AP17, BCT^{+}14, BCM16, HHB15]. Alimentaria [HSBJ12]. Allá [JCA^{+}11, ZBM^{+}10]. Allee [HCLW13, KKH14]. allegations [MSL19]. Allegheny [SJP^{+}13]. Alleles [WGJ13]. Alliances [PDM11]. Allies [Ano19w]. Allocation [RHM10, DKG^{+}15]. Allsopp [Fie15]. along [DMF^{+}18, MFRH12, OGS^{+}19]. Alpine [ER14, Pyk17]. Also [Sou14a]. Alta [BEE^{+}11, SHLK^{+}10]. Alterado [KPL^{+}12]. Alterados [BS10]. Altered [BS10, KPL^{+}12]. Alternativa [HRKMG12]. Alternative [BW13, CT11, GSDE14, HRKMG12, Lor13, McP12, OSF10, PZM14, RGRMPB^{+}13, UDG10, JBH^{+}16, KGF^{+}18, WHR^{+}16]. Alternatives [CBCW18]. Alternativo [CT11]. Alternativos [UDG10]. Altitud [JMP^{+}11]. Altitude [BME^{+}14]. Altitudinal [FMJP11a]. Alto [SSD10, SHC^{+}11]. Altura [Van12]. Always [Mih19]. Amaze [Ban12]. Amazon [BMA^{+}18, NMSF11, NA13, Plo17, RSE12]. Amazonia [DMZ10, MLA^{+}14, DMZ10, RSE12]. Amazonian [BHPM^{+}18, DMF^{+}18, EVF13, FSSW19, GH18, LOME10, LGB^{+}10, OQGR^{+}15]. Amazónica [LOME10]. Amazónicos [LGB^{+}10]. Ambient [OSFFB^{+}11]. Ambiental [AG10, BHK10, Bro10, OSFFB^{+}11, PHC^{+}11]. Ambientales [CGCL11, HSBE12, SMGRJ10]. ambiguous [LMJK^{+}19]. Ambitious [BBC^{+}14]. Amenaza [FP10, RDM10]. Amenazada [DWM^{+}10, MARV^{+}11]. Amenazadas [AWU^{+}11, FLP^{+}11, LN11, MSSM09, WL12, WEC^{+}11]. Amenazados [KTN12, MVB^{+}11, MVB^{+}12, RRCB^{+}11]. Amenazas [DWCW10, FB12, MA10]. America [CMB12, CM17, Cor19, DPS^{+}16, EB15, Mat13, MLRC18, SSRS18, SVN^{+}15, CMB12, HM12]. American [AB13b, BB11a, BCM16, BCLMM19, DHH^{+}12, GLH12, HM12, SM11, TTSZH10, WB14]. Americano [SM11]. Amidst [BBC^{+}14]. among [BGH^{+}15, BKR17, Fli10, FWM^{+}12, GGL^{+}15, GFGR12, HGM19, HSBJ12, NSBJ17a, NSBJ17b, ÖB12, SSVA11, WB14]. Amount [MC14]. Amounts [SDB10]. Amphibian [AEO^{+}13, Bee13, BS10, CLV14, CPB14, GZNC13, HLC13, HAP^{+}13, JLDJ12, JMP^{+}11, MSSM09, OB11b, OSFFB^{+}11, PMS^{+}10, PHC^{+}11, Sch10, SGH^{+}11, SMH^{+}14, SAGSHL10, SDB10, BGH^{+}15, DPGY16, HGP16, MSC^{+}18, SHB^{+}17, SSP^{+}15, TND16]. Amphibian-Decline [OB11b]. Amphibians [AGLRV14, BFHP10, Cat11, FMJP11b, GLH12, LSBB10, SHG^{+}14, WIM^{+}10, WR12, WH13, DP15, IO17, MAG18, NWW^{+}17]. amplification [SHB^{+}17]. amplify [GSSF19]. Análisis [ALL^{+}12, BR11b, CMB12, DDPC12, HM12, MPH^{+}11, MJP^{+}11, ML10, SSK^{+}11, WW10b, BAPP11, JLCK12, MR12, SVH12, SS11, VJMH12, WKF^{+}11, PBH^{+}10a]. analítica [GNWSB10]. Analíticas [BRKP12]. Analog [Col10]. Análogo [Col10]. analogs [LMM^{+}17]. Analyses [ALL^{+}12, EIS14, GMCW13, MR12, PMJ^{+}13, SMTW13, SSK^{+}11, ZCCN13, BRA^{+}19, CCG18, Had15]. Analysis [BAPP11, BAGM13, BR11b, CMB12, CM13, DDPC12, DBR13, ER14, GAFG14, HTH^{+}10, HM12, JLCK12, KKK14, MLLS14, MPH^{+}11, MSF^{+}12, MJP^{+}11, ML10, MDJ14, PBH^{+}10a, PBH^{+}10b, SHWT14, SS11, TNTS13, VJMH12, WW10b, ALA17, DSE15, GDG19, KK18, LB15, MSR19, SEM^{+}15, SJZ19, SCHM18, WHE^{+}17, WBPL15]. Analytic [GNWSB10]. Analytical [BRKP12]. Analyze [Nas17]. Analyzing [SVH12, WKF^{+}11, RS18]. anathema [VBN15]. Anchoring [MAA^{+}10]. Ancient [BME^{+}14, PTN10, WMW^{+}14, WHE^{+}17]. Ancillary [GFG^{+}12]. Anclaje [MAA^{+}10]. Andean [AMO11]. Andes [OPP14, CLV14]. Andinos [AMO11]. anecdotal [MCO^{+}17]. anecdotes [Bai16]. Anfibio [JLDJ12, JMP^{+}11, MSSM09, PMS^{+}10]. Anfibios [BFHP10, BS10, GLH12, LSBB10, OB11b, OSFFB^{+}11, PHC^{+}11, SGH^{+}11, SAGSHL10, SDB10, WIM^{+}10, WR12]. Angeles [WSC^{+}17]. Angelsen [FDBA17]. Angiosperms [LFPT14]. Angling [DC11]. Anidación [HM12, Van12]. Anidan [LB11]. Animal [HMG^{+}13, BTBC^{+}16, CNK^{+}17, GDM^{+}19, HH19a, KBM17, WWWC15]. animal-welfare [HH19a]. Animales [KGS^{+}11, OBB^{+}11]. Animals [Ano16x, KGS^{+}11, MSF^{+}12, OBB^{+}11, BCLMM19, GLK^{+}17]. Annual [MMCUR^{+}12]. Anomalías [BPW12]. Anomalies [BPW12]. Anómalos [RER10]. Anomalous [RER10]. Años [EDB12, MQAMWM12]. Ant [EVF13, FG12, SKCS11, UMK^{+}16]. Antarctic [MMCUR^{+}12]. Antarctica [HICF16]. Antártico [MMCUR^{+}12]. ante [HHA^{+}11]. Anthropocebo [Jac13]. Anthropocene [CDF^{+}12]. Anthropocentric [HRL14]. Anthropogenic [BBP12, ESCF12, HCLW13, HSAB10, HMB^{+}13, MAJ16, PVP^{+}13, RMKM14, SHWT14, WH13, BMF16, CTBR19, DP17, GML^{+}19, SA15, WK15]. Anticompensación [BBBG10]. Antiguos [PTN10]. Antipoaching [VTT^{+}17]. Antropogénico [BBP12]. Antropogénicos [ESCF12]. Ants [FG12, MMPF10, SKCS11, HCH17]. Anual [MMCUR^{+}12]. Anuran [CLRV11, EPD12]. Anurans [FBA^{+}13]. Anuros [CLRV11, EPD12]. Anzuelo [LGK^{+}11]. Apalaches [LPT12]. Aparente [WJ10]. Apex [HJMM14, RBF15]. Aphelocoma [CFBL10]. Aplicación [HNS^{+}11, PHFPP10, WW10b, MA10]. Aplicada [AG10]. Aplicado [NKB10]. Appalachian [CS13, LPT12]. Appalachians [LP13]. Apparent [WSEM13, FWH19]. appeal [GG19]. Application [Ben12, ER14, GSDE14, HNS^{+}11, HCLW13, JHV13, ML14, MA10, PMJ^{+}13, PS13, PHFPP10, PWKB19, SLGB15, WSLW16, WW10b, GMNO15, HWMC15, NMS^{+}18, WML^{+}19]. Applications [LK14, MAB^{+}13, DAA^{+}19]. Applied [AG10, NKB10, BPH^{+}19]. Applying [ARDG17, BCK^{+}19, LdBB^{+}17, SBR^{+}19, RNM^{+}18]. apportion [JRWA15]. Appraisal [MAGA10]. Appreciating [Sha15b]. Approach [CKR^{+}13, CT11, ESCF12, GFM^{+}13, HP14, LOME10, ML14, MC11b, NBNMG13, PA12, PCRD14, SM14, VPS^{+}14, WHB^{+}14, AdBR15, BTM^{+}15, BCA^{+}19, CCS^{+}16, DDv18, FSR^{+}18, JGL^{+}16, MRFvW16, SC19, SKLP16, TMM^{+}18]. Approaches [LW14, Len17, PWRM10, Plo17, Qui14, FTB^{+}16, PNSG18, PHJ^{+}17, PBB^{+}17, WW16]. Approaching [Tru18a]. Appropriate [PCTW10]. Approval [MSZB^{+}14]. Aprendidas [PCTW10]. Aprendizaje [CCBF12]. April [Ano17a]. Aquaculture [LCYJ19]. aquarium [MRPT19]. aquariums [MJG15]. Aquatic [BPA^{+}13, BFHP10, DC11, HMD^{+}10, MAB^{+}13, KCD^{+}17, MDD^{+}18, PTJ^{+}19]. Arañas [GZR12]. Árbol [ZYYH10]. Árboles [SS10a, TAG^{+}10]. Arbustos [CBM^{+}11]. Archaeological [LCT^{+}14]. archaeology [WHE^{+}17]. Archeology [RL13]. Archipelago [ETB10]. Archipiélago [ETB10]. arctic [Pyk17, LSK^{+}15, WBF^{+}19]. arctic-alpine [Pyk17]. Area [AGLRV14, BE10, FJP^{+}11, HBR^{+}14, Hic10, KG10, MGCL14, MRBL^{+}12, ML10, PZM14, Sch14b, SPPN^{+}12, UCFV10, WEC^{+}11, BTM^{+}15, BC16, BLM^{+}16, CVMM17, GDB15, KJWP19, LMJB^{+}19, LTN18, MCM^{+}19, MV18, MVVS15, MBFB18, QKC^{+}19, VMO^{+}18, WHS18, WMFB15, CGCL11, HMIE14, WJ13, Hic10, BE10, CGCL11, KG10, SBT^{+}11]. areal [MBFB18]. Areas [AME^{+}13, ACL^{+}12, BMM^{+}14, BPA14, BLC13, BRCK11, BRKP12, DC11, GJB14, GBMRG^{+}10, GDMF13, HHA^{+}11, KMTK13, LMC^{+}14, LK13, MJ11, NA13, PCTW10, RM11, SMK14, SG10, Sou10, SSD10, UDG10, WT10, WRG^{+}12, ZM12, ANE^{+}16, AHP^{+}15, AGFR18, AC16, Arm18, BCD^{+}16, Ben15c, BBM18, BMAS17, BVAC17, BGP16, BMG^{+}17, CESS^{+}18, CY17, CEdK18, CMG15, CC17, CC18, DSG^{+}17, FBWIO19, GH18, GPB^{+}15, HKFC15, HLE19, HICF16, KTBH17, KHD^{+}15, LLR^{+}15, MAM^{+}15, MBC^{+}16, MAJ16, OHHE16, PCF^{+}19, QKC^{+}19, QCW^{+}18, RBF15, SGB^{+}19, SCA17, SL15, SCHM18, TMM^{+}18, TWB^{+}18, VKM^{+}17, WDV^{+}16, CKKL10, DBK^{+}14, GC10, MCN10, MO11, SSK^{+}11, WRAW10, WGF^{+}12, CKKL10, GC10, MCN10, MO11, SSK^{+}11, WRAW10, ZM12, ACL^{+}12, BRCK11, BRKP12, DC11, FWM^{+}12, FJP^{+}11, GBMRG^{+}10, HHA^{+}11, ML10, PCTW10, RM11]. Áreas [SPPN^{+}12, SSD10, UDG10, UCFV10, WT10, WEC^{+}11, WRG^{+}12]. Arenas [BB10a]. Argentina [MG12]. Argentinas [FG12]. Argentine [FG12]. Argument [KE10]. Argumento [KE10]. Arguments [GGM14]. Áridas [FMF10]. Ark [Arm16]. Arpón [FCGJH11]. Arrastre [HSBJ12]. Arrecifal [BPW12, LGK^{+}11]. Arrecife [Mcc10, NBW^{+}12]. Arrecifes [DFW^{+}10, LW11, MAA^{+}10]. Arrival [LZD^{+}13]. Arroyo [PCBD11]. Arroyos [RTL^{+}12]. Art [Don12]. Artesanales [AS10]. Arthropods [LCFS14]. Artículos [CB10]. Artificial [BS10, HSBE12, LTQ^{+}13, NS18, CZS19, RHR^{+}17]. Artificiales [BS10]. Artisanal [AS10, GD15]. Aseguramiento [MMM^{+}12]. Asentamiento [BV10]. Ash [ECES13]. Asia [BBM13, CPR^{+}16, DUS^{+}16, FTB^{+}16, GHT^{+}15, HSB^{+}16, JCCC14, LKYL15, Nov16, WSLW16, ZFW^{+}11]. Asian [YJL^{+}18, SKH13]. Asignación [RHM10]. Asistida [MMM^{+}12, MVB^{+}12, San10]. Ask [Mih19]. Asociación [MRPPM12, MJA^{+}11]. Asociado [JMP^{+}11, SVH12]. Asociados [CMJ^{+}11]. Aspen [AASK12, EDB12]. aspirations [VMO^{+}18]. assemblage [WMFB15]. Assemblages [AEO^{+}13, CMRH10, CGCL11, DMZ10, GMBM13, SSBN12, SMH^{+}14, SWSM19, TND16]. Assertions [BB11b, Fre11]. Assess [AWKR^{+}11, CWAW14, CM13, DBV13, ESCF12, JLDJ12, LW11, LHDC12, PVP^{+}13, WGJ13, CCS^{+}16, DKL^{+}15, LSA17, MRMH17, SMZZ17]. assessed [FFS^{+}18]. Assessing [BVC^{+}14, BR11b, BSL^{+}19, CB13, DPGY16, DPS^{+}16, EBRS14, Fre11, GRA14, GZJ^{+}13, HRHC16, HLE19, HICF16, KMGCR10, LB15, LGH^{+}12, MTM12, MBL^{+}14, MBG14, NBR^{+}15, NBNMG13, PCW14, RMM^{+}13, RK17, STM^{+}19, TG15, VJE^{+}15, MAB^{+}15]. Assessment [AWU^{+}11, CRNR12, Des19, HWW12, HMB^{+}13, LLW^{+}14, LCG10, MBBB17, MVB^{+}11, MBJWC10, MAGA10, MWC11, PRR^{+}11, PRE^{+}13, RHWS18, RKH^{+}14, SFM^{+}11, VMÁR^{+}13, ALA17, BLR18, CCN16, GBMW15, GHM^{+}15, JBH^{+}16, JWS^{+}18, OHHE16, PVO^{+}16, RTH^{+}15, SLGB15, SJD^{+}19]. Assessments [BMP^{+}12, BWE^{+}13, BJR^{+}11, DTH^{+}14, MA10, SML^{+}13, COZS16, RNM^{+}18, SBR^{+}19, WHK^{+}17]. assignment [MSY^{+}15]. Assisted [MMM^{+}12, MVB^{+}12, San10, Pyk17, WW17]. Associated [CRKM^{+}14, CMJ^{+}11, ECES13, JMP^{+}11, PGC^{+}13b, SVH12, FBWIO19, FHK^{+}15, LBB^{+}16, PAMA17]. Association [MJA^{+}11, MRPPM12, SBB^{+}15]. Associations [AB13b, MRBL^{+}12]. Assumption [PNSG18]. Assumption-versus [PNSG18]. at-risk [BC16]. Atención [HAB^{+}11]. Atlantic [RMKM14, BFF^{+}16, BBB14, KT16, LFPT14, MSMCB18, MPS^{+}16, SRG^{+}15]. Atracción [HBE^{+}10]. Atributos [Ber12, FB12, FP10, SGS^{+}10]. attendance [WFC^{+}18]. Attitudes [CD13, GDMF13, Heb12, KKK14, SBB^{+}14, TNTS13, DSE15, EAPB15]. Attractiveness [HBE^{+}10]. Atún [GOR^{+}12]. Auctions [KWW^{+}14]. Audience [JKJ^{+}19]. augmentation [KGF^{+}18]. August [Ano11a]. Aulas [WD10]. Austral [Arm16]. Australasian [YJL^{+}18]. Australia [Fie15, AWU^{+}11, CVMM17, HKFC15, IBT18, Moo17, Moo19, RBD^{+}13, RB17, WEC^{+}11, WHG^{+}15, WKF^{+}11, YFB15, Fie15]. Australian [PB11, GBB^{+}19a, HH19c, LZR^{+}18, MGG^{+}13, STJ^{+}19, WGLL17]. Australiana [PB11]. authentic [San19]. author [Dar15]. automated [vdWSM^{+}16]. autonomous [BMB^{+}19]. Availability [LCG10, SBU^{+}17]. Avant [BLG^{+}11]. Avant-Garde [BLG^{+}11]. Avanzando [SH11b]. Ave [EG12, MARV^{+}11, Van12]. aversion [TMJ^{+}15]. Aves [BAPP11, BE10, Boy10, CMRH10, EDB12, GS11, GSW10, HM12, HRSK12, HB10, IUM10, Jam11, KCSP11, KTN12, LHDC12, LB11, LDBS11, PAM^{+}10, RPA^{+}12, SA11, SPSS10, SWL^{+}10, WGB^{+}11b, WKF^{+}11, GZA10]. Avian [BAPP11, Boy10, DHD14, EDB12, EBRS14, Jam11, LHDC12, LZD^{+}13, MLA^{+}14, Ore12, RPA^{+}12, SA11, SWL^{+}10, ZLKW13, PCF^{+}19]. Avifauna [NS10, SMGR17]. avifaunal [CWH^{+}17]. Avistamientos [Duf11, SSB^{+}12]. Avoid [KMTK13, EAPB15]. Avoidance [TSH^{+}14]. Avoided [HRBR11]. Avoiding [FBF^{+}14, HP14, VWM^{+}10, MGA16]. Awards [Ano17b, Ano19c, Ano18c]. Awareness [Tru18a, DKL^{+}15]. Away [JCCC14]. axiom [VBN15]. Azules [WHB^{+}11].

B [MAWF12, OSFFB^{+}11]. Back [McP11, RCR^{+}12]. background [VJG^{+}15]. bacterium [EB15, GGB^{+}15]. Bahía [WKF^{+}11]. Baisre [MCH^{+}16]. bait [NJBM15]. Baja [SHLK^{+}10]. Bajo [BBBG10, CVN^{+}10, CBS11]. Balance [Wat11, TXBW19, CKKL10]. Balanceando [Bro10, BEE^{+}11]. Balanced [SSO12]. Balancing [BEE^{+}11, Bro10, Sto12, Zha15]. Ballenas [WHB^{+}11]. Banff [CRG12, HWM10, LP10, SCK13]. Bangladesh [BGAS10]. Bank [JMSM13]. banks [GBD^{+}15]. barcodes [WSLW16]. Barcoding [HJSF^{+}12, GSB^{+}16, WSC^{+}17]. Barcos [DDPC12]. Barda [FEC^{+}10]. Barras [HJSF^{+}12]. Barrera [BPW12, LGK^{+}11]. Barrier [BPW12, CPM^{+}18, LGK^{+}11]. barriers [RCSK15, WHS18]. Basada [LW10, NFF^{+}10]. Basadas [WFM^{+}10]. Basado [ESCF12, GCC^{+}12, GPK^{+}10, MJP^{+}11, MC11b, SCL12]. Base [BV10, NBW^{+}12, RTL^{+}12, UDG10]. Based [BCC^{+}14, EIS14, ESCF12, GCC^{+}12, GPK^{+}10, KRG13, KWW^{+}14, LW10, MJP^{+}11, MC11b, NFF^{+}10, RGRMPB^{+}13, SCL12, SMGRJ10, WFM^{+}10, AG17, AMS^{+}16, BSH^{+}15, BCR^{+}17, BRA^{+}19, BC16, BCSM15, CCCCN18, COZS16, DSG^{+}17, DCP^{+}19, FDBA16, HCH17, HPD16, HLE19, HTH^{+}17b, IRK^{+}18, JMH15, KSH17, LDC^{+}15, MBBB17, MGMC15, MLMP^{+}15, MSY^{+}15, PS15, Pet17, PNSG18, PRNV18, QCW^{+}18, RHR^{+}17, SMF^{+}15, TMM^{+}18, VGAB^{+}19]. Baseline [BGL^{+}14, AC15]. Baselines [Bai13, BV10, BT14, Cla14, NBW^{+}12, TBY^{+}10, WMW^{+}14, PRNV18]. basic [CCB^{+}19a, PVO^{+}16]. Basin [AWU^{+}11, IUM10, MPH^{+}11, WBM13, EDB12, KHGF^{+}17, NMSF11]. Basis [NCC13, UDG10]. Bat [GMBM13, ISA16]. Batrachochytrium [BVC^{+}14, BGH^{+}15, CLRV11, GLH12, MSSM09, OSFFB^{+}11, PGC^{+}13b, SGH^{+}11, SMH^{+}14]. Bats [ATW^{+}11, FCC^{+}11, HM11, SJH13, DMR^{+}17, GLF^{+}18, MF15, SC15]. Bawa [KNK13]. Bay [WKF^{+}11]. Bayesian [ALW^{+}19, BMAS17, CJDZ10, HMM10, JGL^{+}16, SC19, VPMS13]. Bayesian-weighted [SC19]. Bayesianas [HMM10]. Bayesianos [CJDZ10]. be [MBT^{+}17a, RHM10]. beak [FMB^{+}18]. Bears [CRG12, FWM14, Med10, DSE15]. Beckman [Joh11]. Bee [MBGS11, TDCG15]. Beech [GLB^{+}13]. Beehive [KLN^{+}17]. Bees [Gro10]. Beetle [CD13, Fat14, BP16]. Beetles [CMJ^{+}11, EMD^{+}10, GLB^{+}13, SBB^{+}15]. Behavior [Bro10, FCGJH11, Mih19, MSSM10, Sch11, VMW13, BTBC^{+}16, DKL^{+}15, NGB^{+}16, NBA^{+}18]. Behavioral [ESCF12, JGH^{+}11, EEB15, WHEN^{+}16]. Behaviors [CD13, GLK^{+}17]. behind [JD16, LAA^{+}15]. Being [LDBS11, MGMA^{+}14]. belowground [CML^{+}19]. Belt [Hug19]. Beluga [DBP^{+}10]. Ben [MMGN18]. Beneficios [AT11, BMP^{+}12, DEJM10, BR11a]. Benefit [PSD^{+}14, SA14]. Benefits [AHT^{+}15, BR11a, BMP^{+}12, CNvL^{+}15, DEJM10, EBRS14, GBP^{+}13, JVT^{+}13, LCL13, DSG^{+}17, FWC^{+}15, HKFC15, HMB^{+}15, JHAH18, MK18, NWD^{+}16a]. Berger [Joh11, vWWCF15]. Best [BSBM11, LW14, SDT17]. beta [RDS17]. beta-diversity [RDS17]. Better [PMJ^{+}13, Poz17, GMG^{+}19, Had15, HBTMG15]. Between [PFR^{+}14, AP17, ATW^{+}11, BJ16, BB11a, BIF^{+}18, BT14, BRKP12, CZS19, CKKL10, Dre17, DJBB16, Fat14, FMV^{+}19, GRL15, GT10, GLK^{+}17, HNO^{+}18, HRKMG12, LKW^{+}15, LDBS11, MSMCB18, MG12, MCL12, MRMV14, MAWF12, PBH^{+}10a, Pyk17, Rey11, RM14b, RBF15, Rob12, SFAA^{+}18, SMARC^{+}16, SCWN18, SMLB19, TG13, TXBW19, TQD^{+}18, TAP19, VNAT19, VCDF13, WWH^{+}18, Wat11, WWW^{+}12, vECD^{+}18]. Beyond [BC14a, Cla13a, JCA^{+}11, ZBM^{+}10, CBB18b, SBS^{+}15, TAB^{+}18]. Bhutan [Bro10]. Białowieża [SSD^{+}18]. Bias [CW10, MBG14, OB11b, VMO^{+}18, BJM15, FWH19, HHJ^{+}19, MCMP15]. Biased [CWS^{+}14, Pin17]. Biases [DMM14, KJWP19, TBC18]. Bibliometrics [Pau16]. Bienestar [LDBS11]. Big [Bro13, SLH10]. bighead [WCR^{+}15]. Bighorn [JMW^{+}11]. Billed [GCC^{+}12, SSB^{+}12]. Binarios [FP10]. Binary [FP10]. Bioactividad [LAA^{+}10b]. Bioactivity [LAA^{+}10b]. bioblitz [PPM^{+}18]. Biocentric [HRL14]. biocontrol [PBR^{+}16]. Biodegradable [BHSA12]. Biodiversidad [DFW^{+}10, DK10, ERC^{+}10, GBG^{+}10, JCA^{+}11, KG10, LNP10, LOME10, MB10, MAGA10, PBH^{+}10a, SHO^{+}10, WRAW10, WW10b, MPB10]. Biodiversity [BMM^{+}14, BME^{+}14, BGL^{+}14, DFW^{+}10, DK10, DMF^{+}18, Des19, ERC^{+}10, EMJ14, Fat14, Fer19, GBG^{+}10, GHB^{+}13, HFSB13, HTH^{+}10, HDH^{+}10, JIAMS^{+}14, Joh10b, JCA^{+}11, KG10, LNP10, LOME10, MB10, MMK^{+}14, Mcl11, MQ14, MPB10, Moo17, Moo19, MJG15, MDJ14, MAGA10, PBH^{+}10a, PBH^{+}10b, PFA^{+}14b, PFR^{+}14, QM11, RSS^{+}13, RGMR12, Rey11, RBD^{+}13, SABV13, SPN^{+}14, SCSJ13, SS10b, SHO^{+}10, SA14, UIN^{+}13, Ver16, Vit17, WRAW10, WKI^{+}14, WW10b, Zha15, ABMG19, BBM15, BC15, BRA^{+}19, BB15, BMC^{+}18, CML^{+}19, CS17a, CQC^{+}19, CBCW18, CVA15, DTM16, FCS19, FSSW19, FHK^{+}15, GCAF^{+}18, GD15, GS16, GBB^{+}19b, GBBMG19, HCH17, Hen16, HGRS16, HHW^{+}16, IBA16, ISM^{+}18, IO17, JBB^{+}15, JWS^{+}18, JVI^{+}18, KCD^{+}17, KAM18, KBCM15, KRF^{+}19, KK18, LHKA16, LKYL15, LWZL16, LCC^{+}15, MAM^{+}15, MTD16, MSMCB18, MBC^{+}16]. biodiversity [MBV^{+}16, MR16, MGW^{+}15, NM18, NW15, PR17, RB17, SFAA^{+}18, SSRS18, SCWN18, SSW19, SL15, SG18, SJD^{+}19, SURW18, SKPP15, TQD^{+}18, TLT16, VMO^{+}18, WWH^{+}18, WGS^{+}19, vBSC^{+}19]. Bioenergía [HDH^{+}10]. Bioenergy [HDH^{+}10]. Bioenergy-Related [HDH^{+}10]. Biofluorescence [BHAR^{+}18]. Biogeographic [BA13, DP15]. Biogeographical [DMM14]. Biogeography [SNH^{+}14]. Biogladiator [Hel14]. Bioinformatics [Dre11]. Bioko [VCDF13]. Biología [CB10, LH10, WD10, OSC10]. Biológica [GKK^{+}11, Rud11b]. Biological [Arm16, BRS11, BCD^{+}16, CKR^{+}13, Cor16, FB12, GKK^{+}11, HDHL10, JC18, KMTK13, LAA^{+}10a, MMG^{+}13, NCC13, PG12, RL13, Rud11b, SCA17, Whi13, WKI^{+}14, CCCCN18, HLE19, LS15, MBBH16, PS15, WWWC15, vdWSM^{+}16]. biologically [BCD^{+}16]. Biológicos [FB12]. Biologist [AA12, GHKK10, PI11, MMGN18]. Biologists [Gro10, LDB^{+}10]. Biólogo [GHKK10]. Biology [Con11, Die10b, Dre11, HP11, Jen14, LH10, MBG14, OSC10, Pau16, PMP14, RM14a, WD10, BH17, Cha15, HCBB^{+}15, TAB^{+}18, FBB^{+}14, Ano12e, Ano13a, Ano13b, Ano13c, Ano13d, Ano14b, Ano14c, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano18c, Ano18g, Ano18h, Ano18i, Ano19c, Ano19h, Ano19i, Ben15b, Bur13, BJM15, CB10, HRSK13, HDN^{+}14a, JMB16, LZD^{+}14, PFA^{+}14a, Red10a, SGH^{+}13, Sou14b, TAB^{+}18]. Biomasa [GOR^{+}12]. Biomass [GOR^{+}12, MGMC15, CESS^{+}18]. Biomass-based [MGMC15]. Biophilia [SS10b]. Biosecurity [Pau16]. biosentinel [KS17]. Biosphere [VLGRS14]. Biota [Bor19]. biotopes [BMB^{+}19]. Bird [AB13b, CRKM^{+}14, CMRH10, FGG^{+}19, GSW10, IUM10, KCSP11, KTN12, LDBS11, LP13, MRBL^{+}12, MG12, Møl17, PAM^{+}10, PFR^{+}14, SPSS10, Van12, VPS^{+}14, WGB^{+}11b, WFMW^{+}18, AHT^{+}15, BHF^{+}19, DPS^{+}16, FBWIO19, GSB^{+}16, KOA^{+}15, LBB^{+}16, MBC^{+}16, ML15, PAMA17, SB16, SSMC17, VJE^{+}15, ZRM18]. Bird-community [WFMW^{+}18]. bird-conservation [SSMC17]. bird-migration [PAMA17]. birder [SSMC17]. birding [WZC19]. Birds [BPA^{+}13, BE10, BKM14, Cla11, Die10a, EWN^{+}13, GZA10, GS11, Gro10, HM12, HB10, LZD^{+}13, MBL^{+}14, MC14, OPP14, RCK13, Sco16, TSH^{+}14, Vég18, Zac10, EKRW17, ERW18, GBB^{+}19a, HTH^{+}17a, YJL^{+}18]. Bison [DHH^{+}12, RAP16]. Black [AB13b, CWS^{+}14, CRMK18, WLCS14, GLdPL18]. Black-Capped [WLCS14]. Black-Tailed [AB13b]. Blanca [FCC^{+}11, HM11]. Blando [CCBF12]. blind [Dar15]. blindness [BW16]. bloom [BBH17]. Blue [Bar12, WHB^{+}11, HNO^{+}18]. bluefin [SRG^{+}15]. Bluegrass [MMCUR^{+}12]. Boats [DDPC12]. bodied [KSH17]. Body [Ber12, MJA^{+}11, WGR^{+}10, HSV^{+}17, MSY^{+}15]. Body-Size [Ber12]. Bolder [LFMB19, NDB^{+}12a, WDV^{+}16]. Bolivia [PDM11]. Bolivian [PS15, TYA16]. Book [Ano14a, Ano18a, Ano19a, Bra15, CB10, Fie15, Kra12, Löv16, Red10a]. Books [Ano11a, Ano11b, Ano14a, Ano15l, Ano16a, Ano16b, Ano17a, Ano18b, Ano19b]. Boomtowns [BB11b]. Borde [GGG10, PGLB^{+}12]. Border [FEC^{+}10, SK10]. Borderlands [SK10]. Borders [Cla13a]. Boreal [MTN13, PMS^{+}10, Pyk17, SAS^{+}19]. Boreales [PMS^{+}10]. born [WC19]. Boscosos [EMD^{+}10]. Bosque [BEE^{+}11, DWH^{+}11, GS11, HG10, HRBR11, MPB10, Pra10, SMGRJ10, TAG^{+}10, Van12, BMGLW12, LW10]. Bosques [CMRH10, EDB12, ETB10, IP11, IUM10, LB11, LGB^{+}10, MLS12, PBH^{+}10a, SAC12]. Botanic [HFL^{+}11, Hul15]. botanical [CBF^{+}19, HICF16]. Botánicos [HFL^{+}11]. Botella [BAJ^{+}10, HB10]. Both [Whi13]. Bottlenecks [BAJ^{+}10, HB10]. Bottom [GFM^{+}13, HSBJ12]. Bottom-Trawl [HSBJ12]. Boundaries [BS12, FJP^{+}11]. Boundary [CMS^{+}13, NRD^{+}16, PWP^{+}18]. Bowne [LS19]. Brad [LBD^{+}19]. Bradshaw [Hen16, HGRS16]. Brahmaputra [SMS14]. Brasil [TAG^{+}10]. Brasileña [RSE12]. brasiliensis [ATW^{+}11]. Brazil [BMM^{+}14, BPA14, TAG^{+}10]. Brazilian [ATW^{+}11, LFPT14, NFdC^{+}12, RSE12]. Breaking [LSK^{+}14]. Brecha [JDA^{+}10]. Breeding [FWM14, HFK^{+}10, KPL^{+}12, MS12, Rog11, RKH^{+}14, WH13, Alr15, AMS^{+}16, HGP16, JC18, PRB18, QSCM^{+}19, SB16, SMF^{+}15]. Bridge [ATW^{+}11, MQ14, PGBV^{+}13]. Bridge- [ATW^{+}11]. Bridges [CMS^{+}13]. bridging [RIBT18, SLA^{+}19]. Bringing [RCR^{+}12]. British [Sco16]. Broad [BMG^{+}14, Put11]. Broad-Extent [BMG^{+}14]. Broadening [HAV11]. broader [BCA^{+}19]. Bromeliads [CT11]. Bromelias [CT11]. Brook [Hen16, HGRS16]. Brown [Ano16x, OJAV15]. Brussard [Pea18]. Bubble [Tor11]. Buddhism [Dor11]. Buddhist [LWY^{+}14]. budget [DHG^{+}17, Dre17]. budget-constrained [DHG^{+}17]. Buffelgrass [FMF10]. Buffers [EBRS14]. Bug [ER14]. Build [BW10]. Building [RSWM14, SBN^{+}10, VJ15, CLM^{+}15]. Buitres [OTKE12]. Bumble [MBGS11]. burden [CBPM19]. burdens [Arm19]. Bureaucracies [HSRC14]. Burgman [LET19]. Burmese [ÇRG^{+}17]. burn [BHO^{+}18]. burned [PR17]. Bury [VM18]. Bush [Har15]. Bushmeat [AWKR^{+}11, CAB^{+}10, FFM^{+}14, FWM^{+}12, KMGCR10, MJA^{+}11, NJT14, NBNMG13, NGL^{+}14, RMGCR10, VCDF13, VMGB^{+}10, FOF^{+}15, MKR^{+}15]. Bushmeat-Harvesting [RMGCR10]. Business [BG10, GD15]. Businesses [MMG^{+}13]. Bután [Bro10]. Butterflies [Ano19w]. Butterfly [CFO^{+}14, HSU^{+}16, MTM12, MBJWC10, Pek13, VLGRS14, Whe11, BC16, DMB^{+}17, DBV^{+}18]. Bwindi [BMGLW12, BMGLW12]. Bycatch [WB14, MK18, NCKC17, SG18]. bycatch-reduction [MK18].

C [Ano19w, Bra15]. Cables [BAPP11]. Cabo [WRG^{+}12]. Cacería [AWKR^{+}11, KMGCR10, MRRK11, PBK^{+}11, VMGB^{+}10]. Cactaceae [NRRW17]. cactus [GDG19]. Cadenas [DDPC12]. Café [GBG^{+}10, MPB10, LW10]. Calcareous [RBW14]. calculating [JD16]. calculations [CS17a]. Calentamiento [Col10, WK10]. Calibrado [KG10]. Calibrated [KG10, PZM14]. Calidad [KUT11, LHDC12, RC12, SHLK^{+}10]. California [MWC11, ALW^{+}19, HSBJ12, KGG^{+}14, KPL^{+}12, MMPF10, RM11, RJF^{+}13, TG13]. Call [HDHL10, Had15, WZC19]. Cambia [FMJP11a, SKC^{+}12]. Cambiante [RSW11, SBN^{+}10]. Cambio [BB10a, GKK^{+}11, Han10, HFK^{+}10, HHA^{+}11, JMP^{+}11, JGH^{+}11, MRPPM12, Ser10, WIM^{+}10, WGB^{+}11b, WB10, WRG^{+}12, WGR^{+}10, BDH^{+}12, RPA^{+}12, WK10]. Cambios [BV10, Col10, DEJM10, FMJP11a, KG10, EDB12, MRPPM12, RSE12, TBY^{+}10]. Cambodia [CMG15, LBLB14, PGC^{+}13a]. Camera [Møl17]. Cameroon [MJA^{+}11]. Camerún [MJA^{+}11]. Caminos [KCSP11]. camouflage [DP17]. campaign [SMV19]. Campephilus [GCC^{+}12]. Can [BBBG10, Car14, LDB^{+}10, LGM^{+}13, PSD^{+}14, IPA^{+}17, RBE^{+}15]. Canada [BV10, BKM14, CRG12, LMCS17, BV10]. canadensis [JMW^{+}11]. Canal [HTDM17]. cannot [MBT^{+}17a]. Canopy [MPB10]. Canoras [LB11]. Cantidades [SDB10]. Caoba [GT10]. Capacidad [WEC^{+}11]. Capacity [WEC^{+}11, NBR^{+}15, WHK^{+}17]. Cape [HRC10, WRG^{+}12]. capital [DSG^{+}17, LZZ^{+}17, SPD18]. Capítulos [CB10]. Capped [WLCS14]. Caption [Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano12e, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano13d, Ano14b, Ano14c, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano18g, Ano18h, Ano18i, Ano18d, Ano18e, Ano18f, Ano19h, Ano19i, Ano19d, Ano19e, Ano19f, Ano19g, Ben15b, HRSK13, HDN^{+}14a, LZD^{+}14, NDB^{+}12b, PFA^{+}14a, RWB12, SGH^{+}13, Sou14b, WGB^{+}11a, Ano11c]. Captive [RKH^{+}14, Alr15, AMS^{+}16, DPGY16, HGP16, JC18, QSCM^{+}19, WC19]. captive-born [WC19]. Captive-Breeding [RKH^{+}14, AMS^{+}16, JC18]. Captura [KGS^{+}11, LGK^{+}11, RMGCR10]. Captura-Recaptura [KGS^{+}11]. Capturas [BJR^{+}11, MO11]. Capture [KGS^{+}11, WLCW16]. Capture-Recapture [KGS^{+}11]. captures [LCYJ19]. Características [MLS12, MRRK11, WR12]. Carbon [BMM^{+}14, CBS11, RM14b, MGW^{+}15]. Carbono [CBS11]. Care [SRS11]. Career [Con11, ZAP^{+}18]. Careers [BDG^{+}13]. Caribbean [Bai13, GML^{+}19]. Caribou [DWH^{+}11, HWM10, LP10, SW10, TW11]. Caribú [DWH^{+}11]. Carne [CAB^{+}10, FWM^{+}12, MJA^{+}11, RMGCR10, VMGB^{+}10]. Carnivore [BGAS10, DBV13, FC10, HSV^{+}19, SSD^{+}18]. Carnivore-Ungulate [FC10]. Carnivores [MBM^{+}14, Ord11, RM11, SRS11, DPT^{+}18, Har15, vECD^{+}18]. Carnívoro [FC10]. Carnívoros [BGAS10, RM11, SRS11]. carp [WCR^{+}15]. Carpathian [BMG^{+}17]. Carpintero [SSB^{+}12]. Carreteras [MFRH12, SDB10, CBSZ10]. Carroll [Wil12b]. Carroñeros [OTKE12]. carry [DMR^{+}17]. carry-over [DMR^{+}17]. Cascabel [CBSZ10]. Cascadas [SGS^{+}10]. Cascades [HJMM14, SGS^{+}10]. Case [ACB^{+}15, CLM^{+}15, CPC^{+}10, DWCW10, EMD^{+}10, HHDM10, MAA^{+}10, WHB^{+}11, GV16, MF15, Ros15, RKC^{+}19, VTT^{+}17]. Cases [BHK10]. Cashmere [BBM13, BBM15, vWWCF15]. Caso [CPC^{+}10, DWCW10, EMD^{+}10, HHDM10, MAA^{+}10, WHB^{+}11]. Casos [BHK10]. caspia [MPH^{+}11]. Caspian [MPH^{+}11]. cat [LM18, WLL^{+}16]. cat-related [WLL^{+}16]. catastrophic [BBH17]. Catch [LGK^{+}11, RMGCR10, WB14, YFB15]. Catches [BJR^{+}11, MO11]. Catching [Öst14]. Categorías [FB11]. Categories [FB11]. Categorizing [FSAM15, RTT^{+}13]. Catenazzi [FMJP11b]. Caterpillar [WNY13]. Caterpillar-Host [WNY13]. Cats [LCL13, LL14, GTVC16, LSÁL^{+}19, WLL^{+}16]. Cattle [BKM14, DHH^{+}12, MG12, GTC^{+}16]. caught [LH19]. Causa [BFHP10]. Causada [PHC^{+}11]. Causal [SJP^{+}13]. Cause [BFHP10]. Caused [HSAB10]. causes [BHF^{+}19, LPA^{+}17, NCKC17]. Causing [KKK14]. Cautioning [YNR19]. Cautious [CSB^{+}10]. Cautiverio [KE10]. Cave [ATW^{+}11]. Cave-Roosting [ATW^{+}11]. Caves [LFMRD12, Whi10]. Caza [FWM^{+}12]. Cazadores [RMGCR10, VMGB^{+}10]. CB [Ano12g]. celebrates [JMB16]. Celebration [Sco16]. Censos [JDA^{+}10]. Censuses [JDA^{+}10]. Centered [HP14]. Central [FWM^{+}12, SKSS^{+}17, VKKA13, VTT^{+}17, GGL^{+}15, BBM13, CAB^{+}10, CSMG^{+}13, DMZ10, HRKMG12]. Centralidad [CMB12]. Centrality [CMB12]. Centrocercus [BBP12]. centuries [HSU^{+}16]. Century [For11, SPL^{+}11, KMLC19, LSK^{+}15, RV15, TBM^{+}16, Car14]. Cerco [BW10, GOR^{+}12]. Cerrar [JDA^{+}10]. certainty [BSCB18, CBB18a]. Certificación [BR11a]. Certification [BR11a]. cetacean [MRMH17]. Chaco [MG12]. Chains [DDPC12]. Challenge [BCC^{+}14, CGM10, GHB^{+}13, KHD^{+}15, MBT^{+}17b, PR17]. Challenges [BMG^{+}14, EMS^{+}11, GBG^{+}10, Jos12, LW10, OKLP11, PMD^{+}13, RS18, SCL12, Val18, YJL^{+}18, CPR^{+}16, CHN^{+}15, FTB^{+}16, GMV^{+}18, GH18, KBML15, LAA^{+}15, WBB^{+}18]. Change [AH11, AL14, BB10a, BB11a, BWE^{+}13, BDH^{+}12, BGG13, CLT^{+}13, Cla11, CCBP14, CSMG^{+}13, CMG^{+}13, Dol10, Ehr14, GJB14, GKK^{+}11, GFRK12, GRA14, Han10, Han11, HFK^{+}10, HH14, HHA^{+}11, HDHL10, JMP^{+}11, JGH^{+}11, JMSM13, KMH^{+}14, KG10, Lal11, MRPPM12, NMMP14, NLM^{+}13, PS13, PS10a, RPA^{+}12, RP13, SRL^{+}13, Ser10, SML^{+}13, Vég18, WBM13, WIM^{+}10, WGB^{+}11b, WB10, WRG^{+}12, WGR^{+}10, ARDG17, BCR^{+}17, BMF16, CBM^{+}18, HGRS16, JBB^{+}15, KLL^{+}16, LAA^{+}15, LMM^{+}17, NBR^{+}15, NWW^{+}17, PCF^{+}19, RCSK15, SMAG17, TGBB^{+}15, TBC18, WHK^{+}17]. Changed [Ano19w, MBT^{+}17a]. Changes [BLC13, BV10, BSF^{+}14, Col10, DEJM10, EDB12, EPB^{+}14, FMJP11a, HSAB10, MCM^{+}19, PAMA17, RSE12, SKC^{+}12, FBWIO19, MKR^{+}15]. Changing [CFO^{+}14, GAG13, KTHH10, MNW^{+}14, RSW11, Sel16, Sha15a, SBN^{+}10, Sta16, Ben15c, HM16, SSWR17, TWB^{+}18]. Chapters [CB10]. Characteristics [HAP^{+}13, LTQ^{+}13, MLS12, GBB^{+}19b, LS15]. Characterizing [CPB14, HTH^{+}17b, LDD19, LRA^{+}14, VW19]. cheap [Arm18]. Chemical [Pek13]. Chicago [Ano16x, Ano19w]. Chichester [Ano19w]. Chickens [MHG^{+}14]. Children [Jen14]. Chile [GFG^{+}12, OKCG18]. Chilean [WHB^{+}11]. Chilenas [WHB^{+}11]. Chimera [PMMG14]. China [CZS19, HHY13, KOA^{+}15, LP16, LWZL16, LLR^{+}15, LCC^{+}15, LRL^{+}15, LZZ^{+}17, MCM^{+}19, RYW^{+}15, SLWL15, Zha15]. Chinese [Ano18x]. Chinook [HSBE12, JMSM13, WBM13]. Chng [Nat18]. Choice [NJT14, SSMC17]. Choosing [FLW14b, McP12, SML^{+}13]. Christmas [EMJ14]. Chronic [BBP12]. Chytrid [GLH12, MSSM09, PMS^{+}10, Sch10, SSP^{+}15]. Chytridiomycosis [MSSM10, PGC^{+}13b, SHG^{+}14]. Chytridiomycosis-Threatened [SHG^{+}14]. Cichoń [VM18]. Ciencia [HF10, HFL^{+}11, HS10, JGH^{+}11, MBF^{+}12]. Científico [TA10, VLRB12]. Científicos [IP11, Rud11b, GZD11, KUT11]. Cierre [Mcc10]. ciliare [FMF10]. Circle [LSK^{+}14]. Circuit [DAA^{+}19]. Circuit-theory [DAA^{+}19]. Citación [OB11b]. Citas [CB10]. Citation [OB11b, PFN15b, CV16, MCMP15, Oso15]. Citations [CB10]. CITES [RS18, SASO13, ZPJ19]. CITES-listed [RS18]. Cities [MTM12, SCSJ13]. Citizen [DSW16, IP11, JGH^{+}11, KUT11, MTM12, TLT16, BPY^{+}16, BW19, CL16, DMB^{+}17, HPD16, IBT18, JGS^{+}16, LPW16, MCD^{+}16, SKLP16, TRPD16, VLSOS16, WLD16, vdWSM^{+}16]. Citizen-Scientist [KUT11]. City [MF16, SH11a, HCT^{+}17]. Ciudadana [JGH^{+}11]. Ciudadanos [IP11, MTM12, KUT11]. Ciudades [MTM12]. Civic [DSW16]. Civil [TW14]. Clandestina [BM10]. Clandestine [BM10]. clarify [BMAS17]. Clarifying [EAPB15, RM19, SVN^{+}15]. Clasificación [GKK^{+}11, NKB10, WKP11]. Clasificar [DWCW10]. Classic [Ano16x]. Classical [BA13]. Classics [Cla13b]. Classification [GKK^{+}11, WKP11, CCS^{+}18]. classifications [RCARR18]. Classrooms [WD10]. Clave [SGS^{+}10]. Clayoquot [LMCS17]. Clearing [KZG^{+}13]. Clima [FMJP11a, HHDM10, LSBB10, LPT12, RSW11, RTL^{+}12, SKC^{+}12, WR12]. Climate [ACS14, AH11, BB10a, BLC13, BSF^{+}14, BWE^{+}13, BDH^{+}12, BGG13, CLT^{+}13, CBM^{+}18, CFO^{+}14, Cla11, CMG^{+}13, Dol10, EPB^{+}14, FMJP11a, GKK^{+}11, GRA14, GAG13, Han10, Han11, HHDM10, HFK^{+}10, HH19c, HHA^{+}11, HDHL10, JMSM13, KMH^{+}14, KTHH10, Lal11, LSBB10, LPT12, MAB^{+}13, MNW^{+}14, MRPPM12, NLM^{+}13, PS13, PS10a, RSW11, RP13, RTL^{+}12, SRL^{+}13, Ser10, SKC^{+}12, SML^{+}13, Vég18, WBM13, WK10, WGB^{+}11b, WB10, WR12, WRG^{+}12, ARDG17, BBM15, HM16, HGRS16, JBB^{+}15, KLL^{+}16, LMM^{+}17, NWW^{+}17, PCF^{+}19, RCSK15, SMAG17, TGBB^{+}15, TBC18, TBF^{+}17, TWB^{+}18, WHK^{+}17]. Climate-Change [AH11, Dol10, Han10, Lal11, SML^{+}13, JBB^{+}15, KLL^{+}16, SMAG17, WHK^{+}17]. Climate-Change-Induced [GKK^{+}11, RCSK15]. Climate-Induced [EPB^{+}14, RTL^{+}12]. Climate-Smart [HHDM10]. climates [LMM^{+}17]. Climatic [WGF^{+}12, DBV^{+}18, MLRC18, WWH^{+}18]. Climático [BB10a, BDH^{+}12, GKK^{+}11, Han10, HFK^{+}10, HHA^{+}11, MRPPM12, Ser10, WK10, WGB^{+}11b, WB10, WRG^{+}12]. Clonal [Tep12]. Cloning [GGM14]. Closures [FCGJH11, Mcc10]. Clusters [Duf11]. Co [GC10, HPL14, RJF^{+}13, Jør15, LDD19, NRD^{+}16]. Co-Manejo [GC10]. Co-Occurrence [HPL14, LDD19]. Co-Occurring [RJF^{+}13]. co-production [NRD^{+}16]. co-reintroduction [Jør15]. coalitions [CLM^{+}15]. Coarse [BRV10, GBB^{+}15]. coarse-filter [GBB^{+}15]. Coarse-Resolution [BRV10]. Coast [PB11]. Coastal [CKK^{+}14, GFG^{+}12, GPK^{+}10, AC15, BFB^{+}19, CE15, CWH^{+}17, TT16]. coastlines [OGS^{+}19]. Cobertura [FG12, JLCK12, RPA^{+}12]. Cocos [WMFB15]. Código [HJSF^{+}12]. cœrulescens [CFBL10]. coexist [AJ18, PA18]. Coexistence [KKCS10, CGK^{+}19, HSV^{+}19, WW16]. Coexistencia [KKCS10]. coexisting [KLN^{+}17]. Coextinción [MVB^{+}10, MVB^{+}11]. Coextinction [MVB^{+}10, MVB^{+}11]. Coffee [GBG^{+}10, HNDE13, LW10, MPB10, RM14b]. cognition [CRMK18]. Coherent [Ric11]. Colaborativa [LSDK11]. Colapso [CLRV11]. cold [AAvRS19]. Colecciones [SS11]. Colecta [CT11]. Colisiones [BAPP11]. Collaborating [Sch14a]. collaboration [HSV^{+}19, vBSC^{+}19]. Collaborative [GFM^{+}13, LSDK11, CVA15]. Collapse [CLRV11, BIP^{+}15, CFB^{+}17, KOA^{+}15, RNM^{+}18]. Collected [MFRH12, CPD^{+}15, HOC^{+}16]. Collection [CT11]. Collections [SS11, DPGY16]. collective [GMG^{+}19, TLT16, VB19]. Collen [MMGN18]. Collisions [BAPP11]. Colombia [NAB^{+}17, OPP14]. Colonias [HRSK12, JDA^{+}10]. Colonies [HRSK12, JDA^{+}10]. Colonist [LGB^{+}10]. Colonización [FG12, MMM^{+}12, San10]. Colonizadores [LGB^{+}10]. Colonization [FG12, MMM^{+}12, ML15, Pyk17, TMK^{+}16, WHE^{+}17]. Colonization. [San10]. Colony [BMA^{+}14]. color [TMD^{+}17]. Columbia [HSBE12, MPH^{+}11, HSBE12, MPH^{+}11]. Colville [Fie15]. Comanagement [GC10, WJ13, SMY^{+}17]. combatting [FSR^{+}18, MSR19]. Combinación [LH11, MR12]. Combinados [CGCL11, WGB^{+}11b]. Combined [CGCL11, JMSM13, SJD^{+}19, WGB^{+}11b]. Combining [KWW^{+}14, LH11, MR12, Qui14, TBF^{+}17]. Comercial [WGR^{+}10]. Comerciales [NS10]. Comercio [BM10]. Comes [Gon13]. commensal [GGB^{+}15]. Comment [CRNR12, KNK13, Sch10]. Commentaries [Ano16x]. commerce [HHKK15]. Commercial [Lor13, NS10, VCDF13, WGR^{+}10, Har15, JRWA15, LGCF19, NJBM15]. Commitment [AB13a]. Common [GKP13, GPK^{+}10, HM12, Kar14, Rob12, Sou14a, GGB^{+}15, MA19, SMLB19, SSC^{+}19]. common-pool [MA19]. Commons [Kal18, SMLB19]. communal [NWD^{+}16a]. communicate [BFG^{+}16]. Communicating [Cor16]. Communication [HCP^{+}12, KTF^{+}14, MQ14, PMJ^{+}13, DP17]. Communities [AB13b, CRKM^{+}14, EDB12, GSW10, JVT^{+}13, KUT11, LDBS11, MCN10, PVP^{+}13, SK10, TBY^{+}10, CEdK18, GD15]. Community [APL^{+}13, BMGLW12, DC11, GL18, GZNC13, KRG13, LBLB14, LMC^{+}14, LTQ^{+}13, MJP^{+}11, NFF^{+}10, NGB^{+}16, RGRMPB^{+}13, SMGRJ10, TTSZH10, VKKA13, WIM^{+}10, WFM^{+}10, AG17, BCR^{+}17, BGD^{+}15, CQC^{+}19, DSG^{+}17, FWC^{+}15, HSU^{+}16, HOC^{+}16, LBD^{+}15, MBBB17, MSR19, PLH^{+}16, SL15, TQD^{+}18, WHS18, WZWF18, WMFB15, WFMW^{+}18]. Community-Based [KRG13, MJP^{+}11, NFF^{+}10, RGRMPB^{+}13, SMGRJ10, WFM^{+}10, AG17, BCR^{+}17, DSG^{+}17, MBBB17]. community-level [SL15]. Como [LAA^{+}10b, AOLK11, BMGLW12, BFHP10, Col10, CMNM10, DFW^{+}10, FJP^{+}11, GZA10, GZD11, GPK^{+}10, HF10, HRKMG12, MMPF10, MAGA10, NFF^{+}10, ODTS10, PCTW10, RRC^{+}10, RMGCR10, SA11, Van12, VLRB12, WHB^{+}11, WRG^{+}12, ZM12, BE10, BW10, JCA^{+}11]. compact [WÖ16]. Comparación [HSB12, BRCK11, MO11]. Comparative [Fli10, CCS^{+}16]. Comparativos [Fli10]. compare [HBTB15]. Comparing [BRCK11]. Comparison [CB13, GV16, HSB12, MO11, WLL^{+}16]. Compassion [Dor11, WBB^{+}18]. Compassionate [HWS19, DW19, HCA^{+}19, OCR^{+}19, RM19, WBNR15]. Compensaciones [HSBJ12]. Compensation [WS14, RN17]. compensatory [ACW18]. Competencia [WJ10]. Competing [Har11]. Competition [WJ10, WSEM13]. complacency [GWRW18]. Complejidad [HAB^{+}11]. complement [OJAV15]. Complementariedad [LGK^{+}11]. Complementarity [LGK^{+}11, TCP13, CNvL^{+}15]. Complementary [HMG^{+}13, HRL14, NWD^{+}16a, Plo17, ITG^{+}16, TMK^{+}16]. Complex [Ibi17, RSWM14, GLK^{+}17, LSV^{+}19, TAM^{+}19]. Complexity [HAB^{+}11, Wat10, PWP^{+}18]. Compliance [BEA^{+}13, CD13, HGM19, CCG18]. Component [JMW^{+}11]. Componentes [JMW^{+}11]. components [NBR^{+}15, RDS17]. Comportamiento [Bro10, FCGJH11]. Composición [HJSF^{+}12, SAC12, WIM^{+}10]. composite [SWSM19]. Composition [GLB^{+}13, HJSF^{+}12, SAC12, WIM^{+}10, FFS^{+}18, SSE^{+}17]. Comprehending [Kal18]. Comprehension [Mei12]. Comprehensive [HODB10]. Comprensión [WGB^{+}11b]. Comprobación [PAM^{+}10]. Compromises [PBH^{+}10b, KMR^{+}18]. Computer [SCSJ13]. Común [GPK^{+}10]. Comunes [HM12, HWW12]. Comunidad [DC11, TTSZH10, WIM^{+}10]. Comunidades [BMGLW12, EDB12, GSW10, KUT11, LDBS11, MCN10, MJP^{+}11, NFF^{+}10, TBY^{+}10, WFM^{+}10]. Comunitario [SMGRJ10]. con [AOLK11, BFHP10, BPT^{+}11, DDPC12, DWCW10, Duf11, EMD^{+}10, FCGJH11, GOR^{+}12, HRC10, HSAB10, HDH^{+}10, HSB12, JMP^{+}11, KKCS10, LGK^{+}11, LB11, LDBS11, MJA^{+}11, MAA^{+}10, MRPPM12, PA12, PHFPP10, RTL^{+}12, SD11, SPPN^{+}12, SS11, SDB10, UDG10, WJ10, WB10]. Concentración [LW11]. Concentrating [LW11]. Concept [BR11b, BHM^{+}14, MT10, PHFPP10, SBJG13, SBJG14, Agg15, FKRC17, LdBB^{+}17, WHR^{+}16]. Concepto [BR11b, PHFPP10]. Conceptos [LH10]. Concepts [LH10, Val18, WHB^{+}14]. Conceptual [WW10b]. Conceptualization [MAJH^{+}14]. concessions [VTT^{+}17]. Condiciones [HSBE12]. Condition [CGHM13, CMM17, SBG^{+}18]. Conditions [HSBE12]. Condors [KGG^{+}14]. Conductos [VLRB12]. Conductual [JGH^{+}11]. Conectada [SBN^{+}10]. Conectividad [ALL^{+}12, BSBM11, CMB12, JLDJ12, MLS12, ML10, RC12, SBT^{+}11]. Conexión [EMD^{+}10, LDBS11]. Conference [RSWM14, AGS19, BFG^{+}16]. conferences [RIBT18]. conferencing [FSJ^{+}17]. Confiabilidad [PPD^{+}12]. configuration [RSN19]. Conflation [WMS^{+}12]. Conflict [AOLK11, BMGLW12, BGAS10, BAGM13, BT14, GMNO15, KLN^{+}17, KLS^{+}19, LM18, LNNG16, RN17, vECD^{+}18]. conflict-mitigation [KLN^{+}17]. Conflicting [HMG^{+}13, Rob12]. Conflicto [BGAS10]. Conflictos [AOLK11, BMGLW12]. Conflicts [DBR13, DBB^{+}16, EAPB15, HNO^{+}18, NJK^{+}19, Poz18]. Confrontando [AG10]. Confronting [AG10]. confusion [CCS^{+}18]. Congeneric [AWU^{+}11]. Congenérica [AWU^{+}11]. Congo [KHGF^{+}17, MRRK11]. coniferous [SKSS^{+}17]. Conjuntos [UDG10]. connect [EGB18]. Connected [SBN^{+}10, WÖ16]. Connecting [BPN^{+}16, EMD^{+}10, LMM^{+}17]. Connection [LDBS11, BBM15, GLK^{+}17]. connections [SMARC^{+}16, SCWN18]. Connectivity [ALL^{+}12, BSBM11, Ben12, BHM^{+}14, CMB12, CFL14, Han11, JLDJ12, KTHH10, LW14, LGC^{+}13, MLS12, MRBL^{+}12, ML10, NLM^{+}13, OCP^{+}14, RBW14, RC12, SCK13, SBT^{+}11, SMF^{+}15, BGD^{+}15, BPN^{+}16, DAA^{+}19, GDM^{+}19, KBC^{+}18, KM18, LW18, MDD^{+}18, OGS^{+}19, SSWR17]. Conocimiento [GBMRG^{+}10, JGH^{+}11, LSDK11, MBF^{+}12, MAGA10]. Consecuencias [FMF10]. Consenso [ACL^{+}12]. Consensus [ACL^{+}12, DFR^{+}17, NMS^{+}18]. Consequences [BW13, FZ18, FMF10, ACW18, BMF16, KNB^{+}17, PRM^{+}18]. consequentialist [HWS19]. Conservación [AG10, AH12, AMO11, ACE^{+}10, ALL^{+}12, BMGLW12, BPW12, BRV10, BB10b, Ber12, BV10, BG10, BW10, BPT^{+}11, BEE^{+}11, BR11b, Bro10, BRCK11, BDH^{+}12, CND10, CB10, CJDZ10, CWH^{+}10, CLV^{+}10, CKKL10, CBS11, CGM10, DFW^{+}10, DWM^{+}10, EMS^{+}11, FKM^{+}11, GZA10, GZD11, GHKK10, HF10, Han10, HHDM10, HRC10, HSBJ12, HAB^{+}11, HSB12, KKCS10, KCDC10, LSDK11, LH11, LW10, LH10, MBF^{+}12, MC11a, MCL12, MMPF10, MVB^{+}12, MC11b, MBJWC10, MWC11, NFF^{+}10, OSC10, OBBP10, OKLP11, ODTS10, PRR^{+}11, PTN10, PCTW10, PBB^{+}11, RRC^{+}10, RRV10, RHM10, RBC^{+}11, Rud11a, SSVA11, SSO11, SD11, SFM^{+}11, SKC^{+}12, SA11, SHCD11, SSD10, SH11b, TM10, TGD10, TCDR12, UDG10, UCFV10, WFM^{+}10, WRAW10, WB10, WHB^{+}11, WW10b, WD10, GBG^{+}10, RBL10]. conservancies [NWD^{+}16a]. Conservando [BB10a]. Conservar [CMB12, DK10, LOME10, SHO^{+}10]. Conservation [AR10, AME^{+}13, AG10, AH12, AA12, AMO11, ACS14, AH11, ACE^{+}10, ALL^{+}12, AB13a, APR^{+}14, BMGLW12, BPW12, BRV10, BMM^{+}14, BRS11, BMG^{+}14, Ben12, BB10b, BB11a, Ber12, BCC^{+}14, BV10, BG10, BM14, BDG^{+}13, BW10, BPT^{+}11, BCT^{+}14, BME^{+}14, BEE^{+}11, BSF^{+}14, Bor19, BLG^{+}11, BR11b, Bro10, BT14, BRCK11, BGL^{+}14, CND10, CG13, CDF^{+}12, CJDZ10, Car14, CWH^{+}10, CRG12, Cha10, CLV^{+}10, CKKL10, ÇRG^{+}17, Cla13b, CS14, Cla14, Con11, CC14, CMS^{+}13, CWAW14, CS19, Cor19, CD13, CVA15, Cox10, CBS11, CGM10, DFW^{+}10, DBR13, DMR^{+}17, DP17, Der13, Die10a, Die10b, DSW16, DVW^{+}13, DHF16, Dre11, DWM^{+}10, EBL^{+}14, ER14, EMS^{+}11, EPB^{+}14, FKM^{+}11, Fer17, Fie15, Fis12, FBF^{+}14, For11, FBB^{+}14, FLW14a, GKP13, GBG^{+}10, GZA10, GFG^{+}12, GSDE14, GZD11, GRA14, Gor11, GHKK10, GB14]. Conservation [GDMF13, GMCW13, HSRC14, HF10, Hal18, Han10, Han11, HHDM10, Har11, HCLW13, HRC10, HMIE14, HDN^{+}14b, HH14, HSBJ12, HAB^{+}11, HBD^{+}19, HSB12, HMB^{+}15, IBT18, JMF12, Jen14, Jew13, JIAMS^{+}14, Joh10a, Joh11, Jør15, Kar14, KMTK13, KKCS10, Kir14, KCDC10, KH16, KNK13, KB11, LFMRD12, LJA^{+}13, LK14, LSDK11, LKB^{+}10, LMC^{+}14, LRA^{+}14, LKH^{+}13, LH11, Len17, LA13, LDB^{+}10, LWY^{+}14, Lic16, LW10, LH10, LGCF19, Lan11, ML14, Mar10, MBF^{+}12, Mar14, MG12, MA19, MCH^{+}16, MTN13, MT10, MF14, Mih19, MC11a, MCL12, MNW^{+}14, MGMA^{+}14, MMG^{+}13, MMPF10, MVB^{+}12, Mol19, MC11b, MAJH^{+}14, Moo17, Mor12, MBJWC10, MWC11, MGG^{+}13, NBM^{+}12, NFF^{+}10, NMSF11, NKSB^{+}13, NDB^{+}12a, Nov16, NS18, Nut15, OSC10, OPP14, OBBP10, OKLP11, OCP^{+}14, OSF10, Ore12, ODTS10, PDM11]. Conservation [PRR^{+}11, PTN10, PSD^{+}14, PFA^{+}14b, PCTW10, Pau16, Per11, PI11, PCW14, Phi10, PGBV^{+}13, PBB^{+}11, PMMG14, Poz18, Pri10, Pri11, PMP14, PS18, RB13, RSS^{+}13, RSS^{+}12, RGMR12, RPS13, RMRO14, RHWS18, RHT14, RCK13, RRC^{+}10, RGRMPB^{+}13, Ric11, RL13, Rob12, RSWM14, RCR^{+}12, RBL10, RRV10, RHM10, RBC^{+}11, Rud11a, SMK14, SSSD14, SSVA11, SABV13, SH11a, SMM16, SSO11, Sch11, SD11, SPL^{+}11, SBBP11, SFM^{+}11, SKC^{+}12, Sik12, SA11, SML^{+}13, SHCD11, Sou14a, SNH^{+}14, SSD10, SM14, Sul10, SH11b, TM10, TYA16, TW14, TGD10, Tru18a, TCDR12, TTEM14, UDG10, UCFV10, VM14, VPS^{+}14, Ver16, VMÁR^{+}13, VFV^{+}12, Wat10, WFM^{+}10, WRAW10, WD12, Whe11, WT13, WKI^{+}14, Whi10, WB10, WHB^{+}11, WMW^{+}14, WW10b, WSEM13, WWW^{+}12, WD10, Zha15, ZZW^{+}10, AdBR15, AP17, ABMG19, ABG^{+}19]. conservation [AGFR18, ABB^{+}18, AKJ^{+}18, ACB^{+}15, Arm19, AGS19, AAS^{+}17, AHT^{+}15, AWT^{+}15, BLR18, BW16, BDKS15, BH17, Bei19, BMB^{+}15, Ben16, BRK^{+}17, BMB^{+}19, BTBC^{+}16, BCOB15, Blu16, BTM^{+}15, BFG^{+}16, BPH^{+}19, Bra17, BCM16, BCLMM19, BGP16, BB15, BKR17, BBF19, Bus15, CCW^{+}16, CBF^{+}19, CHG^{+}17, CY17, CCG18, Cha15, CCN16, CLM^{+}15, CAB^{+}18, CEdK18, Cla15, CMG15, CC17, CC18, CM17, CLvK^{+}17, CPH^{+}15, CS17b, CNK^{+}17, CHWH18, CCS^{+}16, DAKN^{+}17, DHM^{+}15, DPJM18, DAA^{+}19, DTM16, DLMAG15, DRF^{+}18, DPS^{+}16, DCB^{+}16, DSSW16, DW19, DCP^{+}19, Dys15, EBP^{+}18, EWCG15, ERBM17, FMV^{+}19, FGG^{+}19, FSAM15, FWC^{+}15, FDBA16, FMB^{+}18, FFHA19, FSJ^{+}17, FKRC17, FTB^{+}16, FHK^{+}15, GCAF^{+}18, GLF^{+}18, GWRW18, GD15, GMV^{+}18, GDG19, GG19, GTVC16, GV16, GW17, GPB^{+}15, Had15, HM16, HWS19, HH19a, HFBG18]. conservation [HBTMG15, HCA^{+}19, HPD16, HCBB^{+}15, Hen16, HGRS16, HTE16, HHJ^{+}19, HHB15, HDH^{+}15, HSF17, HPBP16, HAF^{+}17, HICF16, Hul15, IBA16, ISM^{+}18, IO17, IF17, JMH15, JKJ^{+}19, JGS^{+}16, JWS^{+}18, JVI^{+}18, KCD^{+}17, KHB18, KTBH17, KBML15, KBCM15, KP16, KBM17, KF18a, KF18b, KBSX19, KB17, KM18, KT16, LSK^{+}15, LHKA16, LBI^{+}18, LBD^{+}19, LDC^{+}15, LH19, LFS^{+}17, LdBB^{+}17, LKYL15, LKW^{+}15, LLR^{+}15, LCC^{+}15, LRL^{+}15, LGA17, LSV^{+}19, LBD^{+}15, LBB^{+}16, MMGN18, MTD16, MBT^{+}17a, ME16, MBBH16, MCD^{+}19, MR16, MRFvW16, MCMP15, MRPT19, MDSF19, MSC^{+}18, MPP^{+}19, MAB^{+}15, MH16, MBFB18, MWL^{+}15, Mor15, NWD^{+}16b, NAB^{+}17, NMS^{+}18, NRD^{+}16, NYC17, NBR^{+}15, NJK^{+}19, NSBJ17a, NBH^{+}17, NSBJ17b, NWAB19, NGB^{+}16, OSM^{+}17, OHHE16, OCR^{+}19, OJAV15, PNC^{+}15, PMN^{+}16, PPM^{+}18, PFN19, PS15, PWKT18, PFN15a, PRNV18, PG18, PLH^{+}16]. conservation [PSA^{+}17, PWKB19, RSN19, RTB^{+}18, RWF^{+}17, RBF15, RFV^{+}17, RK17, RM19, RIBT18, Ros15, RBE^{+}15, RHCW16, RLB^{+}17, RKC^{+}19, RTGR17, RJA^{+}16, SLGB15, SMV19, SJH19, SFAA^{+}18, San19, SLA^{+}19, SGB^{+}19, SG19, SAB^{+}16, SGB^{+}18, SMAG17, SLWL15, SH16, SCWN18, SBS^{+}15, Sny15, SSC^{+}19, SSMC17, SPD18, SKSS^{+}17, TPL18, THD^{+}18, TXBW19, TAM^{+}19, TAP19, TMJ^{+}15, VTM19, VSA^{+}15, VB19, VJE^{+}15, VJ15, VBN15, WBNR15, WBB^{+}18, Wal17, WDS15, WWH^{+}18, WZC19, WGS^{+}19, WTT^{+}17, WML^{+}19, WA18, WSH^{+}15, WBPL15, WCCT18, WD17, WG15, WWWC15, WSLW16, WHEN^{+}16, YNR19, YJL^{+}18, ZSM18, ZAP^{+}18, Sou13, Bra15, Ano12e, Ano13a, Ano13b, Ano13c, Ano13d, Ano14b, Ano14c, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano17b, Ano17c, Ano17d, Ano17e]. Conservation [Ano17f, Ano17g, Ano17h, Ano18c, Ano18g, Ano18h, Ano18i, Ano19c, Ano19h, Ano19i, Ben15b, Bur13, BJM15, CB10, HRSK13, HDN^{+}14a, JMB16, LZD^{+}14, PFA^{+}14a, Red10a, SGH^{+}13, Sou14b, TAB^{+}18]. conservation-action [LBB^{+}16]. Conservation-Area [UCFV10]. conservation-education [MRPT19]. conservation-genetics [SLA^{+}19]. Conservation-Oriented [Car14]. conservation-policy [CHG^{+}17]. conservationist [HSF17]. Conservationists [Ano19w, HRL14]. Conserve [CMB12, Red12, SHO^{+}10, CLM^{+}15, NGB^{+}16, RDS17, SURW18]. Conserving [BGO^{+}19, BBC^{+}14, BB10a, BHM^{+}14, BDH^{+}12, CNBW16, CKR^{+}13, DK10, ERW18, HDHL10, Kal18, Lin11, LOME10, Moo19, SS10b, ZP18, GBB^{+}15, LAA^{+}15, PR17]. Consideración [MVB^{+}12]. Considerando [HHDM10]. Considerar [CPC^{+}10]. Consideration [AGLRV14]. Considerations [HMG^{+}13, FHK^{+}15, JMH15]. Considering [KB17, MVB^{+}12, SCWN18, BBF19]. Consistency [SSE^{+}17, DPJM18]. Consistent [Boy10]. Consistente [Boy10]. consolidating [GH18]. Conspecific [MAWF12]. Constituency [Pri11]. constrained [DHG^{+}17]. Constraint [Cat11]. Constraints [FMJP11a, LHKA16, MVB^{+}10, MVVS15]. Construcción [SBN^{+}10]. constructed [SKPP15]. constructs [YNR19]. Construir [BW10]. consulting [Fie15]. consumer [NBA^{+}18, VW19]. Contemporáneo [WGR^{+}10]. Contemporáneos [WW10a]. Contemporary [WW10a, WGR^{+}10, ZFW^{+}11, DWM^{+}19]. Contenido [JLCK12]. Content [BAGM13, JLCK12]. Conterminous [RPA^{+}12]. Context [ESCF12, ML14, SGS^{+}10, WFM^{+}10, DUS^{+}16, LKW^{+}15]. Context-Based [ESCF12]. Context-Dependence [SGS^{+}10]. Context-Dependent [ML14]. Contexto [ESCF12, SGS^{+}10, WFM^{+}10]. contexts [CS17a]. contextualize [CBB18b, GML^{+}19]. Continental [RBD^{+}13, KNB^{+}17, RCSK15]. Continental-Scale [RBD^{+}13, KNB^{+}17]. Continente [MMCUR^{+}12]. Continents [Aro11]. Continue [Ban12]. Continuum [CC14, FGG^{+}19]. contra [BAPP11]. contraception [CHWH18]. Contraction [Fat14]. Contracts [LA13]. Contrastando [LGB^{+}10]. Contraste [BJR^{+}11]. Contrasting [BJR^{+}11, LGB^{+}10, GRT^{+}18]. Contrasts [VCDF13]. Contratiempos [OSC10]. Contratos [FKM^{+}11]. Contribuciones [MTM12]. contribute [AKJ^{+}18, RBE^{+}15]. Contribution [CFO^{+}14, CPF13, MJG15, PLH^{+}16, SLWL15, WGLL17]. Contributions [MTM12, SAB^{+}16, BCLMM19, CCB^{+}19b, HLS^{+}15, LZR^{+}18]. Control [AGP^{+}14, DSV17, LMM^{+}10, MFN^{+}11, RSPD13, DFR^{+}17, FS15, NW15, RJA^{+}16]. Controlada [OKLP11]. Controlado [GSW10]. Controlar [LMM^{+}10]. Controlled [GSW10]. Controlling [HWB^{+}16, MDD^{+}18]. Convencionales [FCGJH11]. Convenience [HJSF^{+}12]. Convention [BCD^{+}16, LAA^{+}10a, SCA17, AKJ^{+}18]. conventions [VMO^{+}18, vBSC^{+}19]. Conversion [RWB11, RWB11]. Cool [GRA14]. Cooperación [ÖB12]. Cooperation [ÖB12]. Cooperative [RM14b]. Cooperatives [LA13]. coordination [MV18]. Coprolites [WWW^{+}12]. Copyright [Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano18n, Ano18o]. Coral [BWE^{+}13, BCGP13, CGHM13, DFW^{+}10, HHY13, LW11, MAA^{+}10, Mcc10, VPMS13, WGF^{+}12, Yap11, ANE^{+}16, BH18, CESS^{+}18, KT16, MM15, MGMC15, RBF15, WWdS^{+}16]. corals [AAvRS19]. Cordgrass [MGCL14]. corpora [RCBT18]. Corporal [Ber12, MJA^{+}11, WGR^{+}10]. corporate [ABMG19]. Corporations [Rob12]. Correct [KRS^{+}10]. Corrected [DSW16]. Correcting [SSC^{+}19]. Correction [Ano11c]. Corredores [EMD^{+}10, GNWSB10]. Corregir [KRS^{+}10]. Correlaciones [SPPN^{+}12, FWM^{+}12]. Correlates [FOF^{+}15, FWM^{+}12, SMS14, CBWFL15, PS15, VJE^{+}15]. Correlation [BGZ^{+}15, vWWCF15]. Correlations [RLGG^{+}12, SPPN^{+}12]. Corridor [BC14a, BP13, GNWSB10, MRT^{+}12, DHG^{+}17]. Corridors [EMD^{+}10, GB14, HBD^{+}14, JCCC14, PMS13, Bei19, VNAT19]. Corriente [HSBJ12]. corrupt [Ben15a]. corruption [SBS^{+}15]. Corte [CMRH10]. Cosecha [CBM^{+}11, CT11, GT10, RMGCR10, WW10a]. Cost [BMA^{+}14, BDH^{+}12, CLV^{+}10, CBS11, DSSW16, GSDE14, MARV^{+}11, MNW^{+}14, NM18, PMS13, PMM^{+}14, RHCW16, SA14, AHP^{+}15, BLM^{+}16, EEB15, PG16, RK17, SG18, VNAT19]. Cost-Effective [CLV^{+}10, RHCW16, AHP^{+}15]. Cost-Effectiveness [GSDE14, MARV^{+}11, PMM^{+}14, DSSW16, BLM^{+}16]. cost-efficient [RK17]. Costa [PB11]. Costas [GPK^{+}10]. Costo [BDH^{+}12, CBS11, SKC^{+}12]. Costos [BEE^{+}11, CWH^{+}10, ACE^{+}10, SS12, WRG^{+}12]. Costs [ACE^{+}10, BEE^{+}11, CWH^{+}10, JVT^{+}13, LA13, LCL13, SMTW13, SKC^{+}12, SS12, WRG^{+}12, DKG^{+}15, ISM^{+}18, KMR^{+}18, MGA16, MDD^{+}18]. count [RNE^{+}19]. Counter [LDB^{+}10]. counterfactual [BRA^{+}19]. Countering [LTN18]. Countries [LKH^{+}13, KRF^{+}19]. country [TUHK^{+}17]. Countryside [PZM14]. counts [DMB^{+}17, RNE^{+}19]. Course [NMMP14]. Cover [Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano12e, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano13d, Ano14b, Ano14c, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano18g, Ano18h, Ano18i, Ano18d, Ano18e, Ano18f, Ano19h, Ano19i, Ano19d, Ano19e, Ano19f, Ano19g, Ben15b, FG12, HRSK13, HDN^{+}14a, LZD^{+}14, MRBL^{+}12, NDB^{+}12b, PFA^{+}14a, PRE^{+}13, RWB12, RPA^{+}12, SGH^{+}13, Sou14b, WGB^{+}11a, Ano11c, BVAC17, SBR^{+}19, SJZ19]. Coverage [JLCK12, SCA17, ZAB16]. covered [MAM^{+}15]. Coviability [MDJ14]. Cowbird [WLCS14]. Cowling [Fie15]. Crafting [Fis12]. Crane [BKKM14, HBTB15]. Crayfish [TTSZH10, BSL^{+}19, DLMAG15]. Crayfishes [CLT^{+}13, LO10]. Creación [HSAB10, SHC^{+}11]. Creating [NBW^{+}12, HNO^{+}18]. Creation [SHC^{+}11, WFMW^{+}18]. Creativity [APR^{+}14]. Creatures [Møl17]. Crecimiento [ANTBG12, JMW^{+}11]. Credibility [HPBP16, Mel10, MRFvW16]. Crime [Gor11, Öst14, FSR^{+}18, TAM^{+}19]. Crípticas [WRTB10]. Crisis [Ald12, Car11, Hic10, PS13, DPGY16, DRA^{+}15, MM15, SPD18]. Criteria [ACL^{+}12, Boo14, CFL14, DBV13, ETB10, KMH^{+}14, NCC13, RTT^{+}13, RRCB^{+}11]. Criterios [ACL^{+}12, ETB10, RRCB^{+}11]. Crítica [MWC11]. Critical [HP11, Hur13, LFM17, MWC11, PCRD14, CMMP15, LZZ^{+}17, TGBB^{+}15]. Critically [CWS^{+}14, RLGG^{+}12, SDT17, JLCHNG^{+}17, WD17]. critique [FKRC17]. Crónico [BBP12]. crop [MSMCB18]. Cropland [RWB11]. Crops [BHK10]. cross [WZWF18, FFM^{+}14]. cross-sectoral [WZWF18]. Crossing [FC10, SCK13]. Crowdfunding [GCAF^{+}18]. Cruce [FC10]. crucial [SJH19]. Cryptic [WRTB10, BHAR^{+}18, BFB^{+}19]. Cuando [CWH^{+}10]. Cuantificación [WJ10]. Cuantificar [BAG^{+}11]. Cuantitativa [MBJWC10, PF10]. Cuantitativos [PTN10]. Cuatro [Jam11]. Cuba [SMGR17]. Cucumbers [LAA^{+}10b]. Cuellos [BAJ^{+}10, HB10]. Cuenca [EDB12, AWU^{+}11, MPH^{+}11]. Cuencas [IUM10]. Cuevas [ATW^{+}11]. Cuidado [SRS11]. Culling [HM11, LMM^{+}10]. Cultivating [APR^{+}14, Pri11]. Cultivation [BT14, MRMV14, LGCF19]. Cultivo [HRKMG12, MBJWC10]. Cultivos [BHK10]. Cultura [CND10]. Cultural [BM14, Qui14, WFM^{+}10, GKA^{+}15, NMS^{+}18, WL17a]. Culture [SLSH14, Bus15, MR16]. Cultured [BW13]. Cultures [Har11, VLSOS16]. Culver [Bra15]. cumulative [GHM^{+}15, HH19b, SSRS18, WK15, ZV19]. Cuotas [LGK^{+}11]. Curio [MF16]. Curio-City [MF16]. Current [CMMP15, Col10, CKK^{+}14, GLB^{+}13, MVB^{+}10, SKC^{+}12, LZR^{+}18, LLR^{+}15, PBB^{+}17, VNAT19, HSBJ12]. current-flow [VNAT19]. currently [PA18]. curricula [SJH19]. Customary [FCGJH11]. Cutting [CMRH10].

D [Bra15]. D.C [Ano19w]. Dam [EVF13, IRK^{+}18]. Dam-Induced [EVF13]. Damage [BHK10, KKK14, KTN12]. Damage-Causing [KKK14]. Dams [KPL^{+}12]. Damselfly [SHLK^{+}10]. Daño [BHK10, KTN12]. dark [HRHC16, LdBB^{+}17]. Darkness [Møl17, GDB15]. Darling [AWU^{+}11, AWU^{+}11]. Data [AWKR^{+}11, CFL14, CKK^{+}14, JPT13, KGS^{+}11, KUT11, LHDC12, LRA^{+}14, PBH^{+}10b, PA12, PS10b, SSK^{+}11, SS11, SM14, SA14, UDG10, VMÁR^{+}13, BBM18, BCOB15, BBK^{+}17, BBB^{+}16, BSCB18, CBB18a, CVA15, CPD^{+}15, DBV^{+}18, DMGB^{+}19, Had15, HOC^{+}16, HPD16, KTBH17, LS15, LPW16, MCH^{+}16, MWL^{+}15, PNSG18, PHJ^{+}17, SBU^{+}17, SKLP16, TRPD16, TLT16, VSA^{+}15]. data-based [PNSG18]. data-deficient [BCOB15, BBK^{+}17]. data-driven [PHJ^{+}17]. data-rich [HPD16]. date [BSCB18, CBB18a]. Datos [AWKR^{+}11, KUT11, LHDC12, PA12, SSK^{+}11, SS11, UDG10]. Day [RTL^{+}12]. Dead [OK10]. Deadwood [CMJ^{+}11, KM18, RSN19]. Deadwood-Associated [CMJ^{+}11]. Dealing [BM10]. Debate [SG10, SR11, Die16, HSF17]. debates [Blu16]. Debating [FDBA17]. Debe [GHKK10]. Debido [KG10, SHC^{+}11]. Debris [SHWT14]. Decade [CSMG^{+}13, VLGRS14]. Decades [RPS13]. December [Ano11b, Ano14a, Ano18b, Ano18a, Ano19a]. Deciding [Red12]. Deciduous [Mat13, Mat15, SVN^{+}15]. Decipher [SJP^{+}13]. Decision [CC14, DBR13, GWRW18, HBM^{+}12, LWLC12, MR12, NS12, PBB^{+}11, RTT^{+}13, SSD10, VPMS13, GCC^{+}12, Arm18, CCS^{+}18, DBB^{+}16, EEB15, KMR^{+}18, LBI^{+}18, MBBH16, PG16, TMG^{+}18, VTM19, VB19, YNR19]. Decision-Making [NS12]. Decision-Network [VPMS13]. Decision-Support [SSD10, VTM19]. decision-tree [EEB15]. Decisiones [HBM^{+}12, MR12, PBB^{+}11, SSD10]. Decisions [BRM^{+}14, CPF13, GAG13, SML^{+}13, CPH^{+}15, FRY^{+}17, MWL^{+}15, WDS15]. Declinación [KPL^{+}12, LN11, OTKE12, OB11b, PPD^{+}12, WKF^{+}11, CGM10, JMP^{+}11]. Declinaciones [BFHP10, MBGS11]. Decline [BBM13, CFO^{+}14, CGM10, HJMM14, JDA^{+}10, JMP^{+}11, LN11, MBM^{+}14, OB11b, PPD^{+}12, RP13, SGSR^{+}12, SJP^{+}13, WKF^{+}11, CWH^{+}17, JLCHNG^{+}17, PS15, TDCG15]. Declines [BFHP10, CLV14, GZNC13, HLC13, KPL^{+}12, LDC^{+}13, MBGS11, OTKE12, SKBE15, BHF^{+}19, BIP^{+}15, COZS16, FWH19, PRM^{+}18, RCARR18, SR17, VJE^{+}15, vWWCF15]. Declining [HD14, HLS^{+}18]. Declive [JDA^{+}10]. Deconstructing [HCA^{+}19]. Decoupling [GZR12]. Decreasing [BJM15]. Decreto [BMGLW12]. Deep [Vit17, VLRB12, BCK^{+}19]. Deep-Sea [VLRB12]. Deer [BTBDS12, WNY13]. Defending [CHG^{+}17]. Defensa [Wil12a]. Deficiencies [BSCB18]. deficient [BCOB15, BBK^{+}17]. Definición [BHK10, PTN10, BV10, HSBJ12]. Defining [BV10, BHM^{+}14, HSBJ12, JBB^{+}14, PWRM10, PCRD14, Sel16, SBU^{+}17, SGA^{+}19, TXBW19, TND16]. Definition [BHK10, HAV11, ENW^{+}18]. Deforestación [SWL^{+}10]. Deforestados [RSE12]. Deforestation [HNDE13, SWL^{+}10, VLGRS14, BHPM^{+}18, DMF^{+}18, FSSW19, LPA^{+}17, MGW^{+}15, OQGR^{+}15, TBM^{+}16]. Deforested [RSE12]. Degazettement [BPA14, CVMM17, QKC^{+}19]. Degradation [CRKM^{+}14, VLGRS14, KT16, LPA^{+}17, SBB^{+}15]. degraded [ANE^{+}16]. degrees [BTB^{+}18]. Degu [PMK^{+}10]. del [AT11, AH12, BPW12, BM10, BFHP10, BTBDS12, BEE^{+}11, Boy10, BR11b, Bro10, BRKP12, CMB12, CMRH10, CLRV11, CAB^{+}10, CMJ^{+}11, DWH^{+}11, DBP^{+}10, DWM^{+}10, DEJM10, EMD^{+}10, ERC^{+}10, EMS^{+}11, EPD12, FCC^{+}11, FC10, FMF10, FLP^{+}11, GLH12, GZA10, GSB10, GBMRG^{+}10, GS11, GPK^{+}10, GSW10, HM11, HLB^{+}12, HRSK12, HG10, HWW12, HSBE12, HMM10, JLCK12, JDA^{+}10, JLDJ12, JMW^{+}11, KE10, KPL^{+}12, LW11, LSDK11, LSBB10, LW10, LB11, LPT12, MPH^{+}11, MFN^{+}11, Mcc10, MMM^{+}12, MJP^{+}11, MMPF10, MRRK11, MVB^{+}11, MR12, MPB10, MAGA10, MWC11, ÖB12, PRR^{+}11, PSAZ11, PA12, PCTW10, PHFPP10, PMS^{+}10, PMK^{+}10, RM11, RC12, SVH12, SGS^{+}10, SSK^{+}11, SD11, SPPN^{+}12, SKC^{+}12, SRS11, SSB^{+}12, SMGRJ10, SM11, SAGSHL10, SHO^{+}10, UCFV10, VJMH12, WL12, WIM^{+}10, WEC^{+}11, WFM^{+}10, WK10, WGB^{+}11b]. del [WR12, WKP11, WRG^{+}12, MB10]. Delays [OSC10]. Delfines [DDPC12, KKCS10]. Delon [NWD^{+}16b]. Delta [MWC11, MWC11]. demand [VKM^{+}17, VW19, WGS^{+}19]. Demise [Rob13]. Demografía [GT10]. Demográfica [SM11]. Demográfico [MMM^{+}12, Rog11]. Demográficos [CT11, LHDC12, PA12]. Demographic [CT11, CB13, ER14, JPT13, MMM^{+}12, PA12, RBW14, Rog11, SCK13, SM11, SKH13, ALA17, CCCCN18, WC19]. Demographics [LHDC12, DWCC17]. Demography [GT10, BC16]. Demonio [HLB^{+}12, LMM^{+}10]. demonstrating [MF15]. Demonstration [HF10]. Demoras [OSC10]. Demostración [HF10]. Denali [MFN^{+}11]. dendrobatidis [BVC^{+}14, BGH^{+}15, CLRV11, GLH12, MSSM09, OSFFB^{+}11, PGC^{+}13b, SGH^{+}11, SMH^{+}14]. Dendrocopos [RC12]. denialism [DW19, MSL19]. Densely [Fat14]. Densidad [JDA^{+}10, LB11, SA11]. Densidades [KGS^{+}11]. Densities [KGS^{+}11, FBWIO19, MRMH17]. Density [BCGP13, CGHM13, JDA^{+}10, LB11, MAWF12, SA11, CTBR19, DUS^{+}16, EG17, GDM^{+}19, SBR^{+}19, VSA^{+}15]. Dependence [SGS^{+}10, WLCS14]. Dependencia [SGS^{+}10]. dependency [AC16]. Dependent [ML14, MVB^{+}12, AWHW17]. Dependientes [MVB^{+}12]. Depensation [BBBG10]. Depleted [ETB10]. depletion [FOF^{+}15]. Deportiva [PBK^{+}11]. Depredación [LPT12, SRS11]. Depredador [FC10, Van12]. Depredadores [HSAB10, LPT12, SPSS10, TTSZH10, BB10b]. Depresión [FBE^{+}11, GLNJ10, JMW^{+}11]. Depressed [SRG^{+}15]. Depression [FBE^{+}11, GLNJ10, JMW^{+}11, RKH^{+}14, RLGG^{+}12]. depth [BBH17]. Derelict [BHSA12]. Derivadas [KRS^{+}10]. Derivation [Whi13]. Derived [FWM14, KRS^{+}10, SBG^{+}18]. Derogation [SJH13]. Desarrollar [BSBM11, ETB10]. Desarrollo [BMGLW12, BB10b, Bro10, FMF10, LPT12, MAA^{+}10, MC11a, MBJWC10, OBBP10, SAGSHL10, WT10, WK10]. Desconexión [BAJ^{+}10]. describe [BCD^{+}16]. desde [BV10, GZA10]. Desert [ESD^{+}17, HSAB10, LTQ^{+}13, Lub16, WON^{+}13, BCSM15, LB15, PBR^{+}16, PRB18]. Deserts [Lub17]. Desierto [HSAB10]. Design [Cat18, MC11b, PMJ^{+}13, RBL10, SSK^{+}11, VGAB^{+}19, WLD16, BP16, BLM^{+}16, GDM^{+}19, GPB^{+}15, MVVS15, RB17, TAM^{+}19, WÖ16]. Designación [LCG10, PF10]. Designation [LCG10]. Designations [PF10]. Designing [HHDM10, CL16]. designs [VNAT19]. Desires [LL14]. Desove [WK10]. Después [CMJ^{+}11, EDB12, SS10a]. Destino [PA12]. Desventajas [BRKP12]. Detección [CPC^{+}10]. detect [BTB^{+}18]. detectability [GBMW15]. Detected [MRT^{+}12, LDD19]. Detecting [DMGB^{+}19, KMH^{+}14, LDC^{+}13]. Detection [AL14, CPC^{+}10, VSM^{+}11, FMB^{+}18, MK18, RBHW18, RNE^{+}19]. detections [LMJK^{+}19]. Detectores [VSM^{+}11]. detectors [WFC^{+}18]. Determinación [VSM^{+}11, PF10]. determinant [DSG^{+}17]. Determinants [LBB^{+}16, SMH^{+}14]. Determinar [MPH^{+}11]. Determination [PF10]. Determine [MPH^{+}11, GS16]. determined [GGL^{+}15]. Determining [CC14, NMMP14, NJT14, THD^{+}18, VSM^{+}11, LHKA16]. deterrence [DMGB^{+}19]. Develop [ETB10, RTT^{+}13, ABB^{+}18]. Developed [ER14]. Developing [BSBM11, BCR^{+}17, Boo14, CFL14, MAA^{+}10, WZWF18, WA18]. Development [BMGLW12, BBC^{+}14, BB10b, Bro10, EBL^{+}14, FMF10, HH14, LK13, LPT12, LP13, MHG^{+}14, MC11a, MBL^{+}14, MBJWC10, OBBP10, PG12, PFR^{+}14, Pri19, RHT14, SJH13, WT10, WT13, Zha15, AJ18, FBWIO19, SAB^{+}16, TXBW19, TT16]. Developmental [SAGSHL10, WK10]. Developments [HGP16, Agg15]. devil [RBHW18, HLB^{+}12, HJMM14, LMM^{+}10]. Devising [PCTW10]. Devra [Die10b]. diagnose [WHEN^{+}16]. Diagnóstico [MAGA10]. Dialogue [RSWM14]. dichotomy [SL16]. Diclofenac [MSZB^{+}14, GTC^{+}16]. Die-Off [CBA13]. Dieback [ECES13]. diets [GSB^{+}16]. Díez [Cla15]. Diferencial [SGH^{+}11]. Diferencias [GT10, PBH^{+}10a]. Diferentes [CGCL11, SDB10]. Difference [CZS19, HDH^{+}15]. Differences [BGH^{+}15, GT10, PBH^{+}10a, GRL15]. Different [CGCL11, SDB10, WGJ13, DSSW16, KTBH17, RBK^{+}17, SKPP15]. Differential [SGH^{+}11]. Differentiating [TAP19]. Difficulties [Pel15, WL17b]. digital [BBB^{+}16]. Dignity [BC14b]. dilemma [ABG^{+}19]. Dimensiones [KCDC10, Bro10]. Dimensions [Bro10, KCDC10, LNNG16]. Dinámica [GGG10, MSSM09, RC12, HRSK12, MQAMWM12]. Direct [BKM14]. Directamente [Rog11]. Directions [MVB^{+}10, Red10a]. Directives [RCK13]. Directly [Rog11]. Directrices [GZA10, HHDM10, MVB^{+}10, TBY^{+}10]. Disambiguating [SBJG14]. Disappearing [HHY13]. Disarray [Joh11]. discarded [WHG^{+}15]. Disciplinarity [PMMG14]. discipline [MCD^{+}19]. Disconnection [BAJ^{+}10]. discounting [Arm18]. discovered [SC19]. discretion [LRMWG16]. discussions [BKR17]. Disease [CLT^{+}13, CLV14, DBV13, FCC^{+}11, GAFG14, HLB^{+}12, HSR^{+}13, HJMM14, LMM^{+}10, LW11, MSSM09, OTKE12, PHC^{+}11, SVH12, Sch10, SMH^{+}14, CBM^{+}18, DDv18, FMB^{+}18, SHB^{+}17]. Disease-Driven [PHC^{+}11]. Disease-Induced [HJMM14]. diseases [CBPM19]. Diseño [MC11b, RBL10, SSK^{+}11, HHDM10, MAA^{+}10, PCTW10]. Disentangle [WHB^{+}14, RCBT18]. Disentangling [NCKC17]. Disminuidas [ETB10]. Dispersal [CBSZ10, CFBL10, GZA10, HG10, KM18, LH11, MPH^{+}11, MSF^{+}12, ZYYH10, CTRE^{+}17, HAT^{+}15]. Dispersing [MC14]. Dispersión [CBSZ10, CFBL10, GZA10, HG10, LH11, MPH^{+}11, ZYYH10]. displacement [CBM16, LCYJ19]. Disponibilidad [LCG10, WL12]. Disproportionately [AS13]. Disrupción [SS12]. Disruption [SS12]. disservices [CGK^{+}19, SL16]. Distance [CGHM13, DBK^{+}14, HMIE14, MLS12, MPB10, PRB18, SLB^{+}15]. Distancia [MLS12, MPB10]. Distinct [FB11, Ore12]. Distinctiveness [RDM10]. Distintas [FB11]. Distribución [KRS^{+}10, KPL^{+}12, MFRH12, MRRK11, SBT^{+}11, UDG10, VMGB^{+}10, WGB^{+}11b, CAB^{+}10]. Distribuciones [BRV10, MRPPM12]. Distribution [CLT^{+}13, CAB^{+}10, GRA14, KRS^{+}10, KO12, LMC^{+}14, LKB^{+}14, MLRC18, MRRK11, SMS14, UDG10, VKKA13, GML^{+}19, LHKA16, LMJK^{+}19, TPL18, TLPP17]. Distributions [BRV10, MFRH12, MRPPM12, VMGB^{+}10, WGB^{+}11b, SB16]. Disturbance [FG12, MSF^{+}12, PMM^{+}14, SKCS11, WGF^{+}12, BMG^{+}17, CTBR19, CGE19, SSE^{+}17]. Disturbances [VKKA13]. disturbed [SLS17]. Dive [Vit17]. Divergence [HLS^{+}15]. Diverging [PTJ^{+}19]. Diversas [GLH12, ÖB12]. Diverse [ÖB12, KHK^{+}19]. Diversidad [AMO11, CBSZ10, Fli10, GKK^{+}11, Jam11, PBB^{+}11, RPA^{+}12, Rud11b, SAC12, TM10, VJMH12]. Diversion [WBM13]. Diversity [AMO11, BRS11, BCD^{+}16, CKR^{+}13, CBSZ10, FBA^{+}13, Fli10, FBB^{+}14, GKK^{+}11, HDHL10, Jam11, KMTK13, KO12, LFMRD12, LAA^{+}10a, LKB^{+}14, MMG^{+}13, PG12, Pek13, PBB^{+}11, RM14b, RPA^{+}12, Rud11b, SPN^{+}14, SCA17, SAC12, TM10, VJMH12, Whi13, AHT^{+}15, BdA15, BP16, CQC^{+}19, Dar15, GBD^{+}15, HICF16, LdBB^{+}17, LKYL15, LW18, QCW^{+}18, RDS17, SSW15, WWH^{+}18]. DNA [BME^{+}14, DHH^{+}12, GSB^{+}16, HJSF^{+}12, KGS^{+}11, MSY^{+}15, NJBM15, SMZZ17, WHE^{+}17, WSC^{+}17, WMW^{+}14, WSLW16, WLCW16]. DNA-based [MSY^{+}15]. Do [HMD^{+}10, LDB^{+}10, PA18]. Documenting [TDCG15]. Dodo [Rob13]. Dogs [AB13b, RM11, VSM^{+}11, EB15]. Dolphin [DDPC12, KKCS10]. Domain [Bra15, WZWF18]. Domestic [OBB^{+}11, RM11, SRS11]. Domésticos [OBB^{+}11, RM11, SRS11]. Dominados [AOLK11]. Dominant [Mar10]. Dominated [AME^{+}13, AOLK11, FGG^{+}19, Mor15]. Doñana [GBMRG^{+}10]. Dorada [SHC^{+}11]. Dosel [MPB10]. double [Dar15]. double-blind [Dar15]. doubt [LM18]. Doubting [Hal19]. Down [PMMG14]. Downgrading [BPA14, CVMM17, QKC^{+}19]. Downlisting [WL12]. Downsizing [BPA14, CVMM17, QKC^{+}19]. Downturns [Wit11]. Dr. [MMGN18]. Dragonflies [SS10a]. Dragonfly [SHLK^{+}10]. Drainage [IUM10]. Drainage-Basin [IUM10]. Draining [UIN^{+}13]. Drastic [ZZW^{+}10]. drawn [VTT^{+}17]. Driven [BAJ^{+}10, MT10, PHC^{+}11, DSK^{+}19, JBB^{+}15, MNW^{+}17, PHJ^{+}17]. Drivers [BJMT14, GFRK12, NBNMG13, NGL^{+}14, CCS^{+}16, LPA^{+}17, THD^{+}18, TAM^{+}19, WHS18]. Driving [Mol19]. Drought [AASK12, HAP^{+}13, BMA^{+}18]. Drought-Induced [AASK12]. Dry [BEE^{+}11, HODB10, Pra10]. dryland [AP17]. Drylands [FMF10]. Due [KG10, AGLRV14]. Duetting [VMW13]. Dulceacuícolas [OKLP11, PF10]. Durante [LHDC12, MQAMWM12, MVB^{+}12]. During [CPB14, KGG^{+}14, LHDC12, MVB^{+}12]. Dwelling [AGLRV14]. Dying [Pri19]. Dynamic [LKH^{+}13, YFB15, DTM16, GMNO15, ITG^{+}16, RDS17, WTT^{+}17]. Dynamics [CM13, DUS^{+}16, GGG10, GG13, HRSK12, LMJB^{+}19, MQAMWM12, MSSM09, NGL^{+}14, RC12, WH13, BBH17, DWM^{+}19, GGB^{+}15].

E-commerce [HHKK15]. E.U.A [LPT12, MWC11]. E.U.A. [CVN^{+}10, HSBE12, MPH^{+}11, WT10]. Eagle [MFN^{+}11, KNB^{+}17]. Early [Col10, KS17]. early-warning [KS17]. Earth [KHK^{+}19]. Earthworm [LB11]. Easements [FKM^{+}11, Bra17]. East [WSLW16, GSW10, PB11, YJL^{+}18]. eastern [LKYL15, SMGR17, SVN^{+}15, CLV14, DRF^{+}18, GOR^{+}12, Mat13, MLA^{+}14]. Ebola [PFN15a]. Eclosión [HB10]. ecoacoustics [GS16]. ecocentric [PWKT18]. Ecología [CND10, SH11b]. Ecológica [Boy10, CGM10, ETB10, HFL^{+}11, HSAB10, RDM10, SAB12]. Ecological [AT11, Aro11, AL14, BV10, BCT^{+}14, Boy10, Bra15, BRCK11, Cat18, CSMG^{+}13, CGM10, ETB10, FB11, GT10, GOR^{+}12, GBMRG^{+}10, GZJ^{+}13, HFSB13, HBD^{+}14, HBR^{+}14, HFL^{+}11, HSAB10, HGCT13, KTHH10, MF16, MCL12, PRM^{+}18, RDM10, SK10, SML^{+}13, SAB12, TW11, UCFV10, WT13, ALA17, AC15, AC16, CZS19, CLM^{+}15, FRY^{+}17, GFdM^{+}19, GS16, GW17, HH19b, IPA^{+}17, IRK^{+}18, JMH15, KMR^{+}18, MBBH16, MVVS15, PVO^{+}16, RHPF15, SBB^{+}15, TMG^{+}18, VJE^{+}15, WL17a, WBF^{+}19, Ano19w]. Ecological-Impact [SML^{+}13]. Ecological-Niche [UCFV10]. ecologically [BCD^{+}16, SBU^{+}17]. Ecológicas [FB11, GT10, GOR^{+}12]. Ecológico [GBMRG^{+}10, UCFV10]. Ecológicos [AT11, BV10, BRCK11, MCL12]. Ecologists [Poo13]. Ecology [Bra15, CND10, Ehr14, Fie15, Gon13, GL18, MHG^{+}14, Nov16, Par10, SH11b, Val18, Ven18, Whe11, Bai16, BW19, FSJ^{+}17, KM18, MBFB18, PLH^{+}16, SKSS^{+}17, WSH^{+}15]. Economía [ACE^{+}10, GBMRG^{+}10, RWB11]. Economic [Bro10, GBB^{+}14, GHKK10, HFSB13, SSK^{+}11, SKC^{+}12, SHC^{+}11, SA14, Wit11, CS17a, HNO^{+}18, HKFC15, HMB^{+}15, KMR^{+}18, MVVS15, STM^{+}19]. Económica [GHKK10]. Economically [HPL14]. Económicas [Bro10, Wit11]. Económico [SSK^{+}11, SKC^{+}12, SHC^{+}11]. Economics [ACE^{+}10, Par10, RWB11, DWCC17, IPA^{+}17]. Economists [PI11]. Economy [BL17, GBMRG^{+}10, HNO^{+}18]. Ecoregional [BMP^{+}12]. Ecorregionales [BMP^{+}12]. Ecosistema [AT11, DEJM10, ERC^{+}10, GPK^{+}10, PHFPP10, SKCS11]. Ecosistemas [GZA10, GPK^{+}10, HMD^{+}10, HWW12, RRCB^{+}11, SCL12]. Ecosistémicos [CLV^{+}10, NFF^{+}10]. Ecosystem [AT11, AR10, CLV^{+}10, DVW^{+}13, DEJM10, ERC^{+}10, FGH^{+}13, GPK^{+}10, GFGR12, GBP^{+}13, HGCT13, KAV18, MT10, NFF^{+}10, PRE^{+}13, PHFPP10, PS11, RGMR12, Rey11, SCL12, SLH10, SKCS11, VPMS13, WHB^{+}14, AHP^{+}15, Agg15, AWHW17, BGB17, CPR^{+}16, CGK^{+}19, CCB^{+}19a, CFB^{+}17, CCB^{+}19b, CE15, GKA^{+}15, HCSM15, LMCS17, PVO^{+}16, RNM^{+}18, SL16, TJJB18, TVY^{+}16, VKM^{+}17, WGS^{+}19]. Ecosystem-Approach [WHB^{+}14]. Ecosystem-Based [GPK^{+}10, SCL12]. Ecosystem-Risk [PRE^{+}13]. Ecosystem-Service [Rey11, LMCS17]. Ecosystems [AS13, Fie15, GZA10, GZJ^{+}13, HP11, HMD^{+}10, HWW12, Lin11, MYG^{+}14, MJ11, OSF10, Ric11, RRCB^{+}11, SPL^{+}11, AC15, CBCW18, LWZL16, RDS17]. ecotourism [BRA^{+}19]. Ectotherms [RP13]. Ecuador [PAM^{+}10].

Ecuatorial [AWKR^{+}11]. Edge [GGG10, Ord11, PY13]. Edge-Effect [PY13]. editing [KBSX19]. edition [Ano19w]. editor [Fie15, Ano10g, Ano11i, Ano13e]. editors [Ano16x, Fie15, Ano12a]. eDNA [SJD^{+}19]. education [MRPT19, MCD^{+}16, NMS^{+}18, SPD18]. Efectiva [CFBL10, SFM^{+}11]. Efectivas [BRCK11]. Efectividad [GNWSB10, LNP10, LGH^{+}12, MJP^{+}11, MMPF10, BTBDS12, SPSS10, VSM^{+}11, WRAW10]. Efectivo [HNS^{+}11]. Efectivos [ANTBG12]. Efecto [HLB^{+}12, ANTBG12, FCGJH11, Jam11, KPL^{+}12, LCG10, VMGB^{+}10, WFM^{+}10]. Efectos [AZCT12, BBP12, CW10, EG12, HM12, JMW^{+}11, KCSP11, LW11, MC11a, MMCUR^{+}12, PMS^{+}10, TTSZH10, WK10, WGB^{+}11b, AGAB11, ALL^{+}12, CBM^{+}11, CT11, CGCL11, CMJ^{+}11, CFBL10, DDPC12, DMZ10, EPD12, FG12, FEC^{+}10, GT10, GS11, GSW10, HG10, HSBE12, IUM10, KTN12, LPT12, MLS12, Mcc10, MBGS11, MRRK11, OTKE12, PBK^{+}11, Pra10, PMK^{+}10, PCBD11, RM11, RSW11, Rud11a, SAGSHL10, SKCS11, TAG^{+}10, WIM^{+}10, WSHB11, Wit11, ZBD11]. Effect [AME^{+}13, ANTBG12, DP15, GUHE15, HLB^{+}12, HCLW13, Jac13, Jew13, KKH14, LGC^{+}13, LCG10, PRE^{+}13, Pek13, PY13, QCMB18, RBW14, Rob13, SJH13, TMJ^{+}15, VMGB^{+}10, WFM^{+}10, WGF^{+}12, BMG^{+}17, WDS15]. Effective [ANTBG12, BRCK11, CLV^{+}10, CFBL10, FBF^{+}14, HNS^{+}11, Jos12, NJK^{+}19, SFM^{+}11, AHP^{+}15, GRT^{+}18, NW15, RHCW16, TAM^{+}19]. Effectiveness [BAPP11, BBWK^{+}14, BTBDS12, BVAC17, Bra17, CEdK18, GNWSB10, GSDE14, GB14, HCH17, IO17, LBLB14, LNP10, LGH^{+}12, MARV^{+}11, MJP^{+}11, MMPF10, NA13, PMM^{+}14, QCW^{+}18, RYW^{+}15, SPSS10, SA14, VSM^{+}11, WRAW10, AC16, BLM^{+}16, DSSW16, HICF16, LB15, SURW18, SKPP15, VSA^{+}15, WW17]. Effects [AGAB11, AOPM^{+}19, AASK12, ALL^{+}12, AZCT12, AWT^{+}15, Bee13, BBWK^{+}14, BKM14, BBP12, BP16, BCA^{+}18, BARB17, BRA^{+}19, BBH17, BCGP13, BLM^{+}16, BGG13, CBM^{+}11, CWS^{+}14, CLT^{+}13, CTBR19, CT11, CQC^{+}19, CGHM13, CGCL11, CCBP14, CMJ^{+}11, CD13, CE15, CW10, CFBL10, DDPC12, DHH^{+}12, DMZ10, DWCC17, EVF13, EPD12, EG12, FBA^{+}13, FCGJH11, FBWIO19, FG12, FEC^{+}10, GT10, GLdPL18, GFdM^{+}19, GFM^{+}13, GS11, GLB^{+}13, GG13, GSW10, HBD^{+}14, HTH^{+}10, HM12, HG10, HHJ^{+}19, HSBE12, HMB^{+}13, HNDE13, ISA16, IUM10, Jam11, JMW^{+}11, JMSM13, KMT^{+}17, KO12, KHGF^{+}17, KRL^{+}16, KCSP11, KZG^{+}13, KTN12, KPL^{+}12, LJA^{+}13, LW11, LWZL16, LCFS14, LTQ^{+}13, LW18, LPT12, LP13, MGCL14, MLS12, MLLS14, MSF^{+}12, MYG^{+}14, Mcc10, MHG^{+}14, MBGS11, MC11a, MAWF12, MRRK11, MMCUR^{+}12, ML15, MSSM10, NSBJ17a, NA13]. Effects [OTKE12, OGS^{+}19, PBK^{+}11, PG12, PMS13, PMS^{+}10, Pra10, PMK^{+}10, PCBD11, RSS^{+}13, RM11, RSW11, RCARR18, RSPD13, RJF^{+}13, SC15, SK10, SB16, SAGSHL10, SSP^{+}15, SKCS11, TAG^{+}10, TTSZH10, TLPP17, TMD^{+}17, TVY^{+}16, UMK^{+}16, UIN^{+}13, VMW13, Vég18, VMÁR^{+}13, WBM13, WIM^{+}10, WSHB11, WGS^{+}19, WK10, WWdS^{+}16, WNY13, WGB^{+}11b, WH13, Wit11, ZBD11, ZRM18, BBM18, BCA^{+}19, BSL^{+}19, CBCW18, DMR^{+}17, ESD^{+}17, GSSF19, HH19b, KTBH17, KK18, LS19, MF15, SL15, SJZ19, SSE^{+}17, VMO^{+}18, WHG^{+}15, WC19]. Efficacy [BMC^{+}18, HDN^{+}14b, GSB^{+}16, KLL^{+}16, Les15]. efficiencies [DHG^{+}17]. Efficiency [CLV^{+}10, Dre17]. Efficient [BW10, DFW^{+}10, GPB^{+}15, NBSM12, RK17]. Effort [RMGCR10, WR12, BCM16, RHWS18]. Efforts [PRR^{+}11, BCD^{+}16, LDD19, NBR^{+}15, VW19]. Eficacia [BAPP11]. Eficiencia [CLV^{+}10]. Eficiente [BW10, DFW^{+}10]. Egg [CB13]. Egipcio [LAA^{+}10b]. Egyptian [LAA^{+}10b]. Eight [Ano19w]. Ejercicios [Rud11a]. elasmobranch [WMFB15]. Eléctricos [BAPP11]. Electrocution [DHD14]. Electropositive [BBB14]. Elefantes [Wit11, WRTB10]. Elementos [BB10a]. Elements [SCSJ13]. Elephant [SBS^{+}15, Ben15a, PRM^{+}18]. Elephants [AME^{+}13, EBL^{+}14, Kah10, Wit11, WRTB10, BMF16, KLN^{+}17, SLS17]. Elevation [JMP^{+}11]. Elevational [Cat11, FMJP11a]. Eleven [BMG^{+}14]. Eliciting [MBF^{+}12, GKA^{+}15]. Elsewhere [NS18]. Elusive [BAG^{+}11, KGS^{+}11]. Elusivo [BAG^{+}11]. Elusivos [KGS^{+}11]. Embedded [JMF12]. embedding [RKC^{+}19]. embrace [PWKT18]. Embracing [HAB^{+}11, KHK^{+}19]. Emergence [FCC^{+}11]. Emergencia [FCC^{+}11]. emergency [TWB^{+}18]. emergent [GZR18, PSA^{+}17]. Emergente [SGH^{+}11]. Emerging [KBML15, LPW16, Rey11, SGH^{+}11, Val18, WBPL15]. Emigration [MSSM10]. empires [LET19]. Empirical [GBMRG^{+}10, HCM14, LLW^{+}14, VNAT19, KF18a, TMK^{+}16]. Empírico [GBMRG^{+}10]. Empleo [EMS^{+}11]. Employing [EMS^{+}11]. Empowering [Fra11]. Empresarial [BG10]. Emptiness [Dor11]. enabled [KLL^{+}16]. Encierros [FCGJH11]. Encinos [PTN10]. Encounter [SDB10]. Encouraging [Fra11]. Encroachment [WT10]. Encuentro [SDB10]. Encuestas [DWCW10]. Endangered [ANTBG12, Boo14, BGG13, CWS^{+}14, CVN^{+}10, CFL14, CCM12, Chi10, CD13, DBV13, DWCW10, Duf11, ENW^{+}18, Fli10, Fur19a, GGG10, GAG13, GLNJ10, HMIE14, HB10, JMW^{+}11, KGG^{+}14, LN11, Les15, LRMWG16, MAB^{+}13, NCC13, PS10a, RTT^{+}13, RLGG^{+}12, SRL^{+}13, SGSR^{+}12, SHCD11, SKH13, TG15, VM14, Van12, WL12, Wil17, WLCS14, ZCCN13, ABG^{+}19, ÇRG^{+}17, CCS^{+}18, GGB^{+}15, JLCHNG^{+}17, KP19, LLR^{+}15, LGCF19, ME16, MPS^{+}16, RHCW16, SHB^{+}17, SBU^{+}17, WD17]. endangerment [BBM18]. Endemic [AZT^{+}14, HJM11, MSSM10, PCRD14, SS10a, GML^{+}19, LP16, RBH^{+}15, SMGR17]. Endémicas [SS10a]. Endemism [FB11, SMH^{+}14]. Endemismo [FB11]. Ending [CV16]. Endogamia [Jam11, JMW^{+}11]. Endogámica [GLNJ10]. Enemy [Sal12]. Energético [BB10b]. Energy [BKKM14, BB10b, BB11b, MHG^{+}14, MBL^{+}14, BB15, Die16, FBWIO19, Hen16, HGRS16, KNB^{+}17, SB16]. Enfermedad [FCC^{+}11, HLB^{+}12, LMM^{+}10, MSSM09, PHC^{+}11, SVH12]. Enfermedades [LW11, OTKE12]. Enfoque [DWM^{+}10, LOME10]. Enforcement [NBM^{+}12, CQC^{+}19, DKG^{+}15, DUS^{+}16]. engage [NGB^{+}16, RTB^{+}18]. Engagements [Rob12]. Engaging [DC11, EG13, LBD^{+}15, WLD16].

England [CBA13, FGG^{+}19, SJH13]. Enhance [CLV^{+}10, GC10, VMS^{+}12]. Enhancement [CBF^{+}19, SGE13]. Enhancing [DVW^{+}13, NS10, SPSS10]. enlightened [CS17b]. Enmascaramiento [PHFPP10]. Enough [Sha15b]. Enrichment [KWW^{+}14]. Enriquecen [GC10]. Enriquecimiento [NS10]. Ensambles [CMRH10, CGCL11, DMZ10, SSBN12]. Ensemble [BGB17]. Enseñar [CND10]. entanglements [KRL^{+}16]. Entendiendo [HNS^{+}11, OBBP10, Ser10, TM10]. entire [SWSM19]. Entitlement [McP11]. entre [BMGLW12, BRKP12, CKKL10, Fli10, GT10, HSBJ12, HRKMG12, LDBS11, MLS12, MCL12, ÖB12, PBH^{+}10a, SSVA11]. entrepreneurship [Bus15]. entry [BBB^{+}16]. Envío [OSC10]. Environment [GRA14, BDKS15, DWM^{+}19]. Environmental [AG10, BHK10, BdA15, Bro10, BRM^{+}14, Car11, CGCL11, CCBP14, Dor11, Fer11, Heb12, HSBE12, PG16, PI11, Pin17, PHC^{+}11, SAGSG^{+}14, SMH^{+}14, SMGRJ10, BDKS15, Ben16, BBB^{+}16, CAB^{+}18, CMG15, GBMW15, HW19, Hug19, HN17, KHB18, MCD^{+}16, MGA16, NJBM15, NRRW17, SMZZ17, VB19]. Environments [SRL^{+}13, FCS19, SSC^{+}19]. Epífitas [CT11, MPB10]. Epiphyte [MPB10]. Epiphytes [ECES13]. Epiphytic [CT11]. epistemic [VLSOS16]. Época [Rog11, WK10]. Equality [KNK13, BC15]. Equatorial [GFRK12, VCDF13, AWKR^{+}11]. Equipo [Mcc10, VLRB12]. equitable [GPB^{+}15]. Equity [LBI^{+}18, SG19]. equivalence [ACW18]. Equivalency [HGCT13]. Era [SSBN12, KBML15]. Eradication [BJMT14, OBB^{+}11, AC15, DOC^{+}15]. eradications [HHW^{+}16]. Erie [WCR^{+}15]. erosion [GDB15]. Erradicación [OBB^{+}11]. Errata [Ano11j, Ano12f]. Erratum [Ano12g, Ano13f, Ano13g, Ano14d, Ano14e, Ano14f, Ano14g, Ano15f, Ano15g, Ano16i, Ano17i, Ano17j, Ano17k, Ano19j, Ano19k, DSW16]. Errores [CND10]. Errors [VMÁR^{+}13]. Escala [VMGB^{+}10]. Escalas [CBM^{+}11, KMGCR10, PTN10, PSP11, WB10]. Escarabajos [CMJ^{+}11, EMD^{+}10]. Escuelas [WD10]. Esfuerzo [RMGCR10, WR12]. Esfuerzos [PRR^{+}11]. Espacial [ALL^{+}12, FLP^{+}11, SSK^{+}11]. Espaciales [CBM^{+}11, CJDZ10, CGCL11, HSB12, KGS^{+}11, PSP11, PBB^{+}11]. Espacialmente [BRCK11, HRBR11]. Españoles [MRPPM12]. Especie [BR11b, CMNM10, DWM^{+}10, EG12, JLDJ12, MARV^{+}11, Van12, ZYYH10]. Especies [AWU^{+}11, AG10, BRV10, BHK10, BDH^{+}12, CVN^{+}10, CJDZ10, CMRH10, CLRV11, Chi10, CKKL10, DWCW10, ETB10, FG12, FMJP11a, FMF10, HNS^{+}11, HJSF^{+}12, JMP^{+}11, KG10, LN11, LCG10, MFRH12, MVB^{+}12, MWC11, MSSM09, OKLP11, OSFFB^{+}11, PBH^{+}10a, PTN10, PSAZ11, PAM^{+}10, PF10, RER10, RHM10, San10, SSO11, SD11, SAC12, SS11, SHCD11, SKCS11, TAG^{+}10, TGD10, TA10, TBY^{+}10, VLRB12, VSM^{+}11, WL12, WEC^{+}11, WHB^{+}11, SAB12]. Especies-Área [KG10]. Específicas [ANTBG12]. Especímenes [GCC^{+}12]. Espontánea [ZYYH10]. Essential [KB11]. Estabilidad [SAGSHL10]. Establecimiento [HG10, RRCB^{+}11, WR12]. Establishing [RRCB^{+}11, VNAT19]. Establishment [HG10, WR12]. Estadios [ANTBG12]. Estado [Rud11b].

Estados [LO10, FEC^{+}10, LN11, ML10, RPA^{+}12]. Estar [RHM10]. estate [LMJB^{+}19]. Estatus [BJR^{+}11, HWW12, RDM10, SPPN^{+}12, WHB^{+}11]. Estimación [HRBR11, MVB^{+}10, MVB^{+}11, Ber12, DEJM10, JDA^{+}10]. Estimaciones [KRS^{+}10]. Estimando [HNS^{+}11]. Estimar [CPC^{+}10, KGS^{+}11]. Estimas [WHB^{+}11]. Estimate [HRBR11, KGS^{+}11, WHB^{+}11, MPS^{+}16, RNE^{+}19]. Estimates [KRS^{+}10, EG17, JBH^{+}16, LMJK^{+}19]. Estimating [ACS14, BH18, BGG13, CBB18a, CPC^{+}10, DEJM10, FHBK18, HNS^{+}11, HLHR16, JDA^{+}10, MVB^{+}10, Sch10, SHPR13, SR17, TBB^{+}14, VJG^{+}15, WLCW16, BSCB18, Pet17, SSWR17]. Estimation [Ber12]. estimations [GRT^{+}18]. Estrategias [BRCK11, LH11, MR12]. Estratificación [MPB10]. Estrés [ATW^{+}11, JLDJ12]. Estreses [GKK^{+}11]. Estructura [FLP^{+}11, HSAB10, SHA^{+}10a, SHCD11]. Estructurada [MR12]. Estructuras [FC10]. Estuarina [WGR^{+}10]. Estuarine [WGR^{+}10]. Estudia [BE10]. Estudio [DWCW10, EMD^{+}10, GBMRG^{+}10, MAA^{+}10, TA10, WHB^{+}11]. Estudios [BM10, HHDM10]. Esturión [DBP^{+}10]. Etapas [GLNJ10, WK10]. Eternal [Sha10b]. Ethic [Mar10, SS10b]. Ethical [San10, WZC19, DFR^{+}17, JMH15]. Ethics [HMG^{+}13, San14, WT13]. Ethnoecological [MAGA10]. Éticos [San10]. Etnoecológico [MAGA10]. EU [RCK13]. EUA [Hic10]. Eucalyptus [SBG^{+}18]. Eurasian [RL11, ZMTR15]. Euro [PS13]. Europa [NKB10, PBH^{+}10a]. Europe [BVC^{+}14, CHN^{+}15, DSE15, KLS^{+}19, MAM^{+}15, MV18, MRMV14, NKB10, PBH^{+}10a, RV15, SBB^{+}15, SKSS^{+}17]. European [BV10, BHF^{+}19, CLT^{+}13, DP15, FZ18, HHB15, KHD^{+}15, KRF^{+}19, LKH^{+}13, LKW^{+}15, SFAA^{+}18, TTSZH10]. Europeo [BV10]. Europeos [TTSZH10]. Euthanasia [LCL13]. Evaluación [AWU^{+}11, BJR^{+}11, LCG10, MWC11, SFM^{+}11, AH12, BMP^{+}12, BR11b, HRC10, HWW12, KUT11, KMGCR10, LMM^{+}10, LGH^{+}12, MTM12, MBJWC10, MAGA10, PRR^{+}11, SSD10]. Evaluaciones [BMP^{+}12, MA10]. Evaluar [AWKR^{+}11, ESCF12, JLDJ12, LW11, LHDC12, WHB^{+}11]. Evaluate [CRKM^{+}14, WHB^{+}11, BBM18, NMS^{+}18]. Evaluating [BW13, BMP^{+}12, BHM^{+}14, BGD^{+}15, BGL^{+}14, CC14, GSB^{+}16, HRC10, ITG^{+}16, Jen14, KUT11, MCD^{+}16, MJG15, MAGA10, SSD10, TIB^{+}16, VBN15, WH13, CL16, SDT17, TND16]. Evaluation [AH12, GB14, LMM^{+}10, LH19, LWLC12, MBM^{+}14, MRPT19, PWP^{+}18, SMV19]. evaluations [MRFvW16, SBG^{+}18]. Evelyn [GB11]. Eventos [PMK^{+}10]. Events [PMK^{+}10, ALW^{+}19]. Every [AA12, GHKK10, PI11]. Everything [Mih19]. Evidence [BB11b, BBP12, FBF^{+}14, Fli10, Fre11, LDC^{+}15, LZD^{+}13, LCT^{+}14, MSSM10, SBBP11, SM11, SKH13, Ben16, CAB^{+}18, KF18a, WDS15]. Evidence-based [LDC^{+}15]. Evidencia [Fli10, BBP12]. Evidencias [SM11]. Evitada [HRBR11]. Evolución [RSW11, Van12]. Evolution [Bra15, CG13, Dol10, Fie15, RSW11, Van12, Whi13]. evolutionarily [CS17b]. Evolutionary [HP11, RDM10, TCDR12, CS19, MF15, PBB^{+}15].

Evolutiva [RDM10, TCDR12]. Ex [MS12, MVB^{+}12, ZYYH10, CCW^{+}16]. Examinada [DDPC12]. Examined [DDPC12]. Examining [BIF^{+}18, CCG18, PMM^{+}15]. examples [WK15]. Excelencia [BG10]. Excellence [BG10]. Exceptional [PG18]. exchange [GGL^{+}15, NYC17]. exclusion [QCMB18]. Excretas [VSM^{+}11]. Excursionismo [MFN^{+}11]. exemplary [MMGN18]. Exercises [Rud11a]. existing [DLMAG15]. Éxito [FCGJH11, HM12, MS12, VMS^{+}12, WFM^{+}10]. Exogámica [FBE^{+}11]. Exotic [BBM14, Pra10]. Exótica [Pra10]. Exóticos [SS10a]. Expanding [ANE^{+}16, SCA17, WRG^{+}12, ZAP^{+}18]. Expansion [CFO^{+}14, FMJP11b, WGF^{+}12, WRG^{+}12, HWB^{+}16]. expenditure [JRWA15]. Experiences [JMF12, Mel10]. Experiencing [MQ14]. experiment [KMT^{+}17, ML15, SSMC17, WFMW^{+}18]. Experimental [BBP12, CMRH10, EG12, SJP^{+}13, KGF^{+}18]. Expert [DWCW10, HKB^{+}17, MBF^{+}12, Poz17, HBTB15, SBG^{+}18, WD17, WCR^{+}15]. Expertos [DWCW10, EMS^{+}11, MBF^{+}12]. Experts [EMS^{+}11, HSRC14]. Explain [CEM^{+}13]. Explanations [Fer19]. Explicit [BRCK11, EIS14, HRBR11, Cha15, JHAH18, MPS^{+}16]. Explícita [HRBR11]. Explícitos [BRCK11]. exploitation [BFF^{+}16, KBML15]. Exploración [WRTB10]. Exploration [WRTB10, KBML15]. explore [GGB^{+}15]. Exploring [SCSJ13, Bra15]. Explotación [DMZ10]. exposed [MA19, OGS^{+}19]. Exposing [KMR^{+}18]. Exposure [KGG^{+}14]. extension [LTN18]. Extensiones [ALL^{+}12]. Extensive [RSPD13]. Extent [BMG^{+}14, MC11a, Sou10, BHPM^{+}18, CVMM17, CMM17, HRHC16, HLHR16, JBH^{+}16, MM15]. Extents [ALL^{+}12, CBM^{+}11]. Extinción [FB12, FP10, GCC^{+}12, LO10, MVB^{+}12, PRR^{+}11, PGLB^{+}12, PHC^{+}11, SWL^{+}10, SSB^{+}12, WKP11, SHLK^{+}10]. Extinciones [Boy10, PAM^{+}10]. Extinct [RER10, RHM10, San14, TBB^{+}14, TBY^{+}10, ZYYH10, DBV^{+}18, QSCM^{+}19]. Extinction [AGLRV14, Ano16x, Bai10, Bai13, BWE^{+}13, BCGP13, CCBP14, Fer17, FB12, FP10, GCC^{+}12, HSR^{+}13, HWM10, HJM11, KMH^{+}14, LO10, LFPT14, MMK^{+}14, MVB^{+}12, PRR^{+}11, PGLB^{+}12, PHC^{+}11, RTT^{+}13, RBH^{+}15, SHG^{+}14, SHPR13, SWL^{+}10, SSB^{+}12, SB14, SHLK^{+}10, SKSK14, WKP11, BIF^{+}18, BSCB18, CBB18a, CBB18b, HDH^{+}15, JBH^{+}16, KSH17, MNW^{+}17, ML15, Pyk17, QCMB18, RJS17, RTH^{+}15, SC15, SBB^{+}15, TBM^{+}16, VJG^{+}15]. Extinctions [Boy10, GH10, MLA^{+}14, PAM^{+}10, SW10, CGS^{+}16, JBB^{+}15, RWF^{+}17, SDT17, WGLL17]. Extinta [ZYYH10]. Extintas [RER10, RHM10, TBY^{+}10]. Extirpación [PGLB^{+}12, PB11]. Extirpation [PGLB^{+}12, PB11]. Extirpations [Pri10]. Extracción [LMM^{+}10]. extracting [BMC^{+}18]. Extraída [GOR^{+}12]. Extrapolating [MRMH17]. Extrinsic [Whi13]. Exurban [LPT12, LP13]. Exurbano [LPT12]. Eyvindson [NSBJ17b].

F [Ano16x, Fie15]. Face [HFK^{+}10, HHA^{+}11, HDHL10, CCW^{+}16, HGRS16].

Facebook [PNC^{+}15]. Facets [BB10a]. Facial [HLB^{+}12, LMM^{+}10]. Faciales [HLB^{+}12]. Facilitate [GRA14, LMM^{+}17]. Facilitating [BKR17, RCSK15]. facilities [JC18, SB16]. facto [AG17]. Factores [BMGLW12, GGG10]. Factors [BMGLW12, BGP16, Fre11, GGG10, HFBG18, HPL14, KGG^{+}14, NJT14, SBB^{+}14, ZAB16, LFM17, MCO^{+}17]. Facts [Laz11]. Facultative [OTKE12]. Facultativos [OTKE12]. Fads [RPS13]. failings [WGLL17]. Failure [HB10, CRMK18, CBWFL15, EEB15]. fairly [Arm19]. Falacias [KE10]. Falcons [KTN12]. Fallacies [KE10]. Fallow [BTBDS12]. Faltantes [AG10, NBW^{+}12]. Familiares [FWM^{+}12]. Familiaridad [MS12]. Familiarity [MS12]. family [Gru17, RBH^{+}15]. Farm [UIN^{+}13]. farmers [KLN^{+}17]. Farming [GKB10, HRKMG12, KE10, MBJWC10, NBA^{+}18]. Farmland [Die10a, MDJ14, FGG^{+}19, Poz18]. farmland-dominated [FGG^{+}19]. Farms [Sko19]. Fascinating [Fie15]. Fatal [FCC^{+}11, OCR^{+}19]. fatalities [KNB^{+}17]. Fate [JPT13, DS19]. Fates [PA12]. Fatty [MPH^{+}11]. Faulty [BB11b]. Fauna [BHSA12, CFO^{+}14, SSBN12, HTDM17, WDW^{+}19]. Faunal [SSBN12]. Favor [HSAB10]. Favorece [HSAB10]. Favourable [LJA^{+}13]. Feasibility [LSA17, TTEM14, HAF^{+}17, PRNV18]. feather [FMB^{+}18]. Features [MRRK11, SGS^{+}10, BGP16]. February [Ano19b, Ano16a]. fecal [WLCW16]. Fecundidad [Rog11]. Fecundity [Rog11]. Feedback [MRMV14, vdWSM^{+}16]. Feedbacks [MCL12, JGS^{+}16]. feeding [EWCG15]. felid [DHM^{+}15]. Fence [BW10, FEC^{+}10]. Fences [Qui14, KLN^{+}17]. fencing [RWF^{+}17]. Fenotípica [RSW11]. Feral [LL14]. Fernandez [SPL^{+}11]. Fideicomiso [AH12]. Field [HCM14, KWW^{+}14, KK18, SFKC16]. fight [LFMB19]. Fights [Ibi18]. Fiji [GPB^{+}15, MJP^{+}11, WJ13]. Filipinas [WRAW10, HRKMG12]. Fill [JDA^{+}10]. Filogenética [FP10, PBB^{+}11, WR12]. Filogenético [LOME10]. Filogenéticos [Fli10]. filter [GBB^{+}15]. Filters [FJP^{+}11, Sha15a]. Filtros [FJP^{+}11]. fin [FFS^{+}18]. Finally [Fur18]. Finance [AH11]. financial [CC18]. financing [LFS^{+}17]. Finas [AGAB11]. Finches [ZLKW13]. Finding [Kar14, Wat11]. Finds [Car11]. Fine [SMAG17]. Fine-resolution [SMAG17]. Finest [BBC^{+}14]. Finning [CHHR13]. Fire [BSF^{+}14, GSW10, Mat13, NKSB^{+}13, NA13, SMB^{+}10, SKCS11, WGB^{+}11b, AWHW17, DBB^{+}16, FKRC17, KBCM15, Mat15, SVN^{+}15]. fire-dependent [AWHW17]. Fire-Prone [NKSB^{+}13, SMB^{+}10]. fires [CWS16]. First [Whi13, SSD^{+}18]. Fischer [MBT^{+}17b]. Fish [Bar12, CGCL11, DMZ10, FCGJH11, HPL14, LCT^{+}14, MO11, CEdK18, CBWFL15, HTDM17, HTH^{+}17b, KRF^{+}19, LW18, SJD^{+}19, TPL18, VESHR18]. Fisher [HRKMG12]. Fisheries [AS10, BT14, CGHM13, Fur18, GFG^{+}12, GC10, KKCS10, Lor13, McC14, MJ11, Mcc10, MO11, Öst14, CBM16, DWCC17, JRWA15, LCYJ19, MGMC15, MK18, MDSF19, NCKC17, OJAV15, RKW19, SG18, VGAB^{+}19]. fishers [TMM^{+}18]. Fishery [BJR^{+}11, HSBJ12, LGK^{+}11, SGS^{+}10, SHC^{+}11, WB14, BTTP17, JRWA15]. Fishery-Induced [SGS^{+}10]. Fishes [AWU^{+}11, AMO11, BCGP13, HJM11, PF10, PCRD14, BHPM^{+}18, BCSM15, CESS^{+}18, GHT^{+}15, KSH17, LDD19, LB15]. Fishing [BBB14, BCGP13, CESS^{+}18, FCGJH11, GOR^{+}12, GFGR12, MCN10, Mcc10, ÖB12, TBY^{+}10, WGR^{+}10, CE15, GD15, KRL^{+}16, OKCG18, TGBB^{+}15, WFC^{+}18, WHG^{+}15]. Fishing-gear [CESS^{+}18]. Fisiológico [ATW^{+}11]. Fit [Pri10]. Fitness [BW13, RLGG^{+}12, SS12]. Five [CPR^{+}16]. Fixing [LTQ^{+}13]. Fixism [RFV^{+}17]. Flagship [KCD^{+}17, VPS^{+}14]. Flawed [HTH^{+}10]. flaws [OCR^{+}19]. Fletcher [ABD^{+}17]. Flexibility [EWN^{+}13, LFS^{+}17]. Flexible [HFSB13]. Flight [Ore12]. Flightless [WWW^{+}12]. flights [PRB18]. Floodplain [BP13]. floods [BMA^{+}18]. Flores [CBM^{+}11]. Florida [CFBL10, JLCK12, WLL^{+}16]. Floristic [WRG^{+}12]. Florística [WRG^{+}12]. Flow [EPD12, KPL^{+}12, IRK^{+}18, KP16, KP19, VNAT19, VKM^{+}17, WW17]. Flowers [CBM^{+}11]. Flujo [EPD12]. Flujos [KPL^{+}12]. flunixin [ZMTR15]. fluvial [TPL18]. Flyway [IFP14, YJL^{+}18]. Focal [LLW^{+}14, TND16]. Focas [AGAB11]. Focus [DWM^{+}10, SG18]. Focusing [Wal17]. Following [AEO^{+}13, HJMM14, WGF^{+}12, DSK^{+}19, SR17]. Fondo [HSBJ12]. Food [BGG13, Cop12, HSBJ12, Ibi18, Mcl11, GHM^{+}15, GSSF19]. footprint [BH18]. Forcings [Lal11]. Ford [GZR18]. forecast [WCR^{+}15]. Forero [Cat11]. Forero-Medina [Cat11]. Foresight [CWAW14]. Forest [AGP^{+}14, BMGLW12, BMM^{+}14, BEE^{+}11, CRKM^{+}14, Csó17, DMM14, DFJ^{+}12, EBL^{+}14, GT10, GS11, GLB^{+}13, HTH^{+}10, HG10, HRBR11, HNDE13, Ibi18, KWW^{+}14, LBLB14, LFPT14, LKH^{+}13, MLLS14, MSMCB18, MRBL^{+}12, MPB10, Per11, PFR^{+}14, Pra10, RYW^{+}15, SSD^{+}18, Sel16, SMGRJ10, TAG^{+}10, Van12, VLGRS14, WRTB10, ZBD11, BP16, BVAC17, BARB17, BRA^{+}19, BMF16, BMG^{+}17, CPR^{+}16, CWS16, CMG15, CGE19, EGB18, KHGF^{+}17, KLS^{+}19, LWZL16, LPA^{+}17, Mat15, PRM^{+}18, PSA^{+}17, SLS17, SLWL15, SEM^{+}15, SSE^{+}17, SAS^{+}19, WCCT18, LW10]. forest-conservation [PSA^{+}17]. forest-management [BARB17]. Forest-Regeneration [BEE^{+}11]. Forestal [ZBD11]. Forestales [GT10]. Forested [HD14, ERW18]. Forestry [LBLB14, SGE13, HOC^{+}16]. Forests [AEO^{+}13, APL^{+}13, CMRH10, CS13, DFJ^{+}12, ETB10, GLB^{+}13, HODB10, IP11, IUM10, LB11, LGB^{+}10, Mat13, MTN13, Mic15, NFdC^{+}12, Nov16, PBH^{+}10a, PBH^{+}10b, Put11, RSS^{+}13, RSPD13, SAC12, SMB^{+}10, Sta16, BARB17, BTB^{+}18, CZS19, DRA^{+}15, FTB^{+}16, HSB^{+}16, PR17, PRM^{+}18, SBB^{+}15, SVN^{+}15, SKSS^{+}17, WZWF18]. Forgetting [RPS13]. Form [KZG^{+}13]. Formation [BMG^{+}14]. Fortalecen [Fli10]. Forward [FBF^{+}14]. Fotogrametría [Ber12]. Fouls [CND10]. Foundations [CP14, San10, Ano16x]. Founder [Jam11]. Four [Aro11, FMV^{+}19, Jam11, JBB^{+}15, KRF^{+}19]. Fourth [WT13]. fox [MCO^{+}17]. Foxes [CM13]. Fracaso [HB10]. Fraction [CB13]. Fragment [MRBL^{+}12]. Fragmentación [CFBL10, GZR12, HG10, VJMH12]. Fragmentados [LH11, MLS12, WW10a]. Fragmentation [CFBL10, Csó17, DMM14, EVF13, GZR12, HG10, KZG^{+}13, MLLS14, SHPR13, VJMH12, TPL18]. Fragmented [LH11, MLS12, MHG^{+}14, PY13, WW10a]. Fragmentos [MPB10, TAG^{+}10]. Fragments [AGP^{+}14, MPB10, MC14, TAG^{+}10]. Frame [RM14a, GKA^{+}15]. Frames [Qui14]. Framework [BGAS10, Boo14, HHA^{+}11, MFN^{+}11, NFF^{+}10, PCW14, SD11, AMS^{+}16, CL16, CSR^{+}19, GHM^{+}15, GW17, JMH15, MFHT19, Mor15, TIB^{+}16]. Framing [Agg15, LKW^{+}15, CCS^{+}18]. Franceses [CPC^{+}10]. Frankenstein [OSF10, SPL^{+}11]. Frecuentes [PMK^{+}10]. Free [ATW^{+}11, LM18]. free-ranging [LM18]. Free-Tailed [ATW^{+}11]. French [CPC^{+}10]. Frente [BM10, HLB^{+}12, HFK^{+}10]. Frequency [HAP^{+}13]. Frequent [PMK^{+}10]. Freshwater [OKLP11, PF10, SKSK14, BCA^{+}18, BCA^{+}19, BFB^{+}19, CBWFL15, GHT^{+}15, KSH17, LS19, SDT17, SR17, SURW18, VESHR18]. Friendly [GKB10]. Frog [HPL14, MAWF12, SHSP19, ALW^{+}19, RHCW16, TMD^{+}17]. Frogs [KPL^{+}12, PGC^{+}13b, SAGSHL10]. Front [HLB^{+}12]. Frontera [FEC^{+}10]. Fronteras [ZM12]. Frontier [LK13, MG12, CNK^{+}17, OQGR^{+}15]. Frontiers [ZM12, SSRS18]. Fronts [Fur19a]. frugivorous [GSB^{+}16]. Fuego [CMRH10, GSW10, SKCS11, WGB^{+}11b]. Fuente [Rog11]. Fuente-Sumidero [Rog11]. Fuentes [GZD11]. full [MM15, San19]. Fuller [Ano16x]. Fully [BHSA12]. Función [ETB10, PHFPP10]. Funciona [SHO^{+}10]. Funcional [JLDJ12, MLS12]. Funcionales [SAB12]. Function [BCT^{+}14, ETB10, PHFPP10, ZLKW13, CTRE^{+}17]. Functional [GLB^{+}13, JLDJ12, KZG^{+}13, MLS12, SPN^{+}14, SLS17, SAB12, ACW18, GDM^{+}19, JVI^{+}18, LW18, RJS17, SEM^{+}15, WTT^{+}17]. Functions [HGCT13]. fund [HHW^{+}16, NW15]. Fundador [Jam11]. Fundamentos [San10]. Funding [RPS13, LHKA16, SSMC17]. funds [HFBG18]. fungal [BGH^{+}15, HCBB^{+}15]. fungi [PWKB19]. Fungus [MSSM09, PMS^{+}10, Sch10, SSP^{+}15]. further [IPA^{+}17]. Futuras [MVB^{+}10]. Future [Bur13, Cla13a, Col10, Cop12, CKK^{+}14, CEM^{+}13, EPB^{+}14, Fer11, HP11, JBB^{+}15, Lic16, Mic15, MMM^{+}12, MVB^{+}10, NMSF11, SKC^{+}12, Yap11, ALW^{+}19, BKR17, CBM^{+}18, CE15, DBB^{+}16, DSV17, FDBA16, GML^{+}19, LMM^{+}17, ODAL17, PBB^{+}17, VESHR18]. Futuro [MMM^{+}12, SKC^{+}12]. Futuros [Col10]. Fynbos [Fie15].

G [Fie15]. G. [GB11]. Gabon [CAB^{+}10, CSMG^{+}13, EBL^{+}14, FWM^{+}12, VKKA13, VNAT19, CAB^{+}10, FWM^{+}12]. Gain [JGH^{+}11]. gains [CESS^{+}18, RTGR17]. Galapagos [LZD^{+}13, QSCM^{+}19, ZLKW13, HGCT13]. Gamo [BTBDS12]. Ganado [EDB12, GSW10]. Gap [JDA^{+}10, MQ14, MJP^{+}11, PGBV^{+}13, GDG19, KMR^{+}18, SLA^{+}19]. Gaps [AKJ^{+}18, Bro13, VMÁR^{+}13, RKW19, WBPL15]. Garde [BLG^{+}11]. Garden [SCSJ13]. Gardens [HFL^{+}11, Hul15]. Garnett [KF18a]. gas [MGA16]. Gastos [HSB12]. Gazettement [BMGLW12]. Gear [BBB14, Mcc10, VLRB12, CESS^{+}18, WHG^{+}15]. gene [KP16, KP19, WW17]. Generación [RBC^{+}11]. General [GL18, KAV18]. Generalists [SKSK14]. generalized [SKLP16]. Generated [PMS13]. Generating [Dre17].

Generation [RBC^{+}11, WON^{+}13, FW17, KNB^{+}17, PWKB19]. Generic [NKB10]. Genérico [NKB10]. genes [KBSX19]. Genetic [AWU^{+}11, BAJ^{+}10, CFL14, CBSZ10, CM13, EBL^{+}14, Fli10, HAV11, Jam11, LAA^{+}10a, LKH^{+}13, MMM^{+}12, RBW14, SBT^{+}11, SHA^{+}10a, VJMH12, AMS^{+}16, DBV^{+}18, GRT^{+}18, HM16, HAT^{+}15, NBH^{+}17, SSWR17, WC19]. genetic-based [AMS^{+}16]. Genética [CBSZ10, Fli10, Jam11, SBT^{+}11, SHA^{+}10a, VJMH12]. Genetically [BHK10, QSCM^{+}19]. Genéticamente [BHK10]. Genéticas [AWU^{+}11]. Genético [MMM^{+}12]. Genéticos [BAJ^{+}10]. Genetics [Cha10, LLR^{+}15, SLA^{+}19]. Genome [EBP^{+}18, QCMB18]. genomic [WML^{+}19]. Genomics [Whe11, ÇRG^{+}17, EBP^{+}18, HJH16]. geodiversity [ACB^{+}15, CPH^{+}15, HGGH15, TBF^{+}17]. Geografía [CVN^{+}10]. Geográfica [TCDR12]. Geográficos [LN11]. Geographic [LN11, MCMP15, RTH^{+}15, BJM15, MSY^{+}15, SGA^{+}19]. Geographical [TCDR12]. Geography [CVN^{+}10, MR16]. geological [KT16]. Geomorfología [IUM10]. Geomorphology [IUM10]. Geophysical [ACS14, SSW15]. geopolitics [HBD^{+}19]. Gestión [RBC^{+}11]. Getting [Gru17]. Ghats [GBG^{+}10, GBG^{+}10]. Giant [HGCT13, ZZW^{+}10, LP16, WSH^{+}15]. Gigantes [ZZW^{+}10]. Gilded [SHC^{+}11]. gillnet [NCKC17]. Ginseng [SM11]. Gittleman [Ano16x]. Given [AH11]. Global [AAS^{+}17, BMB^{+}15, BSF^{+}14, BJR^{+}11, BBM14, CML^{+}19, CGHM13, Fur18, GDG19, Han10, HH14, HSB12, JCA^{+}11, KOA^{+}15, LDC^{+}13, LKB^{+}14, MBM^{+}14, ÖB12, RN17, Rud11b, SSW15, SHWT14, Sou10, Sto12, Wat11, ALA17, AKJ^{+}18, BB15, CCB^{+}19b, DPGY16, DHM^{+}15, DWCC17, EGB18, GDB15, GBB^{+}19b, JWS^{+}18, LMJB^{+}19, LKYL15, LFMB19, LGA17, MM15, MBV^{+}16, NBR^{+}15, OHHE16, RKW19, RS18, SCHM18, ZV19, vBSC^{+}19]. Globales [BJR^{+}11, JCA^{+}11]. Globalization [BBM13]. Globally [GJB14, SNH^{+}14]. glyphosate [KS17]. Go [BBBG10, LGM^{+}13]. Goals [AH12, HH14, MO11, SKC^{+}12, SG19]. Goat [GGM14]. Gobi [ESD^{+}17]. Going [McP11]. Golden [KNB^{+}17, MFN^{+}11]. Gone [TBB^{+}14]. Good [BCT^{+}14, LGM^{+}13, Sal12]. goods [SLSH15]. Googling [PMP14]. Gorongosa [CTRE^{+}17]. Governance [BRS11, Jos12, MJ11, RBD^{+}13, GMG^{+}19, MA19, NJK^{+}19, SMLB19, WHS18]. Governing [Ben15c, CGK^{+}19]. Gowdy [PI11]. gradient [DMF^{+}18]. Gradiente [CMB12]. Gradients [CMB12, GS16, RK17]. Gradual [BBWK^{+}14]. Graduate [BDG^{+}13, KTF^{+}14]. Grain [SA14]. Gran [BPW12, BB10b, EDB12, LGK^{+}11, MG12]. Grandes [ALL^{+}12, MTM12]. Grapes [KTN12]. Grappling [SH13]. Grasos [MPH^{+}11]. Grassbird [MGCL14]. Grassland [AB13b, BBC^{+}14, BKM14, HD14, RWB11, RBW14, SB16, ZRM18]. Grasslands [BEE^{+}11, MHG^{+}14, PGC^{+}13a, Sha10b, DPS^{+}16, EB15]. Gravity [RL11]. Gray [CMB12, KMLC19]. Grayling [WK10, WES^{+}13]. Grazing [GSW10, GZJ^{+}13, KMT^{+}17]. Great [AB13b, EDB12, Zac10, BPW12, CPM^{+}18, LGK^{+}11, NJBM15]. Greater [BBP12, LA13, MHG^{+}14, BB10b]. Green [ABB^{+}18, SAGSHL10, SAS^{+}19]. greening [TWZ16]. greening-of-religion [TWZ16]. Griffon [ZMTR15]. Gris [CMB12]. Grizzly [CRG12, FWM14]. Ground [EMD^{+}10, Kar14, LB11, Sou14a, IBA16, SMLB19]. Ground-Nesting [LB11]. Group [Fat14, SEM^{+}15]. Groups [JHV13, SAB12, LSV^{+}19]. Grouse [BBP12]. growing [DRA^{+}15]. Growth [ANTBG12, JMW^{+}11, KZG^{+}13]. Gruesa [BRV10]. Grupos [Duf11, FWM^{+}12, FTSSW12, SAB12]. Gubernamentales [HSB12]. Guiar [BRV10]. Guidance [BVC^{+}14, HHDM10]. Guide [BRV10, BKKM14, BDG^{+}13, MB14, SBBP11, JMH15, SSMC17, TUHK^{+}17, WCCT18]. Guidelines [GZA10, RHT14, TG15]. Guides [HCM14]. Guiding [LH10]. Guinea [AWKR^{+}11, BTB^{+}18, GFRK12, Mel10, Nov16, VCDF13]. Gulf [CTM^{+}14, WHE^{+}17]. Guyana [SMY^{+}17].

H [Ano16x, Fie15]. Habitadas [OBB^{+}11]. Habitat [BFHP10, BKKM14, CMB12, CFL14, CFBL10, EBL^{+}14, GZR12, HSAB10, HMM10, JLDJ12, LHDC12, MNW^{+}17, MMPF10, RC12, RTL^{+}12, Sel16, SGE13, SHLK^{+}10, SKSK14, VJMH12, WGB^{+}11b, ZBD11, BFHP10, GZR12, HSAB10, HMM10, JLDJ12, LHDC12, MMPF10, RC12, RTL^{+}12, SHLK^{+}10, VJMH12, WGB^{+}11b, ZBD11, ARDG17, BPN^{+}16, CMMP15, CMM17, Dre17, HBTB15, HLS^{+}18, HWB^{+}16, IRK^{+}18, JRWA15, JBB^{+}15, KBC^{+}18, LSK^{+}15, MAJ16, MAG18, OSM^{+}17, ROQR^{+}16, SBR^{+}19, SWSM19, TMM^{+}18, TGBB^{+}15, TVY^{+}16, WFMW^{+}18]. habitat-based [TMM^{+}18]. habitat-network [IRK^{+}18]. Habitats [BS10, Bra15, DFW^{+}10, EMD^{+}10, JMSM13, PCRD14, RCK13, Sco16, BS10, CMB12, DFW^{+}10, EMD^{+}10, BBM18, RSN19, WTT^{+}17]. Hacia [MRPPM12, AH12, BSBM11, HHA^{+}11]. Haciendo [BM10]. Haddad [Ano19w]. Halcones [KTN12]. Half [Cox10, KMLC19]. Halmahera [SHA^{+}10a, SHA^{+}10a]. Halme [PBH^{+}10b]. Hamilton [KLL^{+}16]. Hand [LGM^{+}13]. Hand-In-Hand [LGM^{+}13]. Handling [Ano12a]. Hanson [SMB^{+}10]. Harassment [AGAB11]. hardcover [Ano16x, Ano19w, Bra15, Fie15]. Hardening [SK10]. Hardwood [LB11]. Harnessing [LSV^{+}19]. Harvest [CTM^{+}14, CS13, GT10, GBB^{+}14, BHB^{+}19]. Harvested [LGH^{+}12, PR17]. Harvesting [CBM^{+}11, CT11, RMGCR10, SKH13, WW10a]. Hastening [RBD^{+}13]. Hatching [HB10]. Hawaiano [Van12]. Hawaii [HRSK12, LCL13, WLL^{+}16]. Hawaiian [AZT^{+}14, ER14, HRSK12, LL14, RBH^{+}15, SBT^{+}11, Van12]. Hazzah [Dys15]. Health [RMRO14]. Heart [Hal18]. Hearts [Ver16]. Heating [Der13]. Heavily [MTN13]. Hedge [BP13]. Height [Van12]. Heller [FS15]. Help [PGBV^{+}13, RTT^{+}13, Abr15]. Hemisferio [LSBB10]. Hemisphere [Fie15, LSBB10]. Henderson [Ano19w]. Herbivores [VMS^{+}12, RBF15]. Herbívoros [VMS^{+}12]. Here [Löv10]. Heritage [AKJ^{+}18]. Heroic [Tor11].

Herpetofauna [MBG14, SHA^{+}10a, UCFV10, KMT^{+}17, SNK15]. Herramienta [AOLK11, RMGCR10]. Herramientas [WRTB10]. Heterogeneidad [CPC^{+}10, VMGB^{+}10]. Heterogeneity [CPC^{+}10, VMGB^{+}10, HN17, NSBJ17a, NSBJ17b]. Heterogéneos [GZR12]. Heterozygosity [RLGG^{+}12]. Heterozygosity-Fitness [RLGG^{+}12]. Hibernating [FCC^{+}11, ISA16]. Hibridación [TGD10, ZYYH10]. hidden [FCS19]. Hidrotermales [GZD11, VLRB12]. Hierarchical [CJDZ10, FWM14, JGL^{+}16]. hierarchy [SG18]. High [BME^{+}14, BEE^{+}11, LK14, SSD10, SHC^{+}11, SHLK^{+}10, AGFR18, LSA17, MRMH17, PRB18, SCHM18]. High-Altitude [BME^{+}14]. High-Quality [SHLK^{+}10]. High-Resolution [LK14, LSA17]. high-risk [PRB18]. Highly [LJA^{+}13, BCSM15, ERW18, PA18, RDS17, THD^{+}18, WTT^{+}17]. Hiking [MFN^{+}11]. Himalaya [BARB17, PG12]. Himalayan [BRA^{+}19, DRF^{+}18, EKRW17, ERW18]. Hindrance [OSC10]. Hiperabundancia [TAG^{+}10]. Hipótesis [FC10, PAM^{+}10]. Historia [GLNJ10, PMK^{+}10, SH11b]. Historians [Poo13]. Historias [AWU^{+}11]. Historic [GZJ^{+}13]. Historical [AS13, BV10, BTTP17, DWM^{+}19, Duf11, MO11, Bai16, FKRC17, MCH^{+}16, PRNV18, SKSS^{+}17]. historical-fire-regime [FKRC17]. Históricas [BV10, MO11]. Históricos [Duf11]. Histories [AWU^{+}11, WGJ13, AK16, KBCM15]. History [CS13, Cor19, Dre11, GLNJ10, Kal18, Pin17, PMK^{+}10, RB13, RL13, SH11b, AOPM^{+}19, Cha15, DS19, DP15, GRT^{+}18, Mat15, SSD^{+}18, SGA^{+}19, TLPP17, TLT16, WCCT18, YFB15]. Hobbs [FS15]. Hoc [Ano12a]. hochstetteri [GLNJ10]. Hoffman [AR10]. holistic [AMS^{+}16]. Hollywood [SCWN18]. Holocene [Col10]. Holoceno [Col10]. Homage [BLG^{+}11, GB11]. Home [CWS^{+}14, GLJ^{+}16]. homonyms [RCBT18]. Honeybee [JDA^{+}10]. honeybees [CM17]. Hong [FFS^{+}18]. Hongo [MSSM09, PMS^{+}10]. hook [SR17]. Hopcraft [FHK^{+}15]. Hope [RAP16]. Hormigas [FG12, SKCS11, MMPF10]. Hormona [JLDJ12]. Hormone [JLDJ12]. Hormones [AME^{+}13, ML14]. horn [MLMP^{+}15]. horticultural [HLHR16, NRRW17]. Hospedero [BAG^{+}11]. Hospederos [IP11, MVB^{+}12]. Host [BAG^{+}11, IP11, MSSM10, SGH^{+}11, WNY13, BMAS17, CBPM19, SHB^{+}17]. Hostigamiento [AGAB11]. Hosts [MVB^{+}12, SSP^{+}15]. Hot [LFMRD12]. Hotspot [Fat14, MMG^{+}13, BRA^{+}19, CQC^{+}19, LKYL15, LCC^{+}15]. Hotspots [CKPS10, FFM^{+}14, JBB^{+}15]. Households [FWM^{+}12]. Housing [PFR^{+}14]. howler [GFdM^{+}19]. Huéspedes [SGH^{+}11]. Human [AME^{+}13, AA12, AOLK11, BGAS10, BA13, BAGM13, BCGP13, CRG12, CGHM13, CW10, FG12, Fre11, FLP^{+}11, Fur19a, FHK^{+}15, GMBM13, GJB14, HP11, HBM^{+}12, Joh10a, KO12, KCDC10, LDBS11, MSF^{+}12, Mih19, MGMA^{+}14, RMRO14, SMK14, SK10, SSK^{+}11, SBB^{+}14, VKKA13, Vég18, Wat11, ZM12, AC15, BTB^{+}18, CGK^{+}19, DWM^{+}19, GHM^{+}15, GMNO15, GSSF19, HSV^{+}19, HCSM15, KLL^{+}16, LNNG16, MTD16, MRMH17, Mor15, PAMA17, PBB^{+}17, Poz18, RN17, SSRS18, SGA^{+}19, SCHM18, TRPD16, WGS^{+}19, WW16, ZV19]. Human-Carnivore [BGAS10]. Human-Dominated [AME^{+}13, AOLK11, Mor15]. human-enabled [KLL^{+}16]. human-impact [ZV19]. human-induced [PAMA17]. Human-Use [SSK^{+}11]. human-values [HCSM15]. Humana [FG12, FLP^{+}11, HBM^{+}12, ZM12]. Humanas [CW10, KCDC10]. Humano [LDBS11]. Humanos [AOLK11, BGAS10, OBB^{+}11, SSK^{+}11]. Humans [OBB^{+}11, Sul10]. Humedal [ODTS10]. Humpback [RMKM14, GGL^{+}15]. Hundred [MLA^{+}14]. Hunter [GFRK12, RMGCR10, VMGB^{+}10, FWC^{+}15]. Hunters [VCDF13, JKJ^{+}19]. Hunting [AWKR^{+}11, BL17, BEA^{+}13, CSMG^{+}13, FWM^{+}12, GFRK12, HCLW13, JVT^{+}13, KMGCR10, LJA^{+}13, MRRK11, NBM^{+}12, NJT14, NBNMG13, PBK^{+}11, PMMG14, SMLB19, VMGB^{+}10, BGZ^{+}15, DJBB16, HSB^{+}16, JD16, MBBB17, NWD^{+}16a, SMARC^{+}16, SMY^{+}17, TAP19]. Hutchinson [GB11]. Hwange [GDMF13]. Hybrid [RJF^{+}13]. Hybridization [LRMWG16, TGD10, ZYYH10, KLL^{+}16, MH16, QCMB18]. hybrids [JMH15, QSCM^{+}19]. hydrocarbon [KBML15, VTT^{+}17]. Hydroclimatic [BP13]. Hydrological [SAGSG^{+}14, LW18]. hydromorphological [PTJ^{+}19]. Hydropower [PG12]. Hydroprogne [MPH^{+}11]. Hydrothermal [GZD11, VLRB12]. Hyperabundance [TAG^{+}10]. hyperdiverse [RBH^{+}15]. Hypotheses [PAM^{+}10]. Hypothesis [FC10, LLW^{+}14, Rob13, MAG18, TWZ16]. hypothetical [HAF^{+}17]. hypotheticals [RM19].

Iberian [PRR^{+}11, PGLB^{+}12, RCR^{+}12, SGSR^{+}12]. Ibérico [PRR^{+}11, PGLB^{+}12]. ice [LSK^{+}15]. Iconic [JMP^{+}11, QKC^{+}19]. Icónico [JMP^{+}11]. Icons [CG13]. Idea [Mar10]. Identificación [BRCK11, Duf11, MVB^{+}11, RER10, UDG10]. Identificar [HJSF^{+}12, IP11, LH11, MR12]. Identification [BVC^{+}14, BMM^{+}14, MLMP^{+}15, PMS13, BCK^{+}19, CMMP15, MRFvW16, WBPL15]. Identify [AS13, BRCK11, GAFG14, IP11, LMC^{+}14, LH11, MR12, DCP^{+}19, TRPD16, TMK^{+}16, VKM^{+}17]. Identifying [CBWFL15, DRF^{+}18, Duf11, GSSF19, JVI^{+}18, MVB^{+}11, RER10, SLB^{+}15, UDG10, WBF^{+}19, AGFR18, WFSS19]. identities [HTE16]. Ilegal [GSB10, ÖB12]. Illegal [GSB10, LSK^{+}14, NBNMG13, ÖB12, OKCG18, SKH13, BCR^{+}17, CPD^{+}15, DJBB16, KOA^{+}15, LFMB19, MFHT19, MSR19, SMARC^{+}16]. Illustrate [WHB^{+}11]. Illustrated [BHK10]. Ilustrada [BHK10]. Ilustrar [WHB^{+}11]. image [Ano11c]. Imagery [LK14, LSA17]. images [SKLP16]. Immigration [SG10]. Imminent [SW10]. Immune [ZLKW13]. immunity [TMD^{+}17]. Impact [CMRH10, CMG15, DMZ10, JMF12, JBB^{+}14, JBH^{+}16, MGMA^{+}14, MSSM09, NKB10, RBF15, Sch10, SML^{+}13, SMGRJ10, ALA17, Ben15a, GBMW15, GHT^{+}15, HTH^{+}17a, KHB18, MRFvW16, SMV19, SHB^{+}17, SG18, SCHM18, TBM^{+}16, TG15, ZV19]. Impact-Scoring [NKB10]. Impacto [CMRH10, DMZ10, NKB10, MSSM09, SMGRJ10]. Impactos [LSBB10, LGB^{+}10, MCN10]. Impacts [CGE19, Fur19a, HSB^{+}16, LSBB10, Les15, LKB^{+}14, LGB^{+}10, MCN10, SBSF14, BJ16, BTM^{+}15, BGZ^{+}15, DP17, GHM^{+}15, GML^{+}19, GSSF19, Hug19, IRK^{+}18, LGCF19, MRMH17, MAB^{+}15, MGA16, SSRS18, STM^{+}19, SGA^{+}19, TMM^{+}18, WK15]. impediment [FCS19, SAS^{+}19]. impediments [SLB^{+}15]. Impending [CGM10]. Impenetrable [BMGLW12]. Imperative [Kir14]. Imperfect [AZT^{+}14, RBHW18, RNE^{+}19]. imperfectly [LDD19]. imperialism [Bur19]. Imperiled [PF10, BCSM15, HTDM17, HKB^{+}17, LB15]. Imperiled-Species [PF10]. Implement [DC11]. Implementación [DC11, HRC10, OKLP11]. Implementation [BMG^{+}14, LAA^{+}10a, PGBV^{+}13, BPN^{+}16, GW17, KMR^{+}18, LRMWG16, LBB^{+}16, NRD^{+}16, vBSC^{+}19]. Implementing [OKLP11, ENW^{+}18, KHD^{+}15]. Implicaciones [HFK^{+}10, OTKE12, PF10, SBT^{+}11, SHA^{+}10a, AS10, ZZW^{+}10]. Implications [AB13a, AS10, DBK^{+}14, ECES13, HFK^{+}10, LLR^{+}15, MTD16, OTKE12, PF10, RBK^{+}17, RGRMPB^{+}13, SBT^{+}11, SHA^{+}10a, VM14, ZZW^{+}10, BW16, DMR^{+}17, DP17, FSAM15, FDBA17, FMB^{+}18, Jør15, MCH^{+}16, MVVS15, NBA^{+}18, RBH^{+}15]. Implicit [CCS^{+}18]. Importadas [SKCS11]. Importance [BPA^{+}13, BGL^{+}14, CPC^{+}10, HAT^{+}15, Lin11, NWD^{+}16b, RSN19, RJA^{+}16, WL17a, AB15, BBF19, CMM17, EKRW17, TND16, WTT^{+}17]. Importancia [CPC^{+}10]. Important [PFA^{+}14b, GSSF19, MBC^{+}16]. Importante [RER10]. Imported [SKCS11]. Improve [BG10, OCP^{+}14, Ben16, GFJC19, JKJ^{+}19, LMJK^{+}19, NBR^{+}15, OSM^{+}17, RHWS18]. improved [DUS^{+}16, RLL15]. Improving [EWCG15, HSRC14, LKB^{+}10, MRFvW16, RLL15, SAGSG^{+}14, VSA^{+}15, vBSC^{+}19, JGS^{+}16, MAB^{+}15, PBB^{+}17]. Inadaptiva [HBE^{+}10]. Inadvertent [Wil12a, Wil12b]. Inbred [LJA^{+}13]. Inbreeding [CWS^{+}14, GLNJ10, Jam11, JMW^{+}11, RKH^{+}14, RLGG^{+}12]. Incendios [CMJ^{+}11]. Incentive [SHCD11]. Incentives [RHT14, TW14, AJ18, LH19, NM18]. Incentivizing [BEA^{+}13, GD15]. Incentivos [SHCD11]. Incertidumbre [AG10]. incidental [BGP16, MAJ16]. Incierto [SBN^{+}10]. Inciertos [CWH^{+}10, SSB^{+}12]. Incilius [SMM16]. include [ANE^{+}16]. including [WL17a]. inclusion [GPB^{+}15]. Income [CAB^{+}10, Mcc10]. Incorporación [MJP^{+}11, SSK^{+}11, TCDR12]. incorporate [ACB^{+}15]. Incorporating [BM14, BWE^{+}13, CCB^{+}19a, CPH^{+}15, GBMW15, MAB^{+}13, MJP^{+}11, OCP^{+}14, PBB^{+}15, SMTW13, TPL18, TCDR12, MVVS15, WTT^{+}17]. Incorporation [SSK^{+}11]. Increase [BLC13, WES^{+}13, Dar15]. Increased [HB10, CBF^{+}19]. Increases [BBWK^{+}14, HSR^{+}13, HLC13]. Increasing [FBB^{+}14, JMF12]. Incrementar [CLV^{+}10, SPSS10, VMS^{+}12]. Incremento [HRSK12, HB10]. independent [GRL15]. Index [BMA^{+}14, JHV13, SLGB15, JRWA15]. India [AOLK11, GBG^{+}10, NM18, Pri19, SMS14]. Indian [EMJ14, MSY^{+}15, PG12]. Indicador [SA11]. Indicadores [JCA^{+}11, PPD^{+}12]. indicator [SSE^{+}17]. Indicators [CC14, JCA^{+}11, NA13, PMM^{+}14, PPD^{+}12, HLS^{+}18, RNM^{+}18]. Indices [MMK^{+}14, BMC^{+}18]. Indígenas [LGB^{+}10]. Indigenous [AL14, LGB^{+}10, Lyn13, PDM11, GH18, KB17, LZR^{+}18, SMY^{+}17, SGA^{+}19]. Indirect [WNY13]. Individual [LGK^{+}11, PMS^{+}10, WJ10, CMM17, DCP^{+}19]. Individual- [WJ10]. individual-based [DCP^{+}19]. individual-condition [CMM17]. Individuales [LGK^{+}11]. Individuals [BW13, LMM^{+}10, CHWH18, RNE^{+}19, WC19]. Individuos [LMM^{+}10]. Indonesia [CSB^{+}10, CWS16, JC18, LSK^{+}14, MAA^{+}10, MGW^{+}15, Nat18, SHA^{+}10a, WIM^{+}10]. Indonesian [HTH^{+}17a]. Induced [AASK12, EVF13, EPB^{+}14, GKK^{+}11, HJMM14, RTL^{+}12, SGS^{+}10, PAMA17, RCSK15, RHR^{+}17]. Inducida [RTL^{+}12]. Inducidas [SGS^{+}10]. Inducidos [GKK^{+}11]. Industrial [MBL^{+}14]. Industry [JVT^{+}13]. Inequality [TA10]. Inequidad [TA10]. Inertia [ZPJ19]. Infección [OSFFB^{+}11]. Infecciosa [FCC^{+}11]. Infectados [LMM^{+}10]. Infected [LMM^{+}10, SSP^{+}15]. Infection [BVC^{+}14, OSFFB^{+}11]. Infections [SMH^{+}14]. Infectious [FCC^{+}11]. Infer [EBL^{+}14]. Inference [SB14, ALW^{+}19]. Inferences [BRV10, CCBP14]. Inferencias [BRV10]. Inferred [WNY13]. Inferring [FB12, SA15]. Infiriendo [FB12]. Influence [AB13a, BA13, DWM^{+}10, FLP^{+}11, LS15, RC12, SRS11, BTB^{+}18]. Influencia [DWM^{+}10, FLP^{+}11, RC12, SRS11]. Influencing [FKM^{+}11, BGP16, CC18, HFBG18]. Influyen [FKM^{+}11]. Inform [Ber12, HMIE14, MO11, RL13, RBC^{+}11, DBV^{+}18, GTVC16, GW17, LDD19, SBR^{+}19, TPL18, TMG^{+}18, THD^{+}18, VJE^{+}15]. Información [LCG10, MM10, MR12, RMGCR10]. Informal [PAM^{+}10]. Informales [PAM^{+}10]. Informar [Ber12, RBC^{+}11]. Information [Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano16j, Ano16k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano18n, Ano18o, Ano18p, Ano18q, Ano18j, Ano18k, Ano18l, Ano18m, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, LCG10, MM10, MR12, BW19, CBF^{+}19, DCP^{+}19, SMAG17, WML^{+}19]. informative [PS16]. informed [QSCM^{+}19]. Infrastructure [QM11, Ven18, MGA16, RTGR17, SAS^{+}19]. infrastructures [CVA15]. Ingestion [SHWT14, SR17]. Ingreso [CAB^{+}10, Mcc10]. Inhabited [OBB^{+}11]. inhabiting [PWKB19]. Initiative [Hug19, MK18]. Inminente [CGM10]. innovation [GD15]. innovator [LBD^{+}19]. Insect [LDC^{+}13, Whe11, HCT^{+}17]. Insectívoras [IUM10]. Insectivorous [IUM10, BHF^{+}19]. Insectos [HBE^{+}10]. Insects [Csó17, HBE^{+}10, LKB^{+}14, Sha15b]. Inserting [RB13]. Insight [Aro11, SG10]. Insights [Cha15, GDMF13, HSRC14, MAM^{+}15]. inspire [SSC^{+}19]. institutional [HSV^{+}19]. Institutions [Dre11]. Instrumentos [PCTW10]. Instruments [GFG^{+}12, PCTW10]. insurance [RBHW18]. Intact [SSBN12, KHGF^{+}17, RLB^{+}17]. Intactos [SSBN12]. Intake [VCDF13]. Integración [ACE^{+}10, SH11b]. Integrado [MBJWC10]. Integrados [BMGLW12]. Integral [PA12]. integrate [Wal17]. Integrated [BMGLW12, CS14, FRY^{+}17, GDMF13, MBJWC10, PA12, LSV^{+}19, TIB^{+}16].

Integrates [SM14]. Integrating [ACE^{+}10, LRA^{+}14, RL13, WHE^{+}17, WKI^{+}14, CNvL^{+}15]. Integration [MBBH16, SH11b, BTBC^{+}16, WML^{+}19, WBF^{+}19]. integrative [DDv18]. Integridad [CGM10]. Integrity [CRNR12, HNDE13, TW11, CHG^{+}17]. Integrity. [CGM10]. intelligence [VB19]. Intensidad [BRKP12, FP10, SPPN^{+}12]. intensification [BMA^{+}18]. Intensity [BRKP12, SPPN^{+}12]. Intensively [EWN^{+}13]. intentional [MBT^{+}17b]. Interacción [WGB^{+}11b, HRKMG12]. Interacciones [FC10]. Interacting [CCM12, HSBE12, RSW11, BBF19]. Interaction [HRKMG12, WGB^{+}11b]. Interactions [BBB14, FC10, GFGR12, MBH14, MLLS14, PY13, RM14b, AK16, AWT^{+}15, CLM^{+}15, GZR18, GFdM^{+}19, RBF15]. Interactive [WBM13, GSSF19]. Interactivos [RSW11]. Interactuantes [HSBE12]. Interannual [SM11]. Interanual [SM11]. interconnected [Ben15c]. Interdisciplinaria [SCL12]. Interdisciplinary [BC14b, Fis12, SCL12, SH13, PBB^{+}17, WZWF18]. Interesados [FTSSW12]. Intereses [TM10]. Interespecífica [ZYYH10]. Interest [Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano11p, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano14h, Ano14i, Ano14j, Ano14k, Ano14l, Ano15h, Ano15i, Ano15j, Ano15k, Ano16x, Ano16v, Ano16w, Ano17x, Ano17y, Ano17z, Ano17-27, Ano17-28, Ano18r, Ano18s, Ano18t, Ano18u, Ano18v, Ano18w, Ano19w, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Red10b, Ano16y]. Interests [Rob12, TM10]. Intergovernmental [RGMR12]. Internalizing [Sik12]. International [Bai10, DFJ^{+}12, Des19, DFR^{+}17, Har15, Joh10b, FFS^{+}18]. internet [DKL^{+}15, SASO13]. Interpatch [MLS12]. interpretation [Wil17]. Interpreting [RDS17]. Interrumpida [CBSZ10]. Interrupted [CBSZ10]. Intersection [BB11a]. Interspecies [MH16]. Interspecific [MLLS14, ZYYH10, KLL^{+}16]. Intervención [CMRH10, SMGRJ10]. Intervenciones [WFM^{+}10]. Intervention [HP11, Par10, SMGRJ10]. Interventions [BGL^{+}14, SHG^{+}14, WFM^{+}10, CPM^{+}18, HBTMG15, ISM^{+}18, JKJ^{+}19, TAM^{+}19, TAP19, WHEN^{+}16]. Intrinsic [Whi13, VBN15]. Introducción [KTN12, WR12, WD10]. Introduced [HPL14, FMB^{+}18]. Introducidas [Chi10, NKB10]. Introducing [KTN12]. Introduction [AHS12, BHA15, BBGL10, BS15, Bur14, Cin13, DSW16, Din11, ESK19, LPCA16, PGC^{+}13b, Pin17, Ray14, RF11, Wap13, WR12]. Introductions [Ehr14, MRMV14, WFSS19]. introgression [HM16]. Invaded [IP11]. invader [BJ16]. Invadidos [IP11]. Invasion [LO10, Pra10, FJP^{+}11, LO10, Pra10, WT10]. Invasiones [LB11]. Invasions [Arm16, FZ18, FJP^{+}11, LB11, Hul15, WCR^{+}15]. Invasive [BHK10, BLC13, CLT^{+}13, Chi10, EG12, FTSSW12, GG13, Hal19, JMP^{+}11, LCFS14, LWLC12, MGCL14, MBGS11, OBB^{+}11, PSAZ11, Par10, RSPD13, SS10a, DOC^{+}15, DSV17, EAPB15, GRL15, GMG^{+}19, HHW^{+}16, HHKK15, JHAH18, ME16, MDD^{+}18, MSL19, NJBM15, NW15, RJA^{+}16, UMK^{+}16, WHE^{+}17]. invasive-species [JHAH18]. Invasora [EG12]. Invasoras [BHK10, Chi10, JMP^{+}11, PSAZ11]. Invasores [FTSSW12, MBGS11, OBB^{+}11, SS10a]. Inventing [HP14]. Inventory [CKPS10]. Invernante [EG12]. Invernantes [FCC^{+}11]. Inverse [KMTK13]. Inversiones [CLV^{+}10]. Invertebrados [LGH^{+}12, MVB^{+}11, ODTS10]. invertebrate [RBH^{+}15]. Invertebrates [LGH^{+}12, MVB^{+}11, ODTS10, RSPD13]. Invest [CKPS10]. Investigación [EMS^{+}11, RBC^{+}11, SCL12, PSAZ11]. Investigando [BBBG10, FCC^{+}11]. Investigating [BBBG10, FCC^{+}11, HH19b, MFHT19]. Investing [GBP^{+}13]. Investment [SSSD14]. Investments [CLV^{+}10, SFAA^{+}18]. Invierno [LHDC12]. Invisible [WLRR18, WS14]. Involucra [MBJWC10]. Involucrando [DC11]. Involuntaria [Wil12a]. Involving [MBJWC10]. IPBES [LBD^{+}15, PMD^{+}13, SG19]. irreplaceability [MBC^{+}16]. irreplaceable [RHPF15]. ISBN [Ano16x, Ano19w, Bra15, Fie15]. Isla [BV10, SHA^{+}10a]. Island [Ano19w, BV10, Bor19, EMJ14, Fer17, Laz11, NS18, PMM^{+}15, SHA^{+}10a, SNH^{+}14, VCDF13, HAT^{+}15, RJA^{+}16, WA18, Boy10, TQD^{+}18, WMFB15]. Islands [Bor18, BJMT14, DFW^{+}10, OBB^{+}11, SWL^{+}10, WMW^{+}14, AHT^{+}15, DOC^{+}15, DSV17, HHW^{+}16, WFSS19, AS10, LZD^{+}13, LL14, UMK^{+}16]. Islas [OBB^{+}11, SWL^{+}10, AS10, Boy10]. Isolated [ZZW^{+}10, GGB^{+}15, WMFB15]. Isolation [DBK^{+}14, SPPN^{+}12, AC16]. Isotopes [BFB^{+}19]. Issue [Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano16j, Ano16k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano18n, Ano18o, Ano18p, Ano18q, Ano18j, Ano18k, Ano18l, Ano18m, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q]. Issues [RSWM14, CNK^{+}17, HH19a, PSA^{+}17, SG18]. Iterópara [ANTBG12]. Iteroparous [ANTBG12]. IUCN [ABB^{+}18, FW17, KMH^{+}14, RRCB^{+}11, RCARR18, SBR^{+}19, SMM16, TBC18]. Ives [MBT^{+}17b]. Ivory [GCC^{+}12, SSB^{+}12, Ben15a, BHO^{+}18, HGM19, SBS^{+}15]. Ivory-Billed [GCC^{+}12, SSB^{+}12].

J [Ano16x, Fie15]. Jacquet [NWD^{+}16b]. Jaguar [RRV10]. Jaguars [BMM^{+}14, AJ18]. Janssen [Nat18]. January [Ano11a, Ano18a, Ano19a]. January-December [Ano18a, Ano19a]. Japan [KK18]. Jardines [HFL^{+}11]. Jargon [WHB^{+}14]. Jays [CFBL10]. Jerárquicos [CJDZ10]. Joaquin [MWC11, MWC11]. job [LGA17]. John [Ano19w, RTL^{+}12]. Journal [CB10]. Jr. [TWZ16]. Judging [FR16]. judgment [HKB^{+}17, WCR^{+}15]. judgments [Die16]. Judicious [RMRO14]. jure [AG17]. Just [Hal16]. justice [LMCS17]. Juvenile [WBM13, OSM^{+}17].

Karoo [ERC^{+}10, PCTW10, PRE^{+}13]. Keeping [MNW^{+}14]. Kelp [ETB10].

Kemp [Ano19w]. Kendal [GZR18]. Keniano [Mcc10]. Kenya [HDN^{+}14b, KWW^{+}14, SCA17]. Kenyan [Mcc10]. Key [BB15, SGS^{+}10, WHB^{+}14, DSG^{+}17, PA18, WBF^{+}19]. Kibale [LW10]. Killing [KE10, GTC^{+}16]. Kilombero [NJT14]. Kingdom [DOC^{+}15, HGM19]. Kleiman [Die10b]. Klein [PCTW10]. Know [AA12, GHKK10, Mih19, PI11, Gru17]. Knowledge [BC14b, CMS^{+}13, CD13, GBMRG^{+}10, JGH^{+}11, LSDK11, MBF^{+}12, MAGA10, NRD^{+}16, Tru18b, BC15, CBF^{+}19, HBTB15, NYC17, WD17]. Known [Bra15]. Kong [FFS^{+}18]. Korea [LKB^{+}14]. Kovach [MH16]. Kuboń [NJK^{+}19].

L [Ano16x, Ano19w]. Lack [AGLRV14]. Lagartijas [HSAB10]. Lagged [HLC13]. Lago [Hic10]. lags [AOPM^{+}19]. lake [CEdK18, Hic10, WCR^{+}15]. lakes [PRB18, NJBM15]. Lambers [Fie15]. Lambert [BCA^{+}19]. Land [AH12, AH11, BB10a, BSF^{+}14, BRKP12, CKR^{+}13, GMBM13, GJB14, HRC10, JMP^{+}11, KO12, KCDC10, KG10, LK13, Lyn13, MSMCB18, McP11, Mcl11, MC11a, MC11b, PRE^{+}13, RHT14, RPA^{+}12, SPPN^{+}12, SHO^{+}10, WIM^{+}10, WCCT18, ZLKW13, ARDG17, BBH17, CPR^{+}16, CC17, CC18, FBWIO19, FGG^{+}19, FFHA19, GG19, JBB^{+}15, NWAB19, NWW^{+}17, PCF^{+}19, RBH^{+}15, RCSK15, RK17, RHR^{+}17, SBR^{+}19, STM^{+}19, SJZ19, TND16, TRPD16, UMK^{+}16]. land-based [RHR^{+}17]. Land-Conservation [AH11, CC17]. Land-Cover [RPA^{+}12, SBR^{+}19]. Land-Sparing [CKR^{+}13]. Land-Use [BRKP12, JMP^{+}11, KG10, Lyn13, MC11a, RHT14, SPPN^{+}12, SHO^{+}10, WIM^{+}10, MSMCB18, WCCT18, ARDG17, JBB^{+}15, NWW^{+}17, PCF^{+}19, RCSK15]. Landbridge [SWL^{+}10]. Landed [WB14]. landforms [RHPF15]. Landholder [MC11b, TTEM14]. Landholder-Based [MC11b]. landing [JRWA15]. Landmark [WT13]. Landowner [ERBM17, SHCD11]. Landowners [LA13, RRV10, NSBJ17a, NSBJ17b, NWAB19]. lands [EKRW17]. Landscape [ACW18, BHM^{+}14, EBL^{+}14, EVF13, HD14, JLDJ12, MRRK11, MC14, NBSM12, SD11, SMS14, SAS^{+}19, VSM^{+}11, BP16, BGD^{+}15, BPN^{+}16, CTBR19, FFHA19, ITG^{+}16, KMT^{+}17, ML15, PBR^{+}16, RHPF15, THD^{+}18, VESHR18]. landscape-scale [BPN^{+}16, ITG^{+}16, KMT^{+}17, PBR^{+}16]. Landscapes [AME^{+}13, AOLK11, BS10, CKR^{+}13, GBG^{+}10, GZA10, GZR12, LH11, Len17, PY13, Qui14, WW10a, ZFW^{+}11, BLM^{+}16, CNBW16, DMF^{+}18, ERW18, FGG^{+}19, KHGF^{+}17, NWW^{+}17, OQGR^{+}15, SHSP19]. Lane [NS18]. Langosta [SHC^{+}11]. Langostino [TTSZH10]. Langostinos [LO10]. Language [CND10, GPK^{+}10]. Large [ALL^{+}12, Ber12, BBM13, Ber13, BM14, HMB^{+}13, KO12, LK14, Len17, MBH14, MTM12, PVP^{+}13, RMM^{+}13, SHPR13, BIP^{+}15, BMC^{+}18, CSR^{+}19, DPT^{+}18, HSV^{+}19, IRK^{+}18, KRL^{+}16, ML15, RCBT18, RV15, SSD^{+}18, WFMW^{+}18, vECD^{+}18, vWWCF15]. large-dam [IRK^{+}18]. Large-Scale [Ber13, PVP^{+}13, SHPR13, BMC^{+}18, ML15, WFMW^{+}18]. Largest [HMIE14, BHO^{+}18, GLJ^{+}16]. Largo [Boy10, CW10, EDB12, GGG10, MFRH12, TTSZH10]. Larvae [BFHP10, BAG^{+}11]. Larvas [BFHP10, BAG^{+}11]. Last [Ano19w, SSRS18]. Latent [LO10]. lateral [LW18]. laundering [Nat18]. Law [Des19, NBM^{+}12, CQC^{+}19, DUS^{+}16, WGLL17]. Layperson [Poz17]. Lead [KGG^{+}14]. Leader [BLG^{+}11]. Leadership [Tor11]. leakage [BTM^{+}15]. Learned [PCTW10, PMD^{+}13]. Learning [ABD^{+}17, BPY^{+}16, CCBF12, Jen14, MF14, BCK^{+}19, CRMK18, CAB^{+}18, JGS^{+}16, MFHT19, RCBT18, VLSOS16]. Least [PMS13, SG18, VNAT19]. Least-Cost [PMS13, SG18, VNAT19]. Leatherback [GSDE14]. Lebanon [MA10]. Lecciones [OBBP10, PCTW10]. Legacies [CG13]. Legal [Ben15a, MLMP^{+}15]. legged [ALW^{+}19]. Lehtomäki [MBW^{+}19]. Leks [BBP12]. Lemhi [WBM13]. lemurs [BHB^{+}19, GV16]. length [FW17]. Lenguaje [CND10, GPK^{+}10]. Leñosas [VJMH12]. Lens [Con11]. Lenta [PMK^{+}10]. León [Cla15]. Leones [PBK^{+}11]. Leopard [BAGM13, LWY^{+}14, PBK^{+}11, MSY^{+}15, WW16]. Leopardos [AOLK11, PBK^{+}11]. Leopards [AOLK11, JVT^{+}13, SBB^{+}14]. Lessons [AH11, FS15, MDSF19, OBBP10, PCTW10, PS13, SSD^{+}18, AAS^{+}17, HWMC15]. Let [Whe11]. Letal [HS10]. Lethal [HS10, RJF^{+}13, RJA^{+}16]. Level [BBWK^{+}14, BR11a, HRSK12, IFP14, JHV13, NBSM12, NRRW17, RBC^{+}11, TSA^{+}14, WJ10, CCN16, HN17, SL15]. Levels [CGCL11, JLDJ12, WSHB11, TMD^{+}17]. Leverage [LA13]. Leveraging [SPD18]. Líbano [MA10]. Libélulas [SS10a, SHLK^{+}10]. Liberación [BTBDS12]. Libros [CB10]. license [GZR18, KF18a, KF18b]. Licensing [SJH13]. Lichen [MTN13]. Lichens [PTN10]. Lidar [SAC12]. Life [CS13, DS19, Fer17, Fie15, GLNJ10, PMK^{+}10, WGJ13, AOPM^{+}19, AK16, Cha15, GRT^{+}18, RK17, TLPP17]. Life-History [GLNJ10, PMK^{+}10, DS19, GRT^{+}18, TLPP17]. Lifetime [CB13]. Light [GKB10]. lights [RHR^{+}17]. likely [LB15]. Limitaciones [EMS^{+}11, OBBP10]. Limitado [WGR^{+}10]. Limitations [EMS^{+}11, IBA16, OBBP10, RL11, SL16, SFKC16]. Limited [BCLMM19, WGR^{+}10, SMAG17]. Límites [FJP^{+}11]. Limítrofes [RPA^{+}12]. Limits [Alr15, TJJB18]. Lince [PRR^{+}11, PGLB^{+}12]. Line [LGK^{+}11]. Lineages [LZD^{+}13]. Linear [CB13, LGC^{+}13, RTGR17]. Líneas [BV10, NBW^{+}12]. Lines [BAPP11, DHD14, TSH^{+}14]. lining [PFN15a]. Link [WWW^{+}12]. Linkage [CMB12]. Linkages [TQD^{+}18]. Linked [AWU^{+}11]. Linking [HCSM15, LSDK11, MMK^{+}14, NA13, RHPF15, PNC^{+}15]. links [DJBB16, FMV^{+}19]. Lion [BL17, Dys15, HDN^{+}14b, PBK^{+}11, CBM^{+}18, DHF16, EG17]. Lions [Med10]. Líquenes [PTN10]. List [ABB^{+}18, JHV13, KMH^{+}14, RRCB^{+}11, RCARR18, SLGB15, SBR^{+}19, SMM16, TBC18, COZS16]. Lista [PTN10, RRCB^{+}11]. Listas [PAM^{+}10, ZBM^{+}10]. Listed [PTN10, ZCCN13, TG15, RS18]. Listing [RTT^{+}13, ZBM^{+}10]. Lists [PAM^{+}10]. Literacy [CND10, PSD^{+}14]. Literatura [OB11b]. Literature [Dol10, OB11b, Abr15, AGS19, HWMC15].

Little [FBF^{+}14, ERC^{+}10, PRE^{+}13]. Live [OK10, Har15]. Lived [BSF^{+}14, LGH^{+}12, PMK^{+}10, SBJG13, DSK^{+}19]. Livelihoods [DVW^{+}13, VCDF13, CMG15, WZWF18, WHR^{+}16]. Lives [Cox10]. Livestock [EDB12, GSW10, KO12, KMT^{+}17, vECD^{+}18]. Living [GGG10, HBM^{+}12, Ord11, Pel15]. Lizards [BTMF13, HSAB10]. Llegar [BBBG10]. Lluviosos [MLS12]. Lobo [CMB12]. Lobos [ACL^{+}12, CPC^{+}10]. Lobster [SHC^{+}11]. Local [BRV10, BFHP10, CGHM13, DVW^{+}13, EMS^{+}11, GGG10, HSRC14, JVT^{+}13, MLA^{+}14, MAGA10, NGL^{+}14, RHT14, RGRMPB^{+}13, SMS14, SM11, SHO^{+}10, Sto12, SHLK^{+}10, SKSK14, TBY^{+}10, WFM^{+}10, BIF^{+}18, CMG15, EGB18, KNB^{+}17, KB17, NWD^{+}16b, SBU^{+}17, STM^{+}19, TMG^{+}18, WD17]. Local-Scale [NGL^{+}14]. Locales [BRV10, BFHP10, EMS^{+}11, GGG10]. location [VMO^{+}18]. Logged [AS13, EWN^{+}13]. Logging [AEO^{+}13, AS13, CRKM^{+}14, CMRH10, CMJ^{+}11, DMZ10, LSK^{+}14, BGZ^{+}15, KHGF^{+}17]. Logics [WS14]. Logro [SKC^{+}12]. Lolobata [SHA^{+}10a]. Lombrices [LB11]. Long [BSF^{+}14, BHF^{+}19, CW10, FZ18, GGG10, HMIE14, HAP^{+}13, JPT13, LBLB14, LGH^{+}12, MKR^{+}15, PRB18, PMK^{+}10, RMKM14, San14, SBJG13, TTSZH10, WC19, BBM18, CWS16, CMM17, DSK^{+}19, FWH19, ML15, SBU^{+}17, SA15, VMO^{+}18, WFMW^{+}18, OK10]. Long-Distance [HMIE14, PRB18, SLB^{+}15]. Long-Lived [BSF^{+}14, LGH^{+}12, PMK^{+}10, SBJG13, DSK^{+}19]. Long-Range [RMKM14]. Long-Term [CW10, FZ18, GGG10, HAP^{+}13, JPT13, TTSZH10, BHF^{+}19, MKR^{+}15, WC19, BBM18, CWS16, CMM17, FWH19, ML15, SBU^{+}17, SA15, VMO^{+}18, WFMW^{+}18]. Longcore [LDB^{+}10]. Longevo [PMK^{+}10]. Longevos [LGH^{+}12]. Longitudinal [TNTS13, Agg15]. Longleaf [SKCS11]. Longline [WB14, MDSF19]. Look [Cop12, ODAL17, DPJM18]. Looking [AGAB11]. Loop [MRMV14]. López [AR10]. López-Hoffman [AR10]. Loro [AS10]. Loss [AC15, BGL^{+}14, EMJ14, GHB^{+}13, GZR12, HCP^{+}12, HRBR11, Jam11, KMTK13, PGC^{+}13a, RBD^{+}13, RTL^{+}12, SBSF14, VJMH12, ZBD11, BBM18, BRA^{+}19, DMF^{+}18, GBBMG19, JBB^{+}15, LSK^{+}15, LW18, OQGR^{+}15, RV15, ROQR^{+}16, SWSM19, SSW19, SAS^{+}19, TGBB^{+}15]. Losses [KG10, KPL^{+}12, WS14]. Lost [Ano19w, WWW^{+}12, WHG^{+}15, Ano16x]. Love [Hal19]. Low [BBBG10, BMA^{+}14, CBS11, SHLK^{+}10, DUS^{+}16, SCHM18]. Low- [SHLK^{+}10]. Low-Cost [CBS11]. low-density [DUS^{+}16]. Lowland [SMS14, FGG^{+}19]. Lowlands [HRC10]. Lynn [TWZ16]. Lynx [PRR^{+}11, PGLB^{+}12, RCR^{+}12, SGSR^{+}12].

Maasai [DHF16]. Maasailand [Dys15, HDN^{+}14b]. machine [CAB^{+}18, MFHT19, RCBT18]. Machos [CW10, HFK^{+}10]. Macroalgae [PB11]. Macroalgal [DBK^{+}14, WGF^{+}12]. Macroalgas [PB11]. Macroecology [Ano16x]. Macrofungi [MFKMP10]. Macrohongos [MFKMP10]. Madagascar [BHB^{+}19, BDH^{+}12, GBB^{+}14, Nut15, SMGRJ10].

Madera [CMJ^{+}11]. Maderables [GT10]. Maderera [DMZ10]. Mahogany [GT10]. Maine [SHC^{+}11]. Mainland [MMCUR^{+}12, RSPD13]. mainstream [Wal17]. Mainstreaming [BRK^{+}17, HSV^{+}19]. maintaining [BVAC17]. Maintenance [BEE^{+}11]. Major [BMA^{+}18]. makes [HDH^{+}15]. Making [HWMC15, HBM^{+}12, KBC^{+}18, LWLC12, MAA^{+}10, MR12, NS12, PFN19, PBB^{+}11, DBB^{+}16, KMR^{+}18, LBI^{+}18, MBBH16, TMG^{+}18, VB19, YNR19]. Mal [SA11]. Maladaptive [HBE^{+}10]. Malaria [LZD^{+}13, SMGR17]. Male [CW10, HFK^{+}10]. Malpai [SK10]. Mamífero [SS12]. Mamíferos [Ber12, ESCF12, FB12, FP10, ML10, NKB10, OTKE12, OBB^{+}11]. Mammal [ESCF12, SS12, Tru18a, AHT^{+}15, BIP^{+}15, BGD^{+}15, CQC^{+}19, GML^{+}19, LSK^{+}15, MBBB17, MNW^{+}17, RV15, SLB^{+}15]. Mammalian [BC14a, FP10, KKK14, OTKE12, CCS^{+}16]. Mammals [BBWK^{+}14, Ber12, BBM13, BTMF13, FB12, KO12, ML10, NKB10, OBB^{+}11, RLGG^{+}12, SGE13, TSH^{+}14, VKKA13, AP17, Alr15, BGZ^{+}15, DRF^{+}18, HSV^{+}17, LFM17, PBR^{+}16, PS15, TYA16, vWWCF15]. Manage [DC11, MBH14, CS19]. Managed [BSF^{+}14, MTN13, OKLP11, PBH^{+}10a, PBH^{+}10b, CSR^{+}19]. Management [APL^{+}13, BSF^{+}14, CMRH10, CRD^{+}14, CPF13, CD13, CMG^{+}13, CCBF12, DFJ^{+}12, DBP^{+}10, FTSSW12, GLB^{+}13, GPK^{+}10, HM11, HTH^{+}10, HNS^{+}11, HM12, HHA^{+}11, HMB^{+}13, HMM10, HNDE13, IFP14, KRG13, KO12, LL14, MB10, MFN^{+}11, Mat13, MJP^{+}11, MM10, MR12, NMMP14, NS12, NA13, PSAZ11, RM11, RBC^{+}11, Rud11b, SMTW13, SBBP11, SCL12, SAGSG^{+}14, TCP13, TTEM14, WGJ13, WGB^{+}11b, WLCS14, AdBR15, AG17, AWT^{+}15, BDKS15, Ben16, BARB17, CD16, CNvL^{+}15, CC17, CNK^{+}17, DBB^{+}16, EAPB15, FR16, GBB^{+}19a, GHM^{+}15, GMG^{+}19, GFJC19, GH18, HM16, ISM^{+}18, JHAH18, JGS^{+}16, KHB18, LB15, LZR^{+}18, MGMC15, MK18, MWL^{+}15, NYC17, NBR^{+}15, OJAV15, PMM^{+}15, RHWS18, RBK^{+}17, SDT17, VGAB^{+}19, WDS15, WA18, WCCT18, WGLL17, WK15]. Managers [MAA^{+}10, MV18]. Managing [BS12, FCC^{+}11, GZNC13, GAG13, HNO^{+}18, LGK^{+}11, Mat15, MVB^{+}11, Qui14, TW11, vECD^{+}18, CBPM19, TG15, ZP18]. Manejadores [MAA^{+}10]. Manejados [PBH^{+}10a]. Manejando [FCC^{+}11]. Manejo [CMRH10, CCBF12, DC11, DBP^{+}10, FTSSW12, GPK^{+}10, GC10, HNS^{+}11, HM12, HHA^{+}11, HMM10, LGK^{+}11, MB10, MFN^{+}11, MJP^{+}11, MVB^{+}11, MM10, MR12, PSAZ11, RM11, Rud11b, SCL12, WGB^{+}11b, HM11]. Manera [LGH^{+}12]. Manfredo [IF17]. mangrove [FTB^{+}16, WML^{+}19]. Mantenimiento [BEE^{+}11]. Manu [CLRV11]. manuscript [Ano12g]. Manuscripts [OSC10]. Manuscritos [OSC10]. Many [TBB^{+}14]. Map [Bor18, SM14]. Mapas [BSBM11]. Mapeo [CMB12, KCDC10]. Mappes [BC18]. Mapping [BKKM14, CMB12, CHN^{+}15, FFM^{+}14, GJB14, KCDC10, Lyn13, PRE^{+}13, TMG^{+}18, VLSOS16]. Maps [BSBM11, HW19]. Mar [BPW12, HRSK12, LAA^{+}10b]. March [Ano16b]. Marco [BGAS10, HHA^{+}11, MFN^{+}11, NFF^{+}10, SD11]. Marfil [SSB^{+}12]. Mariculture [TSA^{+}14]. Marina [CGCL11, HSBE12]. Marinas [BBBG10, BJR^{+}11, CKKL10, DWCW10, ETB10, GC10, HNS^{+}11, HRSK12, HFK^{+}10, HRKMG12, LGH^{+}12, LGK^{+}11, MCN10, MO11, SSK^{+}11, WL12, WRAW10]. Marine [BBBG10, BBWK^{+}14, BHSA12, BJR^{+}11, BGG13, ESCF12, FCGJH11, Fur18, GFG^{+}12, GFGR12, HNS^{+}11, HCP^{+}12, HJM11, HSBE12, LW11, LGH^{+}12, LGK^{+}11, MJ11, MJP^{+}11, MRMV14, OCP^{+}14, PFA^{+}14b, RMM^{+}13, Ric11, SRL^{+}13, Tru18a, Vit17, WL12, ANE^{+}16, AGFR18, Bai16, BCD^{+}16, Ben15c, BMB^{+}19, BHAR^{+}18, BGP16, CESS^{+}18, CBM16, CE15, DKG^{+}15, DSG^{+}17, GHM^{+}15, GMV^{+}18, GSSF19, GPB^{+}15, HNO^{+}18, HLE19, KJWP19, LSK^{+}15, LFM17, MVVS15, SLA^{+}19, SL15, SG18, SKPP15, TMM^{+}18, VJE^{+}15, WWdS^{+}16, WMFB15, WHG^{+}15, WK15, CKKL10, CGCL11, DBK^{+}14, GC10, HMIE14, MCN10, MO11, SSK^{+}11, WRAW10, WJ13, WGF^{+}12]. marine-protected-area [KJWP19]. Marino [LW11, MJP^{+}11]. Marinos [ESCF12, FCGJH11, LAA^{+}10b]. Mariposas [MTM12, MBJWC10]. Mariscos [HJSF^{+}12]. Marked [BAPP11, FHBK18]. Market [AWKR^{+}11, BM10, BBM13, GBMRG^{+}10, LW10, FFS^{+}18, FDBA16, HAF^{+}17, LGA17]. Market-Based [LW10, FDBA16]. marketing [SMV19]. Markets [BCGP13, CGHM13, FFM^{+}14, Sto12, FOF^{+}15]. Markov [DDPC12]. Mars [Poo13]. Marsh [CBA13, MGCL14, MAWF12]. marshes [CWH^{+}17]. marsupial [KP19]. martens [PMM^{+}15]. Más [JCA^{+}11, ZBM^{+}10, PMK^{+}10, ZZW^{+}10]. Masa [MJA^{+}11]. Mass [MJA^{+}11, HSV^{+}17]. Matando [KE10]. Matanza [HS10]. Matches [SFAA^{+}18]. Matching [WB10]. Matlack [SVN^{+}15]. Matorral [MQAMWM12]. Matrix [CEM^{+}13, KG10, MLS12, MPB10, PZM14, SMTW13]. Matrix-Calibrated [KG10, PZM14]. Matriz [KG10, MLS12, MPB10]. Matters [RER10, HGGH15]. Maximize [TCP13, SCA17]. Maximizing [GBD^{+}15]. May [RHM10]. Mayores [Ber12]. McClenachan [Bai16]. McCoy [SBJG14]. McRae [LBD^{+}19]. Mead [Hic10]. meadows [CFB^{+}17]. Meaningful [EG13, MAA^{+}10]. Means [HF10, Sch11, RIBT18]. Measurable [Boo14]. Measure [FP10, MMPF10, CSR^{+}19, LRL^{+}15, SWSM19]. Measured [CC14, Rog11]. Measures [Bee13, LH11, RHWS18, CMM17, TQD^{+}18]. Measuring [GSB10, HTH^{+}17a, SPPN^{+}12]. Meat [Ald12]. Mechanism [DVW^{+}13]. Mechanisms [Fer17, JGS^{+}16]. Mechanistic [GMNO15]. Media [BAGM13, MGG^{+}13, HLHR16, KMLC19, MFHT19, ZAB16]. Media-Content [BAGM13]. Mediante [BR11b, HM11, MMM^{+}12, MVB^{+}11, PAM^{+}10, SSD10, WGB^{+}11b]. mediated [ZRM18]. Mediating [HPL14]. mediator [MNW^{+}17]. Medición [GSB10, SPPN^{+}12]. Medida [FMJP11a, FP10, MMPF10, Rog11, SKC^{+}12]. Medidas [LH11, WJ10]. Medina [Cat11]. Medio [HF10, VJMH12, ZYYH10]. Mediterranean [Fat14, SAC12, TSA^{+}14]. Mediterráneos [SAC12]. Medium [Str13]. medius [RC12]. Meet [PSAZ11, PPM^{+}18]. Meeting [Ald12]. Meets [Mor12]. Megadiverse [Fie15, LSV^{+}19, TUHK^{+}17]. megafauna [WLRR18]. Mejillones [Hic10]. Mejorar [BG10]. Mejores [BSBM11]. melting [TK15]. Memoria [TBY^{+}10]. Memory [TBY^{+}10, AC15]. Menabe [SMGRJ10]. Mercado [AWKR^{+}11, BM10, GBMRG^{+}10]. Mercados [LW10]. Merchants [LM18]. Merging [SG19]. mesic [Mat15]. mesopredator [MNW^{+}17]. mesopredator-driven [MNW^{+}17]. Message [Str13]. messages [BHO^{+}18]. Meta [BAPP11, BR11b, GNWSB10, HTH^{+}10, HM12, KKK14, KK18, MLLS14, MSF^{+}12, PBH^{+}10a, PBH^{+}10b, SJZ19, VJMH12, DSE15, Had15, SEM^{+}15, BAPP11, JCA^{+}11, PBH^{+}10a, VJMH12, ZBM^{+}10]. Meta-análisis [PBH^{+}10a]. Meta-analítica [GNWSB10]. meta-analyses [Had15]. Meta-Analysis [BAPP11, BR11b, HTH^{+}10, HM12, KKK14, MLLS14, MSF^{+}12, PBH^{+}10a, PBH^{+}10b, VJMH12, KK18, SJZ19, DSE15, SEM^{+}15]. Meta-Analytic [GNWSB10]. metabarcoding [SJD^{+}19]. Metabolism [GTC^{+}16]. metacommunities [RDS17]. Metacommunity [MQAMWM12]. Metacomunidad [MQAMWM12]. Metáforas [CND10]. Metal [BBB14]. Metaphor [Mar10]. Metaphors [CND10]. Metapoblaciones [FLP^{+}11]. Metapopulation [CFL14, SHPR13, VESHR18]. Metapopulations [FLP^{+}11]. Metas [AH12, MO11, SKC^{+}12]. method [AGFR18, CGS^{+}16]. methodology [ZSM18]. Methods [Fli10, GOR^{+}12, Lyn13, PMS13, WHB^{+}11, MRPT19, Pet17, WLCW16, Ano19w]. Método [CT11, ESCF12, MC11b, PA12]. Métodos [WHB^{+}11, Fli10, GOR^{+}12]. Metric [SA11, BLM^{+}16, SBG^{+}18]. Métricas [GOR^{+}12]. Metrics [GBB^{+}19a, GOR^{+}12, WJ10, GV16, HCH17, JBH^{+}16, PG16, SL15]. Mexican [CFL14, UCFV10]. Mexicana [UCFV10]. Mexico [ACL^{+}12, FEC^{+}10, FMF10, JLCHNG^{+}17, RGRMPB^{+}13, RRV10, SK10, SAGSG^{+}14, VLGRS14, ACL^{+}12, FEC^{+}10, FMF10, RRV10]. Microbes [Sik12]. microbiological [DBV^{+}18]. Microbiome [RSS^{+}12]. Micronesia [MWL^{+}15]. Microtargeting [MPP^{+}19]. Middle [RC12]. Migración [MVB^{+}12, ZM12]. Migrant [ZBD11, SKBE15]. Migration [BC14a, BP13, CPB14, HMIE14, MVB^{+}12, MRT^{+}12, SMK14, ZM12, PAMA17, SMF^{+}15]. migrations [SLB^{+}15]. Migratorias [WKF^{+}11]. Migratorios [ZBD11]. Migratory [BCC^{+}14, LZD^{+}13, WKF^{+}11, DAKN^{+}17, HNO^{+}18, KOA^{+}15, YJL^{+}18]. Miller [KLL^{+}16]. million [GBB^{+}15]. Mills [TT16]. mimic [BC18, VM18]. Mimicry [VM14]. Minimize [DBV13, HAT^{+}15]. Minimizing [BW13, EEB15, MNW^{+}14, MDD^{+}18, RWF^{+}17, Hug19]. Minimum [BC16, RM14a, SBJG13, SBJG14, MVVS15]. Mining [SBSF14]. mink [DLMAG15]. misconceptions [SSC^{+}19]. mislabeling [WSC^{+}17]. mismatch [BJ16, WW17]. Mismatches [GMCW13, CML^{+}19, SFAA^{+}18, WHEN^{+}16]. Misrepresentation [Oso15, PFN15b]. Missing [AG10, APL^{+}13, NBW^{+}12]. mission [GLJ^{+}16]. Mississippi [FHBK18]. Mistakes [GKP13]. Misuse [SA11]. Mitigar [AOLK11]. Mitigate [BBWK^{+}14, HMB^{+}13, BTM^{+}15, GSSF19]. Mitigating [AOLK11, GHT^{+}15, SJH13, CBCW18]. Mitigation [Bee13, ACW18, DSK^{+}19, JBB^{+}15, KLN^{+}17, NJK^{+}19, SG18, SNK15, TMK^{+}16]. mitigation-driven [DSK^{+}19]. Mitochondrial [DHH^{+}12]. mixed [BSCB18, CBB18a]. mobile [DCP^{+}19, WTT^{+}17]. mobilization [NYC17]. Model [BG10, DHD14, HCLW13, KG10, MAA^{+}10, PZM14, Rog11, SAC12, SSD10, BC18, GD15, GGB^{+}15, MF15, RBK^{+}17, VM18]. Modelado [GCC^{+}12, KRS^{+}10, UCFV10]. Modelar [SAC12]. modeled [SCHM18]. Modeling [BBWK^{+}14, GCC^{+}12, KRS^{+}10, KGS^{+}11, UCFV10, VPMS13, WW16, AdBR15, BGB17, JGL^{+}16]. Modelo [BG10, KG10, MAA^{+}10, Rog11, SSD10]. Modelos [BRV10, HSB12, KGS^{+}11, MM10, RTL^{+}12, WGB^{+}11b, CJDZ10]. Models [BRV10, CRKM^{+}14, CJDZ10, CB13, CEM^{+}13, FWM14, GFGR12, HSB12, MBL^{+}14, MM10, RL11, SMTW13, SHPR13, SG10, TTEM14, WGB^{+}11b, ALA17, ALW^{+}19, BH18, CBB18b, EEB15, FRY^{+}17, GMNO15, LMJK^{+}19, MLMP^{+}15, NMS^{+}18, PBB^{+}15, RNE^{+}19, SBR^{+}19, TPL18, TMK^{+}16, ZV19]. Modern [LCT^{+}14, MO11, SSBN12]. Moderna [SSBN12]. Modernas [MO11]. modifiable [MBFB18]. Modificadas [MA10]. Modificados [BHK10]. Modification [Hur13]. Modified [BHK10, MA10]. molecular [DDv18]. Monachus [SBT^{+}11]. Monarch [VLGRS14]. Monasteries [LWY^{+}14]. Monetary [DTH^{+}14]. monetization [TJJB18]. money [Sny15]. monitor [BMC^{+}18, SL15]. Monitoreo [CJDZ10, GSB10, MTM12, RMGCR10]. Monitoring [BVC^{+}14, BMB^{+}19, BEA^{+}13, CJDZ10, EIS14, GSB10, GZNC13, Jew13, MTM12, MMK^{+}14, Moo19, NBSM12, PVP^{+}13, RMGCR10, RBHW18, SSSD14, TCP13, VLGRS14, BBB^{+}16, CL16, DSK^{+}19, DHF16, EGB18, JLCHNG^{+}17, JGL^{+}16, MWL^{+}15, OJAV15, WBF^{+}19]. Monk [Bai13, SBT^{+}11]. monkey [GFdM^{+}19]. monoculture [GHT^{+}15]. Montañas [LPT12]. Monte [CAB^{+}10, FWM^{+}12, VMGB^{+}10]. Moral [Kir14, LNNG16, LSÁL^{+}19]. More-Comprehensive [HODB10]. Moreton [WKF^{+}11, WKF^{+}11]. Mortalidad [Wit11]. Mortality [AASK12, Bee13, PGC^{+}13b, Wit11, PAMA17, RHR^{+}17, SR17, TIB^{+}16]. Mosaic [ZBD11, BMB^{+}19]. Mosaico [ZBD11]. Mosaics [NKSB^{+}13]. mosquito [BSL^{+}19]. Mosses [MFKMP10]. Most [ZZW^{+}10]. Motivaciones [DK10, FKM^{+}11]. Motivating [NWAB19]. motivational [CY17]. Motivations [AB13a, DK10, FKM^{+}11, NGB^{+}16]. Mountain [HWM10, LP10]. Mountains [LPT12, SLWL15, BMG^{+}17, EIS14]. Move [HSR^{+}13, Whe11, TAB^{+}18]. Movement [BTMF13, RMKM14, BC16, DCP^{+}19, HAT^{+}15, LMM^{+}17]. Movements [CPB14, EBL^{+}14]. Moving [BC14a, DSW16, FBF^{+}14, JCCC14, Zac10, CBB18b]. Mozambique [JVT^{+}13]. much [MAM^{+}15]. Muerta [CMJ^{+}11]. Muestras [MFRH12]. Muestreo [HS10]. Muestreos [FCGJH11]. Multicriteria [DBR13, DBB^{+}16]. multidimensional [KLN^{+}17]. Multidisciplinary [MAJH^{+}14]. Multiescala [ML10]. Multimethod [JGL^{+}16]. multiobjective [FRY^{+}17, HNO^{+}18]. Multiple [BTBDS12, BPT^{+}11, CBM^{+}11, JHV13, KMGCR10, LZD^{+}13, PA12, PMMG14, RP13, SPN^{+}14, VPMS13, WSHB11, WB10, BBF19, CBPM19, CPM^{+}18, TMK^{+}16, WÖ16]. Múltiples [BTBDS12, BPT^{+}11, CBM^{+}11, CJDZ10, KMGCR10, PTN10, PA12, SD11, WSHB11, WB10]. Multiregion [EBRS14]. Multiscale [ML10, PTN10, SBB^{+}14, GLF^{+}18]. Multispecies [CRKM^{+}14, CJDZ10, IFP14, NBH^{+}17, PTN10, SD11, ARDG17, BGD^{+}15, DHG^{+}17, HLS^{+}18, LB15, MGMC15]. Multistate [FWM14, JGL^{+}16]. Multivariate [PCRD14]. Multiyear [DSK^{+}19]. Mumbai [BAGM13]. Mundo [GZA10, SBN^{+}10]. Murciélagos [ATW^{+}11, FCC^{+}11, HM11]. Murray [AWU^{+}11, AWU^{+}11, Sch10]. Murray-Darling [AWU^{+}11, AWU^{+}11]. Museo [SS11]. Museum [SS11]. Musgos [MFKMP10]. Mussel [Hic10]. Mutación [CW10]. Mutation [CW10]. Mutualisms [Nas17]. Mutualistas [AZCT12]. Mutualistic [EVF13]. Mutualists [AZCT12]. My [Med10]. Myanmar [DRA^{+}15, LPA^{+}17, PSA^{+}17]. mycology [IBT18, MCD^{+}19]. Mysteries [Ban12].

N [Ano19w, Fie15]. Nacional [CLRV11, Hic10, MFN^{+}11, MJP^{+}11, RBC^{+}11, SHA^{+}10a]. Nacionales [ZBM^{+}10]. Namibia [NWD^{+}16a]. Nariz [FCC^{+}11, HM11]. Narratives [Mol19]. Narrow [AGLRV14]. Nation [MAA^{+}10]. National [BPA^{+}13, BBC^{+}14, BCC^{+}14, CLRV11, CRG12, Cla13a, CCB^{+}19b, Cor19, CTRE^{+}17, GDMF13, HCP^{+}12, HWM10, Hic10, JVT^{+}13, JHV13, MFN^{+}11, MF14, MJP^{+}11, RYW^{+}15, RIBT18, RBC^{+}11, SCK13, SW10, SHA^{+}10a, ZBM^{+}10, BARB17, CPD^{+}15, MAM^{+}15]. Nativa [EG12]. Nativas [Chi10, FG12, MMCUR^{+}12, Pra10, SSO11, SKCS11, VLRB12]. Native [AGP^{+}14, AZCT12, AZT^{+}14, CLT^{+}13, Chi10, Cla14, Cla15, EG12, FG12, HPL14, JBB^{+}14, LWLC12, MGCL14, MRMV14, MMCUR^{+}12, Pra10, RSPD13, SSO11, SSO12, SGE13, SBSF14, SLSH14, SLSH15, SKCS11, Van12, VFV^{+}12, VLRB12, WR12, BSL^{+}19, CBWFL15, FMB^{+}18, TPL18, WDW^{+}19]. Nativo [MMCUR^{+}12, Van12]. Nativos [AZCT12, WR12]. Natura [BBM18, KHD^{+}15, MAM^{+}15]. Natural [BRCK11, CMG^{+}13, DTH^{+}14, Dre11, GSB10, GBMRG^{+}10, HH14, JDA^{+}10, Jos12, KRG13, MB14, AG17, CZS19, CL16, CC17, GDB15, GFJC19, HBTMG15, LZZ^{+}17, ML15, NYC17, SPD18, SAS^{+}19, TLT16, WA18, WFMW^{+}18]. Naturales [BRCK11, GSB10, GBMRG^{+}10]. Naturaleza [BM10, HDH^{+}10, LDBS11]. naturalist [MCD^{+}16]. naturalized [GRL15]. Naturally [HWW12]. Naturalmente [HWW12]. Nature [BM10, GLB^{+}13, LKB^{+}10, LDBS11, SLH10, GBB^{+}15, HRHC16, HGGH15, HHB15, LAA^{+}15, LdBB^{+}17, LKW^{+}15, MCM^{+}19, PS16, RYW^{+}15, SA15, VBN15, WÖ16, Ano19w]. Navigating [Heb12]. Navjot [BLG^{+}11]. Near [GLB^{+}13, GDMF13, MFN^{+}11, Arm18, CE15]. near-future [CE15]. Near-to-Nature [GLB^{+}13]. Necesidad [HS10]. Necesidades [PSAZ11]. Necessary [BDG^{+}13]. Need [Fur19b, GLF^{+}18, HS10, Hur13, KF18a, ML14, NAB^{+}17, RKW19, Vit17, Wat11, WHB^{+}14, WON^{+}13, CWS16, JHAH18]. needed [HW19, KCD^{+}17, WDV^{+}16]. Needs [PSAZ11, AP17, CCB^{+}19a, DLMAG15, KB17, MA19, PPM^{+}18, WBF^{+}19]. Negative [HBD^{+}14, MYG^{+}14, MMCUR^{+}12, ZBD11, ESD^{+}17]. Negativos [MMCUR^{+}12, ZBD11]. Neglect [LAA^{+}10a]. Negociables [WW10b]. Neighborhood [SS12]. neighbors [NWAB19]. Neotropical [GMBM13, LFMRD12, MRBL^{+}12]. Nepal [AAS^{+}17]. Nest [HM12, LB11, LPT12]. Nesting [LB11, MHG^{+}14, Van12, ZRM18]. Nests [MFN^{+}11]. Net [BGL^{+}14, GHB^{+}13, SBSF14, CZS19, GBBMG19]. Network [GAFG14, GMCW13, ML10, MSR19, RTL^{+}12, Sou10, SBN^{+}10, VPMS13, WJ13, WRG^{+}12, ARDG17, BMAS17, BLM^{+}16, GFJC19, IRK^{+}18, KHD^{+}15, MAM^{+}15, MVVS15]. Networks [DBK^{+}14, EVF13, HMD^{+}10, HHA^{+}11, HMM10, KCSP11, LGC^{+}13, UCFV10, ARDG17, BTM^{+}15, DAKN^{+}17, DHG^{+}17, HWB^{+}16, ITG^{+}16, TLT16]. Neuter [LDB^{+}10, LCL13]. Nevada [Hic10]. new-conservation [HSF17]. newly [SC19]. News [Mor12, MGG^{+}13, PNC^{+}15]. Newspaper [JLCK12]. Next [WON^{+}13, PWKB19]. next-generation [PWKB19]. Niassa [JVT^{+}13]. Nicaragua [HTDM17]. Niche [UCFV10, LB15]. Niches [HRL14]. Nicho [UCFV10]. Nidos [LB11, LPT12]. Nielsen [Ano19w, ERBM17]. Nigeria [MJA^{+}11]. Niño [PMK^{+}10]. Nivel [BR11a, HRSK12, RBC^{+}11, WJ10]. Niveles [CGCL11, JLDJ12, WSHB11]. NJ [Ano16x, Ano19w]. No [Ano12g, AZCT12, BB10a, BGL^{+}14, CB10, CMRH10, EG12, GT10, GHB^{+}13, GBBMG19, HSB12, LGH^{+}12, LGK^{+}11, MMCUR^{+}12, PBH^{+}10a, PFN15a, SSO11, SBSF14, Van12, VLRB12, WR12, HMD^{+}10, WWdS^{+}16]. No-Take [LGK^{+}11, WWdS^{+}16]. Noise [BBP12, GS11, TMD^{+}17]. nomadic [PRB18, RTH^{+}15]. nomenclature [Les15, ZPJ19]. Non [AZCT12, AZT^{+}14, CB10, Chi10, Cla14, Cla15, DEJM10, EG12, JBB^{+}14, LWLC12, MRMV14, MMCUR^{+}12, SSO11, SSO12, SLSH14, Van12, VFV^{+}12, VLRB12, WR12, BSL^{+}19, TPL18, WDW^{+}19]. Non-Native [AZCT12, AZT^{+}14, Chi10, Cla14, Cla15, EG12, JBB^{+}14, LWLC12, MRMV14, MMCUR^{+}12, SSO11, SSO12, Van12, VFV^{+}12, VLRB12, WR12, SLSH14, BSL^{+}19, TPL18, WDW^{+}19]. Non-Open [CB10]. Non-Use [DEJM10]. Nonacademic [BDG^{+}13]. nonepistemic [BH17]. Nongovernmental [HSB12]. Nonintervention [CMRH10]. Noninvasive [PWRM10]. nonmaterial [GKA^{+}15]. Nonmonetary [DTH^{+}14]. nonmonogamous [TIB^{+}16]. nonparametric [CGS^{+}16]. nonredundancy [SLS17]. nonthreatened [BGO^{+}19]. Nontimber [GT10]. Nord [Ber13]. Noreste [GLH12, TAG^{+}10]. normal [VJG^{+}15]. Normas [HDH^{+}10]. normative [YNR19]. Noroccidental [EDB12]. Norte [CMB12, LB11, MRPPM12, HM12, MRRK11, NFF^{+}10, RM11]. Norteamérica [GLH12]. Norteamericano [TTSZH10]. North [CMB12, Mat13, AB13b, BCM16, BCLMM19, CM17, Cor19, EB15, GLH12, HM12, MLRC18, MPS^{+}16, SSRS18, SVN^{+}15, TTSZH10]. Northeastern [GLH12, SMS14, TAG^{+}10]. Northern [CCM12, Ehr14, HODB10, LB11, SMB^{+}10, CMG15, WHG^{+}15, MRRK11, NFF^{+}10, RM11]. Northward [MRPPM12]. Northwestern [EDB12]. Norway [CSB^{+}10, HAF^{+}17]. Nose [FCC^{+}11, HM11, DMR^{+}17, ISA16, MF15]. Noss [WMS^{+}12]. Noted [Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano11p, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano14h, Ano14i, Ano14j, Ano14k, Ano14l, Ano15h, Ano15i, Ano15j, Ano15k, Ano16x, Ano16v, Ano16w, Ano17x, Ano17y, Ano17z, Ano17-27, Ano17-28, Ano18r, Ano18s, Ano18t, Ano18u, Ano18v, Ano18w, Ano19w, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Red10b, Ano16y]. Novel [BME^{+}14, FCC^{+}11, NBNMG13, AdBR15, AMS^{+}16, GMNO15, NMS^{+}18, TIB^{+}16]. November [Ano17a, Ano16b]. Nuclear [Hen16, BB15, Die16, HGRS16]. Nueva [FCC^{+}11, FP10, AGAB11, HWW12]. Nuevo [ESCF12]. number [MCM^{+}19, SC19]. Numbers [HRKMG12]. Número [HRKMG12]. Nutrias [ETB10]. NY [Bra15].

o [HS10]. Oahu [LCL13]. Oak [SAC12]. Oaks [PTN10]. Objective [Boo14, RM14a]. Objectives [BPW12, BPT^{+}11, BT14, GPB^{+}15, NBH^{+}17]. Objetivos [BPT^{+}11, PTN10, BPW12]. Obligate [BSF^{+}14]. Obscuring [PHFPP10]. Observaciones [KRS^{+}10]. observation [RKW19]. Observations [KRS^{+}10]. Observed [NA13]. Obtained [BJR^{+}11]. Obtención [JGH^{+}11, MBF^{+}12]. Obtenido [BJR^{+}11]. Occidental [LSBB10]. Occidentales [GBG^{+}10, AS10]. Occidente [CMB12, SSD10]. Occupancy [AGLRV14, CRKM^{+}14, EIS14, EPD12, GS11, HD14, HLC13, KRS^{+}10, Lyn13, PSP11, PCBD11, RC12, TG13, ALW^{+}19, BCSM15, GMNO15, MBFB18, SHSP19]. Occupancy-Based [EIS14, BCSM15]. Occupation [HRKMG12]. Occurrence [BA13, HPL14, MMCUR^{+}12, NA13, SS11, ITG^{+}16, JBH^{+}16, LDD19, OSM^{+}17]. Occurring [RJF^{+}13]. Ocean [BWE^{+}13, CKK^{+}14, GFGR12, CE15, HAT^{+}15, DFW^{+}10, EMJ14, GGL^{+}15, GOR^{+}12, LK14, ÖB12, RMKM14]. Oceania [WA18]. Oceanic [CHHR13]. Océano [GOR^{+}12, ÖB12]. Oceans [Fur19a, Fur19b, SCHM18]. October [Ano16a, Ano19b]. Ocupación [EPD12, GS11, HRKMG12, KRS^{+}10, PSP11, PCBD11, RC12]. Ocurrencia [SS11, MMCUR^{+}12]. Oddity [RDM10]. Odonata [SKSK14]. Off [CBA13, Dre17]. Official [AGAB11]. Official-Looking [AGAB11]. Offs [BRKP12, CKKL10, HSBJ12, HAB^{+}11, MG12, CGK^{+}19, Cha15, DHG^{+}17, LBI^{+}18, MSMCB18, MBV^{+}16, VGAB^{+}19, WB14, HAB^{+}11]. Offset [HFSB13, HCH17]. Offsets [GHB^{+}13, QM11, RCK13, CBCW18, RB17]. Offsetting [SBSF14, CS17a, HHW^{+}16, NM18, NW15]. Offshore [LW11, WMW^{+}14, KBML15]. offspring [GLdPL18]. Offtake [GFRK12]. Ogasawara [UMK^{+}16]. Oh [Med10]. Ohlone [CD13]. Oil [FBA^{+}13, LKB^{+}10, WRTB10, GHT^{+}15]. Oil-Palm [FBA^{+}13, GHT^{+}15]. Old [BM14, Lic16, SFKC16, SGB^{+}19]. old-field [SFKC16]. oldest [KLS^{+}19]. Oliveira [SPL^{+}11]. Oliveira-Santos [SPL^{+}11]. omitting [MCH^{+}16]. on-the-ground [IBA16]. One [Bar12, CMNM10, PFA^{+}14b, Poz17, SSRS18]. Ongoing [PG12]. online [PNC^{+}15]. only [COZS16]. onshore [VTT^{+}17]. Ontogeny [CBA13]. Opción [LAA^{+}10b]. Opciones [MARV^{+}11, DBP^{+}10]. Open [Bur19, CB10, EMD^{+}10, FLW14a, FLW14b, Len17, MNW^{+}14, HAT^{+}15]. open-ocean [HAT^{+}15]. Operacionales [HFK^{+}10]. Operational [HFK^{+}10]. Opinion [DWCW10, DWCW10]. Opiniones [Rud11b]. Opinions [Rud11b]. Oportunidad [KCDC10]. Oportunidades [GBG^{+}10, OKLP11]. Oportunistas [KRS^{+}10]. Opportunistic [KRS^{+}10, HBTB15]. Opportunities [GBG^{+}10, GMG^{+}19, OKLP11, PMD^{+}13, SNH^{+}14, AKJ^{+}18, CHN^{+}15, KS17, RLB^{+}17, YJL^{+}18]. Opportunity [CWAW14, CKPS10, KCDC10, LA13, MAJH^{+}14, RMRO14, MDD^{+}18]. Óptima [RHM10]. Optimal [IFP14, KBCM15, MFN^{+}11, RHM10, WÖ16, DHG^{+}17]. optimality [Dre17]. Optimism [BCC^{+}14, CSB^{+}10]. optimistic [ODAL17]. optimization [RBHW18]. Optimize [BMM^{+}14, GDM^{+}19]. optimized [Dre17]. Óptimo [MFN^{+}11]. Opting [Pri12]. Options [BSF^{+}14, DBP^{+}10, LAA^{+}10b, MARV^{+}11, MNW^{+}14, WGJ13]. orangutan [NGB^{+}16]. Orchid [GGG10, LCC^{+}15]. orchids [HLHR16]. order [Gru17]. Oregon [RTL^{+}12, SSD10]. organisms [KM18]. Organización [WSHB11]. Organizaciones [BPT^{+}11, HSB12, ÖB12]. Organization [WSHB11]. Organizations [BPT^{+}11, HSB12, ÖB12, Rob12, GLJ^{+}16, PWP^{+}18]. organizing [RHPF15]. Orientadores [LH10]. Oriental [GOR^{+}12, GSW10, PB11]. Oriented [Car14]. origin [MSY^{+}15, PMM^{+}15]. original [MAJ16]. Ornithochorous [HG10]. Ornitócoras [HG10]. Orquídea [GGG10]. Orthoptera [PVP^{+}13]. Osofsky [PFN15b]. Other [PMD^{+}13, VKKA13, NM18]. Otherwise [WRTB10]. Otras [CMNM10]. Otter [CCM12]. Otters [ETB10, LMCS17]. Our [Ano19w, Fur19a, Fur19b, Hal16, Kah10, Sik12, Tru18b, Abr15]. Outbreaks [SMH^{+}14]. Outbreeding [FBE^{+}11, RKH^{+}14]. Outcomes [HFSB13, HSRC14, HBR^{+}14, OCP^{+}14, VWM^{+}10, AG17, GFJC19, OHHE16, PA18, RHWS18, WBF^{+}19, WW16]. outdoor [GTVC16, WLL^{+}16]. Outreach [JIAMS^{+}14, SMK14, CQC^{+}19]. outsourcing [IBA16]. Overabundance [WNY13]. Overall [JMW^{+}11]. overemphasis [YNR19]. Overestimating [Tep12]. Overfishing [MAA^{+}10]. overharvested [LGCF19]. Overhead [DHD14]. Overlap [RMKM14, LB15]. Overlooked [SSW19, KCD^{+}17]. overseas [DOC^{+}15]. Overwintering [LHDC12]. Ovis [JMW^{+}11]. Owl [SMB^{+}10, TG13]. Owls [HODB10]. Own [Sik12]. owning [GG19]. Oxford [Bra15, Fie15]. oyster [AC15]. Oysters [MRMV14].

P [Ano19w, Fie15]. Pacific [GGL^{+}15, GOR^{+}12, Boy10, CHHR13, DFW^{+}10, FLP^{+}11, GSDE14, NBW^{+}12, TQD^{+}18]. Pacífico [GOR^{+}12, FLP^{+}11, NBW^{+}12]. Pacíficon [Boy10]. Pact [CSB^{+}10]. Pagar [WL12]. page [Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano18n, Ano18o]. pages [GLJ^{+}16]. Pago [CLV^{+}10, SMGRJ10, NFF^{+}10]. Pagos [CBS11]. Paired [MBL^{+}14]. País [MAA^{+}10]. Paisaje [JLDJ12, MRRK11, SD11, VSM^{+}11]. Paisajes [AOLK11, BS10, GBG^{+}10, GZA10, GZR12, LH11, WW10a]. Palenque [PAM^{+}10]. Paleobiology [RL13, WWW^{+}12]. pallid [FHBK18]. Palm [FBA^{+}13, LKB^{+}10, GHT^{+}15]. panda [LP16]. Pandas [ZZW^{+}10, WSH^{+}15]. Pandemic [WNY13]. Panels [BHSA12, HBE^{+}10]. panic [LSÁL^{+}19]. Pantanal [ROQR^{+}16]. Pantera [JLCK12]. Panther [JLCK12]. Panzootic [HM11]. Panzoótico [HM11]. Papel [DWH^{+}11, HFL^{+}11, RRV10, WR12, MS12]. Paper [JCCC14]. paperback [Ano16x, Ano19w]. Papers [Ano16x, CB10, CV16]. Papua [BTB^{+}18, Mel10]. Para [SHCD11, AWU^{+}11, AWKR^{+}11, ACL^{+}12, AS10, AOLK11, BRV10, BAPP11, BB10a, BSBM11, Ber12, BV10, BG10, BS10, BR11b, BRCK11, BAG^{+}11, CJDZ10, CMB12, CLV^{+}10, Chi10, Col10, CCBF12, CGM10, DFW^{+}10, DK10, DWCW10, EMD^{+}10, ESCF12, ETB10, FC10, FJP^{+}11, GZR12, GPK^{+}10, HF10, Han10, HNS^{+}11, HSAB10, HDH^{+}10, HHA^{+}11, HBE^{+}10, HJSF^{+}12, IP11, JDA^{+}10, JLDJ12, KRS^{+}10, KGS^{+}11, KE10, KCDC10, KG10, LMM^{+}10, LW11, LHDC12, LAA^{+}10b, LH11, LNP10, LW10, LH10, LGK^{+}11, LOME10, MPH^{+}11, MFN^{+}11, MAA^{+}10, MO11, ML10, MC11b, MR12, NBW^{+}12, NFF^{+}10, OTKE12, OKLP11, ÖB12, PRR^{+}11, PTN10, PSAZ11, PA12, PCTW10, PF10, PHC^{+}11, RRCB^{+}11, Rog11, RHM10, RBC^{+}11, SBT^{+}11, SD11, SHA^{+}10a, SCL12, SAC12, SPSS10, SHO^{+}10, SH11b, UCFV10, VMS^{+}12, VLRB12, VSM^{+}11, WL12, WEC^{+}11, WRAW10]. para [WHB^{+}11, WRTB10, ZM12]. Paradigm [RGRMPB^{+}13]. Paradigms [DCB^{+}16, Sha10b, WG15]. Paradoja [TTSZH10]. Paradox [TTSZH10, ME16]. paraecologists [SAB^{+}16]. Parameters [PMK^{+}10, JC18]. Parámetros [PMK^{+}10]. parasite [DCB^{+}16, Jør15, SMGR17]. Parasites [LZD^{+}13, MBGS11]. Parasitic [BAG^{+}11, WWW^{+}12]. Parasitology [OK10]. Parásitos [BAG^{+}11, MBGS11]. parataxonomists [SAB^{+}16]. Parche [RC12]. Parches [MLS12, RC12, RSE12, SHLK^{+}10]. Parcial [CMRH10]. Park [BMGLW12, CLRV11, CRG12, GDMF13, LW11, MBH14, CPD^{+}15, CTRE^{+}17, MFN^{+}11, SCK13, SHA^{+}10a]. Parks [BCC^{+}14, Cla13a, Cor19, HWM10, Pri10, SW10]. Parque [LW11, CLRV11, MFN^{+}11, SHA^{+}10a]. Parques [BMGLW12]. Parrot [WWW^{+}12]. Parrotfish [AS10, MYG^{+}14]. parrots [FMB^{+}18]. Partial [CMRH10, MRFvW16]. Participación [FTSSW12, SHCD11]. participants [VLSOS16]. Participation [FTSSW12, SHCD11, WHS18]. Participativo [MAGA10]. Participatory [AJ18, EGB18, MAGA10, AdBR15, BBB^{+}16, NJK^{+}19]. Partir [FB12, MFRH12]. Partnership [Rey11, RHWS18]. parts [MSY^{+}15]. Parúlidos [ZBD11]. Passenger [RJS17]. Passeriformes [KTN12]. Passive [JLCHNG^{+}17]. Past [Cla13a, CEM^{+}13, NCC13, VESHR18, ALW^{+}19, GML^{+}19, KMT^{+}17, ODAL17]. Pastizales [BEE^{+}11, RWB11]. Pasto [MMCUR^{+}12]. pastoralist [ESD^{+}17]. pastoralists [AP17]. Pastoreo [GSW10]. Pastors [O’B11a]. Pasture [FMF10]. Patch [HAP^{+}13, RC12, ALW^{+}19, ZRM18]. Patch-Occupancy [RC12]. Patches [RSE12, SHLK^{+}10]. Patchy [GZR12, PY13]. Path [Dor11, McP12]. Pathogen [GLH12, SGH^{+}11, BMAS17, BGH^{+}15, GGB^{+}15]. Pathogens [SAGSHL10, VMS^{+}12]. pathologies [CC17]. Paths [Ore12]. pathway [Hul15]. pathways [ZAP^{+}18]. Patógeno [GLH12, SGH^{+}11]. Patógenos [SAGSHL10, VMS^{+}12]. Patriotism [O’B11a]. patrol [DMGB^{+}19]. Patrones [GS11, HSB12, MAGA10, ZYYH10, CB10, LN11]. Patterns [AS13, CB10, FWM14, GS11, HMIE14, HD14, HSB12, KGG^{+}14, LN11, MAGA10, PFR^{+}14, TBC18, ZYYH10, GS16, HTE16, ISA16, LDD19, ML15, MGW^{+}15, PAMA17, PTJ^{+}19, RN17, SMARC^{+}16, ZV19]. Pay [WL12, AAvRS19, FWC^{+}15, LH19]. pay-for-release [LH19]. Payment [SMGRJ10, DSSW16]. Payments [CLV^{+}10, CBS11, Fer11, KWW^{+}14, LA13, NFF^{+}10, SAGSG^{+}14, CMG15, PVO^{+}16, TVY^{+}16]. Peak [CGM10]. peatlands [RLB^{+}17]. Peces [AWU^{+}11, AMO11, AS10, CGCL11, DMZ10, FCGJH11, PF10]. Pedagogy [Lal11]. Peer [Wil12b, Dar15]. Peligro [ANTBG12, CVN^{+}10, Chi10, DWCW10, Duf11, Fli10, GGG10, GLNJ10, HB10, LN11, PF10, Van12, WL12]. Peñaranda [TYA16]. penguins [TGBB^{+}15]. Pennisetum [FMF10]. Pennsylvania [HD14]. Pensamiento [CCBF12]. People [EG13, GDMF13, HP14, Kni11, Lin11, MBH14, SCSJ13, SM14, WD12, Abr15, GBBMG19, LKW^{+}15, AL14]. People-Centered [HP14]. peoples [LZR^{+}18]. Pepinos [LAA^{+}10b]. Pequeña [VMGB^{+}10]. Pequeñas [ZZW^{+}10]. Pequeño [ERC^{+}10]. Pequeños [DEJM10]. Perceived [BMP^{+}12, OB11b, WHS18, BCM16, DSG^{+}17]. Percepción [OB11b]. perception [CM17, KRF^{+}19]. perceptions [Ben16, EAPB15, GTVC16, LMCS17, MSC^{+}18, WLL^{+}16]. Perchan [ATW^{+}11]. Percibidos [BMP^{+}12]. Pérdida [GZR12, HRBR11, Jam11, RTL^{+}12, VJMH12, ZBD11]. Pérdidas [KG10, KPL^{+}12]. Perennial [FMF10]. Performance [FW17, KWW^{+}14, LKB^{+}10, OGS^{+}19]. Performance-Based [KWW^{+}14]. Periféricas [TGD10]. Peril [Zac10]. Perilous [Yap11]. Periodicidad [HFK^{+}10]. Periodicity [HFK^{+}10]. Periodística [JLCK12]. Peripheral [TGD10]. Permeability [BTMF13]. Permisos [WW10b]. permit [HAF^{+}17]. Permits [WW10b]. perpetual [TWB^{+}18]. Perros [RM11, VSM^{+}11]. Persa [BTBDS12]. persecuted [DHF16]. Persecution [Vég18]. Persian [BTBDS12]. Persistence [CBM^{+}11, EWN^{+}13, LHDC12, RSW11, SMS14, ALW^{+}19, BGZ^{+}15, CCN16, CMM17, HLS^{+}15, HSV^{+}17, MTD16, PCF^{+}19]. Persistencia [CBM^{+}11, LHDC12, RSW11]. Persistent [WES^{+}13]. Personal [Sco16, GLK^{+}17]. Perspectiva [AMO11, WD10]. Perspectivas [SHO^{+}10].

Perspective [AMO11, Sco16, WD10, GBB^{+}15, HCBB^{+}15, KT16, SKBE15, STJ^{+}19]. Perspectives [CRG12, Lor13, SHO^{+}10, DWM^{+}19, HSF17, ZSM18]. Perth [Fie15]. Perturbación [FG12, SKCS11]. Peru [CLRV11, CLRV11]. perverse [NM18]. Pesca [FCGJH11, GOR^{+}12, LGK^{+}11, ÖB12, WGR^{+}10]. Pescadores [DC11, HRKMG12, MCN10, Mcc10, TBY^{+}10]. Pesquería [SHC^{+}11, HSBJ12]. Pesquerías [AS10, BJR^{+}11, GC10, KKCS10, Mcc10, MO11, SGS^{+}10]. Pest [IP11, ABG^{+}19]. Pest-Invaded [IP11]. pet [GTVC16, HTH^{+}17a]. Peter [Hel14, Pea18]. Peterson [PHJ^{+}17]. Petróleo [CGM10]. Petrolera [WRTB10]. Petroleum [BPA^{+}13, CGM10]. Pets [BBM14]. Phenological [WW17]. Phenotypic [DHH^{+}12, RSW11]. philanthropy [FFHA19, LHKA16]. Philippines [HRKMG12, WRAW10]. philosophy [SJH19]. photo [BCK^{+}19]. Photogrammetry [Ber12]. Photographic [Ano16x]. Phylogenetic [Fli10, FP10, LOME10, PBB^{+}11, WR12, GBD^{+}15, Les15, LKYL15, QCW^{+}18, SMM16]. phylogeny [CCCCN18, GRL15]. Physiological [ATW^{+}11, DMR^{+}17]. Physiology [CLV14]. Pico [CGM10, SSB^{+}12]. Pigeon [RJS17]. Pimm [Pet17]. Pine [CMRH10, SKCS11]. Pino [CMRH10]. Pinos [SKCS11]. Pioneer [Die10b, TAG^{+}10]. Pioneros [TAG^{+}10]. Pipan [Bra15]. Piracy [Dav11]. Pirogénico [MQAMWM12]. Placas [HBE^{+}10]. Place [HPD16]. Place-based [HPD16]. placement [KJWP19]. Places [WKI^{+}14]. Plagas [IP11]. Plague [EB15]. Plains [AB13b]. Plan [BB10a, HN17, Ber13, LET19, MBW^{+}19]. Planet [Hal16, KHK^{+}19]. Planificación [ACE^{+}10, BR11b, CJDZ10, DWM^{+}10, HRC10, MFKMP10, MMPF10, MWC11, PSAZ11, RRC^{+}10, SD11, SFM^{+}11, BPT^{+}11, CWH^{+}10, PTN10, SHO^{+}10]. Planificadores [SHO^{+}10]. Planificar [BB10a]. Planner [SHO^{+}10]. Planning [ACE^{+}10, BPT^{+}11, BR11b, CCW^{+}16, CJDZ10, CRNR12, CWH^{+}10, CKK^{+}14, DFW^{+}10, DWM^{+}10, GMCW13, HMIE14, LRA^{+}14, LGC^{+}13, MFKMP10, MF14, MMG^{+}13, MMPF10, MWC11, NLM^{+}13, PTN10, PSAZ11, Pri10, RMM^{+}13, RRC^{+}10, SD11, SFM^{+}11, SM14, SHO^{+}10, Tru18b, TTEM14, BTM^{+}15, BPH^{+}19, GLF^{+}18, GWRW18, GPB^{+}15, HHJ^{+}19, LKYL15, NAB^{+}17, NBH^{+}17, OSM^{+}17, RWF^{+}17, SMAG17, THD^{+}18, TUHK^{+}17, TG15, VSA^{+}15, WGS^{+}19, WTT^{+}17, WD17]. Plans [LW14, PS10a, Wil12b, ACB^{+}15, GW17, HH19c, NRD^{+}16, RK17, VJ15]. Plant [ANTBG12, BW16, CEM^{+}13, EDB12, EVF13, FB11, Fie15, FJP^{+}11, FMF10, GG13, HCM14, LDBS11, NCC13, Pra10, SAC12, TGD10, WW10a, WWW^{+}12, ZCCN13, GRL15, Hul15, LKYL15, LSV^{+}19, SCA17]. Planta [ANTBG12, Pra10]. Plantaciones [NS10]. Plantas [AZCT12, EDB12, FB11, FJP^{+}11, FMF10, HG10, LDBS11, MMCUR^{+}12, Pra10, SAC12, TGD10, VJMH12, WW10a]. Plantations [FBA^{+}13, NS10, SGE13]. plantings [GSB^{+}16, ITG^{+}16]. Plants [AASK12, AZCT12, AZT^{+}14, BSF^{+}14, KZG^{+}13, LCFS14, MMCUR^{+}12, Pra10, RBW14, Tep12, VJMH12, AOPM^{+}19, AWHW17, BLR18, HHKK15, LRL^{+}15, LGCF19, Pyk17, RBK^{+}17, SLGB15]. Plántulas [HG10]. Plasmodium [LZD^{+}13, SMGR17]. Plasticidad [RSW11]. Plasticity [RSW11]. Platform [RGMR12]. playback [WZC19]. Playera [EG12]. Plazo [CW10, GGG10, TTSZH10]. Pleistocene [OSF10]. Plight [Tho10]. Pluralidad [SSVA11]. Plurality [SSVA11]. Poaching [KRG13, Blu16, BMF16]. Población [BAJ^{+}10, HBM^{+}12, PGLB^{+}12]. Poblacional [ANTBG12, GGG10, HNS^{+}11, HRSK12, JMW^{+}11, KPL^{+}12, RSW11, RC12, SS11, WJ10, WKP11]. Poblacionales [ANTBG12, BFHP10, HB10, KRS^{+}10, NBW^{+}12]. Poblaciones [CBM^{+}11, CMB12, DWH^{+}11, KCSP11, PBK^{+}11, PA12, SPSS10, SHLK^{+}10, TGD10, ZYYH10, ZZW^{+}10]. Pockets [O’B11a]. Poco [HWW12]. Poder [HDH^{+}10]. point [TXBW19]. Poisoning [CRD^{+}14, ZMTR15]. Polarotactic [HBE^{+}10]. Polarotácticos [HBE^{+}10]. Polen [ZYYH10]. Policies [MBH14, PCTW10, Ric11, HHB15, MLMP^{+}15, SH16]. Policy [BM14, Cla13b, FBF^{+}14, FTB^{+}16, Hur13, JMF12, RBC^{+}11, Rud11a, SSSD14, Tör18, Wil12b, Agg15, BPN^{+}16, CHG^{+}17, CM17, CS17b, DPJM18, MWL^{+}15, PWP^{+}18, RTB^{+}18, Ros15, WGLL17]. policy-relevant [Ros15]. Polinizadores [JDA^{+}10, KUT11]. Political [BL17, DPT^{+}18, PSA^{+}17, GZR18]. Políticas [PCTW10, Rud11a, RBC^{+}11]. Politics [Phi10, Pri12, RB13]. Pollen [WMW^{+}14, ZYYH10]. Pollinator [JDA^{+}10, KUT11, LDC^{+}13]. Pollinators [AZT^{+}14, HCT^{+}17]. polluted [SHSP19]. Pollution [TGBB^{+}15, BH18, FS15]. polygon [Pet17]. polygon-based [Pet17]. Pond [UIN^{+}13, WH13, IO17, SSP^{+}15]. Pond-Breeding [WH13]. Ponds [UIN^{+}13]. pool [MA19]. Pooley [Oso15, SH13]. poor [PG16]. Poorly [Bra15, AHT^{+}15]. Populated [Fat14]. Population [AA12, ANTBG12, BFHP10, BCGP13, BHB^{+}19, BAJ^{+}10, BJMT14, CGHM13, CHHR13, CM13, DBV13, EBL^{+}14, GGG10, GGL^{+}15, GG13, HNS^{+}11, HRSK12, HB10, HBM^{+}12, Hur13, JMW^{+}11, KRS^{+}10, KKH14, KPL^{+}12, MDJ14, NBW^{+}12, PGLB^{+}12, PMJ^{+}13, RSW11, RM14a, RC12, SMTW13, SBJG13, SBJG14, SS11, Tep12, VMW13, WJ10, WES^{+}13, WKP11, ZCCN13, AK16, BCA^{+}18, BFB^{+}19, CMM17, COZS16, DMB^{+}17, DHF16, DUS^{+}16, DP15, EBP^{+}18, FWH19, GRT^{+}18, HLS^{+}15, HSV^{+}17, ISA16, KOA^{+}15, KGF^{+}18, LSK^{+}15, LS19, LBB^{+}16, MF15, MLMP^{+}15, PBB^{+}15, RBK^{+}17, RBHW18, RCARR18, RHPF15, SBR^{+}19, SR17, SMZZ17, THD^{+}18]. population-density [SBR^{+}19]. Population-Level [WJ10]. Population-Trend [KRS^{+}10]. population-wide [SMZZ17]. Populations [AGP^{+}14, BW13, Bee13, CBM^{+}11, CMB12, CEM^{+}13, GAFG14, HD14, JPT13, KCSP11, LCT^{+}14, PBK^{+}11, PA12, SCK13, SPSS10, SHLK^{+}10, TGD10, WGJ13, ZYYH10, ZZW^{+}10, BFF^{+}16, Ben15a, BMA^{+}18, BHF^{+}19, BGH^{+}15, BIP^{+}15, CSR^{+}19, DPT^{+}18, DS19, GGL^{+}15, GGB^{+}15, HAT^{+}15, HJH16, HLS^{+}18, HWB^{+}16, KB17, LSA17, MRMH17, SKBE15, WFSS19]. Populations. [DWH^{+}11]. Por [RBL10, AOLK11, BHK10, Ber12, BAJ^{+}10, CLV^{+}10, Chi10, CAB^{+}10, FG12, FBE^{+}11, GOR^{+}12, GKK^{+}11, IP11, JMW^{+}11, KUT11, KPL^{+}12, MSSM09, OSC10, OBB^{+}11, OSFFB^{+}11, PHC^{+}11, RMGCR10, RTL^{+}12, SGS^{+}10, SSK^{+}11, SFM^{+}11, SMGRJ10, VJMH12, WT10, WL12, WGR^{+}10]. Porphyrio [GLNJ10]. Porqué [CND10, JCA^{+}11, RER10]. portion [ENW^{+}18]. Portrayal [GLJ^{+}16, MGG^{+}13]. posed [NRRW17]. Posible [PGLB^{+}12]. Posibles [CW10]. Posicionales [GGG10]. position [BIF^{+}18]. Positional [GGG10]. Positive [Lal11, MYG^{+}14, MRMV14, TTSZH10, ESD^{+}17]. Positives [CC17]. Positivos [TTSZH10]. Possible [CW10, PGLB^{+}12, WES^{+}13]. Post [CMRH10]. Post-Fuego [CMRH10]. Postbreeding [CPB14]. postconflict [NAB^{+}17]. Postfire [CMRH10, CMJ^{+}11]. postnormal [CLvK^{+}17]. Pot [BHSA12, TK15]. Potencial [BBBG10, FB11, SSO11, SA11]. Potenciales [ACL^{+}12, FEC^{+}10, OTKE12]. Potential [ACL^{+}12, BKKM14, BBBG10, BBB14, FB11, FEC^{+}10, GZJ^{+}13, HBD^{+}14, LFMRD12, LKB^{+}14, LWLC12, OTKE12, PG12, SK10, SSO11, SA11, VFV^{+}12, WWW^{+}12, ZCCN13, BPY^{+}16, BHF^{+}19, BW19, HHW^{+}16, IBT18, LZR^{+}18, LSV^{+}19, NW15, VLSOS16, VB19]. Pothole [RWB11]. poverty [CCB^{+}19a, DJBB16]. Power [BAPP11, BGG13, DHD14, EIS14, HDH^{+}10, TCP13, TSH^{+}14, CRMK18]. Pox [ZLKW13]. pp [Ano16x, Ano19w, Bra15, Fie15]. Práctica [HF10, HNS^{+}11, HFL^{+}11, MC11a]. Practical [AB13a, BPH^{+}19, HNS^{+}11, KTF^{+}14, OPP14, PGBV^{+}13, WK15]. Prácticas [BSBM11, GC10]. Practice [CKK^{+}14, Fer17, HF10, HFL^{+}11, HH14, Tör18, CLvK^{+}17, PWP^{+}18, RIBT18, WZWF18]. Practices [BSBM11, GC10, LW14, MC11a, CMMP15, FR16, MRPT19, SDT17, WL17a]. Practitioner [MSC^{+}18, CS19, WDS15]. practitioners [BKR17]. Praderas [FMF10]. Pragmatic [CS14, MB10]. Pragmático [MB10]. Prairie [AB13b, MHG^{+}14, EB15, RWB11]. Prairie-Chickens [MHG^{+}14]. Preaching [Hal16]. Precio [MJA^{+}11]. precipitation [ZRM18]. precise [EG17]. precursor [HPD16]. Predation [LPT12, SRS11]. Predator [BJMT14, HSAB10, HJMM14, LPT12, SPSS10, Van12, BGB17, PBB^{+}17]. Predators [BB10b, LK14, Lin11, TTSZH10, RBF15, RJA^{+}16, WBNR15]. Predecir [KG10]. Predicción [PA12, FBE^{+}11, MFRH12]. Predicciones [PBB^{+}11]. Predict [CEM^{+}13, TTEM14, AB15, CMM17, DMB^{+}17, EEB15, HBTB15, TMM^{+}18]. Predictability [CCCCN18]. Predicting [BKKM14, BCOB15, FBE^{+}11, GFGR12, HW19, IRK^{+}18, KG10, LMC^{+}14, LFPT14, MFRH12, NCC13, OSM^{+}17, PA12, PA18, PS15, RL11, BGB17, SC19]. prediction [KSH17]. Predictions [PBB^{+}11, GRT^{+}18, KAM18]. Predictive [DHD14, ML14]. Predictor [CS13, HSV^{+}17]. Predictores [PSP11]. Predictors [BA13, CWH^{+}17, PSP11, GUHE15]. Preference [MS12, NSBJ17b, NSBJ17a]. Preferences [DTH^{+}14, LL14, ERBM17, RB17, SSMC17]. Preferencia [MS12].

Preferred [SH16]. Preguntas [RBC^{+}11]. Premium [DVW^{+}13]. Preparación [HJSF^{+}12]. preregistration [PFN19]. Presas [KPL^{+}12, VMGB^{+}10]. Presence [AGAB11, FG12, JVT^{+}13, LPT12, MGCL14, MRT^{+}12, VSM^{+}11, COZS16]. presence-only [COZS16]. Presencia [AGAB11, FG12, LPT12, VSM^{+}11]. present [GML^{+}19, KMT^{+}17, VESHR18]. Preserve [Zha15]. Press [Ano16x, Ano19w, Bra15, Fie15]. Pressure [GJB14]. Pressures [PVP^{+}13]. Prevalence [LW11, NBNMG13, OSFFB^{+}11, ZLKW13]. Prevalencia [LW11, OSFFB^{+}11]. Prevenir [PRR^{+}11]. Prevent [PRR^{+}11]. preventing [SDT17]. Prevention [GH10]. previously [BGP16]. prewhaling [MPS^{+}16]. Prey [FC10, HPL14, Sco16, VMGB^{+}10, SBU^{+}17, LBLB14]. Prey-Trap [FC10]. Price [MJA^{+}11]. Pricking [Tor11]. Primary [FMF10]. Primate [NBM^{+}12, SMS14, CTBR19]. Primates [RDM10, Har15, LLR^{+}15]. Prince [PMM^{+}15]. Princeton [Ano16x, Ano19w]. principal [TT16]. principal-agent [TT16]. principalis [GCC^{+}12]. Principles [Whi13, DFR^{+}17, GLJ^{+}16, MA19, PVO^{+}16, WLD16]. Prioridad [BEE^{+}11]. Prioritarias [RBC^{+}11]. Prioritarios [ODTS10]. Priorities [DHM^{+}15, OPP14, AWT^{+}15, Cla15, DAKN^{+}17, DLMAG15, DRF^{+}18, DSV17, DCP^{+}19, EGB18, FMV^{+}19, PS15, TYA16, WA18]. Prioritization [ALL^{+}12, HSB12, KMTK13, Rud11a, TCDR12, UCFV10, ERBM17, GMV^{+}18, NSBJ17a, NSBJ17b]. prioritize [ARDG17, HLS^{+}18, PG16]. Prioritizing [DOC^{+}15, PRNV18, WKI^{+}14, TMJ^{+}15]. Priority [BEE^{+}11, DFJ^{+}12, GKP13, LMC^{+}14, ODTS10, RBC^{+}11, SGB^{+}19, AGFR18, CNvL^{+}15, GDG19, JWS^{+}18, KTBH17, VKM^{+}17]. Priorización [ALL^{+}12, HSB12, Rud11a, TCDR12, UCFV10]. Privadas [KCDC10, PCTW10]. Privados [HRC10, MC11b, SHCD11]. Private [HRC10, KO12, KCDC10, MC11b, PCTW10, SHCD11, CC17, CC18, GG19, NWAB19]. Private-Land [KO12, MC11b, CC18]. privilege [BKR17]. Probabilidad [PMS^{+}10]. Probabilidades [BEE^{+}11]. Probabilities [BEE^{+}11, CCN16]. Probability [FBE^{+}11, PMS^{+}10]. Probar [Rog11]. Problem [EG13, HHY13, Pri10, MBFB18, SSWR17, TT16]. Problematizing [Blu16]. Problems [DSW16, Pet17, SNK15, LPW16, Sny15, WHEN^{+}16]. Proceso [RRC^{+}10]. Procesos [WGB^{+}11b]. Process [BC14b, RRC^{+}10, SSSD14, WGB^{+}11b, LBD^{+}15, Wal17]. Processes [BCT^{+}14, CHG^{+}17, CS19, GFdM^{+}19, PBB^{+}15, TMK^{+}16]. Producción [CGM10]. produce [BW19]. Producer [BR11a]. Producer-Level [BR11a]. Producidos [BAJ^{+}10]. producing [SKLP16]. Production [CB13, Cop12, CGM10, MG12, MSMCB18, NRD^{+}16, vWWCF15]. Productividad [FMF10]. Productivity [FMF10, LWZL16]. Productores [BR11a]. Productos [GT10]. Products [GT10, VW19]. proenvironmental [GLK^{+}17]. Profesionales [SSVA11]. professional [BFG^{+}16]. Professionals [FBB^{+}14, SSVA11]. Profiles [EBL^{+}14]. Profiling [HBTMG15]. Profitability [HSBJ12]. Profits [Dav11]. Profundas [VLRB12]. Program [RYW^{+}15, RKH^{+}14, SCSJ13, SAGSG^{+}14, SHCD11]. Programa [SHCD11]. Programación [KCDC10]. Programas [BG10, IP11, Jam11, JGH^{+}11, MS12, MC11b]. Programs [BMG^{+}14, BG10, HDN^{+}14b, IP11, Jam11, JGH^{+}11, MS12, MC11b, ACW18, AMS^{+}16, CL16, CCS^{+}16, HGP16, KB17, MCD^{+}16, NMS^{+}18]. Progreso [AH12]. Progress [AH12, GH18, VMÁR^{+}13, WSH^{+}15, BLR18, GBB^{+}19a, GBB^{+}19b, WG15]. Project [MBJWC10, LS15]. Projected [LSBB10, RTL^{+}12]. Projecting [JPT13]. projection [MLMP^{+}15]. Projects [BKKM14, BC14b, BGG13, JIAMS^{+}14, MGW^{+}15, PG16, PVO^{+}16, TMJ^{+}15, VLSOS16]. Prolongados [PMK^{+}10]. Prolonged [PMK^{+}10]. promise [PHJ^{+}17, SGB^{+}18]. Promote [PGBV^{+}13, NRD^{+}16]. Promoting [HGRS16, PMN^{+}16, WBNR15]. Prone [NKSB^{+}13, SMB^{+}10]. pronghorn [WLCW16]. Propagación [HSBE12]. Propagation [HSBE12]. property [GZR18, HFBG18]. Propietarios [MC11b, RRV10, SHCD11]. Proporcionada [RMGCR10]. Proporcionados [Chi10]. Proporciones [HFK^{+}10]. Proposals [CRD^{+}14]. Proposed [BHK10, PG12]. Propuesta [BHK10]. Prospects [CKK^{+}14, MV18, KBML15, SSW15]. Protección [CGCL11, SPPN^{+}12, HDH^{+}10]. Protect [BC14a, HDH^{+}10, AAvRS19, Arm18]. Protected [AME^{+}13, BPA14, BRKP12, CKKL10, CGCL11, DC11, DBK^{+}14, FJP^{+}11, GJB14, GDMF13, GC10, HBR^{+}14, HMIE14, HHA^{+}11, KMTK13, LK13, MCN10, MJ11, MO11, ML10, NA13, QKC^{+}19, RM11, SMK14, Sch14b, SG10, SSK^{+}11, SPPN^{+}12, Sou10, WT10, WEC^{+}11, WRAW10, WJ13, WGF^{+}12, WRG^{+}12, ZM12, ANE^{+}16, AC16, Ben15c, BBM18, BTM^{+}15, BVAC17, BGP16, BLM^{+}16, BMG^{+}17, CESS^{+}18, CY17, CEdK18, CMG15, CVMM17, DSG^{+}17, DKL^{+}15, GDB15, GH18, GPB^{+}15, HKFC15, HICF16, KHD^{+}15, KJWP19, LMJB^{+}19, LTN18, LLR^{+}15, MAM^{+}15, MV18, MVVS15, NCKC17, OHHE16, PCF^{+}19, QCW^{+}18, SCA17, SL15, TMM^{+}18, TWB^{+}18, VMO^{+}18, WHS18, WDV^{+}16, WMFB15]. Protected-Area [FJP^{+}11, ML10, SPPN^{+}12, WEC^{+}11, BLM^{+}16, CVMM17, LMJB^{+}19, LTN18, MV18, VMO^{+}18, WHS18]. protected-species [NCKC17]. Protecting [Plo17, TJJB18]. Protection [CGCL11, Fur19b, GRA14, RYW^{+}15, SPPN^{+}12, CE15, HKB^{+}17, MLRC18, PRNV18, SSD^{+}18, SSW15, SBU^{+}17, STM^{+}19]. protective [LP16]. Protegida [CGCL11]. Protegidas [BRKP12, DC11, FJP^{+}11, HHA^{+}11, ML10, RM11, SPPN^{+}12, WT10, WEC^{+}11, WRG^{+}12, ZM12, CKKL10, GC10, MCN10, MO11, SSK^{+}11, WRAW10]. Protein [VCDF13]. Protocol [PMJ^{+}13, GKA^{+}15]. prototyping [VTM19]. Proveedoras [GZA10]. Providing [SGE13]. Provinces [Pek13]. Provision [Chi10, DEJM10]. proxies [FW17]. proximate [LPA^{+}17]. Proximity [BCGP13]. Proxy [LCT^{+}14]. Proyección [LSBB10, RTL^{+}12]. Proyecto [MBJWC10]. Prueba [EG12, CMRH10]. Pseudoreplication [RSS^{+}13]. psychological [GG19, MCO^{+}17]. psychology [SGB^{+}18]. Public [Cor16, JIAMS^{+}14, KRF^{+}19, KBSX19, O’B11a, TM10, Tru18a, WL12, CM17, DKL^{+}15, GTVC16]. Publication [Ano18x, MBG14, AGS19, CVA15, HHJ^{+}19]. publications [BJ16]. Público [WL12]. Públicos [TM10]. published [Ano12g]. Publishing [TAB^{+}18, LET19, MBW^{+}19]. Puede [BBBG10]. Pueden [RHM10]. Puentes [ATW^{+}11]. pumas [AJ18]. Purported [WJ10]. Purpose [BC14b, Sch14a]. Purse [GOR^{+}12]. Purse-Seine [GOR^{+}12]. Putativamente [RER10]. Putatively [RER10]. Putting [Bor18, KBM17, SM14]. Pyrenean [GGM14]. pyrodiversity [KBM17]. Pyrogenic [MQAMWM12].

Q [ZSM18]. Quagga [Hic10]. Qualitative [SMV19]. Quality [BLC13, Jew13, KUT11, LHDC12, RC12, RJF^{+}13, SHLK^{+}10, UIN^{+}13, LS15, LPW16, WWdS^{+}16]. quantification [BCSM15, JHAH18, TBM^{+}16]. Quantify [BAG^{+}11, EPB^{+}14]. Quantifying [ABB^{+}18, BLR18, CBPM19, CBM16, CVMM17, COZS16, DTH^{+}14, GDB15, HCP^{+}12, KJWP19, LZR^{+}18, MBC^{+}16, PNC^{+}15, RTGR17, WJ10, CGS^{+}16, SKLP16]. Quantitative [ENW^{+}18, MBJWC10, PTN10, PF10, RTT^{+}13]. quantitatively [MRMH17]. Quantity [BLC13]. quarter [SKBE15]. Que [GHKK10, JCA^{+}11, ATW^{+}11, FKM^{+}11, FMJP11a, HSAB10, HBM^{+}12, LB11, MBJWC10, RHM10, SKC^{+}12, SHO^{+}10, BBBG10]. Québec [Ber13]. Questioning [CM17, FDBA16]. Questions [PFA^{+}14b, RBC^{+}11, JWS^{+}18]. Quitridio [GLH12, MSSM09, PMS^{+}10]. Quotas [LGK^{+}11, JC18].

R [Ano19w, Fie15]. Rabbit [PBR^{+}16]. Racovitzan [FCS19]. radar [WFC^{+}18]. Radiación [OSFFB^{+}11]. Radiation [MAWF12, OSFFB^{+}11, Sto11]. Rainforest [EWN^{+}13, KZG^{+}13, MC14, NA13, PS10b, BGZ^{+}15, SFKC16]. Rainforests [AS13, MLS12, SMS14]. Ranas [KPL^{+}12, SAGSHL10]. randomized [CCG18]. Range [CWS^{+}14, Cat11, Fat14, FMJP11a, FMJP11b, JDA^{+}10, KPL^{+}12, RMKM14, SBT^{+}11, BIF^{+}18, CBM^{+}18, DSSW16, ENW^{+}18, HWB^{+}16, KAM18, RCSK15, RTH^{+}15, SMF^{+}15]. Range-Wide [KPL^{+}12, SBT^{+}11, SMF^{+}15]. ranger [CPD^{+}15]. ranger-collected [CPD^{+}15]. ranges [Pet17, PNSG18, San19, SGA^{+}19]. ranging [LM18]. Rango [FMJP11a, JDA^{+}10, KPL^{+}12]. Rank [DWCW10, WD17]. rankings [SMM16]. Rapid [AL14, BIP^{+}15, FCC^{+}11, HLC13, MAGA10, PGC^{+}13a, VTM19, KP19, LAA^{+}15]. Rápida [FCC^{+}11, HJSF^{+}12]. Rapidly [TBY^{+}10]. Rápidos [TBY^{+}10]. Raras [TGD10, WHB^{+}11]. Rare [BA13, LTQ^{+}13, Tep12, TGD10, VMW13, WGJ13, WHB^{+}11, AOPM^{+}19, AWHW17, ESD^{+}17, SC19]. Rareza [PF10, RDM10, TCDR12]. Rarity [Fat14, PF10, TCDR12]. Rat [AGP^{+}14]. Rate [CS13, OB11b, WKF^{+}11, SSP^{+}15, VJG^{+}15]. Rates [ANTBG12, HNDE13, JMW^{+}11, SDB10, CCCCN18, CGE19, GLdPL18, MCMP15, RHPF15, SR17, TIB^{+}16]. Ratio [WES^{+}13]. Ratios [HFK^{+}10, GLdPL18]. Rats [RSPD13, DSV17]. Rattlesnakes [CBSZ10]. Re [HP14, NBW^{+}12, PWKT18]. Re-Creating [NBW^{+}12]. re-embrace [PWKT18]. Re-Inventing [HP14]. Reach [BHO^{+}18]. Reaching [O’B11a]. readiness [HH19c]. Reading [Mei12]. Real [GZA10]. Real-World [GZA10]. Realistic [CCM12]. reality [KBC^{+}18]. reassess [SKSS^{+}17]. reassessing [BBK^{+}17]. Reassessment [Mat13]. Recalibrating [TW14]. Recaptura [KGS^{+}11]. Recapture [KGS^{+}11, WLCW16]. recaptures [FHBK18]. Received [Ano11a, Ano11b, Ano14a, Ano15l, Ano16a, Ano16b, Ano17a, Ano18b, Ano19b]. Recently [Ano15l]. Recesiones [Wit11]. Reciente [Ser10]. reckoning [San19]. Reclasificación [WL12]. Reclassification [BPA14]. Recognition [MCD^{+}19]. Recognizing [ZFW^{+}11]. Recolectadas [MFRH12]. Recolectados [LGH^{+}12]. Recommendations [PGBV^{+}13, LSK^{+}15]. reconcile [CPR^{+}16, CCB^{+}19a, DPS^{+}16]. Reconocimiento [HAB^{+}11]. reconstruction [BTTP17, OJAV15]. Record [Kar14, Ano16x]. recorders [vdWSM^{+}16]. Recordings [KB11, BMC^{+}18]. Records [Duf11, FB12, COZS16, SA15]. Recover [CFL14]. Recovering [SLSH14, WL12, DMR^{+}17]. Recovery [AEO^{+}13, Boo14, BJMT14, CVN^{+}10, CRNR12, CCM12, CS13, DWH^{+}11, DBP^{+}10, ETB10, HODB10, KKH14, NCC13, PS10a, SS10a, Wil12b, WGF^{+}12, ZCCN13, ABB^{+}18, AK16, CCN16, CM17, CPM^{+}18, CCS^{+}16, HH19c, LFM17, PBR^{+}16, SFKC16, SEM^{+}15, SJZ19, TG15, TVY^{+}16]. Recreación [Hic10, RM11, NBW^{+}12]. Recreation [Hic10, RM11, JRWA15]. Recreational [DC11, YFB15]. Recreationist [CD13]. Recreativos [DC11]. recruit [NWAB19]. recruitment [AOPM^{+}19, GLdPL18]. Recuento [MB10]. Recuperación [CVN^{+}10, DWH^{+}11, DBP^{+}10, ETB10, WL12, SS10a]. Recursos [GSB10, PSAZ11, RHM10]. Red [Bar12, GOR^{+}12, JHV13, KMH^{+}14, LAA^{+}10b, PTN10, RRCB^{+}11, RCARR18, RTL^{+}12, SLGB15, SBR^{+}19, SMM16, SBN^{+}10, SKCS11, TBC18, WRG^{+}12, ZBM^{+}10, ALW^{+}19, COZS16]. red-legged [ALW^{+}19]. Red-List [JHV13, COZS16]. Red-Listed [PTN10]. REDD [CSB^{+}10, ABD^{+}17, FDBA16, FDBA17, GKB10, HP14, Mel10, MGW^{+}15, NMSF11, VWM^{+}10]. Redefining [Hal18, Kah10]. Redes [HMD^{+}10, HHA^{+}11, KCSP11, UCFV10, HMM10]. Redford [SLH10]. rediscoveries [ZAB16]. Redistribution [MK18]. Reducción [Fli10, MA10, HLB^{+}12, HBE^{+}10, LB11]. Reducciones [ZZW^{+}10]. Reduce [BHSA12, BBB14, ÖB12, VW19]. Reduced [DMZ10, Fli10, HLB^{+}12, LB11, WWdS^{+}16]. Reduced-Impact [DMZ10]. Reducido [DMZ10]. Reducing [BAPP11, GFM^{+}13, HBE^{+}10, SHG^{+}14]. Reducir [BAPP11, ÖB12]. Reduction [MA10, MK18]. Reductions [ZZW^{+}10]. reductive [Pet17]. Reed [SBJG14]. Reef [BCGP13, CGHM13, DFW^{+}10, MAA^{+}10, MYG^{+}14, Mcc10, NBW^{+}12, PCRD14, VPMS13, CESS^{+}18, MM15, MGMC15, BPW12, CPM^{+}18, LGK^{+}11]. Reefs [HHY13, LW11, LCT^{+}14, Yap11, ANE^{+}16, AC15, BH18, KT16, RBF15, WWdS^{+}16]. Reevaluating [CBB18b, HAV11]. Reexamining [LS19, SBJG13]. Refaunation [CTRE^{+}17, OSF11]. Referenced [SA14].

Referencia [MFN^{+}11, NFF^{+}10, SD11]. Refining [MBFB18]. Reflection [FB11]. reflections [NJK^{+}19]. Reflejo [FB11]. Refocusing [SR11]. reform [DRA^{+}15]. Reframing [WHR^{+}16]. Refuge [PHC^{+}11, Rob13, HCT^{+}17]. Refuge-Effect [Rob13]. Refuges [MF14]. Refugia [Chi10, MLRC18]. Refugio [PHC^{+}11]. Refugios [Chi10]. Regarding [LL14, ALW^{+}19, GTVC16, HKB^{+}17]. Regeneración [BEE^{+}11]. Regeneration [BEE^{+}11]. Regime [BSF^{+}14, FKRC17]. Régimen [KPL^{+}12]. Regimes [KPL^{+}12, BARB17, IRK^{+}18]. Region [FFM^{+}14, Fat14, Fie15, NKSB^{+}13, RWB11, WRG^{+}12, RWB11, WRG^{+}12]. Regional [AMO11, BSBM11, CBA13, JMP^{+}11, LDC^{+}13, ISA16, JGL^{+}16]. Regionally [HD14, SBN^{+}10, DBV^{+}18]. Regionalmente [SBN^{+}10]. regions [Bra17, Pyk17]. registered [PFN19]. Registros [Duf11, FB12]. Reglas [GCC^{+}12]. regrowth [SFKC16]. Regulación [EPD12]. Regulating [SASO13]. Regulation [EPD12, TAP19]. Regulations [CHHR13, RHT14, HGM19]. Rehabilitation [BP13, PS18]. reinstatement [CTRE^{+}17]. Reintroducció [ACL^{+}12, BTBDS12, Jam11, VMS^{+}12]. Reintroducciones [PA12]. Reintroduced [CM13]. Reintroduction [ACL^{+}12, BTBDS12, Jam11, KGG^{+}14, MBG14, OSF11, PS11, VMS^{+}12, AMS^{+}16, BGB17, DBV^{+}18, GFdM^{+}19, HGP16, HJH16, Jør15, LDD19, PA18, SSD^{+}18]. Reintroductions [HMG^{+}13, PA12, AOPM^{+}19, CBWFL15]. Reissued [Cla13b]. Relación [LB11, WR12]. Relacionadas [HDH^{+}10]. Relaciones [LDBS11]. Related [HDH^{+}10, WLL^{+}16]. Relatedness [WR12]. Relating [GBB^{+}19b]. Relation [Pyk17, RM14a, TG13, VKKA13]. Relations [Fat14, LDBS11, MAWF12, WNY13, PBB^{+}17]. Relationship [Kah10, PZM14, PFR^{+}14, BIF^{+}18, GLK^{+}17, LKW^{+}15, OJAV15, WWH^{+}18]. Relative [Lin11, LB11, SKCS11, BBH17, KJWP19, MBC^{+}16, SSE^{+}17]. Relativos [SKCS11]. Release [BTBDS12, LCL13, LH19, PS18, WDW^{+}19]. released [BCLMM19]. releasing [WC19]. Relevance [JMF12]. Relevant [Car14, Ros15, SBU^{+}17]. Reliability [PPD^{+}12]. reliable [BW19, HWMC15, PFN19]. relictum [SMGR17]. Religion [Car11, TWZ16]. religious [WDW^{+}19]. Relocation [BSF^{+}14, OKLP11]. Remain [Car14]. Remains [Fle10]. remedies [CWS16]. Remembering [Hel14, MMGN18]. remembrance [Pea18]. Remnant [SMS14]. Remoción [EDB12, SS10a, SPSS10]. Remotas [FWM^{+}12]. Remote [FWM^{+}12, Arm18, RBE^{+}15]. Removal [EDB12, LCL13, SS10a, SPSS10]. removals [IRK^{+}18]. Removed [GOR^{+}12]. Rentabilidad [HSBJ12, MARV^{+}11]. Rentable [CLV^{+}10]. Reorientación [SFM^{+}11]. Reorienting [SFM^{+}11]. Replacement [AZT^{+}14, HGCT13]. Replacements [GZJ^{+}13]. Reply [AR10, BC18, FMJP11b, HODB10, MSSM10, PI11, SPL^{+}11, TW11, Wil12b, BCA^{+}19, FDBA17, IF17, KF18a, LDC^{+}15, MBT^{+}17b, MBW^{+}19, Nat18, PS16, Pet17, SVN^{+}15, SGA^{+}19, TK15]. reported [GLK^{+}17, GBB^{+}19b]. Reportes [RER10]. Reporting [RMGCR10, Had15, ISM^{+}18, KMLC19, MRPT19]. Reports [RER10, PFN19]. Represent [HMD^{+}10, SPN^{+}14, WEC^{+}11].

Representación [LNP10]. Representan [HMD^{+}10]. Representar [WEC^{+}11]. Representation [LNP10, LZZ^{+}17, AB15, BdA15, BSH^{+}15, BGP16]. Representations [BAGM13]. Reproducción [BS10, KE10, MS12, PMK^{+}10]. Reproducing [PMK^{+}10]. Reproduction [BS10, AOPM^{+}19, MBBB17]. Reproductiva [HFK^{+}10, Rog11]. Reproductivas [KPL^{+}12, WR12]. Reproductive [CWS^{+}14, MS12, WR12]. Reproductivo [MS12]. Reptile [NKSB^{+}13, DSK^{+}19, TND16, TLPP17]. Reptiles [MRPPM12, WIM^{+}10, WR12]. requirements [BC16, TLPP17]. requiring [BMAS17]. Rescate [CMJ^{+}11]. Rescue [HAV11, PS18]. Research [DMM14, EMS^{+}11, Fur19a, JMF12, MRPT19, MB14, NS18, PSAZ11, PGBV^{+}13, RBC^{+}11, Rud11a, SABV13, SCL12, SH13, Tru18a, WA18, BJ16, BMAS17, FMV^{+}19, GMG^{+}19, GFJC19, Had15, KMR^{+}18, MCMP15, NYC17, PNC^{+}15, PS18, SAB^{+}16, TAB^{+}18, WBPL15, WGLL17, ZSM18]. research-implementation [KMR^{+}18]. Research-Prioritization [Rud11a]. Researchers [WON^{+}13, RKC^{+}19, SH16]. Resemblance [BC18, VM18]. Resembling [VM14]. Reservas [BJR^{+}11, HMD^{+}10, LGH^{+}12, LGK^{+}11, MFKMP10, SBN^{+}10]. Reservation [MFKMP10]. Reserve [BPA^{+}13, GDM^{+}19, HMD^{+}10, JVT^{+}13, LGC^{+}13, SBN^{+}10, VLGRS14, OGS^{+}19, SMY^{+}17, SMZZ17, SKPP15]. Reserves [LGH^{+}12, LGK^{+}11, Per11, BP16, CBM16, MCM^{+}19, RYW^{+}15, WÖ16, WWdS^{+}16]. Reservoir [SHB^{+}17]. Reservoir-host [SHB^{+}17]. Resettlement [WS14]. residence [GRL15]. Residencial [WT10]. residency [JRWA15]. Residential [WT10]. Resilience [ACS14, CP14, SAB12, Wat10, SRG^{+}15, TQD^{+}18]. Resiliencia [SAB12]. Resilient [Qui14]. Resistance [IP11, LTN18]. Resistencia [IP11]. Resolución [ALL^{+}12, BRV10]. Resoluciones [CGCL11]. Resolution [ALL^{+}12, BRV10, LK14, LSA17, SMAG17]. Resolutions [CGCL11]. resolve [EAPB15]. Resolving [DBB^{+}16, Hul15]. Resource [Bro13, CMG^{+}13, KRG13, MBL^{+}14, PSAZ11, CC17, GFJC19, HM16, HBTMG15, NYC17, WA18]. Resources [GSB10, Jos12, LKH^{+}13, RHM10, AG17, CL16, TDCG15]. Respaldo [SSD10]. responders [PG18]. Response [CRKM^{+}14, Cat11, Col10, GMBM13, GB14, Joh11, KZG^{+}13, LDB^{+}10, OSF11, PBH^{+}10b, PFN15b, SBJG14, SLH10, SM11, SMB^{+}10, SH13, VM18, Van12, WMS^{+}12, WRG^{+}12, WGR^{+}10, ABD^{+}17, Bai16, BGB17, BBM15, BCR^{+}17, CCG18, Cla15, DPGY16, DW19, Dys15, EEB15, ERBM17, FS15, FHK^{+}15, GZR18, Har15, Hen16, HGRS16, KLL^{+}16, LS19, LET19, Mat15, MCH^{+}16, MH16, NWD^{+}16b, NSBJ17b, PTJ^{+}19, PHJ^{+}17, RM19, San19, TYA16, TT16, vWWCF15]. Responses [ESCF12, GLH12, LWLC12, MC14, WRTB10, BHPM^{+}18, FSAM15, GUHE15, HTH^{+}17b, SEM^{+}15, TND16, WFMW^{+}18]. Respuesta [SM11, Van12, WRG^{+}12, WGR^{+}10, Col10]. Respuestas [ESCF12, WRTB10, GLH12]. Restauración [HFL^{+}11]. Restaurado [LHDC12]. restaurants [WSC^{+}17]. Restoration [Aro11, Don12, Gon13, GBP^{+}13, HFL^{+}11, HAP^{+}13, PS11, TUHK^{+}17, Tör18, Val18, VPMS13, AHP^{+}15, EGB18, FRY^{+}17, GSB^{+}16, HTH^{+}17b, IPA^{+}17, ITG^{+}16, MDD^{+}18, PTJ^{+}19, SFKC16, WL17a]. Restore [GZJ^{+}13, HGCT13]. Restored [LHDC12]. Restoring [AGP^{+}14, BHM^{+}14, WFSS19]. Restricciones [FMJP11a, Mcc10, MVB^{+}10]. Restrictions [Mcc10, CESS^{+}18]. Resulting [RSS^{+}13]. Results [BCD^{+}16, PS16]. retail [FFS^{+}18]. retail-market [FFS^{+}18]. retaining [WLD16, vdWSM^{+}16]. Retention [WGJ13]. Reto [CGM10]. Retos [EMS^{+}11, GBG^{+}10, LW10, OKLP11, SCL12]. retracted [CV16]. Retroalimentación [MCL12]. Retrospective [MBM^{+}14]. Return [LDB^{+}10]. Returns [AT11]. Reubicación [OKLP11]. Reveal [AWU^{+}11, HJSF^{+}12, WRTB10, SMARC^{+}16, ZV19]. Revealing [GRT^{+}18]. Revelar [AWU^{+}11, WRTB10]. Reverse [SGSR^{+}12]. Review [Bra15, GNWSB10, Löv16, Wil12b, BSH^{+}15, CY17, Dar15, PBB^{+}17]. Reviewers [Ano14a, Ano16y, Ano18a, Ano19a, RLL15]. Reviews [CPF13, Fie15, Red10a, HWMC15, HW19, WL17b]. Revisión [GNWSB10]. Revisiting [HWM10, MBT^{+}17b, SGB^{+}18, VFV^{+}12]. Reviving [San14]. revolving [HFBG18]. Rewilding [OSF10, CHN^{+}15, STJ^{+}19]. Rhinoceros [CWS^{+}14, AAS^{+}17, GLdPL18, MLMP^{+}15]. Ribereñas [EPD12]. Ribereño [LHDC12]. rice [KK18]. rice-field [KK18]. rich [HPD16]. Richness [CLRV11, FMF10, MRBL^{+}12, PBH^{+}10a, HTE16, LKYL15, SSE^{+}17, TBF^{+}17]. Riesgo [FP10, LCG10, MVB^{+}11, SVH12, WKP11, LO10]. Riesgos [AT11]. Right [HCP^{+}12, MRT^{+}12, BCK^{+}19, MPS^{+}16]. Rights [BRS11, GFG^{+}12, FHK^{+}15, GD15, OKCG18, VGAB^{+}19]. rights-based [VGAB^{+}19]. Río [EPD12, MPH^{+}11, FBF^{+}14, PAM^{+}10, DBP^{+}10, HSBE12, PAM^{+}10]. Ríos [KPL^{+}12]. Riparian [EPD12, LHDC12]. Ripe [GL18]. Riqueza [CLRV11, FMF10, PBH^{+}10a]. Rise [HRSK12, IFP14, HN17, KS17]. Risk [BVC^{+}14, BWE^{+}13, CRNR12, DHD14, EPB^{+}14, FP10, HJM11, KMH^{+}14, KGG^{+}14, KLL^{+}16, LO10, LFPT14, LCG10, MMK^{+}14, MBL^{+}14, MVB^{+}11, PRE^{+}13, RMM^{+}13, RTT^{+}13, RKH^{+}14, SHG^{+}14, SHPR13, SKSK14, WKP11, BIF^{+}18, BC16, BBF19, EAPB15, GTVC16, HDH^{+}15, JBH^{+}16, KSH17, PRB18, PS15, QCMB18, RK17, RBHW18, RNM^{+}18, RTH^{+}15, SC15, SBB^{+}15, TBM^{+}16, TMJ^{+}15, WLL^{+}16]. Risks [AT11, KRG13, SVH12, SMB^{+}10, KAM18]. River [EPD12, HSBE12, KKCS10, KPL^{+}12, MPH^{+}11, RTL^{+}12, WBM13, HTH^{+}17b, IRK^{+}18, KRF^{+}19, DBP^{+}10, FHBK18]. River-Breeding [KPL^{+}12]. River-Flow [EPD12]. Rivers [FFM^{+}14]. Road [Bee13, Hug19, KCSP11, Ven18, HMB^{+}15, Hul15]. roadless [KHGF^{+}17]. roadmap [NYC17]. Roads [BTMF13, CBSZ10, FGH^{+}13, SDB10, VKKA13, FHK^{+}15]. Roadsides [MFRH12]. Roble [SAC12]. Robust [MR12, ARDG17, CGS^{+}16, KAM18, SCHM18, VJ15]. Robustas [MR12].

Rocas [MAA^{+}10]. Rock [MAA^{+}10]. Rocky [EIS14]. Rodent [PMK^{+}10]. Rodents [CRD^{+}14]. Roedor [PMK^{+}10]. Roja [PTN10, RRCB^{+}11]. Rojas [ZBM^{+}10]. Rojo [LAA^{+}10b]. Role [BAGM13, DWH^{+}11, Dre11, HFL^{+}11, HJH16, HP11, LWY^{+}14, MS12, Mol19, RGMR12, RRV10, TW14, WES^{+}13, WWW^{+}12, AK16, BDKS15, BBB^{+}16, BB15, CHWH18, HTE16, IBT18, KF18b, PNC^{+}15, PMM^{+}15, SJH19, SVN^{+}15, Wil17, vdWSM^{+}16]. Roles [HAP^{+}13, RTB^{+}18, WR12, ZFW^{+}11]. roofed [ÇRG^{+}17]. roost [SC15]. Roosting [ATW^{+}11]. roots [PWKT18]. rope [KRL^{+}16]. Rotational [KMGCR10, RBW14]. Rotativo [KMGCR10]. Round [MRT^{+}12]. Roundtable [LKB^{+}10]. routes [HAT^{+}15]. Rs [PS18]. Ruido [BBP12, GS11]. rule [Bei19]. Rules [GCC^{+}12, RTT^{+}13, ACW18]. Rural [CAB^{+}10, FWM^{+}12, MAGA10]. Rurales [FWM^{+}12].

s [Mei12, LET19, MBW^{+}19]. Sabana [VSM^{+}11]. Sabanas [GSW10]. Saber [GHKK10]. sabotage [IF17]. Sacramento [MWC11, MWC11]. sacred [SLWL15]. sacrifices [TWB^{+}18]. Safeguarding [ERC^{+}10]. Sage [BBP12]. Sage-Grouse [BBP12]. sake [MBT^{+}17a]. Salamander [CS13]. Salamanders [PCBD11, RJF^{+}13]. Salamandras [PCBD11]. salinization [SSP^{+}15]. Salmon [Con11, FLP^{+}11, HSBE12, JMSM13, WBM13, FLP^{+}11, HSBE12]. Salmonidae [WES^{+}13, WK10]. Salmónidos [RTL^{+}12]. Salmonids [RTL^{+}12]. Salomón [AS10]. Salt [CBA13, PRB18]. Salvador [RM14b]. Salvage [CMRH10, CMJ^{+}11]. Salvaguardando [ERC^{+}10]. Salvamento [CMRH10]. Salvarlos [KE10]. Same [Fle10]. Samoa [WB14]. Samples [MFRH12]. Sampling [HS10, PWRM10, MBBB17]. San [MWC11]. Sanaga [FFM^{+}14]. Sanctuary [HCP^{+}12]. Sand [LTQ^{+}13]. Sanderson [SGA^{+}19]. Sandplains [Fie15]. Santos [SPL^{+}11]. saola [WD17]. Sapos [PMS^{+}10]. Saproxylic [BP16, GLB^{+}13, SBB^{+}15]. Sargassum [PB11]. satellite [LSA17]. Satisfacer [PSAZ11]. Savanna [VSM^{+}11]. Savannah [AME^{+}13]. Savannas [GSW10]. Save [KE10, PS10b, Sha15b, Tru18b, Gru17, PS10b]. Saving [SLH10]. Scalable [BMA^{+}14]. Scale [AC16, Ber13, CKR^{+}13, GMCW13, NBSM12, NGL^{+}14, PSP11, PVP^{+}13, RBD^{+}13, SHPR13, VMGB^{+}10, BPN^{+}16, BMC^{+}18, CTBR19, GUHE15, ITG^{+}16, KNB^{+}17, KMT^{+}17, ML15, PBR^{+}16, VGAB^{+}19, WFMW^{+}18, WHEN^{+}16]. Scales [KMGCR10, LDC^{+}13, WB10, HH19b, SBU^{+}17]. Scaling [KAM18, Pri10]. Scandinavian [EBP^{+}18]. scarce [TDCG15]. Scat [VSM^{+}11]. Scat-Detection [VSM^{+}11]. Scavengers [OTKE12]. SCB [PMD^{+}13]. Scenarios [RV15, SML^{+}13, JBB^{+}15, KGF^{+}18]. schauinslandi [SBT^{+}11]. Scheduling [KCDC10]. Schemes [PMS13, BDKS15, DSSW16]. Schmidt [MSSM10]. scholarship [SMLB19]. School [Lan11, WD10]. Science [APR^{+}14, Ben12, BC14a, BCC^{+}14, CG13, CMS^{+}13, Cor16, Cor19, DSW16, Don12, DW19, FBB^{+}14, FLW14a, Gor11, HF10, HFL^{+}11, HS10, IP11, Jew13, JGH^{+}11, KTF^{+}14, LGM^{+}13, MBF^{+}12, MAB^{+}13, MQ14, MT10, Mol19, MB14, Mor12, PMMG14, RGMR12, Rey11, RCR^{+}12, TW14, Tru18b, WMS^{+}12, ABMG19, Agg15, AGS19, BPY^{+}16, BFG^{+}16, BPN^{+}16, BW19, CL16, CLvK^{+}17, CS17b, DMB^{+}17, DAA^{+}19, FMV^{+}19, HPD16, HPBP16, IBT18, JGS^{+}16, LPW16, MCD^{+}16, PMN^{+}16, PFN19, PWP^{+}18, RFV^{+}17, RIBT18, RLL15, Ros15, SJH19, SKLP16, TAB^{+}18, TRPD16, TLT16, VLSOS16, WDV^{+}16, WBPL15, WG15, ZAP^{+}18, vdWSM^{+}16]. Science-Based [BCC^{+}14]. sciences [BRK^{+}17, Wal17]. Scientific [CRNR12, Car14, Str13, TA10, VLRB12, Abr15, CHG^{+}17, RTB^{+}18, WDS15]. Scientist [KUT11, Rud11b, MSC^{+}18]. Scientists [Fra11, GZD11, MB14, PSD^{+}14, PWKT18, WLD16, Wil17]. Scoring [NKB10]. Scrub [CFBL10]. Scrub-Jays [CFBL10]. Scrubland [BTMF13]. Se [BBBG10, BE10]. Sea [BPW12, Ban12, CG13, CTM^{+}14, CCM12, DWCW10, ETB10, HRSK12, HFK^{+}10, IFP14, LAA^{+}10b, LMCS17, SHWT14, VLRB12, HN17, LSK^{+}15, LMC^{+}14, LAA^{+}10b]. Sea-Level [HRSK12, IFP14, HN17]. Sea-Surface [BPW12]. Seabird [BMA^{+}14, BJMT14, HRSK12, LMC^{+}14, RHR^{+}17, MDSF19, SBU^{+}17]. Seabirds [SNH^{+}14]. Seafood [HJSF^{+}12, WSC^{+}17]. seagrass [BBH17, CFB^{+}17, JRWA15]. seahorses [ZV19]. Seal [Bai13, SBT^{+}11]. Seals [AGAB11, BFB^{+}19]. search [DKL^{+}15]. Searching [KLS^{+}19]. Seas [Tru18b, MRMH17]. Seascape [OCP^{+}14, OGS^{+}19]. Season [Rog11, WK10]. Seasonal [PGC^{+}13b, ERW18]. Seaweed [HRKMG12]. Seco [Pra10]. Secondary [GH10, WD10]. Secos [BEE^{+}11]. Section [Zha15, DSW16]. sectoral [WZWF18]. Secundarias [WD10]. Securing [MMM^{+}12]. Security [HSBJ12]. Seddon [OSF11]. Sedimentary [BME^{+}14]. Seed [GZA10, HG10, LKYL15, MSF^{+}12, MC14, CTRE^{+}17, GBD^{+}15]. seed-dispersal [CTRE^{+}17]. Seed-Dispersing [MC14]. Seeding [BSF^{+}14]. Seedling [HG10]. Seeing [WD12]. Seeking [Fer19, Sou14a]. segmentation [JKJ^{+}19]. Seguridad [HSBJ12]. Seine [GOR^{+}12]. Selección [BGAS10, CLV^{+}10, HM11, JLDJ12, Boy10, VMS^{+}12]. Select [VPS^{+}14]. Selecting [AHP^{+}15, RNM^{+}18, VMS^{+}12, KTBH17]. Selection [BGAS10, BKKM14, JLDJ12, MBL^{+}14, PBH^{+}10b, SPN^{+}14, BSH^{+}15, FWH19, HFBG18, HJH16, VNAT19, WDW^{+}19]. selection-based [BSH^{+}15]. Selectiva [LMM^{+}10]. Selective [AEO^{+}13, LMM^{+}10, CHWH18]. Selectividad [FP10]. Selectivity [BVC^{+}14, Boy10, FP10]. self [GLK^{+}17, IF17, RHPF15]. self-organizing [RHPF15]. self-reported [GLK^{+}17]. self-sabotage [IF17]. Semi [EMD^{+}10]. Semi-Open [EMD^{+}10]. Semiabiertos [EMD^{+}10]. Semillas [GZA10, HG10]. Señal [FP10]. Señalización [BAPP11]. Sendje [GFRK12]. sense [GLK^{+}17]. Sensibilidad [BE10, DDPC12]. sensing [RBE^{+}15]. sensitivities [HLE19]. Sensitivity [BE10, DDPC12, SMTW13, ALA17, BGH^{+}15, RHPF15, TRPD16]. sensor [WWWC15]. Separación [BFHP10]. Separando [GZR12]. September [Ano11b, Ano14a, Ano18b]. sequencing [EBP^{+}18, PWKB19]. Sequestration [RM14b]. Serengeti [FGH^{+}13, FHK^{+}15, HMB^{+}15, NBNMG13]. series [SBR^{+}19]. Serious [NFdC^{+}12]. Serpents [Tho10]. Serpientes [CBSZ10]. Serrouya [TW11].

Service [Rey11, JRWA15, LMCS17, PVO^{+}16]. Services [AT11, AR10, CLV^{+}10, DVW^{+}13, DEJM10, ERC^{+}10, Fer11, GPK^{+}10, KAV18, MT10, NFF^{+}10, PHFPP10, RGMR12, SAGSG^{+}14, SLH10, SMGRJ10, AHP^{+}15, CZS19, CPR^{+}16, CGK^{+}19, CCB^{+}19a, CMG15, GKA^{+}15, HCSM15, SL16, TJJB18, TVY^{+}16, VKM^{+}17, WGS^{+}19]. Servicios [AT11, CLV^{+}10, DEJM10, ERC^{+}10, GPK^{+}10, NFF^{+}10, PHFPP10, SMGRJ10]. Sesgo [OB11b, CW10]. Sésiles [LGH^{+}12]. Sessile [LGH^{+}12]. Set [APL^{+}13]. Sets [UDG10, VMÁR^{+}13]. Setting [AdBR15, CCM12, DAKN^{+}17, DLMAG15, GKP13, HSV^{+}17, Kar14, OPP14, PTN10, RSWM14, JC18, PS15]. Settings [ACS14]. Settlement [BV10]. Seventy [PFA^{+}14b]. Seventy-One [PFA^{+}14b]. Several [MBH14]. severity [KRL^{+}16]. Sex [CWS^{+}14, HFK^{+}10, WES^{+}13, WGR^{+}10, GLdPL18]. Sex-Biased [CWS^{+}14]. Sex-Limited [WGR^{+}10]. Sexo [WGR^{+}10]. Sexos [HFK^{+}10]. Sexual [BB10b]. Sexuales [BB10b]. Shade [RM14b]. Shady [Cla13a]. shale [MGA16]. shale-gas [MGA16]. Shallow [Bra15, JMSM13]. Shape [Cat18, CS17a]. Shaping [Bur13]. share [KHK^{+}19]. Shared [Sch14a, RTGR17, SPD18]. Sharing [Arm19, FGG^{+}19]. sharing-sparing [FGG^{+}19]. Shark [CHHR13, BFF^{+}16, FFS^{+}18, OSM^{+}17, SH16]. Sharks [CB13, CHHR13, HS10, MGG^{+}13, NBW^{+}12]. Sheep [JMW^{+}11]. shelf [RK17]. sheltering [LP16]. Shepherding [RBW14]. Shift [TSA^{+}14, WK10, MTD16, MBT^{+}17b]. Shifting [Bai13, BT14, Cla14, Sha10b, TBY^{+}10, WMFB15, NM18, NWW^{+}17]. Shifts [FMJP11a, MRPPM12, BMA^{+}18, CBM^{+}18, HSU^{+}16, RCSK15]. Shipping [NS18]. Ships [RMM^{+}13]. Shorebird [EG12, IFP14, SKBE15]. Shorebirds [WKF^{+}11, RB17]. Short [Dol10, JPT13]. Short-Term [JPT13]. Should [AA12, GHKK10, PI11, Abr15, PWKT18]. Shrub [CBM^{+}11, LTQ^{+}13]. Shrubland [MQAMWM12]. Shrubs [CBM^{+}11]. sibling [OJAV15]. Side [PCW14]. sideways [DPJM18]. sierrae [JMW^{+}11]. Sighting [Duf11, FB12, CBB18b]. Sightings [SSB^{+}12, SB14, HBTB15, MCO^{+}17]. Signal [FP10]. signaling [TMD^{+}17]. Significance [BM14, CD13]. significant [BCD^{+}16, ENW^{+}18]. Significativo [MAA^{+}10]. silver [PFN15a, WCR^{+}15]. Silvestre [AWKR^{+}11, BM10, FEC^{+}10, FC10, HMM10, KMGCR10, LW10, MJA^{+}11, RMGCR10, TM10]. Silvestres [CBM^{+}11, PMS^{+}10, Wit11]. similarity [CCCCN18, PRNV18]. Simonetti [TYA16]. Simulating [WGJ13]. Simultánea [EMD^{+}10]. Simultaneous [RNE^{+}19]. Simultaneous-count [RNE^{+}19]. Simultaneously [EMD^{+}10]. Sin [LGK^{+}11]. Since [BV10, LMJB^{+}19]. Síndrome [FCC^{+}11, HM11]. Singing [VMW13]. Single [MBH14, Ore12]. Singularidad [RDM10]. Sink [Rog11, WH13, HLS^{+}15, HLS^{+}18]. Sinojackia [ZYYH10]. Sistema [Han10, KMGCR10, NKB10, RMGCR10, SAC12, WEC^{+}11]. Sistemas [MCL12]. Sistemática [SFM^{+}11]. Sistémico [CCBF12]. Site [BKKM14, FWH19, KRS^{+}10, LHDC12, WDW^{+}19]. Site-Occupancy [KRS^{+}10]. Site-selection [FWH19]. Sites [BTBDS12, GT10, TGD10, AB15, EEB15, GDG19, PRNV18, SMF^{+}15].

Sitio [LHDC12]. Sitios [BTBDS12, GT10, KRS^{+}10, TGD10]. Situ [MS12, MVB^{+}12, ZYYH10, CCW^{+}16]. Six [GKP13]. Size [Ber12, BMA^{+}14, BRKP12, CWS^{+}14, CKKL10, HNS^{+}11, MAWF12, RM14a, RC12, RSE12, VMW13, WGR^{+}10, GRT^{+}18, MVVS15, RTH^{+}15, ZRM18]. Sizes [ANTBG12, Tep12, DSSW16, KAM18]. Skill [APL^{+}13]. Skills [BDG^{+}13]. Skroch [AR10]. Slaughter [HS10]. Slopes [CLV14]. Slow [PMK^{+}10, SFKC16]. Slow-Reproducing [PMK^{+}10]. Small [BTMF13, CKR^{+}13, DEJM10, MBH14, PMM^{+}14, VMGB^{+}10, WGJ13, AP17, KSH17, VGAB^{+}19, WA18]. small-bodied [KSH17]. small-island [WA18]. Small-Scale [CKR^{+}13, VMGB^{+}10, VGAB^{+}19]. Smallest [ZZW^{+}10]. Smallholder [RM14b, SSW19]. Smart [HHDM10]. Smith [Ano16x]. Smooth [VM14]. snail [RBH^{+}15]. Snails [LP10, UMK^{+}16]. Snake [CTM^{+}14, VM14]. Snow [LWY^{+}14, SBB^{+}14, WW16]. Sobre [CGCL11, FMF10, HM12, HG10, KTN12, SRS11, SHO^{+}10, ANTBG12, ALL^{+}12, AZCT12, BAPP11, BBP12, CBM^{+}11, CMRH10, CMJ^{+}11, CFBL10, DDPC12, DMZ10, DWM^{+}10, EMS^{+}11, EPD12, EG12, FCGJH11, FG12, FEC^{+}10, FLP^{+}11, GT10, GS11, GSW10, HSBE12, IUM10, JMW^{+}11, KCSP11, KPL^{+}12, LSBB10, LGH^{+}12, LB11, LGB^{+}10, LCG10, LPT12, MLS12, MCN10, Mcc10, MBGS11, MC11a, MRRK11, MMCUR^{+}12, OTKE12, OB11b, PBK^{+}11, PTN10, PAM^{+}10, PMS^{+}10, Pra10, PMK^{+}10, PCBD11, RM11, RSW11, RMGCR10, RC12, Rud11a, Rud11b, SS12, SMGRJ10, SAGSHL10, SKCS11, TAG^{+}10, TTSZH10, VMGB^{+}10, WIM^{+}10, WSHB11, WFM^{+}10, WK10, WGB^{+}11b, Wit11, ZBD11]. Sobrepesca [MAA^{+}10]. Socavan [CND10]. Social [BB11a, BM14, BCT^{+}14, Bro10, BRCK11, CSMG^{+}13, CKK^{+}14, CCBF12, DSG^{+}17, DTM16, DRA^{+}15, GZR18, GMCW13, HBR^{+}14, Joh11, KCDC10, LRA^{+}14, MCL12, MB14, RB17, SABV13, SAGSG^{+}14, SM14, WSHB11, WKI^{+}14, AC16, BRK^{+}17, Dre17, EAPB15, GFJC19, GW17, HLHR16, KHB18, KTBH17, KF18a, KF18b, LMCS17, MBT^{+}17a, MFHT19, OHHE16, SMV19, SG19, SAB^{+}16, TAB^{+}18, TMG^{+}18, Wal17, WBF^{+}19]. Social-Ecological [MCL12, AC16]. Social-Network [GMCW13, GFJC19]. Sociales [Bro10, BRCK11, KCDC10]. Societies [Joh10a, PMD^{+}13, RTB^{+}18]. Society [Car14, MQ14, TW14, FBB^{+}14]. Socio [HKFC15, MCL12, MVVS15]. Socio-economic [HKFC15, MVVS15]. socioeconomic [MLMP^{+}15]. Sodhi [BLG^{+}11]. Soft [CCBF12]. soil [FSSW19]. Solar [HBE^{+}10]. Solares [HBE^{+}10]. Solitario [SS12]. Solitary [SS12]. Solomon [AS10]. Soluciones [BRKP12]. Solutions [BRKP12, Nat18, KLS^{+}19]. solve [SPD18, WHEN^{+}16]. Solving [EG13, Sny15]. Sombrilla [BR11b]. Some [LH10]. son [CWH^{+}10]. Sonar [BBWK^{+}14]. Song [Møl17]. Songbird [VMW13]. Songbirds [LB11]. Sonidos [ESCF12]. Sonora [FMF10, RRV10]. Sonoran [WON^{+}13, WLCW16]. Sons [Ano19w]. Soporte [PCTW10]. Sostenible [LGH^{+}12]. Sotobosque [Pra10]. Sound [KB11, LMCS17]. Sounds [ESCF12]. soundscapes [BTB^{+}18]. Source [BBWK^{+}14, Rog11, WH13, CCS^{+}18, HJH16, HLS^{+}18, HWB^{+}16, WFSS19]. Source-Sink [WH13]. sources [MCH^{+}16, WHG^{+}15]. South [ERC^{+}10, HRC10, PCTW10, PRE^{+}13, BFF^{+}16, DPS^{+}16, GGL^{+}15, LKB^{+}14, RMKM14]. southeast [DWM^{+}19, CPR^{+}16, DUS^{+}16, FTB^{+}16, GHT^{+}15, HSB^{+}16, JCCC14, ZFW^{+}11]. Southeastern [AS13, CLRV11, RGRMPB^{+}13, HKFC15, LO10]. Southern [Fie15, DRA^{+}15, LPT12, CS13, LK14, LP13, MMPF10, ÖB12]. Southwest [Fie15]. southwestern [BCSM15, LCC^{+}15, SLWL15]. Southwood [Ano19w]. Soviet [BIP^{+}15]. Space [HCP^{+}12, AWHW17, KHGF^{+}17]. Spain [JIAMS^{+}14, ZMTR15]. Spanish [MSZB^{+}14, MRPPM12, SLGB15]. Sparing [CKR^{+}13, FGG^{+}19]. Sparrow [HD14]. Spatial [ALL^{+}12, CBM^{+}11, CJDZ10, CGCL11, FWM14, FLP^{+}11, HSB12, KMTK13, KGS^{+}11, LMC^{+}14, LRA^{+}14, MVVS15, MGW^{+}15, PSP11, PBB^{+}11, RMM^{+}13, SSK^{+}11, VMÁR^{+}13, WWH^{+}18, BJ16, BTTP17, DKG^{+}15, DCP^{+}19, FSR^{+}18, GMV^{+}18, HTE16, NSBJ17a, NBH^{+}17, NSBJ17b, RSN19, SMARC^{+}16, SBU^{+}17, VKM^{+}17]. Spatially [BRCK11, EIS14, HRBR11, RSPD13, SA14, Dre17, JHAH18, MPS^{+}16]. spatio [MKR^{+}15]. spatio-temporal [MKR^{+}15]. Spatiotemporal [CPD^{+}15, KGG^{+}14, SMZZ17]. Spawning [WK10, BTTP17]. spawning-aggregation [BTTP17]. Spear [FCGJH11]. Spear-Fishing [FCGJH11]. spearfishing [YFB15]. Special [Zha15, DSW16]. specialist [CWH^{+}17, RHPF15]. Specialists [SKSK14]. specialization [SC15]. specially [HICF16]. Specie [BRV10, Chi10, FMJP11a, RTT^{+}13, EEB15, HTH^{+}17b, Pet17, PNSG18, San19, WHK^{+}17]. Species [AG10, AGP^{+}14, AZT^{+}14, BHK10, BA13, BCC^{+}14, BLC13, Boo14, BWE^{+}13, BR11b, BHM^{+}14, BDH^{+}12, BGG13, CLT^{+}13, CVN^{+}10, CMRH10, CLRV11, CCM12, CCN16, Chi10, CKKL10, Cla14, CD13, CMNM10, DWCW10, DWM^{+}10, ENW^{+}18, EG12, ETB10, FFM^{+}14, Fat14, FFS^{+}18, FG12, GAG13, Hal19, HNS^{+}11, HMIE14, HSR^{+}13, HGCT13, HJSF^{+}12, JBB^{+}14, JMP^{+}11, JHV13, KG10, LN11, Les15, LRMWG16, LLW^{+}14, LWLC12, LCG10, LP13, MRBL^{+}12, MFRH12, MAB^{+}13, MRMV14, MVB^{+}12, MWC11, MSSM09, NCC13, NBSM12, OKLP11, OCP^{+}14, OSFFB^{+}11, PBH^{+}10a, PSAZ11, Par10, PAM^{+}10, PZM14, PS10a, PF10, RTT^{+}13, RCK13, RER10, RHM10, SASO13, San10, San14, SPN^{+}14, SSO11, SSO12, Sch10, SRL^{+}13, SBJG13, SAC12, SS11, SHCD11, SLSH14, SKCS11, SAB12, TAG^{+}10, TBB^{+}14, TG13, TGD10, TA10, TG15, TSA^{+}14]. Species [TCP13, TBY^{+}10, VPS^{+}14, VFV^{+}12, VLRB12, VSM^{+}11, WL12, WEC^{+}11, WGJ13, Wil17, WHB^{+}11, WSEM13, ZCCN13, ZYYH10, ABG^{+}19, ABB^{+}18, AB15, BGO^{+}19, BdA15, BSH^{+}15, BCOB15, BBK^{+}17, BHAR^{+}18, BH18, BBF19, BC18, CMMP15, CBPM19, Cla15, CGE19, CCS^{+}18, DMB^{+}17, DSSW16, DCP^{+}19, EB15, EAPB15, EWCG15, GRL15, GBMW15, GUHE15, GDG19, GMG^{+}19, GSSF19, HNO^{+}18, HKB^{+}17, HTE16, HH19c, HHW^{+}16, ITG^{+}16, ISA16, JMH15, JHAH18, KCD^{+}17, LDD19, LH19, LZR^{+}18, LKYL15, LMM^{+}17, LMJK^{+}19, LBB^{+}16, ME16, MBBB17, MDD^{+}18, MSL19, NJBM15, NCKC17, NW15, OQGR^{+}15, PA18, PTJ^{+}19, QSCM^{+}19, RSN19, RWF^{+}17, RK17, RTH^{+}15, SC19, SWSM19, SLSH15, SGA^{+}19, SSE^{+}17, TK15, TPL18, TIB^{+}16, TRPD16, TBF^{+}17, TMK^{+}16, VM18, VJG^{+}15, WÖ16, WDW^{+}19, WTT^{+}17, WHE^{+}17, WGLL17, ZAB16, Ano19w]. Species-Area [KG10]. Species-Level [NBSM12, CCN16]. Specific [ANTBG12, KAV18, GPB^{+}15]. Specification [BGL^{+}14]. Specimen [GCC^{+}12]. Specimen-Based [GCC^{+}12]. Spending [HSB12, MR16]. Speziale [TK15]. Spiders [GZR12]. Spies [HODB10]. spillover [AC16]. Spinning [Hal16]. Split [BFHP10]. sponges [BMB^{+}15]. Spontaneous [ZYYH10]. Sport [JVT^{+}13]. Sport-Hunting [JVT^{+}13]. Spotted [HODB10, RC12, SMB^{+}10, TG13]. spp [PB11]. Spread [RL11]. Spring [LP10, JVI^{+}18]. squandering [TDCG15]. squeezing [CBM16]. Stability [SAGSHL10]. Stabilization [JMSM13]. Stage [AEO^{+}13, ANTBG12, GBB^{+}15, HGGH15, LAA^{+}15]. Stage-Specific [ANTBG12]. Stages [GLNJ10, WK10]. Stakeholder [FTSSW12, MT10, GPB^{+}15, MV18, MSR19]. Stakeholder-Driven [MT10]. stakeholder-specific [GPB^{+}15]. Stakeholders [GZD11, LL14, VPMS13]. Stambaugh [Mat15]. Standardized [ISM^{+}18]. Standards [HDH^{+}10, MAB^{+}15, VTT^{+}17]. Start [Löv10]. startup [Bus15]. State [LW14, TWB^{+}18]. Stated [CD13]. statements [GLJ^{+}16]. States [FEC^{+}10, SK10, WA18, BCA^{+}18, BCSM15, LO10, LN11, LWZL16, MC11a, ML10, RPA^{+}12]. Status [BJR^{+}11, CCBP14, CC14, FMJP11b, GRA14, HWW12, LJA^{+}13, RDM10, Rud11b, SPPN^{+}12, WHB^{+}11, BMB^{+}15, BBM18, BCOB15, BCSM15, CCN16, LSK^{+}15, SSW15, SKSS^{+}17]. statutory [Wil17]. Steelhead [BP13]. Steen [BCA^{+}19]. steps [LFMB19]. Stewardship [AB13a, Sto12, Wat11, MCD^{+}16]. stochastic [ALW^{+}19]. Stock [BJR^{+}11]. Stopover [BKKM14, HBTB15]. Stopping [GCC^{+}12]. Stories [LGM^{+}13, RGMR12]. Straight [Kar14]. Strategic [AGP^{+}14, CWAW14, RWF^{+}17]. Strategically [EBRS14]. Strategies [AH11, BW13, BRCK11, DBV13, GFRK12, GSDE14, KTF^{+}14, LH11, MBH14, MR12, WSEM13, DPS^{+}16, GBB^{+}15, TMK^{+}16]. Strategy [GZNC13, SA14]. Stratification [MPB10]. Stream [AGLRV14, LKB^{+}14, PGC^{+}13b, PCBD11, BHPM^{+}18, CBCW18, LDD19, SJZ19]. Stream-Associated [PGC^{+}13b]. Stream-Dwelling [AGLRV14]. Strength [FP10, KRL^{+}16]. Strengthen [Fli10]. Stress [AME^{+}13, ATW^{+}11, JLDJ12, ML14, TMD^{+}17]. Stress-Hormone [JLDJ12]. Stresses [GKK^{+}11]. Stressors [RP13, WH13, HLE19]. stretching [CBM16]. Striking [RMM^{+}13]. Striped [MAWF12]. Structure [CM13, EBL^{+}14, FMF10, FLP^{+}11, HSAB10, SHA^{+}10a, SHCD11, BCA^{+}18, GGL^{+}15, HAT^{+}15, HLHR16, IRK^{+}18, LS19, THD^{+}18, Wal17]. Structured [LWLC12, MR12, WCR^{+}15]. Structures [BCT^{+}14, FC10, SCK13, RBK^{+}17]. Structuring [GAG13, Lin11]. Student [BDG^{+}13]. Students [KTF^{+}14, PGBV^{+}13]. Studied [BE10, AHT^{+}15]. Studies [HHDM10, ACB^{+}15, DSE15, FWH19, KS17]. Study [DWCW10, EMD^{+}10, GBMRG^{+}10, LK14, MAA^{+}10, TA10, WHB^{+}11, Agg15, CLM^{+}15, GV16, MF15, VTT^{+}17]. Studying [JGS^{+}16]. Sturgeon [BBB14, DBP^{+}10, FHBK18]. Subcontinent [MSY^{+}15]. Subestimado [CGM10]. Subject [OB11b]. Subjective [Die16]. Submarines [Mor12]. Submarinos [FCGJH11]. Submission [OSC10]. subnivium [ZP18]. Subsistence [GBB^{+}14, VCDF13, SMARC^{+}16, SMY^{+}17]. substitutes [MAJ16]. Substitution [GH10, MAG18]. Subterranean [Bra15]. Subtropical [KZG^{+}13, PB11]. succeed [LGA17]. Success [BCT^{+}14, CWS^{+}14, FCGJH11, FWM14, Han11, HM12, MS12, Pel15, VMS^{+}12, WFM^{+}10, ABB^{+}18, BCM16, CY17, CBWFL15, CCS^{+}16, IPA^{+}17, JMB16, ZRM18]. Successes [BMG^{+}14]. successful [AAS^{+}17, MSC^{+}18, NWD^{+}16b]. Sucedáneas [MWC11]. Sudáfrica [ERC^{+}10, HRC10]. Suelo [BRKP12, EMD^{+}10, JMP^{+}11, KG10, LB11, MC11a, RPA^{+}12, SPPN^{+}12, SHO^{+}10, WIM^{+}10]. Suitable [BLC13]. Sulawesi [WIM^{+}10]. Sumatran [NGB^{+}16]. Sumidero [Rog11]. Suministro [DEJM10]. summarizing [PNSG18]. Summoning [WBB^{+}18]. Sundabarns [BGAS10]. Sundarbans [BGAS10]. superabundant [KOA^{+}15]. Superficie [BPW12]. Supervivencia [HSBE12, KPL^{+}12, LB11, PMS^{+}10, WSHB11, ZBD11]. supplementary [EWCG15]. Suppliers [GZA10]. Supply [PCW14]. Supply-Side [PCW14]. Support [PCTW10, SSD10, FRY^{+}17, KF18a, MWL^{+}15, VTM19]. suppression [ABG^{+}19]. Sureste [CLRV11, LO10]. Surface [BPW12, MGA16]. Surrogate [CMNM10, LNP10, MWC11, OCP^{+}14, BdA15, SURW18]. Surrogates [DFW^{+}10, MFKMP10, ODTS10, BSH^{+}15, HCH17, IO17, JVI^{+}18, SKPP15]. surrounding [BBH17]. Surveillance [GAFG14, NJBM15, SJD^{+}19]. survey [BTBC^{+}16, BHAR^{+}18, FFS^{+}18]. Surveys [BM10, DWCW10, FCGJH11, LS15]. Survival [HSBE12, KPL^{+}12, LB11, MSSM10, PMS^{+}10, WSHB11, ZBD11, BSL^{+}19, DSK^{+}19, SSP^{+}15, WLCW16]. Survive [Lub16]. Sus [SRS11, MMCUR^{+}12]. Susceptibilidad [VJMH12, SGH^{+}11]. Susceptibility [SGH^{+}11, VJMH12]. Suspected [ZMTR15]. sustain [HGRS16]. Sustainability [AWKR^{+}11, BR11a, CB13, KMGCR10, Lor13, MRRK11, SMY^{+}17, VMGB^{+}10, WT13, BHB^{+}19, DWCC17, GLJ^{+}16, MBBB17, MAB^{+}15]. Sustainable [CRD^{+}14, DFJ^{+}12, KH16, LKB^{+}10, McC14, MO11, Per11, Ric11, MLMP^{+}15, NBA^{+}18]. Sustainable-Use [Per11]. Sustainably [HBM^{+}12]. Sustained [WLCS14, WW16]. Sustentabilidad [AWKR^{+}11, BR11a, KMGCR10, MRRK11, VMGB^{+}10]. Sustentablemente [HBM^{+}12]. Sustentables [MO11]. Sustituta [CMNM10]. Sustitutos [DFW^{+}10, LNP10, ODTS10, MFKMP10]. swans [CRMK18]. Swedish [LJA^{+}13]. Swidden [ZFW^{+}11]. Swift [CM13]. Syndrome [FCC^{+}11, HM11, DMR^{+}17, ISA16, MF15]. Synergies [MBV^{+}16]. Synthesis [Fur18, CAB^{+}18]. System [CSMG^{+}13, GFGR12, Han10, KMGCR10, NKB10, RMGCR10, SAC12, WEC^{+}11, BMAS17, GDB15, MKR^{+}15].

Systematic [CPF13, PFR^{+}14, SFM^{+}11, TTEM14, VPS^{+}14, BTBC^{+}16, BPH^{+}19, CY17, HWMC15, HW19, VSA^{+}15, WL17b]. systematics [LSV^{+}19]. Systemic [NMSF11]. Systems [BRM^{+}14, CCBF12, HFSB13, MCL12, SKPP15, WTT^{+}17].

T [Ano19w]. T. [Bra15]. Tadarida [ATW^{+}11]. Tadpole [MAWF12]. tags [WWWC15]. Tailed [ATW^{+}11, AB13b]. Takahē [GAFG14]. Takahe [GGB^{+}15, GLNJ10]. Take [LGK^{+}11, WWdS^{+}16]. Taking [WD10]. Tala [CMRH10]. Talk [Joh10b]. Tamaño [Ber12, BRKP12, CKKL10, HNS^{+}11, RC12, RSE12, WGR^{+}10]. Tamaños [ANTBG12]. Tan [BBBG10]. Tanzania [PBK^{+}11, BL17, CD16, HBR^{+}14, MBJWC10, NFF^{+}10, NJT14, PBK^{+}11, SMK14]. Tanzanian [MBJWC10]. Target [JCA^{+}11, MJG15, ZBM^{+}10]. Targeted [EBRS14, KP16, KP19]. Targeting [CLV^{+}10, SAGSG^{+}14, SA14, HBTMG15, HWB^{+}16, JKJ^{+}19]. Targets [CCM12, GAFG14, LCT^{+}14, PTN10, GBB^{+}19b, HSV^{+}17, MBV^{+}16, MGMC15, TUHK^{+}17]. Tasa [WKF^{+}11, OB11b]. Tasas [ANTBG12, JMW^{+}11, SRS11, SDB10]. Task [Fle10]. Tasmania [HLB^{+}12, LMM^{+}10, MCO^{+}17]. Tasmanian [HLB^{+}12, HJMM14, LMM^{+}10, RBHW18]. Taxa [Duf11, LNP10, AHT^{+}15, SURW18, SSE^{+}17]. Taxon [GH10]. Taxonomic [BVC^{+}14, DMM14, EMJ14, LN11, SPN^{+}14, JVI^{+}18, QCW^{+}18, SGA^{+}19]. taxonomically [LSV^{+}19]. Taxonómicos [LN11]. taxonomy [CVA15]. Teaching [CND10]. Technique [Jew13, PS11, CCG18]. technologies [PWKB19]. technology [BC15]. teeth [BFB^{+}19]. Tellería [PS16]. Temático [OB11b]. Temperate [AS13, GZA10, IUM10, MLS12, CE15, DPS^{+}16]. Temperatura [BPW12, RTL^{+}12, SM11]. Temperature [BPW12, Cat11, LP13, SM11, WES^{+}13, ZRM18]. Templados [GZA10, IUM10, MLS12]. Temporal [CM13, PFR^{+}14, MKR^{+}15]. temporally [ITG^{+}16]. Temprano [Col10]. Ten [JWS^{+}18, RBE^{+}15]. Ten-year [JWS^{+}18]. Tendencias [BJR^{+}11, KRS^{+}10, PSP11, GBMRG^{+}10]. Tendidos [BAPP11]. Tent [SLH10]. Teoría [GHKK10]. Tepedino [LDC^{+}15]. Term [CW10, FZ18, GGG10, HAP^{+}13, JPT13, TTSZH10, BBM18, BHF^{+}19, CWS16, CMM17, FWH19, MKR^{+}15, ML15, SBU^{+}17, SA15, VMO^{+}18, WFMW^{+}18, WC19]. terminology [FSAM15]. Términos [AT11]. Terms [AT11]. Terns [MPH^{+}11]. Terratenientes [HRC10]. Terrenos [MC11b]. Terrestres [HMD^{+}10]. Terrestrial [HMD^{+}10, PG12, Sou10, CEdK18, MAM^{+}15, PTJ^{+}19, SSW15, TBM^{+}16]. Territorial [GFG^{+}12, TG13, GD15, OKCG18]. Territoriales [BB10a]. territories [DOC^{+}15, GH18]. Test [CMRH10, EG12, Rog11, CBB18b, HLS^{+}18, KGF^{+}18, MAG18]. Testing [CFB^{+}17, HAF^{+}17, PAM^{+}10, RTT^{+}13]. tests [BSH^{+}15, MSY^{+}15]. Tetrápodos [Fli10]. Tetrapods [Fli10]. Text [WBPL15, RCBT18]. Texts [Sto11]. Thailand [CTM^{+}14]. Their [BGG13, GRA14, JDA^{+}10, LA13, MGG^{+}13, O’B11a, RGRMPB^{+}13, RMKM14, SRS11, GDG19, PA18, PWKT18, RLB^{+}17]. Them [KE10, Nas17, Sha15b, CBB18b, Gru17]. Theme [Ore12]. Themes [Val18]. Theories [Laz11]. Theory [GL18, GHKK10, MMK^{+}14, ARDG17, BCR^{+}17, Cha15, DAA^{+}19, GMG^{+}19, HCSM15, LAA^{+}15]. There [McC14]. Thermal [CLV14, FMJP11b, NWW^{+}17]. Think [PI11]. Thinking [CCBF12, CP14, NDB^{+}12a]. Thomas [Hal19]. Threat [AGLRV14, FP10, FGH^{+}13, HSR^{+}13, MSZB^{+}14, MA10, NFdC^{+}12, RDM10, AWT^{+}15, BCM16, CNvL^{+}15, MGW^{+}15, NRRW17, QKC^{+}19, ROQR^{+}16, TMK^{+}16, ZV19]. threat-mitigation [TMK^{+}16]. Threatened [AWU^{+}11, Boo14, BP13, BGG13, DWM^{+}10, FFM^{+}14, FLP^{+}11, GAG13, KTN12, LN11, MARV^{+}11, MYG^{+}14, MVB^{+}11, MVB^{+}12, MSSM09, NCC13, PCRD14, RRCB^{+}11, RJF^{+}13, SHG^{+}14, Sch10, SNH^{+}14, WL12, WEC^{+}11, Arm18, BGD^{+}15, CMMP15, DSK^{+}19, DRF^{+}18, GBMW15, HNO^{+}18, HH19c, LH19, LZR^{+}18, LRL^{+}15, LBB^{+}16, PBR^{+}16, RBK^{+}17, TLPP17, TBF^{+}17, VESHR18]. Threatened-Species [DWM^{+}10]. threatening [TMK^{+}16]. Threats [BHSA12, DBV13, DWCW10, FB12, PS10a, RLB^{+}17, SSRS18, GBB^{+}19a, KAM18, KJWP19, PA18, SMM16, SA15, TGBB^{+}15, TG15, TBC18]. Three [GOR^{+}12, Ore12, RPS13, SG10, WGLL17]. Threshold [BHPM^{+}18]. Thresholds [CC14, OQGR^{+}15, SAB12, AdBR15, CFB^{+}17, MBBH16]. Throughout [Rog11]. thrush [SMF^{+}15]. thumb [Bei19]. thylacine [BSCB18, CBB18a, CBB18b]. Tibetan [LWY^{+}14, SLWL15]. Tiburones [HS10, NBW^{+}12]. Tidal [BGG13]. Tiempo [Boy10]. Tierra [LB11]. Tierras [AH12, FMF10, KCDC10, RWB11, RRV10]. Tiger [CD13, DUS^{+}16]. Tigers [KE10, Med10, CNBW16]. Tigres [KE10]. Tímalo [WK10]. Timber [CBSZ10, CS13]. Time [Arm18, Boy10, Ehr14, GL18, HLC13, RM14a, AOPM^{+}19, GRL15, HW19, HTE16, LAA^{+}15, RDS17, SBR^{+}19, WHEN^{+}16]. Time-Lagged [HLC13]. time-scale [WHEN^{+}16]. timing [BHPM^{+}18]. Timor [LMC^{+}14]. Tipos [FP10]. Toads [GG13, PMS^{+}10, SMM16]. TOC [Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano18p, Ano18q, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q]. today [LMM^{+}17]. Todo [GHKK10]. Tolerance [FMJP11b, KO12, VMS^{+}12, WLL^{+}16]. Tolerancia [VMS^{+}12]. Toma [MR12, PBB^{+}11, SSD10, HBM^{+}12]. Tomados [KUT11]. Tool [AOLK11, BME^{+}14, HCLW13, KB11, Mat13, RMGCR10, BHAR^{+}18, KLN^{+}17]. toolbox [MH16]. Tools [WRTB10, ENW^{+}18, VTM19, WBPL15]. topics [WBPL15]. topography [TBF^{+}17]. Tortoise [GZJ^{+}13, QSCM^{+}19]. Tortoises [HGCT13]. Tortuga [WGR^{+}10]. Tortugas [BBBG10, DWCW10, HFK^{+}10]. Tour [DDPC12]. Tourism [LW11, CC18, DPJM18, NWD^{+}16a]. Tracing [MSY^{+}15]. track [WFC^{+}18, WSC^{+}17]. Tracking [KMLC19, CNK^{+}17, SMF^{+}15]. Tradable [WW10b]. Trade [BM10, BBM14, BRKP12, CKKL10, DHG^{+}17, FMB^{+}18, HSBJ12, HAB^{+}11, MG12, NGL^{+}14, SASO13, WB14, Ben15a, BCR^{+}17, CGK^{+}19, Cha15, Dre17, FFS^{+}18, Har15, HTH^{+}17a, HGM19, HRHC16, HLHR16, HHKK15, LBI^{+}18, LFMB19, MSMCB18, MBV^{+}16, MKR^{+}15, MLMP^{+}15, MFHT19, MSR19, NJBM15, NRRW17, NBA^{+}18, RS18, SMARC^{+}16, SBS^{+}15, VGAB^{+}19, HAB^{+}11]. trade-off [Dre17]. Trade-Offs [BRKP12, CKKL10, HSBJ12, HAB^{+}11, MG12, DHG^{+}17, WB14, CGK^{+}19, Cha15, LBI^{+}18, MSMCB18, MBV^{+}16, VGAB^{+}19, HAB^{+}11]. Traded [FFM^{+}14, MSY^{+}15]. traders [HGM19]. Tradicional [GBMRG^{+}10, WW10a]. Tradicionales [CND10]. Trading [NJT14]. Traditional [CND10, GBMRG^{+}10, Tru18b, WW10a, OJAV15]. Traducción [JMW^{+}11]. Traducir [HF10]. Traffic [GS11, SDB10, TMD^{+}17]. Tráfico [GS11, SDB10]. training [vdWSM^{+}16]. Trait [AWHW17, GRL15, GLB^{+}13, KSH17, GUHE15]. Trait-based [KSH17]. Traits [Ber12, FB12, FP10, KZG^{+}13, NCC13, WR12, DS19, GRT^{+}18, HLE19, HTH^{+}17b]. trajectory [SAS^{+}19]. Trampa [FC10, HSAB10, SHC^{+}11]. Trans [Cla13a]. Transactional [GFM^{+}13]. transboundary [GMV^{+}18, YJL^{+}18]. transcriptomics [HJH16]. Transdisciplinar [RRC^{+}10]. Transdisciplinary [RRC^{+}10]. Transfer [AG10, BC15]. Transferable [LGK^{+}11]. Transferibles [LGK^{+}11]. transformation [ML15]. transformations [CC18]. transformative [BPY^{+}16]. transformer [EB15]. Transforming [CLvK^{+}17, For11, ZFW^{+}11]. Transición [GBMRG^{+}10]. Transition [GBMRG^{+}10, PSA^{+}17]. Translate [HF10]. Translating [JMW^{+}11]. Translocación [AOLK11, DWH^{+}11, HSBE12, MARV^{+}11, SBT^{+}11, VMS^{+}12]. Translocaciones [SVH12, SS12]. Translocated [GAFG14, DS19, GLdPL18]. Translocation [AOLK11, DWH^{+}11, HSBE12, LRL^{+}15, MARV^{+}11, MVB^{+}12, SBT^{+}11, VMS^{+}12, DSK^{+}19, EEB15, SNK15]. Translocations [MBG14, SVH12, SS12, SJP^{+}13, BCM16, BCLMM19]. Transmisión [OTKE12]. Transmission [OTKE12]. transparency [MRFvW16, PMN^{+}16]. Trap [FC10, HSAB10, LDB^{+}10, LCL13, SHC^{+}11]. Trap-Neuter-Release [LCL13]. Trap-Neuter-Return [LDB^{+}10]. trapping [KOA^{+}15]. Traps [CC18, Møl17]. Trasferencia [AG10]. Trawl [HSBJ12]. Trawling [GFM^{+}13]. Treasure [Fur19b]. treatment [Cha15]. Tree [RM14b, TAG^{+}10, ZYYH10, EEB15, MSMCB18, TMD^{+}17]. Trees [BM14, HG10, SS10a]. Trematode [BAG^{+}11]. Trematodos [BAG^{+}11]. Trend [KRS^{+}10, PSP11]. Trends [BFF^{+}16, BJR^{+}11, CHHR13, GBMRG^{+}10, HAP^{+}13, MCO^{+}17, PMP14, VLGRS14, CPD^{+}15, DPGY16, DMB^{+}17, LBB^{+}16, RN17, YFB15, vWWCF15]. Tres [GOR^{+}12]. Trial [KWW^{+}14]. Trials [HCM14]. Triangle [VPMS13]. Tribute [Die10b, LBD^{+}19]. Triggers [NS12]. Tróficas [SGS^{+}10]. Trophic [EWN^{+}13, HJMM14, SGS^{+}10, TSA^{+}14, AK16, RBF15]. Trophy [BL17, BEA^{+}13, FWC^{+}15, HCLW13, PBK^{+}11, JD16]. Tropical [AEO^{+}13, APL^{+}13, AMO11, BME^{+}14, Cat11, CKR^{+}13, DMM14, FBA^{+}13, FMJP11b, GOR^{+}12, LW11, Nov16, NWW^{+}17, Per11, PGC^{+}13b, Pra10, Put11, RSS^{+}13, SWL^{+}10, VSM^{+}11, WIM^{+}10, BVAC17, BGZ^{+}15, BGD^{+}15, BTB^{+}18, CGE19, HSB^{+}16, MAB^{+}15, RLB^{+}17, SLS17, SFKC16, SSW19, SSE^{+}17, GOR^{+}12, WSLW16]. tropical-forest [SSE^{+}17]. Tropicales [AMO11, SWL^{+}10]. Tropics [NS18, SEM^{+}15]. Troubling [CNK^{+}17]. trout [OJAV15]. True [Mar14, LBD^{+}19]. truncation [SRG^{+}15]. Trust [AH12]. trusts [FFHA19]. Tuatara [MMM^{+}12]. Tumor [HLB^{+}12, LMM^{+}10]. Tumores [HLB^{+}12]. tuna [SRG^{+}15]. Tunas [GOR^{+}12]. Turísticas [LW11]. Turísticos [DDPC12]. Turkish [Pek13]. turnover [BH18]. Turtle [CG13, GSDE14, SKH13, WGR^{+}10, ÇRG^{+}17, SDT17, SR17]. Turtles [BBBG10, DWCW10, HFK^{+}10, SHWT14, BCA^{+}18, BCA^{+}19, LS19, WHG^{+}15]. Twitter [BFG^{+}16, PNC^{+}15, PSD^{+}14]. Two [AWU^{+}11, AEO^{+}13, Bar12, CCM12, Fie15, HDN^{+}14b, MTM12, MLA^{+}14, OSFFB^{+}11, Par10, Plo17, RLGG^{+}12, SAGSHL10, VCDF13, GRA14, HSU^{+}16]. Two-Stage [AEO^{+}13]. Types [FP10, GRA14]. Typology [JIAMS^{+}14, WHEN^{+}16].

U.K [Ano19w]. U.S. [BCC^{+}14, CVN^{+}10, ENW^{+}18, EIS14, FR16, FBB^{+}14, FFHA19, HCP^{+}12, Hur13, LW14, Les15, LRMWG16, MAB^{+}13, MF14, MGG^{+}13, O’B11a, PS10a, RTT^{+}13, SRL^{+}13, TG15, WT10, Wil17, ZCCN13]. U.S.A [Ano16x, Ano19w, Bra15, Hic10, LPT12, MWC11, WBM13, MPH^{+}11]. U.S.A. [CS13, HSBE12, RTL^{+}12]. Uganda [BMGLW12]. Ugandan [CPD^{+}15]. UICN [RRCB^{+}11]. UK [Fie15]. Ultraviolet [MAWF12, OSFFB^{+}11, TSH^{+}14]. Ultraviolet-B [MAWF12]. Ultravioleta [OSFFB^{+}11]. Umbrales [SAB12]. Umbrella [BR11b, BHM^{+}14, SPN^{+}14, KCD^{+}17, LP16]. Umbrella-Species [BR11b]. unauthorized [HBTMG15]. Uncertain [CWH^{+}10, Cla13a, Ibi17, SSB^{+}12, SB14, SBN^{+}10, PMM^{+}15]. Uncertainty [AG10, AH11, BP13, EMJ14, KKH14, LRA^{+}14, SMTW13, SCHM18, CCW^{+}16, DAKN^{+}17, HKB^{+}17, RHCW16, RCARR18, WHK^{+}17]. Uncommon [HWW12]. uncritical [Bai16]. Underaddressed [HH19a]. Underestimated [CGM10]. Underestimating [SMB^{+}10]. Underestimation [AGLRV14, RBH^{+}15]. underlying [LPA^{+}17]. Undermine [CND10]. underrepresented [SJH19]. Understand [Fur19a, Abr15, ZSM18]. Understanding [AB13a, BAGM13, GTVC16, HNS^{+}11, HSF17, Hug19, KTBH17, MBG14, MB14, NBA^{+}18, OBBP10, Qui14, SLA^{+}19, Ser10, TM10, TAM^{+}19, WGB^{+}11b, WHG^{+}15, BCA^{+}19, CS19, DJBB16, GBB^{+}19a, GMNO15]. Understandings [RGRMPB^{+}13]. Understory [AASK12, EWN^{+}13, MRBL^{+}12, Pra10, SGE13]. Underwater [FCGJH11, BMB^{+}19]. Undescribed [TBB^{+}14]. undetected [CGS^{+}16]. Unequal [WES^{+}13]. unexpected [PS16]. Unexploited [LCT^{+}14]. Unexplored [LFMRD12, FCS19]. Unfulfilled [PHJ^{+}17]. Ungulado [FC10]. Ungulados [MRRK11]. Ungulate [FC10, MRRK11]. Ungulates [MBM^{+}14, HDH^{+}15]. Unidad [RMGCR10]. Unidades [FB11]. Unidos [FEC^{+}10, LO10, LN11, ML10, RPA^{+}12]. Unintended [BW13, BT14, VWM^{+}10]. Unintentional [CRD^{+}14]. unintentionally [LH19]. Union [HHB15, SFAA^{+}18, BIP^{+}15]. unique [DRA^{+}15]. Unit [RMGCR10, MBFB18]. United [LO10, RPA^{+}12, BCA^{+}18, BCSM15, DOC^{+}15, FEC^{+}10, HGM19, LN11, LWZL16, MC11a, ML10, SK10]. Units [FB11]. Universal [SS10b]. University [Ano16x, Ano19w, Bra15, Fie15]. unknown [BGP16, Mat15]. Unmanaged [PBH^{+}10a, PBH^{+}10b]. unmarked [RNE^{+}19]. Unpacking [BL17]. unprotected [BARB17]. unrecognized [CCS^{+}18]. Unsustainably [LGH^{+}12]. Untangling [AG17, LPA^{+}17]. Untapped [VB19]. Upland [ROQR^{+}16, FBWIO19]. Upper [CLRV11, SMS14]. upstream [SJZ19]. Uptake [TTEM14]. Ural [DBP^{+}10]. Urban [AB13a, CFO^{+}14, DK10, LCL13, LDBS11, SPN^{+}14, Ven18, WW10a, BLM^{+}16, GLF^{+}18, SHSP19, SSC^{+}19]. Urbana [DK10]. Urbanas [LDBS11]. Urbanities [Cat18]. Urbanización [AS10, PCBD11, SDB10]. Urbanization [AS10, BSF^{+}14, PCBD11, SDB10, BCA^{+}18, BCA^{+}19, CBCW18, FSAM15, LS19]. Urbanized [MRT^{+}12, THD^{+}18]. Urbanos [WW10a]. urophasianus [BBP12]. Ursid [SCK13]. US$190.00 [Ano19w]. US$21.00 [Ano19w]. US$24.95 [Ano19w]. US$30.00 [Ano19w]. Use [AWU^{+}11, AS13, ALW^{+}19, AWKR^{+}11, BRV10, BGAS10, BB10a, BHSA12, BBM18, BG10, BME^{+}14, BRKP12, CRKM^{+}14, CMB12, CAB^{+}10, CMNM10, DFW^{+}10, Dar15, DBR13, DBV^{+}18, DEJM10, EBL^{+}14, GMBM13, GSB10, GS16, HMIE14, HBTB15, HJSF^{+}12, JRWA15, JLDJ12, JMP^{+}11, KMTK13, KGS^{+}11, KO12, KG10, LHDC12, LMJK^{+}19, Lyn13, MBH14, MPH^{+}11, Mat13, MC11a, MWL^{+}15, MWC11, OJAV15, Pek13, PVP^{+}13, Per11, RHT14, Ric11, Rog11, SSK^{+}11, SPPN^{+}12, SAC12, SJP^{+}13, SMGRJ10, SHO^{+}10, VKM^{+}17, WIM^{+}10, WFC^{+}18, WMW^{+}14, WCR^{+}15, WRTB10, ZCCN13, ZLKW13, ARDG17, Bai16, BBH17, CPR^{+}16, CAB^{+}18, FBWIO19, GMG^{+}19, Had15, JBB^{+}15, JGL^{+}16, KF18a, MSMCB18, NJBM15, NWW^{+}17, PCF^{+}19, RCSK15, RK17, SCSJ13, SJD^{+}19, SMZZ17, SSE^{+}17, TND16, TRPD16, WCCT18, ZSM18]. Used [HGCT13, JC18]. useful [BW19]. User [GFG^{+}12, GD15, OKCG18]. users [HBTMG15]. uses [SMARC^{+}16]. Using [ABMG19, ALA17, AB15, Ben16, BFG^{+}16, BMAS17, BTB^{+}18, BAG^{+}11, CLV^{+}10, CAB^{+}18, CWAW14, CMM17, DTH^{+}14, DMB^{+}17, DKL^{+}15, DBV13, DWCW10, DCP^{+}19, EBL^{+}14, ETB10, EPB^{+}14, GML^{+}19, GGB^{+}15, GFJC19, GW17, HN17, IP11, LW11, PMM^{+}14, RCBT18, SMARC^{+}16, SBBP11, SML^{+}13, SL15, SSMC17, SMZZ17, SKSS^{+}17, TRPD16, TTEM14, TMK^{+}16, VPS^{+}14, WSC^{+}17, ZV19, AdBR15, DBB^{+}16, HSV^{+}17, LSA17, PG16, SKPP15]. Uso [BRKP12, CAB^{+}10, CMNM10, GSB10, HJSF^{+}12, JMP^{+}11, KGS^{+}11, KG10, MC11a, MWC11, SSK^{+}11, SPPN^{+}12, SA11, SMGRJ10, SHO^{+}10, WIM^{+}10, BRV10, BGAS10, BB10a, DFW^{+}10, JLDJ12]. Utility [CHHR13, EPB^{+}14, WWWC15, SMM16]. Utilización [AWU^{+}11, AWKR^{+}11, BG10, BAG^{+}11, CMB12, CLV^{+}10, DWCW10, ETB10, IP11, LW11, LHDC12, MPH^{+}11, Rog11, SAC12, WRTB10]. Uvas [KTN12].

v25n2 [Ano11c]. Vacíos [MJP^{+}11]. Vadeadoras [WKF^{+}11]. Valid [PZM14]. validating [SKLP16]. validation [TLT16]. Validez [FC10]. Validity [FC10, LCT^{+}14]. Valkonen [BC18]. Valley [NJT14, SMS14]. Valor [AG10, BS10, BR11b, DEJM10, EMD^{+}10, EMS^{+}11, LAA^{+}10b, MR12, San10, SSO11, SSD10, SHC^{+}11, TGD10]. Valoración [AT11, MM10, MAGA10]. Valores [AG10, BRCK11, MB10, SSVA11]. Valuable [HPL14]. Valuation [GBB^{+}14]. Value [AG10, BS10, BR11b, DTH^{+}14, DEJM10, EMD^{+}10, EMS^{+}11, GMCW13, LAA^{+}10b, MR12, PCF^{+}19, SSVA11, San10, SSO11, SSD10, SHC^{+}11, TGD10, TTEM14, VFV^{+}12, WB14, Whi13, CZS19, FSJ^{+}17, JD16, LFS^{+}17, MTD16, MBT^{+}17b, SFKC16, VBN15, WWH^{+}18]. Value-of-Information [MR12]. Values [AG10, BRCK11, MB10, Whi13, WKI^{+}14, WMS^{+}12, AP17, BH17, CY17, CCB^{+}19b, DTM16, EAPB15, GKA^{+}15, HCSM15, JRWA15, JD16, MBT^{+}17a, NWD^{+}16b, SFAA^{+}18]. Valuing [AT11, MM10, HGGH15]. Vancouver [BV10, BV10]. vaquita [JLCHNG^{+}17]. Variabilidad [WKF^{+}11, WKP11]. Variability [WKF^{+}11, WKP11, MA19, RCARR18]. variable [BTB^{+}18]. Variables [CGCL11, GB14, AB15]. Variación [SM11, ATW^{+}11]. Varias [PSP11]. Variation [ATW^{+}11, SM11, ZLKW13, SGA^{+}19]. Varying [GLH12]. vascular [SLGB15]. Vecinal [SS12]. Vector [VLRB12]. Vegetación [Col10]. Vegetal [FG12]. Vegetation [Col10, FG12, FMF10, SBSF14, HCH17]. vegetation-based [HCH17]. vehicle [BMB^{+}19]. Ventajas [BRKP12]. Vents [GZD11, VLRB12]. Venus [Poo13]. Verboom [Fie15]. Verdes [SAGSHL10]. Verge [PGLB^{+}12]. versus [CMRH10, GGG10, HWS19, LP10, PNSG18, SL16]. Vertebrado [BAG^{+}11]. Vertebrados [AZCT12, FTSSW12, SRS11]. Vertebrate [AZCT12, BAG^{+}11, CSR^{+}19, LSA17, WGLL17]. Vertebrates [FTSSW12, PWRM10, PMM^{+}14, SRS11, DOC^{+}15, LP16, MAM^{+}15, QCW^{+}18, TBM^{+}16]. Very [LK14]. vessels [WFC^{+}18]. via [FBB^{+}14, HLHR16, LZD^{+}13, MBL^{+}14]. Viabilidad [SS11]. Viability [DBV13, MDJ14, PMJ^{+}13, SS11, ZCCN13, BHB^{+}19, CC18, MF15, PBB^{+}15]. Viable [RM14a, SLWL15, SBJG13, SBJG14]. Vicious [LSK^{+}14]. Vida [AWKR^{+}11, BM10, FEC^{+}10, FC10, GLNJ10, HMM10, KMGCR10, PMK^{+}10, TM10]. video [SJD^{+}19]. View [Put11, SSO12, LGA17]. Views [Par10, KBSX19]. Village [CAB^{+}10, CSMG^{+}13]. Villages [VCDF13]. Vinculación [LSDK11]. Vínculos [CMB12]. Viñedos [KTN12]. Vineyards [KTN12]. Viper [VM14]. Vireo [WLCS14]. Virtual [SCSJ13, FSJ^{+}17, WLD16]. virus [FMB^{+}18]. Vision [Lan11, TSH^{+}14]. visitors [GLK^{+}17]. Visual [FCGJH11, DP17]. Visuales [FCGJH11, FB12]. Vital [ANTBG12, JMW^{+}11, CGE19, RHPF15]. Vitales [ANTBG12, JMW^{+}11]. Vive [HBM^{+}12]. Viviendo [GGG10]. Vocal [BMA^{+}14, GGL^{+}15]. Volume [Ano12g]. Voluntarios [AGAB11]. Voluntary [AB13a]. volunteer [LS15]. Volunteers [AGAB11]. Vortex [PGLB^{+}12]. Vórtice [PGLB^{+}12]. Voting [BRM^{+}14]. Vulnerability [Fat14, DP15, TMG^{+}18, WHK^{+}17]. Vulnerable [BCGP13, HRSK12, PA18].

Vulnerables [HRSK12]. Vulture [OTKE12, ZMTR15, GTC^{+}16]. vulture-killing [GTC^{+}16]. Vultures [MSZB^{+}14, SGB^{+}19].

Wales [PMM^{+}15]. Walking [Dor11]. Wallach [DW19, RM19]. Wallets [Ver16]. Want [SCSJ13]. Wanted [Mih19]. war [Blu16]. Warblers [ZBD11]. Warming [Col10, Csó17, GFGR12, LKB^{+}14, WK10]. warning [KS17]. warriors [DHF16]. Warshall [Hel14]. Washington [Ano19w, HSBE12]. Water [BS12, JMSM13, RJF^{+}13, UIN^{+}13, WBM13, AAvRS19, BBH17, WWdS^{+}16]. Waterbird [MARV^{+}11]. waterbirds [PRB18]. Watermilfoil [RL11]. Way [KH16]. ways [RBE^{+}15]. web [HRHC16, BGG13]. webs [GHM^{+}15, GSSF19]. Wehrden [BBM15]. weighted [SC19]. Weighting [PMS13]. Wekiu [ER14]. Welfare [HMG^{+}13, HH19a]. Well [LDBS11, MGMA^{+}14]. Well-Being [LDBS11, MGMA^{+}14]. Were [Mih19]. West [BB11a, MKR^{+}15]. Western [Fie15, EB15, GGL^{+}15, KT16, AS10, CMB12, GBG^{+}10, KWW^{+}14, LSBB10, OPP14, SRG^{+}15, SSD10]. Wetland [ODTS10]. wetlands [DKL^{+}15, ROQR^{+}16, SHSP19]. whale [BCK^{+}19, KRL^{+}16, MPS^{+}16]. Whales [HCP^{+}12, HMB^{+}13, MRT^{+}12, Mor12, PS10b, RMM^{+}13, RMKM14, WHB^{+}11, GGL^{+}15]. whaling [HAF^{+}17]. whaling-conservation [HAF^{+}17]. Wheel [HP14]. Wheels [Hal16]. Where [Cox10, MGCL14]. whether [Hul15, VBN15]. which [PS18]. while [MDD^{+}18]. White [FCC^{+}11, HM11, TWZ16, DMR^{+}17, ISA16, MF15]. White-Nose [FCC^{+}11, HM11, DMR^{+}17, ISA16, MF15]. Who [Ano19w]. Whole [GFGR12]. Whole-Ecosystem [GFGR12]. Whooping [BKKM14, HBTB15]. Wicked [DSW16, HHY13]. Wide [KPL^{+}12, SBT^{+}11, SMF^{+}15, SMZZ17]. Widespread [AASK12, HAP^{+}13, Pek13]. Width [CPB14]. widths [Bei19]. Wild [BW13, CBM^{+}11, GGM14, LW10, PMS^{+}10, Wit11, ZYYH10, MBBB17, WC19, ZMTR15, vWWCF15]. Wilderness [LK13, Sel16, AKJ^{+}18]. Wildfire [HLC13]. Wildlife [BM10, CD16, CRD^{+}14, DVW^{+}13, FEC^{+}10, FC10, GKB10, GBB^{+}14, HMM10, KKK14, LW14, MF14, PMS13, PCW14, SCK13, Sko19, TM10, BCR^{+}17, Blu16, BMA^{+}18, CBPM19, CGK^{+}19, DDv18, DFR^{+}17, DJBB16, ESD^{+}17, FOF^{+}15, FSR^{+}18, FSAM15, FWC^{+}15, GMNO15, GGB^{+}15, HRHC16, KBSX19, LSA17, LFMB19, LNNG16, MFHT19, MSR19, Nat18, NBA^{+}18, Poz18, RN17, RS18, SMARC^{+}16, SSRS18, SMLB19, SKLP16, TAM^{+}19, TVY^{+}16, VNAT19, VW19, WBF^{+}19, WSLW16]. Wildlife-Crossing [FC10]. Wildlife-Friendly [GKB10]. Wildlife-Trade [BM10]. wildness [CSR^{+}19]. Wiley [Ano19w]. Wilhere [CRNR12]. Willingness [AAvRS19, WL12, FWC^{+}15]. Wind [BKKM14, MHG^{+}14, MBL^{+}14, Sko19, FBWIO19, KNB^{+}17, SB16]. wind-energy [KNB^{+}17, SB16]. Winning [Ver16]. winter [EKRW17, SKBE15]. winter-quarter [SKBE15]. Wintering [EG12, SMF^{+}15]. Wire [BAPP11]. wisdom [PLH^{+}16]. within [AS13, GJB14, LKYL15]. without [Dav11, FHBK18]. Wittmer [TW11]. Wolf [ACL^{+}12, CMB12, CFL14]. Wolverine [EIS14, EBP^{+}18]. Wolves [CPC^{+}10, LJA^{+}13, SBB^{+}14, TNTS13, DSE15, JGL^{+}16, KMLC19]. wood [PWKB19, SMF^{+}15]. wood-inhabiting [PWKB19]. Wooded [EMD^{+}10]. Woodland [DWH^{+}11, SW10, SBG^{+}18]. Woodlands [EDB12]. Woodpecker [GCC^{+}12, RC12, SSB^{+}12]. Woodrat [SJP^{+}13]. Woody [KZG^{+}13, VJMH12]. work [KBM17, NRD^{+}16]. working [CNBW16, WK15]. Works [Ibi17, SHO^{+}10]. World [BBC^{+}14, Cox10, DTH^{+}14, GZA10, HMIE14, Ibi17, Sel16, SLSH14, SBN^{+}10, Sta16, Ben15c, Ben15a, BHO^{+}18, CLvK^{+}17, DTM16, GLJ^{+}16, HDH^{+}15, Mor15, SSWR17, SSD^{+}18, TK15, AKJ^{+}18, Ano19w, SGB^{+}19]. worldviews [KHK^{+}19]. Would [Löv10]. Wrestling [PWP^{+}18].

xiv [Ano19w]. xx [Ano19w]. xxii [Ano19w]. xylocarpa [ZYYH10].

Yangtze [TBY^{+}10]. Yapu [PS16]. Year [Bai10, DFJ^{+}12, Joh10b, MRT^{+}12, DPGY16, GBB^{+}15, JWS^{+}18]. Year-Round [MRT^{+}12]. Years [EDB12, FBF^{+}14, KGG^{+}14, MQAMWM12, MLA^{+}14, PGC^{+}13b, Tru18a, HTH^{+}17b, PCF^{+}19]. Yellowstone [BB10b]. yield [CS17a].

Zealand [AGAB11, BJMT14, Fer19, HWW12, MBG14, Moo17, RSPD13, WMW^{+}14]. Zelanda [AGAB11, HWW12]. Zimbabwe [GDMF13]. Zonas [EPD12]. zone [BMB^{+}19, HTDM17]. Zones [EPD12, DKG^{+}15]. Zoo [Jen14, DPGY16, GLK^{+}17, MRPT19]. Zoos [Kra12, Alr15, BCLMM19, MJG15].

References

Allendorf:2012:WEC

[AA12] T. D. Allendorf and K. Allendorf. What every conservation biologist should know about human population. Conservation Biology, 26(6):953–955, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aryal:2017:GLS

[AAS^{+}17] Achyut Aryal, Krishna Prasad Acharya, Uttam Babu Shrestha, Maheshwar Dhakal, David Raubenhiemer, and Wendy Wright. Global lessons from successful rhinoceros conservation in Nepal. Conservation Biology, 31(6):1494–1497, December 2017. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anderegg:2012:EWD

[AASK12] William R. L. Anderegg, Leander D. L. Anderegg, Clare Sherman, and Daniel S. Karp. Effects of widespread droughtinduced aspen mortality on understory plants. Conservation Biology, 26(6):1082–1090, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armstrong:2019:WPP

[AAvRS19] Claire W. Armstrong, Margrethe Aanesen, Thomas M. van Rensburg, and Erlend Dancke Sandorf. Willingness to pay to protect cold water corals. Conservation Biology, 33(6):1329– 1337, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Asah:2013:PIU

[AB13a] Stanley T. Asah and Dale J. Blahna. Practical implications of understanding the influence of motivations on commitment to voluntary urban conservation stewardship. Conservation Biology, 27(4):866–875, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Augustine:2013:AGB

[AB13b] David J. Augustine and Bruce W. Baker. Associations of grassland bird communities with black-tailed prairie dogs in the North American Great Plains. Conservation Biology, 27 (2):324–334, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Albuquerque:2015:UAV

[AB15] Fabio Albuquerque and Paul Beier. Using abiotic variables to predict importance of sites for species representation. Conservation Biology, 29(5):1390–1400, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Akcakaya:2018:QSR

[ABB^{+}18] H. Resit Akçakaya, Elizabeth L. Bennett, Thomas M. Brooks, Molly K. Grace, Anna Heath, Simon Hedges, Craig Hilton-Taylor, Michael Hoffmann, David A. Keith, Barney Long, David P. Mallon, Erik Meijaard, E. J. Milner-Gulland, Ana S. L. Rodrigues, Jon Paul Rodriguez, P. J. Stephenson, Simon N. Stuart, and Richard P. Young. Quantifying species recovery and conservation success to develop an IUCN green list of species. Conservation Biology, 32(5):1128–1138, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Angelsen:2017:LRR

[ABD^{+}17] Arild Angelsen, Maria Brockhaus, Amy E. Duchelle, Anne Larson, Christopher Martius, William D. Sunderlin, Louis Verchot, Grace Wong, and Sven Wunder. Learning from REDD+: a response to Fletcher et al. Conservation Biology, 31(3):718–720, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Adler:2019:DPS

[ABG^{+}19] Peter H. Adler, Jeb Barzen, Elmer Gray, Anne Lacy, Richard P. Urbanek, and Sarah J. Converse. The dilemma of pest suppression in the conservation of endangered species. Conservation Biology, 33(4):788–796, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Addison:2019:UCS

[ABMG19] Prue F. E. Addison, Joseph W. Bull, and E. J. Milner-Gulland. Using conservation science to advance corporate biodiversity accountability. Conservation Biology, 33(2):307–318, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Abrams:2015:WWS

[Abr15] R. W. Abrams. Why we should help people understand our scientific literature. Conservation Biology, 29(4):973–974, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alleway:2015:LEB

[AC15] Heidi K. Alleway and Sean D. Connell. Loss of an ecological baseline through the eradication of oyster reefs from coastal ecosystems and human memory. Conservation Biology, 29 (3):795–804, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ament:2016:SDE

[AC16] Judith M. Ament and Graeme S. Cumming. Scale dependency in effectiveness, isolation, and social-ecological spillover of protected areas. Conservation Biology, 30(4):846–855, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anderson:2015:CSC

[ACB^{+}15] M. G. Anderson, P. J. Comer, P. Beier, J. J. Lawler, C. A. Schloss, S. Buttrick, C. M. Albano, and D. P. Faith. Case studies of conservation plans that incorporate geodiversity. Conservation Biology, 29(3):680–691, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Arponen:2010:CIE

[ACE^{+}10] Anni Arponen, Mar Cabeza, Johanna Eklund, Heini Kujala, and Joona Lehtomäki. Costs of integrating economics and conservation planning. Costos de la integración de economía y planificación de la conservación. Conservation Biology, 24(5): 1198–1204, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Araiza:2012:CCP

[ACL^{+}12] Marcela Araiza, Luis Carrillo, Rurik List, Carlos A. López González, Enrique Martínez Meyer, Patricia G. Martínez-Gutiérrez, Oscar Moctezuma, Nahum E. Sánchez-Morales, and Jorge Servín. Consensus on criteria for potential areas for wolf reintroduction in Mexico. Consenso en criterios para Áreas potenciales para la reintroducció de lobos en México. Conservation Biology, 26(4):630–637, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anderson:2014:ECR

[ACS14] Mark G. Anderson, Melissa Clark, and Arlene Olivero Sheldon. Estimating climate resilience for conservation across geophysical settings. Conservation Biology, 28(4):959–970, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Accatino:2018:LCA

[ACW18] Francesco Accatino, Irena F. Creed, and Marian Weber. Landscape consequences of aggregation rules for functional equivalence in compensatory mitigation programs. Conservation Biology, 32(3):694–705, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Addison:2015:SCM

[AdBR15] P. F. E. Addison, K. de Bie, and L. Rumpff. Setting conservation management thresholds using a novel participatory modeling approach. Conservation Biology, 29(5):1411–1422, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Adum:2013:TSR

[AEO^{+}13] Gilbert Baase Adum, Markus Peter Eichhorn, William Oduro, Caleb Ofori-Boateng, and Mark-Oliver Rödel. Two-stage recovery of amphibian assemblages following selective logging of tropical forests. Conservation Biology, 27(2):354–363, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Akter:2010:CUM

[AG10] Sonia Akter and R. Quentin Grafton. Confronting uncertainty and missing values in environmental value transfer as applied to species conservation. Confrontando incertidumbre y valores faltantes en la trasferencia de valor ambiental aplicada a la conservación de especies. Conservation Biology, 24(5):1407–1417, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Agarwala:2017:UOJ

[AG17] Meghna Agarwala and Joshua R. Ginsberg. Untangling outcomes of de jure and de facto community-based management of natural resources. Conservation Biology, 31(6):1232–1246, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Acevedo-Gutierrez:2011:EPO

[AGAB11] Alejandro Acevedo-Gutiérrez, Lisa Acevedo, and Laura Boren. Effects of the presence of official-looking volunteers on harassment of New Zealand fur seals. Efectos de la presencia de voluntarios en el hostigamiento a focas finas de Nueva Zelanda. Conservation Biology, 25(6):623–627, June 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Afan:2018:AMI

[AGFR18] Isabel Afán, Joan Giménez, Manuela G. Forero, and Francisco Ramírez. An adaptive method for identifying marine areas of high conservation priority. Conservation Biology, 32(6):1436– 1447, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aggestam:2015:FEC

[Agg15] Filip Aggestam. Framing the ecosystem concept through a longitudinal study of developments in science and policy. Conservation Biology, 29(4):1052–1064, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Almeida-Gomes:2014:UET

[AGLRV14] Mauricio Almeida-Gomes, Maria Lucia Lorini, Carlos Frederico Duarte Rocha, and Marcus Vinícius Vieira. Underestimation of extinction threat to stream-dwelling amphibians due to lack of consideration of narrow area of occupancy. Conservation Biology, 28(2):616–619, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armstrong:2014:SRC

[AGP^{+}14] Doug P. Armstrong, Nic Gorman, Rhonda Pike, Brigitte Kreigenhofer, Nikki Mcarthur, Susanne Govella, Paul Barrett, and Yvan Richard. Strategic rat control for restoring populations of native species in forest fragments. Conservation Biology, 28(3):713–723, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Arregoitia:2019:CAP

[AGS19] Luis D. Verde Arregoitia and Manuela González-Suárez. From conference abstract to publication in the conservation science literature. Conservation Biology, 33(5):1164–1173, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ando:2011:LFN

[AH11] Amy W. Ando and Lee Hannah. Lessons from finance for new land-conservation strategies given climate-change uncertainty. Conservation Biology, 26(3):412–414, April 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alexander:2012:LTE

[AH12] Louise Alexander and George R. Hess. Land trust evaluation of progress toward conservation goals. Evaluación de fideicomiso de tierras del progreso hacia metas de conservación. Conservation Biology, 26(2):7–12, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Adame:2015:SCE

[AHP^{+}15] M. F. Adame, V. Hermoso, K. Perhans, C. E. Lovelock, and J. A. Herrera-Silveira. Selecting cost-effective areas for restoration of ecosystem services. Conservation Biology, 29(2):493– 502, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Andelman:2012:I

[AHS12] Sandy Andelman, Lee Hannah, and Rebecca Shaw. Introduction. Conservation Biology, 26(3):385, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aslan:2015:BPS

[AHT^{+}15] Clare Aslan, Nick Holmes, Bernie Tershy, Dena Spatz, and Donald A. Croll. Benefits to poorly studied taxa of conservation of bird and mammal diversity on islands. Conservation Biology, 29(1):133–142, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Amit:2018:PDI

[AJ18] Ronit Amit and Susan K. Jacobson. Participatory development of incentives to coexist with jaguars and pumas. Conservation Biology, 32(4):938–948, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Audzijonyte:2016:RLH

[AK16] Asta Audzijonyte and Anna Kuparinen. The role of life histories and trophic interactions in population recovery. Conservation Biology, 30(4):734–743, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allan:2018:GOW

[AKJ^{+}18] James R. Allan, Cyril Kormos, Tilman Jaeger, Oscar Venter, Bastian Bertzky, Yichuan Shi, Brendan Mackey, Remco van Merm, Elena Osipova, and James E. M. Watson. Gaps and opportunities for the World Heritage Convention to contribute to global wilderness conservation. Conservation Biology, 32(1): 116–126, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aswani:2014:IPD

[AL14] Shankar Aswani and Matthew Lauer. Indigenous People’s detection of rapid ecological change. Conservation Biology, 28 (3):820–828, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aiello-Lammens:2017:UGS

[ALA17] Matthew E. Aiello-Lammens and H. Resit Akçakaya. Using global sensitivity analysis of demographic models for ecological impact assessment. Conservation Biology, 31(1):116–125, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aldrich:2012:MMC

[Ald12] Stephen C. Aldrich. Meeting the meat crisis. Conservation Biology, 27(1):752–754, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Arponen:2012:ECS

[ALL^{+}12] Anni Arponen, Joona Lehtomäki, Jarno Leppänen, Erkki Tomppo, and Atte Moilanen. Effects of connectivity and spatial resolution of analyses on conservation prioritization across large extents. Efectos de la conectividad y la resolución espacial de análisis sobre priorización de la conservación en grandes extensiones. Conservation Biology, 26(2):294–304, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alroy:2015:LCB

[Alr15] John Alroy. Limits to captive breeding of mammals in zoos. Conservation Biology, 29(3):926–931, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alcala:2019:USP

[ALW^{+}19] Nicolas Alcala, Alan E. Launer, Michael F. Westphal, Richard Seymour, Esther M. Cole, and Noah A. Rosenberg. Use of stochastic patch occupancy models in the California red-legged frog for Bayesian inference regarding past events and future persistence. Conservation Biology, 33(3):685–696, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ahlering:2013:COP

[AME^{+}13] M. A. Ahlering, J. E. Maldonado, L. S. Eggert, R. C. Fleischer, D. Western, and J. L. Brown. Conservation outside protected areas and the effect of human-dominated landscapes on stress hormones in savannah elephants. Conservation Biology, 27(3):569–575, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anderson:2011:RPD

[AMO11] Elizabeth P. Anderson and Javier A. Maldonado-Ocampo. A regional perspective on the diversity and conservation of tropical Andean fishes. Una perspectiva regional de la diversidad y conservación de peces andinos tropicales. Conservation Biology, 25(1):30–39, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Attard:2016:NHF

[AMS^{+}16] C. R. M. Attard, L. M. Möller, M. Sasaki, M. P. Hammer, C. M. Bice, C. J. Brauer, D. C. Carvalho, J. O. Harris, and L. B. Beheregaray. A novel holistic framework for genetic-based captive-breeding and reintroduction programs. Conservation Biology, 30(5):1060–1069, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Abelson:2016:EMP

[ANE^{+}16] A. Abelson, P. A. Nelson, G. J. Edgar, N. Shashar, D. C. Reed, J. Belmaker, G. Krause, M. W. Beck, E. Brokovich, R. France, and S. D. Gaines. Expanding marine protected areas to include degraded coral reefs. Conservation Biology, 30(6):1182–1191, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2010:CCa

[Ano10a] Anonymous. Cover caption. Conservation Biology, 26(2):i, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2010:CCb

[Ano10b] Anonymous. Cover caption. Conservation Biology, 24(2):i, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2010:CCc

[Ano10c] Anonymous. Cover caption. Conservation Biology, 25(6):i, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2010:CCd

[Ano10d] Anonymous. Cover caption. Conservation Biology, 24(4):i, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2010:CCe

[Ano10e] Anonymous. Cover caption. Conservation Biology, 24(5):i, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2010:CCf

[Ano10f] Anonymous. Cover caption. Conservation Biology, 24(6):i, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2010:EA

[Ano10g] Anonymous. Editor’s acknowledgments. Conservation Biology, 24(6):1439, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2010:NIa

[Ano10h] Anonymous. Noted with interest. Conservation Biology, 24(2): 640, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2010:NIb

[Ano10i] Anonymous. Noted with interest. Conservation Biology, 25 (6):913–914, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2010:NIc

[Ano10j] Anonymous. Noted with interest. Conservation Biology, 24(4): 1168, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2010:NId

[Ano10k] Anonymous. Noted with interest. Conservation Biology, 24 (5):1434, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2010:NIe

[Ano10l] Anonymous. Noted with interest. Conservation Biology, 24 (6):1693, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:BRJ

[Ano11a] Anonymous. Books received (January 2011–August 2011). Conservation Biology, 25(6):1259–1262, December 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:BRS

[Ano11b] Anonymous. Books received (September 2009–December 2010). Conservation Biology, 25(6):648–657, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:CVC

[Ano11c] Anonymous. Correction; [to v25n2 cover image caption]. Conservation Biology, 25(6):i, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:CCa

[Ano11d] Anonymous. Cover caption. Conservation Biology, 25(1):i, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:CCb

[Ano11e] Anonymous. Cover caption. Conservation Biology, 26(3):i, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:CCc

[Ano11f] Anonymous. Cover caption. Conservation Biology, 25(2):i, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:CCd

[Ano11g] Anonymous. Cover caption. Conservation Biology, 25(4):i, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:CCe

[Ano11h] Anonymous. Cover caption. Conservation Biology, 27(3):i, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:EA

[Ano11i] Anonymous. Editor’s acknowledgments. Conservation Biology, 25(6):1071, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:E

[Ano11j] Anonymous. Errata. Conservation Biology, 25(6):1263, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:NIa

[Ano11k] Anonymous. Noted with interest. Conservation Biology, 25 (1):208, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:NIb

[Ano11l] Anonymous. Noted with interest. Conservation Biology, 26(3): 420, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:NIc

[Ano11m] Anonymous. Noted with interest. Conservation Biology, 25(6): 648, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:NId

[Ano11n] Anonymous. Noted with interest. Conservation Biology, 25 (4):852–853, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:NIe

[Ano11o] Anonymous. Noted with interest. Conservation Biology, 27(3): 1063–1064, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2011:NIf

[Ano11p] Anonymous. Noted with interest. Conservation Biology, 25(6): 1258–1259, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2012:AHH

[Ano12a] Anonymous. Ad hoc Handling Editors. Conservation Biology, 26(6):956, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2012:CCa

[Ano12b] Anonymous. Cover caption. Conservation Biology, 26(2):i, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2012:CCb

[Ano12c] Anonymous. Cover caption. Conservation Biology, 27(1):i, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2012:CCc

[Ano12d] Anonymous. Cover caption. Conservation Biology, 26(5):i, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2012:CCCa

[Ano12e] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 26(6):i, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2012:E

[Ano12f] Anonymous. Errata. Conservation Biology, 26(2):378, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2012:EMC

[Ano12g] Anonymous. Erratum for manuscript CB 10–166, published in volume 25, no. 2, 276–284. Conservation Biology, 26(3):581, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [RWB11].

Anonymous:2012:NIa

[Ano12h] Anonymous. Noted with interest. Conservation Biology, 26 (2):194, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2012:NIb

[Ano12i] Anonymous. Noted with interest. Conservation Biology, 26 (2):376–377, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2012:NIc

[Ano12j] Anonymous. Noted with interest. Conservation Biology, 26(3): 580, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2012:NId

[Ano12k] Anonymous. Noted with interest. Conservation Biology, 27 (1):754–755, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2012:NIe

[Ano12l] Anonymous. Noted with interest. Conservation Biology, 26(6): 1160–1161, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2013:CCCa

[Ano13a] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 27(2):i, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2013:CCCb

[Ano13b] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 27(4):i, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2013:CCCc

[Ano13c] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 28(4):i, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2013:CCCd

[Ano13d] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 28(1):i, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2013:EAa

[Ano13e] Anonymous. Editor’s acknowledgments. Conservation Biology, 27(6):1136, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2013:Ea

[Ano13f] Anonymous. Erratum. Conservation Biology, 27(1):238, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2013:Eb

[Ano13g] Anonymous. Erratum. Conservation Biology, 27(3):642, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2013:NIa

[Ano13h] Anonymous. Noted with interest. Conservation Biology, 27 (1):237, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2013:NIb

[Ano13i] Anonymous. Noted with interest. Conservation Biology, 27 (1):1–2, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2013:NIc

[Ano13j] Anonymous. Noted with interest. Conservation Biology, 27 (2):435–436, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2013:NId

[Ano13k] Anonymous. Noted with interest. Conservation Biology, 27(3): 641, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2013:NIe

[Ano13l] Anonymous. Noted with interest. Conservation Biology, 27 (4):893–894, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2013:NIf

[Ano13m] Anonymous. Noted with interest. Conservation Biology, 28(4): 1131–1132, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2013:NIg

[Ano13n] Anonymous. Noted with interest. Conservation Biology, 28(1): 1493–1494, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2014:BRB

[Ano14a] Anonymous. Books received and book reviewers, September 2012–December 2013. Conservation Biology, 28(2):623–626, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2014:CCCa

[Ano14b] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 28(2):i, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2014:CCCb

[Ano14c] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 28(5):i, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2014:Ea

[Ano14d] Anonymous. Erratum. Conservation Biology, 28(1):298, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2014:Eb

[Ano14e] Anonymous. Erratum. Conservation Biology, 28(4):1135, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2014:Ec

[Ano14f] Anonymous. Erratum. Conservation Biology, 28(6):1441, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2014:Ed

[Ano14g] Anonymous. Erratum. Conservation Biology, 28(6):1743, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2014:NIa

[Ano14h] Anonymous. Noted with interest. Conservation Biology, 28(1): 296–297, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2014:NIb

[Ano14i] Anonymous. Noted with interest. Conservation Biology, 28 (2):622–623, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2014:NIc

[Ano14j] Anonymous. Noted with interest. Conservation Biology, 28 (5):887–888, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2014:NId

[Ano14k] Anonymous. Noted with interest. Conservation Biology, 28(4): 1134, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2014:NIe

[Ano14l] Anonymous. Noted with interest. Conservation Biology, 28 (6):1440, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2015:CCCa

[Ano15a] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 29(2):i, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2015:CCCb

[Ano15b] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 29(3):i, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2015:CCCc

[Ano15c] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 29(4):i, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2015:CCCd

[Ano15d] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 29(5):i, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2015:CCCe

[Ano15e] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 29(6):i, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2015:Ea

[Ano15f] Anonymous. Erratum. Conservation Biology, 29(2):607–610, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2015:Eb

[Ano15g] Anonymous. Erratum. Conservation Biology, 29(6):1731, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2015:NIa

[Ano15h] Anonymous. Noted with interest. Conservation Biology, 29(2): 305–306, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2015:NIb

[Ano15i] Anonymous. Noted with interest. Conservation Biology, 29 (2):605–606, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2015:NIc

[Ano15j] Anonymous. Noted with interest. Conservation Biology, 29 (3):970–971, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2015:NId

[Ano15k] Anonymous. Noted with interest. Conservation Biology, 29(4): 1250–1251, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2015:RRB

[Ano15l] Anonymous. Recently received books. Conservation Biology, 29(4):1251–1254, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:BRF

[Ano16a] Anonymous. Books received: (February–October 2015). Conservation Biology, 30(1):236–238, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:BRN

[Ano16b] Anonymous. Books received, November 2015–March 2016. Conservation Biology, 30(4):922–923, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:CCCa

[Ano16c] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 30(1):i, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:CCCb

[Ano16d] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 30(2):1, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:CCCc

[Ano16e] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 30(3):i, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:CCCd

[Ano16f] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 30(4):i, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:CCCe

[Ano16g] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 30(5):i, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:CCCf

[Ano16h] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 30(6):i, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:E

[Ano16i] Anonymous. Erratum. Conservation Biology, 30(6):1372, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:IIa

[Ano16j] Anonymous. Issue information. Conservation Biology, 30(1):1– 2, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:IIb

[Ano16k] Anonymous. Issue information. Conservation Biology, 30(1): 1, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:IICa

[Ano16l] Anonymous. Issue information — copyright page. Conservation Biology, 30(2):241, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:IICb

[Ano16m] Anonymous. Issue information — copyright page. Conservation Biology, 30(3):445, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:IICc

[Ano16n] Anonymous. Issue information — copyright page. Conservation Biology, 30(4):688, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:IICd

[Ano16o] Anonymous. Issue information — copyright page. Conservation Biology, 30(5):927, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:IICe

[Ano16p] Anonymous. Issue information — copyright page. Conservation Biology, 30(6):1147, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:IITa

[Ano16q] Anonymous. Issue information — TOC. Conservation Biology, 30(2):239–240, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:IITb

[Ano16r] Anonymous. Issue information — TOC. Conservation Biology, 30(3):443–444, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:IITc

[Ano16s] Anonymous. Issue information — TOC. Conservation Biology, 30(4):685–687, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:IITd

[Ano16t] Anonymous. Issue information — TOC. Conservation Biology, 30(5):925–926, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:IITe

[Ano16u] Anonymous. Issue information — TOC. Conservation Biology, 30(6):1145–1146, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:NIa

[Ano16v] Anonymous. Noted with interest. Conservation Biology, 30 (5):1143, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:NIb

[Ano16w] Anonymous. Noted with interest. Conservation Biology, 30 (6):1371, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:NIF

[Ano16x] Anonymous. Noted with interest: Foundations of Macroecology — Classic Papers with Commentaries. Smith, F. A., J. L. Gittleman, and J. H. Brown, editors. 2014. The University of Chicago Press, Chicago, IL, U.S.A. 824 pp. $55.00 (paperback). ISBN 13 978-0-226-11547-4. Lost Animals: Extinction and the Photographic Record. Fuller, E. 2014. Princeton University Press, Princeton, NJ, U.S.A. 240 pp. $29.95 (hardcover). ISBN 978-0-691-16137-2. Conservation Biology, 30(1):235, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2016:RNI

[Ano16y] Anonymous. Reviewers for “Noted with Interest”. Conservation Biology, 30(5):1143, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:BRA

[Ano17a] Anonymous. Books received (April–November 2016). Conservation Biology, 31(1):236–238, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:CBA

[Ano17b] Anonymous. Conservation Biology awards. Conservation Biology, 31(3):497–498, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:CCCa

[Ano17c] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 31(1):i, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:CCCb

[Ano17d] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 31(2):i, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:CCCc

[Ano17e] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 31(3):i, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:CCCd

[Ano17f] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 31(4):i, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:CCCe

[Ano17g] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 31(5):i, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:CCCf

[Ano17h] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 31(6):i, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:Ea

[Ano17i] Anonymous. Erratum. Conservation Biology, 31(1):239, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [DSW16].

Anonymous:2017:Eb

[Ano17j] Anonymous. Erratum. Conservation Biology, 31(3):490–491, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:Ec

[Ano17k] Anonymous. Erratum. Conservation Biology, 31(3):738, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:IICa

[Ano17l] Anonymous. Issue information — copyright page. Conservation Biology, 31(1):3, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:IICb

[Ano17m] Anonymous. Issue information — copyright page. Conservation Biology, 31(2):243, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:IICc

[Ano17n] Anonymous. Issue information — copyright page. Conservation Biology, 31(3):495, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:IICd

[Ano17o] Anonymous. Issue information — copyright page. Conservation Biology, 31(4):741, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:IICe

[Ano17p] Anonymous. Issue information — copyright page. Conservation Biology, 31(5):965, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:IICf

[Ano17q] Anonymous. Issue information — copyright page. Conservation Biology, 31(6):1220, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:IITa

[Ano17r] Anonymous. Issue information — TOC. Conservation Biology, 31(1):1–2, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:IITb

[Ano17s] Anonymous. Issue information — TOC. Conservation Biology, 31(2):241–242, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:IITc

[Ano17t] Anonymous. Issue information — TOC. Conservation Biology, 31(3):493–494, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:IITd

[Ano17u] Anonymous. Issue information — TOC. Conservation Biology, 31(4):739–740, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:IITe

[Ano17v] Anonymous. Issue information — TOC. Conservation Biology, 31(5):963–964, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:IITf

[Ano17w] Anonymous. Issue information — TOC. Conservation Biology, 31(6):1217–1219, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:NIa

[Ano17x] Anonymous. Noted with interest. Conservation Biology, 31(1): 232–236, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:NIb

[Ano17y] Anonymous. Noted with interest. Conservation Biology, 31 (3):735–737, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:NIc

[Ano17z] Anonymous. Noted with interest. Conservation Biology, 31 (4):960–961, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:NId

[Ano17-27] Anonymous. Noted with interest. Conservation Biology, 31(5): 1213–1215, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2017:NIe

[Ano17-28] Anonymous. Noted with interest. Conservation Biology, 31(6): 1503–1505, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:BRJ

[Ano18a] Anonymous. Book reviewers, January-December 2017. Conservation Biology, 32(2):508, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:BRD

[Ano18b] Anonymous. Books received, December 2016–September 2017. Conservation Biology, 32(1):254–257, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:CBA

[Ano18c] Anonymous. Conservation Biology 2018 awards. Conservation Biology, 32(4):761, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:CCa

[Ano18d] Anonymous. Cover caption. Conservation Biology, 32(4):i, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:CCb

[Ano18e] Anonymous. Cover caption. Conservation Biology, 32(5):i, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:CCd

[Ano18f] Anonymous. Cover caption. Conservation Biology, 32(6):i, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:CCCa

[Ano18g] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 32(1):i, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:CCCb

[Ano18h] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 32(2):i, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:CCCc

[Ano18i] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 32(3):i, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:IIa

[Ano18j] Anonymous. Issue information. Conservation Biology, 32 (3):509–511, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:IIb

[Ano18k] Anonymous. Issue information. Conservation Biology, 32(4): 757–759, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:IId

[Ano18l] Anonymous. Issue information. Conservation Biology, 32(5): 975–977, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:IIe

[Ano18m] Anonymous. Issue information. Conservation Biology, 32(6): 1217–1219, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:IICa

[Ano18n] Anonymous. Issue information — copyright page. Conservation Biology, 32(1):4, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:IICb

[Ano18o] Anonymous. Issue information — copyright page. Conservation Biology, 32(2):263, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:IITa

[Ano18p] Anonymous. Issue information — TOC. Conservation Biology, 32(1):1–3, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:IITb

[Ano18q] Anonymous. Issue information — TOC. Conservation Biology, 32(2):261–262, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:NIa

[Ano18r] Anonymous. Noted with interest. Conservation Biology, 32(1): 252–254, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:NIb

[Ano18s] Anonymous. Noted with interest. Conservation Biology, 32 (2):506–508, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:NIc

[Ano18t] Anonymous. Noted with interest. Conservation Biology, 32 (3):754–756, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:NId

[Ano18u] Anonymous. Noted with interest. Conservation Biology, 32 (4):971–973, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:NIe

[Ano18v] Anonymous. Noted with interest. Conservation Biology, 33(1): 1215–1216, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:NIf

[Ano18w] Anonymous. Noted with interest. Conservation Biology, 32(6): 1483–1484, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2018:PAC

[Ano18x] Anonymous. Publication of abstracts in Chinese. Conservation Biology, 32(1):5, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:BRJ

[Ano19a] Anonymous. Book reviewers, January-December 2018. Conservation Biology, 33(4):982, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:BRO

[Ano19b] Anonymous. Books received (October 2017–February 2019). Conservation Biology, 33(4):982–986, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:CBA

[Ano19c] Anonymous. Conservation Biology awards. Conservation Biology, 33(3):497, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:CCa

[Ano19d] Anonymous. Cover caption. Conservation Biology, 33(1):i, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:CCb

[Ano19e] Anonymous. Cover caption. Conservation Biology, 33(2):i, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:CCc

[Ano19f] Anonymous. Cover caption. Conservation Biology, 33(3):i, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:CCd

[Ano19g] Anonymous. Cover caption. Conservation Biology, 33(4):i, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:CCCa

[Ano19h] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 33(5):i, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:CCCb

[Ano19i] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 33(6):i, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:Ea

[Ano19j] Anonymous. Erratum. Conservation Biology, 33(2):225, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:Eb

[Ano19k] Anonymous. Erratum. Conservation Biology, 33(5):1229, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:IITa

[Ano19l] Anonymous. Issue information — TOC. Conservation Biology, 33(1):1–3, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:IITb

[Ano19m] Anonymous. Issue information — TOC. Conservation Biology, 33(2):227–229, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:IITc

[Ano19n] Anonymous. Issue information — TOC. Conservation Biology, 33(3):493–495, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:IITd

[Ano19o] Anonymous. Issue information — TOC. Conservation Biology, 33(4):743–745, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:IITe

[Ano19p] Anonymous. Issue information — TOC. Conservation Biology, 33(5):987–990, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:IITf

[Ano19q] Anonymous. Issue information — TOC. Conservation Biology, 33(6):1231–1233, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:NIa

[Ano19r] Anonymous. Noted with interest. Conservation Biology, 33 (2):224, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:NIb

[Ano19s] Anonymous. Noted with interest. Conservation Biology, 33 (2):488–490, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:NIc

[Ano19t] Anonymous. Noted with interest. Conservation Biology, 33 (3):741–742, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:NId

[Ano19u] Anonymous. Noted with interest. Conservation Biology, 33 (4):981–982, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:NIe

[Ano19v] Anonymous. Noted with interest. Conservation Biology, 33(5): 1227–1228, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2019:NIEa

[Ano19w] Anonymous. Noted with interest: Ecological Methods. 4th edition. Henderson, P. A., and T. R. E. Southwood. 2016. John Wiley & Sons, Chichester, U.K. xx + 632 pp. US$190.00 (hardcover). ISBN 978-1-118-89527-6. The Last Butterflies. Haddad, N. 2019. Princeton University Press, Princeton, NJ, U.S.A. xiv + 250 pp. US$24.95 (hardcover). ISBN 978-0-691-16500-4. The Lost Species. Kemp, C. 2017. University of Chicago Press, Chicago, IL, U.S.A. xxii + 250 pp. US$30.00 (hardcover). ISBN 978-0-226-38621-8. Nature’s Allies. Eight Conservationists Who Changed Our World. Nielsen, L. A. 2018. Island Press, Washington, D.C., U.S.A. xiv + 255 pp. US$21.00 (paperback). ISBN 978-1-61091-796-4. Conservation Biology, 33 (6):1453–1455, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Andrello:2012:ESS

[ANTBG12] Marco Andrello, Florence Nicolè, Irène Till-Bottraud, and Oscar E. Gaggiotti. Effect of stage-specific vital rates on population growth rates and effective population sizes in an endangered iteroparous plant. Efecto de las tasas vitales Específicas de estadios sobre las tasas de crecimiento poblacional y los Tamaños poblacionales efectivos en una planta Iterópara en peligro. Conservation Biology, 26(2):208–217, April 2012. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Athreya:2011:TTM

[AOLK11] Vidya Athreya, Morten Odden, John D. C. Linnell, and K. Ullas Karanth. Translocation as a tool for mitigating conflict with leopards in human-dominated landscapes of India. La translocación como una herramienta para mitigar conflictos con leopardos en paisajes dominados por humanos en India. Conservation Biology, 25(1):133–141, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Albrecht:2019:ELH

[AOPM^{+}19] Matthew A. Albrecht, Oyomoare L. Osazuwa-Peters, Joyce Maschinski, Timothy J. Bell, Marlin L. Bowles, William E. Brumback, Janice Duquesnel, Michael Kunz, Jimmy Lange, Kimberlie A. McCue, A. Kathryn McEachern, Sheila Murray, Peggy Olwell, Noel B. Pavlovic, Cheryl L. Peterson, Jennifer Possley, John L. Randall, and Samuel J. Wright. Effects of life history and reproduction on recruitment time lags in reintroductions of rare plants. Conservation Biology, 33(3):601–611, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Addison:2017:ABV

[AP17] Jane Addison and Chris R. Pavey. Alignment between values of dryland pastoralists and conservation needs for small mammals. Conservation Biology, 31(3):331–342, April 2017. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alexiades:2013:MSS

[APL^{+}13] Miguel N. Alexiades, Charles M. Peters, Sarah A. Laird, Citlalli López Binnqüist, and Patricia Negreros Castillo. The missing skill set in community management of tropical forests. Conservation Biology, 27(3):635–637, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aslan:2014:CCC

[APR^{+}14] Clare E. Aslan, Malin L. Pinsky, Maureen E. Ryan, Sara Souther, and Kimberly A. Terrell. Cultivating creativity in conservation science. Conservation Biology, 28(2):345–353, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Adams:2010:ESC

[AR10] William M. Adams and Kent H. Redford. Ecosystem services and conservation: a reply to Skroch and López-Hoffman. Conservation Biology, 26(2):328–329, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Albert:2017:ANT

[ARDG17] Cécile H. Albert, Bronwyn Rayfield, Maria Dumitru, and Andrew Gonzalez. Applying network theory to prioritize multispecies habitat networks that are robust to climate and landuse change. Conservation Biology, 31(6):1383–1396, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armstrong:2016:BIA

[Arm16] Doug P. Armstrong. Biological invasions and the Austral Ark. Conservation Biology, 30(6):1366–1367, December 2016. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armsworth:2018:TDD

[Arm18] Paul R. Armsworth. Time discounting and the decision to protect areas that are near and threatened or remote and cheap to acquire. Conservation Biology, 32(5):1063–1073, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armstrong:2019:SCB

[Arm19] Chris Armstrong. Sharing conservation burdens fairly. Conservation Biology, 33(3):554–560, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aronson:2011:ERI

[Aro11] James Aronson. Ecological restoration in 2010: Insight from four continents. Conservation Biology, 25(1):206–208, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aswani:2010:IUA

[AS10] Shankar Aswani and Armagan Sabetian. Implications of urbanization for artisanal parrotfish fisheries in the Western Solomon Islands. Implicaciones de la urbanización para las pesquerías artesanales de peces loro en las Islas Salomón Occidentales. Conservation Biology, 24(2):520–530, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Albert:2013:UHL

[AS13] David M. Albert and John W. Schoen. Use of historical logging patterns to identify disproportionately logged ecosystems within temperate rainforests of southeastern Alaska. Conservation Biology, 27(4):774–784, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Abson:2011:VES

[AT11] David J. Abson and Mette Termansen. Valuing ecosystem services in terms of ecological risks and returns. Valoración de los servicios del ecosistema en términos de riesgos y beneficios ecológicos. Conservation Biology, 25(2):250–258, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allen:2011:VPS

[ATW^{+}11] Louise C. Allen, Amy S. Turmelle, Eric P. Widmaier, Nickolay I. Hristov, Gary F. Mccracken, and Thomas H. Kunz. Variation in physiological stress between bridge- and cave-roosting Brazilian free-tailed bats. Variación en el Estrés fisiológico de murciélagos Tadarida brasiliensis que perchan en puentes y cuevas. Conservation Biology, 26(3):374–381, April 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ames:2017:TSR

[AWHW17] Gregory M. Ames, Wade A. Wall, Matthew G. Hohmann, and Justin P. Wright. Trait space of rare plants in a fire-dependent ecosystem. Conservation Biology, 31(4):903–911, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allebone-Webb:2011:UMD

[AWKR^{+}11] S. M. Allebone-Webb, N. F. Kümpel, J. Rist, G. Cowlishaw, J. M. Rowcliffe, and E. J. Milner-Gulland. Use of market data to assess bushmeat hunting sustainability in Equatorial Guinea. Utilización de datos de mercado para evaluar la sustentabilidad de la cacería de vida silvestre en Guinea Ecuatorial. Conservation Biology, 25(6):597–606, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Auerbach:2015:ETM

[AWT^{+}15] Nancy A. Auerbach, Kerrie A. Wilson, Ayesha I. T. Tulloch, Jonathan R. Rhodes, Jeffrey O. Hanson, and Hugh P. Possingham. Effects of threat management interactions on conservation priorities. Conservation Biology, 29(6):1626–1635, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Adams:2011:UCA

[AWU^{+}11] M. Adams, S. D. Wedderburn, P. J. Unmack, M. P. Hammer, and J. B. Johnson. Use of congeneric assessment to reveal the linked genetic histories of two threatened fishes in the Murray-Darling Basin, Australia. Utilización de la evaluación Congenérica para revelar las historias Genéticas de dos especies de peces amenazadas en la Cuenca Murray-Darling, Australia. Conservation Biology, 25(4):767–776, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aslan:2012:ENN

[AZCT12] Clare E. Aslan, Erika S. Zavaleta, Don Croll, and Bernie Tershy. Effects of native and non-native vertebrate mutualists on plants efectos de vertebrados mutualistas nativos y no nativos sobre plantas. Conservation Biology, 26(5):778–789, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aslan:2014:IRN

[AZT^{+}14] Clare E. Aslan, Erika S. Zavaleta, Bernie Tershy, Don Croll, and Robert H. Robichaux. Imperfect replacement of native species by non-native species as pollinators of endemic Hawaiian plants. Conservation Biology, 28(2):478–488, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bennett:2013:HIC

[BA13] Joseph R. Bennett and Peter Arcese. Human influence and classical biogeographic predictors of rare species occurrence. Conservation Biology, 27(2):417–421, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Byers:2011:UPT

[BAG^{+}11] James E. Byers, Irit Altman, Andrew M. Grosse, Todd C. Huspeni, and John C. Maerz. Using parasitic trematode larvae to quantify an elusive vertebrate host. Utilización de larvas de trematodos parásitos para cuantificar a un hospedero vertebrado elusivo. Conservation Biology, 25(1):85–93, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bhatia:2013:URR

[BAGM13] Saloni Bhatia, Vidya Athreya, Richard Grenyer, and David W. Macdonald. Understanding the role of representations of human–leopard conflict in Mumbai through media-content analysis. Conservation Biology, 27(3):588–594, June 2013. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baillie:2010:IYE

[Bai10] Jonathan E. M. Baillie. The 2010 international year of extinction. Conservation Biology, 24(5):1432–1434, October 2010.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baisre:2013:SBE

[Bai13] Julio A. Baisre. Shifting baselines and the extinction of the Caribbean monk seal. Conservation Biology, 28(4):927–935, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baisre:2016:UUA

[Bai16] Julio A. Baisre. The uncritical use of anecdotes in marine historical ecology: response to McClenachan et al. Conservation Biology, 30(1):228–229, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Broquet:2010:GBD

[BAJ^{+}10] Thomas Broquet, Sonia Angelone, Julie Jaquiery, Pierre Joly, Jean-Paul Lena, Thierry Lengagne, Sandrine Plenet, Emilien Luquet, and Nicolas Perrin. Genetic bottlenecks driven by population disconnection. Cuellos de botella Genéticos producidos por la Desconexión de la población. Conservation Biology, 24 (6):1596–1605, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ban:2012:MSC

[Ban12] Natalie C. Ban. Mysteries of the sea continue to amaze. Conservation Biology, 26(6):1159–1160, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barrientos:2011:MAE

[BAPP11] Rafael Barrientos, Juan Carlos Alonso, Carlos Ponce, and Carlos Palacín. Meta-analysis of the effectiveness of marked wire in reducing avian collisions with power lines. Meta–Análisis sobre la eficacia de la señalización de los cables para reducir las colisiones de aves contra tendidos eléctricos. Conservation Biology, 25(5):893–903, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barney:2012:OFT

[Bar12] Amanda Barney. One fish or two fish, red fish or blue fish. Conservation Biology, 26(2):372–373, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brandt:2017:ENF

[BARB17] Jodi S. Brandt, Teri Allendorf, Volker Radeloff, and Jeremy Brooks. Effects of national forest-management regimes on unprotected forests of the Himalaya. Conservation Biology, 31 (6):1271–1282, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beier:2010:ULF

[BB10a] Paul Beier and Brian Brost. Use of land facets to plan for climate change: Conserving the arenas, not the actors. Uso de elementos territoriales para planificar para el cambio climático: Conservando las arenas, no los actores. Conservation Biology, 24(3):701–710, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:2010:SPE

[BB10b] Joel Berger and Jon P. Beckmann. Sexual predators, energy development, and conservation in Greater Yellowstone. Depredadores sexuales, desarrollo energético y conservación en Gran Yellowstone. Conservation Biology, 25(6):891–896, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:2011:IBS

[BB11a] Joel Berger and Jon P. Beckmann. The intersection between social change and conservation in the New American West. Conservation Biology, 25(6):633–634, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:2011:EFA

[BB11b] Joel Berger and Jon P. Beckmann. On evidence and faulty assertions about energy boomtowns. Conservation Biology, 25 (6):637–638, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brook:2015:KRN

[BB15] Barry W. Brook and Corey J. A. Bradshaw. Key role for nuclear energy in global biodiversity conservation. Conservation Biology, 29(3):702–712, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bouyoucos:2014:PEM

[BBB14] Ian Bouyoucos, Peter Bushnell, and Richard Brill. Potential for electropositive metal to reduce the interactions of Atlantic sturgeon with fishing gear. Conservation Biology, 28(1):278–282, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brammer:2016:RDD

[BBB^{+}16] Jeremy R. Brammer, Nicolas D. Brunet, A. Cole Burton, Alain Cuerrier, Finn Danielsen, Kanwaljeet Dewan, Thora Martina Herrmann, Micha V. Jackson, Rod Kennett, Guillaume Larocque, Monica Mulrennan, Arun Kumar Pratihast, Marie Saint-Arnaud, Colin Scott, and Murray M. Humphries. The role of digital data entry in participatory environmental monitoring. Conservation Biology, 30(6):1277–1287, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bell:2010:IPD

[BBBG10] C. D. Bell, J. M. Blumenthal, A. C. Broderick, and B. J. Godley. Investigating potential for depensation in marine turtles: How low can you go? Investigando el potencial de anticompensación en tortugas marinas: `?Qué tan bajo se puede llegar? Conservation Biology, 26(2):226–235, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Batsaikhan:2014:CWF

[BBC^{+}14] Nyamsuren Batsaikhan, Bayarbaatar Buuveibaatar, Bazaar Chimed, Oidov Enkhtuya, Davaa Galbrakh, Oyunsaikhan Ganbaatar, Badamjav Lkhagvasuren, Dejid Nandintsetseg, Joel Berger, Justin M. Calabrese, Ann E. Edwards, William F. Fagan, Todd K. Fuller, Michael Heiner, Takehiko Y. Ito, Petra Kaczensky, Peter Leimgruber, Anna Lushchekina, E. J. Milner-Gulland, Thomas Mueller, Martyn G. Murray, Kirk A. Olson, Richard Reading, George B. Schaller, Annagret Stubbe, Michael Stubbe, Chris Walzer, Henrik Von Wehrden, and Tony Whitten. Conserving the world’s finest grassland amidst ambitious national development. Conservation Biology, 28(6):1736– 1739, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burgar:2019:ICM

[BBF19] Joanna M. Burgar, A. Cole Burton, and Jason T. Fisher. The importance of considering multiple interacting species for conservation of species at risk. Conservation Biology, 33(3):709– 715, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Belant:2010:I

[BBGL10] Jerrold L. Belant, Erik A. Beever, John E. Gross, and Joshua J. Lawler. Introduction. Conservation Biology, 24(1):7–9, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Breininger:2017:ESL

[BBH17] David R. Breininger, Robert D. Breininger, and Carlton R. Hall. Effects of surrounding land use and water depth on seagrass dynamics relative to a catastrophic algal bloom. Conservation Biology, 31(1):67–75, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bland:2017:TRD

[BBK^{+}17] Lucie M. Bland, Jon Bielby, Stephen Kearney, C. David L. Orme, James E. M. Watson, and Ben Collen. Toward reassessing data-deficient species. Conservation Biology, 31(3):531– 539, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:2013:GCM

[BBM13] Joel Berger, Bayarbaatar Buuveibaatar, and Charudutt Mishra. Globalization of the cashmere market and the decline of large mammals in Central Asia. Conservation Biology, 27 (4):679–689, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bush:2014:GTE

[BBM14] Emma R. Bush, Sandra E. Baker, and David W. Macdonald. Global trade in exotic pets 2006–2012. Conservation Biology, 28(3):663–676, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:2015:CCB

[BBM15] Joel Berger, Bayarbaatar Buuveibaatar, and Charudutt Mishra. The cashmere connection, biodiversity, and climate: response to von Wehrden et al. 2014. Conservation Biology, 29(1):290–292, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Biro:2018:ULT

[BBM18] Marianna Biró, János Bölöni, and Zsolt Molnár. Use of longterm data to evaluate loss and endangerment status of Natura 2000 habitats and effects of protected areas. Conservation Biology, 32(3):660–671, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blickley:2012:EEE

[BBP12] Jessica L. Blickley, Diane Blackwood, and Gail L. Patricelli. Experimental evidence for the effects of chronic anthropogenic noise on abundance of greater sage-grouse at leks. Evidencia experimental de los efectos de ruido antropogénico crónico sobre la abundancia de Centrocercus urophasianus en leks. Conservation Biology, 26(3):461–471, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Benda-Beckmann:2014:MEG

[BBWK^{+}14] Alexander M. Von Benda-Beckmann, Paul J. Wensveen, Petter H. Kvadsheim, Frans-Peter A. Lam, Patrick J. O. Miller, Peter L. Tyack, and Michael A. Ainslie. Modeling effectiveness of gradual increases in source level to mitigate effects of sonar on marine mammals. Conservation Biology, 28(1):119–128, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:2014:MBS

[BC14a] Joel Berger and Steven L. Cain. Moving beyond science to protect a mammalian migration corridor. Conservation Biology, 28(5):1142–1150, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Black:2014:PPK

[BC14b] Simon A. Black and Jamieson A. Copsey. Purpose, process, knowledge, and dignity in interdisciplinary projects. Conservation Biology, 28(5):1139–1141, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bohm:2015:TEB

[BC15] Monika Böhm and Ben Collen. Toward equality of biodiversity knowledge through technology transfer. Conservation Biology, 29(5):1290–1302, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brown:2016:MAR

[BC16] Leone M. Brown and Elizabeth E. Crone. Minimum area requirements for an at-risk butterfly based on movement and demography. Conservation Biology, 30(1):103–112, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bury:2018:RMS

[BC18] Stanisław Bury and Mariusz Cichoń. Resemblance of a model species and its mimic: Reply to Valkonen and Mappes 2014. Conservation Biology, 32(2):484–486, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [VM14].

Bowne:2018:EUP

[BCA^{+}18] David R. Bowne, Bradley J. Cosentino, Laurel J. Anderson, Christopher P. Bloch, Sandra Cooke, Patrick W. Crumrine, Jason Dallas, Alexandra Doran, Jerald J. Dosch, Daniel L. Druckenbrod, Richard D. Durtsche, Danielle Garneau, Kristen S. Genet, Todd S. Fredericksen, Peter A. Kish, Mary Beth Kolozsvary, Frank T. Kuserk, Erin S. Lindquist, Carol Mankiewicz, James G. March, Timothy J. Muir, K. Greg Murray, Madeline N. Santulli, Frank J. Sicignano, Peter D. Smallwood, Rebecca A. Urban, Kathy Winnett-Murray, and Craig R. Zimmermann. Effects of urbanization on the population structure of freshwater turtles across the United States. Conservation Biology, 32(5):1150–1161, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bowne:2019:BAU

[BCA^{+}19] David R. Bowne, Bradley J. Cosentino, Laurel J. Anderson, Sandra Cooke, Jerald J. Dosch, Danielle Garneau, Todd S. Fredericksen, Mary Beth Kolozsvary, Erin S. Lindquist, James G. March, Rebecca A. Urban, and Craig R. Zimmermann. A broader approach to understanding urbanization effects on freshwater turtles: reply to Lambert and Steen 2019. Conservation Biology, 33(5):1197–1199, October 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:2014:OCS

[BCC^{+}14] Joel Berger, Steven L. Cain, Ellen Cheng, Peter Dratch, Kevin Ellison, John Francis, Herbert C. Frost, Scott Gende, Craig Groves, William A. Karesh, Elaine Leslie, Gary Machlis, Rodrigo A. Medellin, Reed F. Noss, Kent H. Redford, Michael Soukup, David Wilcove, and Steve Zack. Optimism and challenge for science-based conservation of migratory species in and out of U.S. national parks. Conservation Biology, 28(1):4–12, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bax:2016:REC

[BCD^{+}16] Nicholas J. Bax, Jesse Cleary, Ben Donnelly, Daniel C. Dunn, Piers K. Dunstan, Mike Fuller, and Patrick N. Halpin. Results of efforts by the convention on biological diversity to describe ecologically or biologically significant marine areas. Conservation Biology, 31(1):571–581, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brewer:2013:EHP

[BCGP13] T. D. Brewer, J. E. Cinner, A. Green, and R. L. Pressey. Effects of human population density and proximity to markets on coral reef fishes vulnerable to extinction by fishing. Conservation Biology, 27(3):443–452, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bogucki:2019:ADL

[BCK^{+}19] Robert Bogucki, Marek Cygan, Christin Brangwynne Khan, Maciej Klimek, Jan Kanty Milczek, and Marcin Mucha. Applying deep learning to right whale photo identification. Conservation Biology, 33(3):676–684, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brichieri-Colombi:2019:LCR

[BCLMM19] Typhenn A. Brichieri-Colombi, Natasha A. Lloyd, Jana M. McPherson, and Axel Moehrenschlager. Limited contributions of released animals from zoos to North American conservation translocations. Conservation Biology, 33(2):33–39, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brichieri-Colombi:2016:ATE

[BCM16] Typhenn A. Brichieri-Colombi and Axel Moehrenschlager. Alignment of threat, effort, and perceived success in North American conservation translocations. Conservation Biology, 30(6):1159–1172, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bland:2015:PCS

[BCOB15] Lucie M. Bland, Ben Collen, C. David L. Orme, and Jon Bielby. Predicting the conservation status of data-deficient species. Conservation Biology, 29(1):250–259, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Biggs:2017:DTC

[BCR^{+}17] Duan Biggs, Rosie Cooney, Dilys Roe, Holly T. Dublin, James R. Allan, Dan W. S. Challender, and Diane Skinner. Developing a theory of change for a community-based response to illegal wildlife trade. Conservation Biology, 31(1):5–12, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Budy:2015:OBQ

[BCSM15] Phaedra Budy, Mary M. Conner, Nira L. Salant, and William W. Macfarlane. An occupancy-based quantification of the highly imperiled status of desert fishes of the southwestern United States. Conservation Biology, 29(4):1142–1152, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bodin:2014:CSF

[BCT^{+}14] ÖRjan Bodin, Beatrice Crona, Matilda Thyresson, Anna-Lea Golz, and Maria Tengö. Conservation success as a function of good alignment of social and ecological structures and processes. Conservation Biology, 28(6):1371–1379, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beier:2015:EDS

[BdA15] Paul Beier and Fábio Suzart de Albuquerque. Environmental diversity as a surrogate for species representation. Conservation Biology, 29(5):1401–1410, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blickley:2013:GSG

[BDG^{+}13] Jessica L. Blickley, Kristy Deiner, Kelly Garbach, Iara Lacher, Mariah H. Meek, Lauren M. Porensky, Marit L. Wilkerson, Eric M. Winford, and Mark W. Schwartz. Graduate student’s guide to necessary skills for nonacademic conservation careers. Conservation Biology, 27(1):24–34, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Busch:2012:CCC

[BDH^{+}12] Jonah Busch, Radhika Dave, Lee Hannah, Alison Cameron, Andriambolantsoa Rasolohery, Patrick Roehrdanz, and George Schatz. Climate change and the cost of conserving species in Madagascar. Cambio climático y el costo de la conservación de especies en Madagascar. Conservation Biology, 26(3):408– 419, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Batary:2015:RAE

[BDKS15] Péter Batáry, Lynn V. Dicks, David Kleijn, and William J. Sutherland. The role of agri-environment schemes in conservation and environmental management. Conservation Biology, 29(4):1006–1016, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bayard:2010:HAS

[BE10] Trina S. Bayard and Chris S. Elphick. How area sensitivity in birds is studied. Cómo se estudia la sensibilidad al Área en las aves. Conservation Biology, 24(4):938–947, August 2010. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bunnefeld:2013:IMC

[BEA^{+}13] Nils Bunnefeld, Charles T. T. Edwards, Anagaw Atickem, Fetene Hailu, and E. J. Milner-Gulland. Incentivizing monitoring and compliance in trophy hunting. Conservation Biology, 27(6):1344–1354, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bolliger:2011:BFR

[BEE^{+}11] Janine Bolliger, Thomas C. Edwards, Jr., Stefan Eggenberg, Sascha Ismail, Irmi Seidl, and Felix Kienast. Balancing forestregeneration probabilities and maintenance costs in dry grasslands of high conservation priority. Balanceando las probabilidades de Regeneración del bosque y los costos de mantenimiento en pastizales secos de alta prioridad de conservación. Conservation Biology, 25(6):567–576, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beebee:2013:ERM

[Bee13] Trevor J. C. Beebee. Effects of road mortality and mitigation measures on amphibian populations. Conservation Biology, 27 (4):657–668, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beier:2019:RTW

[Bei19] Paul Beier. A rule of thumb for widths of conservation corridors. Conservation Biology, 33(4):976–978, August 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bennett:2012:SAC

[Ben12] Andrew F. Bennett. From the science to the application of connectivity conservation. Conservation Biology, 26(3):577– 578, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bennett:2015:LIT

[Ben15a] Elizabeth L. Bennett. Legal ivory trade in a corrupt world and its impact on African elephant populations. Conservation Biology, 29(1):54–60, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bennett:2015:CCC

[Ben15b] Nathan J. Bennett. Cover caption: Conservation Biology. Conservation Biology, 29(1):i, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bennett:2015:GMP

[Ben15c] Nathan J. Bennett. Governing marine protected areas in an interconnected and changing world. Conservation Biology, 29 (1):303–306, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bennett:2016:UPE

[Ben16] Nathan James Bennett. Using perceptions as evidence to improve conservation and environmental management. Conservation Biology, 31(1):582–592, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:2012:EBS

[Ber12] Joel Berger. Estimation of body-size traits by photogrammetry in large mammals to inform conservation. Estimación de atributos de tamaño corporal por Fotogrametría en mamíferos mayores para informar a la conservación. Conservation Biology, 26(5):769–777, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berteaux:2013:QLS

[Ber13] Dominique Berteaux. Québec’s large-scale Plan Nord. Conservation Biology, 27(2):242–243, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brennan:2019:ITC

[BFB^{+}19] Sean R. Brennan, Diego P. Fernandez, Jennifer M. Burns, Stephanie Aswad, Daniel E. Schindler, and Thure E. Cerling. Isotopes in teeth and a cryptic population of coastal freshwater seals. Conservation Biology, 33(6):1415–1425, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barreto:2016:TES

[BFF^{+}16] Rodrigo Barreto, Francesco Ferretti, Joanna M. Flemming, Alberto Amorim, Humber Andrade, Boris Worm, and Rosangela Lessa. Trends in the exploitation of South Atlantic shark populations. Conservation Biology, 30(4):792–804, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bombaci:2016:UTC

[BFG^{+}16] Sara P. Bombaci, Cooper M. Farr, H. Travis Gallo, Anna M. Mangan, Lani T. Stinson, Monica Kaushik, and Liba Pejchar. Using Twitter to communicate conservation science from a professional conference. Conservation Biology, 30(1):216–225, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Becker:2010:HSC

[BFHP10] C. Guilherme Becker, Carlos R. Fonseca, Célio F. B. Haddad, and Paulo I. Prado. Habitat split as a cause of local population declines of amphibians with aquatic larvae. Separación del hábitat como una causa de declinaciones poblacionales locales de anfibios con larvas acuáticas. Conservation Biology, 26(2): 287–294, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Black:2010:UBE

[BG10] Simon Black and Jim Groombridge. Use of a business excellence model to improve conservation programs. Utilización de un modelo de excelencia empresarial para mejorar programas de conservación. Conservation Biology, 24(6):1448–1458, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barlow:2010:UAS

[BGAS10] Adam C. D. Barlow, Christina J. Greenwood, Ishtiaq U. Ahmad, and James L. D. Smith. Use of an action-selection framework for human-carnivore conflict in the Bangladesh Sundarbans. Uso de un marco de selección de acciones en un conflicto humanos–carnívoros en el Sundabarns de Bangladesh. Conservation Biology, 24(5):1338–1347, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baker:2017:EEM

[BGB17] Christopher M. Baker, Ascelin Gordon, and Michael Bode. Ensemble ecosystem modeling for predicting ecosystem response to predator reintroduction. Conservation Biology, 31(3):376– 384, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brodie:2015:EML

[BGD^{+}15] Jedediah F. Brodie, Anthony J. Giordano, Brett Dickson, Mark Hebblewhite, Henry Bernard, Jayasilan Mohd-Azlan, Jesse Anderson, and Laurentius Ambu. Evaluating multispecies landscape connectivity in a threatened tropical mammal community. Conservation Biology, 29(1):122–132, February 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Busch:2013:EET

[BGG13] D. Shallin Busch, Correigh M. Greene, and Thomas P. Good. Estimating effects of tidal power projects and climate change on threatened and endangered marine species and their food Web. Conservation Biology, 27(6):1190–1200, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bradley:2015:DSF

[BGH^{+}15] Paul W. Bradley, Stephanie S. Gervasi, Jessica Hua, Rickey D. Cothran, Rick A. Relyea, Deanna H. Olson, and Andrew R. Blaustein. Differences in sensitivity to the fungal pathogen Batrachochytrium dendrobatidis among amphibian populations. Conservation Biology, 29(5):1347–1356, October 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bull:2014:IBS

[BGL^{+}14] J. W. Bull, A. Gordon, E. A. Law, K. B. Suttle, and E. J. Milner-Gulland. Importance of baseline specification in evaluating conservation interventions and achieving no net loss of biodiversity. Conservation Biology, 28(3):799–809, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baker:2019:CAN

[BGO^{+}19] David J. Baker, Stephen T. Garnett, James O’Connor, Glenn Ehmke, Rohan H. Clarke, John C. Z. Woinarski, and Melodie A. McGeoch. Conserving the abundance of nonthreatened species. Conservation Biology, 33(2):319–328, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bridge:2016:FII

[BGP16] Tom C. L. Bridge, Alana M. Grech, and Robert L. Pressey. Factors influencing incidental representation of previously unknown conservation features in marine protected areas. Conservation Biology, 30(1):154–165, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brodie:2015:CPH

[BGZ^{+}15] Jedediah F. Brodie, Anthony J. Giordano, Elise F. Zipkin, Henry Bernard, Jayasilan Mohd-Azlan, and Laurentius Ambu. Correlation and persistence of hunting and logging impacts on tropical rainforest mammals. Conservation Biology, 29(1): 110–121, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baumgaertner:2017:NVC

[BH17] Bert Baumgaertner and Wieteke Holthuijzen. On nonepistemic values in conservation biology. Conservation Biology, 31(1):48– 55, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brown:2018:EFP

[BH18] Christopher J. Brown and Richard J. Hamilton. Estimating the footprint of pollution on coral reefs with models of species turnover. Conservation Biology, 32(4):949–958, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beier:2015:I

[BHA15] Paul Beier, Malcolm L. Hunter, and Mark Anderson. Introduction. Conservation Biology, 29(3):613–617, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brauwer:2018:BST

[BHAR^{+}18] Maarten De Brauwer, Jean-Paul A. Hobbs, Rohani Ambo-Rappe, Jamaluddin Jompa, Euan S. Harvey, and Jennifer L. McIlwain. Biofluorescence as a survey tool for cryptic marine species. Conservation Biology, 32(3):706–715, June 2018. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brook:2019:PVH

[BHB^{+}19] Cara E. Brook, James P. Herrera, Cortni Borgerson, Emma C. Fuller, Pascal Andriamahazoarivosoa, B. J. Rodolph Rasolofoniaina, J. L. Rado Ravoavy Randrianasolo, Z. R. Eli Rakotondrafarasata, Hervet J. Randriamady, Andrew P. Dobson, and Christopher D. Golden. Population viability and harvest sustainability for Madagascar lemurs. Conservation Biology, 33 (2):99–111, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bowler:2019:LTD

[BHF^{+}19] Diana E. Bowler, Henning Heldbjerg, Anthony D. Fox, Maaike de Jong, and Katrin Böhning-Gaese. Long-term declines of European insectivorous bird populations and potential causes. Conservation Biology, 33(5):1120–1130, October 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bartz:2010:PDE

[BHK10] Robert Bartz, Ulrich Heink, and Ingo Kowarik. Proposed definition of environmental damage illustrated by the cases of genetically modified crops and invasive species. Propuesta de definición de daño ambiental ilustrada por los casos de cultivos modificados genéticamente y especies invasoras. Conservation Biology, 24(3):675–681, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Breckheimer:2014:DEU

[BHM^{+}14] Ian Breckheimer, Nick M. Haddad, William F. Morris, Anne M. Trainor, William R. Fields, R. Todd Jobe, Brian R. Hudgens, Aaron Moody, and Jeffrey R. Walters. Defining and evaluating the umbrella species concept for conserving and restoring landscape connectivity. Conservation Biology, 28(6):1584–1593, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Braczkowski:2018:RMW

[BHO^{+}18] Alexander Braczkowski, Matthew H. Holden, Christopher O’Bryan, Chi-Yeung Choi, Xiaojing Gan, Nicholas Beesley, Yufang Gao, James Allan, Peter Tyrrell, Daniel Stiles, Peadar Brehony, Revocatus Meney, Henry Brink, Nao Takashina, Ming-Ching Lin, Hsien-Yung Lin, Niki Rust, Severino G. Salmo III, James E. M. Watson, Paula Kahumbu, Martine Maron, Hugh P. Possingham, and Duan Biggs. Reach and messages of the world’s largest ivory burn. Conservation Biology, 32(4): 765–773, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brejao:2018:TRA

[BHPM^{+}18] Gabriel L. Brejão, David J. Hoeinghaus, María Angélica Pérez-Mayorga, Silvio F. B. Ferraz, and Lilian Casatti. Threshold responses of Amazonian stream fishes to timing and extent of deforestation. Conservation Biology, 32(4):860–871, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bilkovic:2012:UFB

[BHSA12] D. M. Bilkovic, K. J. Havens, D. M. Stanhope, and K. T. Angstadt. Use of fully biodegradable panels to reduce derelict pot threats to marine fauna. Conservation Biology, 26(6): 957–966, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boakes:2018:ERB

[BIF^{+}18] Elizabeth H. Boakes, Nicholas J. B. Isaac, Richard A. Fuller, Georgina M. Mace, and Philip J. K. McGowan. Examining the relationship between local extinction risk and position in range. Conservation Biology, 32(1):229–239, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bragina:2015:RDL

[BIP^{+}15] Eugenia V. Bragina, A. R. Ives, A. M. Pidgeon, T. Kuemmerle, L. M. Baskin, Y. P. Gubar, M. Piquer-Rodríguez, N. S. Keuler, V. G. Petrosyan, and V. C. Radeloff. Rapid declines of large mammal populations after the collapse of the Soviet Union. Conservation Biology, 29(3):844–853, June 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bellard:2016:SMB

[BJ16] C. Bellard and J. M. Jeschke. A spatial mismatch between invader impacts and research publications. Conservation Biology, 30(1):230–232, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burgman:2015:DGB

[BJM15] Mark Burgman, Frith Jarrad, and Ellen Main. Decreasing geographic bias in Conservation Biology. Conservation Biology, 29 (5):1255–1256, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buxton:2014:DSP

[BJMT14] Rachel T. Buxton, Christopher Jones, Henrik Moller, and David R. Towns. Drivers of seabird population recovery on New Zealand islands after predator eradication. Conservation Biology, 28(2):333–344, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Branch:2011:CGT

[BJR^{+}11] Trevor A. Branch, Olaf P. Jensen, Daniel Ricard, Yimin Ye, and Ray Hilborn. Contrasting global trends in marine fishery status obtained from catches and from stock assessments. Contraste de las tendencias globales en el estatus de las pesquerías marinas obtenido de capturas y evaluación de reservas. Conservation Biology, 25(4):777–786, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Belaire:2014:PMP

[BKKM14] J. Amy Belaire, Betty J. Kreakie, Timothy Keitt, and Emily Minor. Predicting and mapping potential whooping crane stopover habitat to guide site selection for wind energy projects. Conservation Biology, 28(2):541–550, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bleho:2014:DEC

[BKM14] Barbara I. Bleho, Nicola Koper, and Craig S. Machtans. Direct effects of cattle on grassland birds in Canada. Conservation Biology, 28(3):724–734, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brown:2017:FDA

[BKR17] Holly Milton Brown, Ambika Kamath, and Margaret Rubega. Facilitating discussions about privilege among future conservation practitioners. Conservation Biology, 31(3):727–730, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bluwstein:2017:UPE

[BL17] Jevgeniy Bluwstein and Jens Friis Lund. Unpacking the political economy of lion trophy hunting in Tanzania. Conservation Biology, 31(3):486–487, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bertelsmeier:2013:IQQ

[BLC13] Cleo Bertelsmeier, Gloria M. Luque, and Franck Courchamp. Increase in quantity and quality of suitable areas for invasive species as climate changes. Conservation Biology, 28(1):1458– 1467, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bradshaw:2011:HAG

[BLG^{+}11] Corey J. A. Bradshaw, William F. Laurance, Luke Gibson, Paul R. Ehrlich, and Barry W. Brook. Homage to an avantgarde conservation leader, navjot sodhi. Conservation Biology, 27(3):1056–1058, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burkhalter:2016:ECM

[BLM^{+}16] J. C. Burkhalter, J. L. Lockwood, B. Maslo, K. H. Fenn, and K. Leu. Effects of cost metric on cost-effectiveness of protectedarea network design in urban landscapes. Conservation Biology, 30(2):403–412, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bachman:2018:QPT

[BLR18] Steven P. Bachman, Eimear M. Nic Lughadha, and Malin C. Rivers. Quantifying progress toward a conservation assessment for all plants. Conservation Biology, 32(3):516–524, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bluwstein:2016:PDW

[Blu16] Jevgeniy Bluwstein. Problematizing debates on wildlife conservation and the war on poaching. Conservation Biology, 30 (4):692–693, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barber-Meyer:2010:DCN

[BM10] Shannon M. Barber-Meyer. Dealing with the clandestine nature of wildlife-trade market surveys haciendo frente a la naturaleza clandestina de los estudios de mercado del comercio de vida silvestre. Conservation Biology, 24(4):918–923, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blicharska:2014:ISC

[BM14] Malgorzata Blicharska and Grzegorz Mikusiński. Incorporating social and cultural significance of large old trees in conservation policy. Conservation Biology, 28(6):1558–1567, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Borker:2014:VAL

[BMA^{+}14] Abraham L. Borker, Matthew W. Mckown, Joshua T. Ackerman, Collin A. Eagles-Smith, Bernie R. Tershy, and Donald A. Croll. Vocal activity as a low cost and scalable index of seabird colony size. Conservation Biology, 28(4):1100–1108, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bodmer:2018:MSA

[BMA^{+}18] Richard Bodmer, Pedro Mayor, Miguel Antunez, Kimberlyn Chota, Tula Fang, Pablo Puertas, Marlini Pittet, Maire Kirkland, Mike Walkey, Claudia Rios, Pedro Perez-Peña, Peter Henderson, William Bodmer, Andy Bicerra, Joseph Zegarra, and Emma Docherty. Major shifts in Amazon wildlife populations from recent intensification of floods and drought. Conservation Biology, 32(2):333–344, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bower:2017:UBN

[BMAS17] D. S. Bower, K. Mengersen, R. A. Alford, and L. Schwarzkopf. Using a Bayesian network to clarify areas requiring research in a host–pathogen system. Conservation Biology, 31(6):1373–1382, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bell:2015:GCS

[BMB^{+}15] James J. Bell, Emily McGrath, Andrew Biggerstaff, Tracey Bates, César A. Cárdenas, and Holly Bennett. Global conservation status of sponges. Conservation Biology, 29(1):42–53, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Benoist:2019:MMB

[BMB^{+}19] Noëlie M. A. Benoist, Kirsty J. Morris, Brian J. Bett, Jennifer M. Durden, Veerle A. I. Huvenne, Tim P. Le Bas, Russell B. Wynn, Suzanne J. Ware, and Henry A. Ruhl. Monitoring mosaic biotopes in a marine conservation zone by autonomous underwater vehicle. Conservation Biology, 33(5):1174–1186, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buxton:2018:EEI

[BMC^{+}18] Rachel T. Buxton, Megan F. McKenna, Mary Clapp, Erik Meyer, Erik Stabenau, Lisa M. Angeloni, Kevin Crooks, and George Wittemyer. Efficacy of extracting indices from largescale acoustic recordings to monitor biodiversity. Conservation Biology, 33(1):1174–1184, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boessenkool:2014:UAS

[BME^{+}14] Sanne Boessenkool, Gayle Mcglynn, Laura S. Epp, David Taylor, Manuel Pimentel, Abel Gizaw, Sileshi Nemomissa, Christian Brochmann, and Magnus Popp. Use of ancient sedimentary DNA as a novel conservation tool for high-altitude tropical biodiversity. Conservation Biology, 28(2):446–455, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Breuer:2016:CPA

[BMF16] Thomas Breuer, Fiona Maisels, and Vicki Fishlock. The consequences of poaching and anthropogenic change for forest elephants. Conservation Biology, 30(5):1019–1026, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beever:2014:SCF

[BMG^{+}14] Erik A. Beever, Brady J. Mattsson, Matthew J. Germino, Max Post Van Der Burg, John B. Bradford, and Mark W. Brunson. Successes and challenges from formation to implementation of eleven broad-extent conservation programs. Conservation Biology, 28(2):302–314, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Butsic:2017:EPA

[BMG^{+}17] Van Butsic, Catalina Munteanu, Patrick Griffiths, Jan Knorn, Volker C. Radeloff, Juraj Lieskovský, Daniel Mueller, and Tobias Kuemmerle. The effect of protected areas on forest disturbance in the Carpathian Mountains 1985–2010. Conservation Biology, 31(3):570–580, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baker:2012:PGI

[BMGLW12] Julia Baker, E. J. Milner-Gulland, and Nigel Leader-Williams. Park gazettement and integrated conservation and development as factors in community conflict at Bwindi Impenetrable Forest, Uganda. Decreto de parques y conservación y desarrollo integrados como factores de conflictos entre comunidades en el Bosque Impenetrable Bwindi, Uganda. Conservation Biology, 26(2):160–170, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barros:2014:IAB

[BMM^{+}14] Alan E. De Barros, Ewan A. Macdonald, Marcelo H. Matsumoto, Rogério C. Paula, Sahil Nijhawan, Y. Malhi, and David W. Macdonald. Identification of areas in Brazil that optimize conservation of forest carbon, jaguars, and biodiversity. Conservation Biology, 28(2):580–593, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bottrill:2012:EPB

[BMP^{+}12] Madeleine C. Bottrill, Morena Mills, Robert L. Pressey, Edward T. Game, and Craig Groves. Evaluating perceived benefits of ecoregional assessments. Evaluación de los beneficios percibidos de las evaluaciones ecorregionales. Conservation Biology, 26(5):851–861, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boor:2014:FDO

[Boo14] Gina K. Himes Boor. A framework for developing objective and measurable recovery criteria for threatened and endangered species. Conservation Biology, 28(1):33–43, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Borges:2018:PIM

[Bor18] Paulo A. V. Borges. Putting islands on the map. Conservation Biology, 33(1):1214–1215, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Borges:2019:CIB

[Bor19] Paulo A. V. Borges. Conservation of island biota. Conservation Biology, 33(3):738–739, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boyer:2010:CES

[Boy10] Alison G. Boyer. Consistent ecological selectivity through time in Pacific Island avian extinctions. Selección ecológica consistente a lo largo del tiempo en extinciones de aves en Islas del Pacíficon. Conservation Biology, 24(2):511–519, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boughton:2013:FRH

[BP13] David A. Boughton and Andrew S. Pike. Floodplain rehabilitation as a hedge against hydroclimatic uncertainty in a migration corridor of threatened steelhead. Conservation Biology, 27(6):1158–1168, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bouget:2016:ELD

[BP16] C. Bouget and G. Parmain. Effects of landscape design of forest reserves on saproxylic beetle diversity. Conservation Biology, 30(1):92–102, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bart:2013:INP

[BPA^{+}13] Jonathan Bart, Robert M. Platte, Brad Andres, Stephen Brown, James A. Johnson, and William Larned. Importance of the national petroleum reserve–Alaska for aquatic birds. Conservation Biology, 27(6):1304–1312, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bernard:2014:DDD

[BPA14] E. Bernard, L. A. O. Penna, and E. Araújo. Downgrading, downsizing, degazettement, and reclassification of protected areas in Brazil. Conservation Biology, 28(4):939–950, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Botts:2019:PAA

[BPH^{+}19] Emily A. Botts, Genevieve Pence, Stephen Holness, Kerry Sink, Andrew Skowno, Amanda Driver, Linda R. Harris, Philip Desmet, Boyd Escott, Mervyn Lötter, Jeanne Nel, Tammy Smith, Fahiema Daniels, Samuel Sinclair, Warrick Stewart, and Jeffrey Manuel. Practical actions for applied systematic conservation planning. Conservation Biology, 33(6):1235–1246, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brodie:2016:CSP

[BPN^{+}16] Jedediah F. Brodie, Midori Paxton, Kangayatkarasu Nagulendran, G. Balamurugan, Gopalasamy Reuben Clements, Glen Reynolds, Anuj Jain, and Jason Hon. Connecting science, policy, and implementation for landscape-scale habitat connectivity. Conservation Biology, 30(5):950–961, October 2016. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bode:2011:CPM

[BPT^{+}11] Michael Bode, Will Probert, Will R. Turner, Kerrie A. Wilson, and Oscar Venter. Conservation planning with multiple organizations and objectives. Planificación de la conservación con múltiples organizaciones y objetivos. Conservation Biology, 26 (3):295–304, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ban:2012:COS

[BPW12] Natalie C. Ban, Robert L. Pressey, and Scarla Weeks. Conservation objectives and sea-surface temperature anomalies in the Great Barrier Reef. Objetivos de conservación y Anomalías en la temperatura de la superficie del mar en la Gran Barrera Arrecifal. Conservation Biology, 26(5):799–809, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bela:2016:LTP

[BPY^{+}16] Györgyi Bela, Taru Peltola, Juliette C. Young, Bálint Balázs, Isabelle Arpin, György Pataki, Jennifer Hauck, Eszter Kelemen, Leena Kopperoinen, Ann Van Herzele, Hans Keune, Susanne Hecker, Monika Suškevičs, Helen E. Roy, Pekka Itkonen, Mart Külvik, Miklós László, Corina Basnou, Joan Pino, and Aletta Bonn. Learning and the transformative potential of citizen science. Conservation Biology, 30(5):990–999, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blackman:2011:PLB

[BR11a] Allen Blackman and Jorge Rivera. Producer-level benefits of sustainability certification. Beneficios de la certificación de la sustentabilidad a nivel de productores. Conservation Biology, 25(6):1176–1185, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Branton:2011:AVU

[BR11b] Margaret Branton and John S. Richardson. Assessing the value of the umbrella-species concept for conservation planning with meta-analysis. Evaluación del valor del concepto de especie sombrilla para la planificación de la conservación mediante meta–análisis. Conservation Biology, 25(1):9–20, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brandmayr:2015:BRE

[Bra15] Pietro Brandmayr. Book review: Exploring a Poorly Known Ecological Domain Shallow Subterranean Habitats. Ecology, Evolution, and Conservation. Culver, D. C. and Pipan, T.. 2014. Oxford University Press, NY, U.S.A. 258 pp. $ 99.95 (hardcover). ISBN 978-0-19-964617-3. Conservation Biology, 29(6):1730, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Braza:2017:ECE

[Bra17] Mark Braza. Effectiveness of conservation easements in agricultural regions. Conservation Biology, 31(4):848–859, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brandt:2019:EEF

[BRA^{+}19] Jodi S. Brandt, Volker Radeloff, Teri Allendorf, Van Butsic, and Anand Roopsind. Effects of ecotourism on forest loss in the Himalayan biodiversity hotspot based on counterfactual analyses. Conservation Biology, 33(6):1318–1328, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bryan:2011:CSE

[BRCK11] Brett Anthony Bryan, Christopher Mark Raymond, Neville David Crossman, and Darran King. Comparing spatially explicit ecological and social values for natural areas to identify effective conservation strategies. Comparación de valores ecológicos y sociales espacialmente explícitos de Áreas naturales para la identificación de estrategias de conservación efectivas. Conservation Biology, 25(1):172–181, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bennett:2017:MSS

[BRK^{+}17] Nathan J. Bennett, Robin Roth, Sarah C. Klain, Kai M. A. Chan, Douglas A. Clark, Georgina Cullman, Graham Epstein, Michael Paul Nelson, Richard Stedman, Tara L. Teel, Rebecca E. W. Thomas, Carina Wyborn, Deborah Curran, Alison Greenberg, John Sandlos, and Diogo Veríssimo. Mainstreaming the social sciences in conservation. Conservation Biology, 31 (1):56–66, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Butsic:2012:AST

[BRKP12] Van Butsic, Volker C. Radeloff, Tobias Kuemmerle, and Anna M. Pidgeon. Analytical solutions to trade-offs between size of protected areas and land-use intensity. Soluciones analíticas a las ventajas y desventajas entre el tamaño de Áreas protegidas y la intensidad del uso de suelo. Conservation Biology, 26(5):883–893, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burgman:2014:VSE

[BRM^{+}14] Mark A. Burgman, Helen M. Regan, Lynn A. Maguire, Mark Colyvan, James Justus, Tara G. Martin, and Kris Rothley. Voting systems for environmental decisions. Conservation Biology, 28(2):322–332, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brooks:2010:ESD

[Bro10] Jeremy S. Brooks. Economic and social dimensions of environmental behavior: Balancing conservation and development in Bhutan. Dimensiones económicas y sociales del comportamiento ambiental: Balanceando conservación y desarrollo en Bután. Conservation Biology, 24(6):1499–1509, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brook:2013:BRB

[Bro13] Ryan K. Brook. A big resource with big gaps. Conservation Biology, 27(3):639–641, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bawa:2011:RGC

[BRS11] Kamaljit S. Bawa, Nitin D. Rai, and Navjot S. Sodhi. Rights, governance, and conservation of biological diversity. Conservation Biology, 25(6):639–641, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comment [KNK13].

Barbosa:2010:UCR

[BRV10] A. Márcia Barbosa, Raimundo Real, and J. Mario Vargas. Use of coarse-resolution models of species’ distributions to guide local conservation inferences. Uso de modelos de resolución gruesa de las distribuciones de especies para guiar inferencias de conservación locales. Conservation Biology, 24(5):1378–1387, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brand:2010:VAH

[BS10] Adrianne B. Brand and Joel W. Snodgrass. Value of artificial habitats for amphibian reproduction in altered landscapes. El valor de hábitats artificiales para la reproducción de anfibios en paisajes alterados. Conservation Biology, 26(2):295–301, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bradford:2012:MWA

[BS12] Lori Bradford and Graham E. H. Strickert. Managing water across boundaries. Conservation Biology, 26(5):950–952, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boitani:2015:I

[BS15] Luigi Boitani and William J. Sutherland. Introduction. Conservation Biology, 29(4):975–977, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beier:2011:TBP

[BSBM11] Paul Beier, Wayne Spencer, Robert F. Baldwin, and Brad H. Mcrae. Toward best practices for developing regional connectivity maps. Hacia mejores prácticas para desarrollar mapas de conectividad regional. Conservation Biology, 25(5):879–892, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brook:2018:DEE

[BSCB18] Barry W. Brook, Stephen R. Sleightholme, Cameron R. Campbell, and Jessie C. Buettel. Deficiencies in estimating the extinction date of the thylacine with mixed certainty data. Conservation Biology, 33(1):1195–1197, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bonebrake:2014:FMM

[BSF^{+}14] Timothy C. Bonebrake, Alexandra D. Syphard, Janet Franklin, Kurt E. Anderson, H. Resit Akçakaya, Toni Mizerek, Clark Winchell, and Helen M. Regan. Fire management, managed relocation, and land conservation options for long-lived obligate seeding plants under global changes in climate, urbanization, and fire regime. Conservation Biology, 28(4):1057–1067, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beier:2015:RSB

[BSH^{+}15] Paul Beier, Patricia Sutcliffe, Jan Hjort, Daniel P. Faith, Robert L. Pressey, and Fabio Albuquerque. A review of selection-based tests of abiotic surrogates for species representation. Conservation Biology, 29(3):668–679, June 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bucciarelli:2019:AEN

[BSL^{+}19] Gary M. Bucciarelli, Daniel Suh, Avery Davis Lamb, Dave Roberts, Debra Sharpton, H. Bradley Shaffer, Robert N. Fisher, and Lee B. Kats. Assessing effects of non-native crayfish on mosquito survival. Conservation Biology, 33(2):122–131, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brown:2014:UCS

[BT14] Christopher J. Brown and Rowan Trebilco. Unintended cultivation, shifting baselines, and conflict between objectives for fisheries and conservation. Conservation Biology, 28(3):677– 688, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burivalova:2018:USD

[BTB^{+}18] Zuzana Burivalova, Michael Towsey, Tim Boucher, Anthony Truskinger, Cosmas Apelis, Paul Roe, and Edward T. Game.

Using soundscapes to detect variable degrees of human influence on tropical forests in Papua New Guinea. Conservation Biology, 32(1):205–215, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger-Tal:2016:SSI

[BTBC^{+}16] Oded Berger-Tal, Daniel T. Blumstein, Scott Carroll, Robert N. Fisher, Sarah L. Mesnick, Megan A. Owen, David Saltz, Colleen Cassady St. Claire, and Ronald R. Swaisgood. A systematic survey of the integration of animal behavior into conservation. Conservation Biology, 30(4):744–753, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger-Tal:2012:EMR

[BTBDS12] Oded Berger-Tal, Shirli Bar-David, and David Saltz. Effectiveness of multiple release sites in reintroduction of Persian fallow deer. Efectividad de múltiples sitios de liberación en la reintroducció del gamo Persa. Conservation Biology, 26(2):107–115, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bode:2015:CPA

[BTM^{+}15] Michael Bode, Ayesha I. T. Tulloch, Morena Mills, Oscar Venter, and Amy W. Ando. A conservation planning approach to mitigate the impacts of leakage from protected area networks. Conservation Biology, 29(3):765–774, June 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brehme:2013:PRM

[BTMF13] Cheryl S. Brehme, Jeff A. Tracey, Leroy R. Mcclenaghan, and Robert N. Fisher. Permeability of roads to movement of scrubland lizards and small mammals. Conservation Biology, 27(4): 710–720, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buckley:2017:HSR

[BTTP17] Sarah M. Buckley, Ruth H. Thurstan, Andrew Tobin, and John M. Pandolfi. Historical spatial reconstruction of a spawning-aggregation fishery. Conservation Biology, 31(6): 1322–1332, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burgman:2013:SFC

[Bur13] Mark Burgman. Shaping the future of Conservation Biology. Conservation Biology, 27(4):643, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burgman:2014:I

[Bur14] Mark A. Burgman. Introduction. Conservation Biology, 28(3): 633, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burgman:2019:OAA

[Bur19] Mark Burgman. Open access and academic imperialism. Conservation Biology, 33(1):5–6, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buschke:2015:SCC

[Bus15] Falko T. Buschke. The startup culture of conservation entrepreneurship. Conservation Biology, 29(1):300–302, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bjorkman:2010:DHB

[BV10] Anne D. Bjorkman and Mark Vellend. Defining historical baselines for conservation: Ecological changes since European settlement on Vancouver Island, Canada. Definición de Líneas de base históricas para la conservación: Cambios ecológicos desde el asentamiento europeo en la Isla Vancouver, Canadá. Conservation Biology, 24(6):1559–1568, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bowker:2017:EAT

[BVAC17] J. N. Bowker, A. De Vos, J. M. Ament, and G. S. Cumming. Effectiveness of Africa’s tropical protected areas for maintaining forest cover. Conservation Biology, 31(3):559–569, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Balaz:2014:ARG

[BVC^{+}14] Vojtech Baláž, Judit Vörös, Petr Civiš, Jiri Vojar, Attila Hettyey, Endre Sós, Róbert Dankovics, Robert Jehle, Ditte G. Christiansen, Frances Clare, Matthew C. Fisher, Trenton W. J.

Garner, and Jon Bielby. Assessing risk and guidance on monitoring of Batrachochytrium dendrobatidis in Europe through identification of taxonomic selectivity of infection. Conservation Biology, 28(1):213–223, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bode:2010:HBE

[BW10] Michael Bode and Brendan Wintle. How to build an efficient conservation fence. Cómo construir un cerco de conservación eficiente. Conservation Biology, 24(1):182–188, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baskett:2013:EAS

[BW13] Marissa L. Baskett and Robin S. Waples. Evaluating alternative strategies for minimizing unintended fitness consequences of cultured individuals on wild populations. Conservation Biology, 27(1):83–94, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Balding:2016:PBI

[BW16] Mung Balding and Kathryn J. H. Williams. Plant blindness and the implications for plant conservation. Conservation Biology, 30(6):1192–1199, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brown:2019:PCS

[BW19] Eleanor D. Brown and Byron K. Williams. The potential for citizen science to produce reliable and useful information in ecology. Conservation Biology, 33(3):561–569, June 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brainard:2013:ICO

[BWE^{+}13] Russell E. Brainard, Mariska Weijerman, C. Mark Eakin, Paul Mcelhany, Margaret W. Miller, Matt Patterson, Gregory A. Piniak, Matthew J. Dunlap, and Charles Birkeland. Incorporating climate and ocean change into extinction risk assessments for 82 coral species. Conservation Biology, 27(6):1169– 1178, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coad:2010:DUI

[CAB^{+}10] L. Coad, K. Abernethy, A. Balmford, A. Manica, L. Airey, and E. J. Milner-Gulland. Distribution and use of income from bushmeat in a rural village, Central Gabon. Distribución y uso del ingreso por carne de monte en una aldea rural, Gabón Central. Conservation Biology, 24(6):1510–1518, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cheng:2018:UML

[CAB^{+}18] S. H. Cheng, C. Augustin, A. Bethel, D. Gill, S. Anzaroot, J. Brun, B. DeWilde, R. C. Minnich, R. Garside, Y. J. Masuda, D. C. Miller, D. Wilkie, S. Wongbusarakum, and M. C. McKinnon. Using machine learning to advance synthesis and use of conservation and environmental evidence. Conservation Biology, 32(4):762–764, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carmichael:2011:ECF

[Car11] Cassandra Carmichael. The environmental crisis finds religion. Conservation Biology, 27(3):1059–1060, October 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2014:CCO

[Car14] Carlos Carroll. Can a conservation-oriented scientific society remain relevant in the 21st Century? Conservation Biology, 28 (5):1137–1138, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Catenazzi:2011:TCE

[Cat11] Alessandro Catenazzi. Temperature constraint of elevational range of tropical amphibians: Response to Forero-Medina et al. Conservation Biology, 25(3):425, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Catalano:2018:EDS

[Cat18] Chiara Catalano. Ecological design to shape new urbanities. Conservation Biology, 32(3):752–754, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Calver:2010:PCO

[CB10] Michael C. Calver and J. Stuart Bradley. Patterns of citations of open access and non-open access Conservation Biology journal papers and book chapters. Patrones de citas de artículos y capítulos de libros de biología de la conservación de acceso abierto y acceso no abierto. Conservation Biology, 25(6):872– 880, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chapple:2013:CLD

[CB13] Taylor K. Chapple and Louis W. Botsford. A comparison of linear demographic models and fraction of lifetime egg production for assessing sustainability in sharks. Conservation Biology, 27 (3):560–568, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coverdale:2013:RON

[CBA13] Tyler C. Coverdale, Mark D. Bertness, and Andrew H. Altieri. Regional ontogeny of New England salt marsh die-off. Conservation Biology, 28(4):1041–1048, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carlson:2018:EED

[CBB18a] Colin J. Carlson, Alexander L. Bond, and Kevin R. Burgio. Estimating the extinction date of the thylacine with mixed certainty data. Conservation Biology, 32(2):477–483, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carlson:2018:RSM

[CBB18b] Colin J. Carlson, Alexander L. Bond, and Kevin R. Burgio. Reevaluating sighting models and moving beyond them to test and contextualize the extinction of the thylacine. Conservation Biology, 33(1):1198–1199, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coker:2018:ABO

[CBCW18] Myles E. Coker, Nick R. Bond, Yung En Chee, and Christopher J. Walsh. Alternatives to biodiversity offsets for mitigating the effects of urbanization on stream ecosystems. Conservation Biology, 32(4):789–797, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Canteiro:2019:ECK

[CBF^{+}19] Cátia Canteiro, Laísa Barcelos, Fabiana Filardi, Rafaela Forzza, Laura Green, João Lanna, Paula Leitman, William Milliken, Marli Pires Morim, Kristina Patmore, Sarah Phillips, Barnaby Walker, Marie-Hélène Weech, and Eimear Nic Lughadha. Enhancement of conservation knowledge through increased access to botanical information. Conservation Biology, 33(3):523–533, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cabral:2011:EHF

[CBM^{+}11] Juliano Sarmento Cabral, William J. Bond, Guy F. Midgley, Anthony G. Rebelo, Wilfried Thuiller, and Frank M. Schurr. Effects of harvesting flowers from shrubs on the persistence and abundance of wild shrub populations at multiple spatial extents. Efectos de la cosecha de flores de arbustos sobre la persistencia y abundancia de poblaciones de arbustos silvestres en escalas espaciales múltiples. Conservation Biology, 25(1): 73–84, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chollett:2016:QSS

[CBM16] Iliana Chollett, Stephen J. Box, and Peter J. Mumby. Quantifying the squeezing or stretching of fisheries as they adapt to displacement by marine reserves. Conservation Biology, 30(1): 166–175, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carter:2018:CCD

[CBM^{+}18] Neil H. Carter, Paola Bouley, Sean Moore, Michael Poulos, Jérémy Bouyer, and Stuart L. Pimm. Climate change, disease range shifts, and the future of the Africa lion. Conservation Biology, 33(1):1207–1210, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Canessa:2019:QBM

[CBPM19] Stefano Canessa, Claudio Bozzuto, Frank Pasmans, and An Martel. Quantifying the burden of managing wildlife diseases in multiple host species. Conservation Biology, 33(5): 1131–1140, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crossman:2011:CPL

[CBS11] Neville D. Crossman, Brett A. Bryan, and David M. Summers. Carbon payments and low-cost conservation. Pagos de carbono y conservación de bajo costo. Conservation Biology, 25(4):835– 845, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:2010:RID

[CBSZ10] Rulon W. Clark, William S. Brown, Randy Stechert, and Kelly R. Zamudio. Roads, interrupted dispersal, and genetic diversity in timber rattlesnakes. Carreteras, dispersión interrumpida, y diversidad genética en serpientes de cascabel. Conservation Biology, 24(4):1059–1069, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cochran-Biederman:2015:ICS

[CBWFL15] Jennifer L. Cochran-Biederman, Katherine E. Wyman, William E. French, and Grace L. Loppnow. Identifying correlates of success and failure of native freshwater fish reintroductions. Conservation Biology, 29(1):175–186, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Connors:2014:DDT

[CC14] B. M. Connors and A. B. Cooper. Determining decision thresholds and evaluating indicators when conservation status is measured as a continuum. Conservation Biology, 28(6):1626–1635, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clements:2017:PPN

[CC17] Hayley S. Clements and Graeme S. Cumming. Positives and pathologies of natural resource management on private landconservation areas. Conservation Biology, 31(3):707–717, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clements:2018:TTI

[CC18] Hayley S. Clements and Graeme S. Cumming. Traps and transformations influencing the financial viability of tourism on private-land conservation areas. Conservation Biology, 32 (2):424–436, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chaigneau:2019:IBN

[CCB^{+}19a] Tomas Chaigneau, Sarah Coulthard, Katrina Brown, Tim M. Daw, and Björn Schulte-Herbrüggen. Incorporating basic needs to reconcile poverty and ecosystem services. Conservation Biology, 33(3):655–664, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coristine:2019:NCG

[CCB^{+}19b] Laura E. Coristine, Sheila Colla, Nathan Bennett, Anja M. Carlsson, Christina Davy, Kimberley T. A. Davies, Brett Favaro, D. T. Tyler Flockhart, Kevin Fraser, Diane Orihel, Sarah P. Otto, Wendy Palen, Jean L. Polfus, Oscar Venter, and Adam T. Ford. National contributions to global ecosystem values. Conservation Biology, 33(5):1219–1223, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cundill:2012:SST

[CCBF12] G. Cundill, G. S. Cumming, D. Biggs, and C. Fabricius. Soft systems thinking and social learning for adaptive management. Pensamiento sistémico blando y aprendizaje social para manejo adaptativo. Conservation Biology, 26(2):13–20, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clements:2014:ERE

[CCBP14] Christopher F. Clements, Ben Collen, Tim M. Blackburn, and Owen L. Petchey. Effects of recent environmental change on accuracy of inferences of extinction status. Conservation Biology, 28(4):971–981, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Che-Castaldo:2018:PDR

[CCCCN18] Judy Che-Castaldo, Christian Che-Castaldo, and Maile C. Neel. Predictability of demographic rates based on phylogeny and biological similarity. Conservation Biology, 32(6):1290– 1300, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chang:2018:ECC

[CCG18] Charlotte H. Chang, Maarten J. L. F. Cruyff, and Xingli Giam. Examining conservation compliance with randomized response technique analyses. Conservation Biology, 32(6):1448–1456, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chades:2012:SRR

[CCM12] Iadine Chadès, Janelle M. R. Curtis, and Tara G. Martin. Setting realistic recovery targets for two interacting endangered species, sea otter and northern abalone. Conservation Biology, 26(6):1016–1025, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Che-Castaldo:2016:SLP

[CCN16] Judy P. Che-Castaldo and Maile C. Neel. Species-level persistence probabilities for recovery and conservation status assessment. Conservation Biology, 30(6):1297–1306, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crees:2016:CAA

[CCS^{+}16] Jennifer J. Crees, Amy C. Collins, P. J. Stephenson, Helen M. R. Meredith, Richard P. Young, Caroline Howe, Mark R. Stanley Price, and Samuel T. Turvey. A comparative approach to assess drivers of success in mammalian conservation recovery programs. Conservation Biology, 30(4):694–705, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cummings:2018:IDF

[CCS^{+}18] Jonathan W. Cummings, Sarah J. Converse, David R. Smith, Steve Morey, and Michael C. Runge. Implicit decision framing as an unrecognized source of confusion in endangered species classification. Conservation Biology, 32(6):1246–1254, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Canessa:2016:PES

[CCW^{+}16] Stefano Canessa, Sarah J. Converse, Matt West, Nick Clemann, Graeme Gillespie, Michael McFadden, Aimee J. Silla, Kirsten M. Parris, and Michael A. McCarthy. Planning for ex situ conservation in the face of uncertainty. Conservation Biology, 31(1):599–609, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cornelisse:2013:EKE

[CD13] Tara M. Cornelisse and Timothy P. Duane. Effects of knowledge of an endangered species on recreationists’ attitudes and stated behaviors and the significance of management compliance for ohlone tiger beetle conservation. Conservation Biology, 28(1):1449–1457, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Caro:2016:WWM

[CD16] Tim Caro and Tim R. B. Davenport. Wildlife and wildlife management in Tanzania. Conservation Biology, 30(4):716– 723, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Caro:2012:CA

[CDF^{+}12] Tim Caro, Jack Darwin, Tavis Forrester, Cynthia Ledoux-Bloom, and Caitlin Wells. Conservation in the Anthropocene. Conservation Biology, 26(2):185–188, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cornwall:2015:ENF

[CE15] Christopher E. Cornwall and Tyler D. Eddy. Effects of nearfuture ocean acidification, fishing, and marine protection on a temperate coastal ecosystem. Conservation Biology, 29(1): 207–215, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chu:2018:ETP

[CEdK18] Cindy Chu, Lucy Ellis, and Derrick T. de Kerckhove. Effectiveness of terrestrial protected areas for conservation of lake fish communities. Conservation Biology, 32(3):607–618, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crone:2013:AMM

[CEM^{+}13] Elizabeth E. Crone, Martha M. Ellis, William F. Morris, Amanda Stanley, Timothy Bell, Paulette Bierzychudek, Johan Ehrlén, Thomas N. Kaye, Tiffany M. Knight, Peter Lesica, Gerard Oostermeijer, Pedro F. Quintana-Ascencio, Tamara Ticktin, Teresa Valverde, Jennifer L. Williams, Daniel F. Doak, Rengaian Ganesan, Kathyrn Mceachern, Andrea S. Thorpe, and Eric S. Menges. Ability of matrix models to explain the past and predict the future of plant populations. Conservation Biology, 28(4):968–978, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Campbell:2018:FGR

[CESS^{+}18] Stuart J. Campbell, Graham J. Edgar, Rick D. Stuart-Smith, German Soler, and Amanda E. Bates. Fishing-gear restrictions and biomass gains for coral reef fishes in marine protected areas. Conservation Biology, 32(2):401–410, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Connell:2017:TTE

[CFB^{+}17] Sean D. Connell, Milena Fernandes, Owen W. Burnell, Zoë A. Doubleday, Kingsley J. Griffin, Andrew D. Irving, Jonathan Y. S. Leung, Samuel Owen, Bayden D. Russell, and Laura J. Falkenberg. Testing for thresholds of ecosystem collapse in seagrass meadows. Conservation Biology, 31(5):1196–1201, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coulon:2010:EHF

[CFBL10] AuréLie Coulon, John W. Fitzpatrick, Reed Bowman, and Irby J. Lovette. Effects of habitat fragmentation on effective dispersal of Florida scrub-jays. Efectos de la fragmentación sobre la dispersión efectiva de aphelocoma cœrulescens. Conservation Biology, 24(4):1080–1088, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2014:DMC

[CFL14] Carlos Carroll, Richard J. Fredrickson, and Robert C. Lacy. Developing metapopulation connectivity criteria from genetic and habitat data to recover the endangered Mexican wolf. Conservation Biology, 28(1):76–86, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Casner:2014:CUE

[CFO^{+}14] Kayce L. Casner, Matthew L. Forister, Joshua M. O’Brien, James Thorne, David Waetjen, and Arthur M. Shapiro. Contribution of urban expansion and a changing climate to decline of a butterfly fauna. Conservation Biology, 28(3):773–782, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Campbell:2013:LIE

[CG13] Lisa M. Campbell and Matthew H. Godfrey. Of legacies and icons: Evolution of sea turtle science and conservation. Conservation Biology, 27(4):890–893, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Claudet:2011:CEL

[CGCL11] Joachim Claudet, José Antonio García-Charton, and Philippe Lenfant. Combined effects of levels of protection and environmental variables at different spatial resolutions on fish assemblages in a Marine Protected Area. Efectos combinados de niveles de protección y variables ambientales en resoluciones espaciales diferentes sobre ensambles de peces en un Área marina protegida. Conservation Biology, 25(1):105–114, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cosset:2019:ITF

[CGE19] Cindy C. P. Cosset, James J. Gilroy, and David P. Edwards. Impacts of tropical forest disturbance on species vital rates. Conservation Biology, 33(2):66–75, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cinner:2013:GEL

[CGHM13] Joshua E. Cinner, Nicholas A. J. Graham, Cindy Huchery, and M. Aaron Macneil. Global effects of local human population density and distance to markets on the condition of coral reef fisheries. Conservation Biology, 27(3):453–458, June 2013. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ceausu:2019:GTO

[CGK^{+}19] Silvia Ceauşu, Rose A. Graves, Alexander K. Killion, Jens-Christian Svenning, and Neil H. Carter. Governing trade-offs in ecosystem services and disservices to achieve human–wildlife coexistence. Conservation Biology, 33(3):543–553, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Czucz:2010:IPD

[CGM10] Bálint Czúcz, Joseph P. Gathman, and Guy R. Mcpherson. The impending peak and decline of petroleum production: an underestimated challenge for conservation of ecological integrity.. El inminente pico y Declinación de la Producción de Petróleo: un reto subestimado para la conservación de la integridad ecológica. Conservation Biology, 24(4):948–956, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chisholm:2016:RNM

[CGS^{+}16] Ryan A. Chisholm, Xingli Giam, Keren R. Sadanandan, Tak Fung, and Frank E. Rheindt. A robust nonparametric method for quantifying undetected extinctions. Conservation Biology, 31(1):610–617, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chambers:2010:CG

[Cha10] Steven M. Chambers. Conservation genetics. Conservation Biology, 25(6):911–912, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Charpentier:2015:ILH

[Cha15] Anne Charpentier. Insights from life history theory for an explicit treatment of trade-offs in conservation biology. Conservation Biology, 29(3):738–747, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2017:DSI

[CHG^{+}17] Carlos Carroll, Brett Hartl, Gretchen T. Goldman, Daniel J. Rohlf, Adrian Treves, Jeremy T. Kerr, Euan G. Ritchie, Richard T. Kingsford, Katherine E. Gibbs, Martine Maron, and James E. M. Watson. Defending the scientific integrity of conservation-policy processes. Conservation Biology, 31(5): 967–975, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clarke:2013:PTP

[CHHR13] Shelley C. Clarke, Shelton J. Harley, Simon D. Hoyle, and Joel S. Rice. Population trends in Pacific Oceanic sharks and the utility of regulations on shark finning. Conservation Biology, 27(1):197–209, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chiba:2010:INN

[Chi10] Satoshi Chiba. Invasive non-native species’ provision of refugia for endangered native species refugios para especies nativas en peligro proporcionados por especies introducidas invasoras. Conservation Biology, 24(4):1141–1147, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ceausu:2015:MOC

[CHN^{+}15] Silvia Ceauşu, Max Hofmann, Laetitia M. Navarro, Steve Carver, Peter H. Verburg, and Henrique M. Pereira. Mapping opportunities and challenges for rewilding in Europe. Conservation Biology, 29(4):1017–1027, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cope:2018:RSC

[CHWH18] Holly R. Cope, Carolyn J. Hogg, Peter J. White, and Catherine A. Herbert. A role for selective contraception of individuals in conservation. Conservation Biology, 32(3):546–558, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cinner:2013:I

[Cin13] Joshua Cinner. Introduction. Conservation Biology, 27(3):441– 442, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2010:HBS

[CJDZ10] Carlos Carroll, Devin S. Johnson, Jeffrey R. Dunk, and William J. Zielinski. Hierarchical Bayesian spatial models for multispecies conservation planning and monitoring. Modelos espaciales Bayesianos jerárquicos para la planificación y monitoreo de la conservación de múltiples especies. Conservation Biology, 24(6):1538–1548, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cornu:2014:CPF

[CKK^{+}14] Elodie Le Cornu, John N. Kittinger, J. Zachary Koehn, Elena M. Finkbeiner, and Larry B. Crowder. Current practice and future prospects for social data in coastal and ocean planning. Conservation Biology, 28(4):902–911, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chittaro:2010:TOB

[CKKL10] P. M. Chittaro, I. C. Kaplan, A. Keller, and P. S. Levin. Tradeoffs between species conservation and the size of Marine Protected Areas. Balance entre la conservación de especies y el tamaño de Áreas Marinas Protegidas. Conservation Biology, 26(2):197–206, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cowling:2010:IOI

[CKPS10] Richard M. Cowling, Andrew T. Knight, Sean D. J. Privett, and Gyan Sharma. Invest in opportunity, not inventory of hotspots. Conservation Biology, 24(2):633–635, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chandler:2013:SSL

[CKR^{+}13] Richard B. Chandler, David I. King, Raul Raudales, Richard Trubey, Carlin Chandler, and Víctor Julio Arce Chávez. A small-scale land-sparing approach to conserving biological diversity in tropical agricultural landscapes. Conservation Biology, 27(4):785–795, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chase:2016:FED

[CL16] Sarah K. Chase and Arielle Levine. A framework for evaluating and designing citizen science programs for natural resources monitoring. Conservation Biology, 31(1):456–466, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:2011:CCB

[Cla11] J. Alan Clark. Climate change: for the birds. Conservation Biology, 25(6):1256–1258, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clapperton:2013:BBS

[Cla13a] Jonathan Clapperton. Beyond borders: The shady past and uncertain future of (trans) national parks. Conservation Biology, 28(1):1491–1493, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:2013:CPC

[Cla13b] Douglas A. Clark. Conservation policy classics, reissued. Conservation Biology, 27(1):235–237, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clavero:2014:SBC

[Cla14] Miguel Clavero. Shifting baselines and the conservation of nonnative species. Conservation Biology, 28(6):1434–1436, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clavero:2015:NNS

[Cla15] Miguel Clavero. Non-native species as conservation priorities: response to Díez–León et al. Conservation Biology, 29(3):957– 959, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chen:2015:CSB

[CLM^{+}15] Gao Chen, Shihong Luo, Nianshu Mei, Dingfang Shen, and Weibang Sun. Case study of building of conservation coalitions to conserve ecological interactions. Conservation Biology, 29 (6):1527–1536, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Catenazzi:2011:BDC

[CLRV11] Alessandro Catenazzi, Edgar Lehr, Lily O. Rodriguez, and Vance T. Vredenburg. Batrachochytrium dendrobatidis and the collapse of anuran species richness and abundance in the Upper Manu National Park, Southeastern Peru. Batrachochytrium dendrobatidis y el colapso de la riqueza de especies y abundancia de anuros en el Parque Nacional del Manu, Sureste de Perú. Conservation Biology, 26(3):382–391, April 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Capinha:2013:ECC

[CLT^{+}13] César Capinha, Eric R. Larson, Elena Tricarico, Julian D. Olden, and Francesca Gherardi. Effects of climate change, invasive species, and disease on the distribution of native European crayfishes. Conservation Biology, 27(4):731–740, August 2013.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chen:2010:UCE

[CLV^{+}10] Xiaodong Chen, Frank Lupi, Andrés Viña, Guangming He, and Jianguo Liu. Using cost-effective targeting to enhance the efficiency of conservation investments in payments for ecosystem services. Utilización de selección rentable para incrementar la eficiencia de las inversiones de conservación el pago por servicios ecosistémicos. Conservation Biology, 24(6):1469–1478, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Catenazzi:2014:TPD

[CLV14] Alessandro Catenazzi, Edgar Lehr, and Vance T. Vredenburg. Thermal physiology, disease, and amphibian declines on the Eastern Slopes of the Andes. Conservation Biology, 28(2):509– 517, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Colloff:2017:TCS

[CLvK^{+}17] Matthew J. Colloff, Sandra Lavorel, Lorrae E. van Kerkhoff, Carina A. Wyborn, Ioan Fazey, Russell Gorddard, Georgina M. Mace, Wendy B. Foden, Michael Dunlop, I. Colin Prentice, John Crowley, Paul Leadley, and Patrick Degeorges. Transforming conservation science and practice for a postnormal world. Conservation Biology, 31(5):1008–1017, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cullingham:2013:TAG

[CM13] Catherine I. Cullingham and Axel Moehrenschlager. Temporal analysis of genetic structure to assess population dynamics of reintroduced Swift foxes. Conservation Biology, 28(1): 1389–1398, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Colla:2017:QPP

[CM17] Sheila R. Colla and J. Scott MacIvor. Questioning public perception, conservation policy, and recovery actions for honeybees in North America. Conservation Biology, 31(5):1202–1204, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2012:ULM

[CMB12] Carlos Carroll, Brad H. Mcrae, and Allen Brookes. Use of linkage mapping and centrality analysis across habitat gradients to conserve connectivity of gray wolf populations in Western North America. Utilización del mapeo de Vínculos y el análisis de centralidad en un gradiente de hábitats para conservar la conectividad de poblaciones de lobo gris en el Occidente de Norte América. Conservation Biology, 26(2):78–87, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cross:2013:AAN

[CMG^{+}13] Molly S. Cross, Patrick D. McCarthy, Gregg Garfin, David Gori, and Carolyn A. F. Enquist. Accelerating adaptation of natural resource management to address climate change. Conservation Biology, 27(1):4–13, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clements:2015:IPE

[CMG15] Tom Clements and E. J. Milner-Gulland. Impact of payments for environmental services and protected areas on local livelihoods and forest conservation in northern Cambodia. Conservation Biology, 29(1):78–87, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cobb:2011:EPS

[CMJ^{+}11] T. P. Cobb, J. L. Morissette, J. M. Jacobs, M. J. Koivula, J. R. Spence, and D. W. Langor. Effects of postfire salvage logging on deadwood-associated beetles. Efectos del rescate de madera después de incendios sobre escarabajos asociados a madera muerta. Conservation Biology, 25(1):94–104, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cameron:2019:GMA

[CML^{+}19] Erin K. Cameron, Inês S. Martins, Patrick Lavelle, Jérôme Mathieu, Leho Tedersoo, Mohammad Bahram, Felix Gottschall, Carlos A. Guerra, Jes Hines, Guillaume Patoine, Julia Siebert, Marten Winter, Simone Cesarz, Olga Ferlian, Holger Kreft, Thomas E. Lovejoy, Luca Montanarella, Alberto Orgiazzi, Henrique M. Pereira, Helen R. P. Phillips, Josef Settele, Diana H. Wall, and Nico Eisenhauer. Global mismatches in aboveground and belowground biodiversity. Conservation Biology, 33(5): 1187–1192, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cosgrove:2017:UIC

[CMM17] Anita J. Cosgrove, Todd J. McWhorter, and Martine Maron. Using individual-condition measures to predict the long-term importance of habitat extent for population persistence. Conservation Biology, 31(5):1141–1151, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Camaclang:2015:CPI

[CMMP15] Abbey E. Camaclang, Martine Maron, Tara G. Martin, and Hugh P. Possingham. Current practices in the identification of critical habitat for threatened species. Conservation Biology, 29(2):482–492, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cushman:2010:UAO

[CMNM10] Samuel A. Cushman, Kevin S. Mckelvey, Barry R. Noon, and Kevin Mcgarigal. Use of abundance of one species as a surrogate for abundance of others uso de la abundancia de una especie como sustituta de la abundancia de otras. Conservation Biology, 25(6):830–840, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Castro:2010:ETP

[CMRH10] Jorge Castro, Gregorio Moreno-Rueda, and José A. Hódar. Experimental test of postfire management in pine forests: Impact of salvage logging versus partial cutting and nonintervention on bird–species assemblages. Prueba experimental del manejo post-fuego en bosques de pino: Impacto de la tala de salvamento versus el corte parcial y la no intervención sobre ensambles de especies de aves. Conservation Biology, 25(6):810–819, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cook:2013:ACS

[CMS^{+}13] Carly N. Cook, Michael B. Mascia, Mark W. Schwartz, Hugh P. Possingham, and Richard A. Fuller. Achieving conservation science that bridges the knowledge–action boundary. Conservation Biology, 27(4):669–678, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chanchani:2016:CTW

[CNBW16] Pranav Chanchani, Barry R. Noon, Larissa L. Bailey, and Rekha A. Warrier. Conserving tigers in working landscapes. Conservation Biology, 31(1):649–660, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cachelin:2010:LFT

[CND10] Adrienne Cachelin, Russell Norvell, and Ann Darling. Language fouls in teaching ecology: Why traditional metaphors undermine conservation literacy. Errores de lenguaje al enseñar ecología: Porqué las metáforas tradicionales socavan la cultura de conservación. Conservation Biology, 24(3):669–674, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cooke:2017:TIF

[CNK^{+}17] Steven J. Cooke, Vivian M. Nguyen, Steven T. Kessel, Nigel E. Hussey, Nathan Young, and Adam T. Ford. Troubling issues at the frontier of animal tracking for conservation and management. Conservation Biology, 31(5):1205–1207, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chades:2015:BIC

[CNvL^{+}15] Iadine Chadés, Sam Nicol, Stephen van Leeuwen, Belinda Walters, Jennifer Firn, Andrew Reeson, Tara G. Martin, and Josie Carwardine. Benefits of integrating complementarity into priority threat management. Conservation Biology, 29(2):525–536, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cole:2010:VRE

[Col10] Kenneth L. Cole. Vegetation response to early Holocene warming as an analog for current and future changes. Respuesta de la vegetación al calentamiento en el Holoceno temprano como un análogo para cambios actuales y futuros. Conservation Biology, 24(1):29–37, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Conn:2011:CBT

[Con11] Charles Conn. Conservation biology through the lens of a career in salmon conservation. Conservation Biology, 25(6):1075– 1079, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coppolillo:2012:LFF

[Cop12] Pete Coppolillo. A look at the future of food production. Conservation Biology, 26(2):373–374, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cork:2016:CBS

[Cor16] Susan Catherine Cork. Communicating biological science to the public. Conservation Biology, 30(6):681–683, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cork:2019:HNP

[Cor19] Susan Catherine Cork. History, national parks, and conservation science in North America. Conservation Biology, 33(2): 222–224, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cox:2010:CWH

[Cox10] Linda R. Cox. Conservation where half the world lives. Conservation Biology, 24(2):637–638, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cruickshank:2016:QPD

[COZS16] Sam S. Cruickshank, Arpat Ozgul, Silvia Zumbach, and Benedikt R. Schmidt. Quantifying population declines based on presence-only records for red-list assessments. Conservation Biology, 30(5):1112–1121, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Curtin:2014:FRT

[CP14] Charles G. Curtin and Jessica P. Parker. Foundations of resilience thinking. Conservation Biology, 28(4):912–923, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coster:2014:CWA

[CPB14] Stephanie S. Coster, Jessica S. Veysey Powell, and Kimberly J. Babbitt. Characterizing the width of amphibian movements during postbreeding migration. Conservation Biology, 28(3):

756–762, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cubaynes:2010:IAD

[CPC^{+}10] Sarah Cubaynes, Roger Pradel, Rémi Choquet, Christophe Duchamp, Jean-Michel Gaillard, Jean-Dominique Lebreton, Eric Marboutin, Christian Miquel, Anne-Marie Reboulet, Carole Poillot, Pierre Taberlet, and Olivier Gimenez. Importance of accounting for detection heterogeneity when estimating abundance: the case of French wolves. Importancia de considerar la heterogeneidad de detección al estimar la abundancia: el caso de lobos franceses. Conservation Biology, 24(2):621– 626, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Critchlow:2015:STI

[CPD^{+}15] R. Critchlow, A. J. Plumptre, M. Driciru, A. Rwetsiba, E. J. Stokes, C. Tumwesigye, F. Wanyama, and C. M. Beale. Spatiotemporal trends of illegal activities from ranger-collected data in a Ugandan national park. Conservation Biology, 29 (5):1458–1470, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cook:2013:CSR

[CPF13] Carly N. Cook, Hugh P. Possingham, and Richard A. Fuller. Contribution of systematic reviews to management decisions. Conservation Biology, 28(4):902–915, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Comer:2015:IGC

[CPH^{+}15] Patrick J. Comer, Robert L. Pressey, Malcolm L. Hunter, Jr., Carrie A. Schloss, Steven C. Buttrick, Nicole E. Heller, John M. Tirpak, Daniel P. Faith, Molly S. Cross, and Mark L. Shaffer. Incorporating geodiversity into conservation decisions. Conservation Biology, 29(3):692–701, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Condie:2018:GBR

[CPM^{+}18] Scott A. Condie, Éva E. Plagányi, Elisabetta B. Morello, Karlo Hock, and Roger Beeden. Great Barrier Reef recovery through multiple interventions. Conservation Biology, 32(6):1356–1367, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carrasco:2016:FCR

[CPR^{+}16] L. R. Carrasco, S. K. Papworth, J. Reed, W. S. Symes, A. Ickowitz, T. Clements, K. S.-H. Peh, and T. Sunderland. Five challenges to reconcile agricultural land use and forest ecosystem services in Southeast Asia. Conservation Biology, 30(5): 962–971, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chen:2019:ELE

[CQC^{+}19] Cheng Chen, Rui-Chang Quan, Guanghong Cao, Hongpei Yang, A. Cole Burton, Michael Meitner, and Jedediah F. Brodie. Effects of law enforcement and community outreach on mammal diversity in a biodiversity hotspot. Conservation Biology, 33(3):612–622, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coeurdassier:2014:UWP

[CRD^{+}14] Michael Coeurdassier, Romain Riols, Anouk Decors, Aymeric Mionnet, Fabienne David, Thomas Quintaine, Denis Truchetet, Renaud Scheifler, and Patrick Giraudoux. Unintentional wildlife poisoning and proposals for sustainable management of rodents. Conservation Biology, 28(2):315–321, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chamberlain:2012:HPC

[CRG12] Emily C. Chamberlain, Murray B. Rutherford, and Michael L. Gibeau. Human perspectives and conservation of grizzly bears in Banff National Park, Canada. Conservation Biology, 26 (3):420–431, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cilingir:2017:CGE

[ÇRG^{+}17] F. Gözde Çilingir, Frank E. Rheindt, Kritika M. Garg, Kalyar Platt, Steven G. Platt, and David P. Bickford. Conservation genomics of the endangered Burmese roofed turtle. Conservation Biology, 31(6):1469–1476, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carrillo-Rubio:2014:UMO

[CRKM^{+}14] Eduardo Carrillo-Rubio, Marc Kéry, Stephen J. Morreale, Patrick J. Sullivan, Beth Gardner, Evan G. Cooch, and James P. Lassoie. Use of multispecies occupancy models to evaluate the response of bird communities to forest degradation associated with logging. Conservation Biology, 28(4):1034– 1044, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Catalano:2018:BSC

[CRMK18] Allison S. Catalano, Kent Redford, Richard Margoluis, and Andrew T. Knight. Black swans, cognition, and the power of learning from failure. Conservation Biology, 32(3):584–596, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2012:SIR

[CRNR12] Carlos Carroll, Daniel J. Rohlf, Barry R. Noon, and J. Michael Reed. Scientific integrity in recovery planning and risk assessment: Comment on Wilhere. Conservation Biology, 27(1):743– 745, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Wil12a].

Connette:2013:LHP

[CS13] Grant M. Connette and Raymond D. Semlitsch. Life history as a predictor of Salamander recovery rate from timber harvest in Southern Appalachian forests, U.S.A. Conservation Biology, 28(1):1399–1409, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:2014:TPI

[CS14] Susan G. Clark and Kelly J. Stoner. Toward pragmatic, integrated conservation. Conservation Biology, 28(5):886–887, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carver:2017:HEC

[CS17a] L. Carver and S. Sullivan. How economic contexts shape calculations of yield in biodiversity offsetting. Conservation Biology, 31(5):1053–1065, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cook:2017:ASP

[CS17b] Carly N. Cook and Carla M. Sgrò. Aligning science and policy to achieve evolutionarily enlightened conservation. Conservation Biology, 31(3):501–512, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cook:2019:CPU

[CS19] Carly N. Cook and Carla M. Sgrò. Conservation practitioners’ understanding of how to manage evolutionary processes. Conservation Biology, 33(5):993–1001, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clements:2010:CON

[CSB^{+}10] Gopalasamy Reuben Clements, Jeffrey Sayer, Agni K. Boedhihartono, Oscar Venter, Thomas Lovejoy, Lian Pin Koh, and William F. Laurance. Cautious optimism over Norway– Indonesia REDD Pact. Conservation Biology, 24(6):1437–1438, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coad:2013:SEC

[CSMG^{+}13] L. Coad, J. Schleicher, E. J. Milner-Gulland, T. R. Marthews, M. Starkey, A. Manica, A. Balmford, W. Mbombe, T. R. Diop Bineni, and K. A. Abernethy. Social and ecological change over a decade in a village hunting system, Central Gabon. Conservation Biology, 27(2):270–280, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Csoka:2017:WIF

[Csó17] György Csóka. Warming up insects and forest fragmentation. Conservation Biology, 31(3):487–489, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Child:2019:FMW

[CSR^{+}19] Matthew F. Child, S. A. Jeanetta Selier, Frans G. T. Radloff, W. Andrew Taylor, Michael Hoffmann, Lizanne Nel, R. John Power, Coral Birss, Nicola C. Okes, Michael J. Peel, David Mallon, and Harriet Davies-Mostert. A framework to measure the wildness of managed large vertebrate populations. Conservation Biology, 33(5):1106–1119, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chaparro:2011:DEH

[CT11] Demetria Mondragón Chaparro and Tamara Ticktin. Demographic effects of harvesting epiphytic bromeliads and an alternative approach to collection. Efectos demográficos de la cosecha de bromelias epífitas y un método alternativo de colecta. Conservation Biology, 25(4):797–807, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cavada:2019:EAD

[CTBR19] Nathalie Cavada, Simone Tenan, Claudia Barelli, and Francesco Rovero. Effects of anthropogenic disturbance on primate density at the landscape scale. Conservation Biology, 33 (4):873–882, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cao:2014:SSH

[CTM^{+}14] Nguyen Van Cao, Nguyen Thien Tao, Amelia Moore, Alfred Montoya, Arne Redsted Rasmussen, Kenneth Broad, Harold K. Voris, and Zoltan Takacs. Sea snake harvest in the Gulf of Thailand. Conservation Biology, 28(6):1677–1687, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Correia:2017:RRS

[CTRE^{+}17] Marta Correia, Sérgio Timóteo, Susana Rodríguez-Echeverría, Alban Mazars-Simon, and Ruben Heleno. Refaunation and the reinstatement of the seed-dispersal function in Gorongosa National Park. Conservation Biology, 31(1):76–85, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cosentino:2016:ECR

[CV16] A. Mel Cosentino and Diogo Veríssimo. Ending the citation of retracted papers. Conservation Biology, 31(1):676–678, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Costello:2015:CBT

[CVA15] Mark J. Costello, Bart Vanhoorne, and Ward Appeltans. Conservation of biodiversity through taxonomy, data publication, and collaborative infrastructures. Conservation Biology, 29(4): 1094–1099, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cook:2017:QEP

[CVMM17] Carly N. Cook, Rebecca S. Valkan, Michael B. Mascia, and Melodie A. McGeoch. Quantifying the extent of protectedarea downgrading, downsizing, and degazettement in Australia. Conservation Biology, 31(5):1039–1052, October 2017. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2010:GRU

[CVN^{+}10] Carlos Carroll, John A. Vucetich, Michael P. Nelson, Daniel J. Rohlf, and Michael K. Phillips. Geography and recovery under the U.S. Endangered Species Act. Geografía y recuperación bajo el Acta de Especies en Peligro de E.U.A. Conservation Biology, 24(2):395–403, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cotton:2010:MMB

[CW10] Samuel Cotton and Claus Wedekind. Male mutation bias and possible long-term effects of human activities. Sesgo en la mutación de machos y posibles efectos a largo plazo de actividades humanas. Conservation Biology, 24(5):1190–1197, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cook:2014:USF

[CWAW14] Carly N. Cook, Bonnie C. Wintle, Stephen C. Aldrich, and Brendan A. Wintle. Using strategic foresight to assess conservation opportunity. Conservation Biology, 28(6):1474–1483, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carwardine:2010:CPW

[CWH^{+}10] Josie Carwardine, Kerrie A. Wilson, Stefan A. Hajkowicz, Robert J. Smith, Carissa J. Klein, Matt Watts, and Hugh P. Possingham. Conservation planning when costs are uncertain. Planificación de la conservación cuando los costos son inciertos. Conservation Biology, 24(6):1529–1537, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Correll:2017:PSA

[CWH^{+}17] Maureen D. Correll, Whitney A. Wiest, Thomas P. Hodgman, W. Gregory Shriver, Chris S. Elphick, Brian J. McGill, Kathleen M. O’Brien, and Brian J. Olsen. Predictors of specialist avifaunal decline in coastal marshes. Conservation Biology, 31 (1):172–182, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cain:2014:SBI

[CWS^{+}14] Bradley Cain, Antony B. Wandera, Susan G. Shawcross, W. Edwin Harris, Barry Stevens-Wood, Stephen J. Kemp, Benson Okita-Ouma, and Phillip C. Watts. Sex-biased inbreeding effects on reproductive success and home range size of the critically endangered black rhinoceros. Conservation Biology, 28 (2):594–603, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chisholm:2016:NLT

[CWS16] Ryan A. Chisholm, Lahiru S. Wijedasa, and Tom Swinfield. The need for long-term remedies for Indonesia’s forest fires. Conservation Biology, 30(1):5–6, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cetas:2017:SRM

[CY17] Elijah R. Cetas and Maï Yasué. A systematic review of motivational values and conservation success in and around protected areas. Conservation Biology, 31(1):203–212, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cao:2019:DNV

[CZS19] Shixiong Cao, Junze Zhang, and Wei Su. Difference in the net value of ecological services between natural and artificial forests in China. Conservation Biology, 33(5):1076–1083, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dickson:2019:CTA

[DAA^{+}19] Brett G. Dickson, Christine M. Albano, Ranjan Anantharaman, Paul Beier, Joe Fargione, Tabitha A. Graves, Miranda E. Gray, Kimberly R. Hall, Josh J. Lawler, Paul B. Leonard, Caitlin E. Littlefield, Meredith L. McClure, John Novembre, Carrie A. Schloss, Nathan H. Schumaker, Viral B. Shah, and David M. Theobald. Circuit-theory applications to connectivity science and conservation. Conservation Biology, 33(2):239– 249, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dhanjal-Adams:2017:SCP

[DAKN^{+}17] Kiran L. Dhanjal-Adams, Marcel Klaassen, Sam Nicol, Hugh P. Possingham, Iadine Chadès, and Richard A. Fuller. Setting conservation priorities for migratory networks under uncertainty. Conservation Biology, 31(3):646–656, June 2017. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Darling:2015:UDB

[Dar15] E. S. Darling. Use of double-blind peer review to increase author diversity. Conservation Biology, 29(1):297–299, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davis:2011:TPP

[Dav11] Kate Davis. Toward profits without piracy. Conservation Biology, 25(6):642–644, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Driscoll:2016:RFF

[DBB^{+}16] Don A. Driscoll, Michael Bode, Ross A. Bradstock, David A. Keith, Trent D. Penman, and Owen F. Price. Resolving future fire management conflicts using multicriteria decision making. Conservation Biology, 30(1):196–205, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Durrant:2014:IMI

[DBK^{+}14] Halley M. S. Durrant, Christopher P. Burridge, Brendan P. Kelaher, Neville S. Barrett, Graham J. Edgar, and Melinda A. Coleman. Implications of macroalgal isolation by distance for networks of Marine Protected Areas. Conservation Biology, 28 (2):438–445, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Doukakis:2010:MRO

[DBP^{+}10] Phaedra Doukakis, Elizabeth A. Babcock, Ellen K. Pikitch, Alexei R. Sharov, Mirgaly Baimukhanov, Sagiden Erbulekov, Yelena Bokova, and Akhat Nimatov. Management and recovery options for Ural River beluga sturgeon. Opciones de manejo y recuperación del esturión beluga del Río Ural. Conservation Biology, 24(3):769–777, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davies:2013:UMD

[DBR13] A. L. Davies, R. Bryce, and S. M. Redpath. Use of multicriteria decision analysis to address conservation conflicts. Conservation Biology, 28(4):936–944, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Doak:2013:UPV

[DBV13] Daniel F. Doak, Victoria J. Bakker, and Winston Vickers. Using population viability criteria to assess strategies to minimize disease threats for an endangered carnivore. Conservation Biology, 27(2):303–314, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dinca:2018:UGC

[DBV^{+}18] Vlad Dincă, Zsolt Bálint, Raluca Vodă, Leonardo Dapporto, Paul D. N. Hebert, and Roger Vila. Use of genetic, climatic, and microbiological data to inform reintroduction of a regionally extinct butterfly. Conservation Biology, 32(4):828–837, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Danylchuk:2011:ERA

[DC11] Andy J. Danylchuk and Steven J. Cooke. Engaging the recreational angling community to implement and manage aquatic protected areas. Involucrando a la comunidad de pescadores recreativos en la implementación y manejo de Áreas acuáticas protegidas. Conservation Biology, 25(3):458–464, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dougherty:2016:PPC

[DCB^{+}16] Eric R. Dougherty, Colin J. Carlson, Veronica M. Bueno, Kevin R. Burgio, Carrie A. Cizauskas, Christopher F. Clements, Dana P. Seidel, and Nyeema C. Harris. Paradigms for parasite conservation. Conservation Biology, 30(4):724– 733, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dwyer:2019:UIB

[DCP^{+}19] Ross G. Dwyer, Hamish A. Campbell, Richard D. Pillans, Matthew E. Watts, Barry J. Lyon, Siddeswara M. Guru, Minh N. Dinh, Hugh P. Possingham, and Craig E. Franklin. Using individual-based movement information to identify spatial conservation priorities for mobile species. Conservation Biology, 33(6):1426–1437, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dans:2012:ETB

[DDPC12] Silvana Laura Dans, Mariana Degrati, Susana Noemí Pedraza, and Enrique Alberto Crespo. Effects of tour boats on dolphin activity examined with sensitivity analysis of Markov chains. Efectos de barcos turísticos sobre la actividad de delfines examinada con análisis de sensibilidad de cadenas de Markov. Conservation Biology, 27(1):708–716, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DeCandia:2018:TIM

[DDv18] Alexandra L. DeCandia, Andrew P. Dobson, and Bridgett M. vonHoldt. Toward an integrative molecular approach to wildlife disease. Conservation Biology, 32(4):798–807, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dutton:2010:EVN

[DEJM10] Adam Dutton, Gareth Edwards-Jones, and David W. Macdonald. Estimating the value of non-use benefits from small changes in the provision of ecosystem services. Estimación del valor de los beneficios de cambios pequeños en el suministro de servicios del ecosistema. Conservation Biology, 24(6):1479– 1487, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Derocher:2013:HC

[Der13] Andrew E. Derocher. Heating up conservation. Conservation Biology, 27(3):638–639, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Desrousseaux:2019:AIL

[Des19] Maylis Desrousseaux. Assessment of international law on biodiversity. Conservation Biology, 33(5):1226–1227, October 2019.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dellasala:2012:PAS

[DFJ^{+}12] Dominick A. Dellasala, John M. Fitzgerald, Bengt-Gunnar Jonsson, Jeffrey A. Mcneely, Benjamin Delali Dovie, Martin Dieterich, Patricia Majluf, Simon C. Nemtzov, Owen T. Nevin, E. Christien M. Parsons, and James E. M. Watson. Priority actions for sustainable forest management in the international year of forests. Conservation Biology, 26(3):572–575, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dubois:2017:ICP

[DFR^{+}17] Sara Dubois, Nicole Fenwick, Erin A. Ryan, Liv Baker, Sandra E. Baker, Ngaio J. Beausoleil, Scott Carter, Barbara Cartwright, Federico Costa, Chris Draper, John Griffin, Adam Grogan, Gregg Howald, Bidda Jones, Kate E. Littin, Amanda T. Lombard, David J. Mellor, Daniel Ramp, Catherine A. Schuppli, and David Fraser. International consensus principles for ethical wildlife control. Conservation Biology, 31 (4):753–760, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dalleau:2010:UHS

[DFW^{+}10] Mayeul Dalleau, Serge André Fouët, Colette C. C. Wabnitz, Claude Payri, Laurent Wantiez, Michel Pichon, Kim Friedman, Laurent Vigliola, and Francesca Benzoni. Use of habitats as surrogates of biodiversity for efficient coral reef conservation planning in Pacific Ocean islands. Uso de hábitats como sustitutos de la biodiversidad para la conservación eficiente de arrecifes de coral. Conservation Biology, 24(2):541–552, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dwyer:2014:PMA

[DHD14] J. F. Dwyer, R. E. Harness, and K. Donohue. Predictive model of avian electrocution risk on overhead power lines. Conservation Biology, 28(1):159–168, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dolrenry:2016:CMP

[DHF16] Stephanie Dolrenry, Leela Hazzah, and Laurence G. Frank. Conservation and monitoring of a persecuted African lion population by maasai warriors. Conservation Biology, 31(1):467– 475, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dilkina:2017:TOE

[DHG^{+}17] Bistra Dilkina, Rachel Houtman, Carla P. Gomes, Claire A. Montgomery, Kevin S. McKelvey, Katherine Kendall, Tabitha A. Graves, Richard Bernstein, and Michael K. Schwartz. Tradeoffs and efficiencies in optimal budget-constrained multispecies corridor networks. Conservation Biology, 31(1):192–202, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Derr:2012:PEC

[DHH^{+}12] James N. Derr, Philip W. Hedrick, Natalie D. Halbert, Louis Plough, Lauren K. Dobson, Julie King, Calvin Duncan, David L. Hunter, Noah D. Cohen, and Dennis Hedgecock. Phenotypic effects of cattle mitochondrial DNA in American bison. Conservation Biology, 26(6):1130–1136, December 2012. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dickman:2015:PGF

[DHM^{+}15] Amy J. Dickman, Amy E. Hinks, Ewan A. Macdonald, Dawn Burnham, and David W. Macdonald. Priorities for global felid conservation. Conservation Biology, 29(3):854–864, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dieterich:2010:CFB

[Die10a] Martin Dieterich. Conservation and farmland birds. Conservation Biology, 24(4):1165–1166, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dietz:2010:TDK

[Die10b] Lou Ann Dietz. A tribute to devra kleiman, a pioneer in conservation biology. Conservation Biology, 24(4):1161–1163, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Diesendorf:2016:SJN

[Die16] Mark Diesendorf. Subjective judgments in the nuclear energy debate. Conservation Biology, 31(1):666–669, June 2016. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dinerstein:2011:I

[Din11] Eric Dinerstein. Introduction. Conservation Biology, 25(6): 1112, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Duffy:2016:TNU

[DJBB16] Rosaleen Duffy, Freya A. V. St John, Bram Büscher, and Dan Brockington. Toward a new understanding of the links between poverty and illegal wildlife hunting. Conservation Biology, 30 (1):14–22, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dearborn:2010:MCU

[DK10] Donald C. Dearborn and Salit Kark. Motivations for conserving urban biodiversity motivaciones para conservar la biodiversidad urbana. Conservation Biology, 24(2):432–440, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davis:2015:AEC

[DKG^{+}15] Katrina Davis, Marit Kragt, Stefan Gelcich, Steven Schilizzi, and David Pannell. Accounting for enforcement costs in the spatial allocation of marine zones. Conservation Biology, 29 (1):226–237, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Do:2015:UIS

[DKL^{+}15] Yuno Do, Ji Yoon Kim, Maurice Lineman, Dong-Kyun Kim, and Gea-Jae Joo. Using internet search behavior to assess public awareness of protected wetlands. Conservation Biology, 29 (1):271–279, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Diez-Leon:2015:SPE

[DLMAG15] María Díez-León, Rafael Miranda, Arturo H. Ariño, and David Galicia. Setting priorities for existing conservation needs of crayfish and mink. Conservation Biology, 29(2):599–601, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dennis:2017:UCS

[DMB^{+}17] Emily B. Dennis, Byron J. T. Morgan, Tom M. Brereton, David B. Roy, and Richard Fox. Using citizen science butterfly counts to predict species population trends. Conservation Biology, 31(6):1350–1361, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Decaens:2018:BLA

[DMF^{+}18] Thibaud Decaëns, Marlúcia B. Martins, Alexander Feijoo, Johan Oszwald, Sylvain Dolédec, Jérôme Mathieu, Xavier Arnaud de Sartre, Diego Bonilla, George G. Brown, Yeimmy Andrea Cuellar Criollo, Florence Dubs, Ivaneide S. Furtado, Valérie Gond, Erika Gordillo, Solen Le Clec’h, Raphaël Marichal, Danielle Mitja, Izildinha Miranda de Souza, Catarina Praxedes, Rodolphe Rougerie, Darío H. Ruiz, Joel Tupac Otero, Catalina Sanabria, Alex Velasquez, Luz Elena M. Zararte, and Patrick Lavelle. Biodiversity loss along a gradient of deforestation in Amazonian agricultural landscapes. Conservation Biology, 32(6):1380–1391, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dobson:2019:DDP

[DMGB^{+}19] Andrew D. M. Dobson, E. J. Milner-Gulland, Colin M. Beale, Harriet Ibbett, and Aidan Keane. Detecting deterrence from patrol data. Conservation Biology, 33(3):665–675, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Deikumah:2014:BTB

[DMM14] Justus P. Deikumah, Clive A. Mcalpine, and Martine Maron. Biogeographical and taxonomic biases in tropical forest fragmentation research. Conservation Biology, 28(6):1522–1531, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davy:2017:CIP

[DMR^{+}17] Christina M. Davy, Gabriela F. Mastromonaco, Julia L. Riley, James H. Baxter-Gilbert, Heather Mayberry, and Craig K. R. Willis. Conservation implications of physiological carry-over effects in bats recovering from white-nose syndrome. Conservation Biology, 31(3):615–624, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dias:2010:ERI

[DMZ10] Murilo S. Dias, William E. Magnusson, and Jansen Zuanon. Effects of reduced-impact logging on fish assemblages in Central Amazonia. Efectos de la explotación maderera de impacto reducido sobre ensambles de peces en la Amazonía Central. Conservation Biology, 26(2):278–286, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dawson:2015:PIE

[DOC^{+}15] Jeffrey Dawson, Steffen Oppel, Richard J. Cuthbert, Nick Holmes, Jeremy P. Bird, Stuart H. M. Butchart, Dena R. Spatz, and Bernie Tershy. Prioritizing islands for the eradication of invasive vertebrates in the United Kingdom overseas territories. Conservation Biology, 29(1):143–153, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dolan:2010:SEC

[Dol10] Katie Dolan. The short evolution of climate-change literature. Conservation Biology, 26(2):356–357, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Donlan:2012:ASR

[Don12] C. Josh Donlan. The art and science of restoration. Conservation Biology, 27(1):751–752, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dorje:2011:WPE

[Dor11] H.H.17th Gyalwang Karmapa Ogyen Trinley Dorje. Walking the path of environmental Buddhism through compassion and emptiness. Conservation Biology, 25(6):1094–1097, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dufresnes:2015:EBH

[DP15] Christophe Dufresnes and Nicolas Perrin. Effect of biogeographic history on population vulnerability in European amphibians. Conservation Biology, 29(4):1235–1241, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Delhey:2017:CIA

[DP17] Kaspar Delhey and Anne Peters. Conservation implications of anthropogenic impacts on visual communication and camouflage. Conservation Biology, 31(1):30–39, February 2017. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dawson:2016:AGZ

[DPGY16] Jeff Dawson, Freisha Patel, Richard A. Griffiths, and Richard P. Young. Assessing the global zoo response to the amphibian crisis through 20-year trends in captive collections. Conservation Biology, 30(1):82–91, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dickman:2018:SLC

[DPJM18] Amy Dickman, Craig Packer, Paul J. Johnson, and David W. Macdonald. A sideways look at conservation and consistency in tourism policy. Conservation Biology, 32(3):744–746, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dotta:2016:ASR

[DPS^{+}16] G. Dotta, B. Phalan, T. W. Silva, R. Green, and A. Balmford. Assessing strategies to reconcile agriculture and bird conservation in the temperate grasslands of South America. Conservation Biology, 31(1):618–627, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Darimont:2018:PPL

[DPT^{+}18] Chris T. Darimont, Paul C. Paquet, Adrian Treves, Kyle A. Artelle, and Guillaume Chapron. Political populations of large carnivores. Conservation Biology, 32(3):747–749, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Donald:2015:SRG

[DRA^{+}15] Paul F. Donald, Philip D. Round, Thiri Dai We Aung, Mark Grindley, Rob Steinmetz, Nay Myo Shwe, and Graeme M. Buchanan. Social reform and a growing crisis for southern Myanmar’s unique forests. Conservation Biology, 29(5):1485– 1488, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Drew:2011:RNH

[Dre11] Joshua Drew. The role of natural history institutions and bioinformatics in conservation biology. Conservation Biology, 25(6): 1250–1252, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Drechsler:2017:GSO

[Dre17] Martin Drechsler. Generating spatially optimized habitat in a trade-off between social optimality and budget efficiency. Conservation Biology, 31(1):221–225, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dorji:2018:ICP

[DRF^{+}18] Sangay Dorji, Rajanathan Rajaratnam, Lorena Falconi, Stephen E. Williams, Priyakant Sinha, and Karl Vernes. Identifying conservation priorities for threatened Eastern Himalayan mammals. Conservation Biology, 32(5):1162–1173, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ducatez:2019:LHT

[DS19] Simon Ducatez and Rick Shine. Life-history traits and the fate of translocated populations. Conservation Biology, 33(4):853– 860, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dressel:2015:MAS

[DSE15] S. Dressel, C. Sandström, and G. Ericsson. A meta-analysis of studies on attitudes toward bears and wolves across Europe 1976–2012. Conservation Biology, 29(2):565–574, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Diedrich:2017:SCK

[DSG^{+}17] Amy Diedrich, Natalie Stoeckl, Georgina G. Gurney, Michelle Esparon, and Richard Pollnac. Social capital as a key determinant of perceived benefits of community-based marine protected areas. Conservation Biology, 31(2):311–321, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dickson:2019:MMS

[DSK^{+}19] Brett G. Dickson, Rick D. Scherer, Amanda M. Kissel, Bryan P. Wallace, Kathryn M. Langin, Miranda E. Gray, Amanda F. Scheib, and Bruce Weise. Multiyear monitoring of survival following mitigation-driven translocation of a long-lived threatened reptile. Conservation Biology, 33(5):1094–1105, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Drechsler:2016:CEC

[DSSW16] Martin Drechsler, Henrik G. Smith, Astrid Sturm, and Frank Wätzold. Cost-effectiveness of conservation payment schemes for species with different range sizes. Conservation Biology, 30 (4):894–899, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Duron:2017:CIR

[DSV17] Quiterie Duron, Aaron B. Shiels, and Eric Vidal. Control of invasive rats on islands and priorities for future action. Conservation Biology, 31(4):761–771, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dillon:2016:ISS

[DSW16] Justin Dillon, Robert B. Stevenson, and Arjen E. J. Walss. Introduction to the special section: Moving from citizen to civic science to address wicked conservation problems. Corrected by erratum 12844. Conservation Biology, 30(3):450–455, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See erratum [Ano17i].

Dallimer:2014:QPN

[DTH^{+}14] Martin Dallimer, Dugald Tinch, Nick Hanley, Katherine N. Irvine, James R. Rouquette, Philip H. Warren, Lorraine Maltby, Kevin J. Gaston, and Paul R. Armsworth. Quantifying preferences for the natural world using monetary and nonmonetary assessments of value. Conservation Biology, 28 (2):404–413, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dietsch:2016:SVB

[DTM16] Alia M. Dietsch, Tara L. Teel, and Michael J. Manfredo. Social values and biodiversity conservation in a dynamic world. Conservation Biology, 30(6):1212–1221, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Duffy:2011:ISC

[Duf11] Karl J. Duffy. Identifying sighting clusters of endangered taxa with historical records. Identificación de grupos de avistamientos de taxa en peligro con registros históricos. Conservation Biology, 26(3):392–399, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Duangchantrasiri:2016:DLD

[DUS^{+}16] Somphot Duangchantrasiri, Mayuree Umponjan, Saksit Simcharoen, Anak Pattanavibool, Soontorn Chaiwattana, Sompoch Maneerat, N. Samba Kumar, Devcharan Jathanna, Arjun Srivathsa, and K. Ullas Karanth. Dynamics of a low-density tiger population in Southeast Asia in the context of improved law enforcement. Conservation Biology, 31(1):639–648, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dinerstein:2013:ECE

[DVW^{+}13] Eric Dinerstein, Keshav Varma, Eric Wikramanayake, George Powell, Susan Lumpkin, Robin Naidoo, Mike Korchinsky, Christian Del Valle, Shubash Lohani, John Seidensticker, Dirk Joldersma, Thomas Lovejoy, and Andrey Kushlin. Enhancing conservation, ecosystem services, and local livelihoods through a wildlife premium mechanism. Conservation Biology, 27(1): 14–23, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Driscoll:2019:SDC

[DW19] Don A. Driscoll and Maggie J. Watson. Science denialism and compassionate conservation: response to Wallach et al. 2018. Conservation Biology, 33(4):777–780, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ding:2017:EED

[DWCC17] Qi Ding, Yali Wang, Xinjun Chen, and Yong Chen. Effects of economics and demographics on global fisheries sustainability. Conservation Biology, 31(4):799–808, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Donlan:2010:UEO

[DWCW10] C. Josh Donlan, Dana K. Wingfield, Larry B. Crowder, and Chris Wilcox. Using expert opinion surveys to rank threats to endangered species: A case study with sea turtles. Utilización de encuestas de opinión de expertos para clasificar amenazas a especies en peligro: un caso de estudio con tortugas marinas. Conservation Biology, 24(6):1586–1595, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Decesare:2011:RTR

[DWH^{+}11] Nicholas J. Decesare, Jesse Whittington, Mark Hebblewhite, Hugh Robinson, Mark Bradley, Lalenia Neufeld, and Marco Musiani. The role of translocation in recovery of woodland caribou populations.. El papel de la translocación en la recuperación de poblaciones de caribú del bosque. Conservation Biology, 26(3):365–373, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Drummond:2010:ITS

[DWM^{+}10] Simon P. Drummond, Kerrie A. Wilson, Erik Meijaard, Matthew Watts, Rona Dennis, Lenny Christy, and Hugh P. Possingham. Influence of a threatened-species focus on conservation planning. Influencia del enfoque en una especie amenazada sobre la planificación de la conservación. Conservation Biology, 24(2):441–449, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Douglass:2019:HPC

[DWM^{+}19] Kristina Douglass, Jonathan Walz, Eréndira Quintana Morales, Richard Marcus, Garth Myers, and Jacques Pollini. Historical perspectives on contemporary human–environment dynamics in southeast Africa. Conservation Biology, 33(2):260–274, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dyson:2015:LCM

[Dys15] Philippa Dyson. Lion conservation in Maasailand: response to Hazzah et al. 2014. Conservation Biology, 29(3):939–941, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Estevez:2015:CVR

[EAPB15] Rodrigo A. Estévez, Christopher B. Anderson, J. Cristobal Pizarro, and Mark A. Burgman. Clarifying values, risk perceptions, and attitudes to resolve or avoid social conflicts in invasive species management. Conservation Biology, 29(1):19– 30, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Eads:2015:PBT

[EB15] David A. Eads and Dean E. Biggins. Plague bacterium as a transformer species in prairie dogs and the grasslands of western North America. Conservation Biology, 29(4):1086–1093, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Eggert:2014:UGP

[EBL^{+}14] L. S. Eggert, R. Buij, M. E. Lee, P. Campbell, F. Dallmeier, R. C. Fleischer, A. Alonso, and J. E. Maldonado. Using genetic profiles of African forest elephants to infer population structure, movements, and habitat use in a conservation and development landscape in Gabon. Conservation Biology, 28(1): 107–118, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ekblom:2018:GSC

[EBP^{+}18] Robert Ekblom, Birte Brechlin, Jens Persson, Linnéa Smeds, Malin Johansson, Jessica Magnusson, Øystein Flagstad, and Hans Ellegren. Genome sequencing and conservation genomics in the Scandinavian wolverine population. Conservation Biology, 32(6):1301–1312, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Evans:2014:AMA

[EBRS14] Kristine O. Evans, L. Wes Burger, Jr., Sam Riffell, and Mark D. Smith. Assessing multiregion avian benefits from strategically targeted agricultural buffers. Conservation Biology, 28(4):892– 901, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ellis:2013:IAD

[ECES13] Christopher J. Ellis, Brian J. Coppins, Sally Eaton, and Janet Simkin. Implications of ash dieback for associated epiphytes.

Conservation Biology, 28(4):899–901, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Earnst:2012:CAP

[EDB12] Susan L. Earnst, David S. Dobkin, and Jennifer A. Ballard. Changes in avian and plant communities of aspen woodlands over 12 years after livestock removal in the Northwestern Great Basin. Cambios en las comunidades de aves y plantas en bosques de Álamo a lo largo de 12 años después de la remoción de ganado en la Gran Cuenca Noroccidental. Conservation Biology, 26(5):862–872, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ebrahimi:2015:MCT

[EEB15] Mehregan Ebrahimi, Esmaeil Ebrahimie, and C. Michael Bull. Minimizing the cost of translocation failure with decision-tree models that predict species’ behavioral response in translocation sites. Conservation Biology, 29(4):1208–1216, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Estelle:2012:ETE

[EG12] Veronica Estelle and Edwin D. Grosholz. Experimental test of the effects of a non-native invasive species on a wintering shorebird prueba experimental de los efectos de una especie no nativa invasora sobre una especie de ave playera invernante. Conservation Biology, 26(3):472–481, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Edwards:2013:EPM

[EG13] Felicity N. Edwards and Michael L. Gibeau. Engaging people in meaningful problem solving. Conservation Biology, 27(2):239– 241, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Elliot:2017:TAP

[EG17] Nicholas B. Elliot and Arjun M. Gopalaswamy. Toward accurate and precise estimates of lion density. Conservation Biology, 31(4):934–943, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Evans:2018:PMC

[EGB18] Kristen Evans, Manuel R. Guariguata, and Pedro H. S. Brancalion. Participatory monitoring to connect local and global priorities for forest restoration. Conservation Biology, 32(3): 525–534, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrich:2014:INE

[Ehr14] Dorothee Ehrich. Introductions to northern ecology in a time of change. Conservation Biology, 28(6):1740–1742, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ellis:2014:SEP

[EIS14] Martha M. Ellis, Jacob S. Ivan, and Michael K. Schwartz. Spatially explicit power analyses for occupancy-based monitoring of wolverine in the U.S. Rocky Mountains. Conservation Biology, 28(1):52–62, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Elsen:2017:IAL

[EKRW17] Paul R. Elsen, Ramnarayan Kalyanaraman, Krishnamurthy Ramesh, and David S. Wilcove. The importance of agricultural lands for Himalayan birds in winter. Conservation Biology, 31 (3):416–426, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Eggers:2010:VSO

[EMD^{+}10] Britta Eggers, Andrea Matern, Claudia Drees, Jan Eggers, Werner Härdtle, and Thorsten Assmann. Value of semi-open corridors for simultaneously connecting open and wooded habitats: a case study with ground beetles. El valor de corredores semiabiertos para la conexión simultánea de hábitats abiertos y boscosos: un caso de estudio con escarabajos del suelo. Conservation Biology, 26(2):256–266, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Eldridge:2014:TUL

[EMJ14] Mark D. B. Eldridge, Paul D. Meek, and Rebecca N. Johnson. Taxonomic uncertainty and the loss of biodiversity on Christmas Island, Indian Ocean. Conservation Biology, 28(2):572– 579, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Elbroch:2011:VLC

[EMS^{+}11] Mark Elbroch, Tuyeni H. Mwampamba, Maria J. Santos, Maxine Zylberberg, Louis Liebenberg, James Minye, Christopher Mosser, and Erin Reddy. The value, limitations, and challenges of employing local experts in conservation research. El valor, limitaciones y retos del empleo de expertos locales en la investigación sobre conservación. Conservation Biology, 25(6): 1195–1202, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Earl:2018:QTI

[ENW^{+}18] Julia E. Earl, Sam Nicol, Ruscena Wiederholt, Jay E. Diffendorfer, Darius Semmens, D. T. Tyler Flockhart, Brady J. Mattsson, Gary McCracken, D. Ryan Norris, Wayne E. Thogmartin, and Laura López-Hoffman. Quantitative tools for implementing the new definition of significant portion of the range in the U.S. Endangered Species Act. Conservation Biology, 32 (1):35–49, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Estes:2014:UCA

[EPB^{+}14] Lyndon D. Estes, Lydie-Line Paroz, Bethany A. Bradley, Jonathan M. H. Green, David G. Hole, Stephen Holness, Guy Ziv, Michael G. Oppenheimer, and David S. Wilcove. Using changes in agricultural utility to quantify future climateinduced risk to conservation. Conservation Biology, 28(2):427– 437, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Eskew:2012:ERF

[EPD12] Evan A. Eskew, Steven J. Price, and Michael E. Dorcas. Effects of river-flow regulation on anuran occupancy and abundance in riparian zones. Efectos de la regulación del flujo de un río sobre la ocupación y abundancia de anuros en zonas ribereñas. Conservation Biology, 26(3):504–512, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Eiben:2014:AAD

[ER14] Jesse Eiben and Daniel Rubinoff. Application of agriculturedeveloped demographic analysis for the conservation of the Hawaiian alpine wekiu bug. Conservation Biology, 28(4):1077– 1088, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Eyvindson:2017:LPC

[ERBM17] Kyle Eyvindson, Anna Repo, Daniel Burgas, and Mikko Mönkkönen. Landowner preferences and conservation prioritization: response to Nielsen et al. Conservation Biology, 31 (6):1488–1490, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Egoh:2010:SBE

[ERC^{+}10] Benis N. Egoh, Belinda Reyers, Josie Carwardine, Michael Bode, Patrick J. O’Farrell, Kerrie A. Wilson, Hugh P. Possingham, Mathieu Rouget, Willem De Lange, David M. Richardson, and Richard M. Cowling. Safeguarding biodiversity and ecosystem services in the Little Karoo, South Africa. Salvaguardando la biodiversidad y los servicios del ecosistema en Pequeño Karoo, Sudáfrica. Conservation Biology, 24(4):1021– 1030, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Elsen:2018:CHB

[ERW18] Paul R. Elsen, Krishnamurthy Ramesh, and David S. Wilcove. Conserving Himalayan birds in highly seasonal forested and agricultural landscapes. Conservation Biology, 32(6):1313– 1324, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ellison:2012:NCB

[ESCF12] W. T. Ellison, B. L. Southall, C. W. Clark, and A. S. Frankel. A new context-based approach to assess marine mammal behavioral responses to anthropogenic sounds. Un método nuevo, basado en el contexto, para evaluar respuestas de mamíferos marinos a sonidos Antropogénicos. Conservation Biology, 26 (2):21–28, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ekernas:2017:DPN

[ESD^{+}17] L. Stefan Ekernas, Wesley M. Sarmento, Hannah S. Davie, Richard P. Reading, James Murdoch, Ganchimeg J. Wingard, Sukh Amgalanbaatar, and Joel Berger. Desert pastoralists’ negative and positive effects on rare wildlife in the gobi. Conservation Biology, 31(2):269–277, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ellwood:2019:I

[ESK19] Elizabeth R. Ellwood, Pamela S. Soltis, and Mary L. Klein. Introduction. Conservation Biology, 33(3):498–499, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Estes:2010:UEF

[ETB10] James A. Estes, M. Tim Tinker, and James L. Bodkin. Using ecological function to develop recovery criteria for depleted species: Sea otters and kelp forests in the Aleutian Archipelago. Utilización de la Función ecológica para desarrollar criterios de recuperación para especies disminuidas: Nutrias marinas y bosques de algas en el Archipiélago Aleutiano. Conservation Biology, 25(6):852–860, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Emer:2013:EDI

[EVF13] Carine Emer, Eduardo Martins Venticinque, and Carlos Roberto Fonseca. Effects of dam-induced landscape fragmentation on Amazonian ant–plant mutualistic networks. Conservation Biology, 27(4):763–773, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ewen:2015:ISF

[EWCG15] John G. Ewen, Leila Walker, Stefano Canessa, and Jim J. Groombridge. Improving supplementary feeding in species conservation. Conservation Biology, 29(2):341–349, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Edwards:2013:TFP

[EWN^{+}13] David P. Edwards, Paul Woodcock, Rob J. Newton, Felicity A. Edwards, David J. R. Andrews, Teegan D. S. Docherty, Simon L. Mitchell, Takahiro Ota, Suzan Benedick, Simon H. Bottrell, and Keith C. Hamer. Trophic flexibility and the persistence of understory birds in intensively logged rainforest. Conservation Biology, 28(4):1079–1086, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fattorini:2014:RBS

[Fat14] Simone Fattorini. Relations between species rarity, vulnerability, and range contraction for a beetle group in a densely populated region in the Mediterranean biodiversity hotspot. Conservation Biology, 28(1):169–176, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ferreira:2011:PRD

[FB11] Pedro M. A. Ferreira and Ilsi I. Boldrini. Potential reflection of distinct ecological units in plant endemism categories. Reflejo potencial de unidades ecológicas distintas en las Categorías de endemismo de plantas. Conservation Biology, 25(4):672–679, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fisher:2012:IEM

[FB12] Diana O. Fisher and Simon P. Blomberg. Inferring extinction of mammals from sighting records, threats, and biological traits. Infiriendo la extinción de mamíferos a partir de registros visuales, amenazas y atributos Biológicos. Conservation Biology, 26(2):57–67, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Faruk:2013:EOP

[FBA^{+}13] Aisyah Faruk, Daicus Belabut, Norhayati Ahmad, Robert J. Knell, and Trenton W. J. Garner. Effects of oil-palm plantations on diversity of tropical anurans. Conservation Biology, 27(3):615–624, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Foster:2014:IDU

[FBB^{+}14] Michael J. Foster, Mary E. Blair, Chanda Bennett, Nora Bynum, and Eleanor J. Sterling. Increasing the diversity of U.S. conservation science professionals via the Society for Conservation Biology. Conservation Biology, 28(1):288–291, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Frankham:2011:PPO

[FBE^{+}11] Richard Frankham, Jonathan D. Ballou, Mark D. B. Eldridge, Robert C. Lacy, Katherine Ralls, Michele R. Dudash, and Charles B. Fenster. Predicting the probability of outbreeding depression. Predicción de la depresión por exogámica. Conservation Biology, 25(3):465–475, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fisher:2014:MRF

[FBF^{+}14] Brendan Fisher, Andrew Balmford, Paul J. Ferraro, Louise Glew, Michael Mascia, Robin Naidoo, and Taylor H. Ricketts. Moving Rio forward and avoiding 10 more years with little evidence for effective conservation policy. Conservation Biology, 28(5):880–882, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fernandez-Bellon:2019:EDW

[FBWIO19] Darío Fernández-Bellon, Mark W. Wilson, Sandra Irwin, and John O’Halloran. Effects of development of wind energy and associated changes in land use on bird densities in upland areas. Conservation Biology, 33(2):413–422, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ford:2010:VPT

[FC10] Adam T. Ford and Anthony P. Clevenger. Validity of the preytrap hypothesis for carnivore-ungulate interactions at wildlifecrossing structures. Validez de la hipótesis de la trampa del depredador para interacciones carnívoro–ungulado en estructuras para cruce de vida silvestre. Conservation Biology, 24 (6):1679–1685, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Foley:2011:IMR

[FCC^{+}11] Janet Foley, Deana Clifford, Kevin Castle, Paul Cryan, and Richard S. Ostfeld. Investigating and managing the rapid emergence of white-nose syndrome, a novel, fatal, infectious disease of hibernating bats. Investigando y manejando la rápida emergencia del síndrome de nariz blanca, una enfermedad infecciosa, nueva, fatal, en murciélagos invernantes. Conservation Biology, 25(2):223–231, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Feary:2011:ECM

[FCGJH11] David A. Feary, Joshua E. Cinner, Nicholas A. J. Graham, and Fraser A. Januchowski-Hartley. Effects of customary marine closures on fish behavior, spear-fishing success, and underwater visual surveys. Efecto de encierros marinos convencionales sobre el comportamiento de peces, el Éxito de la pesca con arpón y los muestreos visuales submarinos. Conservation Biology, 26 (3):341–349, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ficetola:2019:RIH

[FCS19] Gentile Francesco Ficetola, Claudia Canedoli, and Fabio Stoch. The racovitzan impediment and the hidden biodiversity of unexplored environments. Conservation Biology, 33(2):214–216, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fletcher:2016:QRF

[FDBA16] Robert Fletcher, Wolfram Dressler, Bram Büscher, and Zachary R. Anderson. Questioning REDD+ and the future of market-based conservation. Conservation Biology, 31(1):673– 675, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fletcher:2017:DRI

[FDBA17] Robert Fletcher, Wolfram Dressler, Bram Büscher, and Zachary R. Anderson. Debating REDD+ and its implications: reply to Angelsen et al. Conservation Biology, 31(3):721–723, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Flesch:2010:PEU

[FEC^{+}10] Aaron D. Flesch, Clinton W. Epps, James W. Cain III, Matt Clark, Paul R. Krausman, and John R. Morgart. Potential effects of the United States–Mexico border fence on wildlife. Efectos potenciales de la barda en la frontera Estados Unidos– México sobre la vida silvestre. Conservation Biology, 24(1): 171–181, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ferraro:2011:FPE

[Fer11] Paul J. Ferraro. The future of payments for environmental services. Conservation Biology, 25(6):1134–1138, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ferrante:2017:EML

[Fer17] Marco Ferrante. Extinction, mechanisms of life, and practice of conservation on an island. Conservation Biology, 31(5): 1208–1209, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ferrante:2019:SEN

[Fer19] Marco Ferrante. Seeking explanations for New Zealand’s biodiversity. Conservation Biology, 33(3):737–738, June 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fovargue:2019:LCP

[FFHA19] Rachel Fovargue, Maria Fisher, Jamal Harris, and Paul R. Armsworth. A landscape of conservation philanthropy for U.S. land trusts. Conservation Biology, 33(2):176–184, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fa:2014:MHT

[FFM^{+}14] Julia E. Fa, Miguel Angel Farfán, Ana Luz Marquez, Jesús Duarte, Janet Nackoney, Amy Hall, Jef Dupain, Sarah Seymour, Paul J. Johnson, David W. Macdonald, and J. Mario Vargas. Mapping hotspots of threatened species traded in bushmeat markets in the Cross–Sanaga Rivers region. Conservation Biology, 28(1):224–233, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fields:2018:SCI

[FFS^{+}18] Andrew T. Fields, Gunter A. Fischer, Stanley K. H. Shea, Huarong Zhang, Debra L. Abercrombie, Kevin A. Feldheim, Elizabeth A. Babcock, and Demian D. Chapman. Species composition of the international shark fin trade assessed through a retail-market survey in Hong Kong. Conservation Biology, 32 (2):376–389, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fitzgerald:2012:EVC

[FG12] Katherine Fitzgerald and Deborah M. Gordon. Effects of vegetation cover, presence of a native ant species, and human disturbance on colonization by Argentine ants. Efectos de la cobertura vegetal, presencia de especies de hormigas nativas y perturbación humana sobre la colonización por hormigas argentinas. Conservation Biology, 26(3):525–538, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Finch:2019:BCL

[FGG^{+}19] Tom Finch, Simon Gillings, Rhys E. Green, Dario Massimino, Will J. Peach, and Andrew Balmford. Bird conservation and the land sharing-sparing continuum in farmland-dominated landscapes of lowland England. Conservation Biology, 33(5): 1045–1055, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fyumagwa:2013:RTS

[FGH^{+}13] R. Fyumagwa, E. Gereta, S. Hassan, J. R. Kideghesho, E. M. Kohi, J. Keyyu, F. Magige, I. M. Mfunda, A. Mwakatobe, J. Ntalwila, J. W. Nyahongo, V. Runyoro, and E. Røskaft. Roads as a threat to the Serengeti ecosystem. Conservation Biology, 28(4):1122–1125, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Friedenberg:2018:EAR

[FHBK18] Nicholas A. Friedenberg, Jan Jeffrey Hoover, Krista Boysen, and K. Jack Killgore. Estimating abundance without recaptures of marked pallid sturgeon in the Mississippi River. Conservation Biology, 32(2):457–465, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fyumagwa:2015:HRC

[FHK^{+}15] R. Fyumagwa, S. Hassan, J. R. Kideghesho, E. M. Kohi, F. Magige, I. M. Mfunda, A. Mwakatobe, J. Ntalwila, J. W. Nyahongo, V. Runyoro, and E. Røskaft. Human rights and conservation of biodiversity considerations associated with roads in the Serengeti: response to Hopcraft et al. Conservation Biology, 29(3):937–938, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fiedler:2015:BRF

[Fie15] Peggy L. Fiedler. Book reviews: The Fascinating Ecology of Two Megadiverse Southern Hemisphere Ecosystems Plant Life on the Sandplains in Southwest Australia. H. Lambers, editor. 2014. University of Western Australia Press, Perth, Australia. 350 pp. $ 69.99 (hardcover). ISBN 13 9871742585642. Fynbos. Ecology, Evolution, and Conservation of a Megadiverse Region. N. Allsopp, J. F. Colville, G. A. Verboom, editors. R. A. Cowling, consulting editor. 2014. Oxford University Press, Oxford, UK. 382 p p. $104.50 (hardcover). ISBN 13 978-0-19-967958-4. Conservation Biology, 29(6):1727–1729, December 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fisher:2012:ICC

[Fis12] William H. Fisher. Interdisciplinary crafting of the conservation agenda. Conservation Biology, 26(5):949–950, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Foxcroft:2011:PAB

[FJP^{+}11] Llewellyn C. Foxcroft, Vojtěch Jarošík, Petr Pyšek, David M. Richardson, and Mathieu Rouget. Protected-area boundaries as filters of plant invasions. Límites de Áreas protegidas como filtros para la invasión de plantas. Conservation Biology, 26 (3):400–405, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Farmer:2011:MIA

[FKM^{+}11] James R. Farmer, Doug Knapp, Vicky J. Meretsky, Charles Chancellor, and Burnell C. Fischer. Motivations influencing the adoption of conservation easements. Motivaciones que influyen en la adopción de contratos de conservación. Conservation Biology, 25(4):827–834, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Freeman:2017:CHF

[FKRC17] Johanna Freeman, Leda Kobziar, Elizabeth White Rose, and Wendell Cropper. A critique of the historical-fire-regime concept in conservation. Conservation Biology, 31(5):976–985, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleishman:2010:TRS

[Fle10] Erica Fleishman. The task remains the same. Conservation Biology, 24(1):1–3, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Flight:2010:PCM

[Fli10]  Patrick A. Flight.  Phylogenetic comparative methods

strengthen evidence for reduced genetic diversity among endangered tetrapods. Métodos filogenéticos comparativos fortalecen la evidencia de la reducción de diversidad genética entre tetrápodos en peligro. Conservation Biology, 24(5):1307–1315, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fullerton:2011:HIS

[FLP^{+}11] Aimee H. Fullerton, Steven T. Lindley, George R. Pess, Blake E. Feist, E. Ashley Steel, and Paul Mcelhany. Human influence on the spatial structure of threatened Pacific salmon metapopulations. Influencia humana sobre la estructura espacial de metapoblaciones de salmón del Pacífico amenazadas. Conservation Biology, 25(5):932–944, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fuller:2014:AOA

[FLW14a] Richard A. Fuller, Jasmine R. Lee, and James E. M. Watson. Achieving open access to conservation science. Conservation Biology, 28(6):1550–1557, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fuller:2014:COA

[FLW14b] Richard A. Fuller, Jasmine R. Lee, and James E. M. Watson. Choosing open access. Conservation Biology, 28(6):1443–1444, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fogell:2018:TCI

[FMB^{+}18] Deborah J. Fogell, Rowan O. Martin, Nancy Bunbury, Becki Lawson, James Sells, Alison M. McKeand, Vikash Tatayah, Cao Tien Trung, and Jim J. Groombridge. Trade and conservation implications of new beak and feather disease virus detection in native and introduced parrots. Conservation Biology, 32(6):1325–1335, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Franklin:2010:CBP

[FMF10] Kimberly Franklin and Francisco Molina-Freaner. Consequences of buffelgrass pasture development for primary productivity, perennial plant richness, and vegetation structure in the drylands of Sonora, Mexico. Consecuencias del desarrollo de praderas de pennisetum ciliare sobre la productividad y riqueza de especies de plantas en tierras Áridas de Sonora, México. Conservation Biology, 24(6):1664–1673, December 2010. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forero-Medina:2011:CSE

[FMJP11a] German Forero-Medina, Lucas Joppa, and Stuart L. Pimm. Constraints to species’ elevational range shifts as climate changes restricciones a los cambios de rango altitudinal de especies a medida que cambia el clima. Conservation Biology, 25(1):163–171, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forero-Medina:2011:TTR

[FMJP11b] German Forero-Medina, Lucas Joppa, and Stuart L. Pimm. Thermal tolerance, range expansion, and status of tropical amphibians: Reply to Catenazzi. Conservation Biology, 25 (3):426–427, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Feger:2019:FPN

[FMV^{+}19] Clément Feger, Laurent Mermet, Bhaskar Vira, Prue F. E. Addison, Richard Barker, Frank Birkin, John Burns, Stuart Cooper, Denis Couvet, Thomas Cuckston, Gretchen C. Daily, Colin Dey, Louise Gallagher, Rosemary Hails, Stephen Jollands, Georgina Mace, Emily Mckenzie, Markus Milne, Paolo Quattrone, Alexandre Rambaud, Shona Russell, Marta Santamaria, and William J. Sutherland. Four priorities for new links between conservation science and accounting research. Conservation Biology, 33(4):972–975, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fa:2015:CBM

[FOF^{+}15] Julia E. Fa, Jesus Olivero, Miguel Á. Farfán, Ana L. Márquez, Jesús Duarte, Janet Nackoney, Amy Hall, Jef Dupain, Sarah Seymour, Paul J. Johnson, David W. Macdonald, Raimundo Real, and Juan M. Vargas. Correlates of bushmeat in markets and depletion of wildlife. Conservation Biology, 29(3):805–815, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forbes:2011:TCC

[For11] Peter Forbes. Transforming conservation for the 21st century. Conservation Biology, 25(2):209–211, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fritz:2010:SME

[FP10] Susanne A. Fritz and Andy Purvis. Selectivity in mammalian extinction risk and threat types: a new measure of phylogenetic signal strength in binary traits. Selectividad en el riesgo de extinción y tipos de amenaza en mamíferos: una nueva medida de la intensidad de la señal filogenética en atributos binarios. Conservation Biology, 24(4):1042–1051, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fischman:2016:JAM

[FR16] Robert L. Fischman and J. B. Ruhl. Judging adaptive management practices of U.S. agencies. Conservation Biology, 30 (2):268–275, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Francis:2011:EES

[Fra11] John Francis. Empowering and encouraging scientists to act. Conservation Biology, 26(3):415–416, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Freudenburg:2011:EAA

[Fre11] William R. Freudenburg. Evidence and assertions in assessing human factors. Conservation Biology, 25(6):635–636, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fraser:2017:IMS

[FRY^{+}17] Hannah Fraser, Libby Rumpff, Jian D. L. Yen, Doug Robinson, and Brendan A. Wintle. Integrated models to support multiobjective ecological restoration decisions. Conservation Biology, 31(6):1418–1427, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fischman:2015:LPC

[FS15] Robert L. Fischman and James Salzman. Lessons from pollution control: response to Heller and Hobbs 2014. Conservation Biology, 29(3):950–952, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [HH14].

Fischer:2015:CWR

[FSAM15] Jason D. Fischer, Sarah C. Schneider, Adam A. Ahlers, and James R. Miller. Categorizing wildlife responses to urbanization and conservation implications of terminology. Conservation Biology, 29(4):1246–1248, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fraser:2017:VVC

[FSJ^{+}17] Hannah Fraser, Kylie Soanes, Stuart A. Jones, Chris S. Jones, and Matthew Malishev. The value of virtual conferencing for ecology and conservation. Conservation Biology, 31(3):540– 546, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Faulkner:2018:SAC

[FSR^{+}18] S. C. Faulkner, M. C. A. Stevens, S. S. Romañach, P. A. Lindsey, and S. C. Le Comber. A spatial approach to combatting wildlife crime. Conservation Biology, 32(3):685–693, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Franco:2019:ADS

[FSSW19] André L. C. Franco, Bruno W. Sobral, Artur L. C. Silva, and Diana H. Wall. Amazonian deforestation and soil biodiversity. Conservation Biology, 33(3):590–600, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Friess:2016:PCA

[FTB^{+}16] Daniel A. Friess, Benjamin S. Thompson, Ben Brown, A. Aldrie Amir, Clint Cameron, Heather J. Koldewey, Sigit D. Sasmito, and Frida Sidik. Policy challenges and approaches for the conservation of mangrove forests in Southeast Asia. Conservation Biology, 30(5):933–949, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ford-Thompson:2012:SPM

[FTSSW12] Adriana E. S. Ford-Thompson, Carolyn Snell, Glen Saunders, and Piran C. L. White. Stakeholder participation in management of invasive vertebrates. Participación de grupos de interesados en el manejo de vertebrados invasores. Conservation Biology, 26(2):345–356, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Furtado:2018:FGS

[Fur18] Miguel Furtado. Finally, a global synthesis on marine fisheries. Conservation Biology, 32(6):1478–1479, December 2018. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Furtado:2019:HIN

[Fur19a] Miguel Furtado. Human impacts on and new research fronts to understand our endangered oceans. Conservation Biology, 33 (5):1224–1226, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Furtado:2019:OOT

[Fur19b] Miguel Furtado. Our oceans: A treasure in need of protection. Conservation Biology, 33(2):221–222, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fung:2017:PIP

[FW17] Han Chi Fung and Robin S. Waples. Performance of IUCN proxies for generation length. Conservation Biology, 31(4):883– 893, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fischer:2015:THW

[FWC^{+}15] Anke Fischer, Yitbarek Tibebe Weldesemaet, Mikołaj Czajkowski, Degu Tadie, and Nick Hanley. Trophy hunters’ willingness to pay for wildlife conservation and community benefits. Conservation Biology, 29(4):1111–1121, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fournier:2019:SSB

[FWH19] Auriel M. V. Fournier, Easton R. White, and Stephen B. Heard. Site-selection bias and apparent population declines in longterm studies. Conservation Biology, 33(6):1370–1379, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Foerster:2012:CBH

[FWM^{+}12] Steffen Foerster, David S. Wilkie, Gilda A. Morelli, Josefien Demmer, Malcolm Starkey, Paul Telfer, Matthew Steil, and Arthur Lewbel. Correlates of bushmeat hunting among remote rural households in Gabon, Central Africa. Correlaciones de la caza de carne de monte en grupos familiares rurales en Áreas remotas de Gabón, África. Conservation Biology, 26(2):335– 344, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fisher:2014:SPB

[FWM14] Jason T. Fisher, Matthew Wheatley, and Darryl Mackenzie. Spatial patterns of breeding success of grizzly bears derived from hierarchical multistate models. Conservation Biology, 28 (6):1249–1259, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Farrell:2018:LTC

[FZ18] Julianne Farrell and Myron P. Zalucki. Long-term consequences of European invasions. Conservation Biology, 32(4):966–968, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grange:2014:NAT

[GAFG14] Zoë L. Grange, Mary Van Andel, Nigel P. French, and Brett D. Gartrell. Network analysis of translocated takahē populations to identify disease surveillance targets. Conservation Biology, 28(2):518–528, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gregory:2013:SDM

[GAG13] Robin Gregory, Joseph Arvai, and Leah R. Gerber. Structuring decisions for managing threatened and endangered species in a changing climate. Conservation Biology, 27(6):1212–1221, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gonzalez:2011:HGE

[GB11] Andrew Gonzalez and Beatrix Beisner. Homage to G. Evelyn Hutchinson. Conservation Biology, 25(6):1253–1254, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gregory:2014:RVE

[GB14] Andrew J. Gregory and Paul Beier. Response variables for evaluation of the effectiveness of conservation corridors. Conservation Biology, 28(3):689–695, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Golden:2014:EVS

[GBB^{+}14] Christopher D. Golden, Matthew H. Bonds, Justin S. Brashares, B. J. Rodolph Rasolofoniaina, and Claire Kremen. Economic valuation of subsistence harvest of wildlife in Madagascar. Conservation Biology, 28(1):234–243, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gill:2015:MYP

[GBB^{+}15] Jacquelyn L. Gill, Jessica L. Blois, Blas Benito, Solomon Dobrowski, Malcolm L. Hunter, Jr., and Jenny L. McGuire. A 2.5-million-year perspective on coarse-filter strategies for conserving nature’s stage. Conservation Biology, 29(3):640–648, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garnett:2019:MPU

[GBB^{+}19a] S. T. Garnett, S. H. M. Butchart, G. B. Baker, E. Bayraktarov, K. L. Buchanan, A. A. Burbidge, A. L. M. Chauvenet, L. Christidis, G. Ehmke, M. Grace, D. G. Hoccom, S. M. Legge, I. Leiper, D. B. Lindenmayer, R. H. Loyn, M. Maron, P. McDonald, P. Menkhorst, H. P. Possingham, J. Radford, A. E. Reside, D. M. Watson, J. E. M. Watson, B. Wintle, J. C. Z. Woinarski, and H. M. Geyle. Metrics of progress in the understanding and management of threats to Australian birds. Conservation Biology, 33(2):456–468, April 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Green:2019:RCG

[GBB^{+}19b] Elizabeth J. Green, Graeme M. Buchanan, Stuart H. M. Butchart, Georgina M. Chandler, Neil D. Burgess, Samantha L. L. Hill, and Richard D. Gregory. Relating characteristics of global biodiversity targets to reported progress. Conservation Biology, 33(6):1360–1369, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Griffiths:2019:NNL

[GBBMG19] Victoria F. Griffiths, Joseph W. Bull, Julia Baker, and E. J. Milner-Gulland. No net loss for people and biodiversity. Conservation Biology, 33(2):76–87, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Griffiths:2015:MPD

[GBD^{+}15] Kate E. Griffiths, Sharon T. Balding, John B. Dickie, Gwilym P. Lewis, Tim R. Pearce, and Richard Grenyer. Maximizing the phylogenetic diversity of seed banks. Conservation Biology, 29(2):370–381, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garcia:2010:BCA

[GBG^{+}10] Claude A. Garcia, Shonil A. Bhagwat, Jaboury Ghazoul, Cheryl D. Nath, Konerira M. Nanaya, Chepudira G. Kushalappa, Yenugula Raghuramulu, Robert Nasi, and Philippe Vaast. Biodiversity conservation in agricultural landscapes: Challenges and opportunities of coffee agroforests in the Western Ghats, India. Conservación de biodiversidad en paisajes Agrícolas: Retos y oportunidades de agrobosques de café en los Ghats Occidentales, India. Conservation Biology, 24(2):479– 488, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gomez-Baggethun:2010:TEK

[GBMRG^{+}10] Erik Gómez-Baggethun, Sara Mingorría, Victoria Reyes-García, Laura Calvet, and Carlos Montes. Traditional ecological knowledge trends in the transition to a market economy: Empirical study in the Doñana natural areas. Tendencias del conocimiento ecológico tradicional en la Transición a una economía de mercado: Estudio empírico en Áreas naturales en Doñana. Conservation Biology, 24(3):721–729, June 2010. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garrard:2015:IDT

[GBMW15] Georgia E. Garrard, Sarah A. Bekessy, Michael A. McCarthy, and Brendan A. Wintle. Incorporating detectability of threatened species into environmental impact assessment. Conservation Biology, 29(1):216–225, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Groot:2013:BIE

[GBP^{+}13] Rudolf S. De Groot, James Blignaut, Sander Van Der Ploeg, James Aronson, Thomas Elmqvist, and Joshua Farley. Benefits of investing in ecosystem restoration. Conservation Biology, 27 (6):1286–1293, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guidetti:2010:CPE

[GC10] Paolo Guidetti and Joachim Claudet. Comanagement practices enhance fisheries in Marine Protected Areas. Las prácticas de co-manejo enriquecen las pesquerías en Áreas Marinas Protegidas. Conservation Biology, 26(2):312–318, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gallo-Cajiao:2018:CBC

[GCAF^{+}18] Eduardo Gallo-Cajiao, Carla Archibald, Rachel Friedman, Rochelle Steven, Richard A. Fuller, Edward T. Game, Tiffany H. Morrison, and Euan G. Ritchie. Crowdfunding biodiversity conservation. Conservation Biology, 32(6):1426–1435, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gotelli:2012:SBM

[GCC^{+}12] Nicholas J. Gotelli, Anne Chao, Robert K. Colwell, Wen-Han Hwang, and Gary R. Graves. Specimen-based modeling, stopping rules, and the extinction of the ivory-billed woodpecker. Modelado basado en especímenes, reglas de decisión y la extinción de Campephilus principalis. Conservation Biology, 26 (2):47–56, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gelcich:2015:IBC

[GD15] Stefan Gelcich and C. Josh Donlan. Incentivizing biodiversity conservation in artisanal fishing communities through territorial user rights and business model innovation. Conservation Biology, 29(4):1076–1085, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gaston:2015:QEN

[GDB15] Kevin J. Gaston, James P. Duffy, and Jonathan Bennie. Quantifying the erosion of natural darkness in the global protected area system. Conservation Biology, 29(4):1132–1141, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goettsch:2019:GGA

[GDG19] Bárbara Goettsch, América Paz Durán, and Kevin J. Gaston. Global gap analysis of cactus species and priority sites for their conservation. Conservation Biology, 33(2):369–376, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gupta:2019:RDO

[GDM^{+}19] Amrita Gupta, Bistra Dilkina, Dana J. Morin, Angela K. Fuller, J. Andrew Royle, Christopher Sutherland, and Carla P. Gomes. Reserve design to optimize functional connectivity and animal density. Conservation Biology, 33(5):1023–1034, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guerbois:2013:IIC

[GDMF13] Chloe Guerbois, Anne-Beatrice Dufour, Godfrey Mtare, and Herve Fritz. Insights for integrated conservation from attitudes of people toward protected areas near Hwange National Park, Zimbabwe. Conservation Biology, 27(4):844–855, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Genes:2019:EHM

[GFdM^{+}19] Luísa Genes, Fernando A. S. Fernandez, Fernando Z. Vaz de Mello, Patrícia da Rosa, Eduardo Fernandez, and Alexandra S. Pires. Effects of howler monkey reintroduction on ecological interactions and processes. Conservation Biology, 33 (2):88–98, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gelcich:2012:TUR

[GFG^{+}12] Stefan Gelcich, Miriam Fernández, Natalio Godoy, Antonio Canepa, Luis Prado, and Juan Carlos Castilla. Territorial user rights for fisheries as ancillary instruments for marine coastal conservation in Chile. Conservation Biology, 26(6):1005–1015, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Griffith:2012:PIA

[GFGR12] Gary P. Griffith, Elizabeth A. Fulton, Rebecca Gorton, and Anthony J. Richardson. Predicting interactions among fishing, ocean warming, and ocean acidification in a marine system with whole-ecosystem models. Conservation Biology, 26 (6):1145–1152, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Groce:2019:USN

[GFJC19] Julie E. Groce, Megan A. Farrelly, Bradley S. Jorgensen, and Carly N. Cook. Using social-network research to improve outcomes in natural resource management. Conservation Biology, 33(2):53–65, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gleason:2013:TCA

[GFM^{+}13] Mary Gleason, Erika M. Feller, Matt Merrifield, Stephen Copps, Rod Fujita, Michael Bell, Steve Rienecke, and Chuck Cook. A transactional and collaborative approach to reducing effects of bottom trawling. Conservation Biology, 27(3):470– 479, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gill:2012:DCH

[GFRK12] David J. C. Gill, Julia E. Fa, J. Marcus Rowcliffe, and Noëlle F. Kümpel. Drivers of change in hunter offtake and hunting strategies in Sendje, Equatorial Guinea. Conservation Biology, 26(6): 1052–1060, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Greenberg:2013:EIP

[GG13] Daniel A. Greenberg and David M. Green. Effects of an invasive plant on population dynamics in toads. Conservation Biology, 28(4):1049–1057, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gooden:2019:PAO

[GG19] Jennifer Gooden and Richard Grenyer. The psychological appeal of owning private land for conservation. Conservation Biology, 33(2):339–350, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grange:2015:UCC

[GGB^{+}15] Zoe L. Grange, Brett D. Gartrell, Patrick J. Biggs, Nicola J. Nelson, Jonathan C. Marshall, Laryssa Howe, Matthew G. M. Balm, and Nigel P. French. Using a common commensal bacterium in endangered takahe as a model to explore pathogen dynamics in isolated wildlife populations. Conservation Biology, 29(5):1327–1336, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garcia:2010:LEL

[GGG10] María B. García, Daniel Goñi, and David Guzmán. Living at the edge: Local versus positional factors in the long-term population dynamics of an endangered orchid. Viviendo en el borde: Factores posicionales versus locales en la dinámica poblacional a largo plazo de una orquídea en peligro. Conservation Biology, 24(5):1219–1229, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garland:2015:PSH

[GGL^{+}15] Ellen C. Garland, Anne W. Goldizen, Matthew S. Lilley, Melinda L. Rekdahl, Claire Garrigue, Rochelle Constantine, Nan Daeschler Hauser, M. Michael Poole, Jooke Robbins, and Michael J. Noad. Population structure of humpback whales in the western and central South Pacific Ocean as determined by vocal exchange among populations. Conservation Biology, 29 (4):1198–1207, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garcia-Gonzalez:2014:AAC

[GGM14] Ricardo García-González and Antoni Margalida. The arguments against cloning the Pyrenean wild goat. Conservation Biology, 28(6):1445–1446, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Griffiths:2010:PSE

[GH10] Christine J. Griffiths and Stephen Harris. Prevention of secondary extinctions through taxon substitution. Conservation Biology, 24(3):645–646, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gullison:2018:PCC

[GH18] Raymond E. Gullison and Jared Hardner. Progress and challenges in consolidating the management of Amazonian protected areas and indigenous territories. Conservation Biology, 32(5):1020–1030, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gardner:2013:BOC

[GHB^{+}13] Toby A. Gardner, Amrei Von Hase, Susie Brownlie, Jonathan M. M. Ekstrom, John D. Pilgrim, Conrad E. Savy, R. T. Theo Stephens, Jo Treweek, Graham T. Ussher, Gerri Ward, and Kerry Ten Kate. Biodiversity offsets and the challenge of achieving no net loss. Conservation Biology, 27(6):1254–1264, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gowdy:2010:WEC

[GHKK10] John Gowdy, Charles Hall, Kent Klitgaard, and Lisi Krall. What every conservation biologist should know about economic theory. Lo que todo biólogo de la conservación debe saber de teoría económica. Conservation Biology, 24(6):1440–1447, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Giakoumi:2015:TFA

[GHM^{+}15] Sylvaine Giakoumi, Benjamin S. Halpern, Loïc N. Michel, Sylvie Gobert, Maria Sini, Charles-François Boudouresque, Maria-Cristina Gambi, Stelios Katsanevakis, Pierre Lejeune, Monica Montefalcone, Gerard Pergent, Christine Pergent-Martini, Pablo Sanchez-Jerez, Branko Velimirov, Salvatrice Vizzini, Arnaud Abadie, Marta Coll, Paolo Guidetti, Fiorenza Micheli, and Hugh P. Possingham. Towards a framework for assessment and management of cumulative human impacts on marine food webs. Conservation Biology, 29(4):1228–1234, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Giam:2015:MIO

[GHT^{+}15] Xingli Giam, Renny K. Hadiaty, Heok Hui Tan, Lynne R. Parenti, Daisy Wowor, Sopian Sauri, Kwek Yan Chong, Darren C. J. Yeo, and David S. Wilcove. Mitigating the impact of oilpalm monoculture on freshwater fishes in Southeast Asia. Conservation Biology, 29(5):1357–1367, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Geldmann:2014:MCH

[GJB14] Jonas Geldmann, Lucas N. Joppa, and Neil D. Burgess. Mapping change in human pressure globally on land and within protected areas. Conservation Biology, 28(6):1604–1616, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gould:2015:PEN

[GKA^{+}15] Rachelle K. Gould, Sarah C. Klain, Nicole M. Ardoin, Terre Satterfield, Ulalia Woodside, Neil Hannahs, Gretchen C. Daily, and Kai M. Chan. A protocol for eliciting nonmaterial values through a cultural ecosystem services frame. Conservation Biology, 29(2):575–586, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ghazoul:2010:RLW

[GKB10] Jaboury Ghazoul, Lian Pin Koh, and Rhett A. Butler. A REDD light for wildlife-friendly farming. Conservation Biology, 24(3):644–645, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Geyer:2011:CCC

[GKK^{+}11] Juliane Geyer, Iris Kiefer, Stefan Kreft, Veronica Chavez, Nick Salafsky, Florian Jeltsch, and Pierre L. Ibisch. Classification of climate-change-induced stresses on biological diversity. Clasificación de estreses inducidos por el cambio climático en la diversidad biológica. Conservation Biology, 25(4):708–715, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Game:2013:SCM

[GKP13] Edward T. Game, Peter Kareiva, and Hugh P. Possingham. Six common mistakes in conservation priority setting. Conservation Biology, 27(3):480–485, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gotzenberger:2018:TRG

[GL18] Lars Götzenberger and Jan Lepš. The time is ripe for general theory in community ecology. Conservation Biology, 32(2):499– 501, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gossner:2013:CNN

[GLB^{+}13] Martin M. Gossner, Thibault Lachat, Jörg Brunet, Gunnar Isacsson, Christophe Bouget, Hervé Brustel, Roland Brandl, Wolfgang W. Weisser, and Jörg Müller. Current near-to-nature forest management effects on functional trait composition of saproxylic beetles in beech forests. Conservation Biology, 27 (3):605–614, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gedir:2018:EAS

[GLdPL18] Jay V. Gedir, Peter R. Law, Pierre du Preez, and Wayne L. Linklater. Effects of age and sex ratios on offspring recruitment rates in translocated black rhinoceros. Conservation Biology, 32(3):628–637, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gallo:2018:NMP

[GLF^{+}18] Travis Gallo, Elizabeth W. Lehrer, Mason Fidino, R. Julia Kilgour, Patrick J. Wolff, and Seth B. Magle. Need for multiscale planning for conservation of urban bats. Conservation Biology, 32(3):638–647, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gahl:2012:VRN

[GLH12] Megan K. Gahl, Joyce E. Longcore, and Jeff E. Houlahan. Varying responses of northeastern North American amphibians to the chytrid pathogen Batrachochytrium dendrobatidis. Respuestas diversas de anfibios del noreste de Norteamérica al quitridio patógeno Batrachochytrium dendrobatidis. Conservation Biology, 26(2):135–141, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garnett:2016:PSP

[GLJ^{+}16] Stephen T. Garnett, Michael J. Lawes, Robyn James, Kristen Bigland, and Kerstin K. Zander. Portrayal of sustainability principles in the mission statements and on home pages of the world’s largest organizations. Conservation Biology, 30(2):297– 307, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grajal:2017:CRB

[GLK^{+}17] Alejandro Grajal, Jerry F. Luebke, Lisa-Anne DeGregoria Kelly, Jennifer Matiasek, Susan Clayton, Bryan T. Karazsia, Carol D. Saunders, Susan R. Goldman, Michael E. Mann, and Ricardo Stanoss. The complex relationship between personal sense of connection to animals and self-reported proenvironmental behaviors by zoo visitors. Conservation Biology, 31 (3):322–330, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grueber:2010:IDA

[GLNJ10] Catherine E. Grueber, Rebecca J. Laws, Shinichi Nakagawa, and Ian G. Jamieson. Inbreeding depression accumulation across life-history stages of the endangered takahe. Acumulación de depresión endogámica en etapas de la historia de vida de porphyrio hochstetteri en peligro. Conservation Biology, 24 (6):1617–1625, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garcia-Morales:2013:RNB

[GMBM13] Rodrigo García-Morales, Ernesto I. Badano, and Claudia E. Moreno. Response of neotropical bat assemblages to human land use. Conservation Biology, 28(4):1096–1106, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guerrero:2013:SMC

[GMCW13] Angela M. Guerrero, Ryan R. J. Mcallister, Jonathan Corcoran, and Kerrie A. Wilson. Scale mismatches, conservation planning, and the value of social-network analyses. Conservation Biology, 27(1):35–44, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Graham:2019:OBU

[GMG^{+}19] Sonia Graham, Alexander L. Metcalf, Nicholas Gill, Rebecca Niemiec, Carlo Moreno, Thomas Bach, Victoria Ikutegbe, Lars Hallstrom, Zhao Ma, and Alice Lubeck. Opportunities for better use of collective action theory in research and governance for invasive species management. Conservation Biology, 33(2):275– 287, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gibson:2019:UPC

[GML^{+}19] L. M. Gibson, A. M. Mychajliw, Y. Leon, E. Rupp, and E. A. Hadly. Using the past to contextualize anthropogenic impacts on the present and future distribution of an endemic Caribbean mammal. Conservation Biology, 33(3):500–510, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goswami:2015:MUH

[GMNO15] Varun R. Goswami, Kamal Medhi, James D. Nichols, and Madan K. Oli. Mechanistic understanding of human–wildlife conflict through a novel application of dynamic occupancy models. Conservation Biology, 29(4):1100–1110, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gissi:2018:ATC

[GMV^{+}18] Elena Gissi, Jennifer McGowan, Chiara Venier, Davide Di Carlo, Francesco Musco, Stefano Menegon, Peter Mackelworth, Tundi Agardy, and Hugh Possingham. Addressing transboundary conservation challenges through marine spatial prioritization. Conservation Biology, 32(5):1107–1117, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gilbert-Norton:2010:MAR

[GNWSB10] Lynne Gilbert-Norton, Ryan Wilson, John R. Stevens, and Karen H. Beard. A meta-analytic review of corridor effectiveness. Una revisión meta-analítica de la efectividad de los corredores. Conservation Biology, 24(3):660–668, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gonzales:2013:REC

[Gon13] Emily Gonzales. Restoration ecology comes of age. Conservation Biology, 27(2):433–435, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gore:2011:SCC

[Gor11] Meredith L. Gore. The science of conservation crime. Conservation Biology, 25(4):659–661, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gerrodette:2012:EMB

[GOR^{+}12] Tim Gerrodette, Robert Olson, Stephen Reilly, George Watters, and William Perrin. Ecological metrics of biomass removed by three methods of purse-seine fishing for tunas in the Eastern Tropical Pacific Ocean. Métricas ecológicas de la biomasa Extraída por tres Métodos de pesca de Atún con red de cerco en el Océano Pacífico Oriental Tropical. Conservation Biology, 26(2):248–256, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gurney:2015:EED

[GPB^{+}15] Georgina G. Gurney, Robert L. Pressey, Natalie C. Ban, Jorge G. Álvarez-Romero, Stacy Jupiter, and Vanessa M. Adams. Efficient and equitable design of marine protected areas in Fiji through inclusion of stakeholder-specific objectives in conservation planning. Conservation Biology, 29(5):1378–1389, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Granek:2010:ESC

[GPK^{+}10] Elise F. Granek, Stephen Polasky, Carrie V. Kappel, Denise J. Reed, David M. Stoms, Evamaria W. Koch, Chris J. Kennedy, Lori A. Cramer, Sally D. Hacker, Edward B. Barbier, Shankar Aswani, Mary Ruckelshaus, Gerardo M. E. Perillo, Brian R. Silliman, Nyawira Muthiga, David Bael, and Eric Wolanski. Ecosystem services as a common language for coastal ecosystem-based management. Los servicios del ecosistema como un lenguaje común para el manejo de costas basado en ecosistemas. Conservation Biology, 26(2):207–216, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gollan:2014:ADP

[GRA14] John R. Gollan, Daniel Ramp, and Michael B. Ashcroft. Assessing the distribution and protection status of two types of cool environment to facilitate their conservation under climate change. Conservation Biology, 28(2):456–466, April 2014. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gallagher:2015:TDB

[GRL15] R. V. Gallagher, R. P. Randall, and M. R. Leishman. Trait differences between naturalized and invasive plant species independent of residence time and phylogeny. Conservation Biology, 29(2):360–369, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grossman:2010:BBB

[Gro10] Richard Grossman. Birds and bees for biologists. Conservation Biology, 24(6):1435–1436, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Greenbaum:2018:RLH

[GRT^{+}18] Gili Greenbaum, Sharon Renan, Alan R. Templeton, Amos Bouskila, David Saltz, Daniel I. Rubenstein, and Shirli Bar-David. Revealing life-history traits by contrasting genetic estimations with predictions of effective population size. Conservation Biology, 32(4):817–827, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gruber:2017:GKF

[Gru17] Thibaud Gruber. Getting to know the family in order to save them. Conservation Biology, 31(6):1501–1503, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goodwin:2011:ETN

[GS11] Sarah E. Goodwin and W. Gregory Shriver. Effects of traffic noise on occupancy patterns of forest birds. Efectos del ruido de tráfico sobre los patrones de ocupación de aves de bosque. Conservation Biology, 26(3):406–411, April 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grant:2016:UED

[GS16] Paul B. C. Grant and Michael J. Samways. Use of ecoacoustics to determine biodiversity patterns across ecological gradients. Conservation Biology, 30(6):1320–1329, December 2016. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gavin:2010:MMI

[GSB10] Michael C. Gavin, Jennifer N. Solomon, and Sara G. Blank. Measuring and monitoring illegal use of natural resources. Medición y monitoreo del uso ilegal de recursos naturales. Conservation Biology, 24(1):89–100, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Galimberti:2016:EER

[GSB^{+}16] A. Galimberti, S. Spinelli, A. Bruno, V. Mezzasalma, F. De Mattia, P. Cortis, and M. Labra. Evaluating the efficacy of restoration plantings through DNA barcoding of frugivorous bird diets. Conservation Biology, 30(4):763–773, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gjertsen:2014:CEA

[GSDE14] Heidi Gjertsen, Dale Squires, Peter H. Dutton, and Tomoharu Eguchi. Cost-effectiveness of alternative conservation strategies with application to the Pacific leatherback turtle. Conservation Biology, 28(1):140–149, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Griffith:2019:IIS

[GSSF19] Gary P. Griffith, Peter G. Strutton, Jayson M. Semmens, and Elizabeth A. Fulton. Identifying important species that amplify or mitigate the interactive effects of human impacts on marine food webs. Conservation Biology, 33(2):403–412, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gregory:2010:ECF

[GSW10] Nathan C. Gregory, Ryan L. Sensenig, and David S. Wilcove. Effects of controlled fire and livestock grazing on bird communities in East African savannas. Efectos del fuego controlado y el pastoreo de ganado sobre comunidades de aves en sabanas de África Oriental. Conservation Biology, 24(6):1606–1616, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gaoue:2010:EHN

[GT10] Orou G. Gaoue and Tamara Ticktin. Effects of harvest of nontimber forest products and ecological differences between sites on the demography of African mahogany. Efectos de la cosecha de productos forestales no maderables y diferencias ecológicas entre sitios sobre la demografía de caoba africana. Conservation Biology, 24(2):605–614, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Galligan:2016:MAC

[GTC^{+}16] T. H. Galligan, M. A. Taggart, R. J. Cuthbert, D. Svobodova, J. Chipangura, D. Alderson, V. M. Prakash, and V. Naidoo. Metabolism of aceclofenac in cattle to vulturekilling diclofenac. Conservation Biology, 30(5):1122–1127, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gramza:2016:UPP

[GTVC16] Ashley Gramza, Tara Teel, Susan VandeWoude, and Kevin Crooks. Understanding public perceptions of risk regarding outdoor pet cats to inform conservation action. Conservation Biology, 30(2):276–286, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gilroy:2015:EST

[GUHE15] James J. Gilroy, Claudia A. Medina Uribe, Torbjørn Haugaasen, and David P. Edwards. Effect of scale on trait predictors of species responses to agriculture. Conservation Biology, 29(2):463–472, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gudde:2016:CCM

[GV16] Renske Gudde and Chris Venditti. Comparison of conservation metrics in a case study of lemurs. Conservation Biology, 30 (6):1347–1356, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guerrero:2017:USE

[GW17] Angela M. Guerrero and Kerrie A. Wilson. Using a social– ecological framework to inform the implementation of conservation plans. Conservation Biology, 31(2):290–301, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gardner:2018:DCC

[GWRW18] Charlie J. Gardner, Patrick O. Waeber, Onja H. Razafindratsima, and Lucienne Wilmé. Decision complacency and conservation planning. Conservation Biology, 32(6):1469–1472, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garcia:2010:BSS

[GZA10] Daniel Garcia, Regino Zamora, and Guillermo C. Amico. Birds as suppliers of seed dispersal in temperate ecosystems: Conservation guidelines from real-world landscapes. Aves como proveedoras de dispersión de semillas en ecosistemas templados: Directrices de conservación desde paisajes del mundo real. Conservation Biology, 24(4):1070–1079, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Godet:2011:SSC

[GZD11] Laurent Godet, Kevin A. Zelnio, and Cindy L. Van Dover. Scientists as stakeholders in conservation of hydrothermal vents. Científicos como actores en la conservación de fuentes hidrotermales. Conservation Biology, 25(2):214–222, April 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Griffiths:2013:APR

[GZJ^{+}13] Christine J. Griffiths, Nicolas Zuël, Carl G. Jones, Zairabee Ahamud, and Stephen Harris. Assessing the potential to restore historic grazing ecosystems with tortoise ecological replacements. Conservation Biology, 27(4):690–700, August 2013. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grant:2013:SMM

[GZNC13] Evan H. Campbell Grant, Elise F. Zipkin, James D. Nichols, and J. Patrick Campbell. A strategy for monitoring and managing declines in an amphibian community. Conservation Biology, 27(6):1245–1253, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gavish:2012:DFH

[GZR12] Yoni Gavish, Yaron Ziv, and Michael L. Rosenzweig. Decoupling fragmentation from habitat loss for spiders in patchy agricultural landscapes. Separando la fragmentación de la pérdida de hábitat para Arañas en paisajes Agrícolas Heterogéneos. Conservation Biology, 26(2):150–159, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garnett:2018:SLE

[GZR18] Stephen T. Garnett, Kerstin K. Zander, and Catherine J. Robinson. Social license as an emergent property of political interactions: response to Kendal and Ford 2017. Conservation Biology, 32(3):734–736, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hirsch:2011:ACT

[HAB^{+}11] Paul D. Hirsch, William M. Adams, J. Peter Brosius, Asim Zia, Nino Bariola, and Juan Luis Dammert. Acknowledging conservation trade-offs and embracing complexity. Reconocimiento los Trade-offs de la conservación y atención a la complejidad. Conservation Biology, 25(2):259–264, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haddaway:2015:CBR

[Had15] Neal R. Haddaway. A call for better reporting of conservation research data for use in meta-analyses. Conservation Biology, 29(4):1242–1245, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Huang:2017:TFH

[HAF^{+}17] Biao Huang, Joshua K. Abbott, Eli P. Fenichel, Rachata Muneepeerakul, Charles Perrings, and Leah R. Gerber. Testing the feasibility of a hypothetical whaling-conservation permit market in Norway. Conservation Biology, 31(4):809–817, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Halley:2016:PPJ

[Hal16] John M. Halley. Preaching for the planet or just spinning our wheels. Conservation Biology, 30(5):1137–1140, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Halley:2018:RHC

[Hal18] John M. Halley. Redefining the heart of conservation. Conservation Biology, 32(2):501–504, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Halley:2019:DTL

[Hal19] John M. Halley. Doubting Thomas and the love of invasive species. Conservation Biology, 33(6):1451–1453, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hannah:2010:GCS

[Han10] Lee Hannah. A global conservation system for climatechange adaptation. Un sistema de conservación global para la adaptación al cambio climático. Conservation Biology, 24(1): 70–77, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hannah:2011:CCC

[Han11] Lee Hannah. Climate change, connectivity, and conservation success. Conservation Biology, 25(6):1139–1142, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hossack:2013:RPC

[HAP^{+}13] Blake R. Hossack, Michael J. Adams, Christopher A. Pearl, Kristine W. Wilson, Evelyn L. Bull, Kristin Lohr, Debra Patla, David S. Pilliod, Jason M. Jones, Kevin K. Wheeler, Samuel P. Mckay, and Paul Stephen Corn. Roles of patch characteristics, drought frequency, and restoration in long-term trends of a widespread amphibian. Conservation Biology, 28(1):1410– 1420, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hardin:2011:CCC

[Har11] Rebecca Hardin. Competing cultures of conservation. Conservation Biology, 25(6):1098–1102, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harrington:2015:ICT

[Har15] Lauren A. Harrington. International commercial trade in live carnivores and primates 2006–2012: response to Bush et al. 2014. Conservation Biology, 29(1):293–296, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harradine:2015:IDR

[HAT^{+}15] E. L. Harradine, M. E. Andrew, J. W. Thomas, R. A. How, L. H. Schmitt, and P. B. S. Spencer. Importance of dispersal routes that minimize open-ocean movement to the genetic structure of island populations. Conservation Biology, 29(6): 1704–1714, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hedrick:2011:RBD

[HAV11] Philip W. Hedrick, Jennifer R. Adams, and John A. Vucetich. Reevaluating and broadening the definition of genetic rescue. Conservation Biology, 25(6):1069–1070, December 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heber:2010:PBI

[HB10] Sol Heber and James V. Briskie. Population bottlenecks and increased hatching failure in endangered birds. Cuellos de botella poblacionales e incremento en el fracaso de la eclosión de aves en peligro. Conservation Biology, 24(6):1674–1678, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haddad:2014:PNE

[HBD^{+}14] Nick M. Haddad, Lars A. Brudvig, Ellen I. Damschen, Daniel M. Evans, Brenda L. Johnson, Douglas J. Levey, John L. Orrock, Julian Resasco, Lauren L. Sullivan, Josh J. Tewksbury, Stephanie A. Wagner, and Aimee J. Weldon. Potential negative ecological effects of corridors. Conservation Biology, 28(5): 1178–1187, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hodgetts:2019:CG

[HBD^{+}19] Timothy Hodgetts, Dawn Burnham, Amy Dickman, Ewan A. Macdonald, and David W. Macdonald. Conservation geopolitics. Conservation Biology, 33(2):250–259, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Horvath:2010:RMA

[HBE^{+}10] Gábor Horváth, Miklós Blahó, Ádám Egri, György Kriska, István Seres, and Bruce Robertson. Reducing the maladaptive attractiveness of solar panels to polarotactic insects. Reducción de la atracción inadaptiva de placas solares para insectos polarotácticos. Conservation Biology, 24(6):1644–1653, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hicks:2012:DMH

[HBM^{+}12] John S. Hicks, Mark A. Burgman, Julian N. Marewski, Fiona Fidler, and Gerd Gigerenzer. Decision making in a human population living sustainably. Toma de decisiones en una población humana que vive sustentablemente. Conservation Biology, 26 (5):760–768, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hall:2014:ESO

[HBR^{+}14] Jaclyn M. Hall, Neil D. Burgess, Salla Rantala, Heini Vihemäki, George Jambiya, Roy E. Gereau, Fortunatus Makonda, Fadhili Njilima, Peter Sumbi, and Adam Kizaji. Ecological and social outcomes of a new protected area in Tanzania. Conservation Biology, 28(6):1512–1521, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hefley:2015:UOS

[HBTB15] Trevor J. Hefley, David M. Baasch, Andrew J. Tyre, and Erin E. Blankenship. Use of opportunistic sightings and expert knowledge to predict and compare whooping crane stopover habitat. Conservation Biology, 29(5):1337–1346, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harrison:2015:PUN

[HBTMG15] Mariel Harrison, Julia Baker, Medard Twinamatsiko, and E. J. Milner-Gulland. Profiling unauthorized natural resource users for better targeting of conservation interventions. Conservation Biology, 29(6):1636–1646, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hayward:2019:DCC

[HCA^{+}19] Matt W. Hayward, Alex Callen, Benjamin L. Allen, Guy Ballard, Femke Broekhuis, Cassandra Bugir, Rohan H. Clarke, John Clulow, Simon Clulow, Jennifer C. Daltry, Harriet T. Davies-Mostert, Peter J. S. Fleming, Andrea S. Griffin, Lachlan G. Howell, Graham I. H. Kerley, Kaya Klop-Toker, Sarah Legge, Tom Major, Ninon Meyer, Robert A. Montgomery, Katherine Moseby, Daniel M. Parker, Stéphanie Périquet, John Read, Robert J. Scanlon, Rebecca Seeto, Craig Shuttleworth, Michael J. Somers, Cottrell T. Tamessar, Katherine Tuft, Rose Upton, Marcia Valenzuela-Molina, Adrian Wayne, Ryan R. Witt, and Wolfgang Wüster. Deconstructing compassionate conservation. Conservation Biology, 33(4):760–768, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heilmann-Clausen:2015:FPC

[HCBB^{+}15] Jacob Heilmann-Clausen, Elizabeth S. Barron, Lynne Boddy, Anders Dahlberg, Gareth W. Griffith, Jenni Nordén, Otso Ovaskainen, Claudia Perini, Beatrice Senn-Irlet, and Panu Halme. A fungal perspective on conservation biology. Conservation Biology, 29(1):61–68, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hanford:2017:EVB

[HCH17] Jayne K. Hanford, Mathew S. Crowther, and Dieter F. Hochuli. Effectiveness of vegetation-based biodiversity offset metrics as surrogates for ants. Conservation Biology, 31(1):161–171, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harris:2013:AAA

[HCLW13] Richard B. Harris, Rosie Cooney, and Nigel Leader-Williams. Application of the anthropogenic Allee effect model to trophy hunting as a conservation tool. Conservation Biology, 28(4): 945–951, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hawthorne:2014:ETP

[HCM14] W. D. Hawthorne, S. Cable, and C. A. M. Marshall. Empirical trials of plant field guides. Conservation Biology, 28(3):654– 662, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hatch:2012:QLA

[HCP^{+}12] Leila T. Hatch, Christopher W. Clark, Sofie M. Van Parijs, Adam S. Frankel, and Dimitri W. Ponirakis. Quantifying loss of acoustic communication space for right whales in and around a U.S. national marine sanctuary. Conservation Biology, 26(6): 983–994, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hicks:2015:LES

[HCSM15] Christina C. Hicks, Joshua E. Cinner, Natalie Stoeckl, and Tim R. McClanahan. Linking ecosystem services and humanvalues theory. Conservation Biology, 29(5):1471–1480, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hall:2017:CRI

[HCT^{+}17] Damon M. Hall, Gerardo R. Camilo, Rebecca K. Tonietto, Jeff Ollerton, Karin Ahrné, Mike Arduser, John S. Ascher, Katherine C. R. Baldock, Robert Fowler, Gordon Frankie, Dave Goulson, Bengt Gunnarsson, Mick E. Hanley, Janet I. Jackson, Gail Langellotto, David Lowenstein, Emily S. Minor, Stacy M. Philpott, Simon G. Potts, Muzafar H. Sirohi, Edward M. Spevak, Graham N. Stone, and Caragh G. Threlfall. The city as a refuge for insect pollinators. Conservation Biology, 31(1):24– 29, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hill:2014:OPR

[HD14] Jason M. Hill and Duane R. Diefenbach. Occupancy patterns of regionally declining grassland sparrow populations in a forested Pennsylvania landscape. Conservation Biology, 28(3):735–744, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hennenberg:2010:PBR

[HDH^{+}10] K. J. Hennenberg, C. Dragisic, S. Haye, J. Hewson, B. Semroc, C. Savy, K. Wiegmann, H. Fehrenbach, and U. R. Fritsche. The power of bioenergy-related standards to protect biodiversity. El poder de las normas para la Protección de la naturaleza relacionadas con la bioenergía. Conservation Biology, 24 (2):412–423, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hoffmann:2015:DCM

[HDH^{+}15] Michael Hoffmann, J. W. Duckworth, Katharine Holmes, David P. Mallon, Ana S. L. Rodrigues, and Simon N. Stuart. The difference conservation makes to extinction risk of the world’s ungulates. Conservation Biology, 29(5):1303–1313, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:2010:CAC

[HDHL10] Malcolm Hunter, Jr., Eric Dinerstein, Jon Hoekstra, and David Lindenmayer. A call to action for conserving biological diversity in the face of climate change. Conservation Biology, 24(5): 1169–1171, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hazzah:2014:CCC

[HDN^{+}14a] Leela Hazzah, Stephanie Dolrenry, Lisa Naughton, Charles T. T. Edwards, Ogeto Mwebi, Fiachra Kearney, and Laurence Frank. Cover caption: Conservation Biology. Conservation Biology, 28(3):i, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hazzah:2014:ETL

[HDN^{+}14b] Leela Hazzah, Stephanie Dolrenry, Lisa Naughton, Charles T. T. Edwards, Ogeto Mwebi, Fiachra Kearney, and Laurence Frank. Efficacy of two lion conservation programs in Maasailand, Kenya. Conservation Biology, 28(3):851–860, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heberlein:2012:NEA

[Heb12] Thomas A. Heberlein. Navigating environmental attitudes. Conservation Biology, 26(4):583–585, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Helfrich:2014:RBP

[Hel14] Joel T. Helfrich. Remembering biogladiator Peter Warshall, 1943–2013. Conservation Biology, 28(5):883–885, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hendrickson:2016:NEB

[Hen16] Ole Hendrickson. Nuclear energy and biodiversity conservation: response to Brook and Bradshaw 2015. Conservation Biology, 31(1):661–662, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hall:2010:DMT

[HF10] John A. Hall and Erica Fleishman. Demonstration as a means to translate conservation science into practice. La demostración como un medio para traducir ciencia de la conservación en práctica. Conservation Biology, 24(1):120–127, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hardy:2018:FIP

[HFBG18] Mathew J. Hardy, James A. Fitzsimons, Sarah A. Bekessy, and Ascelin Gordon. Factors influencing property selection for conservation revolving funds. Conservation Biology, 32(2):276– 286, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hays:2010:BPM

[HFK^{+}10] Graeme C. Hays, Sabrina Fossette, Kostas A. Katselidis, Gail Schofield, and Mike B. Gravenor. Breeding periodicity for male sea turtles, operational sex ratios, and implications in the face of climate change. Periodicidad reproductiva de tortugas marinas machos, proporciones de sexos operacionales e implicaciones frente al cambio climático. Conservation Biology, 24 (6):1636–1643, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hardwick:2011:RBG

[HFL^{+}11] Kate A. Hardwick, Peggy Fiedler, Lyndon C. Lee, Bruce Pavlik, Richard J. Hobbs, James Aronson, Martin Bidartondo, Eric Black, David Coates, Matthew I. Daws, Kingsley Dixon, Stephen Elliott, Kern Ewing, George Gann, David Gibbons, Joachim Gratzfeld, Martin Hamilton, David Hardman, Jim Harris, Pat M. Holmes, Meirion Jones, David Mabberley, Andrew Mackenzie, Carlos Magdalena, Robert Marrs, William Milliken, Anthony Mills, Eimear Nic Lughadha, Margaret Ramsay, Paul Smith, Nigel Taylor, Clare Trivedi, Michael Way, Oliver Whaley, and Stephen D. Hopper. The role of botanic gardens in the science and practice of ecological restoration. El papel de los jardines Botánicos en la ciencia y Práctica de la restauración ecológica. Conservation Biology, 25(2):265– 275, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Habib:2013:EEO

[HFSB13] Thomas J. Habib, Daniel R. Farr, Richard R. Schneider, and Stan Boutin. Economic and ecological outcomes of flexible biodiversity offset systems. Conservation Biology, 27(6):1313– 1323, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Herrera:2010:EFF

[HG10] José Manuel Herrera and Daniel García. Effects of forest fragmentation on seed dispersal and seedling establishment in ornithochorous trees. Efectos de la fragmentación del bosque sobre la dispersión de semillas y establecimiento de plántulas en plantas ornitócoras. Conservation Biology, 24(4):1089–1098, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:2013:EGG

[HGCT13] Elizabeth A. Hunter, James P. Gibbs, Linda J. Cayot, and Washington Tapia. Equivalency of Galápagos giant tortoises used as ecological replacement species to restore ecosystem functions. Conservation Biology, 27(4):701–709, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hjort:2015:WGM

[HGGH15] Jan Hjort, John E. Gordon, Murray Gray, and Malcolm L. Hunter, Jr. Why geodiversity matters in valuing nature’s stage. Conservation Biology, 29(3):630–639, June 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harris:2019:CIT

[HGM19] Lindsey Harris, Meredith Gore, and Morena Mills. Compliance with ivory trade regulations in the United Kingdom among traders. Conservation Biology, 33(4):906–916, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harding:2016:DAC

[HGP16] Gemma Harding, Richard A. Griffiths, and Lissette Pavajeau. Developments in amphibian captive breeding and reintroduction programs. Conservation Biology, 30(2):340–349, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Henle:2016:PNE

[HGRS16] Klaus Henle, Erik Gawel, Irene Ring, and Sebastian Strunz. Promoting nuclear energy to sustain biodiversity conservation in the face of climate change: response to Brook and Bradshaw 2015. Conservation Biology, 31(1):663–665, June 2016.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heller:2014:DNP

[HH14] Nicole E. Heller and Richard J. Hobbs. Development of a natural practice to adapt conservation goals to global change. Conservation Biology, 28(3):696–704, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [FS15].

Hampton:2019:UAW

[HH19a] Jordan O. Hampton and Timothy H. Hyndman. Underaddressed animal-welfare issues in conservation. Conservation Biology, 33(4):803–811, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hodgson:2019:ICE

[HH19b] Emma E. Hodgson and Benjamin S. Halpern. Investigating cumulative effects across ecological scales. Conservation Biology, 33(2):22–32, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hoeppner:2019:CRR

[HH19c] Johanne Malin Hoeppner and Lesley Hughes. Climate readiness of recovery plans for threatened Australian species. Conservation Biology, 33(3):534–542, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hole:2011:TMF

[HHA^{+}11] David G. Hole, Brian Huntley, Julius Arinaitwe, Stuart H. M. Butchart, Yvonne C. Collingham, Lincoln D. C. Fishpool, Deborah J. Pain, and Stephen G. Willis. Toward a management framework for networks of protected areas in the face of climate change. Hacia un marco de manejo para redes de Áreas protegidas ante el cambio climático. Conservation Biology, 26 (3):305–315, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hodge:2015:AAN

[HHB15] Ian Hodge, Jennifer Hauck, and Aletta Bonn. The alignment of agricultural and nature conservation policies in the European union. Conservation Biology, 29(4):996–1005, August 2015.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hansen:2010:DCS

[HHDM10] Lara Hansen, Jennifer Hoffman, Carlos Drews, and Eric Mielbrecht. Designing climate-smart conservation: Guidance and case studies. Diseño de conservación considerando el clima: Directrices y estudios de caso. Conservation Biology, 24(1):63– 69, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hickisch:2019:EPB

[HHJ^{+}19] R. Hickisch, T. Hodgetts, P. J. Johnson, C. Sillero-Zubiri, K. Tockner, and D. W. Macdonald. Effects of publication bias on conservation planning. Conservation Biology, 33(5): 1151–1163, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Humair:2015:CTI

[HHKK15] Franziska Humair, Luc Humair, Fabian Kuhn, and Christoph Kueffer. E-commerce trade in invasive plants. Conservation Biology, 29(6):1658–1665, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holmes:2016:PBO

[HHW^{+}16] N. D. Holmes, G. R. Howald, A. S. Wegmann, C. J. Donlan, M. Finkelstein, and B. Keitt. The potential for biodiversity offsetting to fund invasive species eradications on islands. Conservation Biology, 30(2):425–427, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hughes:2013:WPC

[HHY13] Terry P. Hughes, Hui Huang, and Matthew A. L. Young. The wicked problem of China’s disappearing coral reefs. Conservation Biology, 27(2):261–269, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hickey:2010:QMC

[Hic10] Valerie Hickey. The quagga mussel crisis at Lake Mead National Recreation Area, Nevada (U.S.A.). La crisis de mejillones quagga en el Área de Nacional de Recreación Lago Mead, Nevada (EUA). Conservation Biology, 24(4):931–937, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hughes:2016:AES

[HICF16] Kevin A. Hughes, Louise C. Ireland, Peter Convey, and Andrew H. Fleming. Assessing the effectiveness of specially protected areas for conservation of Antarctica’s botanical diversity. Conservation Biology, 30(1):113–120, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

He:2016:RGT

[HJH16] Xiaoping He, Mattias L. Johansson, and Daniel D. Heath. Role of genomics and transcriptomics in selection of reintroduction source populations. Conservation Biology, 30(5):1010–1018, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hobbs:2011:ERE

[HJM11] J.-P. A. Hobbs, G. P. Jones, and P. L. Munday. Extinction risk in endemic marine fishes. Conservation Biology, 27(3): 1053–1055, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hollings:2014:TCF

[HJMM14] Tracey Hollings, Menna Jones, Nick Mooney, and Hamish Mccallum. Trophic cascades following the disease-induced decline of an apex predator, the Tasmanian Devil. Conservation Biology, 28(1):63–75, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Huxley-Jones:2012:UDB

[HJSF^{+}12] Elizabeth Huxley-Jones, Jennifer L. A. Shaw, Carly Fletcher, Juliette Parnell, and Phillip C. Watts. Use of DNA barcoding to reveal species composition of convenience seafood. El uso de Código de barras de ADN para identificar la Composición de especies de mariscos de Preparación rápida. Conservation Biology, 26(2):367–371, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heeren:2017:EJU

[HKB^{+}17] Alexander Heeren, Gabriel Karns, Jeremy Bruskotter, Eric Toman, Robyn Wilson, and Harmony Szarek. Expert judgment and uncertainty regarding the protection of imperiled species. Conservation Biology, 31(3):657–665, June 2017. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heagney:2015:SEB

[HKFC15] E. C. Heagney, M. Kovac, J. Fountain, and N. Conner. Socioeconomic benefits from protected areas in southeastern Australia. Conservation Biology, 29(6):1647–1657, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hamede:2012:RET

[HLB^{+}12] Rodrigo Hamede, Shelly Lachish, Katherine Belov, Gregory Woods, Alexandre Kreiss, Anne-Maree Pearse, Billie Lazenby, Menna Jones, and Hamish Mccallum. Reduced effect of Tasmanian Devil facial tumor disease at the disease front. Reducción del efecto de la enfermedad de tumores faciales del demonio de Tasmania en el frente de la enfermedad. Conservation Biology, 26(2):124–134, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hossack:2013:RIT

[HLC13] Blake R. Hossack, Winsor H. Lowe, and Paul Stephen Corn. Rapid increases and time-lagged declines in amphibian occupancy after wildfire. Conservation Biology, 27(1):219–228, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hewitt:2019:ASM

[HLE19] Judi E. Hewitt, Carolyn J. Lundquist, and Joanne Ellis. Assessing sensitivities of marine areas to stressors based on biological traits. Conservation Biology, 33(2):142–151, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hinsley:2016:EES

[HLHR16] Amy Hinsley, Tamsin E. Lee, Joseph R. Harrison, and David L. Roberts. Estimating the extent and structure of trade in horticultural orchids via social media. Conservation Biology, 30(5): 1038–1047, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heinrichs:2015:DSC

[HLS^{+}15] Julie A. Heinrichs, Joshua J. Lawler, Nathan H. Schumaker, Chad B. Wilsey, and Darren J. Bender. Divergence in sink contributions to population persistence. Conservation Biology, 29(6):1674–1683, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heinrichs:2018:MTS

[HLS^{+}18] Julie A. Heinrichs, Joshua J. Lawler, Nathan H. Schumaker, Chad B. Wilsey, Kira C. Monroe, and Cameron L. Aldridge. A multispecies test of source–sink indicators to prioritize habitat for declining populations. Conservation Biology, 32(3):648–659, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hallam:2011:MPW

[HM11] Thomas G. Hallam and Gary F. Mccracken. Management of the panzootic white-nose syndrome through culling of bats. Manejo del síndrome de nariz blanca Panzoótico mediante selección de murciélagos. Conservation Biology, 25(1):189–194, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hartway:2012:MAE

[HM12] Cynthia Hartway and L. Scott Mills. A meta-analysis of the effects of common management actions on the nest success of North American birds. Un meta–análisis de los efectos de acciones de manejo comunes sobre el Éxito de anidación de aves de Norte América. Conservation Biology, 26(4):657–666, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hamilton:2016:AIR

[HM16] Jill A. Hamilton and Joshua M. Miller. Adaptive introgression as a resource for management and genetic conservation in a changing climate. Conservation Biology, 30(1):33–41, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hoop:2013:AMM

[HMB^{+}13] Julie M. Van Der Hoop, Michael J. Moore, Susan G. Barco, Timothy V. N. Cole, Pierre-Yves Daoust, Allison G. Henry, Donald F. Mcalpine, William A. Mclellan, Tonya Wimmer, and Andrew R. Solow. Assessment of management to mitigate anthropogenic effects on large whales. Conservation Biology, 27 (1):121–133, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hopcraft:2015:CEB

[HMB^{+}15] J. Grant C. Hopcraft, Simon A. R. Mduma, Markus Borner, Gerald Bigurube, Alain Kijazi, Daniel T. Haydon, William Wakilema, Dennis Rentsch, A. R. E. Sinclair, Andrew Dobson, and James Daudi Lembeli. Conservation and economic benefits of a road around the Serengeti. Conservation Biology, 29(3):932–936, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Herbert:2010:TRN

[HMD^{+}10] Matthew E. Herbert, Peter B. Mcintyre, Patrick J. Doran, J. David Allan, and Robin Abell. Terrestrial reserve networks do not adequately represent aquatic ecosystems. Las redes de reservas terrestres no representan a los ecosistemas acuáticos adecuadamente. Conservation Biology, 24(4):1002–1011, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harrington:2013:CCE

[HMG^{+}13] Lauren A. Harrington, Axel Moehrenschlager, Merryl Gelling, Rob P. D. Atkinson, Joelene Hughes, and David W. Macdonald. Conflicting and complementary ethics of animal welfare considerations in reintroductions. Conservation Biology, 27 (3):486–500, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hays:2014:ULD

[HMIE14] Graeme C. Hays, Jeanne A. Mortimer, Daniel Ierodiaconou, and Nicole Esteban. Use of long-distance migration patterns of an endangered species to inform conservation planning for the world’s largest Marine Protected Area. Conservation Biology, 28(6):1636–1644, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Howes:2010:BNA

[HMM10] Alison L. Howes, Martine Maron, and Clive A. Mcalpine. Bayesian networks and adaptive management of wildlife habitat. Redes Bayesianas y manejo adaptativo del hábitat de vida silvestre. Conservation Biology, 24(4):974–983, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:2017:UEH

[HN17] Elizabeth A. Hunter and Nathan P. Nibbelink. Using environmental heterogeneity to plan for sea-level rise. Conservation Biology, 31(6):1409–1417, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hylander:2013:ECM

[HNDE13] Kristoffer Hylander, Sileshi Nemomissa, Josefien Delrue, and Woldeyohannes Enkosa. Effects of coffee management on deforestation rates and forest integrity. Conservation Biology, 28 (4):1031–1040, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harris:2018:MCB

[HNO^{+}18] Linda R. Harris, Ronel Nel, Herman Oosthuizen, Mike Meÿer, Deon Kotze, Darrell Anders, Steven McCue, and Santosh Bachoo. Managing conflicts between economic activities and threatened migratory marine species toward creating a multiobjective blue economy. Conservation Biology, 32(2):411–423, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hare:2011:UEE

[HNS^{+}11] Matthew P. Hare, Leonard Nunney, Michael K. Schwartz, Daniel E. Ruzzante, Martha Burford, Robin S. Waples, Kristen Ruegg, and Friso Palstra. Understanding and estimating effective population size for practical application in marine species management. Entendiendo y estimando el tamaño poblacional efectivo para aplicación práctica en el manejo de especies marinas. Conservation Biology, 25(3):438–449, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hajjar:2016:DCC

[HOC^{+}16] Reem Hajjar, Johan A. Oldekop, Peter Cronkleton, Emily Etue, Peter Newton, Aaron J. M. Russel, Januarti Sinarra Tjajadi, Wen Zhou, and Arun Agrawal. The data not collected on community forestry. Conservation Biology, 30(6):1357–1362, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hanson:2010:MCR

[HODB10] Chad T. Hanson, Dennis C. Odion, Dominick A. Dellasala, and William L. Baker. More-comprehensive recovery actions for northern spotted owls in dry forests: Reply to Spies et al. Conservation Biology, 26(2):334–337, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [SMB^{+}10].

Hellmann:2011:FHI

[HP11] J. J. Hellmann and M. E. Pfrender. Future human intervention in ecosystems and the critical role for evolutionary biology. Conservation Biology, 25(6):1143–1147, December 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holmes:2014:ARI

[HP14] Ignacia Holmes and Catherine Potvin. Avoiding re-inventing the wheel in a people-centered approach to REDD+. Conservation Biology, 28(6):1380–1393, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Horton:2016:CAC

[HPBP16] Cristi C. Horton, Tarla Rai Peterson, Paulami Banerjee, and Markus J. Peterson. Credibility and advocacy in conservation science. Conservation Biology, 30(1):23–32, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haywood:2016:PBD

[HPD16] Benjamin K. Haywood, Julia K. Parrish, and Jane Dolliver. Place-based and data-rich citizen science as a precursor for conservation action. Conservation Biology, 31(1):476–486, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hartman:2014:FMC

[HPL14] Rosemary Hartman, Karen Pope, and Sharon Lawler. Factors mediating co-occurrence of an economically valuable introduced fish and its native frog prey. Conservation Biology, 28(3):763–772, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Honey-Roses:2011:SEE

[HRBR11] Jordi Honey-Rosés, Kathy Baylis, and M. Isabel Ramírez. A spatially explicit estimate of avoided forest loss. Una estimación espacialmente explícita de la pérdida de bosque evitada. Conservation Biology, 27(3):1032–1043, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hase:2010:EPL

[HRC10] Amrei Von Hase, Mathieu Rouget, and Richard M. Cowling. Evaluating private land conservation in the Cape Lowlands, South Africa. Evaluación de la planificación e implementación de acuerdos de conservación con terratenientes privados en Cape Lowlands, Sudáfrica. Conservation Biology, 24 (5):1182–1189, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harrison:2016:AEN

[HRHC16] Joseph R. Harrison, David L. Roberts, and Julio Hernandez-Castro. Assessing the extent and nature of wildlife trade on the dark web. Conservation Biology, 30(4):900–904, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hill:2012:IBS

[HRKMG12] Nicholas A. O. Hill, J. Marcus Rowcliffe, Heather J. Koldewey, and E. J. Milner-Gulland. The interaction between seaweed farming as an alternative occupation and fisher numbers in the Central Philippines. Interacción entre el cultivo de algas marinas como una ocupación alternativa y el Número de pescadores en Filipinas Central. Conservation Biology, 26(2):324–334, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:2014:CNA

[HRL14] Malcolm L. Hunter, Jr., Kent H. Redford, and David B. Lindenmayer. The complementary niches of anthropocentric and biocentric conservationists. Conservation Biology, 28(3):641– 645, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hatfield:2012:PDH

[HRSK12] Jeff S. Hatfield, Michelle H. Reynolds, Nathaniel E. Seavy, and Crystal M. Krause. Population dynamics of Hawaiian seabird colonies vulnerable to sea-level rise. Dinámica poblacional de colonias de aves marinas vulnerables al incremento del nivel del mar en Hawaii. Conservation Biology, 26(4):667–678, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hatfield:2013:CCC

[HRSK13] Jeff S. Hatfield, Michelle H. Reynolds, Nathaniel E. Seavy, and Crystal M. Krause. Cover caption: Conservation Biology. Conservation Biology, 26(4):i, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heupel:2010:SSN

[HS10] M. R. Heupel and C. A. Simpfendorfer. Science or slaughter: Need for lethal sampling of sharks ciencia o matanza: la necesidad de muestreo letal de tiburones. Conservation Biology, 24(5):1212–1218, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hawlena:2010:ETD

[HSAB10] Dror Hawlena, David Saltz, Zvika Abramsky, and Amos Bouskila. Ecological trap for desert lizards caused by anthropogenic changes in habitat structure that favor predator activity. Creación de una trampa ecológica para lagartijas de desierto con la adición de una estructura de hábitat que favorece la actividad de depredadores. Conservation Biology, 25(6):803– 809, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holmes:2012:CGC

[HSB12] George Holmes, Katherine Scholfield, and Dan Brockington. A comparison of global conservation prioritization models with spatial spending patterns of conservation nongovernmental organizations. Una comparación de modelos de priorización de conservación global con patrones espaciales de gastos de organizaciones de conservación no gubernamentales. Conservation Biology, 26(4):602–609, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harrison:2016:IHT

[HSB^{+}16] Rhett D. Harrison, Rachakonda Sreekar, Jedediah F. Brodie, Sarah Brook, Matthew Luskin, Hannah O’Kelly, Madhu Rao, Brett Scheffers, and Nandini Velho. Impacts of hunting on tropical forests in Southeast Asia. Conservation Biology, 30 (5):972–981, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holsman:2012:IET

[HSBE12] Kirstin K. Holsman, Mark D. Scheuerell, Eric Buhle, and Robert Emmett. Interacting effects of translocation, artificial propagation, and environmental conditions on the marine survival of Chinook salmon from the Columbia River, Washington, U.S.A. Efectos interactuantes de la translocación, Propagación artificial y condiciones ambientales sobre la supervivencia marina de salmón Chinook del Río Columbia, Washington, E.U.A. Conservation Biology, 26(5):912–922, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hilborn:2012:DTO

[HSBJ12] Ray Hilborn, Ian J. Stewart, Trevor A. Branch, and Olaf P. Jensen. Defining trade-offs among conservation, profitability, and food security in the California Current bottom-trawl fishery. Definición de compensaciones entre conservación, rentabilidad y seguridad alimentaria en la Pesquería de arrastre de fondo en la Corriente de California. Conservation Biology, 26 (2):257–268, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holmes:2017:UCP

[HSF17] George Holmes, Chris Sandbrook, and Janet A. Fisher. Understanding conservationists’ perspectives on the new-conservation debate. Conservation Biology, 31(3):353–363, April 2017. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heard:2013:TDI

[HSR^{+}13] Matthew J. Heard, Katherine F. Smith, Kelsey J. Ripp, Melanie Berger, Jane Chen, Justin Dittmeier, Maggie Goter, Stephen T. Mcgarvey, and Elizabeth Ryan. The threat of disease increases as species move toward extinction. Conservation Biology, 28(1):1378–1388, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haenn:2014:ICO

[HSRC14] Nora Haenn, Birgit Schmook, Yol Reyes, and Sophie Calmé. Improving conservation outcomes with insights from local experts and bureaucracies. Conservation Biology, 28(4):951–958, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Habel:2016:BCS

[HSU^{+}16] Jan Christian Habel, Andreas Segerer, Werner Ulrich, Olena Torchyk, Wolfgang W. Weisser, and Thomas Schmitt. Butterfly community shifts over two centuries. Conservation Biology, 30 (4):754–762, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hilbers:2017:SPT

[HSV^{+}17] Jelle P. Hilbers, Luca Santini, Piero Visconti, Aafke M. Schipper, Cecilia Pinto, Carlo Rondinini, and Mark A. J. Huijbregts. Setting population targets for mammals using body mass as a predictor of population persistence. Conservation Biology, 31 (3):385–393, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hartel:2019:MHL

[HSV^{+}19] Tibor Hartel, Ben C. Scheele, Abi Tamim Vanak, Laurenţiu Rozylowicz, John D. C. Linnell, and Euan G. Ritchie. Mainstreaming human and large carnivore coexistence through institutional collaboration. Conservation Biology, 33(6):1256–1265, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harer:2017:IFF

[HTDM17] Andreas Härer, Julián Torres-Dowdall, and Axel Meyer. The imperiled fish fauna in the Nicaragua Canal zone. Conservation Biology, 31(1):86–95, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hewitt:2016:RTS

[HTE16] Judi E. Hewitt, Simon F. Thrush, and Kari E. Ellingsen. The role of time and species identities in spatial patterns of species richness and conservation. Conservation Biology, 30(5): 1080–1088, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Halme:2010:FMA

[HTH^{+}10] Panu Halme, Tero Toivanen, Merja Honkanen, Janne S. Kotiaho, Mikko Mönkkönen, and Jonna Timonen. Flawed metaanalysis of biodiversity effects of forest management. Conservation Biology, 24(4):1154–1156, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harris:2017:MIP

[HTH^{+}17a] J. Berton C. Harris, Morgan W. Tingley, Fangyuan Hua, Ding Li Yong, J. Marion Adeney, Tien Ming Lee, William Marthy, Dewi M. Prawiradilaga, Cagan H. Sekercioglu, Suyadi, Nurul Winarni, and David S. Wilcove. Measuring the impact of the pet trade on Indonesian birds. Conservation Biology, 31 (3):394–405, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hockendorff:2017:CFR

[HTH^{+}17b] Stefanie Höckendorff, Jonathan D. Tonkin, Peter Haase, Margret Bunzel-Drüke, Olaf Zimball, Matthias Scharf, and Stefan Stoll. Characterizing fish responses to a river restoration over 21 years based on species’ traits. Conservation Biology, 31(5): 1098–1108, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hughes:2019:UME

[Hug19] Alice C. Hughes. Understanding and minimizing environmental impacts of the belt and road initiative. Conservation Biology, 33(4):883–894, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hulme:2015:RWB

[Hul15] Philip E. Hulme. Resolving whether botanic gardens are on the road to conservation or a pathway for plant invasions. Conservation Biology, 29(3):816–824, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hurlbert:2013:CNM

[Hur13] Stuart H. Hurlbert. Critical need for modification of U.S. population policy. Conservation Biology, 27(4):887–889, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haddaway:2019:PTN

[HW19] Neal R. Haddaway and Martin J. Westgate. Predicting the time needed for environmental systematic reviews and systematic maps. Conservation Biology, 33(2):434–443, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hock:2016:CRE

[HWB^{+}16] Karlo Hock, Nicholas H. Wolff, Roger Beeden, Jessica Hoey, Scott A. Condie, Kenneth R. N. Anthony, Hugh P. Possingham, and Peter J. Mumby. Controlling range expansion in habitat networks by adaptively targeting source populations. Conservation Biology, 30(4):856–866, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hebblewhite:2010:REN

[HWM10] M. Hebblewhite, C. White, and M. Musiani. Revisiting extinction in national parks: Mountain caribou in Banff. Conservation Biology, 26(2):341–344, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haddaway:2015:MLR

[HWMC15] N. R. Haddaway, P. Woodcock, B. Macura, and A. Collins. Making literature reviews more reliable through application of lessons from systematic reviews. Conservation Biology, 29(6): 1596–1605, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hampton:2019:CVC

[HWS19] Jordan O. Hampton, Bruce Warburton, and Peter Sandøe. Compassionate versus consequentialist conservation. Conservation Biology, 33(4):751–759, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holdaway:2012:SAN

[HWW12] Robert J. Holdaway, Susan K. Wiser, and Peter A. Williams. Status assessment of New Zealand’s naturally uncommon ecosystems. Evaluación del estatus de ecosistemas naturalmente poco comunes en Nueva Zelanda. Conservation Biology, 26(4):619–629, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Iacona:2016:LOG

[IBA16] Gwenllian D. Iacona, Michael Bode, and Paul R. Armsworth. Limitations of outsourcing on-the-ground biodiversity conservation. Conservation Biology, 30(6):1245–1254, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ibisch:2017:WWU

[Ibi17] Pierre L. Ibisch. What works in an uncertain and complex world. Conservation Biology, 31(5):1211–1213, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ibisch:2018:FFF

[Ibi18] Pierre L. Ibisch. Food for forest fights. Conservation Biology, 32(6):1479–1483, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Irga:2018:CMA

[IBT18] Peter J. Irga, Katherine Barker, and Fraser R. Torpy. Conservation mycology in Australia and the potential role of citizen science. Conservation Biology, 32(5):1031–1037, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ives:2017:SSC

[IF17] Christopher D. Ives and Joern Fischer. The self-sabotage of conservation: reply to Manfredo et al. Conservation Biology, 31(6):1483–1485, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Iwamura:2014:OMM

[IFP14] Takuya Iwamura, Richard A. Fuller, and Hugh P. Possingham. Optimal management of a multispecies shorebird flyway under sea-level rise. Conservation Biology, 28(6):1710–1720, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ilg:2017:EAB

[IO17] Christiane Ilg and Beat Oertli. Effectiveness of amphibians as biodiversity surrogates in pond conservation. Conservation Biology, 31(3):437–445, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ingwell:2011:UCS

[IP11] Laura L. Ingwell and Evan L. Preisser. Using citizen science programs to identify host resistance in pest-invaded forests. Utilización de programas científicos ciudadanos para identificar la resistencia de hospederos en bosques invadidos por plagas. Conservation Biology, 25(1):182–188, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Iftekhar:2017:HEC

[IPA^{+}17] Md Sayed Iftekhar, Maksym Polyakov, Dean Ansell, Fiona Gibson, and Geoffrey M. Kay. How economics can further the success of ecological restoration. Conservation Biology, 31(2):261– 268, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ishiyama:2018:PEI

[IRK^{+}18] Nobuo Ishiyama, Masahiro Ryo, Taiga Kataoka, Shigeya Nagayama, Masanao Sueyoshi, Akira Terui, Terutaka Mori, Takumi Akasaka, and Futoshi Nakamura. Predicting the ecological impacts of large-dam removals on a river network based on habitat-network structure and flow regimes. Conservation Biology, 32(6):1403–1413, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ingersoll:2016:EWN

[ISA16] Thomas E. Ingersoll, Brent J. Sewall, and Sybill K. Amelon. Effects of white-nose syndrome on regional population patterns of 3 hibernating bat species. Conservation Biology, 30 (5):1048–1059, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Iacona:2018:SRC

[ISM^{+}18] Gwenllian D. Iacona, William J. Sutherland, Bonnie Mappin, Vanessa M. Adams, Paul R. Armsworth, Tim Coleshaw, Carly Cook, Ian Craigie, Lynn V. Dicks, James A. Fitzsimons, Jennifer McGowan, Andrew J. Plumptre, Tal Polak, Andrew S. Pullin, Jeremy Ringma, Ian Rushworth, Andrea Santangeli, Annette Stewart, Ayesha Tulloch, Jessica C. Walsh, and Hugh P. Possingham. Standardized reporting of the costs of management interventions for biodiversity conservation. Conservation Biology, 32(5):979–988, October 2018. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ikin:2016:ECN

[ITG^{+}16] Karen Ikin, Ayesha Tulloch, Philip Gibbons, Dean Ansell, Julian Seddon, and David Lindenmayer. Evaluating complementary networks of restoration plantings for landscape-scale occurrence of temporally dynamic species. Conservation Biology, 30(5):1027–1037, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Iwata:2010:EDB

[IUM10] Tomoya Iwata, Jotaro Urabe, and Hiromune Mitsuhashi. Effects of drainage-basin geomorphology on insectivorous bird abundance in temperate forests. Efectos de la geomorfología de cuencas sobre la abundancia de aves insectívoras en bosques templados. Conservation Biology, 24(5):1278–1289, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jacquet:2013:AE

[Jac13] Jennifer Jacquet. The anthropocebo effect. Conservation Biology, 28(4):898–899, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jamieson:2011:FEI

[Jam11] Ian G. Jamieson. Founder effects, inbreeding, and loss of genetic diversity in four avian reintroduction programs. Efecto fundador, endogamia y pérdida de diversidad genética en cuatro programas de reintroducció de aves. Conservation Biology, 25(1):115–123, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jeschke:2014:DIN

[JBB^{+}14] Jonathan M. Jeschke, Sven Bacher, Tim M. Blackburn, Jaimie T. A. Dick, Franz Essl, Thomas Evans, Mirijam Gaertner, Philip E. Hulme, Ingolf Kühn, Agata Mrugała, Jan Pergl, Petr Pyšek, Wolfgang Rabitsch, Anthony Ricciardi, David M. Richardson, Agnieszka Sendek, Montserrat Vilà, Marten Winter, and Sabrina Kumschick. Defining the impact of non-native species. Conservation Biology, 28(5):1188–1194, October 2014.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jantz:2015:FHL

[JBB^{+}15] Samuel M. Jantz, Brian Barker, Thomas M. Brooks, Louise P. Chini, Qiongyu Huang, Rachel M. Moore, Jacob Noel, and George C. Hurtt. Future habitat loss and extinctions driven by land-use change in biodiversity hotspots under four scenarios of climate-change mitigation. Conservation Biology, 29(4):1122– 1131, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Joppa:2016:IAM

[JBH^{+}16] Lucas N. Joppa, Stuart H. M. Butchart, Michael Hoffmann, Steve P. Bachman, H. Resit Akçakaya, Justin F. Moat, Monika Böhm, Robert A. Holland, Adrian Newton, Beth Polidoro, and Adrian Hughes. Impact of alternative metrics on estimates of extent of occurrence for extinction risk assessment. Conservation Biology, 30(2):362–370, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Janssen:2018:BPU

[JC18] Jordi Janssen and Serene C. L. Chng. Biological parameters used in setting captive-breeding quotas for Indonesia’s breeding facilities. Conservation Biology, 32(1):18–25, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See reply [Nat18].

Jones:2011:WWH

[JCA^{+}11] Julia P. G. Jones, Ben Collen, Giles Atkinson, Peter W. J. Baxter, Philip Bubb, Janine B. Illian, Todd E. Katzner, Aidan Keane, Jonathan Loh, Eve Mcdonald-Madden, Emily Nicholson, Henrique M. Pereira, Hugh P. Possingham, Andrew S. Pullin, Ana S. L. Rodrigues, Viviana Ruiz-Gutierrez, Matthew Sommerville, and E. J. Milner-Gulland. The why, what, and how of global biodiversity indicators beyond the 2010 target. El porqué, qué y Cómo de los indicadores globales de biodiversidad más allá de la Meta 2010. Conservation Biology, 25 (3):450–457, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jain:2014:MPC

[JCCC14] Anuj Jain, Kwek Yan Chong, Marcus Aik Hwee Chua, and Gopalasamy Reuben Clements. Moving away from paper corridors in Southeast Asia. Conservation Biology, 28(4):889–891, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jacquet:2016:VBC

[JD16] Jennifer Jacquet and Nicolas Delon. The values behind calculating the value of trophy hunting. Conservation Biology, 30 (4):910–911, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jaffe:2010:EDH

[JDA^{+}10] Rodolfo Jaffé, Vincent Dietemann, Mike H. Allsopp, Cecilia Costa, Robin M. Crewe, Raffaele Dall’Olio, Pilar De La Rúa, Mogbel A. A. El-Niweiri, Ingemar Fries, Nikola Kezic, Michael S. Meusel, Robert J. Paxton, Taher Shaibi, Eckart Stolle, and Robin F. A. Moritz. Estimating the density of honeybee colonies across their natural range to fill the gap in pollinator decline censuses. Estimación de la densidad de colonias de abejas en su rango natural para cerrar la brecha en los censos del declive de polinizadores. Conservation Biology, 24(2):583– 593, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jensen:2014:ECC

[Jen14] Eric Jensen. Evaluating Children’s conservation biology learning at the zoo. Conservation Biology, 28(4):1004–1011, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jewell:2013:EMT

[Jew13] Zoe Jewell. Effect of monitoring technique on quality of conservation science. Conservation Biology, 27(3):501–508, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jordan:2011:KGB

[JGH^{+}11] Rebecca C. Jordan, Steven A. Gray, David V. Howe, Wesley R. Brooks, and Joan G. Ehrenfeld. Knowledge gain and behavioral change in citizen–science programs. Obtención de conocimiento y cambio conductual en los programas de ciencia–ciudadana. Conservation Biology, 25(6):1148–1154, December 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jimenez:2016:MMB

[JGL^{+}16] José Jiménez, Emilio J. García, Luis Llaneza, Vicente Palacios, Luis Mariano González, Francisco García-Domínguez, Jaime Múñoz-Igualada, and José Vicente López-Bao. Multimethod, multistate Bayesian hierarchical modeling approach for use in regional monitoring of wolves. Conservation Biology, 30(4): 883–893, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jordan:2016:SCS

[JGS^{+}16] Rebecca Jordan, Steven Gray, Amanda Sorensen, Greg Newman, David Mellor, Greg Newman, Cindy Hmelo-Silver, Shannon LaDeau, Dawn Biehler, and Alycia Crall. Studying citizen science through adaptive management and learning feedbacks as mechanisms for improving conservation. Conservation Biology, 31(1):487–495, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Januchowski-Hartley:2018:NSE

[JHAH18] Stephanie R. Januchowski-Hartley, Vanessa M. Adams, and Virgilio Hermoso. The need for spatially explicit quantification of benefits in invasive-species management. Conservation Biology, 32(2):287–293, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Juslen:2013:ARL

[JHV13] Aino Juslén, Esko Hyvärinen, and Laura K. Virtanen. Application of the red-list index at a national level for multiple species groups. Conservation Biology, 27(2):398–406, April 2013. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jimenez:2014:TPO

[JIAMS^{+}14] Amanda Jiménez, Irene Iniesta-Arandia, Maria Muñoz-Santos, Berta Martín-López, Susan K. Jacobson, and Javier Benayas. Typology of public outreach for biodiversity conservation projects in Spain. Conservation Biology, 28(3):829–840, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jones:2019:ASI

[JKJ^{+}19] Sorrel Jones, Aidan Keane, Freya St John, Juliet Vickery, and Sarah Papworth. Audience segmentation to improve targeting of conservation interventions for hunters. Conservation Biology, 33(4):895–905, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jaramillo-Legorreta:2017:PAM

[JLCHNG^{+}17] Armando Jaramillo-Legorreta, Gustavo Cardenas-Hinojosa, Edwyna Nieto-Garcia, Lorenzo Rojas-Bracho, Jay Ver Hoef, Jeffrey Moore, Nicholas Tregenza, Jay Barlow, Tim Gerrodette, Len Thomas, and Barbara Taylor. Passive acoustic monitoring of the decline of Mexico’s critically endangered vaquita. Conservation Biology, 31(1):183–191, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jacobson:2012:CAN

[JLCK12] Susan K. Jacobson, Cynthia Langin, J. Stuart Carlton, and Lynda Lee Kaid. Content analysis of newspaper coverage of the Florida panther. Análisis del contenido de la cobertura periodística de la pantera de Florida. Conservation Biology, 26 (2):171–179, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Janin:2012:USH

[JLDJ12] Agnès Janin, Jean-Paul Léna, Sandrine Deblois, and Pierre Joly. Use of stress-hormone levels and habitat selection to assess functional connectivity of a landscape for an amphibian. Uso de niveles de hormona del Estrés y selección de hábitat para evaluar la conectividad funcional de un paisaje para una especie de anfibio. Conservation Biology, 26(5):923–931, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jarrad:2016:CBC

[JMB16] Frith Jarrad, Ellen Main, and Mark Burgman. Conservation Biology celebrates success. Conservation Biology, 30(5):929– 930, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jenkins:2012:ICI

[JMF12] Lekelia D. Jenkins, Sara M. Maxwell, and Erik Fisher. Increasing conservation impact and policy relevance of research through embedded experiences. Conservation Biology, 27(1): 740–742, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jackiw:2015:FGC

[JMH15] Raeya N. Jackiw, Ghada Mandil, and Heather A. Hager. A framework to guide the conservation of species hybrids based on ethical and ecological considerations. Conservation Biology, 29(4):1040–1051, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnson:2011:RDI

[JMP^{+}11] Pieter T. J. Johnson, Valerie J. Mckenzie, Anna C. Peterson, Jacob L. Kerby, Jennifer Brown, Andrew R. Blaustein, and Tina Jackson. Regional decline of an iconic amphibian associated with elevation, land-use change, and invasive species. Declinación regional de un anfibio icónico asociado con la altitud, el cambio de uso de suelo y especies invasoras. Conservation Biology, 25(6):556–566, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jorgensen:2013:CEC

[JMSM13] Jeffrey C. Jorgensen, Michelle M. Mcclure, Mindi B. Sheer, and Nancy L. Munn. Combined effects of climate change and bank stabilization on shallow water habitats of Chinook salmon. Conservation Biology, 27(6):1201–1211, December 2013. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnson:2011:TEI

[JMW^{+}11] Heather E. Johnson, L. Scott Mills, John D. Wehausen, Thomas R. Stephenson, and Gordon Luikart. Translating effects of inbreeding depression on component vital rates to overall population growth in endangered bighorn sheep. Traducción de los efectos de la depresión por endogamia sobre tasas vitales componentes del crecimiento poblacional de Ovis canadensis sierrae. Conservation Biology, 25(6):1240–1249, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johns:2010:AHS

[Joh10a] David Johns. Adapting human societies to conservation. Conservation Biology, 24(3):641–643, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johns:2010:IYB

[Joh10b] David Johns. The international year of biodiversity–from talk to action. Conservation Biology, 26(2):338–340, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnson:2011:CSD

[Joh11] Jerry Johnson. Conservation and social disarray: Response to Berger and Beckman. Conservation Biology, 25(6):631–632, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jorgensen:2015:CIP

[Jør15] Dolly Jørgensen. Conservation implications of parasite coreintroduction. Conservation Biology, 29(2):602–604, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Joseph:2012:CEG

[Jos12] Liana Joseph. Challenges in effective governance of natural resources. Conservation Biology, 26(3):578–579, June 2012. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jakalaniemi:2013:AST

[JPT13] Anne Jäkäläniemi, Heini Postila, and Juha Tuomi. Accuracy of short-term demographic data in projecting long-term fate of populations. Conservation Biology, 27(3):552–559, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jackson:2015:USR

[JRWA15] Emma L. Jackson, § Siân E. Rees, Catherine Wilding, and Martin J. Attrill. Use of a seagrass residency index to apportion commercial fishery landing values and recreation fisheries expenditure to seagrass habitat service. Conservation Biology, 29(3):899–909, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jyvasjarvi:2018:ITF

[JVI^{+}18] Jussi Jyväsjärvi, Risto Virtanen, Jari Ilmonen, Lauri Paasivirta, and Timo Muotka. Identifying taxonomic and functional surrogates for spring biodiversity conservation. Conservation Biology, 32(4):883–893, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jorge:2013:CBP

[JVT^{+}13] Agostinho A. Jorge, Abi T. Vanak, Maria Thaker, Colleen Begg, and Rob Slotow. Costs and benefits of the presence of leopards to the sport-hunting industry and local communities in Niassa National Reserve, Mozambique. Conservation Biology, 27(4):832–843, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jucker:2018:TYA

[JWS^{+}18] Tommaso Jucker, Bonnie Wintle, Gorm Shackelford, Pierre Bocquillon, Jan Laurens Geffert, Tim Kasoar, Eszter Kovacs, Hannah S. Mumby, Chloé Orland, Judith Schleicher, Eleanor R. Tew, Aiora Zabala, Tatsuya Amano, Alexandra Bell, Boris Bongalov, Josephine M. Chambers, Colleen Corrigan, América P. Durán, Leslie-Anne Duvic-Paoli, Caroline Emilson, Erik J. S. Emilson, Jéssica Fonseca da Silva, Emma E. Garnett, Elizabeth J. Green, Miriam K. Guth, Andrew Hacket-Pain, Amy Hinsley, Javier Igea, Martina Kunz, Sarah H. Luke, William Lynam, Philip A. Martin, Matheus H. Nunes, Nancy Ockendon, Aly Pavitt, Charlotte L. R. Payne, Victoria Plutshack, Tim T. Rademacher, Rebecca J. Robertson, David C. Rose, Anca Serban, Benno I. Simmons, Catherine Tayleur, Claire F. R. Wordley, and Nibedita Mukherjee. Ten-year assessment of the 100 priority questions for global biodiversity conservation. Conservation Biology, 32(6):1457–1463, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kahumbu:2010:ROR

[Kah10] Paula Kahumbu. Redefining our relationship with elephants. Conservation Biology, 24(2):636–637, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kallimanis:2018:CHC

[Kal18] Athanasios S. Kallimanis. Comprehending history for conserving the commons. Conservation Biology, 32(2):498–499, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keith:2018:SRS

[KAM18] David A. Keith, H. Resit Akçakaya, and Nicholas J. Murray. Scaling range sizes to threats for robust predictions of risks to biodiversity. Conservation Biology, 32(2):322–332, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kareiva:2014:NCS

[Kar14] Peter Kareiva. New conservation: Setting the record straight and finding common ground. Conservation Biology, 28(3):634– 636, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kiss:2018:GSE

[KAV18] Márton Kiss, Ildikó Arany, and Ágnes Vári. The general and the specific of ecosystem services. Conservation Biology, 33(1): 1212–1214, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kroodsma:2011:SRE

[KB11] Donald Kroodsma and Gregory F. Budney. Sound recordings — an essential tool for conservation. Conservation Biology, 25 (4):851–852, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kohler:2017:CNI

[KB17] Florent Kohler and Eduardo S. Brondizio. Considering the needs of indigenous and local populations in conservation programs. Conservation Biology, 31(2):245–251, April 2017. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keeley:2018:MHC

[KBC^{+}18] Annika T. H. Keeley, Galli Basson, D. Richard Cameron, Nicole E. Heller, Patrick R. Huber, Carrie A. Schloss, James H. Thorne, and Adina M. Merenlender. Making habitat connectivity a reality. Conservation Biology, 32(6):1221–1232, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kelly:2015:OFH

[KBCM15] Luke T. Kelly, Andrew F. Bennett, Michael F. Clarke, and Michael A. McCarthy. Optimal fire histories for biodiversity conservation. Conservation Biology, 29(2):473–481, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kelly:2017:PPW

[KBM17] Luke T. Kelly, Lluís Brotons, and Michael A. McCarthy. Putting pyrodiversity to work for animal conservation. Conservation Biology, 31(4):952–955, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kark:2015:ECC

[KBML15] Salit Kark, Eran Brokovich, Tessa Mazor, and Noam Levin. Emerging conservation challenges and prospects in an era of offshore hydrocarbon exploration and exploitation. Conservation Biology, 29(6):1573–1585, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kohl:2019:PVA

[KBSX19] P. A. Kohl, D. Brossard, D. A. Scheufele, and M. A. Xenos. Public views about editing genes in wildlife for conservation. Conservation Biology, 33(6):1286–1295, December 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kalinkat:2017:FUS

[KCD^{+}17] Gregor Kalinkat, Juliano S. Cabral, William Darwall, G. Francesco Ficetola, Judith L. Fisher, Darren P. Giling, Marie-Pierre Gosselin, Hans-Peter Grossart, Sonja C. Jähnig, Jonathan M. Jeschke, Klaus Knopf, Stefano Larsen, Gabriela Onandia, Marlene Pätzig, Wolf-Christian Saul, Gabriel Singer, Erik Sperfeld, and Ivan Jarić. Flagship umbrella species needed for the conservation of overlooked aquatic biodiversity. Conservation Biology, 31(3):481–485, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knight:2010:MHS

[KCDC10] Andrew T. Knight, Richard M. Cowling, Mark Difford, and Bruce M. Campbell. Mapping human and social dimensions of conservation opportunity for the scheduling of conservation action on private land. Mapeo de dimensiones humanas y sociales de la oportunidad de conservación para la programación de acciones de conservación en tierras privadas. Conservation Biology, 24(5):1348–1358, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kociolek:2011:ERN

[KCSP11] A. V. Kociolek, A. P. Clevenger, C. C. St. Clair, and D. S. Proppe. Effects of road networks on bird populations efectos de las redes de caminos sobre poblaciones de aves. Conservation Biology, 25(2):241–249, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kirkpatrick:2010:KTS

[KE10] R. Craig Kirkpatrick and Lucy Emerton. Killing tigers to save them: Fallacies of the farming argument. Matando tigres para salvarlos: Falacias del argumento de la reproducción en cautiverio. Conservation Biology, 24(3):655–659, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kendal:2018:NEE

[KF18a] Dave Kendal and Rebecca M. Ford. Need for empirical evidence to support use of social license in conservation: reply to Garnett et al. Conservation Biology, 32(3):737–739, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kendal:2018:RSL

[KF18b] Dave Kendal and Rebecca M. Ford. The role of social license in conservation. Conservation Biology, 32(2):493–495, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Koh:2010:MCS

[KG10] Lian Pin Koh and Jaboury Ghazoul. A matrix-calibrated species-area model for predicting biodiversity losses due to land-use change. Un modelo especies-Área calibrado para la matriz para predecir las Pérdidas de biodiversidad debido a cambios en el uso de suelo. Conservation Biology, 24(4):994– 1001, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kronenberger:2018:ETA

[KGF^{+}18] John A. Kronenberger, Jill C. Gerberich, Sarah W. Fitzpatrick, E. Dale Broder, Lisa M. Angeloni, and W. Chris Funk. An experimental test of alternative population augmentation scenarios. Conservation Biology, 32(4):838–848, August 2018. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kelly:2014:SPR

[KGG^{+}14] Terra R. Kelly, Jesse Grantham, Daniel George, Alacia Welch, Joseph Brandt, L. Joseph Burnett, Kelly J. Sorenson, Matthew Johnson, Robert Poppenga, David Moen, James Rasico, James W. Rivers, Carie Battistone, and Christine K. Johnson. Spatiotemporal patterns and risk factors for lead exposure in endangered California condors during 15 years of reintroduction. Conservation Biology, 28(6):1721–1730, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kery:2011:USC

[KGS^{+}11] Marc Kéry, Beth Gardner, Tabea Stoeckle, Darius Weber, and J. Andrew Royle. Use of spatial capture-recapture modeling and DNA data to estimate densities of elusive animals uso de modelos espaciales de captura-recaptura y ADN para estimar densidades de animales elusivos. Conservation Biology, 26(3):356–364, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kovacs-Hostyanszki:2016:WSA

[KH16] Anikó Kovács-Hostyánszki. On the way of sustainable agriculture conservation agriculture. Conservation Biology, 30(6): 1367–1369, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kaplan-Hallam:2018:ASI

[KHB18] Maery Kaplan-Hallam and Nathan J. Bennett. Adaptive social impact management for conservation and environmental management. Conservation Biology, 32(2):304–314, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kati:2015:CIE

[KHD^{+}15] Vassiliki Kati, Tasos Hovardas, Martin Dieterich, Pierre L. Ibisch, Barbara Mihok, and Nuria Selva. The challenge of implementing the European network of protected areas Natura 2000. Conservation Biology, 29(1):260–270, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kleinschroth:2017:ELR

[KHGF^{+}17] Fritz Kleinschroth, John R. Healey, Sylvie Gourlet-Fleury, Frédéric Mortier, and Radu S. Stoica. Effects of logging on roadless space in intact forest landscapes of the Congo Basin. Conservation Biology, 31(3):469–480, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kohler:2019:EDW

[KHK^{+}19] F. Kohler, T. G. Holland, J. S. Kotiaho, M. Desrousseaux, and M. D. Potts. Embracing diverse worldviews to share planet Earth. Conservation Biology, 33(5):1014–1022, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kirby:2014:NCM

[Kir14] Kathryn R. Kirby. “New conservation” as a moral imperative. Conservation Biology, 28(3):639–640, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kuempel:2019:QBM

[KJWP19] Caitlin D. Kuempel, Kendall R. Jones, James E. M. Watson, and Hugh P. Possingham. Quantifying biases in marineprotected-area placement relative to abatable threats. Conservation Biology, 33(6):1350–1359, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Koshida:2018:MAE

[KK18] Chieko Koshida and Naoki Katayama. Meta-analysis of the effects of rice-field abandonment on biodiversity in Japan. Conservation Biology, 32(6):1392–1402, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kelkar:2010:CFR

[KKCS10] Nachiket Kelkar, Jagdish Krishnaswamy, Sunil Choudhary, and Dipani Sutaria. Coexistence of fisheries with river dolphin conservation. Coexistencia de pesquerías con la conservación de delfines. Conservation Biology, 24(4):1130–1140, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kuparinen:2014:AEU

[KKH14] Anna Kuparinen, David M. Keith, and Jeffrey A. Hutchings. Allee effect and the uncertainty of population recovery. Conservation Biology, 28(3):790–798, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kansky:2014:MAA

[KKK14] Ruth Kansky, Martin Kidd, and Andrew T. Knight. Metaanalysis of attitudes toward damage-causing mammalian wildlife. Conservation Biology, 28(4):924–938, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kovach:2016:REH

[KLL^{+}16] Ryan P. Kovach, Gordon Luikart, Winsor H. Lowe, Matthew C. Boyer, and Clint C. Muhlfeld. Risk and efficacy of humanenabled interspecific hybridization for climate-change adaptation: response to Hamilton and Miller (2016). Conservation Biology, 30(2):428–430, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

King:2017:BFM

[KLN^{+}17] Lucy E. King, Fredrick Lala, Hesron Nzumu, Emmanuel Mwambingu, and Iain Douglas-Hamilton. Beehive fences as a multidimensional conflict-mitigation tool for farmers coexisting with elephants. Conservation Biology, 31(4):743–752, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kubon:2019:SSC

[KLS^{+}19] Maciej Kuboń, Agnieszka E. Latawiec, Fabio Rubio Scarano, Adam Drosik, Bernardo B. N. Strassburg, Włodzimierz Grzebieniowski, and Juan Gomes Bastos. Searching for solutions to the conflict over Europe’s oldest forest. Conservation Biology, 33(2):476–479, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Komonen:2018:DED

[KM18] Atte Komonen and Jörg Müller. Dispersal ecology of deadwood organisms and connectivity conservation. Conservation Biology, 32(3):535–545, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kumpel:2010:ASM

[KMGCR10] Noëlle F. Kümpel, E. J. Milner-Gulland, Guy Cowlishaw, and J. Marcus Rowcliffe. Assessing sustainability at multiple scales in a rotational bushmeat hunting system. Evaluación de la sustentabilidad en múltiples escalas en un sistema rotativo de cacería de vida silvestre. Conservation Biology, 25(6):861–871, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keith:2014:DER

[KMH^{+}14] David A. Keith, Michael Mahony, Harry Hines, Jane Elith, Tracey J. Regan, John B. Baumgartner, David Hunter, Geoffrey W. Heard, Nicola J. Mitchell, Kirsten M. Parris, Trent Penman, Ben Scheele, Christopher C. Simpson, Reid Tingley, Christopher R. Tracy, Matt West, and H. Resit Akçakaya. Detecting extinction risk from climate change by IUCN red list criteria. Conservation Biology, 28(3):810–819, June 2014. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Killion:2019:THC

[KMLC19] Alexander K. Killion, Tracy Melvin, Eric Lindquist, and Neil H. Carter. Tracking a half century of media reporting on gray wolves. Conservation Biology, 33(3):645–654, June 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kareksela:2018:EEE

[KMR^{+}18] Santtu Kareksela, Atte Moilanen, Olli Ristaniemi, Reima Välivaara, and Janne S. Kotiaho. Exposing ecological and economic costs of the research-implementation gap and compromises in decision making. Conservation Biology, 32(1):9–17, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kay:2017:EPP

[KMT^{+}17] Geoffrey M. Kay, Alessio Mortelliti, Ayesha Tulloch, Philip Barton, Daniel Florance, Saul A. Cunningham, and David B. Lindenmayer. Effects of past and present livestock grazing on herpetofauna in a landscape-scale experiment. Conservation Biology, 31(3):446–458, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kareksela:2013:UIS

[KMTK13] Santtu Kareksela, Atte Moilanen, Seppo Tuominen, and Janne S. Kotiaho. Use of inverse spatial conservation prioritization to avoid biological diversity loss outside protected areas. Conservation Biology, 27(6):1294–1303, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Katzner:2017:GEF

[KNB^{+}17] Todd E. Katzner, David M. Nelson, Melissa A. Braham, Jacqueline M. Doyle, Nadia B. Fernandez, Adam E. Duerr, Peter H. Bloom, Matthew C. Fitzpatrick, Tricia A. Miller, Renee C. E. Culver, Loan Braswell, and J. Andrew DeWoody. Golden eagle fatalities and the continental-scale consequences of local wind-energy generation. Conservation Biology, 31(3):406–415, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knight:2011:WAP

[Kni11] Andrew T. Knight. What about the people? Conservation Biology, 26(3):416–418, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Krishnadas:2013:ECC

[KNK13] Meghna Krishnadas, Tarun Nair, and Divya Karnad. Equality in conservation: Comment on Bawa et al. 2011. Conservation Biology, 27(2):422–424, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [BRS11].

Kinnaird:2012:EPL

[KO12] Margaret F. Kinnaird and Timothy G. O’Brien. Effects of private-land use, livestock management, and human tolerance on diversity, distribution, and abundance of large African mammals. Conservation Biology, 26(6):1026–1039, December 2012.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kamp:2015:GPC

[KOA^{+}15] Johannes Kamp, Steffen Oppel, Alexandr A. Ananin, Yurii A. Durnev, Sergey N. Gashev, Norbert Hölzel, Alexandr L. Mishchenko, Jorma Pessa, Sergey M. Smirenski, Evgenii G. Strelnikov, Sami Timonen, Kolja Wolanska, and Simba Chan. Global population collapse in a superabundant migratory bird and illegal trapping in China. Conservation Biology, 29(6): 1684–1694, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kelly:2016:TGF

[KP16] Ella Kelly and Ben L. Phillips. Targeted gene flow for conservation. Conservation Biology, 30(2):259–267, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kelly:2019:TGF

[KP19] Ella Kelly and Ben L. Phillips. Targeted gene flow and rapid adaptation in an endangered marsupial. Conservation Biology, 33(2):112–121, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kupferberg:2012:EFR

[KPL^{+}12] Sarah J. Kupferberg, Wendy J. Palen, Amy J. Lind, Steve Bobzien, Alessandro Catenazzi, Joe Drennan, and Mary E. Power. Effects of flow regimes altered by dams on survival, population declines, and range-wide losses of California riverbreeding frogs. Efecto del Régimen de flujos alterado por presas sobre la supervivencia, declinación poblacional y Pérdidas en el rango de distribución de ranas reproductivas en ríos de California. Conservation Biology, 26(3):513–524, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kranz:2012:BAZ

[Kra12] Karl R. Kranz. A book for all zoos. Conservation Biology, 26 (3):576, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kochalski:2019:PPR

[KRF^{+}19] Sophia Kochalski, Carsten Riepe, Marie Fujitani, Øystein Aas, and Robert Arlinghaus. Public perception of river fish biodiversity in four European countries. Conservation Biology, 33(2): 164–175, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kahler:2013:PRC

[KRG13] Jessica S. Kahler, Gary J. Roloff, and Meredith L. Gore. Poaching risks in community-based natural resource management. Conservation Biology, 27(1):177–186, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knowlton:2016:EFR

[KRL^{+}16] Amy R. Knowlton, Jooke Robbins, Scott Landry, Henry A. McKenna, Scott D. Kraus, and Timothy B. Werner. Effects of fishing rope strength on the severity of large whale entanglements. Conservation Biology, 30(2):318–328, April 2016. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kery:2010:SOD

[KRS^{+}10] Marc Kéry, J. Andrew Royle, Hans Schmid, Michael Schaub, Bernard Volet, Guido Häfliger, and Niklaus Zbinden. Siteoccupancy distribution modeling to correct population-trend estimates derived from opportunistic observations. Modelado de la distribución de la ocupación de sitios para corregir estimaciones de tendencias poblacionales derivadas de observaciones oportunistas. Conservation Biology, 24(5):1388–1397, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kissane:2017:RGN

[KS17] Zoe Kissane and Jill M. Shephard. The rise of glyphosate and new opportunities for biosentinel early-warning studies. Conservation Biology, 31(6):1293–1300, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kopf:2017:TBP

[KSH17] R. Keller Kopf, Casey Shaw, and Paul Humphries. Traitbased prediction of extinction risk of small-bodied freshwater fishes. Conservation Biology, 31(3):581–591, June 2017. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kuffner:2016:GPD

[KT16] Ilsa B. Kuffner and Lauren T. Toth. A geological perspective on the degradation and conservation of western Atlantic coral reefs. Conservation Biology, 30(4):706–715, August 2016. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Karimi:2017:UED

[KTBH17] Azadeh Karimi, Ayesha I. T. Tulloch, Greg Brown, and Marc Hockings. Understanding the effects of different social data on selecting priority conservation areas. Conservation Biology, 31 (6):1439–1449, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kuehne:2014:PSC

[KTF^{+}14] Lauren M. Kuehne, Laura A. Twardochleb, Keith J. Fritschie, Meryl C. Mims, David J. Lawrence, Polly P. Gibson, Ben Stewart-Koster, and Julian D. Olden. Practical science communication strategies for graduate students. Conservation Biology, 28(6):1225–1235, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Krosby:2010:ECC

[KTHH10] Meade Krosby, Joshua Tewksbury, Nick M. Haddad, and Jonathan Hoekstra. Ecological connectivity for a changing climate. Conservation Biology, 24(6):1686–1689, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kross:2012:EIT

[KTN12] Sara M. Kross, Jason M. Tylianakis, and Ximena J. Nelson. Effects of introducing threatened falcons into vineyards on abundance of passeriformes and bird damage to grapes. Efectos de la introducción de halcones amenazados en Viñedos sobre la abundancia de passeriformes y el daño de aves a uvas. Conservation Biology, 26(2):142–149, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kremen:2011:EQC

[KUT11] C. Kremen, K. S. Ullman, and R. W. Thorp. Evaluating the quality of citizen-scientist data on pollinator communities. Evaluación de la calidad de datos de comunidades de polinizadores tomados por Ciudadanos–Científicos. Conservation Biology, 25(6):607–617, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Khalumba:2014:CAP

[KWW^{+}14] Mercelyne Khalumba, Tobias WüNscher, Sven Wunder, Mirjam BüDenbender, and Karin Holm-MüLler. Combining auctions and performance-based payments in a forest enrichment field trial in Western Kenya. Conservation Biology, 28(5):861– 866, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kooyman:2013:EGF

[KZG^{+}13] R. M. Kooyman, A. E. Zanne, R. V. Gallagher, W. Cornwell, M. Rossetto, P. O’Connor, E. A. Parkes, C. F. Catterall, S. W. Laffan, and C. H. Lusk. Effects of growth form and functional traits on response of woody plants to clearing and fragmentation of subtropical rainforest. Conservation Biology, 28(1): 1468–1477, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

LopezTerceroCaamano:2011:VCS

[Lan11] José Antonio López Tercero Caamaño. A vision of conservation from school. Conservation Biology, 25(6):1091–1093, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lennox:2013:ALT

[LA13] Gareth D. Lennox and Paul R. Armsworth. The ability of landowners and their cooperatives to leverage payments greater than opportunity costs from conservation contracts. Conservation Biology, 27(3):625–634, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laikre:2010:NGD

[LAA^{+}10a] Linda Laikre, Fred W. Allendorf, Laurel C. Aroner, C. Scott Baker, David P. Gregovich, Michael M. Hansen, Jennifer A. Jackson, Katherine C. Kendall, Kevin Mckelvey, Maile C. Neel, Isabelle Olivieri, Nils Ryman, Michael K. Schwartz, Ruth Short Bull, Jeffrey B. Stetz, David A. Tallmon, Barbara L. Taylor, Christina D. Vojta, Donald M. Waller, and Robin S. Waples. Neglect of genetic diversity in implementation of the convention on biological diversity. Conservation Biology, 24(1):86–88, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lawrence:2010:BOV

[LAA^{+}10b] A. J. Lawrence, R. Afifi, M. Ahmed, S. Khalifa, and T. Paget. Bioactivity as an options value of sea cucumbers in the Egyptian Red Sea. La bioactividad como un valor de opción para pepinos marinos en el Mar Rojo Egipcio. Conservation Biology, 26(2):217–225, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lawler:2015:TBC

[LAA^{+}15] Joshua J. Lawler, David D. Ackerly, Christine M. Albano, Mark G. Anderson, Solomon Z. Dobrowski, Jacquelyn L. Gill, Nicole E. Heller, Robert L. Pressey, Eric W. Sanderson, and Stuart B. Weiss. The theory behind, and the challenges of, conserving nature’s stage in a time of rapid change. Conservation Biology, 29(3):618–629, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lalley:2011:PFC

[Lal11] Jennifer S. Lalley. Positive forcings in climate-change pedagogy. Conservation Biology, 25(4):850–851, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lazell:2011:IFT

[Laz11] James Lazell. Island facts and theories. Conservation Biology, 25(6):644–646, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Loss:2011:RDN

[LB11] Scott R. Loss and Robert B. Blair. Reduced density and nest survival of ground-nesting songbirds relative to earthworm invasions in northern hardwood forests. Reducción de la densidad y supervivencia de nidos de aves canoras que anidan sobre el suelo en Relación con invasiones de lombrices de tierra en bosques del norte. Conservation Biology, 27(3):983–992, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laub:2015:ALE

[LB15] Brian G. Laub and Phaedra Budy. Assessing the likely effectiveness of multispecies management for imperiled desert fishes with niche overlap analysis. Conservation Biology, 29(4):1153– 1163, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Luther:2016:DBC

[LBB^{+}16] David A. Luther, Thomas M. Brooks, Stuart H. M. Butchart, Matt W. Hayward, Marieke E. Kester, John Lamoreux, and Amy Upgren. Determinants of bird conservation-action implementation and associated population trends of threatened species. Conservation Biology, 30(6):1338–1346, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lundquist:2015:ECC

[LBD^{+}15] Carolyn J. Lundquist, András Báldi, Martin Dieterich, Kyle Gracey, Eszter Krasznai Kovacs, Judith Schleicher, Teuta Skorin, Eleanor Sterling, and Bengt-Gunnar Jonsson. Engaging the conservation community in the IPBES process. Conservation Biology, 29(6):1493–1495, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lawler:2019:TTC

[LBD^{+}19] Joshua Lawler, Paul Beier, Brett Dickson, Joseph Fargione, John Novembre, and David Theobald. A tribute to a true conservation innovator, Brad McRae, 1966–2017. Conservation Biology, 33(2):480–482, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Law:2018:ETO

[LBI^{+}18] Elizabeth A. Law, Nathan J. Bennett, Christopher D. Ives, Rachel Friedman, Katrina J. Davis, Carla Archibald, and Kerrie A. Wilson. Equity trade-offs in conservation decision making. Conservation Biology, 32(2):294–303, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lambrick:2014:ECF

[LBLB14] Frances H. Lambrick, Nick D. Brown, Anna Lawrence, and Daniel P. Bebber. Effectiveness of community forestry in Prey Long Forest, Cambodia. Conservation Biology, 28(2):372–381, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Liu:2015:OCB

[LCC^{+}15] Qiang Liu, Jin Chen, Richard T. Corlett, XuLi Fan, Dongli Yu, HongPei Yang, and JiangYun Gao. Orchid conservation in the biodiversity hotspot of southwestern China. Conservation Biology, 29(6):1563–1572, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Litt:2014:EIP

[LCFS14] Andrea R. Litt, Erin E. Cord, Timothy E. Fulbright, and Greta L. Schuster. Effects of invasive plants on arthropods. Conservation Biology, 28(6):1532–1549, December 2014. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lukey:2010:EIA

[LCG10] James R. Lukey, Stephen S. Crawford, and Daniel Gillis. Effect of information availability on assessment and designation of species at risk. Efecto de la disponibilidad de información sobre la evaluación y designación de especies en riesgo. Conservation Biology, 24(5):1398–1406, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lohr:2013:CBT

[LCL13] Cheryl A. Lohr, Linda J. Cox, and Christopher A. Lepczyk. Costs and benefits of trap-neuter-release and euthanasia for removal of urban cats in Oahu, Hawaii. Conservation Biology, 27(1):64–73, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Longenecker:2014:AEV

[LCT^{+}14] Ken Longenecker, Yvonne L. Chan, Robert J. Toonen, David B. Carlon, Terry L. Hunt, Alan M. Friedlander, and Edward E. Demartini. Archaeological evidence of validity of fish populations on unexploited reefs as proxy targets for modern populations. Conservation Biology, 28(6):1322–1330, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Longo:2019:ADF

[LCYJ19] Stefano B. Longo, Brett Clark, Richard York, and Andrew K. Jorgenson. Aquaculture and the displacement of fisheries captures. Conservation Biology, 33(4):832–841, August 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lepczyk:2010:WCB

[LDB^{+}10] Christopher A. Lepczyk, Nico Dauphiné, David M. Bird, Sheila Conant, Robert J. Cooper, David C. Duffy, Pamela Jo Hatley, Peter P. Marra, Elizabeth Stone, and Stanley A. Temple. What conservation biologists can do to counter trap-neuter-return: Response to Longcore et al. Conservation Biology, 24(2):627– 629, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lewis:2017:ADD

[LdBB^{+}17] Rob J. Lewis, Francesco de Bello, Jonathan A. Bennett, Pavel Fibich, Genevieve E. Finerty, Lars Götzenberger, Inga Hiiesalu, Liis Kasari, Jan Lepš, Maria Májeková, Ondřej Mudrák, Kersti Riibak, Argo Ronk, Terezie Rychtecká, Alena Vitová, and Meelis Pärtel. Applying the dark diversity concept to nature conservation. Conservation Biology, 31(1):40–47, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Luck:2011:RBU

[LDBS11] Gary W. Luck, Penny Davidson, Dianne Boxall, and Lisa Smallbone. Relations between urban bird and plant communities and human well-being and connection to nature. Relaciones entre las comunidades de aves y plantas urbanas y el bienestar humano y la conexión con la naturaleza. Conservation Biology, 25(4):816–826, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lebuhn:2013:DIP

[LDC^{+}13] Gretchen Lebuhn, Sam Droege, Edward F. Connor, Barbara Gemmill-Herren, Simon G. Potts, Robert L. Minckley, Terry Griswold, Robert Jean, Emanuel Kula, David W. Roubik, Jim Cane, Karen W. Wright, Gordon Frankie, and Frank Parker. Detecting insect pollinator declines on regional and global scales. Conservation Biology, 27(1):113–120, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lebuhn:2015:EBC

[LDC^{+}15] Gretchen Lebuhn, Sam Droege, Edward F. Connor, Barbara Gemmill-Herren, Simon G. Potts, Robert L. Minckley, Robert P. Jean, Emanuel Kula, David W. Roubik, Karen W.

Wright, Gordon Frankie, and Frank Parker. Evidence-based conservation: reply to Tepedino et al. Conservation Biology, 29(1):283–285, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lamothe:2019:CSC

[LDD19] Karl A. Lamothe, Alan J. Dextrase, and D. Andrew R. Drake. Characterizing species co-occurrence patterns of imperfectly detected stream fishes to inform species reintroduction efforts. Conservation Biology, 33(6):1392–1403, December 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lengyel:2017:NAC

[Len17] Szabolcs Lengyel. New approaches in the conservation of large open landscapes. Conservation Biology, 31(1):231–232, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leslie:2015:IPN

[Les15] Matthew S. Leslie. Impacts of phylogenetic nomenclature on the efficacy of the U.S. Endangered Species Act. Conservation Biology, 29(1):69–77, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lehtomaki:2019:PAP

[LET19] Joona Lehtomäki, Johanna Eklund, and Tuuli Toivonen. Plan S and academic publishing empires: response to Burgman et al. 2018. Conservation Biology, 33(5):1200–1202, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lotze:2017:CFR

[LFM17] Heike K. Lotze, Joanna Mills Flemming, and Anna M. Magera. Critical factors for the recovery of marine mammals. Conservation Biology, 31(6):1301–1311, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lopes:2019:BSF

[LFMB19] Ricardo Jorge Lopes, Juliana Machado Ferreira, and Nadia Moraes-Barros. Bolder steps to fight global wildlife illegal trade. Conservation Biology, 33(2):7–8, February 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ladle:2012:UDC

[LFMRD12] Richard J. Ladle, João V. L. Firmino, Ana C. M. Malhado, and Armando Rodríguez-Durán. Unexplored diversity and conservation potential of neotropical hot caves. Conservation Biology, 26(6):978–982, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leao:2014:PER

[LFPT14] Tarciso C. C. Leão, Carlos R. Fonseca, Carlos A. Peres, and Marcelo Tabarelli. Predicting extinction risk of Brazilian Atlantic forest angiosperms. Conservation Biology, 28(6):1349– 1359, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lennox:2017:VFC

[LFS^{+}17] Gareth D. Lennox, Joseph Fargione, Sacha Spector, Gwyn Williams, and Paul R. Armsworth. The value of flexibility in conservation financing. Conservation Biology, 31(3):666–674, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lucas:2017:VGC

[LGA17] Jane Lucas, Evan Gora, and Alfonso Alonso. A view of the global conservation job market and how to succeed in it. Conservation Biology, 31(6):1223–1231, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lu:2010:CCI

[LGB^{+}10] Flora Lu, Clark Gray, Richard E. Bilsborrow, Carlos F. Mena, Christine M. Erlien, Jason Bremner, Alisson Barbieri, and Stephen J. Walsh. Contrasting colonist and indigenous impacts on Amazonian forests. Contrastando impactos de colonizadores e indígenas sobre bosques amazónicos. Conservation Biology, 25(6):881–885, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lentini:2013:EPC

[LGC^{+}13] Pia E. Lentini, Philip Gibbons, Josie Carwardine, Joern Fischer, Michael Drielsma, and Tara G. Martin. Effect of planning for connectivity on linear reserve networks. Conservation Biology, 27(4):796–807, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Liu:2019:CIC

[LGCF19] Hong Liu, Stephan W. Gale, Mang Lung Cheuk, and Gunter A. Fischer. Conservation impacts of commercial cultivation of endangered and overharvested plants. Conservation Biology, 33 (2):288–299, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Linares:2012:AEM

[LGH^{+}12] Cristina Linares, Joaquim Garrabou, Bernat Hereu, David Diaz, Christian Marschal, Enric Sala, and Mikel Zabala. Assessing the effectiveness of marine reserves on unsustainably harvested long-lived sessile invertebrates. Evaluación de la efectividad de las reservas marinas sobre invertebrados Sésiles longevos recolectados de manera no sostenible. Conservation Biology, 26(2):88–96, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Little:2011:CNT

[LGK^{+}11] L. R. Little, R. Q. Grafton, T. Kompas, A. D. M. Smith, A. E. Punt, and B. D. Mapstone. Complementarity of no-take marine reserves and individual transferable catch quotas for managing the line fishery of the Great Barrier Reef. Complementariedad de reservas marinas sin captura y cuotas de captura individuales transferibles para el manejo de la pesca con anzuelo en la Gran Barrera Arrecifal. Conservation Biology, 26(3):333– 340, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leslie:2013:HGS

[LGM^{+}13] Heather M. Leslie, Erica Goldman, Karen L. Mcleod, Leila Sievanen, Hari Balasubramanian, Richard Cudney-Bueno, Amanda Feuerstein, Nancy Knowlton, Kai Lee, Richard Pollnac, and Jameal F. Samhouri. How good science and stories can go hand-in-hand. Conservation Biology, 28(4):1126–1129, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindenmayer:2010:SGC

[LH10] David Lindenmayer and Malcolm Hunter. Some guiding concepts for conservation biology. Algunos conceptos orientadores para la biología de la conservación. Conservation Biology, 24 (6):1459–1468, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leidner:2011:CMD

[LH11] Allison K. Leidner and Nick M. Haddad. Combining measures of dispersal to identify conservation strategies in fragmented landscapes. Combinación de medidas de dispersión para identificar estrategias de conservación en paisajes fragmentados. Conservation Biology, 27(3):1022–1031, October 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leduc:2019:EPR

[LH19] Antoine O. H. C. Leduc and Nigel E. Hussey. Evaluation of payfor-release conservation incentives for unintentionally caught threatened species. Conservation Biology, 33(4):953–961, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Latta:2012:UDA

[LHDC12] Steven C. Latta, Christine A. Howell, Mark D. Dettling, and Renée L. Cormier. Use of data on avian demographics and site persistence during overwintering to assess quality of restored riparian habitat. Utilización de datos demográficos de aves y de persistencia de sitio durante el invierno para evaluar la calidad de hábitat ribereño restaurado. Conservation Biology, 26 (3):482–492, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Larson:2016:CPD

[LHKA16] Eric R. Larson, Stephen Howell, Peter Kareiva, and Paul R. Armsworth. Constraints of philanthropy on determining the distribution of biodiversity conservation funding. Conservation Biology, 30(1):206–215, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lichtenstein:2016:OCN

[Lic16] Gabriela Lichtenstein. The old conservation, the new conservation, and the future of conservation. Conservation Biology, 30 (5):1135–1137, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Linnell:2011:RIP

[Lin11] John D. C. Linnell. The relative importance of predators and people in structuring and conserving ecosystems. Conservation Biology, 25(6):646–647, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laikre:2013:HEF

[LJA^{+}13] Linda Laikre, Mija Jansson, Fred W. Allendorf, Sven Jakobsson, and Nils Ryman. Hunting effects on favourable conservation status of highly inbred Swedish wolves. Conservation Biology, 27(2):248–253, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leroux:2013:LDA

[LK13] Shawn J. Leroux and Jeremy T. Kerr. Land development in and around protected areas at the wilderness frontier. Conservation Biology, 27(1):166–176, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Larue:2014:AVH

[LK14] Michelle A. Larue and Joseph Knight. Applications of very high-resolution imagery in the study and conservation of large predators in the Southern Ocean. Conservation Biology, 28(6): 1731–1735, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurance:2010:IPR

[LKB^{+}10] William F. Laurance, Lian P. Koh, Rhett Butler, Navjot S. Sodhi, Corey J. A. Bradshaw, J. David Neidel, Hazel Consunji, and Javier Mateo Vega. Improving the performance of the roundtable on sustainable palm oil for nature conservation. Conservation Biology, 24(2):377–381, April 2010. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Li:2014:PIG

[LKB^{+}14] Fengqing Li, Yong-Su Kwon, Mi-Jung Bae, Namil Chung, Tae-Sung Kwon, and Young-Seuk Park. Potential impacts of global warming on the diversity and distribution of stream insects in South Korea. Conservation Biology, 28(2):498–508, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lefevre:2013:DCF

[LKH^{+}13] François Lefèvre, Jarkko Koskela, Jason Hubert, Hojka Kraigher, Roman Longauer, Ditte C. Olrik, Silvio Schüler, Michele Bozzano, Paraskevi Alizoti, Remigijus Bakys, Cathleen Baldwin, Dalibor Ballian, Sanna Black-Samuelsson, Dagmar Bednarova, Sándor Bordács, Eric Collin, Bart De Cuyper, Sven M. G. De Vries, Thröstur Eysteinsson, Josef Frýdl, Michaela Haverkamp, Mladen Ivankovic, Heino Konrad, Czesław Koziol, Tiit Maaten, Eduardo Notivol Paino, Hikmet Öztürk, Ivanova Denitsa Pandeva, Gheorghe Parnuta, Andrej Pilipovič, Dragos Postolache, Cathal Ryan, Arne Steffenrem, Maria Carolina Varela, Federico Vessella, Roman T. Volosyanchuk, Marjana Westergren, Frank Wolter, Leena Yrjänä, and Inga Zarinja. Dynamic conservation of forest genetic resources in 33 European countries. Conservation Biology, 27(2):373–384, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Linnell:2015:FRB

[LKW^{+}15] John D. C. Linnell, Petra Kaczensky, Ulrich Wotschikowsky, Nicolas Lescureux, and Luigi Boitani. Framing the relationship between people and nature in the context of European conservation. Conservation Biology, 29(4):978–985, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Li:2015:SPP

[LKYL15] Rong Li, Nathan J. B. Kraft, Haiying Yu, and Heng Li. Seed plant phylogenetic diversity and species richness in conservation planning within a global biodiversity hotspot in eastern Asia. Conservation Biology, 29(6):1552–1562, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lohr:2014:DMP

[LL14] Cheryl A. Lohr and Christopher A. Lepczyk. Desires and management preferences of stakeholders regarding feral cats in the Hawaiian Islands. Conservation Biology, 28(2):392–403, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Liu:2015:IGC

[LLR^{+}15] Zhijin Liu, Guangjian Liu, Christian Roos, Ziming Wang, ZuoFu Xiang, Pingfen Zhu, Boshi Wang, Baoping Ren, Fanglei Shi, Huijuan Pan, and Ming Li. Implications of genetics and current protected areas for conservation of 5 endangered primates in China. Conservation Biology, 29(6):1508–1517, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindenmayer:2014:EAF

[LLW^{+}14] D. B. Lindenmayer, P. W. Lane, M. J. Westgate, M. Crane, D. Michael, S. Okada, and P. S. Barton. An empirical assessment of the focal species hypothesis. Conservation Biology, 28 (6):1594–1603, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Loss:2018:MDF

[LM18] Scott R. Loss and Peter P. Marra. Merchants of doubt in the free-ranging cat conflict. Conservation Biology, 32(2):265–266, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lavers:2014:PSD

[LMC^{+}14] Jennifer L. Lavers, Mark G. R. Miller, Michael J. Carter, George Swann, and Rohan H. Clarke. Predicting the spatial distribution of a seabird community to identify priority conservation areas in the Timor Sea. Conservation Biology, 28(6): 1699–1709, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Levine:2017:SOS

[LMCS17] Jordan Levine, Michael Muthukrishna, Kai M. A. Chan, and Terre Satterfield. Sea otters, social justice, and ecosystemservice perceptions in Clayoquot Sound, Canada. Conservation Biology, 31(3):343–352, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lewis:2019:DGP

[LMJB^{+}19] Edward Lewis, Brian MacSharry, Diego Juffe-Bignoli, Nyeema Harris, Georgina Burrows, Naomi Kingston, and Neil D. Burgess. Dynamics in the global protected-area estate since 2004. Conservation Biology, 33(3):570–579, June 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Louvrier:2019:UAD

[LMJK^{+}19] Julie Louvrier, Anja Molinari-Jobin, Marc Kéry, Thierry Chambert, David Miller, Fridolin Zimmermann, Eric Marboutin, Paolo Molinari, Oliver Müeller, Rok Černe, and Olivier Gimenez. Use of ambiguous detections to improve estimates from species distribution models. Conservation Biology, 33(2): 185–195, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lachish:2010:ESC

[LMM^{+}10] Shelly Lachish, Hamish Mccallum, Dydee Mann, Chrissy E. Pukk, and Menna E. Jones. Evaluation of selective culling of infected individuals to control Tasmanian Devil facial tumor disease. Evaluación de la Extracción selectiva de individuos infectados para controlar la enfermedad de tumor facial en el demonio de Tasmania. Conservation Biology, 25(6):841–851, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Littlefield:2017:CTC

[LMM^{+}17] Caitlin E. Littlefield, Brad H. McRae, Julia L. Michalak, Joshua J. Lawler, and Carlos Carroll. Connecting today’s climates to future climate analogs to facilitate movement of species under climate change. Conservation Biology, 31(6): 1397–1408, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leidner:2011:TGP

[LN11] Allison K. Leidner and Maile C. Neel. Taxonomic and geographic patterns of decline for threatened and endangered species in the United States. Patrones Taxonómicos y Geográficos de la declinación de especies amenazadas y en peligro en los Estados Unidos. Conservation Biology, 25(4):716–725, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lute:2016:MDH

[LNNG16] Michelle L. Lute, Carlos David Navarrete, Michael Paul Nelson, and Meredith L. Gore. Moral dimensions of human–wildlife conflict. Conservation Biology, 30(6):1200–1211, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lewandowski:2010:EST

[LNP10] Adam S. Lewandowski, Reed F. Noss, and David R. Parsons. The effectiveness of surrogate taxa for the representation of biodiversity. La efectividad de taxa sustitutos para la representación de biodiversidad. Conservation Biology, 24(5): 1367–1377, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Larson:2010:LEI

[LO10] Eric R. Larson and Julian D. Olden. Latent extinction and invasion risk of crayfishes in the Southeastern United States. Riesgo de extinción e invasión de langostinos en el Sureste de Estados Unidos. Conservation Biology, 24(4):1099–1110, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lopez-Osorio:2010:PAC

[LOME10] Federico López-Osorio and Daniel Rafael Miranda-Esquivel. A phylogenetic approach to conserving Amazonian biodiversity. Un enfoque Filogenético para conservar la biodiversidad Amazónica. Conservation Biology, 24(5):1359–1366, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Loring:2013:APS

[Lor13] Philip A. Loring. Alternative perspectives on the sustainability of Alaska’s commercial fisheries. Conservation Biology, 27(1): 55–63, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lovei:2010:WYS

[Löv10] Gabor L. Lövei. Would you start from here? Conservation Biology, 24(5):1430–1431, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lovei:2016:BRA

[Löv16] Gabor L. Lövei. Book review access. Conservation Biology, 30(4):918, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lepitzki:2010:BMC

[LP10] Dwayne A. W. Lepitzki and Charlie Pacas. Banff’s mountain caribou versus spring snails. Conservation Biology, 24(2):364– 365, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lumpkin:2013:EED

[LP13] Heather A. Lumpkin and Scott M. Pearson. Effects of exurban development and temperature on bird species in the Southern Appalachians. Conservation Biology, 28(4):1069–1078, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Li:2016:CEV

[LP16] Binbin V. Li and Stuart L. Pimm. China’s endemic vertebrates sheltering under the protective umbrella of the giant panda. Conservation Biology, 30(2):329–339, April 2016. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lim:2017:UPC

[LPA^{+}17] Cheng Ling Lim, Graham W. Prescott, Jose Don T. De Alban, Alan D. Ziegler, and Edward L. Webb. Untangling the proximate causes and underlying drivers of deforestation and forest degradation in Myanmar. Conservation Biology, 31(6): 1362–1372, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lynam:2016:I

[LPCA16] Antony J. Lynam, Lindsay Porter, and Ahimsa Campos-Arceiz. Introduction. Conservation Biology, 30(5):931–932, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lumpkin:2012:ECE

[LPT12] Heather A. Lumpkin, Scott M. Pearson, and Monica G. Turner. Effects of climate and exurban development on nest predation and predator presence in the southern Appalachian Mountains (U.S.A.). Efectos del clima y el desarrollo exurbano sobre la Depredación de nidos y la presencia de depredadores en el sur de las Montañas Apalaches (E.U.A.). Conservation Biology, 27 (1):679–688, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lukyanenko:2016:EPD

[LPW16] Roman Lukyanenko, Jeffrey Parsons, and Yolanda F. Wiersma. Emerging problems of data quality in citizen science. Conservation Biology, 30(3):447–449, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lechner:2014:CSU

[LRA^{+}14] A. M. Lechner, C. M. Raymond, V. M. Adams, M. Polyakov, A. Gordon, J. R. Rhodes, M. Mills, A. Stein, C. D. Ives, and E. C. Lefroy. Characterizing spatial uncertainty when integrating social data in conservation planning. Conservation Biology, 28(6):1497–1511, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Liu:2015:TTP

[LRL^{+}15] Hong Liu, Hai Ren, Qiang Liu, Xiang Ying Wen, Michael Maunder, and JiangYun Gao. Translocation of threatened plants as a conservation measure in China. Conservation Biology, 29(6):1537–1551, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lind-Riehl:2016:HAD

[LRMWG16] Jennifer F. Lind-Riehl, Audrey L. Mayer, Adam M. Wellstead, and Oliver Gailing. Hybridization, agency discretion, and implementation of the U.S. Endangered Species Act. Conservation Biology, 30(6):1288–1296, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lewandowski:2015:IVP

[LS15] Eva Lewandowski and Hannah Specht. Influence of volunteer and project characteristics on data quality of biological surveys. Conservation Biology, 29(3):713–723, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lambert:2019:REU

[LS19] Max R. Lambert and David A. Steen. Reexamining effects of urbanization on population structure of freshwater turtles: response to Bowne et al. 2018. Conservation Biology, 33(5): 1193–1196, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

LaRue:2017:FUH

[LSA17] Michelle A. LaRue, Seth Stapleton, and Morgan Anderson. Feasibility of using high-resolution satellite imagery to assess vertebrate wildlife populations. Conservation Biology, 31(1): 213–220, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lynn:2019:MPC

[LSÁL^{+}19] William S. Lynn, Francisco Santiago-Ávila, Joann Lindenmayer, John Hadidian, Arian Wallach, and Barbara J. King. A moral panic over cats. Conservation Biology, 33(4):769–776, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lawler:2010:PCI

[LSBB10] Joshua J. Lawler, Sarah L. Shafer, Betsy A. Bancroft, and Andrew R. Blaustein. Projected climate impacts for the amphibians of the Western Hemisphere. Proyección de impactos del clima sobre los anfibios del Hemisferio Occidental. Conservation Biology, 24(1):38–50, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lauber:2011:LKA

[LSDK11] T. Bruce Lauber, Richard C. Stedman, Daniel J. Decker, and Barbara A. Knuth. Linking knowledge to action in collaborative conservation. Vinculación del conocimiento y la acción en conservación colaborativa. Conservation Biology, 25(6):1186– 1194, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Linkie:2014:BVC

[LSK^{+}14] Matthew Linkie, Sean Sloan, Rahmad Kasia, Dedy Kiswayadi, and Wahdi Azmi. Breaking the vicious circle of illegal logging in Indonesia. Conservation Biology, 28(4):1023–1033, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laidre:2015:AMM

[LSK^{+}15] Kristin L. Laidre, Harry Stern, Kit M. Kovacs, Lloyd Lowry, Sue E. Moore, Eric V. Regehr, Steven H. Ferguson, Øystein Wiig, Peter Boveng, Robyn P. Angliss, Erik W. Born, Dennis Litovka, Lori Quakenbush, Christian Lydersen, Dag Vongraven, and Fernando Ugarte. Arctic marine mammal population status, sea ice habitat loss, and conservation recommendations for the 21st century. Conservation Biology, 29(3):724– 737, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lughadha:2019:HPI

[LSV^{+}19] Eimear M. Nic Lughadha, Vanessa Graziele Staggemeier, Thais N. C. Vasconcelos, Barnaby E. Walker, Cátia Canteiro, and Eve J. Lucas. Harnessing the potential of integrated systematics for conservation of taxonomically complex, megadiverse plant groups. Conservation Biology, 33(3):511–522, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindenmayer:2018:CRP

[LTN18] David Lindenmayer, Simon Thorn, and Reed Noss. Countering resistance to protected-area extension. Conservation Biology, 32(2):315–321, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Liu:2013:EAS

[LTQ^{+}13] Huiliang Liu, Ye Tao, Dong Qiu, Daoyuan Zhang, and Yongkuan Zhang. Effects of artificial sand fixing on community characteristics of a rare desert shrub. Conservation Biology, 28 (4):1011–1019, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lubin:2016:HSD

[Lub16] Yael Lubin. How to survive in the desert. Conservation Biology, 30(1):233–237, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lubin:2017:AAD

[Lub17] Yael Lubin. All about deserts. Conservation Biology, 31 (3):733–735, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lilieholm:2010:KFW

[LW10] Robert J. Lilieholm and W. Paul Weatherly. Kibale Forest wild coffee: Challenges to market-based conservation in Africa.

Café silvestre del Bosque Kibale: Retos para la conservación basada en mercados en África. Conservation Biology, 24(4): 924–930, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lamb:2011:UCD

[LW11] Joleah B. Lamb and Bette L. Willis. Using coral disease prevalence to assess the effects of concentrating tourism activities on offshore reefs in a tropical marine park. Utilización de la prevalencia de enfermedades del coral para evaluar los efectos de la Concentración de actividades Turísticas en arrecifes en un parque marino tropical. Conservation Biology, 27(3):1044–1052, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lacher:2014:WCA

[LW14] Iara Lacher and Marit L. Wilkerson. Wildlife connectivity approaches and best practices in U.S. state wildlife action plans. Conservation Biology, 28(1):13–21, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Liu:2018:ELL

[LW18] Xueqin Liu and Hongzhu Wang. Effects of loss of lateral hydrological connectivity on fish functional diversity. Conservation Biology, 32(6):1336–1345, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Liu:2012:EPR

[LWLC12] Shuang Liu, Terry Walshe, Graham Long, and David Cook. Evaluation of potential responses to invasive non-native species with structured decision making. Conservation Biology, 26 (3):539–546, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Li:2014:RTB

[LWY^{+}14] Juan Li, Dajun Wang, Hang Yin, Duojie Zhaxi, Zhala Jiagong, George B. Schaller, Charudutt Mishra, Thomas M. Mccarthy, Hao Wang, Lan Wu, Lingyun Xiao, Lamao Basang, Yuguang Zhang, Yunyun Zhou, and Zhi Lu. Role of Tibetan Buddhist monasteries in snow leopard conservation. Conservation Biology, 28(1):87–94, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Liang:2016:EPB

[LWZL16] Jingjing Liang, James V. Watson, Mo Zhou, and Xiangdong Lei. Effects of productivity on biodiversity in forest ecosystems across the United States and China. Conservation Biology, 30 (2):308–317, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lynch:2013:MIL

[Lyn13] Amanda H. Lynch. Methods for indigenous land-use and occupancy mapping. Conservation Biology, 28(4):1130–1131, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Levin:2013:MLA

[LZD^{+}13] I. I. Levin, P. Zwiers, S. L. Deem, E. A. Geest, J. M. Higashiguchi, T. A. Iezhova, G. Jiménez-Uzcátegui, D. H. Kim, J. P. Morton, N. G. Perlut, R. B. Renfrew, E. H. R. Sari, G. Valkiunas, and P. G. Parker. Multiple lineages of avian malaria parasites (Plasmodium) in the Galapagos Islands and evidence for arrival via migratory birds. Conservation Biology, 27(6):1366–1377, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Levin:2014:CCC

[LZD^{+}14] I. I. Levin, P. Zwiers, S. L. Deem, E. A. Geest, J. M. Higashiguchi, T. A. Iezhova, G. Jiménez-Uzcátegui, D. H. Kim, J. P. Morton, N. G. Perlut, R. B. Renfrew, E. H. R. Sari, G. Valkiunas, and P. G. Parker. Cover caption: Conservation Biology. Conservation Biology, 27(6):i, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leiper:2018:QCP

[LZR^{+}18] Ian Leiper, Kerstin K. Zander, Cathy J. Robinson, Josie Carwadine, Bradley J. Moggridge, and Stephen T. Garnett. Quantifying current and potential contributions of Australian indigenous peoples to threatened species management. Conservation Biology, 32(5):1038–1047, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lu:2017:RCN

[LZZ^{+}17] Yihe Lü, Liwei Zhang, Yuan Zeng, Bojie Fu, Charlotte Whitham, Shuguang Liu, and Bingfang Wu. Representation of critical natural capital in China. Conservation Biology, 31 (4):894–902, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Matar:2010:AMT

[MA10] Diane A. Matar and Brandon P. Anthony. Application of modified threat reduction assessments in Lebanon. Aplicación de evaluaciones modificadas de reducción de amenazas en Líbano. Conservation Biology, 24(5):1174–1181, October 2010. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McClanahan:2019:CNE

[MA19] Tim R. McClanahan and Caroline A. Abunge. Conservation needs exposed by variability in common-pool governance principles. Conservation Biology, 33(4):917–929, August 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maynard:2010:MMM

[MAA^{+}10] Jeffrey A. Maynard, Kenneth R. N. Anthony, Siham Afatta, Nancy Dahl-Tacconi, and Ove Hoegh-Guldberg. Making a model meaningful to coral reef managers in a developing nation: a case study of overfishing and rock anchoring in Indonesia. Diseño de un modelo significativo para manejadores de arrecifes de coral en un País en desarrollo: un estudio de caso de sobrepesca y anclaje con rocas en Indonesia. Conservation Biology, 24(5):1316–1326, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcclure:2013:ICS

[MAB^{+}13] Michelle M. Mcclure, Michael Alexander, Diane Borggaard, David Boughton, Lisa Crozier, Roger Griffis, Jeffrey C. Jorgensen, Steven T. Lindley, Janet Nye, Melanie J. Rowland, Erin E. Seney, Amy Snover, Christopher Toole, and Kyle Van Houtan. Incorporating climate science in applications of the U.S. Endangered Species Act for aquatic species. Conservation Biology, 27(6):1222–1233, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Milder:2015:AAI

[MAB^{+}15] Jeffrey C. Milder, Margaret Arbuthnot, Allen Blackman, Sharon E. Brooks, Daniele Giovannucci, Lee Gross, Elizabeth T. Kennedy, Kristin Komives, Eric F. Lambin, Audrey Lee, Daniel Meyer, Peter Newton, Ben Phalan, Götz Schroth, Bambi Semroc, Henk Van Rikxoort, and Michal Zrust. An agenda for assessing and improving conservation impacts of sustainability standards in tropical agriculture. Conservation Biology, 29(2):309–320, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martinez-Abrain:2018:TSH

[MAG18] Alejandro Martínez-Abraín and Pedro Galán. A test of the substitution–habitat hypothesis in amphibians. Conservation Biology, 32(3):725–730, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mueller:2010:ERP

[MAGA10] Jocelyn G. Mueller, Issoufou Hassane Bil Assanou, Iro Dan Guimbo, and Astier M. Almedom. Evaluating rapid participatory rural appraisal as an assessment of ethnoecological knowledge and local biodiversity patterns. Evaluación del Diagnóstico rural participativo como una Valoración del conocimiento etnoecológico y los patrones de biodiversidad local. Conservation Biology, 24(1):140–150, February 2010. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martinez-Abrain:2016:AAI

[MAJ16] Alejandro Martínez-Abraín and Juan Jiménez. Anthropogenic areas as incidental substitutes for original habitat. Conservation Biology, 31(1):593–598, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moon:2014:MCC

[MAJH^{+}14] Katie Moon, Vanessa M. Adams, Stephanie R. Januchowski-Hartley, Maksym Polyakov, Morena Mills, Duan Biggs, Andrew T. Knight, Edward T. Game, and Christopher M. Raymond. A multidisciplinary conceptualization of conservation opportunity. Conservation Biology, 28(6):1484–1496, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maiorano:2015:HMB

[MAM^{+}15] L. Maiorano, G. Amori, A. Montemaggiori, C. Rondinini, L. Santini, S. Saura, and L. Boitani. On how much biodiversity is covered in Europe by national protected areas and by the Natura 2000 network: insights from terrestrial vertebrates.

Conservation Biology, 29(4):986–995, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martin:2010:MID

[Mar10] Laura Martin. Metaphor and the idea of a dominant conservation ethic. Conservation Biology, 24(5):1172–1173, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marvier:2014:NCT

[Mar14] Michelle Marvier. New conservation is true conservation. Conservation Biology, 28(1):1–3, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martinez-Abrain:2011:CET

[MARV^{+}11] Alejandro Martínez-Abraín, Helen M. Regan, Covadonga Viedma, Elena Villuendas, Miguel Angel Bartolomé, Juan Antonio Gómez, and Daniel Oro. Cost-effectiveness of translocation options for a threatened waterbird. Rentabilidad de las opciones de translocación de una especie de ave Acuática amenazada. Conservation Biology, 25(4):726–735, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Matlack:2013:RUF

[Mat13] Glenn R. Matlack. Reassessment of the use of fire as a management tool in deciduous forests of Eastern North America. Conservation Biology, 28(4):916–926, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Matlack:2015:MFM

[Mat15] Glenn R. Matlack. Managing fire in the mesic deciduous forest when fire history is unknown: response to Stambaugh et al. Conservation Biology, 29(3):947–949, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mitchell:2012:RBC

[MAWF12] Toby Mitchell, Lesley A. Alton, Craig R. White, and Craig E. Franklin. Relations between conspecific density and effects of ultraviolet-B radiation on tadpole size in the striped marsh frog. Conservation Biology, 26(6):1112–1120, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maris:2010:AMA

[MB10] Virginie Maris and Arnaud BéChet. From adaptive management to adjustive management: A pragmatic account of biodiversity values. Del manejo adaptativo al manejo ajustativo: Un recuento pragmático de los valores de la biodiversidad. Conservation Biology, 24(4):966–973, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moon:2014:GUS

[MB14] Katie Moon and Deborah Blackman. A guide to understanding social science research for natural scientists. Conservation Biology, 28(5):1167–1177, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mayor:2017:AMR

[MBBB17] Pedro Mayor, Hani El Bizri, Richard E. Bodmer, and Mark Bowler. Assessment of mammal reproduction for hunting sustainability through community-based sampling of species in the wild. Conservation Biology, 31(4):912–923, August 2017. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mavrommati:2016:IEB

[MBBH16] Georgia Mavrommati, Kostas Bithas, Mark E. Borsuk, and Richard B. Howarth. Integration of ecological–biological thresholds in conservation decision making. Conservation Biology, 30(6):1173–1181, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marco:2016:QRI

[MBC^{+}16] Moreno Di Marco, Thomas Brooks, Annabelle Cuttelod, Lincoln D. C. Fishpool, Carlo Rondinini, Robert J. Smith, Leon Bennun, Stuart H. M. Butchart, Simon Ferrier, Ruud P. B. Foppen, Lucas Joppa, Diego Juffe-Bignoli, Andrew T. Knight, John F. Lamoreux, Penny F. Langhammer, Ian May, Hugh P. Possingham, Piero Visconti, James E. M. Watson, and Stephen Woodley. Quantifying the relative irreplaceability of important bird and biodiversity areas. Conservation Biology, 30(2):392– 402, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martin:2012:EEK

[MBF^{+}12] Tara G. Martin, Mark A. Burgman, Fiona Fidler, Petra M. Kuhnert, Samantha Low-Choy, Marissa Mcbride, and Kerrie Mengersen. Eliciting expert knowledge in conservation science. Obtención de conocimiento de expertos en ciencia de la conservación. Conservation Biology, 26(2):29–38, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moat:2018:RAO

[MBFB18] Justin Moat, Steven P. Bachman, Richard Field, and Doreen S. Boyd. Refining area of occupancy to address the modifiable areal unit problem in ecology and conservation. Conservation Biology, 32(6):1278–1289, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:2014:UPB

[MBG14] Kimberly A. Miller, Trent P. Bell, and Jennifer M. Germano. Understanding publication bias in reintroduction biology by assessing translocations of New Zealand’s herpetofauna. Conservation Biology, 28(4):1045–1056, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meeus:2011:EIP

[MBGS11] Ivan Meeus, Mark J. F. Brown, Dirk C. De Graaf, and Guy Smagghe. Effects of invasive parasites on bumble bee declines. Efectos de parásitos invasores sobre declinaciones de abejorros. Conservation Biology, 25(4):662–671, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mackenzie:2014:USL

[MBH14] Catrina A. Mackenzie, Timothy D. Baird, and Joel Hartter. Use of single large or several small policies as strategies to manage people–park interactions. Conservation Biology, 28(6): 1645–1656, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morgan-Brown:2010:QAT

[MBJWC10] Theron Morgan-Brown, Susan K. Jacobson, Kenneth Wald, and Brian Child. Quantitative assessment of a Tanzanian integrated conservation and development project involving butterfly farming. Evaluación cuantitativa de un proyecto integrado de conservación y desarrollo que involucra el cultivo de mariposas en Tanzania. Conservation Biology, 24(2):563–572, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:2014:ARB

[MBL^{+}14] Tricia A. Miller, Robert P. Brooks, Michael Lanzone, David Brandes, Jeff Cooper, Kieran O’Malley, Charles Maisonneuve, Junior Tremblay, Adam Duerr, and Todd Katzner. Assessing risk to birds from industrial wind energy development via paired resource selection models. Conservation Biology, 28 (3):745–755, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marco:2014:REG

[MBM^{+}14] M. Di Marco, L. Boitani, D. Mallon, M. Hoffmann, A. Iacucci, E. Meijaard, P. Visconti, J. Schipper, and C. Rondinini. A retrospective evaluation of the global decline of carnivores and ungulates. Conservation Biology, 28(4):1109–1118, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Manfredo:2017:WSV

[MBT^{+}17a] Michael J. Manfredo, Jeremy T. Bruskotter, Tara L. Teel, David Fulton, Shalom H. Schwartz, Robert Arlinghaus, Shigehiro Oishi, Ayse K. Uskul, Kent Redford, Shinobu Kitayama, and Leeann Sullivan. Why social values cannot be changed for the sake of conservation. Conservation Biology, 31(4):772– 780, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Manfredo:2017:RCI

[MBT^{+}17b] Michael J. Manfredo, Jeremy T. Bruskotter, Tara L. Teel, David C. Fulton, Shigehiro Oishi, Ayse K. Uskul, Kent H. Redford, Shalom H. Schwartz, Robert Arlinghaus, Shinobu Kitayama, and Leeann Sullivan. Revisiting the challenge of intentional value shift: reply to Ives and Fischer. Conservation Biology, 31(6):1486–1487, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marco:2016:STO

[MBV^{+}16] Moreno Di Marco, Stuart H. M. Butchart, Piero Visconti, Graeme M. Buchanan, Gentile F. Ficetola, and Carlo Rondinini. Synergies and trade-offs in achieving global biodiversity targets. Conservation Biology, 30(1):189–195, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McCarthy:2019:PPR

[MBW^{+}19] Michael A. McCarthy, Mark A. Burgman, Fuwen Wei, Frith C. Jarrad, Carlo Rondinini, Carolina Murcia, Helene D. Marsh, H. Resit Akçakaya, Karen J. Esler, Edward T. Game, and Mark W. Schwartz. Plan S and publishing: reply to Lehtomäki et al. 2019. Conservation Biology, 33(5):1203–1204, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Milder:2011:CDP

[MC11a] Jeffrey C. Milder and Story Clark. Conservation development practices, extent, and land-use effects in the United States. Práctica, alcance y efectos sobre el uso de suelo de la conservación y desarrollo. Conservation Biology, 25(4):697–707, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moon:2011:LBA

[MC11b] Katie Moon and Chris Cocklin. A landholder-based approach to the design of private-land conservation programs. Un método basado en propietarios para el diseño de programas de conservación en terrenos privados. Conservation Biology, 25(3):493– 503, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moran:2014:RSD

[MC14] Cath Moran and Carla P. Catterall. Responses of seeddispersing birds to amount of rainforest in the landscape around fragments. Conservation Biology, 28(2):551–560, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcclanahan:2010:EFC

[Mcc10] Timothy R. Mcclanahan. Effects of fisheries closures and gear restrictions on fishing income in a Kenyan coral reef. Efectos del cierre de pesquerías y de restricciones de equipo sobre el ingreso de pescadores en un arrecife de coral Keniano. Conservation Biology, 24(6):1519–1528, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McConney:2014:SFW

[McC14] Patrick McConney. Sustainable fisheries — are we there yet? Conservation Biology, 28(1):292–293, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Merenlender:2016:EEE

[MCD^{+}16] Adina M. Merenlender, Alycia W. Crall, Sabrina Drill, Michelle Prysby, and Heidi Ballard. Evaluating environmental education, citizen science, and stewardship through naturalist programs. Conservation Biology, 30(6):1255–1265, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

May:2019:RDC

[MCD^{+}19] Tom W. May, Jerry A. Cooper, Anders Dahlberg, Giuliana Furci, David W. Minter, Gregory M. Mueller, Alison Pouliot, and Zhuliang Yang. Recognition of the discipline of conservation mycology. Conservation Biology, 33(3):733–736, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McClenachan:2016:CIO

[MCH^{+}16] Loren McClenachan, Andrew B. Cooper, Marah Hardt, Matthew McKenzie, and Joshua A. Drew. Conservation implications of omitting historical data sources: response to Baisre. Conservation Biology, 30(1):226–227, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mclaughlin:2011:LFB

[Mcl11] David W. Mclaughlin. Land, food, and biodiversity. Conservation Biology, 25(6):1117–1120, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:2012:FBC

[MCL12] Brian W. Miller, Susan C. Caplow, and Paul W. Leslie. Feedbacks between conservation and social-ecological systems. Retroalimentación entre conservación y sistemas socio– ecológicos. Conservation Biology, 26(2):218–227, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ma:2019:CAN

[MCM^{+}19] Zhijun Ma, Ying Chen, David S. Melville, Jun Fan, Jianguo Liu, Jinwei Dong, Kun Tan, Xuefei Cheng, Richard A. Fuller, Xiangming Xiao, and Bo Li. Changes in area and number of nature reserves in China. Conservation Biology, 33(5):1066–1075, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meijaard:2015:GBC

[MCMP15] Erik Meijaard, Marcel Cardillo, Emily M. Meijaard, and Hugh P. Possingham. Geographic bias in citation rates of conservation research. Conservation Biology, 29(3):920–925, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mascia:2010:IMP

[MCN10] Michael B. Mascia, C. Anne Claus, and Robin Naidoo. Impacts of Marine Protected Areas on fishing communities. Impactos de las Áreas Marinas Protegidas sobre comunidades de pescadores. Conservation Biology, 24(5):1424–1429, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marks:2017:TAF

[MCO^{+}17] Clive A. Marks, Malcolm Clark, David Obendorf, Graham P. Hall, Inês Soares, and Filipe Pereira. Trends in anecdotal fox sightings in Tasmania accounted for by psychological factors. Conservation Biology, 31(6):1450–1458, December 2017. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McPherson:2011:GBL

[McP11] Guy McPherson. Going back to the land in the age of entitlement. Conservation Biology, 25(5):855–857, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McPherson:2012:CAP

[McP12] Guy R. McPherson. Choosing an alternative path. Conservation Biology, 26(3):383–384, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Milt:2018:MOC

[MDD^{+}18] Austin W. Milt, Matthew W. Diebel, Patrick J. Doran, Michael C. Ferris, Matthew Herbert, Mary L. Khoury, Allison T. Moody, Thomas M. Neeson, Jared Ross, Ted Treska, Jesse R. O’Hanley, Lisa Walter, Steven R. Wangen, Eugene Yacobson, and Peter B. McIntyre. Minimizing opportunity costs to aquatic connectivity restoration while controlling an invasive species. Conservation Biology, 32(4):894–904, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mouysset:2014:PVA

[MDJ14] L. Mouysset, L. Doyen, and F. Jiguet. From population viability analysis to coviability of farmland biodiversity and agriculture. Conservation Biology, 28(1):187–201, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Melvin:2019:LSC

[MDSF19] Edward F. Melvin, Kimberly S. Dietrich, Robert M. Suryan, and Shannon M. Fitzgerald. Lessons from seabird conservation in Alaskan longline fisheries. Conservation Biology, 33(4):842– 852, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marchetti:2016:CPE

[ME16] Michael P. Marchetti and Tag Engstrom. The conservation paradox of endangered and invasive species. Conservation Biology, 30(2):434–437, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Medellin:2010:LTB

[Med10] Rodrigo A. Medellin. Of lions, and tigers, and bears, oh my! Conservation Biology, 25(6):910, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meine:2012:RC

[Mei12] Curt Meine. Reading(s) for comprehension. Conservation Biology, 27(1):749–751, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Melick:2010:CRE

[Mel10] David Melick. Credibility of REDD and experiences from Papua New Guinea. Conservation Biology, 24(2):359–361, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meretsky:2014:LCP

[MF14] Vicky J. Meretsky and Robert L. Fischman. Learning from conservation planning for the U.S. National Wildlife Refuges. Conservation Biology, 28(6):1415–1427, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maslo:2015:CSB

[MF15] Brooke Maslo and Nina H. Fefferman. A case study of bats and white-nose syndrome demonstrating how to model population viability with evolutionary effects. Conservation Biology, 29(4): 1176–1185, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

MacGregor-Fors:2016:ECC

[MF16] Ian MacGregor-Fors. Ecological curio-city. Conservation Biology, 30(4):920–922, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Minin:2019:FII

[MFHT19] Enrico Di Minin, Christoph Fink, Tuomo Hiippala, and Henrikki Tenkanen. A framework for investigating illegal wildlife trade on social media with machine learning. Conservation Biology, 33(2):210–213, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcmullan-Fisher:2010:SMM

[MFKMP10] Sapphire J. M. Mcmullan-Fisher, Jamie B. Kirkpatrick, Tom W. May, and Emma J. Pharo. Surrogates for macrofungi and mosses in reservation planning. Sustitutos de macrohongos y musgos en la planificación de reservas. Conservation Biology, 24(3):730–736, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martin:2011:AMF

[MFN^{+}11] Julien Martin, Paul L. Fackler, James D. Nichols, Michael C. Runge, Carol L. Mcintyre, Bruce L. Lubow, Maggie C. Mccluskie, and Joel A. Schmutz. An adaptive-management framework for optimal control of hiking near golden eagle nests in Denali National Park. Un marco de referencia de manejo adaptativo para el control Óptimo del excursionismo en el Parque Nacional Denali. Conservation Biology, 26(3):316–323, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mccarthy:2012:PSD

[MFRH12] Kyle P. Mccarthy, Robert J. Fletcher, Jr., Christopher T. Rota, and Richard L. Hutto. Predicting species distributions from samples collected along roadsides. Predicción de la distribución de especies a partir de muestras recolectadas a lo largo de carreteras. Conservation Biology, 26(2):68–77, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mastrangelo:2012:TOB

[MG12] Matias E. Mastrangelo and Michael C. Gavin. Trade-offs between cattle production and bird conservation in an agricultural frontier of the Gran Chaco of Argentina. Conservation Biology, 26(6):1040–1051, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Milt:2016:CAE

[MGA16] Austin W. Milt, Tamara D. Gagnolet, and Paul R. Armsworth. The costs of avoiding environmental impacts from shale-gas surface infrastructure. Conservation Biology, 30(6):1151–1158, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ma:2014:EIC

[MGCL14] Zhijun Ma, Xiaojing Gan, Chi-Yeung Choi, and Bo Li. Effects of invasive cordgrass on presence of marsh grassbird in an area where it is not native. Conservation Biology, 28(1):150–158, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Muter:2013:AUN

[MGG^{+}13] Bret A. Muter, Meredith L. Gore, Katie S. Gledhill, Christopher Lamont, and Charlie Huveneers. Australian and U.S. news media portrayal of sharks and their conservation. Conservation Biology, 27(1):187–196, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Milner-Gulland:2014:AIC

[MGMA^{+}14] E. J. Milner-Gulland, J. A. Mcgregor, M. Agarwala, G. Atkinson, P. Bevan, T. Clements, T. Daw, K. Homewood, N. Kumpel, J. Lewis, S. Mourato, B. Palmer Fry, M. Redshaw, J. M. Rowcliffe, S. Suon, G. Wallace, H. Washington, and D. Wilkie. Accounting for the impact of conservation on human well-being. Conservation Biology, 28(5):1160–1166, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McClanahan:2015:BBT

[MGMC15] T. R. McClanahan, N. A. J. Graham, M. A. MacNeil, and J. E. Cinner. Biomass-based targets and the management of multispecies coral reef fisheries. Conservation Biology, 29(2):409– 417, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murray:2015:SPC

[MGW^{+}15] Josil P. Murray, Richard Grenyer, Sven Wunder, Niels Raes, and Julia P. G. Jones. Spatial patterns of carbon, biodiversity, deforestation threat, and REDD+ projects in Indonesia. Conservation Biology, 29(5):1434–1445, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:2016:IHC

[MH16] Joshua M. Miller and Jill A. Hamilton. Interspecies hybridization in the conservation toolbox: response to Kovach et al. (2016). Conservation Biology, 30(2):431–433, April 2016. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcnew:2014:EWE

[MHG^{+}14] Lance B. Mcnew, Lyla M. Hunt, Andrew J. Gregory, Samantha M. Wisely, and Brett K. Sandercock. Effects of wind energy development on nesting ecology of greater prairie-chickens in fragmented grasslands. Conservation Biology, 28(4):1089– 1099, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Michel:2015:FF

[Mic15] Nicole L. Michel. The future of forests. Conservation Biology, 29(3):968–970, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mihok:2019:EYA

[Mih19] Barbara Mihók. Everything you always wanted to know about conservation and human behavior but were afraid to ask. Conservation Biology, 33(3):739–741, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mccay:2011:MPA

[MJ11] Bonnie J. Mccay and Peter J. S. Jones. Marine protected areas and the governance of marine ecosystems and fisheries. Conservation Biology, 25(6):1130–1133, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Macdonald:2011:ABM

[MJA^{+}11] David W. Macdonald, Paul J. Johnson, Lise Albrechtsen, Adam Dutton, Sarah Seymour, Jef Dupain, Amy Hall, and Julia E. Fa. Association of body mass with price of bushmeat in Nigeria and Cameroon. Asociación de la masa corporal con el precio de carne silvestre en Nigeria y Camerún. Conservation Biology, 25(6):1220–1228, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moss:2015:ECZ

[MJG15] Andrew Moss, Eric Jensen, and Markus Gusset. Evaluating the contribution of zoos and aquariums to Aichi Biodiversity Target 1. Conservation Biology, 29(2):537–544, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mills:2011:IEC

[MJP^{+}11] Morena Mills, Stacy D. Jupiter, Robert L. Pressey, Natalie C. Ban, and James Comley. Incorporating effectiveness of community-based management in a national marine gap analysis for Fiji. Incorporación de la efectividad del manejo basado en comunidades en un análisis de vacíos nacional marino en Fiji. Conservation Biology, 25(6):1155–1164, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McClanahan:2018:RBD

[MK18] T. R. McClanahan and J. K. Kosgei. Redistribution of benefits but not detection in a fisheries bycatch-reduction management initiative. Conservation Biology, 32(1):159–170, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McNamara:2015:LTS

[MKR^{+}15] J. McNamara, J. M. Kusimi, J. M. Rowcliffe, G. Cowlishaw, A. Brenyah, and E. J. Milner-Gulland. Long-term spatiotemporal changes in a West African bushmeat trade system. Conservation Biology, 29(5):1446–1457, October 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Minor:2010:MNA

[ML10] Emily S. Minor and Todd R. Lookingbill. A multiscale network analysis of protected-area connectivity for mammals in the United States. Un análisis multiescala de la conectividad de Áreas protegidas para mamíferos en los Estados Unidos. Conservation Biology, 24(6):1549–1558, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Madliger:2014:NPC

[ML14] Christine L. Madliger and Oliver P. Love. The need for a predictive, context-dependent approach to the application of stress hormones in conservation. Conservation Biology, 28(1): 283–287, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mortelliti:2015:ELT

[ML15] Alessio Mortelliti and David B. Lindenmayer. Effects of landscape transformation on bird colonization and extinction patterns in a large-scale, long-term natural experiment. Conservation Biology, 29(5):1314–1326, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moura:2014:THY

[MLA^{+}14] Nárgila G. Moura, Alexander C. Lees, Alexandre Aleixo, Jos Barlow, Sidnei M. Dantas, Joice Ferreira, Maria De Fátima C. Lima, and Toby A. Gardner. Two hundred years of local avian extinctions in Eastern Amazonia. Conservation Biology, 28(6):

1271–1281, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Magrach:2014:MAE

[MLLS14] Ainhoa Magrach, William F. Laurance, Asier R. Larrinaga, and Luis Santamaria. Meta-analysis of the effects of forest fragmentation on interspecific interactions. Conservation Biology, 28 (6):1342–1348, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Minin:2015:IPS

[MLMP^{+}15] Enrico Di Minin, Jussi Laitila, Federico Montesino-Pouzols, Nigel Leader-Williams, Rob Slotow, Peter S. Goodman, Anthony J. Conway, and Atte Moilanen. Identification of policies for a sustainable legal trade in rhinoceros horn based on population projection and socioeconomic models. Conservation Biology, 29(2):545–555, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Michalak:2018:DPC

[MLRC18] Julia L. Michalak, Joshua J. Lawler, David R. Roberts, and Carlos Carroll. Distribution and protection of climatic refugia in North America. Conservation Biology, 32(6):1414–1425, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Magrach:2012:EMC

[MLS12] Ainhoa Magrach, Asier R. Larrinaga, and Luis Santamaría. Effects of matrix characteristics and interpatch distance on functional connectivity in fragmented temperate rainforests. Efectos de las características de la matriz y la distancia entre parches sobre la conectividad funcional en bosques lluviosos templados fragmentados. Conservation Biology, 26(2):238–247, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moore:2010:VIA

[MM10] Alana L. Moore and Michael A. Mccarthy. On valuing information in adaptive-management models. Valoración de la información en modelos de manejo adaptativo. Conservation Biology, 24(4):984–993, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Madin:2015:FEG

[MM15] Joshua S. Madin and Elizabeth M. P. Madin. The full extent of the global coral reef crisis. Conservation Biology, 29 (6):1724–1726, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Molina-Montenegro:2012:ONN

[MMCUR^{+}12] Marco A. Molina-Montenegro, Fernando Carrasco-Urra, Cristian Rodrigo, Peter Convey, Fernando Valladares, and Ernesto Gianoli. Occurrence of the non-native annual bluegrass on the Antarctic mainland and its negative effects on native plants. Ocurrencia de pasto anual no nativo en el Continente Antártico y sus efectos negativos sobre plantas nativas. Conservation Biology, 27(1):717–723, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Minin:2013:CBC

[MMG^{+}13] Enrico Di Minin, Douglas Craig Macmillan, Peter Styan Goodman, Boyd Escott, Rob Slotow, and Atte Moilanen. Conservation businesses and conservation planning in a biological diversity hotspot. Conservation Biology, 27(4):808–820, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mace:2018:RDB

[MMGN18] Georgina M. Mace, E. J. Milner-Gulland, and Emily Nicholson. Remembering Dr. Ben Collen, an exemplary conservation biologist. Conservation Biology, 32(6):1473–1475, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mccarthy:2014:LIB

[MMK^{+}14] Michael A. Mccarthy, Alana L. Moore, Jochen Krauss, John W. Morgan, and Christopher F. Clements. Linking indices for biodiversity monitoring to extinction risk theory. Conservation Biology, 28(6):1575–1583, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:2012:SDG

[MMM^{+}12] Kimberly A. Miller, Hilary C. Miller, Jennifer A. Moore, Nicola J. Mitchell, Alison Cree, Fred W. Allendorf, Stephen D.

Sarre, Susan N. Keall, and Nicola J. Nelson. Securing the demographic and genetic future of tuatara through assisted colonization. Aseguramiento del futuro demográfico y genético de la tuatara mediante colonización asistida. Conservation Biology, 26(5):790–798, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mitrovich:2010:AME

[MMPF10] Milan J. Mitrovich, Tritia Matsuda, Krista H. Pease, and Robert N. Fisher. Ants as a measure of effectiveness of habitat conservation planning in Southern California. Hormigas como una medida de la efectividad de la planificación de la conservación del hábitat en el Sur de California. Conservation Biology, 24(5):1239–1248, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mills:2014:MCK

[MNW^{+}14] Morena Mills, Sam Nicol, Jessie A. Wells, José J. Lahoz-Monfort, Brendan Wintle, Michael Bode, Martin Wardrop, Terry Walshe, William J. M. Probert, Michael C. Runge, Hugh P. Possingham, and Eve Mcdonald Madden. Minimizing the cost of keeping options open for conservation in a changing climate. Conservation Biology, 28(3):646–653, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McDonald:2017:HMM

[MNW^{+}17] Peter J. McDonald, Catherine E. M. Nano, Simon J. Ward, Alistair Stewart, Chris R. Pavey, Gary W. Luck, and Chris R. Dickman. Habitat as a mediator of mesopredator-driven mammal extinction. Conservation Biology, 31(5):1183–1191, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcclanahan:2011:CMH

[MO11] Timothy R. Mcclanahan and Johnstone O. Omukoto. Comparison of modern and historical fish catches (AD 750–1400) to inform goals for Marine Protected Areas and sustainable fisheries. Comparación de capturas modernas e históricas (750–1400 AD) para metas en Áreas Marinas Protegidas y pesquerías sustentables. Conservation Biology, 25(5):945–955, October 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moller:2017:CTB

[Møl17] Ian Max Møller. Camera traps, bird song, and creatures of darkness. Conservation Biology, 31(4):958–960, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Molnar:2019:RND

[Mol19] Zsolt Molnár. Role of narratives in driving conservation science. Conservation Biology, 33(4):979–980, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moorhouse:2017:BCA

[Moo17] Ron Moorhouse. Biodiversity conservation in Australia and New Zealand. Conservation Biology, 31(1):229–231, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moorhouse:2019:MCB

[Moo19] Ron Moorhouse. Monitoring and conserving biodiversity in Australia. Conservation Biology, 33(2):485–486, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moretti:2012:SWS

[Mor12] David Moretti. Science on whales, submarines, and conservation meets the news. Conservation Biology, 26(3):379–381, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morrison:2015:FCH

[Mor15] Scott A. Morrison. A framework for conservation in a humandominated world. Conservation Biology, 29(3):960–964, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moorhead:2010:EBC

[MPB10] Leigh C. Moorhead, Stacy M. Philpott, and Peter Bichier. Epiphyte biodiversity in the coffee agricultural matrix: Canopy stratification and distance from forest fragments. Biodiversidad de epífitas en la matriz agrícola de café: Estratificación del dosel y distancia a los fragmentos de bosque. Conservation Biology, 24(3):737–746, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maranto:2011:UFA

[MPH^{+}11] Christina J. Maranto, Julia K. Parrish, David P. Herman, André E. Punt, Julian D. Olden, Michael T. Brett, and Daniel D. Roby. Use of fatty acid analysis to determine dispersal of Caspian terns in the Columbia River Basin, U.S.A. Utilización de análisis de Ácidos grasos para determinar la dispersión de Hydroprogne caspia en la Cuenca del Río Columbia, E.U.A. Conservation Biology, 25(4):736–746, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Metcalf:2019:MC

[MPP^{+}19] Alexander L. Metcalf, Conor N. Phelan, Cassandra Pallai, Michael Norton, Ben Yuhas, James C. Finley, and Allyson Muth. Microtargeting for conservation. Conservation Biology, 33(5):1141–1150, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Monsarrat:2016:SEE

[MPS^{+}16] Sophie Monsarrat, M. Grazia Pennino, Tim D. Smith, Randall R. Reeves, Christine N. Meynard, David M. Kaplan, and Ana S. L. Rodrigues. A spatially explicit estimate of the prewhaling abundance of the endangered North Atlantic right whale. Conservation Biology, 30(4):783–791, August 2016. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meinard:2014:EBB

[MQ14] Yves Meinard and Fabien Quétier. Experiencing biodiversity as a bridge over the science–society communication gap. Conservation Biology, 28(3):705–712, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:2012:MDY

[MQAMWM12] Timothy J. Miller, Pedro F. Quintana-Ascencio, Satya Maliakal-Witt, and Eric S. Menges. Metacommunity dynamics over 16 years in a pyrogenic shrubland. Dinámica de la metacomunidad durante 16 años en un matorral pirogénico. Conservation Biology, 26(2):357–366, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moore:2012:CSD

[MR12] Joslin L. Moore and Michael C. Runge. Combining structured decision making and value-of-information analyses to identify robust management strategies. Combinación de toma de decisiones estructurada y Análisis del valor de la información para identificar estrategias de manejo robustas. Conservation Biology, 26(5):810–820, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McClanahan:2016:GCS

[MR16] T. R. McClanahan and P. S. Rankin. Geography of conservation spending, biodiversity, and culture. Conservation Biology, 30(5):1089–1101, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martensen:2012:AFC

[MRBL^{+}12] Alexandre Camargo Martensen, Milton Cezar Ribeiro, Cristina Banks-Leite, Paulo Inácio Prado, and Jean Paul Metzger. Associations of forest cover, fragment area, and connectivity with neotropical understory bird species richness and abundance. Conservation Biology, 26(6):1100–1111, December 2012. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McConnachie:2016:ICT

[MRFvW16] Matthew M. McConnachie, Claudia Romero, Paul J. Ferraro, and Brian W. van Wilgen. Improving credibility and transparency of conservation impact evaluations through the partial identification approach. Conservation Biology, 30(2):371–381, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mannocci:2017:ECD

[MRMH17] Laura Mannocci, Jason J. Roberts, David L. Miller, and Patrick N. Halpin. Extrapolating cetacean densities to quantitatively assess human impacts on populations in the high seas. Conservation Biology, 31(3):601–614, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mineur:2014:PFL

[MRMV14] Frederic Mineur, Auguste Le Roux, Christine A. Maggs, and Marc Verlaque. Positive feedback loop between introductions of non-native marine species and cultivation of oysters in Europe. Conservation Biology, 28(6):1667–1676, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moreno-Rueda:2012:NSD

[MRPPM12] Gregorio Moreno-Rueda, Juan M. Pleguezuelos, Manuel Pizarro, and Albert Montori. Northward shifts of the distributions of Spanish reptiles in association with climate change. Cambios hacia el norte en las distribuciones de reptiles españoles en asociación con el cambio climático. Conservation Biology, 26(2):278–283, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mellish:2019:RMR

[MRPT19] Sarah Mellish, Jillian C. Ryan, Elissa L. Pearson, and Michelle R. Tuckey. Research methods and reporting practices in zoo and aquarium conservation-education evaluation. Conservation Biology, 33(2):40–52, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mockrin:2011:ELF

[MRRK11] Miranda H. Mockrin, Robert F. Rockwell, Kent H. Redford, and Nicholas S. Keuler. Effects of landscape features on the distribution and sustainability of ungulate hunting in Northern Congo. Efectos de las características del paisaje sobre la distribución y sustentabilidad de la cacería de ungulados en el Norte de Congo. Conservation Biology, 25(3):514–525, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morano:2012:ADY

[MRT^{+}12] Janelle L. Morano, Aaron N. Rice, Jamey T. Tielens, Bobbi J. Estabrook, Anita Murray, Bethany L. Roberts, and Christopher W. Clark. Acoustically detected year-round presence of right whales in an urbanized migration corridor. Conservation Biology, 27(1):698–707, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martin:2012:RFP

[MS12] Meghan S. Martin and David J. Shepherdson. Role of familiarity and preference in reproductive success in ex situ breeding programs. Papel de la familiaridad y la preferencia en el Éxito reproductivo en programas de reproducción ex situ. Conservation Biology, 26(4):649–656, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meredith:2018:PSP

[MSC^{+}18] Helen M. R. Meredith, Freya A. V. St. John, Ben Collen, Simon A. Black, and Richard A. Griffiths. Practitioner and scientist perceptions of successful amphibian conservation. Conservation Biology, 32(2):366–375, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Markl:2012:MAE

[MSF^{+}12] Julia S. Markl, Matthias Schleuning, Pierre Michel Forget, Pedro Jordano, Joanna E. Lambert, Anna Traveset, S. Joseph Wright, and Katrin Böhning-Gaese. Meta-analysis of the effects of human disturbance on seed dispersal by animals. Conservation Biology, 26(6):1072–1081, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Munro:2019:AIS

[MSL19] David Munro, Jamie Steer, and Wayne Linklater. On allegations of invasive species denialism. Conservation Biology, 33 (4):797–802, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marcilio-Silva:2018:LUT

[MSMCB18] Vinícius Marcilio-Silva, Márcia C. M. Marques, and Jeannine Cavender-Bares. Land-use trade-offs between tree biodiversity and crop production in the Atlantic forest. Conservation Biology, 32(5):1074–1084, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moshier:2019:NAS

[MSR19] Andrea Moshier, Janna Steadman, and David L. Roberts. Network analysis of a stakeholder community combatting illegal wildlife trade. Conservation Biology, 33(6):1307–1317, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murray:2009:IDD

[MSSM09] Kris A. Murray, Lee F. Skerratt, Rick Speare, and Hamish Mccallum. Impact and dynamics of disease in species threatened by the amphibian chytrid fungus, Batrachochytrium dendrobatidis. Impacto y dinámica de la enfermedad en especies amenazadas por el hongo quitridio anfibio Batrachochytrium dendrobatidis. Conservation Biology, 23(5):1242–1252, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comment [Sch10].

Murray:2010:EEE

[MSSM10] Kris A. Murray, Lee F. Skerratt, Rick Speare, and Hamish Mccallum. Evidence of effects of endemic chytridiomycosis on host survival, behavior, and emigration: Reply to Schmidt. Conservation Biology, 25(6):900–902, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mondol:2015:TGO

[MSY^{+}15] Samrat Mondol, Vanjulavalli Sridhar, Prasanjeet Yadav, Sanjay Gubbi, and Uma Ramakrishnan. Tracing the geographic origin of traded leopard body parts in the Indian Subcontinent with DNA-based assignment tests. Conservation Biology, 29 (2):556–564, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Margalida:2014:DAT

[MSZB^{+}14] Antoni Margalida, José A. Sánchez-Zapata, Guillermo Blanco, Fernando Hiraldo, and José A. Donázar. Diclofenac approval as a threat to Spanish vultures. Conservation Biology, 28(3):631– 632, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Menzel:2010:ESS

[MT10] Susanne Menzel and Jack Teng. Ecosystem services as a stakeholder-driven concept for conservation science. Conservation Biology, 25(6):907–909, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Manfredo:2016:IHV

[MTD16] Michael J. Manfredo, Tara L. Teel, and Alia M. Dietsch. Implications of human value shift and persistence for biodiversity conservation. Conservation Biology, 30(2):287–296, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Matteson:2012:ACC

[MTM12] K. C. Matteson, D. J. Taron, and E. S. Minor. Assessing citizen contributions to butterfly monitoring in two large cities. Evaluación de contribuciones de ciudadanos al monitoreo de mariposas en dos ciudades grandes. Conservation Biology, 26 (3):557–564, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcmullin:2013:LCH

[MTN13] Richard Troy Mcmullin, Ian D. Thompson, and Steven G. Newmaster. Lichen conservation in heavily managed boreal forests. Conservation Biology, 28(4):1020–1030, October 2013. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mattsson:2018:PSC

[MV18] Brady J. Mattsson and Harald Vacik. Prospects for stakeholder coordination by protected-area managers in Europe. Conservation Biology, 32(1):98–108, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moir:2010:CCF

[MVB^{+}10] Melinda L. Moir, Peter A. Vesk, Karl E. C. Brennan, David A. Keith, Lesley Hughes, and Michael A. Mccarthy. Current constraints and future directions in estimating coextinction. Restricciones actuales y directrices futuras en la estimación de la coextinción. Conservation Biology, 24(3):682–690, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moir:2011:IMT

[MVB^{+}11] Melinda L. Moir, Peter A. Vesk, Karl E. C. Brennan, David A. Keith, Michael A. Mccarthy, and Lesley Hughes. Identifying and managing threatened invertebrates through assessment of coextinction risk. Identificación y manejo de invertebrados amenazados mediante la estimación del riesgo de coextinción. Conservation Biology, 25(4):787–796, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moir:2012:CED

[MVB^{+}12] Melinda L. Moir, Peter A. Vesk, Karl E. C. Brennan, Robert Poulin, Lesley Hughes, David A. Keith, Michael A. Mccarthy, and David J. Coates. Considering extinction of dependent species during translocation, ex situ conservation, and assisted migration of threatened hosts. Consideración de la extinción de especies dependientes durante la conservación ex situ y migración asistida de hospederos amenazados. Conservation Biology, 26(2):199–207, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Metcalfe:2015:SSE

[MVVS15] Kristian Metcalfe, Gregory Vaughan, Sandrine Vaz, and Robert J. Smith. Spatial, socio-economic, and ecological implications of incorporating minimum size constraints in marine protected area network design. Conservation Biology, 29(6): 1615–1625, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murphy:2011:CAU

[MWC11] Dennis D. Murphy, Paul S. Weiland, and Kenneth W. Cummins. A critical assessment of the use of surrogate species in conservation planning in the Sacramento–San Joaquin Delta, California (U.S.A.). Una evaluación crítica del uso de especies sucedáneas en la planificación de la conservación en el Delta Sacramento–San Joaquín, California (E.U.A.). Conservation Biology, 25(5):873–878, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Montambault:2015:UMD

[MWL^{+}15] Jensen Reitz Montambault, Supin Wongbusarakum, Trina Leberer, Eugene Joseph, Wayne Andrew, Fran Castro, Brooke Nevitt, Yimnang Golbuu, Noelle W. Oldiais, Craig R. Groves, Willy Kostka, and Peter Houk. Use of monitoring data to support conservation management and policy decisions in Micronesia. Conservation Biology, 29(5):1279–1289, October 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mccauley:2014:PNE

[MYG^{+}14] Douglas J. Mccauley, Hillary S. Young, Roger Guevara, Gareth J. Williams, Eleanor A. Power, Robert B. Dunbar, Douglas W. Bird, William H. Durham, and Fiorenza Micheli. Positive and negative effects of a threatened parrotfish on reef ecosystems. Conservation Biology, 28(6):1312–1321, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nolte:2013:LME

[NA13] Christoph Nolte and Arun Agrawal. Linking management effectiveness indicators to observed effects of protected areas on fire occurrence in the Amazon rainforest. Conservation Biology, 27(1):155–165, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Negret:2017:NCP

[NAB^{+}17] Pablo Jose Negret, James Allan, Alexander Braczkowski, Martine Maron, and James E. M. Watson. Need for conservation planning in postconflict Colombia. Conservation Biology, 31 (3):499–500, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nash:2017:MHA

[Nas17] David Nash. Mutualisms and how to analyze them. Conservation Biology, 31(6):1498–1499, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Natusch:2018:SWL

[Nat18] Daniel James Deans Natusch. Solutions to wildlife laundering in Indonesia: reply to Janssen and Chng 2018. Conservation Biology, 32(3):731–733, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [JC18].

Nuno:2018:UIC

[NBA^{+}18] A. Nuno, J. M. Blumenthal, T. J. Austin, J. Bothwell, G. Ebanks-Petrie, B. J. Godley, and A. C. Broderick. Understanding implications of consumer behavior for wildlife farming and sustainable wildlife trade. Conservation Biology, 32 (2):390–400, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nielsen:2017:MGO

[NBH^{+}17] Erica S. Nielsen, Maria Beger, Romina Henriques, Kimberly A. Selkoe, and Sophie von der Heyden. Multispecies genetic objectives in spatial conservation planning. Conservation Biology, 31 (4):872–882, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

NGoran:2012:HLE

[NBM^{+}12] Paul K. N’Goran, Christophe Boesch, Roger Mundry, Eliezer K. N’Goran, Ilka Herbinger, Fabrice A. Yapi, and Hjalmar S. Kühl. Hunting, law enforcement, and African primate conservation. Conservation Biology, 26(3):565–571, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nuno:2013:NAA

[NBNMG13] Ana Nuno, Nils Bunnefeld, Loiruck C. Naiman, and E. J. Milner-Gulland. A novel approach to assessing the prevalence and drivers of illegal bushmeat hunting in the Serengeti. Conservation Biology, 27(6):1355–1365, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nicotra:2015:ACA

[NBR^{+}15] Adrienne B. Nicotra, Erik A. Beever, Amanda L. Robertson, Gretchen E. Hofmann, and John O’Leary. Assessing the components of adaptive capacity to improve conservation and management efforts under global change. Conservation Biology, 29 (5):1268–1278, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noon:2012:ESL

[NBSM12] Barry R. Noon, Larissa L. Bailey, Thomas D. Sisk, and Kevin S. Mckelvey. Efficient species-level monitoring at the landscape scale. Conservation Biology, 26(3):432–441, June 2012. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nadon:2012:RCM

[NBW^{+}12] Marc O. Nadon, Julia K. Baum, Ivor D. Williams, Jana M. Mcpherson, Brian J. Zgliczynski, Benjamin L. Richards, Robert E. Schroeder, and Russell E. Brainard. Re-creating missing population baselines for Pacific reef sharks. Recreación de las Líneas de base poblacionales faltantes para tiburones de arrecife en el Pacífico. Conservation Biology, 26(3):493–503, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Neel:2013:PRC

[NCC13] Maile C. Neel and Judy P. Che-Castaldo. Predicting recovery criteria for threatened and endangered plant species on the basis of past abundances and biological traits. Conservation Biology, 27(2):385–397, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Northridge:2017:DCP

[NCKC17] Simon Northridge, Alex Coram, Al Kingston, and Rory Crawford. Disentangling the causes of protected-species bycatch in gillnet fisheries. Conservation Biology, 31(3):686–695, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:2012:BTC

[NDB^{+}12a] Reed F. Noss, Andrew P. Dobson, Robert Baldwin, Paul Beier, Cory R. Davis, Dominick A. Dellasala, John Francis, Harvey Locke, Katarzyna Nowak, Roel Lopez, Conrad Reining, Stephen C. Trombulak, and Gary Tabor. Bolder thinking for conservation. Conservation Biology, 26(1):1–4, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:2012:CC

[NDB^{+}12b] Reed F. Noss, Andrew P. Dobson, Robert Baldwin, Paul Beier, Cory R. Davis, Dominick A. Dellasala, John Francis, Harvey Locke, Katarzyna Nowak, Roel Lopez, Conrad Reining, Stephen C. Trombulak, and Gary Tabor. Cover caption. Conservation Biology, 26(1):i, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nazareno:2012:SNT

[NFdC^{+}12] Alison G. Nazareno, Juliana M. Feres, Dulcineia de Carvalho, Alexandre M. Sebbenn, Thomas E. Lovejoy, and William F. Laurance. Serious new threat to Brazilian forests. Conservation Biology, 26(2):5–6, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nelson:2010:PES

[NFF^{+}10] Fred Nelson, Charles Foley, Lara S. Foley, Abraham Leposo, Edward Loure, David Peterson, Mike Peterson, Thad Peterson, Hassan Sachedina, and Andrew Williams. Payments for ecosystem services as a framework for community-based conservation in Northern Tanzania. Pago de los servicios ecosistémicos como marco de referencia para la conservación basada en comunidades en el Norte de Tanzania. Conservation Biology, 24 (1):78–85, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nilsson:2016:CME

[NGB^{+}16] Danielle Nilsson, Galina Gramotnev, Greg Baxter, James R. A. Butler, Serge A. Wich, and Clive A. McAlpine. Community motivations to engage in conservation behavior to conserve the Sumatran orangutan. Conservation Biology, 30(4):816–826, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nyaki:2014:LSD

[NGL^{+}14] Angela Nyaki, Steven A. Gray, Christopher A. Lepczyk, Jeffrey C. Skibins, and Dennis Rentsch. Local-scale dynamics and local drivers of bushmeat trade. Conservation Biology, 28(6): 1403–1414, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nathan:2015:UED

[NJBM15] Lucas R. Nathan, Christopher L. Jerde, Michelle L. Budny, and Andrew R. Mahon. The use of environmental DNA in invasive species surveillance of the Great Lakes commercial bait trade. Conservation Biology, 29(2):430–439, April 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Niedzialkowski:2019:EMC

[NJK^{+}19] Krzysztof Niedziałkowski, Bogdan Jaroszewicz, Rafał Kowalczyk, Dries P. J. Kuijper, Grzegorz Mikusiński, Nuria Selva, Wiesław Walankiewicz, and Tomasz Wesołowski. Effective mitigation of conservation conflicts and participatory governance: reflections on Kuboń et al. Conservation Biology, 33(4):962– 965, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nielsen:2014:FDC

[NJT14] Martin Reinhardt Nielsen, Jette Bredahl Jacobsen, and Bo Jellesmark Thorsen. Factors determining the choice of hunting and trading bushmeat in the Kilombero Valley, Tanzania. Conservation Biology, 28(2):382–391, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nentwig:2010:GIS

[NKB10] Wolfgang Nentwig, Elfi Kühnel, and Sven Bacher. A generic impact-scoring system applied to alien mammals in Europe. Un sistema Genérico de Clasificación de impacto aplicado a mamíferos introducidas en europa. Conservation Biology, 26 (2):302–311, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nimmo:2013:FMR

[NKSB^{+}13] D. G. Nimmo, L. T. Kelly, L. M. Spence-Bailey, S. J. Watson, R. S. Taylor, M. F. Clarke, and A. F. Bennett. Fire mosaics and reptile conservation in a fire-prone region. Conservation Biology, 27(2):345–353, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nunez:2013:CPA

[NLM^{+}13] Tristan A. Nuñez, Joshua J. Lawler, Brad H. Mcrae, D. John Pierce, Meade B. Krosby, Darren M. Kavanagh, Peter H. Singleton, and Joshua J. Tewksbury. Connectivity planning to address climate change. Conservation Biology, 27(2):407–416, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Narain:2018:CSO

[NM18] Divya Narain and Martine Maron. Cost shifting and other perverse incentives in biodiversity offsetting in India. Conservation Biology, 32(4):782–788, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ng:2014:DWC

[NMMP14] Chooi Fei Ng, Michael A. Mccarthy, Tara G. Martin, and Hugh P. Possingham. Determining when to change course in management actions. Conservation Biology, 28(6):1617–1625, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nekaris:2018:NAC

[NMS^{+}18] K. A. I. Nekaris, Sharon McCabe, Denise Spaan, Muhammad Imron Ali, and Vincent Nijman. A novel application of cultural consensus models to evaluate conservation education programs. Conservation Biology, 32(2):466–476, April 2018.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nepstad:2011:SCR

[NMSF11] Daniel C. Nepstad, David G. Mcgrath, and Britaldo Soares-Filho. Systemic conservation, REDD, and the future of the Amazon Basin. Conservation Biology, 25(6):1113–1116, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Novotny:2016:ECA

[Nov16] Vojtech Novotny. Ecology, conservation, and adventure of the tropical forests in Asia and New Guinea. Conservation Biology, 30(6):679–681, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nel:2016:KCP

[NRD^{+}16] Jeanne L. Nel, Dirk J. Roux, Amanda Driver, Liesl Hill, Ashton C. Maherry, Kate Snaddon, Chantel R. Petersen, Lindie B. Smith-Adao, Heidi Van Deventer, and Belinda Reyers. Knowledge co-production and boundary work to promote implementation of conservation plans. Conservation Biology, 30(1): 176–188, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Novoa:2017:LET

[NRRW17] Ana Novoa, Johannes J. Le Roux, David M. Richardson, and John R. U. Wilson. Level of environmental threat posed by horticultural trade in cactaceae. Conservation Biology, 31(5): 1066–1075, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Najera:2010:EAC

[NS10] Andrea Nájera and Javier A. Simonetti. Enhancing avifauna in commercial plantations. Enriquecimiento de la avifauna en plantaciones comerciales. Conservation Biology, 26(2):319–324, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nie:2012:DMT

[NS12] Martin A. Nie and Courtney A. Schultz. Decision-making triggers in adaptive management. Conservation Biology, 26(6):

1137–1144, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Novotny:2018:RCA

[NS18] Vojtech Novotny and Katerina Sam. Research and conservation on an artificial island in a shipping lane and elsewhere in the tropics. Conservation Biology, 32(3):750–752, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nielsen:2017:EPH

[NSBJ17a] Anne Sofie Elberg Nielsen, Niels Strange, Hans Henrik Bruun, and Jette Bredahl Jacobsen. Effects of preference heterogeneity among landowners on spatial conservation prioritization. Conservation Biology, 31(3):675–685, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nielsen:2017:PHA

[NSBJ17b] Anne Sofie Elberg Nielsen, Niels Strange, Hans Henrik Bruun, and Jette Bredahl Jacobsen. Preference heterogeneity among landowners and spatial conservation prioritization: response to Eyvindson et al. 2017. Conservation Biology, 31(6):1491–1493, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nuttman:2015:CM

[Nut15] Clive Nuttman. Conservation in Madagascar. Conservation Biology, 29(5):1490–1492, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Norton:2015:PBO

[NW15] David A. Norton and Bruce Warburton. The potential for biodiversity offsetting to fund effective invasive species control. Conservation Biology, 29(1):5–11, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Niemiec:2019:MLR

[NWAB19] R. M. Niemiec, R. Willer, N. M. Ardoin, and F. K. Brewer. Motivating landowners to recruit neighbors for private land conservation. Conservation Biology, 33(4):930–941, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Naidoo:2016:CBT

[NWD^{+}16a] Robin Naidoo, L. Chris Weaver, Richard W. Diggle, Greenwell Matongo, Greg Stuart-Hill, and Chris Thouless. Complementary benefits of tourism and hunting to communal conservancies in Namibia. Conservation Biology, 31(1):628–638, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Naidoo:2016:ILV

[NWD^{+}16b] Robin Naidoo, L. Chris Weaver, Richard W. Diggle, Greenwell Matongo, Greg Stuart-Hill, and Chris Thouless. Importance of local values to successful conservation: response to Jacquet and Delon. Conservation Biology, 30(4):912–913, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nowakowski:2017:TAS

[NWW^{+}17] A. Justin Nowakowski, James I. Watling, Steven M. Whitfield, Brian D. Todd, David J. Kurz, and Maureen A. Donnelly. Tropical amphibians in shifting thermal landscapes under landuse and climate change. Conservation Biology, 31(1):96–105, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nguyen:2017:RKE

[NYC17] Vivian M. Nguyen, Nathan Young, and Steven J. Cooke. A roadmap for knowledge exchange and mobilization research in conservation and natural resource management. Conservation Biology, 31(4):789–798, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

OBrien:2011:RUP

[O’B11a] Paul O’Brien. Reaching the U.S. public through their patriotism, pastors, and pockets. Conservation Biology, 25(6):1108– 1111, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ohmer:2011:CRP

[OB11b] Michel E. Ohmer and Phillip J. Bishop. Citation rate and perceived subject bias in the amphibian-decline literature. Tasa de citación y percepción de sesgo temático en la literatura sobre declinación de anfibios. Conservation Biology, 25(1):195–199, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Osterblom:2012:GCA

[ÖB12] Henrik Österblom and Örjan Bodin. Global cooperation among diverse organizations to reduce illegal fishing in the Southern Ocean. Cooperación global entre organizaciones diversas para reducir la pesca ilegal en el Océano del Sur. Conservation Biology, 26(4):638–648, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oppel:2011:EIM

[OBB^{+}11] Steffen Oppel, Brent M. Beaven, Mark Bolton, Juliet Vickery, and Thomas W. Bodey. Eradication of invasive mammals on islands inhabited by humans and domestic animals. Erradicación de mamíferos invasores en islas habitadas por humanos y animales domésticos. Conservation Biology, 25(2):232–240, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oldekop:2010:ULL

[OBBP10] Johan A. Oldekop, Anthony J. Bebbington, Dan Brockington, and Richard F. Preziosi. Understanding the lessons and limitations of conservation and development. Entendiendo las lecciones y limitaciones de la conservación y el desarrollo. Conservation Biology, 24(2):461–469, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Olds:2014:ISS

[OCP^{+}14] Andrew D. Olds, Rod M. Connolly, Kylie A. Pitt, Paul S. Maxwell, Shankar Aswani, and Simon Albert. Incorporating surrogate species and seascape connectivity to improve marine conservation outcomes. Conservation Biology, 28(4):982–991, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oommen:2019:FFC

[OCR^{+}19] Meera Anna Oommen, Rosie Cooney, Madhuri Ramesh, Michael Archer, Daniel Brockington, Bram Buscher, Robert Fletcher, Daniel J. D. Natusch, Abi T. Vanak, Grahame Webb, and Kartik Shanker. The fatal flaws of compassionate conservation. Conservation Biology, 33(4):784–787, August 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

ODea:2017:LPO

[ODAL17] Aaron O’Dea, Erin M. Dillon, Andrew H. Altieri, and Mauro L. Lepore. Look to the past for an optimistic future. Conservation Biology, 31(6):1221–1222, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ormerod:2010:PWI

[ODTS10] S. J. Ormerod, Isabelle Durance, Aurelie Terrier, and Alisa M. Swanson. Priority wetland invertebrates as conservation surrogates. Invertebrados de humedal prioritarios como sustitutos de conservación. Conservation Biology, 24(2):573–582, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ortodossi:2019:ESC

[OGS^{+}19] Nicholas L. Ortodossi, Ben L. Gilby, Thomas A. Schlacher, Rod M. Connolly, Nicholas A. Yabsley, Christopher J. Henderson, and Andrew D. Olds. Effects of seascape connectivity on reserve performance along exposed coastlines. Conservation Biology, 33(3):580–589, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oldekop:2016:GAS

[OHHE16] J. A. Oldekop, G. Holmes, W. E. Harris, and K. L. Evans. A global assessment of the social and conservation outcomes of protected areas. Conservation Biology, 30(1):133–141, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ozerov:2015:USR

[OJAV15] Mikhail Ozerov, Tauno Jürgenstein, Tutku Aykanat, and Anti Vasemägi. Use of sibling relationship reconstruction to complement traditional monitoring in fisheries management and conservation of brown trout. Conservation Biology, 29(4):1164– 1175, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ostfeld:2010:PDL

[OK10] Richard S. Ostfeld and Felicia Keesing. Parasitology is dead. Long live parasitology! Conservation Biology, 24(6):1690–1691, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oyanedel:2018:IFT

[OKCG18] Rodrigo Oyanedel, Andres Keim, Juan Carlos Castilla, and Stefan Gelcich. Illegal fishing and territorial user rights in Chile. Conservation Biology, 32(3):619–627, June 2018. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Olden:2011:COI

[OKLP11] Julian D. Olden, Mark J. Kennard, Joshua J. Lawler, and N. Leroy Poff. Challenges and opportunities in implementing managed relocation for conservation of freshwater species. Retos y oportunidades para la implementación de la Reubicación controlada para la conservación de especies dulceacuícolas. Conservation Biology, 25(1):40–47, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ocampo-Penuela:2014:SPC

[OPP14] Natalia Ocampo-Peñuela and Stuart L. Pimm. Setting practical conservation priorities for birds in the Western Andes of Colombia. Conservation Biology, 28(6):1260–1270, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ochoa-Quintero:2015:TSL

[OQGR^{+}15] Jose Manuel Ochoa-Quintero, Toby A. Gardner, Isabel Rosa, Silvio Frosini de Barros Ferraz, and William J. Sutherland. Thresholds of species loss in Amazonian deforestation frontier landscapes. Conservation Biology, 29(2):440–451, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ordenana:2011:CLE

[Ord11] Miguel A. Ordeñana. Carnivores living on the edge. Conservation Biology, 27(3):1062–1063, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oren:2012:SAC

[Ore12] David C. Oren. A single avian conservation theme, three distinct flight paths. Conservation Biology, 26(2):192–194, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

ODonnell:2010:HCB

[OSC10] Ryan P. O’Donnell, Sarah R. Supp, and Stephanie M. Cobbold. Hindrance of conservation biology by delays in the submission of manuscripts. Contratiempos de la Biología de la conservación por demoras en el envío de manuscritos. Conservation Biology, 24(2):615–620, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oliveira-Santos:2010:PRF

[OSF10] Luiz G. R. Oliveira-Santos and Fernando A. S. Fernandez. Pleistocene rewilding, Frankenstein ecosystems, and an alternative conservation agenda. Conservation Biology, 24(1):4–5, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oliveira-Santos:2011:RRR

[OSF11] Luiz G. R. Oliveira-Santos and Fernando A. S. Fernandez. Reintroduction and refaunation: Response to Seddon et al. Conservation Biology, 25(2):213, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ortiz-Santaliestra:2011:AUB

[OSFFB^{+}11] Manuel E. Ortiz-Santaliestra, Matthew C. Fisher, Saioa Fernández-Beaskoetxea, María J. Fernández-Benéitez, and Jaime Bosch. Ambient ultraviolet B radiation and prevalence of infection by Batrachochytrium dendrobatidis in two amphibian species. Radiación ultravioleta B ambiental y prevalencia de infección por Batrachochytrium dendrobatidis en dos especies de anfibios. Conservation Biology, 27(3):975–982, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oh:2017:POJ

[OSM^{+}17] Beverly Z. L. Oh, Ana M. M. Sequeira, Mark G. Meekan, Jonathan L. W. Ruppert, and Jessica J. Meeuwig. Predicting occurrence of juvenile shark habitat to improve conservation planning. Conservation Biology, 31(3):635–645, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Osofsky:2015:MCP

[Oso15] Steven A. Osofsky. Misrepresentation by citation in Pooley et al. Conservation Biology, 29(4):1037–1038, August 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [PFN15b].

Osterblom:2014:CFC

[Öst14] Henrik Österblom. Catching up on fisheries crime. Conservation Biology, 28(5):877–879, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ogada:2012:EVD

[OTKE12] D. L. Ogada, M. E. Torchin, M. F. Kinnaird, and V. O. Ezenwa. Effects of vulture declines on facultative scavengers and potential implications for mammalian disease transmission. Efectos de la declinación de buitres sobre Carroñeros facultativos e implicaciones potenciales para la transmisión de enfermedades a mamíferos. Conservation Biology, 26(3):453–460, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parlato:2012:IAP

[PA12] Elizabeth H. Parlato and Doug P. Armstrong. An integrated approach for predicting fates of reintroductions with demographic data from multiple populations. Un método integral para la predicción del destino de reintroducciones con datos demográficos de múltiples poblaciones. Conservation Biology, 26(2):97–106, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parlato:2018:PRO

[PA18] Elizabeth H. Parlato and Doug P. Armstrong. Predicting reintroduction outcomes for highly vulnerable species that do not currently coexist with their key threats. Conservation Biology, 32(6):1346–1355, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pearson:2010:THB

[PAM^{+}10] David L. Pearson, Corey Devin Anderson, Brian R. Mitchell, Michael S. Rosenberg, Ronald Navarrete, and Paul Coopmans. Testing hypotheses of bird extinctions at Rio Palenque, Ecuador, with informal species lists. Comprobación de hipótesis sobre las extinciones de aves en Río Palenque, Ecuador mediante listas informales de especies. Conservation Biology, 24 (2):500–510, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Palacin:2017:CBM

[PAMA17] Carlos Palacín, Juan C. Alonso, Carlos A. Martín, and Javier A. Alonso. Changes in bird-migration patterns associated with human-induced mortality. Conservation Biology, 31(1):106–115, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parkes:2010:TVE

[Par10] John Parkes. Two views on the ecology and economics of intervention against invasive species. Conservation Biology, 24(5): 1431–1432, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pautasso:2016:BBC

[Pau16] Marco Pautasso. Biosecurity, bibliometrics, and conservation biology. Conservation Biology, 30(2):440–442, April 2016. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phillips:2011:EMS

[PB11] Julie A. Phillips and Judith K. Blackshaw. Extirpation of macroalgae (Sargassum spp.) on the subtropical East Australian Coast. Extirpación de macroalgas (Sargassum spp.) en la Costa Australiana Oriental subtropical. Conservation Biology, 25(5):913–921, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pio:2011:SPP

[PBB^{+}11] Dorothea V. Pio, Olivier Broennimann, Timothy G. Barraclough, Gail Reeves, Anthony G. Rebelo, Wilfried Thuiller, Antoine Guisan, and Nicolas Salamin. Spatial predictions of phylogenetic diversity in conservation decision making. Predicciones espaciales de la diversidad filogenética en la toma de decisiones de conservación. Conservation Biology, 25(6):1229– 1239, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pierson:2015:IEP

[PBB^{+}15] Jennifer C. Pierson, Steven R. Beissinger, Jason G. Bragg, David J. Coates, J. Gerard B. Oostermeijer, Paul Sunnucks, Nathan H. Schumaker, Meredith V. Trotter, and Andrew G. Young. Incorporating evolutionary processes into population viability models. Conservation Biology, 29(3):755–764, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pooley:2017:IRC

[PBB^{+}17] S. Pooley, M. Barua, W. Beinart, A. Dickman, G. Holmes, J. Lorimer, A. J. Loveridge, D. W. Macdonald, G. Marvin, S. Redpath, C. Sillero-Zubiri, A. Zimmermann, and E. J. Milner-Gulland. An interdisciplinary review of current and future approaches to improving human–predator relations. Conservation Biology, 31(3):513–523, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paillet:2010:BDB

[PBH^{+}10a] Yoan Paillet, Laurent Bergès, Joakim Hjältén, Péter Ódor, Catherine Avon, Markus Bernhardt-Römermann, Rienk-Jan Bijlsma, Luc De Bruyn, Marc Fuhr, Ulf Grandin, Robert Kanka, Lars Lundin, Sandra Luque, Tibor Magura, Silvia Matesanz, Ilona Mészáros, M.-Teresa Sebastià, Wolfgang Schmidt, Tibor Standovár, Béla Tóthmérész, Anneli Uotila, Fernando Valladares, Kai Vellak, and Risto Virtanen. Biodiversity differences between managed and unmanaged forests: Meta-analysis of species richness in Europe. Diferencias en la biodiversidad entre bosques manejados y no manejados: Metaanálisis de la riqueza de especies en Europa. Conservation Biology, 24(1):101–112, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paillet:2010:CDS

[PBH^{+}10b] Yoan Paillet, Laurent Bergès, Joakim Hjältén, Péter Ódor, Catherine Avon, Markus Bernhardt-Römermann, Rienk-Jan Bijlsma, Luc De Bruyn, Marc Fuhr, Ulf Grandin, Robert Kanka, Lars Lundin, Sandra Luque, Tibor Magura, Silvia Matesanz, Ilona Mészáros, M.-Teresa Sebastià, Wolfgang Schmidt, Tibor Standovár, Béla Tóthmérész, Anneli Uotila, Fernando Valladares, Kai Vellak, and Risto Virtanen. Compromises in data selection in a meta-analysis of biodiversity in managed and unmanaged forests: Response to Halme et al. Conservation Biology, 24(4):1157–1160, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Packer:2011:ETH

[PBK^{+}11] C. Packer, H. Brink, B. M. Kissui, H. Maliti, H. Kushnir, and T. Caro. Effects of trophy hunting on lion and leopard populations in Tanzania. Efectos de la cacería deportiva sobre poblaciones de leones y leopardos en tanzania. Conservation Biology, 25(1):142–153, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pedler:2016:RBL

[PBR^{+}16] Reece D. Pedler, Robert Brandle, John L. Read, Richard Southgate, Peter Bird, and Katherine E. Moseby. Rabbit biocontrol and landscape-scale recovery of threatened desert mammals. Conservation Biology, 30(4):774–782, August 2016. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Price:2011:EUO

[PCBD11] Steven J. Price, Kristen K. Cecala, Robert A. Browne, and Michael E. Dorcas. Effects of urbanization on occupancy of stream salamanders. Efectos de la urbanización sobre la ocupación de salamandras de arroyo. Conservation Biology, 25 (6):547–555, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peach:2019:VPA

[PCF^{+}19] Michelle A. Peach, Jonathan B. Cohen, Jacqueline L. Frair, Benjamin Zuckerberg, Patrick Sullivan, William F. Porter, and Corey Lang. Value of protected areas to avian persistence across 20 years of climate and land-use change. Conservation Biology, 33(2):423–433, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Purcell:2014:DCH

[PCRD14] Steven W. Purcell, K. Robert Clarke, Kelvin Rushworth, and Steven J. Dalton. Defining critical habitats of threatened and endemic reef fishes with a multivariate approach. Conservation Biology, 28(6):1688–1698, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pasquini:2010:DAP

[PCTW10] Lorena Pasquini, Richard M. Cowling, Chasca Twyman, and John Wainwright. Devising appropriate policies and instruments in support of private conservation areas: Lessons learned from the Klein Karoo, South Africa. Diseño de políticas e instrumentos adecuados como soporte para las Áreas de conservación privadas: Lecciones aprendidas de los Klein Karoo, África del Sur. Conservation Biology, 24(2):470–478, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phelps:2014:FAS

[PCW14] J. Phelps, L. R. Carrasco, and E. L. Webb. A framework for assessing supply-side wildlife conservation. Conservation Biology, 28(1):244–257, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Painter:2011:IAC

[PDM11] R. L. E. Painter, A. Duran, and E. Miro. Indigenous alliances for conservation in Bolivia. Conservation Biology, 25(6):1084– 1086, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peacock:2018:RPB

[Pea18] Mary Peacock. In remembrance of Peter Brussard. Conservation Biology, 32(6):1476–1477, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pekin:2013:EWA

[Pek13] Burak K. Pekin. Effect of widespread agricultural chemical use on butterfly diversity across Turkish provinces. Conservation Biology, 28(1):1439–1448, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pellikka:2015:DLS

[Pel15] Jani Pellikka. Difficulties of living with success. Conservation Biology, 29(4):1249–1250, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peres:2011:CSU

[Per11] Carlos A. Peres. Conservation in sustainable-use tropical forest reserves. Conservation Biology, 25(6):1124–1129, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peterson:2017:PRP

[Pet17] A. Townsend Peterson. Problems with reductive, polygonbased methods for estimating species’ ranges: reply to Pimm et al. 2017. Conservation Biology, 31(4):948–951, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pritt:2010:QDR

[PF10] Jeremy J. Pritt and Emmanuel A. Frimpong. Quantitative determination of rarity of freshwater fishes and implications for imperiled-species designations. Determinación cuantitativa de la rareza de peces dulceacuícolas e implicaciones para la designación de especies en peligro. Conservation Biology, 24 (5):1249–1258, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parsons:2014:CCC

[PFA^{+}14a] E. C. M. Parsons, Brett Favaro, A. Alonso Aguirre, Amy L. Bauer, Louise K. Blight, John A. Cigliano, Melinda A. Coleman, Isabelle M. Côté, Megan Draheim, Stephen Fletcher, Melissa M. Foley, Rebecca Jefferson, Miranda C. Jones, Brendan P. Kelaher, Carolyn J. Lundquist, Julie-Beth Mccarthy, Anne Nelson, Katheryn Patterson, Leslie Walsh, Andrew J. Wright, and William J. Sutherland. Cover caption: Conservation Biology. Conservation Biology, 28(5):i, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parsons:2014:SOI

[PFA^{+}14b] E. C. M. Parsons, Brett Favaro, A. Alonso Aguirre, Amy L. Bauer, Louise K. Blight, John A. Cigliano, Melinda A. Coleman, Isabelle M. Côté, Megan Draheim, Stephen Fletcher, Melissa M. Foley, Rebecca Jefferson, Miranda C. Jones, Brendan P. Kelaher, Carolyn J. Lundquist, Julie-Beth Mccarthy, Anne Nelson, Katheryn Patterson, Leslie Walsh, Andrew J. Wright, and William J. Sutherland. Seventy-one important questions for the conservation of marine biodiversity. Conservation Biology, 28(5):1206–1214, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pooley:2015:NCS

[PFN15a] Simon Pooley, Julia E. Fa, and Robert Nasi. No conservation silver lining to Ebola. Conservation Biology, 29(3):965–967, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pooley:2015:ROM

[PFN15b] Simon Pooley, Julia E. Fa, and Robert Nasi. Response to Osofsky’s “Misrepresentation by citation”. Conservation Biology, 29(4):1039, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Oso15].

Parker:2019:MCS

[PFN19] Timothy Parker, Hannah Fraser, and Shinichi Nakagawa. Making conservation science more reliable with preregistration and registered reports. Conservation Biology, 33(4):747–750, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pidgeon:2014:STP

[PFR^{+}14] Anna M. Pidgeon, Curtis H. Flather, Volker C. Radeloff, Christopher A. Lepczyk, Nicholas S. Keuler, Eric M. Wood, Susan I. Stewart, and Roger B. Hammer. Systematic temporal patterns in the relationship between housing development and forest bird biodiversity. Conservation Biology, 28(6): 1291–1301, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pandit:2012:PEO

[PG12] Maharaj K. Pandit and R. Edward Grumbine. Potential effects of ongoing and proposed hydropower development on terrestrial biological diversity in the Indian Himalaya. Conservation Biology, 26(6):1061–1071, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pannell:2016:ECU

[PG16] David J. Pannell and Fiona L. Gibson. Environmental cost of using poor decision metrics to prioritize environmental projects. Conservation Biology, 30(2):382–391, April 2016. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Post:2018:ERC

[PG18] Gerald Post and Jonas Geldmann. Exceptional responders in conservation. Conservation Biology, 32(3):576–583, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pietri:2013:PRH

[PGBV^{+}13] Diana M. Pietri, Georgina G. Gurney, Nancy Benitez-Vina, Audrey Kuklok, Sara M. Maxwell, Libby Whiting, Michael A. Vina, and Lekelia D. Jenkins. Practical recommendations to help students bridge the research–implementation gap and promote conservation. Conservation Biology, 28(4):958–967, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Packman:2013:RLC

[PGC^{+}13a] Charlotte E. Packman, Thomas N. E. Gray, Nigel J. Collar, Tom D. Evans, Robert N. Van Zalinge, Son Virak, Andrew A. Lovett, and Paul M. Dolman. Rapid loss of Cambodia’s grasslands. Conservation Biology, 27(2):245–247, April 2013. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phillott:2013:CSM

[PGC^{+}13b] Andrea D. Phillott, Laura F. Grogan, Scott D. Cashins, Keith R. Mcdonald, Lee Berger, and Lee F. Skerratt. Chytridiomycosis and seasonal mortality of tropical stream-associated frogs 15 years after introduction of Batrachochytrium dendrobatidis. Conservation Biology, 28(4):1058–1068, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Palomares:2012:PEV

[PGLB^{+}12] Francisco Palomares, José Antonio Godoy, José Vicente López-Bao, Alejandro Rodríguez, Severine Roques, Mireia Casas-Marce, Eloy Revilla, and Miguel Delibes. Possible extinction vortex for a population of Iberian lynx on the verge of extirpation. Posible vórtice de extinción en una población de lince Ibérico al borde de la extirpación. Conservation Biology, 27 (1):689–697, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Puschendorf:2011:ERD

[PHC^{+}11] Robert Puschendorf, Conrad J. Hoskin, Scott D. Cashins, Keith Mcdonald, Lee F. Skerratt, Jeremy Vanderwal, and Ross A. Alford. Environmental refuge from disease-driven amphibian extinction. Refugio ambiental para la extinción de anfibios causada por enfermedad. Conservation Biology, 25(5): 956–964, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peterson:2010:OEF

[PHFPP10] Markus J. Peterson, Damon M. Hall, Andrea M. Feldpausch-Parker, and Tarla Rai Peterson. Obscuring ecosystem function with application of the ecosystem services concept. Enmascaramiento de la Función del ecosistema con la aplicación del concepto de servicios del ecosistema. Conservation Biology, 24(1):113–119, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phillips:2010:CP

[Phi10] Mike Phillips. Conservation and politics. Conservation Biology, 26(2):353–354, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pimm:2017:UPD

[PHJ^{+}17] Stuart L. Pimm, Grant Harris, Clinton N. Jenkins, Natalia Ocampo-Peñuela, and Binbin V. Li. Unfulfilled promise of data-driven approaches: response to Peterson et al. Conservation Biology, 31(4):944–947, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Petrolia:2011:WEE

[PI11] Daniel R. Petrolia and Matthew G. Interis. What environmental economists think every conservation biologist should know: Reply to Gowdy et al. Conservation Biology, 25(6):628–630, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pinke:2017:BIN

[Pin17] Zsolt Pinke. A biased introduction to the “New environmental history”. Conservation Biology, 31(6):1499–1501, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Predavec:2016:CCW

[PLH^{+}16] Martin Predavec, Daniel Lunney, Ben Hope, Eleanor Stalenberg, Ian Shannon, Mathew S. Crowther, and Indrie Miller. The contribution of community wisdom to conservation ecology. Conservation Biology, 31(1):496–505, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Plotkin:2017:TCA

[Plo17] Mark Plotkin. Two complementary approaches to protecting the Amazon. Conservation Biology, 31(4):956–958, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peer:2013:IOC

[PMD^{+}13] Guy Pe’er, Jeffrey A. McNeely, Martin Dieterich, Bengt-Gunnar Jonsson, Nuria Selva, John M. Fitzgerald, and Carsten Nesshöver. IPBES: Opportunities and challenges for SCB and other learned societies. Conservation Biology, 27(1):1–3, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

PeEr:2013:PBD

[PMJ^{+}13] Guy Pe’Er, Yiannis G. Matsinos, Karin Johst, Kamila W. Franz, Camille Turlure, Viktoriia Radchuk, Agnieszka H. Malinowska, Janelle M. R. Curtis, Ilona Naujokaitis-Lewis, Brendan A. Wintle, and Klaus Henle. A protocol for better design, application, and communication of population viability analyses. Conservation Biology, 27(4):644–656, August 2013. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Previtali:2010:EMF

[PMK^{+}10] M. Andrea Previtali, Peter L. Meserve, Douglas A. Kelt, W. Bryan Milstead, and Julio R. Gutierrez. Effects of more frequent and prolonged El Niño events on life-history parameters of the degu, a long-lived and slow-reproducing rodent. Efectos de eventos El Niño más frecuentes y prolongados sobre Parámetros de historia de vida del degu, un roedor longevo y de reproducción lenta. Conservation Biology, 24(1):18–28, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peck:2014:CEU

[PMM^{+}14] Mika Robert Peck, Simon T. Maddock, Jorge Noe Morales, Hugolino Oñate, Paola Mafla-Endara, Vanessa Aguirre Peñafiel, Omar Torres-Carvajal, Wilmer E. Pozo-Rivera, Xavier A. Cueva-Arroyo, and Bryony A. Tolhurst. Costeffectiveness of using small vertebrates as indicators of disturbance. Conservation Biology, 28(6):1331–1341, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pauli:2015:EUO

[PMM^{+}15] Jonathan N. Pauli, Wynne E. Moss, Philip J. Manlick, Emily D. Fountain, Rebecca Kirby, Sean M. Sultaire, Paula L. Perrig, Jorge E. Mendoza, John W. Pokallus, and Timothy H. Heaton. Examining the uncertain origin and management role of martens on Prince of Wales Island, Alaska. Conservation Biology, 29(5):1257–1267, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pooley:2014:HCM

[PMMG14] Simon P. Pooley, J. Andrew Mendelsohn, and E. J. Milner-Gulland. Hunting down the Chimera of multiple disciplinarity in conservation science. Conservation Biology, 28(1):22–32, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parker:2016:PTC

[PMN^{+}16] T. H. Parker, E. Main, S. Nakagawa, J. Gurevitch, F. Jarrad, and M. Burgman. Promoting transparency in conservation science. Conservation Biology, 30(6):1149–1150, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Proulx:2014:GTC

[PMP14] Raphaël Proulx, Philippe Massicotte, and Marc Pépino. Googling trends in conservation biology. Conservation Biology, 28(1):44–51, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pilliod:2010:EAC

[PMS^{+}10] David S. Pilliod, Erin Muths, Rick D. Scherer, Paul E. Bartelt, Paul Stephen Corn, Blake R. Hossack, Brad A. Lambert, Rebecca Mccaffery, and Christopher Gaughan. Effects of amphibian chytrid fungus on individual survival probability in wild boreal toads efectos del hongo quitridio anfibio sobre la probabilidad de supervivencia individual en sapos boreales silvestres. Conservation Biology, 24(5):1259–1267, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parks:2013:EWS

[PMS13] Sean A. Parks, Kevin S. Mckelvey, and Michael K. Schwartz. Effects of weighting schemes on the identification of wildlife corridors generated with least-cost methods. Conservation Biology, 27(1):145–154, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Papworth:2015:QRO

[PNC^{+}15] S. K. Papworth, T. P. L. Nghiem, D. Chimalakonda, M. R. C. Posa, L. S. Wijedasa, D. Bickford, and L. R. Carrasco. Quantifying the role of online news in linking conservation research to Facebook and Twitter. Conservation Biology, 29(3):825– 833, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peterson:2018:AVD

[PNSG18] A. Townsend Peterson, Adolfo G. Navarro-Sigüenza, and Alejandro Gordillo. Assumption-versus data-based approaches to summarizing species’ ranges. Conservation Biology, 32(3):568– 575, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pooley:2013:HVE

[Poo13] Simon Pooley. Historians are from Venus, ecologists are from Mars. Conservation Biology, 28(1):1481–1483, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pozsgai:2017:OEB

[Poz17] Gabor Pozsgai. One for the expert, and a better one for the layperson. Conservation Biology, 31(5):1209–1211, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pozsgai:2018:CHW

[Poz18] Gabor Pozsgai. Conservation and human–wildlife conflicts on farmland. Conservation Biology, 32(1):250–252, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Porszt:2012:RID

[PPD^{+}12] Erin J. Porszt, Randall M. Peterman, Nicholas K. Dulvy, Andrew B. Cooper, and James R. Irvine. Reliability of indicators of decline in abundance. Confiabilidad de indicadores de declinación de abundancia. Conservation Biology, 26(5):894–904, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parker:2018:ABM

[PPM^{+}18] Sophie S. Parker, Gregory B. Pauly, James Moore, Naomi S. Fraga, John J. Knapp, Zachary Principe, Brian V. Brown, John M. Randall, Brian S. Cohen, and Thomas A. Wake. Adapting the bioblitz to meet conservation needs. Conservation Biology, 32(5):1007–1019, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pons:2017:CCB

[PR17] Pere Pons and Josep Rost. The challenge of conserving biodiversity in harvested burned forests. Conservation Biology, 31 (1):226–228, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Prasad:2010:EEP

[Pra10] Ayesha E. Prasad. Effects of an exotic plant invasion on native understory plants in a tropical dry forest. Efectos de la invasión de una planta exótica sobre plantas nativas en el sotobosque de un bosque tropical seco. Conservation Biology, 24(3):747–757, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pedler:2018:LDF

[PRB18] Reece D. Pedler, Raoul F. H. Ribot, and Andrew T. D. Bennett. Long-distance flights and high-risk breeding by nomadic waterbirds on desert salt lakes. Conservation Biology, 32(1): 216–228, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Payet:2013:ELC

[PRE^{+}13] Karine Payet, Mathieu Rouget, Karen J. Esler, Belinda Reyers, Tony Rebelo, Mark W. Thompson, and Jan H. J. Vlok. Effect of land cover and ecosystem mapping on ecosystem-risk assessment in the Little Karoo, South Africa. Conservation Biology, 27(3):531–541, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Prior:2010:EPS

[Pri10] Kent Prior. Extirpations from parks: Scaling conservation planning to fit the problem. Conservation Biology, 24(3):646– 648, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pritchard:2011:CCC

[Pri11] Lowell Pritchard. Cultivating a constituency for conservation. Conservation Biology, 25(6):1103–1107, December 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pritchard:2012:PO

[Pri12] Sara B. Pritchard. The politics of opting out. Conservation Biology, 26(3):382–383, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Price:2019:IDD

[Pri19] Trevor Price. India’s dying for development. Conservation Biology, 33(2):483–485, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Poulsen:2018:ECF

[PRM^{+}18] John R. Poulsen, Cooper Rosin, Amelia Meier, Emily Mills, Chase L. Nuñez, Sally E. Koerner, Emily Blanchard, Jennifer Callejas, Sarah Moore, and Mark Sowers. Ecological consequences of forest elephant declines for afrotropical forests. Conservation Biology, 32(3):559–567, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Popejoy:2018:PSC

[PRNV18] Traci Popejoy, Charles R. Randklev, Thomas M. Neeson, and Caryn C. Vaughn. Prioritizing sites for conservation based on similarity to historical baselines and feasibility of protection.

Conservation Biology, 32(5):1118–1127, October 2018. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Palomares:2011:ACE

[PRR^{+}11] Francisco Palomares, Alejandro Rodríguez, Eloy Revilla, José Vicente López-Bao, and Javier Calzada. Assessment of the conservation efforts to prevent extinction of the Iberian lynx. Evaluación de los esfuerzos de conservación para prevenir la extinción del lince Ibérico. Conservation Biology, 25(1):4–8, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Povilitis:2010:ACC

[PS10a] Anthony Povilitis and Kierán Suckling. Addressing climate change threats to endangered species in U.S. recovery plans. Conservation Biology, 24(2):372–376, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pullin:2010:SWS

[PS10b] Andrew S. Pullin and Nick Salafsky. Save the whales? Save the rainforest? Save the data! Conservation Biology, 24(4):915– 917, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Polak:2011:RER

[PS11] Tal Polak and David Saltz. Reintroduction as an ecosystem restoration technique. Conservation Biology, 25(3):424, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peh:2013:ALE

[PS13] Kelvin S.-H. Peh and Franziska Schrodt. Application of lessons from the Euro crisis to climate change. Conservation Biology, 27(3):439–440, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Penaranda:2015:PSC

[PS15] Diego A. Peñaranda and Javier A. Simonetti. Predicting and setting conservation priorities for Bolivian mammals based on biological correlates of the risk of decline. Conservation Biology, 29(3):834–843, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [TYA16].

Penaranda:2016:INU

[PS16] Diego A. Peñaranda and Javier A. Simonetti. The informative nature of unexpected results: reply to Tellería and Yapu– Alcázar. Conservation Biology, 30(4):907–909, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pyke:2018:CRW

[PS18] Graham H. Pyke and Judit K. Szabo. Conservation and the 4 Rs, which are rescue, rehabilitation, release, and research. Conservation Biology, 32(1):50–59, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Prescott:2017:PTE

[PSA^{+}17] Graham W. Prescott, William J. Sutherland, Daniel Aguirre, Matthew Baird, Vicky Bowman, Jake Brunner, Grant M. Connette, Martin Cosier, David Dapice, Jose Don T. De Alban, Alex Diment, Julia Fogerite, Jefferson Fox, Win Hlaing, Saw Htun, Jack Hurd, Katherine LaJeunesse Connette, Felicia Lasmana, Cheng Ling Lim, Antony Lynam, Aye Chan Maung, Benjamin McCarron, John F. McCarthy, William J. McShea, Frank Momberg, Myat Su Mon, Than Myint, Robert Oberndorf, Thaung Naing Oo, Jacob Phelps, Madhu Rao, Dietrich Schmidt-Vogt, Hugh Speechly, Oliver Springate-Baginski, Robert Steinmetz, Kirk Talbott, Maung Maung Than, Tint Lwin Thaung, Salai Cung Lian Thawng, Kyaw Min Thein, Shwe Thein, Robert Tizard, Tony Whitten, Guy Williams, Trevor Wilson, Kevin Woods, Alan D. Ziegler, Michal Zrust, and Edward L. Webb. Political transition and emergent forest-conservation issues in Myanmar. Conservation Biology, 31(6):1257–1270, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Papes:2011:ISR

[PSAZ11] M. Papeş, M. Sällström, T. R. Asplund, and M. J. Vander Zanden. Invasive species research to meet the needs of resource management and planning. Investigación de especies invasoras para satisfacer las necesidades del manejo y planificación de recursos. Conservation Biology, 25(5):867–872, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parsons:2014:HTL

[PSD^{+}14] E. C. M. Parsons, D. S. Shiffman, E. S. Darling, N. Spillman, and A. J. Wright. How Twitter literacy can benefit conservation scientists. Conservation Biology, 28(2):299–301, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Patten:2011:POT

[PSP11] Michael A. Patten and Brenda D. Smith-Patten. Predictors of occupancy trend across spatial scale. Predictores de tendencias de ocupación en varias escalas espaciales. Conservation Biology, 25(6):1203–1211, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pilotto:2019:DRP

[PTJ^{+}19] Francesca Pilotto, Jonathan D. Tonkin, Kathrin Januschke, Armin W. Lorenz, Jonas Jourdan, Andrea Sundermann, Daniel Hering, Stefan Stoll, and Peter Haase. Diverging response patterns of terrestrial and aquatic species to hydromorphological restoration. Conservation Biology, 33(2):132–141, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paltto:2010:MMC

[PTN10] Heidi Paltto, Ingrid Thomasson, and Björn Nordén. Multispecies and multiscale conservation planning: Setting quantitative targets for red-listed lichens on ancient oaks. Planificación de la conservación para múltiples especies y escalas: Definición de objetivos cuantitativos para Líquenes sobre encinos antiguos en la lista roja. Conservation Biology, 24(3):758–768, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Putz:2011:BVT

[Put11] Francis E. Putz. Broad view of tropical forests. Conservation Biology, 25(4):849–850, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Prager:2016:AAB

[PVO^{+}16] C. M. Prager, A. Varga, P. Olmsted, J. C. Ingram, M. Cattau, C. Freund, R. Wynn-Grant, and S. Naeem. An assessment of adherence to basic ecological principles by payments for ecosystem service projects. Conservation Biology, 30(4): 836–845, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Penone:2013:ULS

[PVP^{+}13] Caterina Penone, Isabelle Le Viol, Vincent Pellissier, Jean-François Julien, Yves Bas, and Christian Kerbiriou. Use of large-scale acoustic monitoring to assess anthropogenic pressures on orthoptera communities. Conservation Biology, 28(4): 979–987, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Purahong:2019:ANG

[PWKB19] Witoon Purahong, Tesfaye Wubet, Dirk Krüger, and François Buscot. Application of next-generation sequencing technologies to conservation of wood-inhabiting fungi. Conservation Biology, 33(3):716–724, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Piccolo:2018:WCS

[PWKT18] John J. Piccolo, Haydn Washington, Helen Kopnina, and Bron Taylor. Why conservation scientists should re-embrace their ecocentric roots. Conservation Biology, 32(4):959–961, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pitt:2018:WCE

[PWP^{+}18] R. Pitt, C. Wyborn, G. Page, J. Hutton, M. Virah Sawmy, M. Ryan, and L. Gallagher. Wrestling with the complexity of evaluation for organizations at the boundary of science, policy, and practice. Conservation Biology, 32(5):998–1006, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pauli:2010:DNA

[PWRM10] Jonathan N. Pauli, John P. Whiteman, Meghan D. Riley, and Arthur D. Middleton. Defining noninvasive approaches for sampling of vertebrates. Conservation Biology, 26(2):349–352, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Porensky:2013:EEI

[PY13] Lauren M. Porensky and Truman P. Young. Edge-effect interactions in fragmented and patchy landscapes. Conservation Biology, 27(3):509–519, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pykala:2017:RBE

[Pyk17] Juha Pykälä. Relation between extinction and assisted colonization of plants in the arctic-alpine and boreal regions. Conservation Biology, 31(3):524–530, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pereira:2014:CSA

[PZM14] Henrique Miguel Pereira, Guy Ziv, and Murilo Miranda. Countryside species–area relationship as a valid alternative to the matrix-calibrated species–area model. Conservation Biology, 28(5):874–876, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quilodran:2018:EHG

[QCMB18] Claudio S. Quilodrán, Mathias Currat, and Juan I. Montoya-Burgos. Effect of hybridization with genome exclusion on extinction risk. Conservation Biology, 32(5):1139–1149, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quan:2018:EPA

[QCW^{+}18] Qing Quan, Xianli Che, Yongjie Wu, Yuchun Wu, Qiang Zhang, Min Zhang, and Fasheng Zou. Effectiveness of protected areas for vertebrates based on taxonomic and phylogenetic diversity. Conservation Biology, 32(2):355–365, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Qin:2019:PAD

[QKC^{+}19] Siyu Qin, Rachel E. Golden Kroner, Carly Cook, Anteneh T. Tesfaw, Rowan Braybrook, Carlos Manuel Rodriguez, Claire Poelking, and Michael B. Mascia. Protected area downgrading, downsizing, and degazettement as a threat to iconic protected areas. Conservation Biology, 33(6):1275–1285, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quintero:2011:BOI

[QM11] Juan David Quintero and Aradhna Mathur. Biodiversity offsets and infrastructure. Conservation Biology, 25(6):1121–1123, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quinzin:2019:GIC

[QSCM^{+}19] Maud C. Quinzin, Jonathan Sandoval-Castillo, Joshua M. Miller, Luciano B. Beheregaray, Michael A. Russello, Elizabeth A. Hunter, James P. Gibbs, Washington Tapia, Freddy Villalva, and Adalgisa Caccone. Genetically informed captive breeding of hybrids of an extinct species of Galapagos tortoise. Conservation Biology, 33(6):1404–1414, December 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quinlan:2014:FFC

[Qui14] Allyson Quinlan. Frames and fences: Combining approaches to understanding and managing resilient cultural landscapes. Conservation Biology, 28(1):293–296, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2016:HB

[RAP16] Kent H. Redford, Keith Aune, and Glenn Plumb. Hope is a bison. Conservation Biology, 30(4):689–691, August 2016. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Raymond:2014:I

[Ray14] Christopher M. Raymond. Introduction. Conservation Biology, 28(6):1447–1450, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rai:2013:IPH

[RB13] Nitin D. Rai and Kamaljit S. Bawa. Inserting politics and history in conservation. Conservation Biology, 27(2):425–428, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rogers:2017:SPD

[RB17] Abbie A. Rogers and Michael P. Burton. Social preferences for the design of biodiversity offsets for shorebirds in Australia. Conservation Biology, 31(4):828–836, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rudd:2011:GPR

[RBC^{+}11] Murray A. Rudd, Karen F. Beazley, Steven J. Cooke, Erica Fleishman, Daniel E. Lane, Michael B. Mascia, Robin Roth, Gary Tabor, Jiselle A. Bakker, Teresa Bellefontaine, Dominique Berteaux, Bernard Cantin, Keith G. Chaulk, Kathryn Cunningham, Rod Dobell, Eleanor Fast, Nadia Ferrara, C. Scott Findlay, Lars K. Hallstrom, Thomas Hammond, Luise Hermanutz, Jeffrey A. Hutchings, Kathryn E. Lindsay, Tim J. Marta, Vivian M. Nguyen, Greg Northey, Kent Prior, Saudiel Ramirez-Sanchez, Jake Rice, Darren J. H. Sleep, Nora D. Szabo, Geneviève Trottier, Jean-Patrick Toussaint, and Jean-Philippe Veilleux. Generation of priority research questions to inform conservation policy and management at a national level. Generación de preguntas de investigación prioritarias para informar a las Políticas y gestión de la conservación a nivel nacional. Conservation Biology, 25(3):476–484, June 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ritchie:2013:CSG

[RBD^{+}13] Euan G. Ritchie, Corey J. A. Bradshaw, Chris R. Dickman, Richard Hobbs, Christopher N. Johnson, Emma L. Johnston, William F. Laurance, David Lindenmayer, Michael A. Mc-Carthy, Dale G. Nimmo, Hugh H. Possingham, Robert L. Pressey, David M. Watson, and John Woinarski. Continentalscale governance and the hastening of loss of Australia’s biodiversity. Conservation Biology, 27(6):1133–1135, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rose:2015:TWR

[RBE^{+}15] Robert A. Rose, Dirck Byler, J. Ron Eastman, Erica Fleishman, Gary Geller, Scott Goetz, Liane Guild, Healy Hamilton, Matt Hansen, Rachel Headley, Jennifer Hewson, Ned Horning, Beth A. Kaplin, Nadine Laporte, Allison Leidner, Peter Leimgruber, Jeffrey Morisette, John Musinsky, Lilian Pintea, Ana Prados, Volker C. Radeloff, Mary Rowen, Sassan Saatchi, Steve Schill, Karyn Tabor, Woody Turner, Anthony Vodacek, James Vogelmann, Martin Wegmann, David Wilkie, and Cara Wilson. Ten ways remote sensing can contribute to conservation. Conservation Biology, 29(2):350–359, April 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rizzari:2015:ICA

[RBF15] Justin R. Rizzari, Brock J. Bergseth, and Ashley J. Frisch. Impact of conservation areas on trophic interactions between apex predators and herbivores on coral reefs. Conservation Biology, 29(2):418–429, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Regnier:2015:EHE

[RBH^{+}15] Claire Régnier, Philippe Bouchet, Kenneth A. Hayes, Norine W. Yeung, Carl C. Christensen, Daniel J. D. Chung, Benoît Fontaine, and Robert H. Cowie. Extinction in a hyperdiverse endemic Hawaiian land snail family and implications for the underestimation of invertebrate extinction. Conservation Biology, 29(6):1715–1723, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rout:2018:MID

[RBHW18] Tracy M. Rout, Christopher M. Baker, Stewart Huxtable, and Brendan A. Wintle. Monitoring, imperfect detection, and risk optimization of a Tasmanian devil insurance population. Conservation Biology, 32(2):267–275, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Regan:2017:IDP

[RBK^{+}17] Helen M. Regan, Clara I. Bohórquez, David A. Keith, Tracey J. Regan, and Kurt E. Anderson. Implications of different population model structures for management of threatened plants. Conservation Biology, 31(3):459–468, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Root-Bernstein:2010:CDC

[RBL10] Meredith Root-Bernstein and Richard J. Ladle. Conservation by design. Conservación por diseño. Conservation Biology, 24 (5):1205–1211, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rico:2014:ERS

[RBW14] Yessica Rico, Hans Juergen Boehmer, and Helene H. Wagner. Effect of rotational shepherding on demographic and genetic connectivity of calcareous grassland plants. Conservation Biology, 28(2):467–477, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robles:2012:IHQ

[RC12] Hugo Robles and Carlos Ciudad. Influence of habitat quality, population size, patch size, and connectivity on patchoccupancy dynamics of the middle spotted woodpecker. Influencia de la calidad del hábitat, el tamaño poblacional, tamaño del parche y la conectividad sobre la dinámica de ocupación de parches de Dendrocopos medius. Conservation Biology, 26 (2):284–293, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rueda-Cediel:2018:EUV

[RCARR18] Pamela Rueda-Cediel, Kurt E. Anderson, Tracey J. Regan, and Helen M. Regan. Effects of uncertainty and variability on population declines and IUCN red list classifications. Conservation Biology, 32(4):916–925, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roll:2018:UML

[RCBT18] Uri Roll, Ricardo A. Correia, and Oded Berger-Tal. Using machine learning to disentangle homonyms in large text corpora. Conservation Biology, 32(3):716–724, June 2018. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Regnery:2013:OCS

[RCK13] Baptiste Regnery, Denis Couvet, and Christian Kerbiriou. Offsets and conservation of the species of the EU habitats and birds directives. Conservation Biology, 27(6):1335–1343, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodriguez:2012:BSB

[RCR^{+}12] Alejandro Rodríguez, Javier Calzada, Eloy Revilla, José Vicente López-Bao, and Francisco Palomares. Bringing science back to the conservation of the Iberian lynx. Conservation Biology, 27(1):737–739, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robillard:2015:FCC

[RCSK15] Cassandra M. Robillard, Laura E. Coristine, Rosana N. Soares, and Jeremy T. Kerr. Facilitating climate-change-induced range shifts across continental land-use barriers. Conservation Biology, 29(6):1586–1595, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redding:2010:EDT

[RDM10] David W. Redding, Curt V. Dewolff, and Arne Ø. Mooers. Evolutionary distinctiveness, threat status, and ecological oddity in primates. Singularidad evolutiva, estatus de amenaza y rareza ecológica en primates. Conservation Biology, 24(4):1052–1058, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ruhi:2017:IBD

[RDS17] Albert Ruhí, Thibault Datry, and John L. Sabo. Interpreting beta-diversity components over time to conserve metacommunities in highly dynamic ecosystems. Conservation Biology, 31 (6):1459–1468, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2010:NDB

[Red10a] Kent H. Redford. New directions for book reviews in Conservation Biology. Conservation Biology, 26(2):353, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2010:NI

[Red10b] Kent H. Redford. Noted with interest. Conservation Biology, 26(2):353, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2012:DWC

[Red12] Kent H. Redford. Deciding what to conserve. Conservation Biology, 26(5):948–949, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roberts:2010:IAR

[RER10] David L. Roberts, Chris S. Elphick, and J. Michael Reed. Identifying anomalous reports of putatively extinct species and why it matters. Identificación de reportes anómalos de especies putativamente extintas y porqué es importante. Conservation Biology, 24(1):189–196, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reyers:2011:EPB

[Rey11] Belinda Reyers. An emerging partnership between biodiversity and ecosystem-service science. Conservation Biology, 26 (3):418–420, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2011:I

[RF11] Kent H. Redford and Erica Fleishman. Introduction. Conservation Biology, 25(6):1072–1074, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robert:2017:FCS

[RFV^{+}17] Alexandre Robert, Colin Fontaine, Simon Veron, Anne-Christine Monnet, Marine Legrand, Joanne Clavel, Stéphane Chantepie, Denis Couvet, Frédéric Ducarme, Benoît Fontaine, Frédéric Jiguet, Isabelle le Viol, Jonathan Rolland, François Sarrazin, Céline Teplitsky, and Maud Mouchet. Fixism and conservation science. Conservation Biology, 31(4):781–788, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2012:CSC

[RGMR12] Kent H. Redford, Craig Groves, Rodrigo A. Medellin, and John G. Robinson. Conservation stories, conservation science, and the role of the intergovernmental platform on biodiversity and ecosystem services. Conservation Biology, 26(5):757–759, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reyes-Garcia:2013:LUC

[RGRMPB^{+}13] Victoria Reyes-Garcia, Isabel Ruiz-Mallen, Luciana Porter-Bolland, Eduardo Garcia-Frapolli, Edward A. Ellis, Maria-Elena Mendez, Diana J. Pritchard, and María-Consuelo Sanchez-Gonzalez. Local understandings of conservation in southeastern Mexico and their implications for communitybased conservation as an alternative paradigm. Conservation Biology, 27(4):856–865, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rose:2016:CEC

[RHCW16] Lucy E. Rose, Geoffrey W. Heard, Yung En Chee, and Brendan A. Wintle. Cost-effective conservation of an endangered frog under uncertainty. Conservation Biology, 30(2):350–361, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rout:2010:OAC

[RHM10] Tracy M. Rout, Dean Heinze, and Michael A. Mccarthy. Optimal allocation of conservation resources to species that may be extinct. Asignación Óptima de recursos para la conservación de especies que pueden estar extintas. Conservation Biology, 24(4):1111–1118, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ryberg:2015:LIL

[RHPF15] Wade A. Ryberg, Michael T. Hill, Charles W. Painter, and Lee A. Fitzgerald. Linking irreplaceable landforms in a selforganizing landscape to sensitivity of population vital rates for an ecological specialist. Conservation Biology, 29(3):888–898, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodriguez:2017:SMI

[RHR^{+}17] Airam Rodríguez, Nick D. Holmes, Peter G. Ryan, Kerry-Jayne Wilson, Lucie Faulquier, Yovana Murillo, André F. Raine, Jay F. Penniman, Verónica Neves, Beneharo Rodríguez, Juan J. Negro, André Chiaradia, Peter Dann, Tracy Anderson, Benjamin Metzger, Masaki Shirai, Lorna Deppe, Jennifer Wheeler, Peter Hodum, Catia Gouveia, Vanda Carmo, Gilberto P. Carreira, Luis Delgado-Alburqueque, Carlos Guerra-Correa, François-Xavier Couzi, Marc Travers, and Matthieu Le Corre. Seabird mortality induced by land-based artificial lights. Conservation Biology, 31(5):986–1001, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:2014:GIC

[RHT14] Sarah E. Reed, Jodi A. Hilty, and David M. Theobald. Guidelines and incentives for conservation development in local landuse regulations. Conservation Biology, 28(1):258–268, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2018:ACM

[RHWS18] Kent H. Redford, Kristin B. Hulvey, Matthew A. Williamson, and Mark W. Schwartz. Assessment of the conservation measures Partnership’s effort to improve conservation outcomes through adaptive management. Conservation Biology, 32(4): 926–937, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roll:2018:NCS

[RIBT18] Uri Roll, Takuya Iwamura, and Oded Berger-Tal. National conservation science conferences as a means of bridging conservation science and practice. Conservation Biology, 33(1): 1200–1202, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rice:2011:ACP

[Ric11] Jake C. Rice. Achieving coherent policies for conservation and sustainable use of marine ecosystems. Conservation Biology, 25(6):1065–1068, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Russell:2016:ILC

[RJA^{+}16] James C. Russell, Holly P. Jones, Doug P. Armstrong, Franck Courchamp, Peter J. Kappes, Philip J. Seddon, Steffen Oppel, Mark J. Rauzon, Phil E. Cowan, Gérard Rocamora, Piero Genovesi, Elsa Bonnaud, Bradford S. Keitt, Nick D. Holmes, and Bernie R. Tershy. Importance of lethal control of invasive predators for island conservation. Conservation Biology, 31(1):670–672, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ryan:2013:LEW

[RJF^{+}13] Maureen E. Ryan, Jarrett R. Johnson, Benjamin M. Fitzpatrick, Linda J. Lowenstine, Angela M. Picco, and H. Bradley Shaffer. Lethal effects of water quality on threatened California salamanders but not on co-occurring hybrid salamanders. Conservation Biology, 27(1):95–102, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roberts:2017:FEP

[RJS17] David L. Roberts, Ivan Jarić, and Andrew R. Solow. On the functional extinction of the passenger pigeon. Conservation Biology, 31(5):1192–1195, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robillard:2017:ASL

[RK17] Cassandra M. Robillard and Jeremy T. Kerr. Assessing the shelf life of cost-efficient conservation plans for species at risk across gradients of agricultural land use. Conservation Biology, 31(4):837–847, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roux:2019:CER

[RKC^{+}19] Dirk J. Roux, Richard T. Kingsford, Carly N. Cook, Jane Carruthers, Katrina Dickson, and Marc Hockings. The case for embedding researchers in conservation agencies. Conservation Biology, 33(6):1266–1274, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rollinson:2014:RAI

[RKH^{+}14] Njal Rollinson, Dave M. Keith, Aimee Lee S. Houde, Paul V. Debes, Meghan C. Mcbride, and Jeffrey A. Hutchings. Risk assessment of inbreeding and outbreeding depression in a captivebreeding program. Conservation Biology, 28(2):529–540, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roberson:2019:NAG

[RKW19] Leslie A. Roberson, Jeremy J. Kiszka, and James E. M. Watson. Need to address gaps in global fisheries observation. Conservation Biology, 33(4):966–968, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rothlisberger:2011:LGM

[RL11] John D. Rothlisberger and David M. Lodge. Limitations of gravity models in predicting the spread of Eurasian watermilfoil. Conservation Biology, 25(1):64–72, February 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rick:2013:IPA

[RL13] Torben C. Rick and Rowan Lockwood. Integrating paleobiology, archeology, and history to inform biological conservation. Conservation Biology, 27(1):45–54, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roucoux:2017:TIT

[RLB^{+}17] K. H. Roucoux, I. T. Lawson, T. R. Baker, D. Del Castillo Torres, F. C. Draper, O. Lähteenoja, M. P. Gilmore, E. N. Honorio Coronado, T. J. Kelly, E. T. A. Mitchard, and C. F. Vriesendorp. Threats to intact tropical peatlands and opportunities for their conservation. Conservation Biology, 31(6): 1283–1292, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ruiz-Lopez:2012:HFC

[RLGG^{+}12] María José Ruiz-López, Natalia Gañan, José Antonio Godoy, Ana Del Olmo, Julian Garde, Gerardo Espeso, Astrid Vargas, Fernando Martinez, Eduardo R. S. Roldán, and Montserrat Gomendio. Heterozygosity-fitness correlations and inbreeding depression in two critically endangered mammals. Conservation Biology, 26(6):1121–1129, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rolls:2015:IST

[RLL15] Robert J. Rolls, Catherine Leigh, and Simone D. Langhans. Improving science through improved acknowledgment of reviewers. Conservation Biology, 29(2):307–308, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:2011:EMD

[RM11] Sarah E. Reed and Adina M. Merenlender. Effects of management of domestic dogs and recreation on carnivores in protected areas in Northern California. Efectos del manejo de perros domésticos y recreación sobre carnívoros en Áreas protegidas en el Norte de California. Conservation Biology, 25(3):504– 513, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:2014:RMV

[RM14a] J. Michael Reed and Earl D. Mccoy. Relation of minimum viable population size to biology, time frame, and objective. Conservation Biology, 28(5):867–870, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Richards:2014:IBC

[RM14b] Meryl Breton Richards and V. Ernesto Méndez. Interactions between carbon sequestration and shade tree diversity in a smallholder coffee cooperative in El Salvador. Conservation Biology, 28(2):489–497, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rohwer:2019:CCC

[RM19] Yasha Rohwer and Emma Marris. Clarifying compassionate conservation with hypotheticals: response to Wallach et al. 2018. Conservation Biology, 33(4):781–783, August 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rist:2010:HRC

[RMGCR10] Janna Rist, E. J. Milner-Gulland, Guy Cowlishaw, and Marcus Rowcliffe. Hunter reporting of catch per unit effort as a monitoring tool in a bushmeat-harvesting system. Información sobre la captura por unidad de esfuerzo proporcionada por cazadores como una herramienta de monitoreo en un sistema de cosecha de carne silvestre. Conservation Biology, 24(2):489–499, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rosenbaum:2014:LRM

[RMKM14] Howard C. Rosenbaum, Sara M. Maxwell, Francine Kershaw, and Bruce Mate. Long-range movement of humpback whales and their overlap with anthropogenic activity in the South Atlantic Ocean. Conservation Biology, 28(2):604–615, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redfern:2013:ARS

[RMM^{+}13] J. V. Redfern, M. F. Mckenna, T. J. Moore, J. Calambokidis, M. L. Deangelis, E. A. Becker, J. Barlow, K. A. Forney, P. C. Fiedler, and S. J. Chivers. Assessing the risk of ships striking large whales in marine spatial planning. Conservation Biology, 27(2):292–302, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2014:HHJ

[RMRO14] Kent H. Redford, Samuel S. Myers, Taylor H. Ricketts, and Steven A. Osofsky. Human health as a judicious conservation opportunity. Conservation Biology, 28(3):627–629, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ravenelle:2017:GPT

[RN17] Jeremy Ravenelle and Philip J. Nyhus. Global patterns and trends in human–wildlife conflict compensation. Conservation Biology, 31(6):1247–1256, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ryan:2019:SCM

[RNE^{+}19] Gerard Edward Ryan, Emily Nicholson, Jonathan C. Eames, Thomas N. E. Gray, Robin Loveridge, Simon P. Mahood, Phearun Sum, and Michael A. McCarthy. Simultaneous-count models to estimate abundance from counts of unmarked individuals with imperfect detection. Conservation Biology, 33 (3):697–708, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rowland:2018:SAI

[RNM^{+}18] Jessica A. Rowland, Emily Nicholson, Nicholas J. Murray, David A. Keith, Rebecca E. Lester, and Lucie M. Bland. Selecting and applying indicators of ecosystem collapse for risk assessments. Conservation Biology, 32(6):1233–1245, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:2012:CCI

[Rob12] John G. Robinson. Common and conflicting interests in the engagements between conservation organizations and corporations. Conservation Biology, 26(6):967–977, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roberts:2013:REH

[Rob13] David L. Roberts. Refuge-effect hypothesis and the demise of the dodo. Conservation Biology, 28(1):1478–1480, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rogers:2011:UFM

[Rog11] Christopher M. Rogers. Use of fecundity measured directly throughout the breeding season to test a source–sink demographic model. Utilización de la fecundidad medida directamente en la Época reproductiva para probar un modelo demográfico de fuente-sumidero. Conservation Biology, 25(6):

1212–1219, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roque:2016:UHL

[ROQR^{+}16] Fabio O. Roque, Jose Ochoa-Quintero, Danilo B. Ribeiro, Larissa S. M. Sugai, Raul Costa-Pereira, Reinaldo Lourival, and Gilad Bino. Upland habitat loss as a threat to pantanal wetlands. Conservation Biology, 30(5):1131–1134, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rose:2015:CPR

[Ros15] David C. Rose. The case for policy-relevant conservation science. Conservation Biology, 29(3):748–754, June 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rohr:2013:CCM

[RP13] Jason R. Rohr and Brent D. Palmer. Climate change, multiple stressors, and the decline of ectotherms. Conservation Biology, 27(4):741–751, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rittenhouse:2012:LCC

[RPA^{+}12] Chadwick D. Rittenhouse, Anna M. Pidgeon, Thomas P. Albright, Patrick D. Culbert, Murray K. Clayton, Curtis H. Flather, Jeffrey G. Masek, and Volker C. Radeloff. Land-cover change and avian diversity in the Conterminous United States. Cambio de cobertura de suelo y diversidad de aves en los Estados Unidos limítrofes. Conservation Biology, 26(5):821–829, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2013:FFF

[RPS13] Kent H. Redford, Christine Padoch, and Terry Sunderland. Fads, funding, and forgetting in three decades of conservation. Conservation Biology, 27(3):437–438, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reyers:2010:CPT

[RRC^{+}10] Belinda Reyers, Dirk J. Roux, Richard M. Cowling, Aimee E. Ginsburg, Jeanne L. Nel, and Patrick O’ Farrell. Conservation planning as a transdisciplinary process. La planificación de la conservación como un proceso transdisciplinar. Conservation Biology, 24(4):957–965, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodriguez:2011:EIR

[RRCB^{+}11] Jon Paul Rodríguez, Kathryn M. Rodríguez-Clark, Jonathan E. M. Baillie, Neville Ash, John Benson, Timothy Boucher, Claire Brown, Neil D. Burgess, Ben Collen, Michael Jennings, David A. Keith, Emily Nicholson, Carmen Revenga, Belinda Reyers, Mathieu Rouget, Tammy Smith, Mark Spalding, Andrew Taber, Matt Walpole, Irene Zager, and Tara Zamin. Establishing IUCN red list criteria for threatened ecosystems establecimiento de criterios para la lista roja de UICN de ecosistemas amenazados. Conservation Biology, 25(1):21–29, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rosas-Rosas:2010:RLJ

[RRV10] Octavio C. Rosas-Rosas and Raul Valdez. The role of landowners in jaguar conservation in Sonora, Mexico. El papel de propietarios de tierras en la conservación de jaguar en Sonora, México. Conservation Biology, 24(2):366–371, April 2010. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:2018:CAG

[RS18] Janine E. Robinson and Pablo Sinovas. Challenges of analyzing the global trade in CITES-listed wildlife. Conservation Biology, 33(1):1203–1206, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rosa:2012:CSD

[RSE12] Isabel M. D. Rosa, Carlos Souza, Jr., and Robert M. Ewers. Changes in size of deforested patches in the Brazilian Amazon. Cambios en el tamaño de parches deforestados en la Amazonía Brasileña. Conservation Biology, 26(5):932–937, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ranius:2019:ISC

[RSN19] Thomas Ranius, Tord Snäll, and Jenni Nordén. Importance of spatial configuration of deadwood habitats in species conservation. Conservation Biology, 33(5):1205–1207, October 2019.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ruscoe:2013:ESE

[RSPD13] Wendy A. Ruscoe, Peter J. Sweetapple, Mike Perry, and Richard P. Duncan. Effects of spatially extensive control of invasive rats on abundance of native invertebrates in mainland New Zealand forests. Conservation Biology, 27(1):74–82, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2012:CM

[RSS^{+}12] Kent H. Redford, Julia A. Segre, Nick Salafsky, Carlos Martinez del Rio, and Denise McAloose. Conservation and the microbiome. Conservation Biology, 26(2):195–197, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ramage:2013:PTF

[RSS^{+}13] Benjamin S. Ramage, Douglas Sheil, Hannah M. W. Salim, Christine Fletcher, Nur-Zati A. Mustafa, Joann C. Luruthusamay, Rhett D. Harrison, Elizabeth Butod, Ahmad Dzamir Dzulkiply, Abd Rahman Kassim, and Matthew D. Potts. Pseudoreplication in tropical forests and the resulting effects on biodiversity conservation. Conservation Biology, 27(2):364– 372, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:2011:IEP

[RSW11] Thomas E. Reed, Daniel E. Schindler, and Robin S. Waples. Interacting effects of phenotypic plasticity and evolution on population persistence in a changing climate. Efectos interactivos de la plasticidad Fenotípica y evolución sobre la persistencia poblacional en un clima cambiante. Conservation Biology, 25 (1):56–63, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rock:2014:BDC

[RSWM14] Jenny Rock, Andrew Sparrow, Rob Wass, and Henrik Moller. Building dialogue on complex conservation issues in a conference setting. Conservation Biology, 28(6):1428–1433, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:2018:RSS

[RTB^{+}18] Sarah E. Reed, Sarah L. Thomas, Angela T. Bednarek, Dominick A. DellaSala, Megan C. Evans, Carolyn Lundquist, Michael B. Mascia, Tsitsi Y. McPherson, and James E. M. Watson. Roles for scientific societies to engage with conservation policy. Conservation Biology, 32(3):513–515, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Runge:2017:QCG

[RTGR17] Claire A. Runge, Ayesha I. T. Tulloch, Ascelin Gordon, and Jonathan R. Rhodes. Quantifying the conservation gains from shared access to linear infrastructure. Conservation Biology, 31(6):1428–1438, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Runge:2015:GRS

[RTH^{+}15] Claire A. Runge, Ayesha Tulloch, Edd Hammill, Hugh P. Possingham, and Richard A. Fuller. Geographic range size and extinction risk assessment in nomadic species. Conservation Biology, 29(3):865–876, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ruesch:2012:PCI

[RTL^{+}12] Aaron S. Ruesch, Christian E. Torgersen, Joshua J. Lawler, Julian D. Olden, Erin E. Peterson, Carol J. Volk, and David J. Lawrence. Projected climate-induced habitat loss for salmonids in the John Day River Network, Oregon, U.S.A.. Proyección de la pérdida de hábitat de salmónidos inducida por el clima con base en un red de modelos de la temperatura de arroyos. Conservation Biology, 26(5):873–882, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Regan:2013:TDR

[RTT^{+}13] Tracey J. Regan, Barbara L. Taylor, Grant G. Thompson, Jean Fitts Cochrane, Katherine Ralls, Michael C. Runge, and Richard Merrick. Testing decision rules for categorizing species’ extinction risk to help develop quantitative listing criteria for the U.S. Endangered Species Act. Conservation Biology, 27 (4):821–831, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rudd:2011:HRP

[Rud11a] Murray A. Rudd. How research-prioritization exercises affect conservation policy. Efectos de los ejercicios de priorización sobre las políticas de conservación. Conservation Biology, 25(5): 860–866, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rudd:2011:SOG

[Rud11b] Murray A. Rudd. Scientists’ opinions on the global status and management of biological diversity. Opiniones de científicos sobre el estado global y manejo de la diversidad Biológica. Conservation Biology, 25(6):1165–1175, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rondinini:2015:SLM

[RV15] Carlo Rondinini and Piero Visconti. Scenarios of large mammal loss in Europe for the 21^{st} century. Conservation Biology, 29 (4):1028–1036, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rashford:2011:EGC

[RWB11] Benjamin S. Rashford, Johann A. Walker, and Christopher T. Bastian. Economics of grassland conversion to cropland in the Prairie Pothole region. Economía de la conversión de pastizales a tierras agrícolas en la región Prairie Pothole. Conservation Biology, 25(2):276–284, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See erratum [Ano12g].

Rashford:2012:CC

[RWB12] Benjamin S. Rashford, Johann A. Walker, and Christopher T. Bastian. Cover caption. Conservation Biology, 25(2):i, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ringma:2017:MSE

[RWF^{+}17] Jeremy L. Ringma, Brendan Wintle, Richard A. Fuller, Diana Fisher, and Michael Bode. Minimizing species extinctions through strategic planning for conservation fencing. Conservation Biology, 31(5):1029–1038, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ren:2015:ECN

[RYW^{+}15] Guopeng Ren, Stephen S. Young, Lin Wang, Wei Wang, Yongcheng Long, Ruidong Wu, Junsheng Li, Jianguo Zhu, and Douglas W. Yu. Effectiveness of China’s national forest protection program and nature reserves. Conservation Biology, 29(5): 1368–1377, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Skagen:2011:PMA

[SA11] Susan K. Skagen and Amy A. Yackel Adams. Potential misuse of avian density as a conservation metric. Potencial mal uso de la densidad de aves como un indicador de conservación. Conservation Biology, 25(1):48–55, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sutton:2014:GSR

[SA14] N. J. Sutton and P. R. Armsworth. The grain of spatially referenced economic cost and biodiversity benefit data and the effectiveness of a cost targeting strategy. Conservation Biology, 28(6):1451–1461, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shoemaker:2015:INA

[SA15] Kevin T. Shoemaker and H. Resit Akçakaya. Inferring the nature of anthropogenic threats from long-term abundance records. Conservation Biology, 29(1):238–249, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sundstrom:2012:SFG

[SAB12] Shana M. Sundstrom, Craig R. Allen, and Chris Barichievy. Species, functional groups, and thresholds in ecological resilience. Especies, grupos funcionales y umbrales en resiliencia ecológica. Conservation Biology, 26(2):305–314, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schmiedel:2016:CPP

[SAB^{+}16] Ute Schmiedel, Yoseph Araya, Maria Ieda Bortolotto, Linda Boeckenhoff, Winnie Hallwachs, Daniel Janzen, Shekhar S. Kolipaka, Vojtech Novotny, Matilda Palm, Marc Parfondry, Athanasios Smanis, and Pagi Toko. Contributions of paraecologists and parataxonomists to research, conservation, and social development. Conservation Biology, 31(1):506–519, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sandbrook:2013:SRB

[SABV13] Chris Sandbrook, William M. Adams, Bram Büscher, and Bhaskar Vira. Social research and biodiversity conservation. Conservation Biology, 28(1):1487–1490, December 2013. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simonson:2012:UAL

[SAC12] William D. Simonson, Harriet D. Allen, and David A. Coomes. Use of an airborne lidar system to model plant species composition and diversity of Mediterranean oak forests. Utilización de un sistema lidar aerotransportado para modelar la Composición y diversidad de especies de plantas en bosques Mediterráneos de roble. Conservation Biology, 26(5):840–850, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sims:2014:IES

[SAGSG^{+}14] Katharine R. E. Sims, Jennifer M. Alix-Garcia, Elizabeth Shapiro-Garza, Leah R. Fine, Volker C. Radeloff, Glen Aronson, Selene Castillo, Carlos Ramirez-Reyes, and Patricia Yañez-Pagans. Improving environmental and social targeting through adaptive management in Mexico’s payments for hydrological services program. Conservation Biology, 28(5):1151–1159, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

St-Amour:2010:ETA

[SAGSHL10] Valérie St-Amour, Trenton W. J. Garner, Albrecht I. Schulte-Hostedde, and David Lesbarrères. Effects of two amphibian pathogens on the developmental stability of green frogs. Efectos de dos patógenos de anfibios sobre la estabilidad del desarrollo de ranas verdes. Conservation Biology, 25(6):788–794, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salafsky:2012:EG

[Sal12] Nick Salafsky. The enemy of the good. Conservation Biology, 26(2):190–192, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sandler:2010:VSE

[San10] Ronald Sandler. The value of species and the ethical foundations of assisted colonization.. El valor de las especies y los fundamentos Éticos de la colonización asistida. Conservation Biology, 24(2):424–431, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sandler:2014:ERL

[San14] Ronald Sandler. The ethics of reviving long extinct species. Conservation Biology, 28(2):354–360, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanderson:2019:FAR

[San19] Eric W. Sanderson. A full and authentic reckoning of species’ ranges for conservation: response to Akçakaya et al. 2018. Conservation Biology, 33(5):1208–1210, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See reply [SGA^{+}19].

Svensson:2019:LTN

[SAS^{+}19] Johan Svensson, Jon Andersson, Per Sandström, Grzegorz Mikusiński, and Bengt Gunnar Jonsson. Landscape trajectory of natural boreal forest loss as an impediment to green infrastructure. Conservation Biology, 33(2):152–163, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sajeva:2013:RIT

[SASO13] Maurizio Sajeva, Claudio Augugliaro, Matthew J. Smith, and Elisabetta Oddo. Regulating Internet trade in CITES species. Conservation Biology, 27(2):429–430, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Solow:2014:USI

[SB14] Andrew R. Solow and Andrew R. Beet. On uncertain sightings and inference about extinction. Conservation Biology, 28(4):

1119–1123, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shaffer:2016:EWE

[SB16] Jill A. Shaffer and Deborah A. Buhl. Effects of wind-energy facilities on breeding grassland bird distributions. Conservation Biology, 30(1):59–71, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Suryawanshi:2014:MFA

[SBB^{+}14] Kulbhushansingh R. Suryawanshi, Saloni Bhatia, Yash Veer Bhatnagar, Stephen Redpath, and Charudutt Mishra. Multiscale factors affecting human attitudes toward snow leopards and wolves. Conservation Biology, 28(6):1657–1666, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seibold:2015:AER

[SBB^{+}15] Sebastian Seibold, Roland Brandl, Jörn Buse, Torsten Hothorn, Jürgen Schmidl, Simon Thorn, and Jörg Müller. Association of extinction risk of saproxylic beetles with ecological degradation of forests in Europe. Conservation Biology, 29(2):382–390, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Segan:2011:UCE

[SBBP11] Daniel B. Segan, Madeleine C. Bottrill, Peter W. J. Baxter, and Hugh P. Possingham. Using conservation evidence to guide management. Conservation Biology, 25(1):200–202, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sinclair:2018:CME

[SBG^{+}18] Steve J. Sinclair, Matthew J. Bruce, Peter Griffioen, Amanda Dodd, and Matthew D. White. A condition metric for Eucalyptus woodland derived from expert evaluations. Conservation Biology, 32(1):195–204, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shoemaker:2013:RMV

[SBJG13] Kevin T. Shoemaker, Alvin R. Breisch, Jesse W. Jaycox, and James P. Gibbs. Reexamining the minimum viable population concept for long-lived species. Conservation Biology, 27(3):542– 551, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shoemaker:2014:DMV

[SBJG14] Kevin T. Shoemaker, Alvin R. Breisch, Jesse W. Jaycox, and James P. Gibbs. Disambiguating the minimum viable population concept: Response to Reed and McCoy. Conservation Biology, 28(5):871–873, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Spring:2010:BRC

[SBN^{+}10] Daniel Spring, Jiri Baum, Ralph Mac Nally, Michael Mackenzie, A. Sanchez-Azofeifa, and James R. Thomson. Building a regionally connected reserve network in a changing and uncertain world. Construcción de una red de reservas conectada regionalmente en un mundo cambiante e incierto. Conservation Biology, 24(3):691–700, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Santini:2019:AHP

[SBR^{+}19] Luca Santini, Stuart H. M. Butchart, Carlo Rondinini, Ana Benítez-López, Jelle P. Hilbers, Aafke M. Schipper, Mirza Cengic, Joseph A. Tobias, and Mark A. J. Huijbregts. Applying habitat and population-density models to land-cover time series to inform IUCN red list assessments. Conservation Biology, 33(5):1084–1093, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:2015:ECC

[SBS^{+}15] Robert J. Smith, Duan Biggs, Freya A. V. St. John, Michael ’t Sas-Rolfes, and Robert Barrington. Elephant conservation and corruption beyond the ivory trade. Conservation Biology, 29(3):953–956, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sonter:2014:OIM

[SBSF14] L. J. Sonter, D. J. Barrett, and B. S. Soares-Filho. Offsetting the impacts of mining to achieve no net loss of native vegetation. Conservation Biology, 28(4):1068–1076, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schultz:2011:RWG

[SBT^{+}11] Jennifer K. Schultz, Jason D. Baker, Robert J. Toonen, Albert L. Harting, and Brian W. Bowen. Range-wide genetic connectivity of the Hawaiian monk seal and implications for translocation. Conectividad genética en el Área de distribución de monachus schauinslandi e implicaciones para la translocación. Conservation Biology, 25(1):124–132, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sherley:2017:DER

[SBU^{+}17] Richard B. Sherley, Philna Botha, Les G. Underhill, Peter G. Ryan, Danie van Zyl, Andrew C. Cockcroft, Robert J. M. Crawford, Bruce M. Dyer, and Timothée R. Cook. Defining ecologically relevant scales for spatial protection with long-term data on an endangered seabird and local prey availability. Conservation Biology, 31(6):1312–1321, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sagot:2015:ERS

[SC15] Maria Sagot and Gloriana Chaverri. Effects of roost specialization on extinction risk in bats. Conservation Biology, 29(6): 1666–1673, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shen:2019:BWA

[SC19] Tsung-Jen Shen and Youhua Chen. A Bayesian-weighted approach to predicting the number of newly discovered rare species. Conservation Biology, 33(2):444–455, April 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scherer:2017:EKP

[SCA17] Laura Scherer, Michael Curran, and Miguel Alvarez. Expanding Kenya’s protected areas under the convention on biological diversity to maximize coverage of plant diversity. Conservation Biology, 31(2):302–310, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schmidt:2010:EID

[Sch10] Benedikt R. Schmidt. Estimating the impact of disease in species threatened by amphibian chytrid fungus: Comment on Murray et al. Conservation Biology, 25(6):897–899, June 2010.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [MSSM09].

Schultz:2011:CMB

[Sch11] P. Wesley Schultz. Conservation means behavior. Conservation Biology, 25(6):1080–1083, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schaefer:2014:CSP

[Sch14a] Jame Schaefer. Collaborating for a shared purpose. Conservation Biology, 28(4):1132–1134, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schmidt:2014:APA

[Sch14b] Aimee Schmidt. After the protected area. Conservation Biology, 28(2):620–622, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stock:2018:UAR

[SCHM18] Andy Stock, Larry B. Crowder, Benjamin S. Halpern, and Fiorenza Micheli. Uncertainty analysis and robust areas of high and low modeled human impact on the global oceans. Conservation Biology, 32(6):1368–1379, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sawaya:2013:DCU

[SCK13] Michael A. Sawaya, Anthony P. Clevenger, and Steven T. Kalinowski. Demographic connectivity for ursid populations at wildlife crossing structures in Banff National Park. Conservation Biology, 27(4):721–730, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sievanen:2012:CIR

[SCL12] Leila Sievanen, Lisa M. Campbell, and Heather M. Leslie. Challenges to interdisciplinary research in ecosystem-based management. Retos para la investigación interdisciplinaria en el manejo basado en ecosistemas. Conservation Biology, 26(2):315–323, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scott:2016:CBH

[Sco16] Michael Scott. A celebration of British habitats and a personal perspective on its birds of prey. Conservation Biology, 30 (2):438–440, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shwartz:2013:VGC

[SCSJ13] Assaf Shwartz, Helene Cheval, Laurent Simon, and Romain Julliard. Virtual garden computer program for use in exploring the elements of biodiversity people want in cities. Conservation Biology, 27(4):876–886, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Silk:2018:CCB

[SCWN18] Matthew J. Silk, Sarah L. Crowley, Anna J. Woodhead, and Ana Nuno. Considering connections between Hollywood and biodiversity conservation. Conservation Biology, 32(3):597– 606, June 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schwenk:2011:MFL

[SD11] W. Scott Schwenk and Therese M. Donovan. A multispecies framework for landscape conservation planning. Un marco de referencia con múltiples especies para la planificación de la conservación del paisaje. Conservation Biology, 27(3):1010–1021, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sutherland:2010:AER

[SDB10] Ronald W. Sutherland, Phillip R. Dunning, and Whit M. Baker. Amphibian encounter rates on roads with different amounts of traffic and urbanization. Tasas de encuentro de anfibios en carreteras con diferentes cantidades de tráfico y urbanización. Conservation Biology, 24(6):1626–1635, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Spencer:2017:CEB

[SDT17] R.-J. Spencer, J. U. Van Dyke, and Michael B. Thompson. Critically evaluating best management practices for preventing freshwater turtle extinctions. Conservation Biology, 31(6): 1340–1349, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Selva:2016:DFH

[Sel16] Nuria Selva. Defining forest, habitat, and wilderness in a changing world. Conservation Biology, 30(5):1140–1142, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Spake:2015:MAF

[SEM^{+}15] Rebecca Spake, Thomas H. G. Ezard, Philip A. Martin, Adrian C. Newton, and C. Patrick Doncaster. A meta-analysis of functional group responses to forest recovery outside of the tropics. Conservation Biology, 29(6):1695–1703, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Serreze:2010:URC

[Ser10] Mark C. Serreze. Understanding recent climate change. Entendiendo el cambio climático reciente. Conservation Biology, 24 (1):10–17, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanchez-Fernandez:2018:MMB

[SFAA^{+}18] David Sánchez-Fernández, Pedro Abellán, Pedro Aragón, Sara Varela, and Mar Cabeza. Matches and mismatches between conservation investments and biodiversity values in the European union. Conservation Biology, 32(1):109–115, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shoo:2016:SRT

[SFKC16] Luke P. Shoo, Kylie Freebody, John Kanowski, and Carla P. Catterall. Slow recovery of tropical old-field rainforest regrowth and the value and limitations of active restoration. Conservation Biology, 30(1):121–132, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sewall:2011:RSC

[SFM^{+}11] Brent J. Sewall, Amy L. Freestone, Mohamed F. E. Moutui, Nassuri Toilibou, Ishaka Saïd, Saindou M. Toumani, Daoud Attoumane, and Cheikh M. Iboura. Reorienting systematic conservation assessment for effective conservation planning. Reorientación de la evaluación Sistemática de la conservación por la planificación efectiva de la conservación. Conservation Biology, 25(4):688–696, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scholte:2010:DIT

[SG10] Paul Scholte and Wouter T. De Groot. From debate to insight: Three models of immigration to protected areas. Conservation Biology, 24(2):630–632, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Squires:2018:LCB

[SG18] Dale Squires and Serge Garcia. The least-cost biodiversity impact mitigation hierarchy with a focus on marine fisheries and bycatch issues. Conservation Biology, 32(5):989–997, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sarkki:2019:MSE

[SG19] Simo Sarkki and Nicolás Acosta García. Merging social equity and conservation goals in IPBES. Conservation Biology, 33(5): 1214–1218, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stephenson:2019:DIR

[SGA^{+}19] P. J. Stephenson, Molly K. Grace, H. Resit Akçakaya, Ana S. L. Rodrigues, Barney Long, David P. Mallon, Erik Meijaard, Jon Paul Rodriguez, Richard P. Young, Thomas M. Brooks, and Craig Hilton-Taylor. Defining the indigenous ranges of species to account for geographic and taxonomic variation in the history of human impacts: reply to Sanderson 2019. Conservation Biology, 33(5):1211–1213, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [San19].

Selinske:2018:RPC

[SGB^{+}18] Matthew J. Selinske, Georgia E. Garrard, Sarah A. Bekessy, Ascelin Gordon, Alexander M. Kusmanoff, and Fiona Fidler. Revisiting the promise of conservation psychology. Conservation Biology, 32(6):1464–1468, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Santangeli:2019:PAC

[SGB^{+}19] Andrea Santangeli, Marco Girardello, Evan Buechley, Andre Botha, Enrico Di Minin, and Atte Moilanen. Priority areas for conservation of Old World vultures. Conservation Biology, 33 (5):1056–1065, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simonetti:2013:PHN

[SGE13] Javier A. Simonetti, Audrey A. Grez, and Cristián F. Estades. Providing habitat for native mammals through understory enhancement in forestry plantations. Conservation Biology, 28 (4):1117–1121, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Searle:2011:DHS

[SGH^{+}11] C. L. Searle, S. S. Gervasi, J. Hua, J. I. Hammond, R. A. Relyea, D. H. Olson, and A. R. Blaustein. Differential host susceptibility to Batrachochytrium dendrobatidis, an emerging amphibian pathogen. Susceptibilidad diferencial de huéspedes a Batrachochytrium dendrobatidis, un patógeno de anfibios emergente. Conservation Biology, 25(5):965–974, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Searle:2013:CCC

[SGH^{+}13] C. L. Searle, S. S. Gervasi, J. Hua, J. I. Hammond, R. A. Relyea, D. H. Olson, and A. R. Blaustein. Cover caption: Conservation Biology. Conservation Biology, 25(5):i, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salomon:2010:KFC

[SGS^{+}10] Anne K. Salomon, Sarah K. Gaichas, Nick T. Shears, Jennifer E. Smith, Elizabeth M. P. Madin, and Steven D. Gaines. Key features and context-dependence of fishery-induced trophic cascades. Atributos clave y dependencia del contexto de cascadas tróficas inducidas por pesquerías. Conservation Biology, 24(2):382–394, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simon:2012:RDE

[SGSR^{+}12] Miguel A. Simón, José M. Gil-Sánchez, Gema Ruiz, Germán Garrote, Emil B. Mccain, Leonardo Fernández, Marcos López-Parra, Eva Rojas, Rafael Arenas-Rojas, Teresa Del Rey, Maribel García-Tardío, and Guillermo López. Reverse of the decline of the endangered Iberian lynx. Conservation Biology, 27(1): 731–736, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanderson:2011:CC

[SH11a] Eric W. Sanderson and Amanda Huron. Conservation in the city. Conservation Biology, 25(3):421–423, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Szabo:2011:AIH

[SH11b] Péter Szabó and Radim Hédl. Advancing the integration of history and ecology for conservation. Avanzando en la integración de la historia y la ecología para la conservación. Conservation Biology, 25(4):680–687, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Szabo:2013:GIR

[SH13] Péter Szabó and Radim Hédl. Grappling with interdisciplinary research: Response to Pooley. Conservation Biology, 28(1): 1484–1486, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shiffman:2016:PCP

[SH16] David S. Shiffman and Neil Hammerschlag. Preferred conservation policies of shark researchers. Conservation Biology, 30 (4):805–815, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Setiadi:2010:GSH

[SHA^{+}10a] Mohammad Iqbal Setiadi, Amir Hamidy, Zainal Abidin, Dwi Susanto, Rafe M. Brown, A. Townsend Peterson, Xingong Li, and Ben J. Evans. Genetic structure of herpetofauna on Halmahera Island, Indonesia: Implications for Aketajawe-Lolobata National Park. Estructura genética de la herpetofauna de la Isla Halmahera, Indonesia: Implicaciones para el Parque Nacional Aketajawe-Lolobata. Conservation Biology, 24(2):553– 562, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shaw:2010:EGS

[Sha10b] James H. Shaw. Eternal grasslands, shifting paradigms. Conservation Biology, 24(6):1692–1693, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shaffer:2015:CF

[Sha15a] Mark Shaffer. Changing filters. Conservation Biology, 29 (3):611–612, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shaw:2015:AIE

[Sha15b] Scott R. Shaw. Appreciating insects enough to save them. Conservation Biology, 29(5):1489–1490, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scheele:2017:RHA

[SHB^{+}17] Ben C. Scheele, David A. Hunter, Laura A. Brannelly, Lee F. Skerratt, and Don A. Driscoll. Reservoir-host amplification of disease impact in an endangered amphibian. Conservation Biology, 31(3):592–600, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Steneck:2011:CGT

[SHC^{+}11] R. S. Steneck, T. P. Hughes, J. E. Cinner, W. N. Adger, S. N. Arnold, F. Berkes, S. A. Boudreau, K. Brown, C. Folke, L. Gunderson, P. Olsson, M. Scheffer, E. Stephenson, B. Walker, J. Wilson, and B. Worm. Creation of a gilded trap by the high economic value of the Maine lobster fishery. Creación de una trampa dorada debido al alto valor económico de la pesquería de langosta en Maine. Conservation Biology, 25(5):904– 912, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sorice:2011:ISP

[SHCD11] Michael G. Sorice, Wolfgang Haider, J. Richard Conner, and Robert B. Ditton. Incentive structure of and private landowner participation in an endangered species conservation program. Estructura de los incentivos para propietarios privados y su Participación en un programa de conservación de especies.

Conservation Biology, 25(6):587–596, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scheele:2014:IRE

[SHG^{+}14] Ben C. Scheele, David A. Hunter, Laura F. Grogan, Lee Berger, Jon E. Kolby, Michael S. Mcfadden, Gerry Marantelli, Lee F. Skerratt, and Don A. Driscoll. Interventions for reducing extinction risk in chytridiomycosis-threatened amphibians. Conservation Biology, 28(5):1195–1205, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Suhonen:2010:LED

[SHLK^{+}10] Jukka Suhonen, Milla Hilli-Lukkarinen, Esa Korkeamäki, Markku Kuitunen, Johanna Kullas, Jouni Penttinen, and Jukka Salmela. Local extinction of dragonfly and damselfly populations in low- and high-quality habitat patches. Extinción local de poblaciones de libélulas en parches de hábitat de alta y baja calidad. Conservation Biology, 24(4):1148–1153, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stokes:2010:LLU

[SHO^{+}10] David L. Stokes, Marian F. Hanson, Deborah D. Oaks, Jaime E. Straub, and Aileen V. Ponio. Local land-use planning to conserve biodiversity: Planners’ perspectives on what works. Planificación local del uso de suelo para conservar la biodiversidad: Perspectivas de planificadores sobre lo que funciona. Conservation Biology, 24(2):450–460, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schnell:2013:EER

[SHPR13] Jessica K. Schnell, Grant M. Harris, Stuart L. Pimm, and Gareth J. Russell. Estimating extinction risk with metapopulation models of large-scale fragmentation. Conservation Biology, 27(3):520–530, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sievers:2019:FOP

[SHSP19] Michael Sievers, Robin Hale, Stephen E. Swearer, and Kirsten M. Parris. Frog occupancy of polluted wetlands in urban landscapes. Conservation Biology, 33(2):389–402, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schuyler:2014:GAA

[SHWT14] Qamar Schuyler, Britta Denise Hardesty, Chris Wilcox, and Kathy Townsend. Global analysis of anthropogenic debris ingestion by sea turtles. Conservation Biology, 28(1):129–139, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sikes:2012:ICT

[Sik12] Benjamin A. Sikes. Internalizing conservation through our own microbes. Conservation Biology, 26(2):198, April 2012. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stat:2019:CUE

[SJD^{+}19] Michael Stat, Jeffrey John, Joseph D. DiBattista, Stephen J. Newman, Michael Bunce, and Euan S. Harvey. Combined use of eDNA metabarcoding and video surveillance for the assessment of fish biodiversity. Conservation Biology, 33(2):196–205, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stone:2013:MED

[SJH13] Emma Louise Stone, Gareth Jones, and Stephen Harris. Mitigating the effect of development on bats in England with derogation licensing. Conservation Biology, 27(6):1324–1334, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saltz:2019:CUR

[SJH19] David Saltz, James Justus, and Buddy Huffaker. The crucial but underrepresented role of philosophy in conservation science curricula. Conservation Biology, 33(2):217–220, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smyser:2013:UET

[SJP^{+}13] Timothy J. Smyser, Scott A. Johnson, L. Kristen Page, Cassie M. Hudson, and Olin E. Rhodes, Jr. Use of experimental translocations of allegheny woodrat to decipher causal agents of decline. Conservation Biology, 27(4):752–762, August 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stanford:2019:MAE

[SJZ19] Bronwen Stanford, Holly Jones, and Erika Zavaleta. Metaanalysis of the effects of upstream land cover on stream recovery. Conservation Biology, 33(2):351–360, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sayre:2010:PEU

[SK10] Nathan F. Sayre and Richard L. Knight. Potential effects of United States–Mexico border hardening on ecological and human communities in the Malpai Borderlands. Conservation Biology, 26(2):345–348, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simmons:2015:DMS

[SKBE15] Robert E. Simmons, Holger Kolberg, Rod Braby, and Birgit Erni. Declines in migrant shorebird populations from a winterquarter perspective. Conservation Biology, 29(3):877–887, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shaw:2012:ECA

[SKC^{+}12] M. Rebecca Shaw, Kirk Klausmeyer, D. Richard Cameron, Jason Mackenzie, and Patrick Roehrdanz. Economic costs of achieving current conservation goals in the future as climate changes. Costo económico en el futuro del logro de metas de conservación actuales a medida que cambia el clima. Conservation Biology, 26(3):385–396, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stuble:2011:RED

[SKCS11] Katharine L. Stuble, L. Katherine Kirkman, C. Ronald Carroll, and Nathan J. Sanders. Relative effects of disturbance on red imported fire ants and native ant species in a longleaf pine ecosystem. Efectos relativos de la perturbación sobre especies de hormigas de fuego importadas y nativas en un ecosistema de pinos. Conservation Biology, 25(6):618–622, June 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sung:2013:DEI

[SKH13] Yik-Hei Sung, Nancy E. Karraker, and Billy C. H. Hau. Demographic evidence of illegal harvesting of an endangered Asian turtle. Conservation Biology, 28(1):1421–1428, December 2013.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Swanson:2016:GAP

[SKLP16] Alexandra Swanson, Margaret Kosmala, Chris Lintott, and Craig Packer. A generalized approach for producing, quantifying, and validating citizen science data from wildlife images. Conservation Biology, 31(1):520–531, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Skov:2019:WWF

[Sko19] Henrik Skov. Wildlife and wind farms. Conservation Biology, 33(2):486–488, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sutcliffe:2015:EMR

[SKPP15] P. R. Sutcliffe, C. J. Klein, C. R. Pitcher, and H. P. Possingham. The effectiveness of marine reserve systems constructed using different surrogates of biodiversity. Conservation Biology, 29 (3):657–667, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Suhonen:2014:RLE

[SKSK14] Jukka Suhonen, Esa Korkeamäki, Jukka Salmela, and Markku Kuitunen. Risk of local extinction of odonata freshwater habitat generalists and specialists. Conservation Biology, 28(3): 783–789, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Szabo:2017:UHE

[SKSS^{+}17] Péter Szabó, Petr Kuneš, Helena Svobodová-Svitavská, Markéta Gabriela Švarcová, Lucie Křížová, Silvie Suchánková, Jana Müllerová, and Radim Hédl. Using historical ecology to reassess the conservation status of coniferous forests in central Europe. Conservation Biology, 31(1):150–160, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soykan:2015:UCL

[SL15] Candan U. Soykan and Rebecca L. Lewison. Using communitylevel metrics to monitor the effects of marine protected areas on biodiversity. Conservation Biology, 29(3):775–783, June 2015.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saunders:2016:LES

[SL16] Manu E. Saunders and Gary W. Luck. Limitations of the ecosystem services versus disservices dichotomy. Conservation Biology, 30(6):1363–1365, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sandstrom:2019:UBC

[SLA^{+}19] Annica Sandström, Carina Lundmark, Klas Andersson, Kerstin Johannesson, and Linda Laikre. Understanding and bridging the conservation-genetics gap in marine conservation. Conservation Biology, 33(3):725–728, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seidler:2015:IIL

[SLB^{+}15] Renee G. Seidler, Ryan A. Long, Joel Berger, Scott Bergen, and Jon P. Beckmann. Identifying impediments to longdistance mammal migrations. Conservation Biology, 29(1): 99–109, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saiz:2015:ARL

[SLGB15] Juan Carlos Moreno Saiz, Felipe Domínguez Lozano, Manuel Marrero Gómez, and Ángel Bañares Baudet. Application of the red list index for conservation assessment of Spanish vascular plants. Conservation Biology, 29(3):910–919, June 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Skroch:2010:SNU

[SLH10] Matt Skroch and Laura López-Hoffman. Saving nature under the big tent of ecosystem services: a response to Adams and Redford. Conservation Biology, 26(2):325–327, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sekar:2017:FNE

[SLS17] Nitin Sekar, Chia-Lo Lee, and Raman Sukumar. Functional nonredundancy of elephants in a disturbed tropical forest. Conservation Biology, 31(5):1152–1162, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Speziale:2014:RNC

[SLSH14] Karina L. Speziale, Sergio A. Lambertucci, Cintia P. Souto, and Fernando Hiraldo. Recovering native culture in a world of nonnative species. Conservation Biology, 28(4):1129–1131, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Speziale:2015:NSG

[SLSH15] Karina L. Speziale, Sergio A. Lambertucci, Cintia P. Souto, and Fernando Hiraldo. Native species as goods. Conservation Biology, 29(2):596–598, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shen:2015:VCT

[SLWL15] Xiaoli Shen, Sheng Li, Dajun Wang, and Zhi Lu. Viable contribution of Tibetan sacred mountains in southwestern China to forest conservation. Conservation Biology, 29(6):1518–1526, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Souther:2011:ELA

[SM11] Sara Souther and James B. Mcgraw. Evidence of local adaptation in the demographic response of American ginseng to interannual temperature variation. Evidencias de adaptación local en la respuesta demográfica del ginseng americano a la variación de temperatura interanual. Conservation Biology, 25(5):922– 931, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stephanson:2014:PPM

[SM14] Sheri L. Stephanson and Michael B. Mascia. Putting people on the map through an approach that integrates social data in conservation planning. Conservation Biology, 28(6):1236–1248, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shah:2017:FRC

[SMAG17] Payal Shah, Mindy L. Mallory, Amy W. Ando, and Glenn R. Guntenspergen. Fine-resolution conservation planning with limited climate-change information. Conservation Biology, 31 (2):278–289, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanchez-Mercado:2016:USP

[SMARC^{+}16] Ada Sánchez-Mercado, Marianne Asmüssen, Kathryn M. Rodríguez-Clark, Jon Paul Rodríguez, and Wlodzimierz Jedrzejewski. Using spatial patterns in illegal wildlife uses to reveal connections between subsistence hunting and trade. Conservation Biology, 30(6):1222–1232, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Spies:2010:URN

[SMB^{+}10] Thomas A. Spies, Jay D. Miller, Joseph B. Buchanan, John F. Lehmkuhl, Jerry F. Franklin, Sean P. Healey, Paul F. Hessburg, Hugh D. Safford, Warren B. Cohen, Rebecca S. H. Kennedy, Eric E. Knapp, James K. Agee, and Melinda Moeur. Underestimating risks to the northern spotted owl in fire-prone forests: Response to Hanson et al. Conservation Biology, 26 (2):330–333, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See reply [HODB10].

Stanley:2015:CWT

[SMF^{+}15] Calandra Q. Stanley, Emily A. McKinnon, Kevin C. Fraser, Maggie P. Macpherson, Garth Casbourn, Lyle Friesen, Peter P. Marra, Colin Studds, T. Brandt Ryder, Nora E. Diggs, and Bridget J. M. Stutchbury. Connectivity of wood thrush breeding, wintering, and migration sites based on range-wide tracking. Conservation Biology, 29(1):164–174, February 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soares:2017:MPP

[SMGR17] Letícia Soares, Peter Marra, Lindsey Gray, and Robert E. Ricklefs. The malaria parasite Plasmodium relictum in the endemic avifauna of eastern Cuba. Conservation Biology, 31(6):1477– 1482, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sommerville:2010:ICB

[SMGRJ10] Matthew Sommerville, E. J. Milner-Gulland, Michael Rahajaharison, and Julia P. G. Jones. Impact of a community-based payment for environmental services intervention on forest use in Menabe, Madagascar. Impacto del pago comunitario por intervención de servicios ambientales sobre uso del bosque en Menabe, Madagascar. Conservation Biology, 24(6):1488–1498, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sluijs:2014:EDR

[SMH^{+}14] Annemarieke Spitzen-Van Der Sluijs, An Martel, Caspar A. Hallmann, Wilbert Bosman, Trenton W. J. Garner, Pascale Van Rooij, Robert Jooris, Freddy Haesebrouck, and Frank Pasmans. Environmental determinants of recent endemism of Batrachochytrium dendrobatidis infections in amphibian assemblages in the absence of disease outbreaks. Conservation Biology, 28(6):1302–1311, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salerno:2014:HMP

[SMK14] Jonathan D. Salerno, Monique Borgerhoff Mulder, and Shawn C. Kefauver. Human migration, protected areas, and conservation outreach in Tanzania. Conservation Biology, 28 (3):841–850, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Snover:2013:CUC

[SML^{+}13] Amy K. Snover, Nathan J. Mantua, Jeremy S. Littell, Michael A. Alexander, Michelle M. Mcclure, and Janet Nye. Choosing and using climate-change scenarios for ecologicalimpact assessments and conservation decisions. Conservation Biology, 27(6):1147–1157, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:2019:HCG

[SMLB19] Hillary Smith, Sergio Marrocoli, Alejandro Garcia Lozano, and Xavier Basurto. Hunting for common ground between wildlife governance and commons scholarship. Conservation Biology, 33 (2):9–21, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schachat:2016:CTP

[SMM16] Sandra R. Schachat, Daniel G. Mulcahy, and Joseph R. Mendelson III. Conservation threats and the phylogenetic utility of IUCN red list rankings in Incilius toads. Conservation Biology, 30(1):72–81, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sharma:2014:LLC

[SMS14] Narayan Sharma, M. D. Madhusudan, and Anindya Sinha. Local and landscape correlates of primate distribution and persistence in the remnant lowland rainforests of the Upper Brahmaputra Valley, Northeastern India. Conservation Biology, 28(1): 95–106, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salomon:2013:IUM

[SMTW13] Yacov Salomon, Michael A. Mccarthy, Peter Taylor, and Brendan A. Wintle. Incorporating uncertainty of management costs in sensitivity analyses of matrix population models. Conservation Biology, 27(1):134–144, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salazar:2019:QIE

[SMV19] Gabby Salazar, Morena Mills, and Diogo Veríssimo. Qualitative impact evaluation of a social marketing campaign for conservation. Conservation Biology, 33(3):634–644, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shaffer:2017:SCS

[SMY^{+}17] Christopher A. Shaffer, Marissa S. Milstein, Charakura Yukuma, Elisha Marawanaru, and Phillip Suse. Sustainability and comanagement of subsistence hunting in an indigenous reserve in Guyana. Conservation Biology, 31(5):1119–1131, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stewart:2017:UED

[SMZZ17] Kathryn Stewart, Hongjuan Ma, Jinsong Zheng, and Jianfu Zhao. Using environmental DNA to assess population-wide spatiotemporal reserve use. Conservation Biology, 31(5):1173– 1182, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Spatz:2014:BGT

[SNH^{+}14] Dena R. Spatz, Kelly M. Newton, Reina Heinz, Bernie Tershy, Nick D. Holmes, Stuart H. M. Butchart, and Donald A. Croll. The biogeography of globally threatened seabirds and island conservation opportunities. Conservation Biology, 28(6):

1282–1290, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sullivan:2015:PMT

[SNK15] Brian K. Sullivan, Erika M. Nowak, and Matthew A. Kwiatkowski. Problems with mitigation translocation of herpetofauna. Conservation Biology, 29(1):12–18, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Snyder:2015:SCM

[Sny15] Brian F. Snyder. Solving conservation’s money problems. Conservation Biology, 29(1):1–2, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soutullo:2010:EGN

[Sou10] Alvaro Soutullo. Extent of the global network of terrestrial protected areas. Conservation Biology, 24(2):362–363, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soule:2013:NC

[Sou13] Michael Soulé. The “New Conservation”. Conservation Biology, 27(5):895–897, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soule:2014:SCG

[Sou14a] Michael Soule. Also seeking common ground in conservation. Conservation Biology, 28(3):637–638, June 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soule:2014:CCC

[Sou14b] Michael Soulé. Cover caption: Conservation Biology. Conservation Biology, 27(5):i, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stevenson:2018:LNC

[SPD18] Kathryn T. Stevenson, M. Nils Peterson, and Robert R. Dunn. Leveraging natural capital to solve the shared education and conservation crisis. Conservation Biology, 32(2):490–492, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seddon:2011:FEC

[SPL^{+}11] Philip J. Seddon, Mark Stanley Price, Frederic Launay, Mike Maunder, Pritpal Soorae, Sanjay Molur, Doug Armstrong, Mike Jordan, Sarah Dalrymple, and Piero Genovesi. Frankenstein ecosystems and 21st century conservation agendas: Reply to Oliveira-Santos and Fernandez. Conservation Biology, 25(2): 212, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sattler:2014:SMU

[SPN^{+}14] T. Sattler, G. B. Pezzatti, M. P. Nobis, M. K. Obrist, T. Roth, and M. Moretti. Selection of multiple umbrella species for functional and taxonomic diversity to represent urban biodiversity. Conservation Biology, 28(2):414–426, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seiferling:2012:MPA

[SPPN^{+}12] Ian S. Seiferling, Raphaël Proulx, Pedro R. Peres-Neto, Lenore Fahrig, and Christian Messier. Measuring protected-area isolation and correlations of isolation with land-use intensity and protection status. Medición del aislamiento de Áreas protegidas y correlaciones del aislamiento con la intensidad de uso del suelo y el estatus de protección. Conservation Biology, 26(4): 610–618, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:2010:EPR

[SPSS10] Rebecca K. Smith, Andrew S. Pullin, Gavin B. Stewart, and William J. Sutherland. Effectiveness of predator removal for enhancing bird populations. Efectividad de la remoción de depredadores para incrementar las poblaciones de aves. Conservation Biology, 25(6):820–829, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scott:2011:RDA

[SR11] J. Michael Scott and Janet L. Rachlow. Refocusing the debate about advocacy. Conservation Biology, 25(1):1–3, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Steen:2017:EFT

[SR17] David A. Steen and Orin J. Robinson, Jr. Estimating freshwater turtle mortality rates and population declines following hook ingestion. Conservation Biology, 31(6):1333–1339, December 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Secor:2015:DRB

[SRG^{+}15] D. H. Secor, J. R. Rooker, B. I. Gahagan, M. R. Siskey, and R. W. Wingate. Depressed resilience of bluefin tuna in the Western Atlantic and age truncation. Conservation Biology, 29(2):400–408, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seney:2013:CCM

[SRL^{+}13] Erin E. Seney, Melanie J. Rowland, Ruth Ann Lowery, Roger B. Griffis, and Michelle M. Mcclure. Climate change, marine environments, and the U.S. Endangered Species Act. Conservation Biology, 27(6):1138–1146, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Silva-Rodriguez:2011:ICD

[SRS11] Eduardo A. Silva-Rodríguez and Kathryn E. Sieving. Influence of care of domestic carnivores on their predation on vertebrates. Influencia del cuidado de carnívoros domésticos sobre sus tasas de Depredación en vertebrados. Conservation Biology, 25(4): 808–815, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Samways:2010:RED

[SS10a] Michael J. Samways and Norma J. Sharratt. Recovery of endemic dragonflies after removal of invasive alien trees. Recuperación de libélulas endémicas después de la remoción de Árboles exóticos invasores. Conservation Biology, 26(2):267– 277, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simaika:2010:BUE

[SS10b] John P. Simaika and Michael J. Samways. Biophilia as a universal ethic for conserving biodiversity. Conservation Biology, 25(6):903–906, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Skarpaas:2011:PVA

[SS11] Olav Skarpaas and Odd E. Stabbetorp. Population viability analysis with species occurrence data from museum collections. Análisis de viabilidad poblacional con datos de ocurrencia de especies de colecciones de museo. Conservation Biology, 25 (6):577–586, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shier:2012:FCN

[SS12] Debra M. Shier and Ronald R. Swaisgood. Fitness costs of neighborhood disruption in translocations of a solitary mammal. Costos sobre la adaptabilidad de la disrupción vecinal en translocaciones de un mamífero solitario. Conservation Biology, 26(2):116–123, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Solow:2012:USE

[SSB^{+}12] Andrew Solow, Woollcott Smith, Mark Burgman, Tracy Rout, Brendan Wintle, and David Roberts. Uncertain sightings and the extinction of the ivory-billed woodpecker. Avistamientos inciertos y la extinción del carpintero pico de marfil. Conservation Biology, 26(2):180–184, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanjayan:2012:IFA

[SSBN12] M. Sanjayan, Leah H. Samberg, Timothy Boucher, and Jesse Newby. Intact faunal assemblages in the modern era ensambles de fauna intactos en la era moderna. Conservation Biology, 27 (1):724–730, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soanes:2019:CCM

[SSC^{+}19] Kylie Soanes, Michael Sievers, Yung En Chee, Nicholas S. G. Williams, Manisha Bhardwaj, Adrian J. Marshall, and Kirsten M. Parris. Correcting common misconceptions to inspire conservation action in urban environments. Conservation Biology, 33(2):300–306, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Staus:2010:EAH

[SSD10] Nancy L. Staus, James R. Strittholt, and Dominick A. Dellasala. Evaluating areas of high conservation value in Western Oregon with a decision-support model. Evaluación de Áreas de alto valor de conservación en el Occidente de Oregon mediante un modelo de respaldo a la toma de decisiones. Conservation Biology, 24(3):711–720, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Samojlik:2018:LBZ

[SSD^{+}18] Tomasz Samojlik, Nuria Selva, Piotr Daszkiewicz, Anastasia Fedotova, Adam Wajrak, and Dries Pieter Jan Kuijper. Lessons from Białowieża forest on the history of protection and the world’s first reintroduction of a large carnivore. Conservation Biology, 32(4):808–816, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stork:2017:CET

[SSE^{+}17] N. E. Stork, D. S. Srivastava, P. Eggleton, M. Hodda, G. Lawson, R. R. B. Leakey, and A. D. Watt. Consistency of effects of tropical-forest disturbance on species composition and richness relative to use of indicator taxa. Conservation Biology, 31 (4):924–933, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scholz:2011:ISE

[SSK^{+}11] Astrid J. Scholz, Charles Steinback, Sarah A. Kruse, Mike Mertens, and Howard Silverman. Incorporation of spatial and economic analyses of human-use data in the design of Marine Protected Areas. Incorporación del análisis espacial y económico de datos de uso por humanos al diseño de Áreas Marinas Protegidas. Conservation Biology, 25(3):485–492, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Steven:2017:UCE

[SSMC17] Rochelle Steven, James C. R. Smart, Clare Morrison, and J. Guy Castley. Using a choice experiment and birder preferences to guide bird-conservation funding. Conservation Biology, 31(4):818–827, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schlaepfer:2011:PCV

[SSO11] Martin A. Schlaepfer, Dov F. Sax, and Julian D. Olden. The potential conservation value of non-native species. El valor de conservación potencial de especies no nativas. Conservation Biology, 25(3):428–437, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schlaepfer:2012:TMB

[SSO12] Martin A. Schlaepfer, Dov F. Sax, and Julian D. Olden. Toward a more balanced view of non-native species. Conservation Biology, 26(6):1156–1158, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stockwell:2015:EPS

[SSP^{+}15] Michelle Pirrie Stockwell, Lachlan James Storrie, Carla Jean Pollard, John Clulow, and Michael Joseph Mahony. Effects of pond salinization on survival rate of amphibian hosts infected with the chytrid fungus. Conservation Biology, 29(2):391–399, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shackelford:2018:TBC

[SSRS18] Nancy Shackelford, Rachel J. Standish, William Ripple, and Brian M. Starzomski. Threats to biodiversity from cumulative human impacts in one of North America’s last wildlife frontiers. Conservation Biology, 32(3):672–684, June 2018. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanchirico:2014:IPP

[SSSD14] James N. Sanchirico, Michael R. Springborn, Mark W. Schwartz, and Angela N. Doerr. Investment and the policy process in conservation monitoring. Conservation Biology, 28 (2):361–371, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sandbrook:2011:VPA

[SSVA11] Chris Sandbrook, Ivan R. Scales, Bhaskar Vira, and William M. Adams. Value plurality among conservation professionals. Pluralidad de valores entre profesionales de la conservación. Conservation Biology, 26(3):285–294, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanderson:2015:GSP

[SSW15] Eric W. Sanderson, Daniel B. Segan, and James E. M. Watson. Global status of and prospects for protection of terrestrial geophysical diversity. Conservation Biology, 29(3):649–656, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Socolar:2019:OBL

[SSW19] Jacob B. Socolar, Elvis H. Valderrama Sandoval, and David S. Wilcove. Overlooked biodiversity loss in tropical smallholder agriculture. Conservation Biology, 33(6):1338–1349, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Samarasin:2017:PER

[SSWR17] Pasan Samarasin, Brian J. Shuter, Stephen I. Wright, and F. Helen Rodd. The problem of estimating recent genetic connectivity in a changing world. Conservation Biology, 31(1): 126–135, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Standovar:2016:FCW

[Sta16] Tibor Standovár. Forests in a changing world. Conservation Biology, 30(4):918–920, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sweeney:2019:APR

[STJ^{+}19] Oisín F. Sweeney, John Turnbull, Menna Jones, Mike Letnic, Thomas M. Newsome, and Andy Sharp. An Australian perspective on rewilding. Conservation Biology, 33(4):812–820, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sims:2019:ALE

[STM^{+}19] Katharine R. E. Sims, Jonathan R. Thompson, Spencer R. Meyer, Christoph Nolte, and Joshua S. Plisinski. Assessing the local economic impacts of land protection. Conservation Biology, 33(5):1035–1044, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stokes:2011:ART

[Sto11] David L. Stokes. An adaptive radiation of texts. Conservation Biology, 25(6):1254–1256, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stolton:2012:BGA

[Sto12] Sue Stolton. Balancing global agricultural markets with local stewardship. Conservation Biology, 26(2):189–190, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Struzik:2013:SAM

[Str13] Ed Struzik. Scientific alert: The medium is the message. Conservation Biology, 27(2):431–433, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sullivan:2010:HC

[Sul10] William Sullivan. Humans and conservation. Conservation Biology, 26(2):354–356, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stewart:2018:EST

[SURW18] David R. Stewart, Zachary E. Underwood, Frank J. Rahel, and Annika W. Walters. The effectiveness of surrogate taxa to conserve freshwater biodiversity. Conservation Biology, 32 (1):183–194, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sainsbury:2012:ADR

[SVH12] Anthony W. Sainsbury and Rebecca J. Vaughan-Higgins. Analyzing disease risks associated with translocations. Análisis del riesgo de enfermedad asociado a translocaciones. Conservation Biology, 26(3):442–452, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stambaugh:2015:CRF

[SVN^{+}15] Michael C. Stambaugh, J. Morgan Varner, Reed F. Noss, Daniel C. Dey, Norman L. Christensen, Robert F. Baldwin, Richard P. Guyette, Brice B. Hanberry, Craig A. Harper, Sam G. Lindblom, and Thomas A. Waldrop. Clarifying the role of fire in the deciduous forests of eastern North America: reply to Matlack. Conservation Biology, 29(3):942–946, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Serrouya:2010:IEW

[SW10] Robert Serrouya and Heiko U. Wittmer. Imminent extinctions of woodland caribou from national parks. Conservation Biology, 24(2):363–364, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sodhi:2010:DAE

[SWL^{+}10] Navjot S. Sodhi, David S. Wilcove, Tien Ming Lee, Cagan H. Sekercioglu, R. Subaraj, Henry Bernard, Ding Li Yong, Susan L. H. Lim, Dewi M. Prawiradilaga, and Barry W. Brook. Deforestation and avian extinction on tropical landbridge islands. Deforestación y la extinción de aves en islas tropicales. Conservation Biology, 24(5):1290–1298, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simmonds:2019:CMH

[SWSM19] Jeremy S. Simmonds, James E. M. Watson, Alvaro Salazar, and Martine Maron. A composite measure of habitat loss for entire assemblages of species. Conservation Biology, 33(6):1438–1447, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trimble:2010:SIS

[TA10] Morgan J. Trimble and Rudi J. Van Aarde. Species inequality in scientific study. Inequidad de especies en el estudio científico. Conservation Biology, 25(6):886–890, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Teel:2018:PSS

[TAB^{+}18] Tara L. Teel, Christopher B. Anderson, Mark A. Burgman, Joshua Cinner, Douglas Clark, Rodrigo A. Estévez, Julia P. G. Jones, Tim R. McClanahan, Mark S. Reed, Chris Sandbrook, and Freya A. V. St. John. Publishing social science research in Conservation Biology to move beyond biology. Conservation Biology, 32(1):6–8, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tabarelli:2010:EPT

[TAG^{+}10] Marcelo Tabarelli, Antonio V. Aguiar, Luciana C. Girão, Carlos A. Peres, and Ariadna V. Lopes. Effects of pioneer tree species hyperabundance on forest fragments in Northeastern Brazil. Efectos de la hiperabundancia de especies de Árboles pioneros sobre fragmentos de bosque en el Noreste de Brasil. Conservation Biology, 24(6):1654–1663, December 2010. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Travers:2019:UCD

[TAM^{+}19] Henry Travers, Lucy J. Archer, Geoffrey Mwedde, Dilys Roe, Julia Baker, Andrew J. Plumptre, Aggrey Rwetsiba, and E. J. Milner-Gulland. Understanding complex drivers of wildlife crime to design effective conservation interventions. Conservation Biology, 33(6):1296–1306, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Treves:2019:DBR

[TAP19] Adrian Treves, Kyle A. Artelle, and Paul C. Paquet. Differentiating between regulation and hunting as conservation interventions. Conservation Biology, 33(2):472–475, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tedesco:2014:EHM

[TBB^{+}14] P. A. Tedesco, R. Bigorne, A. E. Bogan, X. Giam, C. Jézéquel, and B. Hugueny. Estimating how many undescribed species have gone extinct. Conservation Biology, 28(6):1360–1370, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trull:2018:PBC

[TBC18] Nicholas Trull, Monika Böhm, and Jamie Carr. Patterns and biases of climate change threats in the IUCN red list. Conservation Biology, 32(1):135–147, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tukiainen:2017:CGC

[TBF^{+}17] Helena Tukiainen, Joseph J. Bailey, Richard Field, Katja Kangas, and Jan Hjort. Combining geodiversity with climate and topography to account for threatened species richness. Conservation Biology, 31(3):364–375, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tracewski:2016:TQI

[TBM^{+}16] Łukasz Tracewski, Stuart H. M. Butchart, Moreno Di Marco, Gentile F. Ficetola, Carlo Rondinini, Andy Symes, Hannah Wheatley, Alison E. Beresford, and Graeme M. Buchanan. Toward quantification of the impact of 21st-century deforestation on the extinction risk of terrestrial vertebrates. Conservation Biology, 30(5):1070–1079, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Turvey:2010:RSB

[TBY^{+}10] Samuel T. Turvey, Leigh A. Barrett, Hao Yujiang, Zhang Lei, Zhang Xinqiao, Wang Xianyan, Huang Yadong, Zhou Kaiya, Tom Hart, and Wang Ding. Rapidly shifting baselines in Yangtze fishing communities and local memory of extinct species. Cambios Rápidos en las directrices en comunidades de pescadores de Yangtze y en la memoria local de especies extintas. Conservation Biology, 25(6):778–787, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tucker:2012:IGE

[TCDR12] Caroline M. Tucker, Marc W. Cadotte, T. Jonathan Davies, and Tony G. Rebelo. Incorporating geographical and evolutionary rarity into conservation prioritization. Incorporación de la rareza geográfica y evolutiva en la priorización de la conservación. Conservation Biology, 26(4):593–601, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tulloch:2013:ACM

[TCP13] Ayesha I. T. Tulloch, Iadine Chadès, and Hugh P. Possingham. Accounting for complementarity to maximize monitoring power for species management. Conservation Biology, 28(4):988–999, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tepedino:2015:DBD

[TDCG15] V. J. Tepedino, Susan Durham, Sydney A. Cameron, and Karen Goodell. Documenting bee decline or squandering scarce resources. Conservation Biology, 29(1):280–282, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tepedino:2012:OPS

[Tep12] V. J. Tepedino. Overestimating population sizes of rare clonal plants. Conservation Biology, 26(5):945–947, October 2012.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tempel:2013:RBO

[TG13] Douglas J. Tempel and R. J. Gutiérrez. Relation between occupancy and abundance for a territorial species, the California spotted owl. Conservation Biology, 28(4):1087–1095, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Troyer:2015:AIU

[TG15] Caitlin M. Troyer and Leah R. Gerber. Assessing the impact of the U.S. Endangered Species Act recovery planning guidelines on managing threats for listed species. Conservation Biology, 29(5):1423–1433, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trathan:2015:PHL

[TGBB^{+}15] Phil N. Trathan, Pablo García-Borboroglu, Dee Boersma, Charles-André Bost, Robert J. M. Crawford, Glenn T. Crossin, Richard J. Cuthbert, Peter Dann, Lloyd Spencer Davis, Santiago De La Puente, Ursula Ellenberg, Heather J. Lynch, Thomas Mattern, Klemens Pütz, Philip J. Seddon, Wayne Trivelpiece, and Barbara Wienecke. Pollution, habitat loss, fishing, and climate change as critical threats to penguins. Conservation Biology, 29(1):31–41, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thompson:2010:CVS

[TGD10] John D. Thompson, Myriam Gaudeul, and Max Debussche. Conservation value of sites of hybridization in peripheral populations of rare plant species. Valor de conservación de sitios de Hibridación en poblaciones periféricas de especies de plantas raras. Conservation Biology, 26(2):236–245, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomassen:2018:DDP

[THD^{+}18] Henri A. Thomassen, Ryan J. Harrigan, Kathleen Semple Delaney, Seth P. D. Riley, Laurel E. K. Serieys, Katherine Pease, Robert K. Wayne, and Thomas B. Smith. Determining the drivers of population structure in a highly urbanized landscape to inform conservation planning. Conservation Biology, 32(1):

148–158, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thorbjarnarson:2010:PS

[Tho10] John Thorbjarnarson. The plight of serpents. Conservation Biology, 24(2):638–639, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tenan:2016:EMR

[TIB^{+}16] Simone Tenan, Aaron Iemma, Natalia Bragalanti, Paolo Pedrini, Marta De Barba, Ettore Randi, Claudio Groff, and Meritxell Genovart. Evaluating mortality rates with a novel integrated framework for nonmonogamous species. Conservation Biology, 30(6):1307–1319, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Temel:2018:LMP

[TJJB18] Julia Temel, Aled Jones, Nikoleta Jones, and Lenke Balint. Limits of monetization in protecting ecosystem services. Conservation Biology, 32(5):1048–1062, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tassin:2015:MPW

[TK15] Jacques Tassin and Christian A. Kull. A melting pot world of species: reply to Speziale et al. Conservation Biology, 29 (2):593–595, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thompson:2017:ELH

[TLPP17] Denise M. Thompson, Day B. Ligon, Jason C. Patton, and Monica Papeş. Effects of life-history requirements on the distribution of a threatened reptile. Conservation Biology, 31 (3):427–436, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Turnhout:2016:CSN

[TLT16] Esther Turnhout, Anna Lawrence, and Sander Turnhout. Citizen science networks in natural history and the collective validation of biodiversity data. Conservation Biology, 31(1):532– 539, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Teel:2010:UDP

[TM10] Tara L. Teel and Michael J. Manfredo. Understanding the diversity of public interests in wildlife conservation. Entendiendo la diversidad de intereses Públicos en la conservación de vida silvestre. Conservation Biology, 24(1):128–139, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Troianowski:2017:ETN

[TMD^{+}17] Mathieu Troïanowski, Nathalie Mondy, Adeline Dumet, Caroline Arcanjo, and Thierry Lengagne. Effects of traffic noise on tree frog stress levels, immunity, and color signaling. Conservation Biology, 31(5):1132–1140, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thiault:2018:MSE

[TMG^{+}18] Lauric Thiault, Paul Marshall, Stefan Gelcich, Antoine Collin, Frédérique Chlous, and Joachim Claudet. Mapping social– ecological vulnerability to inform local decision making. Conservation Biology, 32(2):447–456, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tulloch:2015:ERA

[TMJ^{+}15] Ayesha I. T. Tulloch, Richard F. Maloney, Liana N. Joseph, Joseph R. Bennett, Martina M. I. Di Fonzo, William J. M. Probert, Shaun M. O’Connor, Jodie P. Densem, and Hugh P. Possingham. Effect of risk aversion on prioritizing conservation projects. Conservation Biology, 29(2):513–524, April 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tulloch:2016:UEM

[TMK^{+}16] Ayesha I. T. Tulloch, Alessio Mortelliti, Geoffrey M. Kay, Daniel Florance, and David Lindenmayer. Using empirical models of species colonization under multiple threatening processes to identify complementary threat-mitigation strategies. Conservation Biology, 30(4):867–882, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Teixeira:2018:HBA

[TMM^{+}18] João B. Teixeira, Rodrigo L. Moura, Morena Mills, Carissa Klein, Christopher J. Brown, Vanessa M. Adams, Hedley Grantham, Matthew Watts, Deborah Faria, Gilberto M.

Amado-Filho, Alex C. Bastos, Reinaldo Lourival, and Hugh P. Possingham. A habitat-based approach to predict impacts of marine protected areas on fishers. Conservation Biology, 32(5): 1096–1106, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thompson:2016:IDF

[TND16] Michelle E. Thompson, A. Justin Nowakowski, and Maureen A. Donnelly. The importance of defining focal assemblages when evaluating amphibian and reptile responses to land use. Conservation Biology, 30(2):249–258, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Treves:2013:LAA

[TNTS13] Adrian Treves, Lisa Naughton-Treves, and Victoria Shelley. Longitudinal analysis of attitudes toward wolves. Conservation Biology, 27(2):315–323, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Torres:2011:PBH

[Tor11] Roselinde Torres. Pricking the bubble of heroic leadership. Conservation Biology, 27(3):1060–1061, October 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Torok:2018:PPR

[Tör18] Péter Török. Policy and practice in restoration. Conservation Biology, 32(1):249–250, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Taylor:2018:IFN

[TPL18] Andrew T. Taylor, Monica Papeş, and James M. Long. Incorporating fragmentation and non-native species into distribution models to inform fluvial fish conservation. Conservation Biology, 32(1):171–182, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ticktin:2018:LBM

[TQD^{+}18] Tamara Ticktin, Shimona Quazi, Rachel Dacks, Mesulame Tora, Ashley McGuigan, Zoe Hastings, and Alivereti Naikatini. Linkages between measures of biodiversity and community resilience in Pacific Island agroforests. Conservation Biology, 32 (5):1085–1095, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Todd:2016:UCS

[TRPD16] Brian D. Todd, Jonathan P. Rose, Steven J. Price, and Michael E. Dorcas. Using citizen science data to identify the sensitivity of species to human land use. Conservation Biology, 30(6):1266–1276, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trukhanova:2018:AMM

[Tru18a] Irina S. Trukhanova. Approaching marine mammal conservation through years of research and public awareness. Conservation Biology, 32(4):968–969, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trukhanova:2018:PST

[Tru18b] Irina S. Trukhanova. Planning, science, and traditional knowledge to save our seas. Conservation Biology, 33(1):1211–1212, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tsikliras:2014:STL

[TSA^{+}14] Athanassios C. Tsikliras, Konstantinos I. Stergiou, Nikolaos Adamopoulos, Daniel Pauly, and Eleni Mente. Shift in trophic level of Mediterranean mariculture species. Conservation Biology, 28(4):1124–1128, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tyler:2014:UVA

[TSH^{+}14] Nicholas Tyler, Karl-Arne Stokkan, Chris Hogg, Christian Nellemann, Arnt-Inge Vistnes, and Glen Jeffery. Ultraviolet vision and avoidance of power lines in birds and mammals. Conservation Biology, 28(3):630–631, June 2014. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Truong:2016:PAP

[TT16] Chi Truong and Stefan Trück. The principal-agent problem in coastal development: response to Mills et al. Conservation Biology, 30(5):1128–1130, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tulloch:2014:VUF

[TTEM14] Ayesha I. T. Tulloch, Vivitskaia J. D. Tulloch, Megan C. Evans, and Morena Mills. The value of using feasibility models in systematic conservation planning to predict landholder management uptake. Conservation Biology, 28(6):1462–1473, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tablado:2010:PLT

[TTSZH10] Zulima Tablado, José L. Tella, José A. Sánchez-Zapata, and Fernando Hiraldo. The paradox of the long-term positive effects of a North American crayfish on a European community of predators la paradoja de los efectos positivos a largo plazo de un langostino norteamericano sobre una comunidad de depredadores europeos. Conservation Biology, 24(5):1230– 1238, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tobon:2017:RPG

[TUHK^{+}17] Wolke Tobón, Tania Urquiza-Haas, Patricia Koleff, Matthias Schröter, Rubén Ortega-Álvarez, Julio Campo, Roberto Lindig-Cisneros, José Sarukhán, and Aletta Bonn. Restoration planning to guide Aichi targets in a megadiverse country. Conservation Biology, 31(5):1086–1097, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tuanmu:2016:EPE

[TVY^{+}16] Mao-Ning Tuanmu, Andrés Viña, Wu Yang, Xiaodong Chen, Ashton M. Shortridge, and Jianguo Liu. Effects of payments for ecosystem services on wildlife habitat recovery. Conservation Biology, 30(4):827–835, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Theberge:2011:MCE

[TW11] Jeannette Theberge and Gregg Walker. Managing for caribou and ecological integrity: Reply to Serrouya and Wittmer 2010. Conservation Biology, 25(5):858–859, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tewksbury:2014:RCS

[TW14] Joshua Tewksbury and Gernot Wagner. The role of civil society in recalibrating conservation science incentives. Conservation Biology, 28(6):1437–1439, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Twidwell:2018:PSE

[TWB^{+}18] Dirac Twidwell, Carissa L. Wonkka, Christine H. Bielski, Craig R. Allen, David G. Angeler, Jacob Drozda, Ahjond S. Garmestani, Julia Johnson, Larkin A. Powell, and Caleb P. Roberts. The perpetual state of emergency that sacrifices protected areas in a changing climate. Conservation Biology, 32 (4):905–915, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Taylor:2016:LWJ

[TWZ16] Bron Taylor, Gretel Van Wieren, and Bernard Daley Zaleha. Lynn White, Jr. and the greening-of-religion hypothesis. Conservation Biology, 30(5):1000–1009, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tian:2019:DBP

[TXBW19] Dexin Tian, Yan Xie, Anthony D. Barnosky, and Fuwen Wei. Defining the balance point between conservation and development. Conservation Biology, 33(2):231–238, April 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Telleria:2016:CPB

[TYA16] Luciana Tellería and Mariela Yapu-Alcázar. Conservation priorities for Bolivian mammals: response to Peñaranda and Simonetti. Conservation Biology, 30(4):905–906, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [PS15].

Urbina-Cardona:2010:ENM

[UCFV10] J. Nicolás Urbina-Cardona and Oscar Flores-Villela. Ecologicalniche modeling and prioritization of conservation-area networks for Mexican herpetofauna. Modelado del nicho ecológico y priorización de redes de Áreas de conservación para la herpetofauna mexicana. Conservation Biology, 24(4):1031–1041, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Underwood:2010:ICA

[UDG10] Jared G. Underwood, Caterina D’Agrosa, and Leah R. Gerber. Identifying conservation areas on the basis of alternative distribution data sets. Identificación de Áreas de conservación con base en conjuntos alternativos de datos de distribución.

Conservation Biology, 24(1):162–170, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Usio:2013:EPD

[UIN^{+}13] Nisikawa Usio, Miho Imada, Megumi Nakagawa, Munemitsu Akasaka, and Noriko Takamura. Effects of pond draining on biodiversity and water quality of farm ponds. Conservation Biology, 28(1):1429–1438, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Uchida:2016:EIA

[UMK^{+}16] Shota Uchida, Hideaki Mori, Tsubasa Kojima, Kayo Hayama, Yuko Sakairi, and Satoshi Chiba. Effects of an invasive ant on land snails in the Ogasawara Islands. Conservation Biology, 30(6):1330–1337, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Valko:2018:CCE

[Val18] Orsolya Valkó. Concepts, challenges, and emerging themes of restoration ecology. Conservation Biology, 32(4):969–971, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vanderwerf:2012:ENH

[Van12] Eric A. Vanderwerf. Evolution of nesting height in an endangered Hawaiian forest bird in response to a non-native predator. Evolución de la altura de anidación en una especie de ave de bosque hawaiano en peligro como respuesta a un depredador no nativo. Conservation Biology, 26(5):905–911, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vercammen:2019:UPC

[VB19] Ans Vercammen and Mark Burgman. Untapped potential of collective intelligence in conservation and environmental decision making. Conservation Biology, 33(6):1247–1255, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vucetich:2015:EWN

[VBN15] John A. Vucetich, Jeremy T. Bruskotter, and Michael Paul Nelson. Evaluating whether nature’s intrinsic value is an axiom of or anathema to conservation. Conservation Biology, 29(2):321– 332, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

vonBieberstein:2019:ICI

[vBSC^{+}19] Katharina Rogalla von Bieberstein, Elsa Sattout, Mads Christensen, Balakrishna Pisupati, Neil D. Burgess, Jerry Harrison, and Jonas Geldmann. Improving collaboration in the implementation of global biodiversity conventions. Conservation Biology, 33(4):821–831, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vega:2013:CLP

[VCDF13] María Grande Vega, Bruno Carpinetti, Jesús Duarte, and Julia E. Fa. Contrasts in livelihoods and protein intake between commercial and subsistence bushmeat hunters in two villages on Bioko Island, Equatorial Guinea. Conservation Biology, 27 (3):576–587, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

vanderWal:2016:RAF

[vdWSM^{+}16] René van der Wal, Nirwan Sharma, Chris Mellish, Annie Robinson, and Advaith Siddharthan. The role of automated feedback in training and retaining biological recorders for citizen science. Conservation Biology, 31(1):550–561, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

vanEeden:2018:MCB

[vECD^{+}18] Lily M. van Eeden, Mathew S. Crowther, Chris R. Dickman, David W. Macdonald, William J. Ripple, Euan G. Ritchie, and Thomas M. Newsome. Managing conflict between large carnivores and livestock. Conservation Biology, 32(1):26–34, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vegvari:2018:ECC

[Vég18] Zsolt Végvári. The effects of climate change and human persecution of birds. Conservation Biology, 32(2):496–498, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Venn:2018:EUR

[Ven18] Stephen Venn. The ecology of urban and road infrastructure. Conservation Biology, 32(2):504–506, April 2018. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Verissimo:2016:WHW

[Ver16] Diogo Veríssimo. Winning hearts and wallets for biodiversity conservation. Conservation Biology, 30(6):1369–1370, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vera-Escalona:2018:PPF

[VESHR18] Iván Vera-Escalona, Shreeram Senthivasan, Evelyn Habit, and Daniel E. Ruzzante. Past, present, and future of a freshwater fish metapopulation in a threatened landscape. Conservation Biology, 32(4):849–859, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vitule:2012:RPC

[VFV^{+}12] Jean Ricardo Simões Vitule, Carolina A. Freire, Diego P. Vazquez, Martin A. Nuñez, and Daniel Simberloff. Revisiting the potential conservation value of non-native species. Conservation Biology, 26(6):1153–1155, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Viana:2019:DTO

[VGAB^{+}19] Daniel F. Viana, Stefan Gelcich, Erendira Aceves-Bueno, Becky Twohey, and Steven D. Gaines. Design trade-offs in rightsbased management of small-scale fisheries. Conservation Biology, 33(2):361–368, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vitule:2017:NDD

[Vit17] Jean Ricardo Simões Vitule. The need to dive deep in marine biodiversity. Conservation Biology, 31(3):731–733, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Visconti:2015:BRC

[VJ15] Piero Visconti and Lucas Joppa. Building robust conservation plans. Conservation Biology, 29(2):503–512, April 2015. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vilchis:2015:AEC

[VJE^{+}15] L. Ignacio Vilchis, Christine K. Johnson, Joseph R. Evenson, Scott F. Pearson, Karen L. Barry, Peter Davidson, Martin G. Raphael, and Joseph K. Gaydos. Assessing ecological correlates of marine bird declines to inform marine conservation. Conservation Biology, 29(1):154–163, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vos:2015:ENB

[VJG^{+}15] Jurriaan M. De Vos, Lucas N. Joppa, John L. Gittleman, Patrick R. Stephens, and Stuart L. Pimm. Estimating the normal background rate of species extinction. Conservation Biology, 29(2):452–462, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vranckx:2012:MAS

[VJMH12] Guy Vranckx, Hans Jacquemyn, Bart Muys, and Olivier Honnay. Meta-analysis of susceptibility of woody plants to loss of genetic diversity through habitat fragmentation. Meta–Análisis de la susceptibilidad de plantas Leñosas a la pérdida de diversidad genética por medio de la fragmentación del hábitat. Conservation Biology, 26(2):228–237, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vanthomme:2013:DCM

[VKKA13] Hadrien Vanthomme, Joseph Kolowski, Lisa Korte, and Alfonso Alonso. Distribution of a community of mammals in relation to roads and other human disturbances in Gabon, Central Africa. Conservation Biology, 27(2):281–291, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Verhagen:2017:UDS

[VKM^{+}17] Willem Verhagen, Aija S. Kukkala, Atte Moilanen, Astrid J. A. van Teeffelen, and Peter H. Verburg. Use of demand for and spatial flow of ecosystem services to identify priority areas. Conservation Biology, 31(4):860–871, August 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vidal:2014:TDF

[VLGRS14] Omar Vidal, José López-García, and Eduardo Rendón-Salinas. Trends in deforestation and forest degradation after a decade of monitoring in the Monarch butterfly Biosphere reserve in Mexico. Conservation Biology, 28(1):177–186, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Voight:2012:SGV

[VLRB12] Janet R. Voight, Raymond W. Lee, Abigail J. Reft, and Amanda E. Bates. Scientific gear as a vector for non-native species at deep-sea hydrothermal vents. Equipo científico como un vector para especies no nativas en conductos hidrotermales de aguas profundas. Conservation Biology, 26(5):938–942, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vallabh:2016:MEC

[VLSOS16] Priya Vallabh, Heila Lotz-Sisitka, Rob O’Donoghue, and Ingrid Schudel. Mapping epistemic cultures and learning potential of participants in citizen science projects. Conservation Biology, 31(1):540–549, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Valkonen:2014:RVI

[VM14] Janne K. Valkonen and Johanna Mappes. Resembling a viper: Implications of mimicry for conservation of the endangered smooth snake. Conservation Biology, 28(6):1568–1574, December 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See reply [BC18].

Valkonen:2018:RMS

[VM18] Janne K. Valkonen and Johanna Mappes. Resemblance of a model species and its mimic: Response to Bury and Cichoń. Conservation Biology, 32(2):487–489, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Visconti:2013:EEG

[VMÁR^{+}13] P. Visconti, M. Di Marco, J. G. Álvarez-Romero, S. R. Januchowski-Hartley, R. L. Pressey, R. Weeks, and C. Rondinini. Effects of errors and gaps in spatial data sets on assessment of conservation progress. Conservation Biology, 28(4): 1000–1010, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vliet:2010:ESS

[VMGB^{+}10] Nathalie Van Vliet, E. J. Milner-Gulland, François Bousquet, Mehdi Saqalli, and Robert Nasi. Effect of small-scale heterogeneity of prey and hunter distributions on the sustainability of bushmeat hunting. Efecto de la heterogeneidad a pequeña escala de la distribución de presas y cazadores sobre la sustentabilidad de la cacería de carne de monte. Conservation Biology, 24(5):1327–1337, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Venter:2018:BPA

[VMO^{+}18] Oscar Venter, Ainhoa Magrach, Nick Outram, Carissa Joy Klein, Hugh P. Possingham, Moreno Di Marco, and James E. M. Watson. Bias in protected-area location and its effects on long-term aspirations of biodiversity conventions. Conservation Biology, 32(1):127–134, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Venesky:2012:STA

[VMS^{+}12] Matthew D. Venesky, Joseph R. Mendelson III, Brittany F. Sears, Peter Stiling, and Jason R. Rohr. Selecting for tolerance against pathogens and herbivores to enhance success of reintroduction and translocation. Selección de tolerancia a patógenos y Herbívoros para incrementar el Éxito de la reintroducció y la translocación. Conservation Biology, 26(4):586–592, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Valderrama:2013:EPS

[VMW13] Sandra V. Valderrama, Laura E. Molles, and Joseph R. Waas. Effects of population size on singing behavior of a rare duetting songbird. Conservation Biology, 27(1):210–218, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vanthomme:2019:ESB

[VNAT19] Hadrien P. A. Vanthomme, Brave S. Nzamba, Alfonso Alonso, and Angelique F. Todd. Empirical selection between least-cost and current-flow designs for establishing wildlife corridors in Gabon. Conservation Biology, 33(2):329–338, April 2019. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Varkey:2013:BDN

[VPMS13] Divya A. Varkey, Tony J. Pitcher, Murdoch K. Mcallister, and Rashid S. Sumaila. Bayesian decision-network modeling of multiple stakeholders for reef ecosystem restoration in the coral triangle. Conservation Biology, 27(3):459–469, June 2013. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Verissimo:2014:USA

[VPS^{+}14] Diogo Veríssimo, Tatiana Pongiluppi, Maria Cintia M. Santos, Pedro F. Develey, Iain Fraser, Robert J. Smith, and Douglas C. Macmilan. Using a systematic approach to select flagship species for bird conservation. Conservation Biology, 28(1): 269–277, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Veloz:2015:IES

[VSA^{+}15] Samuel Veloz, Leonardo Salas, Bob Altman, John Alexander, Dennis Jongsomjit, Nathan Elliott, and Grant Ballard. Improving effectiveness of systematic conservation planning with density data. Conservation Biology, 29(4):1217–1227, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vynne:2011:ESD

[VSM^{+}11] Carly Vynne, John R. Skalski, Ricardo B. Machado, Martha J. Groom, Anah T. A. Jácomo, Jader Marinho-Filho, Mario B. Ramos Neto, Cristina Pomilla, Leandro Silveira, Heath Smith, and Samuel K. Wasser. Effectiveness of scat-detection dogs in determining species presence in a tropical savanna landscape. Efectividad de perros detectores de excretas para la determinación de la presencia de especies en un paisaje de sabana tropical. Conservation Biology, 25(1):154–162, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Valle:2019:RPD

[VTM19] Denis Valle, Kok Ben Toh, and Justin Millar. Rapid prototyping of decision-support tools for conservation. Conservation Biology, 33(6):1448–1450, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vanthomme:2017:ASO

[VTT^{+}17] Hadrien P. A. Vanthomme, Elie Tobi, Angelique F. Todd, Lisa Korte, and Alfonso Alonso. Antipoaching standards in onshore hydrocarbon concessions drawn from a central African case study. Conservation Biology, 31(3):696–706, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Verissimo:2019:CER

[VW19] Diogo Veríssimo and Anita K. Y. Wan. Characterizing efforts to reduce consumer demand for wildlife products. Conservation Biology, 33(3):623–633, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Venter:2010:AUO

[VWM^{+}10] Oscar Venter, James E. M. Watson, Erik Meijaard, William F. Laurance, and Hugh P. Possingham. Avoiding unintended outcomes from REDD. Conservation Biology, 24(1):5–6, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

vonWehrden:2015:CTC

[vWWCF15] Henrik von Wehrden, Karsten Wesche, Oyundari Chuluunkhuyag, and Pascal Fust. Correlation of trends in cashmere production and declines of large wild mammals: response to Berger et al. 2013. Conservation Biology, 29(1):286–289, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weeks:2018:RPC

[WA18] R. Weeks and V. M. Adams. Research priorities for conservation and natural resource management in Oceania’s smallisland developing states. Conservation Biology, 32(1):72–83, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wallen:2017:FSP

[Wal17] Kenneth E. Wallen. Focusing on structure and process to integrate and mainstream the social sciences in conservation. Conservation Biology, 31(3):724–726, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waples:2013:I

[Wap13] Robin S. Waples. Introduction. Conservation Biology, 27(6): 1137, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watson:2010:CRC

[Wat10] James Watson. Complexity, resilience, and conservation. Conservation Biology, 24(4):1166–1167, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waterhouse:2011:FBB

[Wat11] Jon Waterhouse. Finding balance between human need and global stewardship. Conservation Biology, 25(6):1087–1090, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wiens:2010:MMS

[WB10] John A. Wiens and Dominique Bachelet. Matching the multiple scales of conservation with the multiple scales of climate change. Adecuación de las múltiples escalas de conservación con las múltiples escalas de cambio climático. Conservation Biology, 24(1):51–62, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watson:2014:TOA

[WB14] Jordan T. Watson and Keith A. Bigelow. Trade-offs among catch, bycatch, and landed value in the American Samoa longline fishery. Conservation Biology, 28(4):1012–1022, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wallach:2018:SCA

[WBB^{+}18] Arian D. Wallach, Marc Bekoff, Chelsea Batavia, Michael Paul Nelson, and Daniel Ramp. Summoning compassion to address the challenges of conservation. Conservation Biology, 32(6): 1255–1265, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wheeler:2019:IKN

[WBF^{+}19] Helen C. Wheeler, Dominique Berteaux, Chris Furgal, Kevin Cazelles, Nigel G. Yoccoz, and David Grémillet. Identifying key needs for the integration of social–ecological outcomes in Arctic wildlife monitoring. Conservation Biology, 33(4):861– 872, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walters:2013:IEW

[WBM13] Annika W. Walters, Krista K. Bartz, and Michelle M. Mcclure. Interactive effects of water diversion and climate change for juvenile Chinook salmon in the Lemhi River Basin (U.S.A.). Conservation Biology, 27(6):1179–1189, December 2013. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wallach:2015:PPC

[WBNR15] Arian D. Wallach, Marc Bekoff, Michael Paul Nelson, and Daniel Ramp. Promoting predators and compassionate conservation. Conservation Biology, 29(5):1481–1484, October 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Westgate:2015:TAT

[WBPL15] Martin J. Westgate, Philip S. Barton, Jennifer C. Pierson, and David B. Lindenmayer. Text analysis tools for identification of emerging topics and research gaps in conservation science. Conservation Biology, 29(6):1606–1614, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Willoughby:2019:LTD

[WC19] Janna R. Willoughby and Mark R. Christie. Long-term demographic and genetic effects of releasing captive-born individuals into the wild. Conservation Biology, 33(2):377–388, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whitlock:2018:LUH

[WCCT18] Cathy Whitlock, Daniele Colombaroli, Marco Conedera, and Willy Tinner. Land-use history as a guide for forest conservation and management. Conservation Biology, 32(1):84–97, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wittmann:2015:USE

[WCR^{+}15] Marion E. Wittmann, Roger M. Cooke, John D. Rothlisberger, Edward S. Rutherford, Hongyan Zhang, Doran M. Mason, and David M. Lodge. Use of structured expert judgment to forecast invasions by bighead and silver carp in Lake Erie. Conservation Biology, 29(1):187–197, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wyner:2010:TCB

[WD10] Yael Wyner and Rob Desalle. Taking the conservation biology perspective to secondary school classrooms. Introducción de la perspectiva de la biología de la conservación a las aulas de escuelas secundarias. Conservation Biology, 24(3):649–654, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Welch-Devine:2012:SPC

[WD12] Meredith Welch-Devine. Seeing people in conservation. Conservation Biology, 26(2):374–376, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilkinson:2017:RAL

[WD17] Nicholas M. Wilkinson and Luong Van Duc. Rank aggregation of local expert knowledge for conservation planning of the critically endangered saola. Conservation Biology, 31(3):625– 634, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walsh:2015:ESE

[WDS15] Jessica C. Walsh, Lynn V. Dicks, and William J. Sutherland. The effect of scientific evidence on conservation practitioners’ management decisions. Conservation Biology, 29(1):88–98, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watson:2016:BSN

[WDV^{+}16] James E. M. Watson, Emily S. Darling, Oscar Venter, Martine Maron, Joe Walston, Hugh P. Possingham, Nigel Dudley, Marc Hockings, Megan Barnes, and Thomas M. Brooks. Bolder science needed now for protected areas. Conservation Biology, 30 (2):243–248, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wasserman:2019:SSS

[WDW^{+}19] Ryan J. Wasserman, Jaimie T. A. Dick, Rebecca J. Welch, Tatenda Dalu, and Kit Magellan. Site and species selection for religious release of non-native fauna. Conservation Biology, 33 (4):969–971, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watson:2011:CAP

[WEC^{+}11] James E. M. Watson, Megan C. Evans, Josie Carwardine, Richard A. Fuller, Liana N. Joseph, Dan B. Segan, Martin F. J. Taylor, R. J. Fensham, and Hugh P. Possingham. The capacity of Australia’s protected-area system to represent threatened species. La capacidad del sistema de Áreas protegidas de Australia para representar especies amenazadas. Conservation Biology, 26(3):324–332, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wedekind:2013:PUS

[WES^{+}13] Claus Wedekind, Guillaume Evanno, Tamás Székely, Manuel Pompini, Olivier Darbellay, and Joachim Guthruf. Persistent unequal sex ratio in a population of grayling (salmonidae) and possible role of temperature increase. Conservation Biology, 27 (1):229–234, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weimerskirch:2018:URD

[WFC^{+}18] H. Weimerskirch, D. P. Filippi, J. Collet, S. M. Waugh, and S. C. Patrick. Use of radar detectors to track attendance of albatrosses at fishing vessels. Conservation Biology, 32(1): 240–245, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waylen:2010:ELC

[WFM^{+}10] Kerry A. Waylen, Anke Fischer, Philip J. K. Mcgowan, Simon J. Thirgood, and E. J. Milner-Gulland. Effect of local cultural context on the success of community-based conservation interventions. Efecto del contexto cultural local sobre el Éxito de intervenciones de conservación basadas en comunidades. Conservation Biology, 24(4):1119–1129, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whytock:2018:BCR

[WFMW^{+}18] Robin C. Whytock, Elisa Fuentes-Montemayor, Kevin Watts, Patanjaly Barbosa De Andrade, Rory T. Whytock, Paul French, Nicholas A. Macgregor, and Kirsty J. Park. Birdcommunity responses to habitat creation in a long-term, largescale natural experiment. Conservation Biology, 32(2):345–354, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wauchope:2019:RII

[WFSS19] Hannah S. Wauchope, Richard A. Fuller, Danielle F. Shanahan, and Justine D. Shaw. Restoring islands and identifying source populations for introductions. Conservation Biology, 33 (3):729–732, June 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Williams:2015:PPC

[WG15] Sophie J. Williams and Holly L. Gordon. Paradigms and progress in conservation science. Conservation Biology, 29 (1):3–4, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

White:2011:CC

[WGB^{+}11a] Joseph D. White, Kevin J. Gutzwiller, Wylie C. Barrow, Lori Johnson-Randall, Lisa Zygo, and Pamela Swint. Cover caption. Conservation Biology, 25(3):i, December 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

White:2011:UIE

[WGB^{+}11b] Joseph D. White, Kevin J. Gutzwiller, Wylie C. Barrow, Lori Johnson-Randall, Lisa Zygo, and Pamela Swint. Understanding interaction effects of climate change and fire management on bird distributions through combined process and habitat models. Comprensión de los efectos de la interacción del cambio climático y el manejo de fuego sobre la distribución de aves mediante modelos combinados de procesos y hábitat. Conservation Biology, 26(2):536–546, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilson:2012:EME

[WGF^{+}12] Shaun K. Wilson, Nicholas A. J. Graham, Rebecca Fisher, Jan Robinson, Kirsty Nash, Karen Chong-Seng, Nicholas V. C. Polunin, Riaz Aumeeruddy, and Rodney Quatre. Effect of macroalgal expansion and Marine Protected Areas on coral recovery following a climatic disturbance. Conservation Biology, 26(6):995–1004, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weiser:2013:SRR

[WGJ13] Emily L. Weiser, Catherine E. Grueber, and Ian G. Jamieson. Simulating retention of rare alleles in small populations to assess management options for species with different life histories. Conservation Biology, 27(2):335–344, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Woinarski:2017:CPL

[WGLL17] John C. Z. Woinarski, Stephen T. Garnett, Sarah M. Legge, and David B. Lindenmayer. The contribution of policy, law, management, research, and advocacy failings to the recent extinctions of three Australian vertebrate species. Conservation Biology, 31(1):13–23, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wolak:2010:CSL

[WGR^{+}10] Matthew E. Wolak, George W. Gilchrist, Victoria A. Ruzicka, Daniel M. Nally, and Randolph M. Chambers. A contemporary, sex-limited change in body size of an estuarine turtle in response to commercial fishing. Un cambio contemporáneo, limitado por el sexo, en el tamaño corporal de una tortuga estuarina en respuesta a la pesca comercial. Conservation Biology, 24 (5):1268–1277, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watson:2019:EHD

[WGS^{+}19] Keri B. Watson, Gillian L. Galford, Laura J. Sonter, Insu Koh, and Taylor H. Ricketts. Effects of human demand on conservation planning for biodiversity and ecosystem services. Conservation Biology, 33(4):942–952, August 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Willson:2013:EEA

[WH13] John D. Willson and William A. Hopkins. Evaluating the effects of anthropogenic stressors on source-sink dynamics in pond-breeding amphibians. Conservation Biology, 27(3):595– 604, June 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Williams:2011:CBW

[WHB^{+}11] Rob Williams, Sharon L. Hedley, Trevor A. Branch, Mark V. Bravington, Alexandre N. Zerbini, and Ken P. Findlay. Chilean blue whales as a case study to illustrate methods to estimate abundance and evaluate conservation status of rare species. Ballenas azules Chilenas como caso de estudio para ilustrar métodos para estimas la abundancia y evaluar el estatus de conservación de especies raras. Conservation Biology, 25(3):526– 535, June 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waylen:2014:NDK

[WHB^{+}14] K. A. Waylen, E. J. Hastings, E. A. Banks, K. L. Holstead, R. J. Irvine, and K. L. Blackstock. The need to disentangle key concepts from ecosystem-approach jargon. Conservation Biology, 28(6):1215–1224, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wheat:2011:LMT

[Whe11] Christopher W. Wheat. Let’s move toward genomics for butterfly ecology and insect conservation. Conservation Biology, 25(1):203–205, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

West:2017:IAA

[WHE^{+}17] Catherine West, Courtney A. Hofman, Steve Ebbert, John Martin, Sabrina Shirazi, Samantha Dunning, and Jesus E. Maldonado. Integrating archaeology and ancient DNA analysis to address invasive species colonization in the Gulf of Alaska. Conservation Biology, 31(5):1163–1172, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilson:2016:TTS

[WHEN^{+}16] Robyn S. Wilson, David J. Hardisty, Rebecca S. Epanchin-Niell, Michael C. Runge, Kathryn L. Cottingham, Dean L. Urban, Lynn A. Maguire, Alan Hastings, Peter J. Mumby, and Debra P. C. Peters. A typology of time-scale mismatches and behavioral interventions to diagnose and solve conservation problems. Conservation Biology, 30(1):42–49, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilcox:2015:USE

[WHG^{+}15] Chris Wilcox, Grace Heathcote, Jennifer Goldberg, Riki Gunn, David Peel, and Britta Denise Hardesty. Understanding the sources and effects of abandoned, lost, and discarded fishing gear on marine turtles in northern Australia. Conservation Biology, 29(1):198–206, February 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whitten:2010:CC

[Whi10] Tony Whitten. Caves and conservation. Conservation Biology, 25(6):912–913, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

White:2013:DEV

[Whi13] Peter S. White. Derivation of the extrinsic values of biological diversity from its intrinsic value and of both from the first principles of evolution. Conservation Biology, 27(6):1279–1285, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wade:2017:AAC

[WHK^{+}17] Alisa A. Wade, Brian K. Hand, Ryan P. Kovach, Gordon Luikart, Diane C. Whited, and Clint C. Muhlfeld. Accounting for adaptive capacity and uncertainty in assessments of species’ climate-change vulnerability. Conservation Biology, 31 (1):136–149, February 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wright:2016:RCA

[WHR^{+}16] Juliet H. Wright, Nicholas A. O. Hill, Dilys Roe, J. Marcus Rowcliffe, Noëlle F. Kümpel, Mike Day, Francesca Booker, and E. J. Milner-Gulland. Reframing the concept of alternative livelihoods. Conservation Biology, 30(1):7–13, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ward:2018:PBD

[WHS18] Caroline Ward, George Holmes, and Lindsay Stringer. Perceived barriers to and drivers of community participation in protected-area governance. Conservation Biology, 32(2):437– 446, April 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilhere:2012:IAD

[Wil12a] George F. Wilhere. Inadvertent advocacy defensa involuntaria. Conservation Biology, 26(2):39–46, February 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comment [CRNR12].

Wilhere:2012:IPA

[Wil12b] George F. Wilhere. Inadvertent policy advocacy in peer review of recovery plans: Reply to Carroll et al. Conservation Biology, 27(1):746–748, August 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilhere:2017:RSS

[Wil17] George F. Wilhere. The role of scientists in statutory interpretation of the U.S. Endangered Species Act. Conservation Biology, 31(2):252–260, April 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wanger:2010:ELU

[WIM^{+}10] Thomas C. Wanger, Djoko T. Iskandar, Iris Motzke, Barry W. Brook, Navjot S. Sodhi, Yann Clough, and Teja Tscharntke. Effects of land-use change on community composition of tropical amphibians and reptiles in Sulawesi, Indonesia. Efectos del cambio de uso de suelo sobre la Composición de la comunidad de anfibios y reptiles en Sulawesi, Indonesia. Conservation Biology, 25(6):795–802, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wittemyer:2011:EED

[Wit11] George Wittemyer. Effects of economic downturns on mortality of wild African elephants. Efectos de las recesiones económicas sobre la mortalidad de elefantes africanos silvestres. Conservation Biology, 27(3):1002–1009, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walters:2010:QPC

[WJ10] Eric L. Walters and Frances C. James. Quantifying purported competition with individual- and population-level metrics. Cuantificación de la competencia aparente con medidas a nivel individual y poblacional. Conservation Biology, 24(6): 1569–1577, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weeks:2013:ACM

[WJ13] Rebecca Weeks and Stacy D. Jupiter. Adaptive comanagement of a Marine Protected Area network in Fiji. Conservation Biology, 27(6):1234–1244, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wedekind:2010:SSS

[WK10] Claus Wedekind and Christoph Küng. Shift of spawning season and effects of climate warming on developmental stages of a grayling (Salmonidae). Cambio en la Época de desove y efectos del calentamiento climático sobre etapas de desarrollo de un tímalo (Salmonidae). Conservation Biology, 24(5):1418–1423, October 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wright:2015:PMC

[WK15] Andrew J. Wright and Line A. Kyhn. Practical management of cumulative anthropogenic impacts with working marine examples. Conservation Biology, 29(2):333–340, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilson:2011:AVR

[WKF^{+}11] Howard B. Wilson, Bruce E. Kendall, Richard A. Fuller, David A. Milton, and Hugh P. Possingham. Analyzing variability and the rate of decline of migratory shorebirds in Moreton Bay, Australia. Análisis de la variabilidad y tasa de declinación de aves vadeadoras migratorias en la Bahía Moreton, Australia. Conservation Biology, 25(4):758–766, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whitehead:2014:IBS

[WKI^{+}14] Amy L. Whitehead, Heini Kujala, Christopher D. Ives, Ascelin Gordon, Pia E. Lentini, Brendan A. Wintle, Emily Nicholson, and Christopher M. Raymond. Integrating biological and social values when prioritizing places for biodiversity conservation. Conservation Biology, 28(4):992–1003, August 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilson:2011:VPA

[WKP11] Howard B. Wilson, Bruce E. Kendall, and Hugh P. Possingham. Variability in population abundance and the classification of extinction risk. Variabilidad de la abundancia poblacional y la Clasificación del riesgo de extinción. Conservation Biology, 25 (4):747–757, August 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wallmo:2012:PWP

[WL12] Kristy Wallmo and Daniel K. Lew. Public willingness to pay for recovering and downlisting threatened and endangered marine species. Disponibilidad del Público para pagar por la recuperación y reclasificación de especies marinas amenazadas y en peligro. Conservation Biology, 26(5):830–839, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wehi:2017:IIC

[WL17a] Priscilla M. Wehi and Janice M. Lord. Importance of including cultural practices in ecological restoration. Conservation Biology, 31(5):1109–1118, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Westgate:2017:DSR

[WL17b] Martin J. Westgate and David B. Lindenmayer. The difficulties of systematic reviews. Conservation Biology, 31(5):1002–1007, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilsey:2014:DEB

[WLCS14] Chad B. Wilsey, Joshua J. Lawler, David Cimprich, and Nathan H. Schumaker. Dependence of the endangered blackcapped vireo on sustained cowbird management. Conservation Biology, 28(2):561–571, April 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Woodruff:2016:ESP

[WLCW16] Susannah P. Woodruff, Paul M. Lukacs, David Christianson, and Lisette P. Waits. Estimating Sonoran pronghorn abundance and survival with fecal DNA and capture–recapture methods. Conservation Biology, 30(5):1102–1111, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wald:2016:DPE

[WLD16] Dara M. Wald, Justin Longo, and A. R. Dobell. Design principles for engaging and retaining virtual citizen scientists.

Conservation Biology, 31(1):562–570, June 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wald:2016:CCR

[WLL^{+}16] Dara M. Wald, Cheryl A. Lohr, Christopher A. Lepczyk, Susan K. Jacobson, and Linda J. Cox. A comparison of cat-related risk perceptions and tolerance for outdoor cats in Florida and Hawaii. Conservation Biology, 30(6):1233–1244, December 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wallach:2018:IM

[WLRR18] Arian D. Wallach, Erick J. Lundgren, William J. Ripple, and Daniel Ramp. Invisible megafauna. Conservation Biology, 32 (4):962–965, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

White:2015:SEC

[WMFB15] Easton R. White, Mark C. Myers, Joanna Mills Flemming, and Julia K. Baum. Shifting elasmobranch community assemblage at Cocos Island — an isolated marine protected area. Conservation Biology, 29(4):1186–1197, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wee:2019:IAG

[WML^{+}19] Alison K. S. Wee, Gustavo M. Mori, Catarina F. Lira, Juan Núñez-Farfán, Koji Takayama, Leanne Faulks, Suhua Shi, Yoshiaki Tsuda, Yoshihisa Suyama, Takashi Yamamoto, Takaya Iwasaki, Yukio Nagano, Zhengzhen Wang, Shin Watanabe, and Tadashi Kajita. The integration and application of genomic information in mangrove conservation. Conservation Biology, 33(2):206–209, February 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilhere:2012:CVS

[WMS^{+}12] George F. Wilhere, Lynn A. Maguire, J. Michael Scott, Janet L. Rachlow, Dale D. Goble, and Leona K. Svancara. Conflation of values and science: Response to Noss et al. Conservation Biology, 26(5):943–944, October 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilmshurst:2014:UPA

[WMW^{+}14] Janet M. Wilmshurst, Neville T. Moar, Jamie R. Wood, Peter J. Bellingham, Amy M. Findlater, James J. Robinson, and Clive Stone. Use of pollen and ancient DNA as conservation baselines for offshore islands in New Zealand. Conservation Biology, 28(1):202–212, February 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wheatall:2013:IEP

[WNY13] Laura Wheatall, Tim Nuttle, and Ellen Yerger. Indirect effects of pandemic deer overabundance inferred from caterpillarhost relations. Conservation Biology, 28(4):1107–1116, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wang:2016:ODC

[WÖ16] Yicheng Wang and Hayri Önal. Optimal design of compact and connected nature reserves for multiple species. Conservation Biology, 30(2):413–424, April 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilder:2013:NNG

[WON^{+}13] Benjamin T. Wilder, Carolyn O’Meara, Nemer Narchi, Alfonso Medel Narváez, and Octavio Aburto-Oropeza. The need for a next generation of Sonoran desert researchers. Conservation Biology, 27(2):243–245, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilgen:2012:RCP

[WR12] Nicola J. Van Wilgen and David M. Richardson. The roles of climate, phylogenetic relatedness, introduction effort, and reproductive traits in the establishment of non-native reptiles and amphibians. El papel del clima, la Relación filogenética, el esfuerzo de introducción y las características reproductivas en el establecimiento de reptiles y anfibios no nativos. Conservation Biology, 26(2):267–277, April 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weeks:2010:EMP

[WRAW10] Rebecca Weeks, Garry R. Russ, Angel C. Alcala, and Alan T. White. Effectiveness of Marine Protected Areas in the Philippines for biodiversity conservation. Efectividad de las Áreas Marinas Protegidas en las filipinas para la conservación de biodiversidad. Conservation Biology, 24(2):531–540, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wise:2012:CEN

[WRG^{+}12] Russell M. Wise, Belinda Reyers, Chris Guo, Guy F. Midgley, and Willem De Lange. Costs of expanding the network of protected areas as a response to climate change in the Cape Floristic region. Costos de la expansión de la red de Áreas protegidas como una respuesta al cambio climático en la Región Florística del Cabo. Conservation Biology, 26(3):397–407, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wrege:2010:UAT

[WRTB10] Peter H. Wrege, Elizabeth D. Rowland, Bruce G. Thompson, and Nikolas Batruch. Use of acoustic tools to reveal otherwise cryptic responses of forest elephants to oil exploration. Utilización de herramientas Acústicas para revelar respuestas Crípticas de elefantes a la exploración petrolera. Conservation Biology, 24(6):1578–1585, December 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Witter:2014:ILL

[WS14] Rebecca Witter and Terre Satterfield. Invisible losses and the logics of resettlement compensation. Conservation Biology, 28 (6):1394–1402, October 2014. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Willette:2017:UDB

[WSC^{+}17] Demian A. Willette, Sara E. Simmonds, Samantha H. Cheng, Sofia Esteves, Tonya L. Kane, Hayley Nuetzel, Nicholas Pilaud, Rita Rachmawati, and Paul H. Barber. Using DNA barcoding to track seafood mislabeling in Los angeles restaurants. Conservation Biology, 31(5):1076–1085, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wittmer:2013:CSS

[WSEM13] Heiko U. Wittmer, Robert Serrouya, L. Mark Elbroch, and Andrew J. Marshall. Conservation strategies for species affected by apparent competition. Conservation Biology, 27(2):254–260, April 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wei:2015:PEC

[WSH^{+}15] Fuwen Wei, Ronald Swaisgood, Yibo Hu, Yonggang Nie, Li Yan, Zejun Zhang, Dunwu Qi, and Lifeng Zhu. Progress in the ecology and conservation of giant pandas. Conservation Biology, 29(6):1497–1507, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ward:2011:EML

[WSHB11] Eric J. Ward, Brice X. Semmens, Elizabeth E. Holmes, and Ken C. Balcomb III. Effects of multiple levels of social organization on survival and abundance. Efectos de niveles múltiples de organización social sobre la supervivencia y abundancia. Conservation Biology, 26(3):350–355, April 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilson:2016:ADB

[WSLW16] John-James Wilson, Kong-Wah Sing, Ping-Shin Lee, and Alison K. S. Wee. Application of DNA barcodes in wildlife conservation in Tropical East Asia. Conservation Biology, 30(5): 982–989, October 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wade:2010:RDE

[WT10] Alisa A. Wade and David M. Theobald. Residential development encroachment on U.S. protected areas. Invasión de Áreas protegidas en E.U.A. por el desarrollo residencial. Conservation Biology, 24(1):151–161, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

White:2013:ESF

[WT13] Peter S. White and Julie P. Tuttle. Ecological sustainability as the fourth landmark in the development of conservation ethics. Conservation Biology, 28(4):952–957, October 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Webb:2017:IIF

[WTT^{+}17] Matthew H. Webb, Aleks Terauds, Ayesha Tulloch, Phil Bell, Dejan Stojanovic, and Robert Heinsohn. The importance of incorporating functional habitats into conservation planning for highly mobile species in dynamic systems. Conservation Biology, 31(5):1018–1028, October 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wehi:2010:TPH

[WW10a] Priscilla M. Wehi and William L. Wehi. Traditional plant harvesting in contemporary fragmented and urban landscapes. Cosecha tradicional de plantas en paisajes contemporáneos fragmentados y urbanos. Conservation Biology, 24(2):594–604, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wissel:2010:CAA

[WW10b] Silvia Wissel and Frank Wätzold. A conceptual analysis of the application of tradable permits to biodiversity conservation. Un análisis conceptual de la aplicación de permisos negociables en la conservación de la biodiversidad. Conservation Biology, 24 (2):404–411, April 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilman:2016:MOA

[WW16] Elizabeth A. Wilman and Elspeth N. Wilman. Modeling outcomes of approaches to sustained human and snow leopard coexistence. Conservation Biology, 30(1):50–58, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wadgymar:2017:PME

[WW17] Susana M. Wadgymar and Arthur E. Weis. Phenological mismatch and the effectiveness of assisted gene flow. Conservation Biology, 31(3):547–558, June 2017. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wenger:2016:ERW

[WWdS^{+}16] Amelia S. Wenger, David H. Williamson, Eduardo T. da Silva, Daniela M. Ceccarelli, Nicola K. Browne, Caroline Petus, and Michelle J. Devlin. Effects of reduced water quality on coral reefs in and out of no-take marine reserves. Conservation Biology, 30(1):142–153, February 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wang:2018:SRB

[WWH^{+}18] Junjun Wang, Ruidong Wu, Daming He, Feiling Yang, Peijun Hu, Shiwei Lin, Wei Wu, Yixin Diao, and Yang Guo. Spatial relationship between climatic diversity and biodiversity conservation value. Conservation Biology, 32(6):1266–1277, December 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wood:2012:LLB

[WWW^{+}12] Jamie R. Wood, Janet M. Wilmshurst, Trevor H. Worthy, Avi S. Holzapfel, and Alan Cooper. A lost link between a flightless parrot and a parasitic plant and the potential role of coprolites in conservation paleobiology. Conservation Biology, 26(6):1091–1099, December 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilson:2015:UBS

[WWWC15] A. D. M. Wilson, M. Wikelski, R. P. Wilson, and S. J. Cooke. Utility of biological sensor tags in animal conservation. Conservation Biology, 29(4):1065–1075, August 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watson:2019:EBC

[WZC19] David M. Watson, Elizabeth Znidersic, and Michael D. Craig. Ethical birding call playback and conservation. Conservation Biology, 33(2):469–471, April 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watkins:2018:DIC

[WZWF18] Cristy Watkins, Jennifer Zavaleta, Sarah Wilson, and Scott Francisco. Developing an interdisciplinary and cross-sectoral community of practice in the domain of forests and livelihoods. Conservation Biology, 32(1):60–71, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yap:2011:PFC

[Yap11] Helen T. Yap. A perilous future for coral reefs. Conservation Biology, 25(1):205–206, February 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Young:2015:DCT

[YFB15] Matthew A. L. Young, Simon Foale, and David R. Bellwood. Dynamic catch trends in the history of recreational spearfishing in Australia. Conservation Biology, 29(3):784–794, June 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yong:2018:COT

[YJL^{+}18] Ding Li Yong, Anuj Jain, Yang Liu, Muhammad Iqbal, Chang-Yong Choi, Nicola J. Crockford, Spike Millington, and Jennifer Provencher. Challenges and opportunities for transboundary conservation of migratory birds in the East Asian–Australasian flyway. Conservation Biology, 32(3):740–743, June 2018. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yanco:2019:CAO

[YNR19] Esty Yanco, Michael Paul Nelson, and Daniel Ramp. Cautioning against overemphasis of normative constructs in conservation decision making. Conservation Biology, 33(5):1002–1013, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zablocki:2016:FAM

[ZAB16] John Zablocki, Siddharth Arora, and Maan Barua. Factors affecting media coverage of species rediscoveries. Conservation Biology, 30(4):914–917, August 2016. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zack:2010:MBA

[Zac10] Steve Zack. Moving birds around, at great peril. Conservation Biology, 24(4):1164–1165, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zavaleta:2018:ECP

[ZAP^{+}18] Erika Zavaleta, Clare Aslan, Wendy Palen, Tom Sisk, Maureen E. Ryan, and Brett G. Dickson. Expanding career pathways in conservation science. Conservation Biology, 32(1): 246–248, February 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zitske:2011:NEH

[ZBD11] B. P. Zitske, M. G. Betts, and A. W. Diamond. Negative effects of habitat loss on survival of migrant warblers in a forest mosaic. Efectos negativos de la pérdida de hábitat sobre la supervivencia de Parúlidos migratorios en un mosaico forestal. Conservation Biology, 27(3):993–1001, October 2011. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zamin:2010:NRL

[ZBM^{+}10] Tara J. Zamin, Jonathan E. M. Baillie, Rebecca M. Miller, Jon Paul Rodríguez, Ana Ardid, and Ben Collen. National red listing beyond the 2010 target. Listas rojas nacionales más allá de la Meta 2010. Conservation Biology, 24(4):1012–1020, August 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zeigler:2013:APU

[ZCCN13] Sara L. Zeigler, Judy P. Che-Castaldo, and Maile C. Neel. Actual and potential use of population viability analyses in recovery of plant species listed under the U.S. Endangered Species Act. Conservation Biology, 27(6):1265–1278, December 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ziegler:2011:RCR

[ZFW^{+}11] Alan D. Ziegler, Jeff M. Fox, Edward L. Webb, Christine Padoch, Steve J. Leisz, Rob A. Cramb, Ole Mertz, Thilde B. Bruun, and Tran Duc Vien. Recognizing contemporary roles of swidden agriculture in transforming landscapes of Southeast Asia. Conservation Biology, 25(4):846–848, August 2011. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhang:2015:SSB

[Zha15] Li Zhang. Special section: Balancing conservation and development to preserve China’s biodiversity. Conservation Biology, 29 (6):1496, December 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zylberberg:2013:VLU

[ZLKW13] Maxine Zylberberg, Kelly A. Lee, Kirk C. Klasing, and Martin Wikelski. Variation with land use of immune function and prevalence of avian pox in Galapagos finches. Conservation Biology, 27(1):103–112, February 2013. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zommers:2012:PAF

[ZM12] Zinta Zommers and David W. Macdonald. Protected areas as frontiers for human migration. Áreas protegidas como fronteras para la migración humana. Conservation Biology, 26(3):547– 556, June 2012. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zorrilla:2015:SFP

[ZMTR15] Irene Zorrilla, Rosa Martinez, Mark A. Taggart, and Ngaio Richards. Suspected flunixin poisoning of a wild Eurasian griffon vulture from Spain. Conservation Biology, 29(2):587–592, April 2015. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zuckerberg:2018:CMS

[ZP18] Benjamin Zuckerberg and Jonathan N. Pauli. Conserving and managing the subnivium. Conservation Biology, 32(4):774–781, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zeng:2019:ICN

[ZPJ19] Yan Zeng, Xiaoge Ping, and Zhigang Jiang. Inertia in CITES nomenclature. Conservation Biology, 33(5):991–992, October 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zuckerberg:2018:ETP

[ZRM18] Benjamin Zuckerberg, Christine A. Ribic, and Lisa A. Mc-Cauley. Effects of temperature and precipitation on grassland bird nesting success as mediated by patch size. Conservation Biology, 32(4):872–882, August 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zabala:2018:WHU

[ZSM18] Aiora Zabala, Chris Sandbrook, and Nibedita Mukherjee. When and how to use Q methodology to understand perspectives in conservation research. Conservation Biology, 33(1): 1185–1194, October 2018. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhang:2019:UCH

[ZV19] Xiong Zhang and Amanda C. J. Vincent. Using cumulative human-impact models to reveal global threat patterns for seahorses. Conservation Biology, 33(6):1380–1391, December 2019. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhang:2010:SIH

[ZYYH10] Jin-Ju Zhang, Qi-Gang Ye, Xiao-Hong Yao, and Hong-Wen Huang. Spontaneous interspecific hybridization and patterns of pollen dispersal in ex situ populations of a tree species (Sinojackia xylocarpa) that is extinct in the wild. Hibridación interespecífica espontánea y patrones de dispersión de polen en poblaciones ex situ de una especie de Árbol (Sinojackia xylocarpa) extinta en su medio. Conservation Biology, 26(2): 246–255, February 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhu:2010:CID

[ZZW^{+}10] Lifeng Zhu, Xiangjiang Zhan, Hua Wu, Shanning Zhang, Tao Meng, Michael W. Bruford, and Fuwen Wei. Conservation implications of drastic reductions in the smallest and most isolated populations of giant pandas. Implicaciones de las reducciones en las poblaciones de pandas gigantes más pequeñas y aisladas. Conservation Biology, 24(5):1299–1306, October 2010. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).